A Bibliography of Publications in The Computer Journal: 1958–1999

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

01 March 2018 Version 1.70

Title word cross-reference

1 − 2 [OW80]. 2 [HD72, ILPP98, MF88, PW94, TH82, Zak84]. 3 [Dob84, Zak84]. 4 [Lu89, Pik81]. 7 [MN90]. link [Che95a]. A [PW69]. A_{k} [Cha93a]. Ax = λBx [Fet65, PW69]. B [PW69]. B_{k} [Roh75]. χ^{2} [PFW74]. d [RDT80, EHH86]. e [Ber69, Sal68]. ℓ_{p} [BRH70]. F [MSS93]. F (X, Y) = 0 [Sut76a]. h [Cha93a]. k [CLC92, Lin72b, LOKM99, PHS96, She95, Bay76, Er83a]. K_{n}(x + iy) [Gar78]. L [Bra94]. L_{1} [OW71, SKY73]. L_{p} [FGH71]. LL^{T} [OK63]. LR [Fra61, GZ72, Wil65c]. m [AD95, AO80b]. μ [BG94]. N [Bel70b, Far78, Hun72, Phi67c, BJ80, LOKM99, Man71b, Ull77, Var80b, Wat81, Har76, Woo77a]. n^{th} [Cam73, Mac73]. O(1) [Tak99]. p [Dub93]. π [Pri95]. QR [Fra61, Fra62, GRB68, OK63, Wat08, Wil65d, Wil65c]. S [Jor73]. s^{2} [PFW74]. σ_{1} [DM64, Eva64, Lyn64]. T [Er87b]. τ [BDG99, Riv68, Wra66]. X [Nar96]. x_{n,ν} ≡ cos(νπ/n), ν = 0(1)n [Sal72, Sal73]. y′ = f(x)y + g(x) [Day65]. Y_{n}(x + iy) [Gar78]. Z [CW98].

-adic [Dub93]. -Ary [Er87b, CLC92, LOKM99]. -bases [Cha93a]. -Bit [Pik81, MN90]. -calculus [Pri95]. -clusters [Lin72b]. -code [MSS93]. -component [TH82]. -cubes [LOKM99]. -D [ILPP98]. -dimensional [PHS96, Wat81, Phi67c]. -Formula [Bay76]. -functions [Jor73]. -gram [Ull77]. -Inhibition [RDT80]. -Level [Dob84].

-method [Riv68]. -methods [Wra66]. -ordered [Eva64, Lyn64]. -Parallel [Woo77a]. -partition [AO80b]. -peg [Lu89]. -restricted [BDG99]. -Set [She95]. -sided [BJ80]. -Space [Har76]. -system [Bra94]. -Tree [Che95a]. -Way [AD95].

1 [Boo78, Cra86, DO72, EAL66, HO78, Loo59, Mac70b, Mac71, Mer70, Obr78b, Pit78, RT78, TS68, Var67]. 100s [Wil67b]. 105 [Ano79b]. 11 [WD80]. 117 [WH87]. 121 [Wic87]. 13 [HJBH67]. 17 [Var62]. 19 [Rud67b]. 1905 [Ham66a]. 1909’ [WR72]. 1960 [Her61]. 1962 [BT63, GT63, Hou63, Nee63, PCSG63, RP63, Var62]. 1963 [BT63, PCSG63]. 1967 [Var67]. 1968 [Alb71]. 1969 [Alb71]. 1970 [Sla71, TCB71, WY71, YFM71]. 1971 [Gal71, Wre71]. 1972 [Bro74b, BS73b, Cha73, GS73, HR73, RL73, Ull73, WS73]. 1973 [Ano73c, IL74, NLW74, RL73]. 1974 [NLW74]. 1975 [FJA80]. 1976 [Ano75a]. 1977 [Bra78, FE80, NQ80]. 1978 [Cor79, FJA80, WD80]. 1979 [BB80a, BL81a, Eme80, FE80, FJA80, Kit74, NQ80, Pau80, VP80, WD80]. 1980 [CF81, JR81, NPCA81]. 1981 [FSWL82, Kil81, NPCA81]. 1990 [HV92]. 1991 [HV92].

2 [Bai88, BH63, Boo79c, Gea64, Ove67c, PHJ67, WHM71]. 2000 [Tho98]. 26 [Hil69, Woo70g]. 2900 [IS81]. 2nd [CF81].

3 [Boo79d, Pal62, Ran79, RP63]. 3-stock [Dou58c]. 34 [Ham69d]. 35 [Cob69, PP69]. 3D [Ric71]. 3rd [Ano75a].

4 [BW69a, Bea74]. 40 [Sco69]. 405 [Liv59]. 42 [Sal70b]. 43 [Woo70f]. 44 [Fie70, Red71]. 46 [KL71, Lil71]. 47 [Mux70, Sal70c, vR70b]. 48 [GE71].

50 [BB71].

60 [BBG^{+}63, BBG^{+}76, DHW76, DHW78, Dij62, Dun62b, Dun63, GS63, Hoc62, McE67, Ryd64, Sam63, Woo60]. 61 [Kro71]. 66 [McC72]. 67 [SM71]. 68 [Bae70, GP80, HH79a, ON79, Pag79]. 68-Rt [ON79]. 69 [Lav70].

7 [PHJ67]. 71 [Dew72a]. 74 [Hos74]. 77 [SW91a, VP80].

80 [HW89]. 802 [Coo60b]. ’89 [Coo89a].

9 [Dun63]. 95 [MWR98, WB96].

A.D.I. [FM66]. A.D.P. [Lav64, Pol60, Str61]. Abadie [TCB71]. Abate [Dur74]. ability [KV63]. ABS12 [Hoc62]. abscissas [OSS69]. absolute [Dun73, HW74, KM94, Par74, Van73]. Abstract [BL81b, CGK80, Kac91, Les83a, Mai85, MTW88, Rob88, YP86, CFL84, FCHK96, Fur89, Rv88, WS86, Yel89]. Abstraction [FM86b, Bur79, JL80, KGC85, Ree80b]. Abstractions [Jok90, Sho81, Wel80, SC87]. abuse [Cha78]. Accelerate [Pro74]. Accelerated [Aha65]. Accelerating [Wri63, Car61]. acceleration [Gou68]. Acceptance [BS61, Sha72a, Pai65]. Access [CDCL97, DL97, GP88, Gun70, HS71, KS82, Kie88, LW85b, MSNW99, MK82, SW84, WW73, WZ99, BT71, BBG^{+}71, Fra68, Har70, HSW92, MT70, O’S88, O’S94, Pez76, WR67, Wil67a, WN68, Win59a, ZR88]. Access-Control [KS82]. Accesses [KP92, WA91]. Accessing [CN82, Mul87, WN81]. accommodate [SW91a]. according [Boo67g, Boo67h, Hil69, Woo70g]. Accounting [Cas63, Moo63, Hop63, Kol83, Sym71, Wil63b]. accumulator [Col63a]. accuracy [Hys72, MP62]. Achieving [Lyo85, ZS96]. Acid [CCL87]. ACLE [PBR^{+}90]. ACM [Har80]. Acquisition [Var76b, AB84, KC88, Pin77]. Action [CGRJ92, WHF83, MKP94]. Actions [MWR98, HRW95]. Active [Gee91, HTW89]. Activity [Tul85, BWM88, Tur68]. actor [KL89]. Actual [HC78, PRLW93]. acyclic [Cha90]. Ada [BM86, Bro84a, Hun85, MWR98, RB85, SW83, WB96, Wic82, Wic87]. Adaptation [Cha62]. Adapted [Ver87]. Adapting [SB99]. Adaptive [Akl87, AA68, DS99, FA88, Hie98, SP91, BM88, CC93, DP76, Haz78, MC95, Mak89, Oli72, Sri91, Wit80]. adder [Fen87, Sal87]. Adding [BP92, LE86]. Addition [War94]. Additions [AM98]. Additive [MP68]. Address [Par88, Col63a, Gea64, Las68, Mic64, SAA88]. Addressing [Bra85, Bra86a, CEM72, EAL66, RKA92, Bre93, CFL84, Elc65]. Adele [BM94a]. Adequacy [DW88]. ADI [Had71, Lew71b, ML65]. adic [Dub93]. adjusting [She92]. Adjustment [O’R70, Hio67, Whi72]. Admin [PPB84a, PPB84b, PPB83]. Administration [NT79, Aki88, Sam65, WKRJ60]. Admissible [Mak91a, Er89a]. Admission [CW96a]. adoption [AB84]. adopts [Hin60]. Advanced [Lew71a, NQ80, WW83, Bai88, Pri64]. advances [Mur83]. advantages [Sal72]. adversity [Mus66b]. advice [BM80]. aeroplane [Tay61]. Affine [LF97, Skl81, Bra94]. after [SPMD83]. against [Hie98, OL68]. age [Geo63]. Ageing [DGM68]. agglomerative [EHW89]. Aggregate [Bow84]. Aggregate-Incomplete [Bow84]. aggregation [Rib95]. agreement [BDG99]. agricultural [AS63]. AI [Bru92, Che95b]. Aid [DP88, Mul81, WB99, BD61, Bow73, Bro69b, FFMR66, HM80a, HK67, KH81, Tom82, Tra82, Tra84]. Aided [FR80, Kin69b, Wat74a, Gol69]. aids [Geo63]. Air [Dav63, LS97, Lan65].

Air-Storage [LS97]. Aircraft [Moo59]. aired [Gil73]. Aitken [Eva63, Eva64]. ALAM [d’I69]. ALEC [NF76]. Algebra [BBK87, Bar84, Bar85, BBP94, BKLL95, Sar90, BGMW96, Bar67, BBB68, BBF70, BBH70, Bur67, CCB95, Cli62c, FZ94, HH95, KU74, UK74, ZD71]. Algebraic [Bir89, Can93a, EMCO92, EMO92, HK88, Jel79, Nas75, TD77, Ull67, Wil60, d’I69, BF72, Boh68, Fit73, FR63, FM85, Hit78, Hus61, LM77a, LS78, MR94, McC93, Ros67, SG78]. Algebraically [Fos86]. Algébraiques [Wil62a, Wil63a]. Algebras [GH95, HM95a, Ret95, Rib95, Ser95]. ALGOL [Dun62a, SS72a, Bro74b, PA90, BBG^{+}63, BBG^{+}76, Bae70, Bar65b, Bar64, CHH73, DHW76, DHW78, Dij62, Dun62b, Dun63, FK68, GP80, GS63, HP71, HH79a, Hig63, Hil71, Hoa62b, Hoa63, Hoc62, HB67, Jon72, Lin72a, McE67, Mil68, Pul64, Ryd64, Sam63, Sam65, SAHS71, Sip69, Tan78, Tay64, Tru66, Val74, Wat63, Wic72, Woo60, Zar69, DM85, ON79, Pag79]. ALGOL-60 [BBG^{+}76]. Algol60 [Wil78a]. Algols [FJA80]. Algorithm [Abb80, Ada69, AHM74, AM97, Bel70a, BE73, Ber76, BGO94, BGR86, Boo67k, Boo67e, Boo67g, Boo67i, Boo67j, Boo67b, Boo67h, Boo67a, Boo67c, Boo67d, Boo67f, Boo68a, Boo79b, Boo79a, Bre71, Bro79a, Bro69c, BW75b, BF78, CL83, Chr73a, Chr71a, CD80, CHPS85, CW70, Cra72, CM93, Dak65, Day69, Day76, Day77, DdS98, DA77, Del77, Dew71, DS90, Dob91, DB81, Er83a, Er84c, Er88b, EA70, Eva71, ES77, Far78, FB75, FG77, For75, Fro81, Gar78, GC87, GRB68, Ham91a, Hat84, Haz78, HPR88, HPB84, HAAZ99, HP67, Hil75, Hil77, Hil79c, Hil79b, HP79, HWW87, Hos70, HKA72, HK72, HM77, HGMM75, HKPT99, Hun85, IZP81, IP86, Jen71, Jon72, JBJ84, Kro70a, Kro70b, Kro70c, Kro70d, Kro70e, Le 69, Les83b, Lig73, Lil70, Lit70, Mak68a, Man71b, McC71a]. Algorithm [McC71b, McC78, McC82, MH94, McL78, NR92, Obr78a, OP71, Øst79, Ove67a, Ove67b, PTV97, PP98, Pie82, PP79, Pik79, PFW74, Pit67, Pou93, Pre68, PD79, Ree79b, Ree79a, Roh76, Roh73, Roh74, RL75, Rud67a, Sal70a, Sal71a, SMR98, She92, SM97, SW91a, Smi70a, Spä68, Spä69b, Spä69a, Spo85, SHCP91, SM88, Sut76a, Sut76b, TI75, Til69, Van73, Van91, VR81b, WT79, Wat08, Wat11, Wat67, Wil85a, Wil83b, Woo69d, Woo70e, Woo70d, Woo70c, Woo71c, Woo71b, Woo92, XY90, Zak84, Zak77, Zam70, Zam74, ZOG85, vR70a, van70, Akk73, AJ70, Alm66, AM65, Alw71, Ano80b, Aus76, Aus71, Bae62, BJS74, Bar78a, BE79, BH68, Bar65a, BMT85, Bra75, Bro83, BP69, CDHL93, Cha97, Che95d, CLW98, Com85, Cox71, CD85, Dav68, Def77, Deu69]. algorithm [DMS^{+}89, ET88, EL89, Er88a, Er88c, Er89b, EN82, Eva72, Eva75, EH76, GM72, Gol78a, Goo75, GZ72, GP96a, Ham67a, HD71, HL76a, HR70, Hit78, Hor93, HCH96b, Hut86a, ILPP98, Iga80, Jon70, KR89, KK94, Kni68, Lau91, LS78, LMS73, Lut80, Maj94, May66, MR76a, MB66, MR76b, Mul75, Mux70, NSW87, OW69, OW71, PG82, PK85, Ric71, Rob63, Roh91, Ruž91, Rya66, SS96a, Sal70c, Sal74, SE82, Sib73, SG78, Sne80, SAA88, Tak96, Tem69, Thi81, TU96, VR81a, War94, Wes80a, Wil65d, Wil76a, Wil86, YJ94, Yul68, dS77, vR70b, van71, Lee74, Mar74b, Ano79b, BB71, Boo68b, Boo78, Boo79d, Boo79c, Cob69, Dew72a, Dew74, Fie70, GE71, Ham69d, HJBH67, HW74, HO78, Hos74, KL71, Lil71, McC72, Mux70, Obr78b, Ove67c, Par74, PHJ67]. Algorithm [Pit78, PP69, Ran79, Red71, RT78, Rud67b, SM71, Sal70c, Sal70b, Sco69, WH87, Wic87, Woo70f, vR70b]. Algorithmic [BBG^{+}76, V’y99, vSHBT98, BBG^{+}63, Oul87]. algorithmically [DP66]. Algorithms [Akl87, Ano78, BE82, Boo67l, Boo67m, Boo67n, BF81, Bri85, CO76, CR87, CV86, CZL97, Cle59, DS91, Day78, DM85, Dro87, GH75, Ham67b, Ham67c, Ham67d, Ham67e, Ham68a, Ham68b, Ham68c, Ham69a, Ham69b, Ham69c, Ham70a, Ham70b, Ham70c, Ham70d, Ham71a, Ham71b, Har92, HS71, HM83, Jar70, Jen99, Joh74, JES98, Kho93, Man93, MLR90, Mit76, Mur85, NS87, NR79, Pha90, Phi68a, Ree84a, Ric88, RV78, Roy78, SNH76, Sam76, She71, She72a, She72b, She73a, She73b, She74, She79, Suf90, Tak99, TCB71, Var68, VR92, WT79, Wit80, Atk80, BM90b, Can93b, Che95c, Cho95, CP66b, FG90, FL80, FCHK96, Fle70, GP80, GF91, Gol78b, GP96b, GHR^{+}93, HM96, Lep95, LoD71, Mad97, ME88b, Mur83, Oxl89, Par96a, PR64]. algorithms [RS69, Sab72, Sta71, Tza73, Wal77, Wil65c, WM84, Ham66b, Hil69, Kro71, Wil70, Woo70g, O’K85]. Alignment [LF97]. All-sharing [HLK94]. Allen [Nee63]. allied [Scr64a]. allocated [Dav82]. Allocating [Zmu75a, Zmu75b, KH81]. Allocation [Bas86, BGR86, CC87, CCC92, CD82, McI82, NR79, Pfl84, Ree87, Rob77, SZ74, Aus71, Cam71, CC88, CLC92, IJ62, Lam63, Leu82a, Ree79c, Ree80a, Ree83, Ree84b, Wan80]. almost [Fos68a]. ALP [CW62]. alphabetically [Far78]. Alphabets [TM95]. Alphacode [DH60]. alphanumeric [Tho64]. alternate [Jon66]. Alternating [MM70, MM71, Eva65a, FM64, Had68]. Alternative [DC90a, Hou86, Sch85, Bak64, MD87, Par78, Zar80]. Alternatives [Fry71]. Ambiguity [Sch85, MV88]. America [GG59]. Amoeba [Shi90]. Amoeba-Prolog [Shi90]. Among [BT81, Kin69a]. AMOS [Ber59]. analog [SWZ93]. Analogue [Her61, BBS66, Bea68, BW69b, Cra67, DP66, Fir61, Fra61, HB66, Lan65, Lew65a, Mea73, RWT64, RW68, Tho61, Tho68a, Tiz60, Vin59]. Analogy [Mic97]. analyse [LW66a]. Analyser [Wat76a, Mar69, Sim69, Wat77a]. Analyses [Ree84a, O’K85]. analysi [LW65]. Analysing [Jar68, LE86, Par75a, Tow78].

Analysis [ABD84, AB78, Aky92, AHL92a, BS86, Boo68b, BF81, CS77, CC87, CDZ74, CY95, CW96b, CL99a, CK94, Cla77a, CGK80, DP88, EW80, FS94, FSW85, FR98, HLK94, HD82, IL74, Jos70, Kin69b, Lar85, Leu83a, LW85a, LCH99, Lom88, Mic91, Mil81a, PW81a, Pag84, PL74, PP98, Par75c, Par82, PB76, PB78, PAW86, Pfl84, Phi70, Ric88, She99, Ste77b, SAA88, VR92, WL96, War96, Wat75a, Wil67b, XY90, ZE88, ZOG85, AB84, AHL92b, Bae66, BCH91, Bec58, Cam73, Che95c, Coo83, Fin93, FK68, Gea80, Gen69, Gha72, GE71, Gri62, HV78, Ham80a, HM80a, HR70, HGL93, JC86, KP60, Leu82b, LC93, LS93a, Lit70, Liv60, Lut80, Mac73, MG85, MR77, Mar76, Mar91a, Mid77, Moo59, Mum80, OW66, Osb64a, Par75d, Pen66, RT76, SI80, Sau72]. analysis [Sco82, SFD92, SW68, Sla62, Smi66b, Spu88, TW64, Var80a, WJW95, WW79, WM95, WHF82, Wri74, YS60, YS61, YGS63, dSA92, Bro74b]. Analyst [BS87, AT70, Fel69, Fel70]. Analytic [BD81, CL99a, FS81, Ros67, Sch65, Zve75, Ede65]. Analytical [Sto86, Wen59, Ran71, WR72]. analyzed [Fos64]. analyzers [CND66, Wal64]. Analyzing [Kuh92]. anatomy [LM77a]. AND-Parallel [NL90]. and/or [Kol83]. Andrews [Ull73]. Anecdotes [SHH81]. angular [Ede65]. Animated [Rep72]. Animating [BS99]. Animation [PW94]. Anne [NQ80]. Annotated [SS80]. Annual [RP63, Ano73c]. ANSI [BP81]. Answers [FR98]. Anthony [VP80]. any [Wat77b]. APL [Ear75, Hol78, JK80, LL75, OL78]. appears [OM64]. applicability [Ros82]. applicable [Tay70a, Woo69c]. Application [CBH^{+}94, Cla78, Dea79, Flo76, FH97, GH95, HW77, IB96, KP79, KBTB97, NM90, Par75a, Sel88, Sme92, SA81, TR93, Vin59, Wis64, Ake66, ABN61, Bar65c, BF72, CAC82, Col63b, CR63, FB75, For66, FGR67, God75, GF87, HK59, HB58a, HB58b, HCH96b, Hur73, JI93, LL90, MC95, OS65, OM64, Per76, PW67, SWA97, Smi66b, Stu71, TS64, Tho62, TB63, Wat81, WP67, Wri59]. application-oriented [For66, FGR67]. Applications [All88, BMRS90, EHH86, GV99a, JLV99, LR78, Mel95, SD82, SM88, WS73, BP94, Cam97, FK68, Gea61, GK96, Gos60, Har88, Kol83, Pai60, RL73, RP63]. Applied [BES83, FJA80, Eva64, Lew65a, Lyn64, Par78, ZD71]. Applying [LS93a, LW93, VH99, Che95b, Pla61]. Appraisal [Cle92]. appreciation [MM69]. Approach [AMN81, ABC^{+}83, Ben82, BR86, Bro84a, CT94, CM79, Che86, CT92, Er83c, Er84b, FE80, FW86, Har92, JSKCG94, Jon71b, KB87, LO99, Mar91b, MMR^{+}98, ST82, Tse87a, UVS87, VVW96, Wal76, WT84, WL87, Wil84, ZSM99, AO80b, Ano63, AHL92b, BM90a, Bak94, Ben72, BD80, Bra76, CBB^{+}76, Cuc93, DP73, Er82b, Fle70, For66, FGR67, Fun85, FM85, Gea59, GlM68, Gre67, Har70, HM80b, Hum62, JC88, Lar64, Lar95, Lle73, Mar91a, May64, MHLS93, MW68, PW74a, Pen76, Pen77, PL94a, PL94b, Pro80b, Qui83, Roh75, Ros66, Rus86, Sch94, ST89, Spu88, SL73, Thr69, Ull65, WN88, Win75]. Approaches [Bar71, CH99, Ole91, Coo83, EWH86, Gea80, JJ67, KH87, RDPS95, Var80a, WHF82]. appropriate [EWH86]. approximants [HGMM75]. Approximate [BWM88, HKPT99, TDW^{+}99, Akm91, APD99, BLPL92, Wes80b]. approximating [Cox71]. Approximation [McL78, OP73a, VG99, YFM71, Ash68, BY66, BMT85, Cob69, CP66a, Eve63, For72, KR89, Kei79, KL71, Lil70, Lil71, MR76a, MR76b, OP73b, Osb64b, OW67, OW69, OW71, PP69, Pre68, Rei71, Riv68, SKY73, Tee63b, Tem69, ZL73, Sal70a]. Approximations [Bel63, Har76, BRH70, CP66b, CO66, Dun73, FGH71, Jac73, OP75, Osb63, Osb65, Phi68a, Phi68b, Pow67, Wat73]. April [Var62]. Arbitrary [Cra72, HPR88, McL74, ON90, Pag82b, AO87]. ARC [Kel98]. ARCA [MHS85]. architect [FFMR66]. Architects [IHH95]. Architectural [Ole91, PGM84, New66]. Architecture [AE90a, Bow84, BWW81, Bro74a, DAODT85, MRW90, Rey90, TGE96, Tur92, UHS85, WW83, ATL93, BBKR89, BB80b, CMW98, CFL84, FH77, KVKP95, MMH87, ME80, WS74]. Architectures [BCV93, DT96, Nai97, Tre82, CFS89, TAP89]. archiving [Wig75]. arcs [Abb80]. arcsin [Mak68a]. Area [GP88, IEH^{+}85, KA98, Mul81, Smi71, Lew71b, VV94, vSHBT98]. Areas [Ada69, Bae75, OW83b, dS80]. argument [OL68]. Ariadne [Cra86]. Ariadne-1 [Cra86]. Arithmetic [Aki77, Goo77, Goo80, GH73, RL75, SD82, Vel82, CL66, Dew71, Dew72a, FN93, How67b, Jon72, RL72, Spo85, Dew74]. Arithmetical [Ham69d]. Arjuna [DPSW89]. Army [Sla60, TS64]. ARP [MZ96]. ARPL [HR74]. Arrangement [BENM98, Bea65, Gil66, RW67]. arrangements [SRS94, VR81a]. Array [Bow90, PS96, Boo67g, Boo67h, Boo67f, BW73b, DW77, Hil69, Hor93, LC90, Meg90, ME90, Woo70g]. Arrays [Bla82, Cha91, HM93, TL97, ZBS81, Cha92, Fos68a, HV74, Iga80, Kro70d, ME88a, Mul87, Smi70a, Tay76]. art [Dou59b, Gol59]. Arthur [WD80]. articulated [LP73b]. Artificial [Bur80, BMPS71]. artificially [ON94]. Arts [WCH78]. Ary [Er87b, CLC92, LOKM99]. Aspects [F^{+}73, CL78, Dem60, Dru63, Lew71a, Mum80, Pie71]. assembler [Gea64]. Assembly [WD80, Bow73, BM60, BM61, Cle67, LWS69a]. assembly-language [Bow73]. assertion [YG93]. assertion-guided [YG93]. Assessing [Rze85, SW79]. Assessment [LW83, Mae78, ABBS83, BD61, JP69, Par78]. Assignment [CP64, ME90, Pag63].

Assisted

[PW94, Smi83, Geo66, WW79, LM88]. Associated [Mar94, GL64]. Association [AM92, LW65, EDS86]. Associative [Eke95, Eme80, Lea78, Lea86, BL72, CWH93, MMH87, Par96a]. Associative-Commutative [Eke95]. assurance [Gil64, Sut60]. asymmetric [Duk66, Jon66, Rac65]. Asymmetrical [BZN98]. Asymptotic [Cha68, Jar67b]. Asynchronous [ZSP99, BBP98, BKK78, Che64, Lew71a]. Atlas [BT71, BRC66, BMR67a, CP62, Dol66, EAL66, HB67, HPS61, HJW62, KHPS61, McE67, MR67, RC70, SW68, d’I69]. Atomic [MWR98]. Atoms [WJ61]. Atre [NPCA81]. Attachment [PW81b]. attempt [WY62]. Attitudes [EDS86, JC86, Lem80]. Attribute [BDS82, BW78, DC90b, Fro92, PK86, TUW98, WM83, JO92, KC88]. Attributes [Der79, Wu96]. Audit [Cha75, Tho60]. Auditing [Cha81, Flo74, Goo64, dP60, Tho60]. auditor [Fri65, dP60]. augmenting [Hut86a]. Australian [CF96]. Authenticating [LC94]. authentication [ZH93]. Author [Ano97a, Boo67l, Boo67m, Boo67n]. Authoring [BGS99]. authorities [Mal59]. Authorization [TS64]. Authors [Ano82b, Ano81a]. AUTO [TCL89]. AUTO-DFD [TCL89]. autocode [BM63, Bro58a, Bro58b, CMRW72, CF59, CW62, Gib61, YGS63, And66, BRC66, BMR67a, Gea61, Gib68, McE67, RC70]. Autodescriptivity [BRS91]. Automata [Ano75a, FH98, Gol93, OM91, SY72]. Automate [JES98]. Automated [GN94, PSK96, SN75, VW96, CZW94, FCHK96]. Automatic [AEH75, Cho82, Clo86, DN64, Fai58, Gil65, Han70, HW77, Hou63, KP97, Mea73, Mui58, O’N92, Pai60, RP63, SI80, Sch82a, WL72, Ano60, BWZ71, Chr73a, CP61, CP71, FR63, Fra67, FW83, GL70, Gib60, Gil59, Hin60, Inc87, JP69, JS71, Lei62, Lyn69, LK77, McC71a, McC72, Rai87, RB86, Ros60, Tho58a, Tow77, Ull77, Val71, WF78, d’A62, vdV70]. automatically [Par75d]. Automating [CGRJ92]. Automation [GT63]. automaton [OF93]. automaton-based [OF93]. AUTOSTAT [DM60]. Autumnal [FM86a]. Availability [TK99, Mar76]. Available [CM91, Gri92, TF96]. Average [Bra85, Bra86a, Dei95, Li98]. AVL [OW80]. avoid [Cam58]. Award [HT81]. Ax=b [Loi68]. Axially [Abb64, Ran68a]. Axiomatic [FW86, Pra84, Wri74]. axiomatising [Gib88]. axioms [FZ94]. axis [McC71b, SM71].

B [HR73, Kit74, NPCA81, Dou59b, BD84, BKMK90, Cha92, Che95a, LG86, MZ86, Mul81, Sri91, TB86, ZOG85]. B-Splines [MZ86]. B-Tree [ZOG85, Cha92, TB86]. B-Trees [BD84, LG86, Mul81, BKMK90, Sri91]. Babbage [SHH81]. Bacchus [Ede63]. Backbone [RLD94]. background [Rob63, Spä68]. backing [Cha62]. Backtrack [Hel84]. Backtracking [Irv84, Er87a]. Backus [CCH73, Roh68]. Backus-Naur [CCH73]. Backward [MS92, Rya66]. Bacon [NLW74]. Bairstow [Han62, Wat67]. Balanced [Cha91, EA83, GOW83, WW76, ES77]. Balancing [Day76, Gra71, GIB90, Wil78b, BM93, HLK94, Smy91]. Balfour [VP80]. band [PW69, Rei67, Zam70]. banded [EH76]. Bandwidth [WL96, AM65, Jen71]. bank [Hin60]. banking [BSH70]. banks [Hin61]. Barron [NPCA81]. Bartee [PCSG63]. Base [AK81, BHW78, CS77, CGRJ92, Dav78, Dav81, Dea79, Eri74, Flo76, Fro85, GKS79, HW77, NPCA81, PDY74, PSD^{+}81, PJ80, SM88, Str72, VP80, Wel80, BJ78, CCL72, GW88, Mil81b, Pro80b, Tay90]. Based [AM99, BK92, BKLL95, Cha87, DS99, DW88, DRT96, ES81, FSWL82, FR98, JCCJ99, Lal99, LW83, Lom88, MK85, MW97, Nar96, OM88, Ric88, SCC^{+}96, TUW98, TK99, WL96, WC86, Yok97, YP86, Azm90, BWM88, Bar65a, BBG95, BMNS87, BKM78, Bre93, Cam78, Cha92, CW96a, CLW98, CFL84, Cra72, DP66, Du96, ESK95, Fun84, Fur89, GGH^{+}87, GHR^{+}93, JES98, KL89, Ken80, KGC85, KVKP95, Lan76, MKP94, MV88, Mar91b, Mic82, MT86, Mul92, OF93, PC78, PL94a, PL94b, Ros66, SP91, TWMG93, TAP89, Tur78, WJW95, YY89, vBFvdW96, KM74, MBC^{+}86]. Baseline [LC98]. Bases [CO76, CEM72, Sar96, Wil72, Cha93a, Don78, MMH87, NS77, Sch76]. BASIC [Bar72a, Sil72, BC73, Coo68, FZ94, Ham69d, Sol68, Har70, MR94, Ull65, YS61, Bul72, Lee72, Ili69, Wil78a]. Basis [Mac81, PD96a, AS63, Dou65, Mar82]. batch [HG71, Sto72]. Batched [MKH86, MKB87, Kol78]. batchwise [You60]. Bayesian [FH97, Wu96, van96]. BCL [Hou69a, Hou69b]. BCPL [Bro82]. BCS [Har80, Dak73]. BD [DS91]. Be [AC84, BS73a, Fin60, Hil67, PJ89]. Beaumont [Ano76]. become [Mil95, MMG91]. BEDSOCS [ER80, SB81]. Beginners [Gal92a]. beginning [Cam97]. Behavior [Joh74]. Behaviour [And90, Jos70, Abb70, BM84, CM89, MES97]. Behavioural [Les83a, Cla94]. Behaviours [HJ99]. Behind [Rob75]. Belief [CGRJ92, van96]. Benchmark [Wal76, CW76]. Benchmarks [DBCRW92]. Bendix [Dic91]. benefits [Pai69, WP69]. Bessel [Gar78, Sto68]. Best [Cob69, Fle68, PP69, Pre68, Rob77, BRH70, CP87, FGH71, OW67, Phi68b, Wat73]. best-fit [CP87]. beta [Luk99, Pha90]. Beta-splines [Pha90]. Between [Bra86a, Bru92, Ste77a, Boo67c, Boo67d, Bra85, Cor65, HO93, Nic66, Obr67, Tiz60, Wak93, WCL71, Wil78d]. beware [BRS91]. Beyond [IS81]. Bézier [For72]. bi [dS77]. bi-directional [dS77]. bibliographical [YAH80]. Bibliography [SS80, DGM68]. Bickley [MM69]. Bidirectional [FK90]. Biharmonic [BW70a, FM64, Gre67, Tee63b]. Binary [AM90, And90, Bri85, EA83, Fro82, Har95, Les83b, Mah86, Mak91a, MKH86, Pal86, PP98, PB76, PB78, Sha81, Sip88, Spr92, V’y99, WW76, Wei93, Wym73, XTU97, ZOG85, CC88, Chr71b, CM89, Day76, Er89b, ES77, FL80, FN93, FP82, Gil58a, Hat84, Hur68, Hur73, Mäk91b, OV92, Pal88, PW74b, RM89, Sal87, SGR96, Sou77, Ull77]. Binary-Relational [Fro82]. Binary-Residue [AM90]. binary-ripple [Sal87]. Binding [ABM88, LM77b]. biology [CRS^{+}93]. Bipartite [Eke95]. Biquinary [Bro74c]. Bird [GS73]. bisection [BE79, SE82]. Bishop [Wil67b]. Bit [Pik81, BG94, LMS73, MN90]. bivalent [Chr73b, Mot80]. blackboard [Cra86]. BLEND [MP85, Sha82, SPMD83]. BLEND-LINC [SPMD83]. Blind [Fis76]. BLM [Ili69]. Block [Edw77, Hou86, ID73, Lop94, PTV97, SH97, TH82, Car66, Fen96, FNK94, Han94, Lee78b, OS65]. Block-length [ID73]. Block-Structured [Hou86, Lop94, Han94]. Blocked [CN82]. Blocking [KW84, MS74, NV77, Wat71]. blocks [Sam65]. Blood [Sca85]. Bloom [Mul87, She89]. Blue [BKMT96]. board [RHS89, VW95]. Book [Alb71, Ano96a, Ano96b, BB80a, BL81a, BT63, Bra78, Bro74b, BS73b, Bur80, CF81, Cha73, Cla92, Cor79, Den80, Eme80, FSWL82, FE80, FJA80, Fry71, GS73, Ham91b, HV92, HR73, Her61, Hou63, IL74, JR81, Kil81, Kit74, Llo74, MdlIDS99, Nee63, NLW74, NQ80, NPCA81, Pau80, PCSG63, RL73, RP63, RLT99, San99, Sla71, TCB71, Ull73, VP80, WS73, WY71, Web70, Wil62a, Wil63a, Wil67b, Wil70, WD80, Wre71, YFM71, Bon71]. Boolean [BML94, Bla82, Edw72, LT79, Mar69, Miy90, Wil77d, ZD71, ZD73]. Boor [GT63]. Both [SNH76, Bae66, OSS69, Pen66]. Bottom [UVS87]. Bottom-Up [UVS87]. Bounce [NI70, Osb71]. Bounce-and-skip [NI70, Osb71]. bound [WP67, MMB87]. Boundaries [BW70a, LW66b]. Boundary [BW70a, Tay70a, AH69, Bic68, Bic69, Cas75, Eva77, Fyf69, Fyf70, Has61, JM68, KM66, Osb69, Pri68, Sne70, Tay70b, TF96, Thi81, Zaf71]. boundary-restricted [TF96]. boundary-value [Pri68]. Bounded [CD82, DE86, WC98]. Bounding [FK93]. Bounds [Han70, KP98, LM68, MSNW99, SNH76, Wil61b]. Boyer [Tak96]. Boyserf [Ree79b]. BPL [Wil82b]. Bracketing [Sza73, Cox70, SC71]. Brain [WY71, dF66]. Brains [WY71]. Branch [JES98, ME98, Won99, MC95, Tho80]. branches [Wig81]. branching [HRW95, Ret95]. break [AB84]. break-even [AB84]. Breaking [HM83]. Brelsford [FSWL82]. Brent [SL78]. Bresenham [PG82]. bridge [Cha97]. bridge- [Cha97]. Bringing [Par75b]. Britain [Gol59, TB63]. British [Aga96, Bon71, CK95, Dou59b, Pla61, Ros70, VW95]. BRMQ [Azm90]. Broadcast [FLW99, KA98]. Broadcasting [BENM98, PHS96]. Brodie [JR81]. brother [OW80]. Brown [NQ80, Bur80]. Browsing [LSV89, DB88, MP85]. Bruijn [VR92]. Bubble [Wri83]. Bucket [Lyo85]. Buckley [NQ80]. Buddy [CR87, PBH86]. Buffer [DE86, LV79, DL76]. Buffered [HPB84, WH87, PM90]. Buffering [Wri82]. Building [BL91, Cor79, God96, NCR97, Sar96, Sla71, Wai70, Kec87]. Bull [NLW74]. Bulletin [Ano73a]. Bulterman [Er88c]. bureau [AF76, Dou65, Kin62, Lew65b]. Burrows [Fen96]. Bus [WL96, Mar82, WWN91]. Bus-Based [WL96]. bus-connected [WWN91]. Business [FE80, LO99, Mum85, PS93, Ste77a, Bar58, Ben70, Cam97, Che62b, CT74, Coa74, DG58, EDS86, Fis74b, Gea58, Gea59, Gea61, Gea73, Lic89, Pai60, Poo91, Wri59]. bypass [LOKM99]. byte [BS89].

C [Fry71, IL74, Llo74, PCSG63, Ris99b, Ull73, Wre71, KMK95, SAM88]. C-ODA [KMK95]. Cache [FB99, Lop94, SM97]. Caching [FLW99, LS97]. CAD [Mar73, Mar74a]. CAL [GF87]. calculate [WD79]. Calculating [Oxl89, Pow64, Pow65, Abb70, CP66b, Goo61, Pec70, Pha90, Tay70a, Zan67]. Calculation [GRB68, Hon71, Pow68, Sam72a, Wil58c, Wil58b, Wil65b, And69, Ber69, Bir89, Eve63, FGH71, Hun85, May66, Sal68, SAA88, SD62, Wic87]. Calculational [SB99]. Calculations [AB90, God75, MMB87, Pal63]. Calculator [BS99, Thi95, Gol94]. calculators [BL77]. Calculus [Pai99, All66a, Pak67, Pri95, Tur78, WS73]. Calibration [GP81]. Call [Ano75a, Ano82a, ZG99]. calling [BB80b, ME80]. Calls [Ano80a, HR96, PJ86]. CAM [Wat76a]. CAM02 [Wat77a]. Cambridge [Wil67b, BBH70, JJ67]. Campey [Hos71]. CAMPS [LM88]. can [Fin60, PJ89]. Canadian [God75]. candidate [SS96a]. cannot [RHS89]. Canonical [DNV81, Dee80]. Capabilities [FC87, SW79]. capability [APW86, CFL84, MT86]. capability-based [MT86]. capacitated [Lig73]. Capacity [PCSG63, CWH93]. car [Lei62]. Carberry [BL81a]. Cards [BDB92]. Carlo [Akm91, CR63, KB62, KP79, Mus59]. carousel [SRS94]. carry [Fen87, Sal87]. carry-completion-detection [Sal87]. carte [LSV89]. Cartesian [CC88]. Cartoon [Rep72]. cartridges [SRS94]. Cascade [Sco74]. Case [CCL87, DH97, HZ95, Joh74, Lar82, LZ95, Mum85, Rob77, SW79, VR92, Wre71, Cla88, Erk84, Lee78a, LP73a, Mid77, OW69, Osb93, Tem69, TB82, dSA92]. catalogues [Bir81]. categorical [TLF96, TL92]. Categoricity [Les83a]. Causal [KA98, RG95]. Causality [BKLL95]. Causality-Based [BKLL95]. cautious [Haz78, Qui83]. cavity [Gre69]. CCS [CHM94]. CD [HLK94]. CDC [BBG^{+}71]. cell [Hol95]. Cellular [DS99, ZBS81, DW77]. census [TB63]. Central [Jen99, Fos71]. Centralised [Dav78]. Centralized [IP86]. centre [Buc60, Muc82, TS64]. Centrenet [IEH^{+}85]. centres [Jam81]. Centroid [Mot85]. Certain [Bro74e, FD88, Hun70, AHM74, Bak64, BM61, Coo66, EF63b, Eva72, EH76, Fyf70, Goo75, GM65, HM77, Jar66a, Jar67b, LoD70, Lyn65, MP65a, MP66, Sal70b, Spä69b, Wil59]. Certification [Boo78, Boo79d, Boo79c, Ran79, RT78]. CFG [HKPW89]. CGI [Tom82]. Chain [Mer70, She95, Woo70e]. Chained [Bay73, CS77, Sur86]. Chainmail [Ree87]. Chains [Bra88, Leu83a, Lor76, MKP87]. Challenge [NQ80]. Chalmers [AJ89]. Change [Mul81, Tre70, Chr73a, Pap60]. change-ringing [Pap60]. Changes [Kuh92, Cur95, Ano73c]. changing [Bri80, Ham80b]. Channel [DRT96, RM89]. channels [LOKM99]. Character [GB61, Hin61, MN61, MN90, Rey77, Sha65, WW82, YAH80, Bau65, CCL72, CP61, Gri80, New61, Pai66, cGR61, Ros64b, Sau72, Str61]. characterisation [HRW95]. Characterise [Fro85]. characteristic [Hop71]. Characteristics [Lle87, Lyn72, RS88]. Characterization [Mic63]. Characters [AG98, Coo68]. Charles [HV92, SHH81]. Chart [PG81, OB68a]. charts [Dun79, How67a, MR94]. Chebyshev [Sal73, AEH76, Bea65, BS71, Cha67, Cha70, CN63, Coo67, Eg69, Ell63, Fox62, Gou68, KS71, Nor64, OP73a, Oli69, OW67, Sal72, Scr65, Scr70, Sne70, Tee63a, Wol69, Wri64, Wri63]. Check [Bro74c, ES75, How75, War86]. checked [Cur95]. checker [HP71]. Checking [ABM88, Cow86, Dun78, Hig76a, And74, Bro74d, CY95, Hou83, LEN63]. checks [HK73]. Chemical [GWWB98, Cam69, Gal59, KM74, You60]. chemin [Fos64]. chemin-de-fer [Fos64]. Chess [Mic81, Pro74, Bel70a, BJ79, BB71, BKM78, BM80, BW75b, RHS89, WBB72]. chi [PFW74]. chi-squared [PFW74]. Chilton [BT71]. China [GGH^{+}87]. chip [JI93]. Chisholm [HGMM75]. Choice [Edw77, JS71, Sto86, BSH70, WCL71]. choose [EWH86]. Choosing [Roh78, Wil78c, cGR61]. chordal [Bro95]. Chow [Pau80, Hur73]. Christoph [YFM71]. Chromatic [Chr71a, Col82, Col67, Mit76, SNH76, WP67]. Chuen [Pau80]. Chuen-Yen [Pau80]. Church [PR98]. Cie [Wil62a, Wil63a]. Cilk [VH99]. Ciphers [GN94, HM83]. Circle [NP90]. Circuit [AM98, Mic97, AL72, SWZ93, ZD71]. Circuits [PL74, TD77, YT79, AA68, BH68, Ben72, CDHL93, DZW75, OS65, PM90]. circular [Lew71b, MP65a, MP66]. Circulating [WI84]. circulation [Dow71]. cited [Chu68]. City [PTV97]. CLAM [dRC74]. Clashes [DC90a, Lah91, Stu84]. Clason [GT63]. Class [CD82, HW92, Has75, Nar96, WL96, Wen59, Che95d, Fox67, Gro68, HH94a, IW86, LoD70, PW74a]. Classes [Bae75, Er89a, FD88]. Classical [EMO92, Gal92a, van86]. Classification [BW73a, BW75a, JW74, LJ98, Oka98, Ros64a, WB68, Wal98, BW70b, CD80, Con66, Hal68, JW72, JS71, JJ67, LW65, LW66a, LW67c, MD87, Mur98, Sal71b]. Classifications [HAM98, JS68, WCL71]. classificatory [AW85, LW67a, LW67b, LW68, LW71]. clause [Pad91]. Clauses [Ber87a]. Clear [GGH^{+}87]. Clenshaw [Hun70, OS68, Oli72]. CLEO [Ric64]. Client [PRS96, TU96]. Client-Server [PRS96, TU96]. Clinical [Gor78, CBB^{+}76]. clock [CF88]. Clocks [JCCJ99]. Closed [Bra88, Jac73, Tha66]. Closure [FD88, Weg74]. Clubs [Rob90]. clues [SD86]. clump [JJ67]. clump-finding [JJ67]. Cluster [FR98, SW68, SWZ93, Sco82, Sib73, WJW95]. cluster-based [WJW95]. Clustering [BMSA88, FM86b, FR98, HR97, Mir98, Mur85, EHW89, Mur83, Mur98, Mux70, Ree84b, Roh73, Sal70c, Sal71a, vR70a, van70, vR70b, LW67b]. Clusters [CW70, FR98, SAM88, GP96a, Lin72b, Roh75]. CMCM [TL92]. CMOS [Fen87]. CMSR [BG71]. Co [Bra88, WG83, MES97, RW61]. co-diagonalization [RW61]. Co-expressions [WG83]. Co-Relationship [Bra88]. co-simulation [MES97]. Coalesced [CV86]. coated [Lep95]. Cobol [Eme80, Ano73a, Cli62a, Gil69, Gol64, Hum62, Kin62, LM77b, Tri74, Tri76, Tri78, Tri80, Tri83, TL85a, Wil77a, HH79b]. CODASYL [Aki88, HKT75, Sta74, GE85]. Code [AD76, Bro79b, CCC92, EW80, Miy90, ME98, NC76, Vel82, Akm86, Bem74, B– D96, CH60, DP73, Daw73, Mac74, MSS93, PW93, cGR61, Ros64b, War74, Ros70]. codec [Wak93]. Codes [Bro74c, Bro74e, Lle83, War86, Gob75, Lle85, RM89]. Codewords [Mak91a]. Codiagonal [Wil58c, SF61]. CODIL [Rey71b, Rey71a, Rey90]. Coding [FH97, Pik81, TM95, WD99b, Ber59, Bra63, Col63a, FK90, FN93, FNK94, Ker63, Pai60, PG82, Tee67, YAH80]. codings [Mäk91b]. Coefficients [Han70, Cha67, Coo67, Gen69, O’B59b, SS63a]. Cognitive [VVW96, BWM88]. coherence [TF96]. Colin [Cha73]. Collaboration [VCB93]. collaborative [Ski96]. Collection [BS73b, FN78, MS91, Wil70, BPSS72, Hay79, NSW87, Tem72, Tur78, Wri61a]. collections [BKMK90]. collector [Weg72]. College [BF71, Pri64]. Collision [Sam81]. collocation [Wri64]. Coloring [Joh74]. Colour [Er84c, FS87, Er84a]. Colouring [DB81, Sel88, Wil85a, EL89, WM84, Woo69c]. Colours [Wil85a]. column [Tew67]. Columns [Boo67b]. COMA [DSB97, HLK94]. COMA/CD [HLK94]. Comandos [CBH^{+}94, KMR94]. COMAR [HKPW89]. combination [Phi74, Tee63a, Vin59]. combinational [Ben72, CDHL93]. Combinations [Akl87]. Combinator [Bur87a, TL92]. Combinatorial [FM83, Tak99, CRS^{+}93, Lyn65]. combinators [Cla94]. combined [Lar64]. Combining [Daw73, TCM94, THR82, Wie95]. comes [Mus67]. Comm [Gol73]. Comma [Lle83, Lle85]. Comma-Free [Lle83, Lle85]. Command [MA89, MMG91, Tes77]. Commands [Sch77b]. Comment [BK83, Coh01, Mäk01, O’K85, TAC01, Rei81, Wil82a]. Comments [Bre81, Deu69, Mic81, Sab72, Fra70, Kin67b, New61]. commerce [Bag61b, Dou58c, Dou58b, Dou58a, Dou59a]. Commercial [GS73, Gun70, OBLP75, Cut59, Gri66, Ham81b, Lew65b, Pai65, Pla61, d’A62, Gol59, HSP66]. commissioning [Ano60]. committee [Smi63]. Committing [LPSC99]. Common [MJ75]. communicating [Kel98, Nat86]. Communication [Aky92, Bru92, CGRJ92, LF97, Nai97, PB86, Pri78, Sil81, BBP98, Bri67, FH77, JBD73a, JBD73b, Mar82, PR64]. Communication-Minimal [LF97]. Communications [AM99, BB80a, Par88, Pea63]. Commutative [Eke95]. Compact [Sto86, Jen66, Sim69]. compactifying [Weg72]. Compacting [FN78]. Compaction [Bre82, Col87, HW67, QY89].

Comparative [Cla77a, Laz81, MB96, Ole91, OKMS96, Val74, Wil91b, Fin93, KH87]. Comparing [Min88, LK77, MMG91]. Comparison [AHL92a, Cro75, DF87, EHW89, Jar68, JW82, Lom88, MK82, Nob91, Pra88, Ric88, SL78, TC92, VW96, AL72, BML94, Box65, Box66a, DGM68, GGS89, Goo61, LMS73, MD71, RD80, Tan78, Wic82, Wie95, Win69, ZH93]. Comparison-Based [Lom88, Ric88]. Compatible [BL90]. Competition [NL90, WR67]. Competitive [Rom99, HM96]. Compilation [AMN81, Ber87a, BR90, Els79, Smi83, BY89, Gea64, Rya66, Wil89]. Compile [Bur72, CZL97, DT96, Hex67, Mux73]. Compile-Time [Bur72, CZL97, DT96, Mux73, Hex67]. Compiler [AD76, BL90, BW78, BMR67a, Der79, Mil70, Paa91, PB74, SS81, AJ89, BY89, BL77, BHY89, BMR67b, BC67, Cam78, CMRW72, Cor65, CGH^{+}68, Cuf72, ELR78, Fin68, For65, Jon65, MR67, Per76, Ryd64, SA73, Sch65, SC87, Ste74, Tay68, BMR67a]. compiler-compilation [BY89]. Compilers [Bur72, Mux73, CP62, Wic72]. Compiling [Hus61, B– D96, PR77, RL72, Wis68]. COMPL [FS66]. complainants [Hal72]. Complete [Pag81, Bra94, Def77, WC94]. Completeness [CL99b, FZ94, Mer90]. completer [Zar80]. completion [Dic91, Sal87]. Complex [BGR86, Boo68a, Boo68b, KBTB97, LZ95, RL75, Rud67a, Rud67b, SMY95, ZBS81, vtv92, BM80, Che65, CL66, Gui68, HM80a, Jar70, Lar64, LQ96, Mak68a, Sea69, Str59, WM95, Wri63]. Complexity [BR86, CCD86, CCS87, FvLMS98, GV99a, GV99b, MK85, Pra84, Ree84a, Ris99b, She99, VG99, V’y99, WD99a, XY90, tHO99, Cuc93, HRS93, Nav87, Und63, Weg93, O’K85]. Component [UVS87, TH82]. Components [Fro85, Cha97]. Composite [FJA80, McL85]. Composition [CR97, SS96b, Tre70, Gil63]. Compositional [Roj95, RL89]. Comprehension [WB99]. Comprehensive [Wat79, Sol60]. Compressed [AF88, Woo84, CMW98, Lan65, Mul87]. Compressing [WZ99, FN93]. Compression [APD99, BD82, CH87, DdS98, FM83, FNK94, FH98, GW73, How97, Lea78, Lle87, MJ75, MW97, Mof97, NMW97, Pik81, SH97, THHS93, TWMG93, Wel72, Yok97, BM90b, BM89, CC93, DP76, Fen96, Fra67, HM98, JM96, Mak89, Rai87, TR93, Tur75, WBM^{+}94]. computability [Goo85a, PR98]. Computation [BY79, BB69, Er83c, Han62, Her61, Mar94, Rac65, Sto73, WY71, BL72, BW69b, Bra76, Che95b, Cra66a, Cro75, DP66, Duk66, ET75, HK73, Jon66, Lew65a, ME88a, NPW68, Roh75, Sch94, Tak96]. Computational [Alb71, BT63, Coo75, FFP99, Hon93b, MS70, Pau80, WS73, Wil70, BS75, CL78, DF90, FM66, HD71, Jac71, Jam89, Pap68, Per65, Ral61, Tew68]. Computations [AM99, JW74, JJ86, Coo66, HH94a]. Compute [Bra85, Bra86a, Rya66, SS96a]. Computed [Wil61b, Loi68]. Computer [AE90a, AM91, Arm78, Bag62, Bei72, BL81a, BN78, Bur66, CF81, CD81, Cav83, Cha75, Cha78, Cha81, Coh01, CW82, Cor79, Dak73, Daw74, Dou58a, Dou59a, Eas76, Eme80, EM81b, FSWL82, FR80, Fle75, FJA80, FS81, GK84, Geo66, GG86, GS73, Gol69, The91a, The91b, HR73, Hug61, IHH95, IL74, JP76, JJ86, JR81, Kin69b, KTW^{+}73, KP79, LW65, LW66a, LV79, Laz81, Leh73, Li98, Mäk01, Mic97, MN90, NLW74, O’B70, Pal63, PW94, Pie74, PBR^{+}90, Pro74, RL73, Sco74, SZ74, SS75, Smi64, Smi83, TS64, TAC98, TAC01, TF87, Tur92, UHS85, Var76b, WJ74, WS73, Wat74a, Wat74b, Web72, Wil91a, WW79, Wil90, WCH78, Woo82, Zmu75a, Zmu75b, AS63, ABN61, Bae66, BBG^{+}71, Bar58, Bar65c, Bar67, Bea68, Bec58, Bel70a, BS61, BD61]. computer [BB71, BCR72, BP82, Bri79, BB59, Bri67, BF71, Bro58b, Bry59, Buc60, Buc65, BSH70, BK83, Cap86, Car64, Cha62, CP79, Che62a, Che62b, Che64, CF96, Cle60, Coa74, CW71, Col64a, CBB^{+}76, Com75a, Com75b, Coo60a, Cor65, Cut59, DN64, Dol66, Dou59c, Dou65, Dow71, Ede63, Er86b, FFMR66, Fin60, Fir61, For66, FGR67, Fos64, FdVR58, FL68, Fox78, GP73, Gil63, Gil58c, Gil58d, Gla69, GGH^{+}87, Goo61, Gos63, GB61, Gro68, oH59, HM80a, HK67, Hea81, HK59, Hil67, HB58a, HB58b, HM80b, HG67, JC88, Jam81, Jan63, JN65, Joh61, Kec87, Kni68, KH81, Lan76, Lan65, LR67, LEN63, Lew65b, Loo59, Mar73, Mar74a, Mar79, MR85, McG77, McO66, Mea73, MP65b, MFM64, MW68, Moo59, MTPL88, MES97, ML65, OB68a]. computer [Ong69, Pag62, Pal62, Pan70, dP60, Pen66, Pla61, PW67, Red60, Ric73, RWT64, Rig64, RW61, SW68, Smi66b, Smi68, SS72c, SS72d, SW91b, Sut60, Tay61, Tho61, Tho62, TB63, Tur79, WWN91, War74, War60, WY62, Wat82, Wil58a, Wil61a, Wil67a, WN68, WO75, Win59a, Win59b, WH68, WN81, Wri59, YR58, You63, HL76b, LM88, VW95]. Computer-Aided [Kin69b, Wat74a, Gol69]. Computer-Assisted [Smi83, LM88]. Computer-Based [FSWL82]. Computer-Oriented [Bag62, ML65]. Computerised [BD81]. Computerized [F^{+}73]. Computers [All75, BD61, Ben85, BHRS69, BHRS70, BS79, Dou58c, Dou58b, FRS^{+}87, GS73, Gol73, GT63, HRS68, KV62, Mas63, Pap60, Pla64, RP63, Tre70, WY71, Yat62, Aga96, AB78, Ben70, Cha63, Cle67, Col63a, Cra67, Dou61, Edm60, Gal59, Gea61, Gol59, Gow62, GG59, Gri66, Hil77, HP77, JI93, Kec87, Mar58, Mel66, OS65, dP60, Pin61, RW68, cGR61, Sow87, Sym71, Tak59, Tho68a, Vin59, Bro58a, Dou59b, Web70]. Computing [Abb80, Aga96, BRH70, Bar78b, Bow81, Can93a, Cla77a, Col82, Er83a, FSWL82, FKA^{+}93, Fis74a, FLW99, Fox61, Gar78, GS78, Ham93, Kit74, Kum99, LCH99, Muc82, PTV97, Par75a, PCSG63, RS88, SB70, Sla66, THR82, VW95, Wat81, WN86, Ahm73, Akk73, Bag61b, Bur87b, Cam97, Cha93a, Cha97, Cla60, Cox70, EAL66, Er82a, Fin93, Fur92, Gen69, Gil65, IWT73, Lai94, Lau75, LC93, Noo84, Nud69, Pit65, PL68, Pri64, Pro83, SC71, SM94, Stu71, Tho60, WM69, Wol90, Woo68, Bel74, Bro73a, CF81]. Concept [Eri74, Sar96, SA81, BPG87, Fun84, HCH96b, SL73]. Concepts [FKT90, Fro85, Hou75, Bra76, HS82]. Conceptual [AM92, BK75, FL85, Fro83, Sha81, TLF96, WB98, Min88, SC87]. concerning [WLDC71]. Concrete [BBK87]. Concurrency [CL99a, FKT90, HM84, IP86, LH95, PPB84b, Wet81, GF91, Mad97, Rv88]. Concurrent [BIY88, CG92, GP90, HW89, Hul86, KS85, MMR^{+}98, PBF91, Rom99, Shi85, Ski91, Tho93, Wat91b, Bur87b, KL89, Pon91, Shr75, Wal77, KS82, Sil79]. condition [BS73a, Gra87, HM95b, Zaf71]. conditional [Tur79]. conditioned [Oom93]. Conditions [Goo77, Man74, San75, Wil79, CP66a, Has61, JM68, KM66, Kin69a, Osb69, Sne70, Tay70b, Twi75, Wol71]. conduction [Bir68, MP62, Osb69]. Conference [Ano75a, Coo89a, Dak73, Ano73c]. configuration [AM65, MBC^{+}86]. Conflicts [LPSC99, Mad98, TK95]. congruence [Tho58b]. congruential [DR67, Ful76, MP68]. conjectures [Cut59]. Conjugate [FR64, Cla77b]. Connected [UVS87, Cha97, WWN91]. Connection [Bra88]. Connectivity [Wil91b, BP79]. Connolly [Her61]. Conquer [dHM93]. Consequences [LM77b, Wil91a]. conservative [RM69]. consideration [Hil67, Hur79]. Considerations [Hel93, cGR61, Wri83, AB84, Bon69, Gil66, Pag60, PGM84]. considered [Lip80]. CONSIM [Nel79]. Consistency [BKK84, Flo74, HKT75, LW85a, Lev96, Che90].

Consistent [Cho99, UHS85]. consisting [Wil68b]. console [PL68]. Constant [BKMT96, Hor93, Mul85, HL92, Jar67b, O’B59b, SS63a]. constants [Tee67]. constituencies [HW74, Par74, Van73]. constituent [JC86]. Constrained [Mur85, PS90, Wri82, BJS74, Box65, Box66a, Bra75, Gui68, Hon93a, LH95, Oom93, Tho72]. Constraint [Bol97, Orm91, CRS^{+}93, FG90, Imb93, YG93]. constraint-driven [YG93]. Constraint-oriented [Bol97]. Constraints [Cha93b, CY96, Orm88, Sho81, vtv92, AJ70, HB70, Ord64, PPB84c, SWA97, WTW91]. construct [Lo84]. Constructing [FL89, Fro92, GS79, YAB92, Alm66, EL82]. Construction [EA83, LO75, Roj95, SS80, And66, Bar65c, BF71, BS79, Che95d, CO66, ES77, FL68, GV91, How67a, JS68, Lin72b, Lle85, MHF93, OS65, WP75, WL72, YG93]. Constructive [Nip91, Wal76, Ric73]. Constructs [DE86]. contact [Abb70]. Containing [KC97, Scr64b, WC85]. Contemporary [DL97, FSW85]. Content [AA99, CEM72]. Content-Addressing [CEM72]. contents [PW93]. Context [AMN81, FH98, MW97, Ram85, TWMG93, EDS86, Gha72, Goo75, LL70, SEP89, Woo70a]. Context-Based [MW97, TWMG93]. context-free [Goo75, LL70, SEP89, Woo70a]. Contexts [CT97]. contextual [BBG95]. contextually [EM71]. Contiguity [Mur85]. Contiguity-Constrained [Mur85]. Continuation [SL78]. Continuations [All88, All90]. Continuity [Man74, San75, RV93]. Continuous [LS93b, BM59]. Contour [Cra72, RB86]. Contouring [Sch82a, ST81, Suf90, Sut76b, RB86]. Contours [McL74]. contributions [Dru63, Pol60, Str60]. Control [Bar74, BPR74, BL62, CW96a, CL99a, ER86a, EM81b, Fra76, Ger70, GP88, GT63, GN83, HR73, Hou63, IP86, KS82, KP79, LH95, Lei62, MK85, Mil70, Nel79, PPB84b, THR82, TF87, Zmu75a, Zmu75b, tHO99, Bar72b, Bri67, Bry59, Cle66, Com75a, Com75b, Dak75, Dou58c, Dou59a, EMM72, For66, FGR67, Fra68, FK86, GF91, Gol78b, GE69, Han94, HK59, Hon74, Hug61, HSW92, JP69, Ken80, Lam63, Leh73, Mad97, MT83, NI70, O’S88, O’S94, Osb71, Pal62, Rv88, SP82, TS64, YAH80, Den80, SS63b]. controllability [GE69]. Controlled [YS91, Dow71, KK79, Pri77]. Controller [SA81]. Controlling [O’S88, MP95]. Controversy [WLDC71]. convention [Ros63b]. Conventions [HH80]. Convergence [Bar61, Bre71, CWH93, Loi72, Wat68, Wil65c, Aha65, CP66b, Erd65, Eva65a, GL70, Jam87, Lyn69, Scr70, Sto68]. convergent [FP63, MS70, PM72]. converging [Jam89]. Conversational [Nel79, Bar71, Col66b, Cuf72]. Conversion [CCD86, SW83, Ste86, Tiz60, WO78, BP94, Boo74a, Boo74b, Ham74, KJ75, Par89]. Converters [AM90, Par92]. Converting [SS75]. Convex [BH69, GS79, Col66a, Cox71]. Convexity [YFM71]. Convolution [McC78]. Cooper [RS69]. Cooperative [BBF^{+}94, MR96, Rom99, SBRS94]. Coordinate [Cra72, Ter73]. Coordination [DRT96, Gee91]. copies [KH81]. Coping [BR86]. Copper [Coc97]. Copying [Har81]. Copyright [BM94b]. Core [BGMW96]. Corecs [PPB84a]. Corporate [Ham80a, WJ74, Car86]. Corps [TS64, Sla60]. Correct [BPTV92, Bra80]. correctable [TB86]. Correcting [Bro74e]. Correction [WT84, And75, ES75, How75, Sub85, Ull77]. corrections [Sac62]. Corrective [JSKCG94]. Correctness [BG94, B– D96, LO75, Lew84, Sch77b, Bak94, MP89, O’S88, PJ86, Smi72]. corrector [Ano80b, For61, Hal67, Ral61, Sne80, Wes80a]. Correctors [Rod70]. Correlated [Wal64, Wal98]. correlation [Jac71]. Correspondence [AT70, All71, And69, And74, And75, AMKM66, Art75, Bac71, Bae66, Bar72a, Bar64, Bar66, Bar73, Bar78b, Bau65, Bea74, Bel74, BM81, Bem74, Ber69, Bic69, Boo74a, BSPI65, Box66b, Bra75, Bra72, Bre81, BVH74, Bro74d, BH69, BJ80, BBdD66, Car70, Cha72, CL66, CT74, Col67, Coo72a, CG65, Coo72b, Cr081, Cra67, Cro75, Dak73, DD70, Daw74, Deu69, Dew72b, Duk66, DW73, Dun78, DI78, Ecc71, ET75, Ehr72, Eva66, ES75, FT73, Fel69, Fel70, Fel72, Fin72, Fis74b, Fis74a, Fla72, Fra70, Fre74, Gal71, Gal73, GL70, GT72, GW73, Gla73, Gol78a, Gol78b, Goo68, GB69, Gri92, Hal72, Ham74, HL70, HT81, HT78, HV74, Hay79, HLOS65, Hig67, Hil67, Hil72, Hil80, Hit78, Hon71, Hop78, How75, HSP66, Hun81, IWT73, Jac79, JL80]. Correspondence [JD59, Jon71a, Kei79, Ker63, Kin71, Kio81, KTW^{+}73, Kli75, Lan71, LWS69b, Lee74, Lew72, LS78, Mac74, Mar74b, Mee72, Mer70, ME80, MR82, MT72, MR65, Mul75, Mux73, Myl66, O’B59a, O’B68b, Obr67, Osb71, OL68, Pak67, Pal72, Pal73, Pap68, PH73, PB69, Pit74, PS74, PW74b, Red71, Ree80b, Rob80, Sab72, Sal73, Sam63, Sam66, San75, SA79, SS72a, Scr72, Sel72, Sil72, Sne80, Sol68, SC71, SB81, Str74, Tay64, Tay68, Tay80, Tay81, Tee67, Tem69, Tem72, Tho80, Tho68a, Tho68b, TT73, Val71, Wal71, Wat82, Wel70, WNT73, WS74, WP69, Whi72, Wig81, WR72, Wod78, Wri66a, dD66, Bar69, Ben70, Bir81, Bow73, Bro73a, Chu68, Gui68, Jar70, Kin67b, Lem80, O’C69, Sam68, Var67, Wil67c, WF78, Zar80]. COSEC [PL68]. Cosets [PPB84a]. COSMOS [AA99, O’B67]. COSMOS-Content [AA99]. Cost [BK92, CN82, Lea86, Par75b, ABBS83, Bon69, PS92, Sym71]. Cost-Effective [Lea86]. Cost-Effectiveness [BK92]. costs [Chr73b]. Cotes [Sac62]. Could [AC84]. counter [Col64b, Sow87]. Counterfactual [NBL95]. Counting [Axf90, HW74, Lun75, MS91, Par74, Van73, Akm91]. countries [Noo84]. coupled [AHM74]. Course [Gal92a, Maz78, NQ80]. courses [Bro73a, Fis74a]. Courtauld [Cra66b, O’B67]. covariance [Goo61]. cover [RHS89]. Covering [SM88]. Cowper [RWT64]. CPL [BBH^{+}63]. CPL1 [CGH^{+}68]. CPM [Ste77b]. CPU [ET75, LY79, MMB87]. CPU-bound [MMB87]. Crank [Dav67, KM66, dD66]. Crank-Nicholson [Dav67]. Crash [Gal92a]. Creation [Sin97, Noo84]. Creations [HV92]. Creed [GS63, Sam63]. CRESTS [Cra66b]. Criteria [DW88, KH81, Lom88, Rze85, Wil78c, GL70, Lyn69, Par78, Pow67, Van73, WLDC71, HW74, Par74]. Criterion [Orm91, Tse87b]. Critical [Mae78, Wic92, BM92, DGM68, DCHR76, Mil95, WM95, dSA92]. Critique [Nau79, Par78]. CRL [BG94]. Cros [KPP^{+}95]. cross [BL77, Jac71]. cross-correlation [Jac71]. Crossword [Ber87a, BY89, BR90, SS81, Smi83, Wil89, SD86]. crosswords [WW79]. Crout [Tew69a]. CRT [Gol69, Rep72]. crunching [GIB90]. Cryptanalysis [GN94]. cryptic [SD86, WW79]. Cryptosystem [LCL92]. crystallographer [Mat67]. CSL [Bux66]. CSP [Hul86]. cubature [Mei66]. Cubes [KP98, QW96, LOKM99]. Cubic [AH69, Bic68, Bic69, For75, Fyf69, Fyf70, HK72, Pha90]. culture [PRLW93]. cumulative [Not70, Smi66a]. Currency [LW85a]. Current [Cha81, Gil59, JJ67, Tri80, d’A62, Box66a, Coo83, WHF82]. Curricula [Ham93, Fin93]. Curtis [OS68, Oli72]. Curve [Cle60, MZ86, Sut76a, Ada69, CW61, Coo72b, Ecc71, Gro71, Hay69, Le 69, McC70b, McC71a, McC71b, McC72, OSS69, OP71, SM71, Sab76, Sne66]. Curve-Smoothing [MZ86]. Curves [BT81, Col87, FS87, GM83, Man74, McL78, San75, BH69, Kei79, Kiz66, QY89, Sab72, Wri66a]. Custom [HAAZ99]. cutting [AO80a, ILPP98]. Cweb [Thi86]. CyberCrime [BGR98]. Cybernetic [FE80]. Cycle [DM83, Had90, Har84b, PB76, PB78, Ros82]. Cycle-Sort [Had90]. cycles [Hun60, Ros82]. Cyclic [Axf90, Bro74d, Er84b, Hil79a, Pag84, HK73, Ged96]. Cyclide [McL85]. cylindrical [McC93].

D [FSWL82, IL74, JR81, NLW74, NPCA81, PCSG63, RL73, Ris99b, VP80, Wil67b, Wil70, ILPP98, PW94]. D. [FSWL82]. Dahl [Llo74]. Dakin [Gla73]. dama [Fin60]. damped [Til69]. Daniel [WY71]. DAP [CHPS85, DM85, KB87, SHCP91]. DAP-Algol [DM85]. DAPLEX [PG90]. Dashed [Ber76]. Data [AE90a, AK81, BB80a, BIY88, Bel84, BY79, Bir81, BHW78, Bow84, BL81b, BP81, CS77, CO76, CCC92, Che95a, CW96b, Cle92, CH87, CEM72, Cro74, Dav78, Dav81, Dav63, DW88, DdS98, Dea74b, Dea79, Dea82, EMM72, Eri74, Eth71, FM86b, Flo76, Fra67, GS73, Gol85, Gor78, GP90, GKS79, GL94, Ham77, HK92, Har76, HW77, HC78, HAM98, HM84, IS81, JBD73a, JBD73b, Kin69b, KSD77, KA86, Lea77, LF97, Les83a, LW85a, Lle87, Llo64, LT61, Mai85, MS92, McL74, McL76, Med70, Mic91, Mir98, MVL92, MB96, MSS93, NLW74, NPCA81, NF94, OD91, Oka98, Orm91, Pad73, Par82, PDY74, PB86, PAW86, PSD^{+}81, PPB83, Pri78, PJ80, Ros70, Ros93, Sad92, Sar90, Sch82a, Sha95, Sho81, Sou76, SW79, Ste77a, Str72, SK95, ST82]. Data [THR82, VP80, Wal98, Wat91a, Wel80, Wil61a, Wil72, WKRJ60, Wri61a, Yok97, YP86, YY89, Zah81, AF76, Ano63, BM90b, Bak94, BJ78, BM89, BFH61, Bre93, BPSS72, Bur79, CS75, CC93, CLW98, CPR75, CCL72, CBB^{+}76, Dac63, DP76, Dea74a, Dee80, Don78, Dou59c, DM60, FCHK96, FE69, Fri65, Gat65, Gil64, GM72, God96, Ham81b, Ham91a, Hay79, HKPW89, Hin60, Hun59, Hut86a, Inc87, Jar70, JL80, Ken85, KGC85, KH87, Kol83, LW67c, LW71, Lee78a, Lic89, Lon61, MG85, Mak89, Mal59, Mar73, MP79, Mil81b, Min88, Nau79, NS77, OP71, ON94, Par77, Par78, dP60, Pin77, Pit65, PS74, Pro80b, QY89, Ree80b, Rob79, RB86, Rus63, Sch76, Sla60, Smi63, Sol68, Spä68, Spu88, SD73, Str61, Sub85, TCL89, Tem72, TLF96, Tho60, Til69]. data [TP89, Tur75, Tur78, Var79, VV94, Wil61c, Wil63b, Wis68, WS86, ZR88, SS63b]. data-compression [Mak89]. data-display [Mar73]. Data-Flow [PB86]. data-processing [Hun59]. Database [ABM88, BSH90, BJ98, Bow84, CN82, CCL87, Dav76, Dea78, DNV81, DAODT85, DBCRW92, HKT75, IP86, JR81, KU99, Ker76, KP92, LPSC99, LS97, LW83, LCL92, LB92, Mai85, Mar92, Orm88, Pou92, PF86, Rob75, She85, Sta74, SMY92, YCC92, Aki88, Azm90, Ber87b, BMT85, Bro82, Che95c, CMW98, Dea76, DTIR88, ET88, Fun84, GF91, GE85, HK82, HS82, JO74, KM94, LB91, LC95, MB90, Par75d, Rv88, SS96a, Seg87, TU96, WWK^{+}90, WS86, YAB92, ZR88]. database/administration [Aki88]. Databases [AE90b, BMSA88, BCV93, Bra86b, Bra88, Col97, Flo74, GWWB98, Kie88, LH95, LW85a, Mar91b, Pou93, Sch83, BM90a, Car86, CAC82, CG99, Dee85, FW83, JKW94, KS91, KCP99, LSV89, LL90, Mad98, Sad92, WN88]. DATAFAIR [Lav70]. dataflow [PK85, Sow87]. Datagram [Pri78]. datalog [AGNS91]. Date [Ran74]. Date-Sensitive [Ran74]. dates [Hil77]. Davidon [KL71, Lil70, Lil71]. Davies [PCSG63, Hos71]. Davis [Eme80]. Day [Cha73, Var67]. DBMS [CS85, PPRB89, PPB84c]. DBTG [HKT75]. Deadlock [TF87, BN76, Che95c, SP91]. deadlock-free [SP91]. deadly [Lle73]. dealing [Bro60]. Debugging [CG92, JP69, Pie74, Bar69, Bul72, KK79, Sil72]. decade [Wil58a]. decay [Bra94]. Decentralising [Vau95]. Decentralized [Li98]. Decidable [CHM94]. Decimal [And74, And75, Bro74c, Bro74e, SD82]. Decision [Har95, Kin67a, Lew84, Mae78, Sch85, SS75, Van91, Ben70, Car70, Gri68, Kin69a, Kin71, KJ75, LW66b, MV88, PW74b, TB82, Wym73]. Decisions [Coc97, Par75b, AB84]. declarative [BMRS90]. Decomposition [Den80, DK74, DD91, FW86, Nas75, NS87, PG81, BH68, BHS65, DĎ88a, HM95a, Lai94, McC93, O’B59b, WV70, Wil82a, Zam70]. dedicated [CP79]. Deduction [Pau85]. Deductive [BSH90, Sad92, WWK^{+}90]. Default [KS91]. defer [Sri91]. Deferred [GlM68, LM77b, May64]. Defined [Gol85, GN83, Har84a, Har76, Has75, Pow67, Wat73]. Defining [Mum85, Mar74a]. definite [CP71, JB94, Jen71, MT61, MP65a, MP66, OS69, Oli71, Val71, Zam70]. Definition [Fro83, Geh83, Kin69b, MM84, Pit74, PPB83, ST82, Bar72a, Bar75, BJ78, Cam78, ESK95, Gal87, Gol78b, Hou69a, Lee72, Thr69]. Definitions [Pag76, Pag81, Bur79, PS92]. Degeneracy [DW73, Sch94]. Degradable [Che95a]. degrading [CPV93]. degree [Bro95, Cox71, MF88, Sco69, Spä69a]. Degrees [Mah86, Atk93]. Delaunay [Wat81]. Delay [Bas86, GE69, War74]. Delays [CD82, KV63, RW68, Tho68a]. Deletion [CV86, Ull77]. Delphi [Clo87]. demand [Wil78b]. DeMarco [TP89]. DeMasi [HR73]. Demonstrating [PW93]. Dendrogram [Roh74]. Denison [Wre71]. Denotational [All83, MA89, Pag79, All85]. Department [GG86]. Dependable [BL91, DJ95]. dependence [WCL71]. dependences [JC97]. Dependencies [Bra86b, Bra88, HKPT99, Jaj86, HW88, Sad92]. Dependency [CW96b, CAC82, O’S88]. depending [Tom67]. DEPOT [LMP79]. Depth [FK93, Cha90, Er89a]. depth-first [Cha90, Er89a]. Derivation [Dro87, Roh87, WW82, FCHK96]. Derivative [MM70, All66b, KM66, Mil60, OP73b, Scr64b, Zaf71]. Derivatives [Fle65, Pow64, Pow65, DW73, KL71, LM63a, Lil70, Lil71, MM71, OP75, Tay70b, Zan67]. Derived [Har95, NE67]. Derrick [Eme80]. descent [FP63]. described [PJ89]. Describing [RBCM97, SWB87, JC86]. Description [DBH85, MM78, Mot79, Mot80, Mot85, Ree67, Zve75, BG79, God96, HPS61, ML65, Nau93, Sau72, Uck85, Wat91c]. Descriptive [BS86, GGS89]. descriptor [All82]. Descriptors [BB81, IS81]. Design [ADK^{+}93, AB90, AM98, ACDP91, BML94, BW69a, BK75, Bri80, CF81, CS77, CD81, CS85, Cla92, Das91, Dav81, DNV81, DMW91, FR80, GM83, Gol82, GP81, Ham91b, HS83, Hel93, HAAZ99, Hil76, Hol75, HM84, Hut86b, Jel79, JK80, LS93b, Mac70a, Mac81, Mil81a, Moo79, NM90, NPCA81, Ole91, Orm91, Pai99, Par75b, Par82, PS96, PCSG63, Ran86, Rep72, Rob90, Sco74, SWB87, SBRS94, TL85a, TL85b, UC90, Wat74a, Wat74b, Wel80, Woo86, vFV92, AL72, BH68, BWM88, BD61, BBKR89, BMT85, Bur66, BKK78, Cam69, CS75, CCJ91, CS95, Cur95, Dem60, EH88, Eth71, FKLM76, Gea80, Gri62, HB58a, HB58b, HO93, JC86, JL80, KL77, KGC85, LBC^{+}94, Lan76, Lar95, Lav64, Lew71a, LC90, MV88, MT86, Mum84, New66, Nud77]. design [OS65, Pen76, Pen77, Pin61, Ree80b, RB85, Ros82, Rus86, Sal87, SI80, Teu87, TAP89, Vin64, Wat72a, Wat72b, Wil67a, WN68, Wis64, ZD71, Le 88, vSHBT98, BB80a, Eme80, FJA80, HR73, Her61, IL74, JR81, Var80a]. design/drafting [CS75]. designational [Wat63]. designed [Aus71]. Designing [Nai97, WWK^{+}90, Gri66, Kol83, LQ96]. Designs [Edw78, Clo86]. Desirable [Pri64]. destination [Mit89b]. Detail [Sch77a].

Detailed [LW83]. Detecting

[AG98, Jan88, LC98, VW96, YT79, DF90]. Detection [Bro74c, EW80, WT84, WC98, Ben72, BD80, Che95c, Com75b, Cze88, Sal87]. Determination [SC81, Can93b, Car61, HK82, SE82]. determines [Sal74]. Determining [BF78, NV77, vtv92, Boo67f, Bye67, Eva75, Hil69, Lan76, Rya66, Woo70g]. Determinism [SS92]. Deterministic [CC74, Hie98, JB94, Woo71a]. detonation [Pal63]. detranslation [HM80b]. DEUCE [BB59, DH60, Rob59, Red60]. Develop [Nob91]. developed [Bro58a]. Developing [Sel88, BWM88, MR96, Noo84]. Development [BL90, BPTV92, BTD99, Dea82, DM83, DM85, EMCO92, The91b, Har80, HOH91, Nob91, Tak59, Tre91, WB99, AFWH88, AWH91, BM92, BM94a, BD72, GK96, oH59, Har88, LLR73, Lum64, Poo91, PL68, RB85, RHHR98, Ros82, SC87, Smi64, SW91b, Tyl88, Wil58a, Wil67c]. Developments [MW97, Smi70b, Ano73c, d’A62]. device [FS94, Fos71]. Devices [Sam81, Pez76]. devolved [Car86]. Dewey [Foo86]. DFD [GGS89, TCL89]. Diagnosable [Lom88]. Diagnosis [BZN98, CH99, BH63, BL71, Ler69]. diagnostic [Pan70, WL72]. Diagnostics [Bur72, Mux73, PB74]. Diagonal [Wil62b, Bak64, Cay61, CL83, EF63b]. Diagonalisation [Sea69]. diagonalization [RW61]. Diagrams [DP88, Har95, Wil77b, WO78, Min88, TP89]. Diakoptics [Lai94]. dialects [HB67]. dialogue [BMPS71, HS82, RB85]. Diameter [KP98, MF88]. DiCE [VCB93]. dictation [Sha94]. Dictionaries [FM83]. Dictionary [GN88, GT63, Kin69b, MVL92, Nar96, Zah81, MG85, Spu88]. Difference [CM79, BW71, Eva63, Eva65b, Eva66, Had71, KL71, Lil70, Lil71, LS93a, MP62, Osb63, Osb64a, Osb64b, Osb65, Ran68a, Tee63b, Wil76b, Zaf71]. differencing [ME88b]. Different [Sch77a]. differentiable [BH69]. Differential [Alb71, BD81, Nor76, SB70, SS72b, WY71, All66b, BWZ71, BS71, Cay61, Chr73a, CN63, CG65, Dav67, Day65, Day66, Day67, Eg69, Eng69, Eva65b, Eva66, For61, Fox62, Fox63, Gib60, Hai69, Has61, Hop71, JS67, Kar66, KS71, Mak68b, Mar58, May64, Mil60, Nor64, O’B59b, Oli69, Osb63, Osb64a, Osb64b, OM64, Osb65, OW68, PC64, RM69, Ros63a, SS63a, Scr64b, Scr65, Sne70, Tay70a, Tho68b, Twi75, Wal64, Whe59, Wil76a, WP68, Wol71, Wri64]. differentiation [FR63, Hun61, Sch65]. Difficult [Bel84]. diffraction [MP65a, MP66]. Diffusion [LW85b, McC70a, Tay70b]. Digital [CF81, Coo75, DBH85, HB66, IL74, KP79, Lyn72, PCSG63, RL73, RP63, Sto73, WL87, Aga96, AS63, Bar65c, BBS66, Bec58, BL71, Bri79, Cla60, Cle60, Cro75, DP66, Edm60, Fin60, FNK94, GP73, Gil58c, Gil58d, GB61, JN65, Joh61, Lar95, Mar58, Moo59, Mor89, Pit67, Tak59, Tay61, Tiz60, Vin59, Wal64, Win59b, Win75, Wri59, Edw78]. digitised [AO87, Lut80, SHCP91]. Digits [Bro74c, Ber69, ES75, How75, Rea70, Sal68]. Digraph [Bay76]. digraphs [Cha90, MF88]. Dijkstra [Llo74, Gal73]. dimension [Com85]. Dimensional [Ham91a, McL76, Che65, Ken85, Kro70d, Lun71, PHS96, Phi67c, Wat81]. Dimensions [EHH86, HAM98, MM70, YFM71, Bel70b, CD69, Hal68, Hun72, MM71, MLB63, Phi67b]. Direct [Bra63, Gun70, Ker63, Osb65, WD74, All66a, Jen61, Mad80, MMG91, Pak67, Pal69, Pez76, SK76, Spo85]. direct-access [Pez76]. Directed [Bur72, HS83, Mux73, NN97, Sil81, BC67, EMM72, PJ86, PR77, Ree67, Sch65, Wil76a]. directing [NI70, Osb71]. Direction [MM70, Eva65a, FM64, Had68, MM71, MP65b, Rya66]. direction-independent [Rya66]. directional [dS77]. director [Ree75, WBB72]. Directory [Zve80, DTIR88, KVKP95]. directory-based [KVKP95]. diregular [MF88]. Dirichlet [Bow81, FM66, GS78, JM68]. Disc [BK79b, CC87, CCC92, HS71, Huz79, Kie88, Leu83a, LW83, LW85b, PW81a, SR76, Str72, Wat75a, Wat75b, Aus71, CC88, CLC92, Leu82b, Leu83b, Per80]. Disciplines [MK82]. Disclosure [BV92]. Discontinuities [HCC74, O’R70, Whi72]. Discourse [AHL92a, Mar91a, AHL92b]. Discovering [HKPT99, YG93]. Discrete [Gre70, Med70, Sch82b, SKY73, Coo72b, Ecc71, McC70b, OW71, PW74a, PC78]. Discrete-Time [Med70]. discretization [PC64, Wat68]. Discretizer [Wu96]. discs [Dei95]. Discussion [Ano63, Bar69, Bar78b, Ben70, Bir81, Bow73, Bro73a, Chu68, Cla99, Daw99, Gui68, Jar70, Kin67b, Lau75, Lem80, O’C69, Ris99b, Sam68, She99, Var62, Var67, Var79, Var80a, Wil67c, WF78, Zar80]. disentangle [Woo70e]. disequations [Imb93]. Disk [TL97, Ros66, TB94]. disk-based [Ros66]. disparate [Par75d]. Dispatching [Li98]. Display [Van91, BPSS72, Gla69, Jon79, Mar73, Sip69, Tem72, Wis68]. Displaying [NW87, Sab76]. Dissemination [AM91, HT78, Vau95]. Dissimilarity [GWWB98]. Distance [Bra85, Bra86a, PTV97, Cha63, Jac73, LoD71]. distant [HKA72, Hos74]. Distillation [Sto73, Cro75]. distinct [Far78]. distinguish [Ros63b]. distinguishing [KRK73]. Distributed [AE90b, AM99, BIY88, BR91, BHW78, CBH^{+}94, CNTR97, CR97, CS95, Das91, Dav78, Dav81, DAODT85, DRT96, FLW99, Gee91, HPR88, Her91, HCH96a, Hut93, IP86, Jen99, KC97, KLB91, KBTB97, LH95, LPSC99, LW85a, LCH99, Mah84, MS91, MLR90, MWR98, MB96, NH89, NCR97, PB86, RBCM97, RG95, Sil82, SMY92, Vau95, Ver91, BBP98, BPG87, BBG95, Ber87b, BBKR89, BP82, Bur87b, CP79, Cha97, Cho95, CFS89, DJ95, DTIR88, DPT94, ET88, GF91, ILPP98, KMR94, LC93, Mad98, Mar82, Mea73, MB90, MT86, NDD96, Nat86, NPW68, RV93, Rv88, Ski96, SZNM94, TCCH94, Tay90, TF96, TB82]. Distributed-Application [KBTB97]. Distributed-Memory [AM99]. Distributing [DMW91]. Distribution [Har84b, Mah86, Ash68, Bak94, BB69, But61, Com85, McC78, Par75d, PFW74, Ree79c, Ree80a, Ree83, Tur79]. Distributions [ON90, SB85, YA61]. Distributive [McC82]. Divide [dHM93]. Divide-and-Conquer [dHM93]. Divider [PB78]. Division [HAAZ99, PS96, Ahm72, Var77]. Division-free [PS96]. divisionless [Ara94]. divisive [LW71]. divisors [MST64]. DMC [BM89]. Do [Nef65, Bae64]. Do-it-yourself [Nef65]. Document [Mac81, WK77, ZSM99, All82, ATL93, EHW89, Fur89, Hin61, Sv83, Str61]. Documentation [CW82, Kin69b, Vin64, Dou61]. Documents [Lal99, LJ98, Bro89, Mul87, SAHS71, She89]. does [Hut84, MMG91]. Doesn’t [Ver91]. DOLPHIN [Col70]. Domain [Coo75, Lai94, LL90]. Domains [NW74]. Domolki [HR70]. DOPLs [Lee82]. dotted [DF90]. Double [Alp71, Wri82, Che95d, GRB68, Hon71, ME88a, Meg90, OOW91, Sam72a]. double-ended [Che95d, OOW91]. double-length [Hon71, Sam72a]. double-pipeline [ME88a]. Doubly [CS77, Sur86, NO92b, Oli72]. doubly-adaptive [Oli72]. Doubly-Chained [CS77]. doubly-linked [NO92b]. Douglas [CF81, FM64]. Douglas-Rachford [FM64]. Dowe [Cla99, Daw99, Ris99b]. Down [ABD84, MG95, Bar87, Dow69, HW88, Lau91, Woo71a]. Downdating [WB98]. Dr. [Fox87]. drafting [CS75]. Drawability [WC94]. Drawing [Ber76, Gri73, McL74, Smi71, Sut76a, Pit67]. Drawings [PT90]. Drawn [Les83b]. Driven [Pad73, FE69, SEP89, YG93]. Drivers [MH94]. DROL [DPT94]. drown [Mil95]. DSA [BS86]. DSTL [Edw78]. dual [Kro70e, Tur68]. Dubner [Dur74]. Dumping [Fro81, KV62]. dumps [Bro73c]. Durand [Wil62a, Wil63a]. duration [LC95]. during [Aus71, Rya66, Smi66b]. duties [WR86]. Dynamic [AB84, BGR86, Bul72, Che86, CTH99, CH87, Gee91, GF91, GS73, McI82, Mul85, Pfl84, RL82, Sil72, TUW98, TW77, TCB71, Wan80, Abb70, Aus71, BMT95, Cam71, Che95c, CPR75, DC84, Dav84, DO72, Fro81, IJ62, KM74, Lan65, Leu82a, MW68, Par75d, Ree84b, Sla64a, Wie95]. Dynamically [NR79, Dav82]. dynamics [NPW68, Wil73].

E.D.P. [Hic60]. E.S.P. [Wil64b]. each [Boo67f, Hil69, Woo70g]. Early [Cor62, The91a, Hic60]. earthworks [Oxl89]. Easter [Boo79d, Hil79c, Ran79]. ECLAIR [HW92]. econometric [BS75]. Econometrics [TCB71]. economic [BS79, Gal59, Mor89]. Economics [KV62, Sla71, Web70]. Economy [Wol78, Sla66]. ECOOP [Coo89a]. ed [CF81, FSWL82]. edge [YJ94]. edge-solvable [YJ94]. Edged [WW82]. Edinburgh [Ano75a, BMPS71, WW73]. edited [Eme80, JR81, NQ80, NPCA81, PCSG63, RL73, RP63, YFM71]. Editing [AMN81]. Editor [Par67, RW62, Sla71, TCB71, Str65b]. Editorial [Ano76, Ano78, Ano79a, Coo92a, Fis74b, GV99b, Gea73, Ham73, Ham81a, Ham84, Ham93, Her91, Hon93b, Hut93, KP97, Kum99, Mel95, Mof97, NF94, Pag73, SK93, TT73, Wat91b, Wat91a, Bar73, Gal71, GT72, Ham80a, MT72]. edits [SAHS71].

EDSAC [BH63]. EDSIM [PC78]. EDSIM-event [PC78]. Educating [PS93]. education [Osb93]. EFDM [KA86]. Effect [RM89, GL70, KV63, Kol83, LC95, Lyn69, Mal61]. Effective [CH99, Lea86, TI75]. Effectiveness [BK92, Ben70]. Effects [Kuh92, Sv83]. efficiencies [Jam89]. Efficiency [BR90, Fle75, Hil67, LW83, TL98, Ruž91]. Efficient [AM91, BDB92, Cze88, DS90, Dur74, EA83, FH97, GP90, Ham91a, Har81, Har82, Has75, HAAZ99, HKPT99, Kle97, Mak91a, Man93, MSNW99, MB90, Mul85, Øst79, PTV97, Pow64, Ram85, SSDS89, She95, Sip88, VH99, ZSP99, Ano80b, BMT95, BKMK90, CP87, Che95d, CLW98, Def77, DF90, Er87a, GP96a, HH94a, HV74, Ken85, KK94, KPT96, Meg90, Par75d, RW61, SS96a, Sal74, Sal87, Sam65, Sib73, Sne80, TU96, Wes80a, Win75, YJ91]. Egdon [Bar66, Sam66, BHJV66]. Eiffel [Coo89b, NO92a]. eigenfunctions [BW71, Bye67]. Eigenproblem [Nas75, Wil60]. eigensolution [CHPS85]. Eigensystems [Wil61b]. Eigenvalue [BES83, Boh68, Erd65, GM65, Osb64c, OM64, Zak77]. Eigenvalues [PW69, SC81, BW71, Bye67, CL83, CJ70, Eva75, FJ66, GRB68, Kai72, Par67, Sea71, SE82, Tay70a, Wil65d, Wri63]. Eigenvectors [Tee63a, Wil58c, Wil58b, CJ70, Kai72, Sea71]. Eight [RHS89]. EL1 [Weg74]. elastic [Sza73]. Eldon [WHM71]. election [MM60]. electoral [Hil75]. Electre [ER86a]. Electric [Dun62a, Dun62b, Abb70, KP60, RW67]. Electrical [Mic97]. electricity [Bur66]. Electroencephalograms [Tow78]. electron [Vin59]. Electronic [BB89, Gat65, GS73, GG59, KV62, Lor76, OD91, Edm60, Fri65, Geo63, GGH^{+}87, Goo61, Mit89b, Sha82, SPMD83, TB63, YR58]. Element [She85, ST81, Boo67f, DHW78, Hil69, Thr69, Woo70g]. Elemental [Abe85]. Elementary [BT63, MST64, Gou70]. Elements [Ili69, Boo67k, Boo67g, Boo67i, Boo67j, Boo67h, Boo67f, CL83, Con66, HW67, Hil69, Ros64a, Woo70g]. Eliminating [HY81, Ram99]. Elimination [Bir79, Cha93b, Dub93, Hon93b, LW93, Wil62b, Can93b, Cuc93, Hon93a, Imb93, LS93a]. eliminations [SF61]. ElipSys [CRS^{+}93]. ELLA [BGMW96]. Elliott [Coo60b, Hoa62b, Hoa63, Liv59, Wil64b]. Ellipses [Smi71, Pit67]. elliptic [Eva63]. Elsevier [GT63]. EMAS [MRW75, Ree75, WW73, Wig75]. Embedded [Bro84b, HKT95]. embedding [Bre93, Bro95]. embrace [Lle73]. EMPA [BBG95]. Empirical [IH86, PW81a, VR92, Wei91, All59, EDS86, Loi68, RT76]. Employing [Tyl88]. encapsulated [Bur79]. Encoded [OW83a]. Encoding [MN90, Sto86, Bae84, CP87, KR89]. Encodings [GW73, Wel72]. encountered [MP65a, MP66, Smi66b]. encryption [YY89]. end [Boo67e, Boo67c, BKM78, WN81]. end-game [BKM78]. ended [Che95d, OOW91]. Ends [van96]. energy [Jon66, Rac65]. enforce [Bak94]. Engineering [BR86, DH97, Fal81, FKA^{+}93, Her61, Sta94, Ver91, Bry59, Hur79, Lar95, Ler69, MR77, MCC^{+}95, PD96b, Rus86, Tyl88, War94]. Engineers [TCB71]. engines [Wil76b]. English [Dun62a, Ker63, Bra63, Dun62b, HSP66, Pai65, VM94, WN81]. Enhance [IB96]. Enhanced [Par88, QW96]. Enhancement [TL85a]. Enhancements [Zmu75a, Zmu75b]. Enhancing [TL85b, ON94]. Enough [Lan85, NN97, AFWH88]. Enquiry [Sta85, Smi68]. Ensembles [GIB90]. Entities [Sha81, Azm90, Com85]. Entity [DP88, FM86b, Par82, Ros82, Rob79, Var79]. Entity-Relationship [DP88, Par82]. entity/event [Rob79, Var79]. entrance [Hon74]. Entrant [BW70a, GP90]. Entropy [Wil78c, Fun84, Fun85, Wil78d]. entry [Car70, Kin69a, Kin71]. Enumerating [Akl87, Er85, Pal86]. Enumeration [SM88, XTU97, BP69]. Environment [BIY88, DS99, FLW99, HJ99, NMM91, YS91, Bar69, Bri80, DB88, DPT94, DO72, Mad97, MMG91, Pon91, SZNM94, UPT89]. Environmental [GL94, GW88]. Environments [BBF^{+}94, CL99a, KS85, MN90, PRS96, RS88, TH98, Ver91, GF87, MCC^{+}95]. EOL [Luk67]. equally [OP71]. Equation [SL78, Wat67, Bir68, Bye65, Day65, Del77, Eva65b, Eva66, FM64, Fet65, Had68, Had71, JS67, Jam87, JM68, Lew71b, Mal61, Man71a, Mil60, MP62, Osb69, Ran68b, Shi89, Tay70b, Tho68b, Zaf71]. equational [FZ94]. Equations [BD81, Bro69d, DAEH81, Fle68, GH73, GH75, MM70, Nor76, Phi70, SB70, SS72b, Abb80, All66b, AW71, AHM74, BFMW64, Bak64, Bar65a, BSD60, BWZ71, BE73, BS71, Boy60, Bro69c, Car66, Cay61, Chr73a, Cla77b, CN63, CG65, DM64, Dav67, Day66, Day67, DS67, DA77, Eg69, Ell63, Eng69, Eva63, EF63b, Eva64, Eva72, EH76, Fie70, For61, Fox62, Fox63, GJM68, Gib60, Hai69, Has61, Hon93a, Hop71, HM77, Hur68, IZP81, JS67, JN65, Jar66a, Jar66b, Jar67b, Jen66, Jen71, JBJ84, Kar66, KS71, Lew72, LC90, LW66b, Lin69, LM68, Lyn64, Mak68b, Mar58, MT61, May64, MM71, MP67, Nor64, O’B59b, O’B68b, Oli69, Osb63, Osb64a, Osb64b, OM64, Osb65, OW68, PC64, PD79, Rei67, RD80, RM69, Ros63a, SWZ93, SS63a, Scr64b, Scr65, SK76]. equations [Sne70, Tay70a, Tew69b, Tew72, Twi75, Wat73, Wat68, Whe59, Wil76a, Wil59, WP68, Wol69, Wol71, Wri64, Wri63, Zam74, Alb71, WY71, Wil62a, Wil63a]. equi [HKA72, Hos74]. equi-distant [HKA72, Hos74]. equiangular [Sib78]. equidistant [For75, Hos70]. Equifrequent [YW82, CCL72]. Equipment [Wil90, GS63, Hop63, Sam63]. Equivalence [RS69, BM90a, Coo66, Lyn64, PW93]. Eratosthenes [Hoa72]. Eric [Ano73b]. Errata [Cam73, Dun62a]. Erratum [Ano79b, Ano80b, Boo74b, MP66, Sav70, Sea71, Wat72a, Zmu75a]. erroneous [Kol83]. Error [BDS82, Bro74c, Bro74e, Bur72, Eng69, Han70, KS85, Kiz66, MS92, Mul81, Mux73, OS68, Ram99, WT84, Wil61b, dW64, Wri66a, ABBS83, And75, BH63, Cha68, CY95, ES75, Gen69, Hai69, How75, Jar67b, Loi68, OSS69, OP75, Oli69, Osb64a, Phi68b, Scr64a, Smi66a]. error-correction [And75]. Errors [Bel76, Bro74e, Cha70, Lei86, YFM71, BD80, Jac73, PC64, Pow67, Ull77, Wal64]. ES-IS [MZ96]. essay [Las68]. essential [JC86]. established [Wil67c]. Establishing [TH98, Ada76]. establishment [Bry59, Tay61]. Estimate [LV79, Phi68b, Hai69, Kol78, Loi68, Smi66a]. estimates [Cha68, Eng69, dW64]. Estimating [BK79a, CN82, Gos63, LW85b, Mit76, Wat75b, Pez76, Wat73]. Estimation [Cha70, Eva65a, Scr64a, TDW^{+}99, Cha67, CW71, HD71, Jon70, OS68, Oli69, Ran68b, Tza73]. ETAOIN [Kli75]. Ethernet [Fen90, Par88, TH98]. Euclidean [Wri74]. Euler [Scr72, Wyn71]. European [Coo89a, Fin93]. evaluate [MB90]. Evaluating [CTW94, Fle65, HG67, Hun70, RS87, GE69, Gro68, Tew68]. Evaluation [AHL92a, BPR74, Cav83, CCL98, Coo75, CO72, IH86, KU99, Lan76, Lan64, Mah84, Mak68a, MP65a, MP66, MLR90, MZ96, MB96, Pap82, PS92, PB74, PRS96, Roy78, TK99, YW82, AEH76, Ber87b, CPV93, Cha92, CJ70, Coo67, CR63, CP68, CP71, Del68, GR75, GE71, HPLG79, Hav95, Lit70, LP73a, Mil60, MB66, Moj77, OS69, Oli71, OKMS96, Rei81, Sea71, Val71, Wat59, Wes80b, Win69, YG93]. Evaluative [Mic81]. even [AB84, Com85]. Event [Ree84a, Nud69, O’K85, Rob79, Var79, PC78]. Events [PA90, Cla94]. EVERYBODY [Hig63]. Evidential [Lal99]. Evolution [AM99, CK95, Pin61, ST89, FE80]. Evolution-Based [AM99]. evolutionary [SL73]. exact [Bra94]. examination [AW85, Col64a, DCHR76, FL68, OF93, Stu84, TP91, Woo68]. Examinations [Ong69, Gro68]. Example [Muc82, Mum85, Par82]. Examples [Age95, Wil77a, GF87]. exception [PA90]. Exceptions [LO99, Ram99]. Exchange [ZSM99, CP66b]. exchanges [DD91]. Excluding [Kam77]. Exclusion [HPR88, TK99, PJ89, YJ91, YJ94]. Executable [Bro79b, Fro92, LP92, GCS89]. Executing [LPSC99, PG90]. Executing-Committing [LPSC99]. Execution [GP90, NL90, Tre92, WD74, WC86, Gal87, PS92, Teu87, Wal77]. Executions [KBTB97]. Executive [BC73, Ros66]. Exegesis [Sma88]. Exercise [HS83, Akm91, Oxl89]. Exercises [Wre71, TS68]. exhibit [Par75d]. existential [Can93b, HRS93]. Expanding [Nor76]. Expansion [DdS98, Cha67, Coo67, Sv83, You63]. expansions [Sto68]. Expected [And90, Bri85, Lar82, Mah86, Wei91]. expenditure [Red60]. Experience [BM88, BMR67b, Dac63, FR87, Lum64, ON79, Red60, SA81, Tay61, Wil63b, Ayr64, Cor62, Dij62, Fis76, FGR67, Gle62, Gol64, Kin62, Mil63, MS70, Nef65, Pla61, Ric64, Sut60, YR58]. Experiences [BHS65, KMK95, Thi86, VR81b, Gri59, Hic60, Liv59, Sla64b]. experiment [CF88, Lew71b, LK76, PRO80a]. Experimental [BC73, Mil70, YP86, Aki88, Gea61, Hen78, HH79a, LKK71, New66, SS72c, SS72d]. Experiments [GP81, HM83, Mic63, Mic82, Abb64, KRK73, YGS63]. Expert [BM81, Mic80, Rob90, BWM88].

Explaining [CM89]. Explanation

[NMW97]. Explicit

[Woo82, Bir68, BK83, Eva65b, Eva66]. explicitly [Scr64b]. exploit [Ros70]. Exploiting [Ben85, HKT95, HH95, ME98, MCC^{+}95, Ire93]. Exploration [Gor78, AWH91]. Exploratory [Tre91]. explosive [Pal63]. exponential [Lip80, Spä68]. exponentially [Til69]. Exposing [NN97]. expressing [Tho62]. Expression [DE86, Gol93]. Expressions [Fok96, HD82, Lan64, Vel82, Col63a, Hou83, Hus61, Rei81, RL72, Sip69, SY72, WG83, Wat63]. Expressive [Bol97]. Expressiveness [MH98]. Extended [Cle66, Fen90, KA86, KS97, MVL92, Ros93, WM83, Wil79, Azm90, BM63, BML94, HL76a, Hou83, WWK^{+}90, And66]. Extending [CZW94, Hie96, Hie97, Tes77]. extensibility [Lay85]. Extensible [HW92, NF76, NF80, FKLM76, Mol71]. Extension [Pra88, Gal87, Hoc62, Hur68, OS65]. Extensions [Ehr72, Fin72, For61, Mee72, Yul68]. external [dP60, TH82, Tho60]. extra [HL92]. extract [Cra66b]. Extraction [Gow58, VV94]. extrapolated [Eva63, Eva64, Had71]. Extrapolation [Vit75, dW64, Wri63]. extremal [Cha93a]. Extreme [Par96a]. Extreme-value [Par96a]. extremum [Bae62].

F [BL81a, FE80, Web70]. F.R.S. [Fox87]. faces [Jar75]. facilitate [HKT95]. Facilitating [Fro85]. Facilities [PPB83, RC70, Bro58b, BMR67a, LR67, Nap67, PL68, Pri64, Wil61c, You63, For66]. Facility [TL85a, Azm90, BD72, Cam78, PPRB89, Uck85]. FACT [Cli62b, Lum64]. Factor [TI75, BA88, Car61, Eva62, Eva65a, LM68]. factored [Kom74, Woo69b, Woo70b]. factorial [RS69]. Factoring [MD84]. factorisation [HSW92]. Factorising [MJ75, Alw71]. Factors [MS74, NV77, Le 88]. factory [Bry59]. Faculty [Sel88, Alm69, DD70]. Failure [The91a, TL97, HGL93, MR77]. failures [LC95]. Fair [ZSM99]. fallacies [Wat72b, Wat72a]. Families [FS87]. Family [HO93, MWHC96, MP79, Mei66, Ram70, Red60]. Fast [AM98, Boo67k, Dav92, Er83a, GC87, Ham91a, HL92, MH77, Mad80, OF93, Tay76, Ull88, UPT89, WZ99, Coo89c, DA77, DMS^{+}89, Er88a, Eva77, Fen87, Wis64, vR70a]. fast-carry [Fen87]. faster [GV91]. Fat [NW93]. Fault [AE90b, BENM98, BL71, CH99, Com75b, GIB90, JES98, MS91, MS92, WFD98, DJ95, GP96b, OS65]. Fault-based [JES98]. Fault-Secure [WFD98]. Fault-Tolerant [BENM98, MS91, MS92, AE90b, GIB90]. Faults [BKL74, Hil71, LC98, SS72a, YT79]. faulty [GP96a]. Favour [DM83]. FC [DB90]. FDDI [CDCL97, RLD94]. FDDI-M [CDCL97]. FDL [Pou92]. Feature [Aga96, FSW85, SCC^{+}96, Tay96, VVW96]. Features [MRW90, Wil78a, BBH^{+}63, BRC66, FH77, LT71, Par75d]. February [Gal71]. Federating [Col97]. FEDOS [She85]. Feedback [GK73, RV93, GE69, Sv83]. fer [Fos64]. Ferranti [FdVR58, Gri59, Hun59]. Fewer [Pri81, Lin72a]. Fibonacci [Boo79c, Er83a, Ove67c, PHJ67, PP79, QW96]. field [Aha65, Gea64, Wil61a]. Fields [Dub93, WN86, Ran68a]. FIFO [BBP98]. figures [BJ80]. Fike [Roh76]. File [FB99, GW73, HK80, Hel93, Lar85, LW85c, MHS85, Par75b, SR76, SW84, SH72, Sur86, Wat72b, Wel72, WZ99, Wri83, All82, Aus71, BBG^{+}71, Bon71, DC84, Dav84, KCP99, KH81, MKP87, Rai87, TH82, Wil78d, Wis68, Wat72a]. Files [Bra86a, BHK85, CC87, Edw77, Huz79, ID73, KW84, Lar82, NV77, NW93, Ran74, Wat75a, BMT95, Bra85, BKMK90, CC88, CLC92, Cro80, Kol78, Wat71]. Filing [Fra69, Wil64a, Col70, SS72d]. filled [Dav82, Par89]. Filling [Col87]. films [Liv59]. filter [BMT95, She89]. Filtering [BGS99, Hun72]. Filters [Han78, Lyn72, Bea68, Mul87]. Find [Kro70a]. Finding [Bre71, CM91, Cha90, CM93, Gri92, Jen99, JP86, Nic66, Pow64, SNH76, Wei93, CL83, Che65, FJ66, Jam89, JJ67, Kai72, KL71, Lar64, Lil70, Lil71, Par67, Pow62, Rei66, Ros60]. Findler [Eme80]. Finite [Gol93, Ham77, Hie97, Hun70, Lle83, V’y99, Woo77a, YFM71, BW71, DL76, Eva65b, Eva66, Hie98, Lle73, Osb63, Osb64a, Osb64b, Osb65, Ran68a, Tee63b, TR93, Thr69, She85]. finite-difference [Eva65b, Eva66, Osb63, Osb64a, Osb64b, Osb65, Ran68a, Tee63b]. Finite-State [Hie97, TR93]. firm [BHRS69, BHRS70, HRS68]. firms [Lic89]. First [Ree87, Rob77, SA81, YCC92, Yel89, All66b, AW71, BFMW64, Bar58, Cha90, Che62a, Chr73a, Cox71, Er89a, Lew72, Lin69, Mil60, O’B59b, OP73b, RM69, Rya63, Scr64b, Sut60, War60, Wol71]. First-Fit [Ree87]. First-Order [YCC92, RM69]. Fisher [Eme80]. Fit [Pag82b, Ree87, Rob77, TI75, Cam71, CP87, Leu82a, Ree79c, Ree80a, Ree83]. Fitting [Oxl85, Sou76, Spä68, Til69, Cad61, CW61, Cle60, Gro71, Hay69, McC71b, OSS69, OP71, SM71]. Five [Wic72, Wil85a, Chu68]. Fixed [Bla82, Nie65, PW81a, Smi71, TB94, BP79, She89, Sto72]. fixed-size [She89]. Flat [DSB97]. Flexibility [BP92, Bol97, ID73, Rey71a]. Flexible [Cha91, Cha92, IHH95, SMR98, SZ74, ZSM99, Zmu75a, Zmu75b, ESK95, GF87]. Floating [RS88, Wic92, SS63b, Wic89]. Floating-Point [RS88, SS63b]. Flocoder [MKL71]. Flow [AE90a, Deo83, GP90, PB86, PG81, THR82, Wil77b, WO78, Gre69, How67a, Mid77, Min88, NI70, OB68a, Osb71, TCL89, TP89]. Flowchart [Wil83a, Wil82a]. Flowcharting [Mil81a]. Flowers [Aga96]. Fluid [Pau80, Sta85, Abb64]. flying [GB61]. flying-spot [GB61]. FOCUS [BBG^{+}71]. folding [EH76, Lun71]. following [Dak73]. Force [Dav63]. Forecasting [MM60, Hio67, Mui58]. forecasts [Gea58]. Foresight [Tay96]. forests [Win60]. Foreword [Mur98]. Form [DP88, Geh83, JW74, Wil62b, WO78, CCH73, Gil58a, Gla69, Hou83, Jac73, LL70, Mid77, Oul87, Ser95, Sim69, SF61, Woo69a, Woo70a, Dix85, Roh68]. Formal [BM92, BPTV92, BS99, Coo92a, Coo92b, Coo89c, DF87, EMCO92, Gal92b, Gha72, HRW95, Kuh92, Ler69, MM84, MMR^{+}98, MVL92, PD96a, Rob88, Sta94, TK95, Tse87a, VM94, War96, WB99, vFV92, ATL93, Bar72a, Cur95, FM85, GK96, Lar95, Lee72, Lei87, Nap67, RDPS95, SFD92, TP89, TL85c, War94, Wic89]. Formalisms [Fro85, Sch77b]. Format [BP81, Car70, HKPW89, Kin69a, Kin71]. Formats [Zve75, Zar69]. Formatting [Cho82, Sip69]. Formed [CCD86, Er83b]. forming [Col63b]. Forms [FR80]. Formula [Bay76, Ada76, CC88, Cha68, Cle59, JS67, KM66, OS68, Ran68a, Sym71]. Formulae [AEH75, Bar75, BJ78, Bar61, FR63, Jam89, LM63a, Mei66, MP62, Oli71, Pie71, Ste67, dW64]. formulas [Hon93a, Sac62, Tha66]. Formulation [BGR86, TPvdW96]. Formulations [BT81]. Forsythe [Hit78, LS78, SG78]. Forsythe-Moler [Hit78, LS78, SG78]. Fort [Tay70b]. Fort-Frankel [Tay70b]. FORTRAN [BP81, Cha72, Cha73, Dew72b, Ehr72, Fla72, Mee72, SS72a, FJA80, VP80, Art75, Bar71, BL90, BLY70, Bur79, Cha71, CL83, CPR75, CHPS85, Coo72a, DCHR76, Fin72, Gle62, Hal72, HV74, Hea68, HS69, HPB84, Hil71, Hou71, IZP81, Mac70b, Mac71, MP65b, PC78, Pul64, RT76, Sal71b, SA73, SW91a, SHCP91, TH64, Tay68, Tay76, WH87, CC74, HL70, NC76, Rey77, Sch82b, SW83, Sta74, Wel70]. forum [Ano63]. Forward [MS92, Tay96, Rya66]. Forwarding [CCL98]. Foundation [RBCM97, BGMW96, TP89, VG73]. Foundations [BL81a, EMO92, Gea58]. Four [BT81, Tho58a, AW85, Bro95, Wie95, YR58]. Fourier [Bel63, Boo68a, Boo68b, Dol66, Gen69, Ham91a, MH77, Mat63, Sto68]. Fourth [DP88, O’R70, Ede63, Eva65b, Eva66, Fyf70, Kar66, Whi72]. fourth-generation [Ede63]. fourth-order [Eva65b, Eva66]. FP [TAP89]. FP-based [TAP89]. Fractal [HV92, HM98, JM96]. fractions [How59]. Fragmentation [Leu83a, Rob77, Aus71, Leu82b]. fragments [All82, Lev65]. Frame [YP86]. Framework [BL91, CTH99, Coh01, Mäk01, TAC98, TAC01, CS95, CCB95, EWH86, GK96, KH87, LQ96, Mil81b, Noo84, Ros82]. Francis [HR73, Wat11]. Frankel [Tay70b]. fraud [Cha78]. Fredholm [BFMW64, DA77, Ell63, RD80]. Free [BP81, Lle83, MH77, Ree79c, Ree80a, Ree83, Gob75, Goo75, KPT96, LL70, Lle85, Pac88, PS96, SEP89, SP91, Woo70a, YJ91]. Free-Format [BP81]. free-text [Gob75]. French [McC71a, McC72]. French-curve [McC71a, McC72]. frequency [Mea73, Par75d]. frequently [MP65a, MP66]. friend [Mat67]. Friendly [SLH83]. front [WN81]. Full [Bur75a, Mar91b, Cuf72, You60, Dix85]. full-scale [You60]. Full-text [Mar91b]. fully [Par96a, Yel89]. Function [AM92, Aki77, Bre71, FR64, Fle65, GP88, NM65, Pow64, dRC74, Boy63, Bra94, CP79, Cha67, Cox70, Dun73, FB75, HB70, Hor60, Jar66b, KL71, Lil70, Lil71, Mil60, MR65, OP73b, Phi74, Pow62, RS69, Ros60, SC71, Til69, Tom67, Zan67, Zar80]. Functional [Axf90, Bur87b, Bur88, HM93, Har82, HK88, Hug89, HKPT99, JS89, KA86, Orm88, Pou92, Pou93, Rin75, SS92, SK95, Top82, BHY89, CAC82, Cla94, EB89, FCHK96, FL89, Gol94, Mar91a, Pey89, Stu71, Wat68]. Functionality [Sha86]. Functions [Cra72, CM93, Has75, Jar68, Kno75, Miy90, Pow65, Ram99, Sha81, Sou76, Suf90, dHM93, All59, BY66, BRH70, Bea68, Che90, Che65, Cox71, CO66, Er82a, EM71, Gar78, Goo61, GL64, Hil71, Hur73, Jac73, JN65, Jor73, Lar64, McC78, MS97, Mul92, Ord64, SS72a, WV70, Wat73, TCB71]. Fundamental [Lei86, Tho62]. Fundamentals [FSWL82, Wil90, WD80]. FunZ [SC95]. Further [Bro58b, Fra70, Goo85b, Woo71a]. Fusion [SM97, AW85]. Future [Ham66b, Bro73a, Cut59, Fis74a, War60, Wil61c].

G [Den80, FJA80, HR73, MM69, NLW74, NQ80, RP63, Sla71, Wil67b]. Galerkin [DA77, Del77]. game [Bra80, BKM78, EDS86, Fin60, Fos64, Mic63, Mic82]. game-learning [Mic63, Mic82]. game-theoretic [Fos64]. GAMM [WD79]. Gamma [Fal81, Pal62]. Garbage [FN78, MS91, NSW87, Weg72]. gas [Lan65]. Gate [Edw78, KK94, Sal87]. Gates [EH78, Fen87, RDT80, Win75]. Gauss [Cha68, Cha70]. Gaussian [Bar61, FR63, HCH96a, KP98, Pie71, Tom67]. gear [Lei62]. General [Can93a, Has74, LC93, MK94, TK99, Alw71, And66, Bra74, Bri79, BKK87, Dak75, Dol66, EA70, Eva71, GE69, LW67a, LW67b, Mar74a, MP67, Par96a, PC78, Ran68a, Ros63a, Rus63, Sne70, SS72c, Str65a, Und63, YS60]. general-purpose [Und63]. Generalisation [Gou70, Par75d]. Generalised [BHW78, CC87, DAODT85, OP73a, Zak77, BJS74, Bra75, Dee80, Had68, Pin77, PM72, RS69, Scr72, Weg72, Woo70a, Wyn71]. generalization [Er84a]. Generalized [Er84c, FD88, Fle68, FM64, Hat84, Imb93, Tew68, WWN91]. Generalizing [HTW89]. generatable [Er89a]. Generate [PP98, CDHL93, VR81a]. generated [Ful76]. Generating [BM59, CW70, CF89, Day77, Dix85, Er88b, McL85, ON90, Roh76, Roh78, SM88, Top82, Wil77b, Bra94, DMS^{+}89, Er83b, Er88a, Er89b, FL80, Iga80, Pan70, Tho58b]. Generation [AEH75, AD76, BHK85, Far78, GB83, HGMM75, LB86, Mak91a, Mid87, MF82, OS70, OS71, RW68, SY72, Spo85, Tak99, Tho68a, VR81b, Vel82, Woo77b, YW82, YT79, Atk93, Ben72, Ede63, Edm60, Fik75, FR63, IW86, Inc87, OB68a, Pag71, Roh90, SD86, Tow77, Var77, WF78]. Generative [RW74, WBM^{+}94]. Generator [SS72b, BW69a, Gra87, HM95b, Hun60, MWC70]. Generators [All90, Gri88, LB86, NP90, Woo77a, DR67, MP68]. Generic [Sch77b]. Genetic [JES98]. Gentle [CF81]. Geoffrion [BP69]. Geographic [Abe85, GL94, LG86]. Geometric [FR80, SCC^{+}96, WT84, FB75, Oxl89, PL94a, Sch94]. Geometry [Bra86b, DF90, MP65a, MP66, Ric73]. German [Muc82]. getting [Atk84]. Gill [WD80]. Gilmour [FE80]. Girling [HR73]. Giro [Gra67]. given [Tur79]. Givens [RW62, RW61]. giving [Woo70d]. Gladwell [Wil67b]. Glass [BL81a, Coc97]. Global [WC98, Pri77]. globe [vSHBT98]. go [LR72, TW64]. goal [UPT89]. goal-oriented [UPT89]. goals [Lan76]. Goblin [Wil88]. Gödel [Goo68]. Goertzel [Gen69]. GOLD [Ove67a]. Golden [Hil79b]. Goldfarb [HV92]. GOLDSEC [Ove67b]. good [Mus67]. Goodman [RP63]. GOTO [Gal73, Pag82a, WC85, PD96a]. GOTOs [Tay81]. Goursat [Day66]. government [BS61]. Gozinto [DK74]. GPSS [Sch82b]. GPSS-Fortran [Sch82b]. gracefully [CPV93]. Grader [Ber66]. gradient [Cla77b, Ram70]. Gradients [FR64, Pec70]. grading [TS68]. gradual [Pai65]. graduated [Dru63, Pol60]. gram [Ull77].

Grammar [Sto86, Wil85b, XTU97, CND66, EB89, Mar91a]. Grammar-Oriented [XTU97]. Grammars [BDS82, BW78, DC90b, Fro92, Has74, HS83, NMW97, PK86, WM83, KC88, LL70, Pag82a, Ter73, Woo70a]. Graph [AD95, BENM98, Chr71a, DB81, DK74, Eke95, Joh74, KLB91, Mau73, Mit76, SNH76, SCC^{+}96, Azm90, BBKR89, EL89, GP80, WP67, Wil76a, Woo69c]. Graph-Based [SCC^{+}96]. Graph-Reducer [KLB91]. Graphic [GM83, WCH78, CS75, Gla69]. Graphical [BJ78, EM71, Gor78, Mac70a, CCH73, HKT95, Lar64, New68, Pon91, YJ91, ZM95]. Graphically [CR83, PJ89]. Graphics [CDZ74, Col87, CW82, JR81, MN90, NM90, Woo82, FS94, Gro73, Mar73, Mar74a, Rey83, Ric73, BK83]. Graphs [ADD86, Axf90, FvLMS98, GH94, Wil85a, NE67, WC94, WM84, YJ94]. Gray [O’B59a, CCC92]. Great [TB63]. greatest [Ros60]. Greedy [BGO94]. Greeting [MZ96]. Gregory [Wil70]. Grid [Du96, FA88, Bre93, GHR^{+}93]. Grid-based [Du96, GHR^{+}93]. Grids [Bra86b, Boy63, GHR^{+}93]. grist [Mus67]. Gros [O’B59a]. ground [Tur75]. Group [WCL71, YW82, Sta74]. Group-size [WCL71]. grouping [CO72, Moj77]. groups [Atk80, Azm90]. Growth [LW85c, Bra94, Und63]. growth-decay [Bra94]. Guaranteed [Bre71]. Guaranteeing [BRP93]. Guard [NW93, BS84]. guide [Fin93, BLR94]. Guidebook [BL81a]. guided [YG93]. Guidelines [Hil76, Wal82].

H [Alb71, BL81a, Bro74b, BS73b, Bur80, Cor79, Fox87, HR73, Kil81, NLW74, YFM71, Cor62]. Hagen [Kil81]. Hairy [KV95]. Halstead [Pra88]. halting [Hut84]. Handbook [FE80, HV92]. Handling [Cro74, Hil77, LO99, Mad98, Rey77, Sam81, Tri74, BBG^{+}71, Gow62, Hin61, Mil68, PA90, Str61, WTW91, WN88]. Hanoi [Wod78, Er82b, Er84a, Er84b, Er84c, GC87, Ged96, Hay77, Lu89, Roh87]. Hao [PCSG63]. hard [TB94]. Hardware [Laz81, MS92, MG95, GS63, HP77, Mar76, MES97, OL68, PK85, Sam63]. hardware/software [HP77, MES97]. Hardwired [BK75]. Harmonic [Bec58]. Harry [Ull73]. Hartley [NQ80]. Harvard [RP63]. HASE [IHH95]. Hash [AK74, Bay73, BD84, Bra85, Bra86a, CM93, HD72, Lar82, Lyo85, Mul92, SD78, Sam81, CLW98, Eck74, Fro81, Gob75, Mad80]. hash-based [CLW98, Mul92]. Hashing [BD82, CL86, CC87, CV86, CK94, CHK85, Kno75, Lar85, MWHC96, Mul85, OT91, PH73, RL82, CCJ91, CLC92, Cze98]. Having [Han70, Bea68, Cha67, SH72]. HDG [KLB91]. HDG-Machine [KLB91]. Head [BK79b, PW81a]. headed [Dei95]. Heads [PW81a]. Health [Ker76]. Heap [Luk99]. Heaps [vW93]. HEAPSORT [Weg93, XY90]. heat [Bir68, MP62, Osb69, Zaf71]. HEC [Woo59]. Height [GOW83, Wei93]. Height-Ratio-Balanced [GOW83]. heights [MW68]. held [Var67]. helicopter [Har59]. help [GF87]. Herbert [BS73b, VP80]. Hermitian [Gou70, Ham91a]. Hessenberg [Han62]. Heterogeneity [Col97, Ver91]. Heterogeneous [CZL97, Lal99, Li98, BM90a, BS84, CS95, SZNM94]. Heuristic [Ben82, DT96, Mic71, NV77, She92, AO80a, Bur66, Hut86a, Wym73, Yul68, dS77]. heuristics [NO92b]. hi [Lic89]. hi-tech [Lic89]. Hidden [Jon71b, Wil77c, Ric71, Tom82]. HIDEL [ACDP91]. Hierarchic [BW73a, Mac91, EHW89, JS68, SR74, vR70a]. Hierarchical [ACDP91, DS99, Gru96, Kau92, KVKP95, Moj77, NMW97, Roh73, Vau95, Coo89c, CO72, LW66a, Mur83, Wil78b, LW67a]. Hierarchically [Lei83]. Hierarchy [MM84, FK86, KVW89]. High [Bac71, CMW98, ER71, Geh83, Ger70, Hil72, IEH^{+}85, MP62, OT91, Tre82, TL85b, Wal71, Ahm72, BPG87, BLY70, BP79, Bra72, Dav67, Hav95, Hop71, Ken80, Mak68b, Man71a, Mar58, Oli71, Pal72, Pal63, Wal82]. high-explosive [Pal63]. High-Level [Ger70, TL85b, BPG87, BLY70, Hav95, Wal82]. high-order [Dav67, Hop71, Man71a, Oli71]. Higher [Har92, Hen88, Mel95, Ram99, von95, CWH93, Cho95, LM63a, Osb93]. Higher-Order [Har92, Ram99, von95, Hen88, CWH93, Cho95]. Highly [KLB91, TF96]. Hill [NPCA81]. Himmelbett [Bel73, Fre74, Str74]. Hints [Bra77]. Historical [Sar90, Mor89]. History [Aga96, The91a, Les83b, Lay85]. Hit [Wat76b, Pez76]. Hoare [Llo74]. HOL [Age95, GM95, Har95, Tre92, von95]. Holder [Zah81]. holds [PR98]. holes [KVW89]. HOPE [WS86]. Hopfield [CWH93]. horizons [Ben70]. horizontally [Seg87]. Horn [Ber87a, Pad91]. Horner [Gil58a]. Horsley [BB80a]. house [AF76]. Householder [RW62, Rei67, Wil60]. HPP [KA98]. Hubs [TH98]. Huffman [DdS98, FK90]. Hull [GS79]. Human [FE80, Lic89, MTPL88, Rey83, Tul85, vO78, BA88, dF66, JC88, LBC^{+}94, Mic82, SW91b]. Human-computer [MTPL88, JC88]. Hutchison [FSWL82]. Hybrid [AE90a, AM98, BP92, Coo75, Gol85, Gom81, KP79, NR92, NPW68, SB70, Sto73, Wu97, ZSP99, Car64, Cro75, Fox78, Jac71, NSW87]. hydro [KP60]. hydro-electric [KP60]. hyperbolae [Pit67]. Hyperbolas [Smi71]. Hyperbolic [SB70, Day66, GlM68, Hop71, Twi75]. Hypercube [BGO94, GIB90]. Hypercube-Labeling [BGO94]. Hypercubes [She95, Bre93, GP96a, GP96b, LM63b]. Hypergraph [OM88]. Hypermedia [BV92, Ski96]. hypermesh [OKMS96]. hypermeshes [LOKM99]. Hypernet [Kau92, Mar94]. Hypertext [MH98, PSK96, Rit89]. Hypertext-moving [Rit89]. Hyphens [Pri81]. Hypothesis [Ris99a]. Hyvärinen [TCB71].

I-Pigs [Pon91]. I.C.T. [Ham66a]. I.S.O. [Ros64b]. I/O [LV79, MK82, Per80, Sil79]. Ibbett [Eme80]. IBM [Pal73, Sha86]. IC [HGL93]. ICC [GT65]. ICL [CK95, HP77, IS81]. ICN [AHL92a]. Icon [Gri80, Gri90, WG83]. Idea [Coc97]. Ideas [Rob75, AWH91, Wol90]. Identification [BDB92, Hol75, Tse87a, Wha90, CCL72, O’S88, Tho64]. Identifiers [dHZ97]. Identifying [Tse87b]. identities [Bir89]. ideographic [WN81]. IFIP [GT65]. II [Wil63a, All85, HH79a, HPS61, LW67b, Mic82, Pen77, Tra84]. III [Lew63, Pin61, Ree83]. ill [Oom93]. ill-conditioned [Oom93]. ILLIAC [Gea64]. illustrative [Kro70b, Kro70c, Kro70d, Kro70e]. ILP [Kil81, ME98]. Image [AF88, How97, Meg96, MW97, OW83b, SH97, TWMG93, DW77, JM96, OF93]. Images [OW83a, AO87, DF90, FN93, FNK94, GV91, Lut80]. Imbedding [Sco70]. immediate [HRW95]. IMP [Ste74]. Impact [Col97, LE86, ME98, Dou61, Ham80b, Ham81b, Hea81, Jam81, Wat82, You63]. imperative [Bai88]. Imperfect [LCH99]. Implement [All90]. Implementation [Akm91, BSH90, BKMK90, CS85, Cha91, Dan92, DAT87, DS90, Dun62a, Dun62b, ER80, FE69, FA88, Gil73, GP90, Has75, HAAZ99, Hou75, Kac91, LPS94, MA89, MN90, MSS93, NF80, Pou92, Sip88, SB81, TL85a, Ull88, WF89, dRC74, dHZ97, vFV92, Bar75, BML94, Bra74, Bro69b, BMRS90, CCB95, CFL84, Cze88, Er87a, FK68, Ham66a, Hie98, Hou69a, LMP79, McD77, OL78, PRLW93, RB85, Ros82, SSDS89, Sma88, TB86, ZR88, FGR67, NPCA81]. Implementations [HKK91, Hul86, Cho99, Pey89]. Implementing [FRS^{+}87, KRR79, MRW90, GP80, Mak89]. implementors [ELR78]. implicant [Iga80]. implications [PRLW93, Smi86]. Implicit [Hai69, Nor76, SK85, Shi89, BP69, CL78, Che95d, Had68, JKW94, Mak68b, Ros63a, Wol71, Zaf71]. implied [EM71]. Imply [Ver91]. importance [Rey71a]. Important [Sha86]. Impossible [SHH81, AMKM66, BSPI65, HLOS65, Str65b]. Imprecise [Lev96, Che95b]. Impressions [Dak73]. Improve [HR97]. Improved [AW71, Can93b, CW70, DS91, Edw78, Lew72, MM84, ME88a, WTW91, ZOG85, FM66, GN88, Leu82a, Pal69, Sal71a]. Improvement [DL76, Dur74, TM94, Aus76, PL94a]. Improvements [Bun97, EH78, Roh76, Fen96]. Improving [Bag61a, KSD77, LY79, SM97, TL97, Wes80b, Fos68b, Scr70, War74, Weg93]. in-house [AF76]. In/One [CCD86]. Inaccuracy [JP76]. inadequate [GL70, Lyn69]. Inclusion [CW82]. income [God75]. incompatible [WR86]. Incomplete [Bow84, SK95, KS91, WN88]. incompleteness [MV88]. Incompressibility [JLV99]. incompressible [Abb64]. Increasing [BR86, Laz81]. Incremental [AMN81, Bre82, Bro84a, Ear80, HH94b, PW81b, Gri73, McC71a, McC72, vdWHvB98]. Independence [BL81b]. Independent [KBTB97, CCJ91, Rya66]. Index [Col82, HK85b, JW82, NT79, NS77, PF86, WK77, Zve80, All82, Ano97a, Ano97b, BMT85, Bon71, BC67, Fro81, Kol78, McD77, Wri60, Wri61b]. Indexed [Huz79]. Indexes [HK85a, Ing74, LP73b, Mul87]. Indexing [Foo86, HK80, JO92, MM84, SB85, Smi70a, TUW98, Wat75a, WB98]. Indicators [FR87, VVW96, Er86b]. Indirection [Gol85]. Individual [YT79, Le 88, Ada76]. Inducing [CD82]. Induction [Bur69b, Sol99, Wu97, Bra77]. industrial [Hoc62, Hol95]. Industry [The91a, Wil91a, Bag61b, Gal59, Sla71]. inequalities [McC93, Wil77d]. inequality [AJ70]. Infection [Coh01, TAC98, TAC01, Mäk01]. Inference [Clo87, HW88, UC90, Ull67]. INFERNO [Qui83]. influence [Win59a]. Informal [Atk76, FM85]. Information [Abe85, AM91, Ano73a, Ano81a, Ano82b, AFWH88, AWH91, Bla92, BHW78, BW73a, BW75a, BV92, Cha87, Cli62c, CCS87, CT92, F^{+}73, FSW85, FFP99, Fuh92, Geh83, Ham91b, HW92, Har80, HN88, HM70, IL74, IB96, Kni93, Lan85, Lev99, Mac91, MJ75, MHLS93, Nee63, NLW74, NH89, NBL95, OD91, Ole91, Rey90, Rze85, Sca85, Sch77a, Sch83, Sme92, Sta85, SK95, TCB71, Tul85, Vau95, VS87, WB68, WB98, WHF83, Wor94, van86, AB84, Ben70, Bon69, BG71, Cap86, Cla88, CH60, Col66a, Fai58, Gat65, H. 68, HT78, HGL93, KS91, LW68, Lan76, MR96, Mil95, PRLW93, RSS87, Ros82, SI80, Sca89, Spu88, TPvdW96, WN88, van71, vdWHvB98]. information-statistic [LW68]. Information-Theoretic [CCS87]. inherent [Roh90]. Inheritance [Hut89, KL89]. inhibition [RDT80]. Initial [Bae64, Bar75, Sco70, Car64, Hop71, Twi75, Wol71]. initial-value [Car64]. Initialising [BP69]. initiation [Pal63]. injection [DJ95]. inner [Bro73b, Wil68a]. Innovation [The91a]. Input [BP81, Dun63, Pad73, BS75, Coo68, DI78, FdVR58, Gla69, Hoa63, Hop78, Ing76, New68, Sol68, SS72d, TS64, Tay76, Wil76a, Zar69]. input-output [BS75]. input/output [Hoa63, SS72d, Tay76, Zar69]. inputs [Sal74]. inquiries [Wri61b]. inquiry [Sto72]. Inscribed [Smi71]. insertion [Ull77, Zak84]. Insight [Sam72b]. Instability [Wil62b, KM66]. installation [Rus63]. Instances [Mai85, CK93]. Institute [Ano74]. institution [Osb93]. instruction [Bae84, BS89, Geo66, Ibb72, OW66]. instructional [Rus86, WO75]. insurance [Dru63, MFM64, Ano60]. Insuring [All75]. Integer [Dak65, TCB71, FL80, Jon72, Law69, Pro83, Wil77d, Wil89]. Integers [Bro74e, WZ99, Bel70b, MD84]. Integral [DAEH81, Phi70, AW71, BFMW64, Eg69, Ell63, FR63, GE71, JS67, Lew72, Lin69, Lit70, RD80, SY72, Wat73, Wol69]. Integrals [AEH75, AEH76, CR63, CP68, CP71, Del68, GL64, MP65a, MP66, McC78, OS69, Oli71, Val71]. integrands [Fox67, Sha72b]. Integrated [AGNS91, BSH90, FKA^{+}93, Hop63, SFD92, WM95, Cut59, Dav63, LC90, Mac70b]. Integrating [AE90b, Pai99, SW91b, Mar58, Wal64, Whe59]. Integration [CG65, CT92, DM83, Har79, HCC74, HG71, LM63b, RV78, BS71, Day65, Don78, Fox67, Gib60, GM72, Hai69, HKPW89, Phi67c, Phi67b, Ral61, RM69, Tha66, Tom67, Wri66b]. integrations [Ede65]. Integrity [Dav76, Fra69, Wil72, FW83, PPB84c, WTW91]. integro [Day67, Eg69, WP68]. integro-differential [Day67, WP68]. Integrodifferential [Phi70, Wol69]. intelligence [BMPS71, Bro73c, Bur80]. Intelligent [Kie88, KSD77, TH98, RSS87, TCL89]. intentions [Lo84]. inter [RV93]. inter-media [RV93]. Interacting [CR83]. Interaction [Cor65, NC76, PR98, SW91b, Wak93]. Interactive [BY79, BTD99, CDZ74, ER80, For72, GK84, GP73, Gor78, Hil76, HJ99, Mac70a, SLH83, BWM88, CS75, God75, Gro73, Mar73, Mar74a, MR76b, PW74a, Pin77, Pon91, Rey83, SB81, Stu71, VM93]. interchanges [Hea63b]. Intercode [Ber59]. Interconnect [Tho93]. interconnected [MCL89]. Interconnecting [ZE88]. Interconnection [LC98, Wil91b, HH94a, OKMS96, PM90, SP91]. interconnections [Wis64]. Interface [AK81, Dix85, Fro85, HW77, PJ80, Sta74, Ste77a, Azm90, GE85, GF87, HP77, Jam80, MTPL88, Shi89]. Interfaces [Hop81, NM90, SLH83, BM88, Dee80, RB85]. Interfacing [NO92a]. Interference [Han94, Qui83]. interior [BW71]. Interlingual [Ric58]. Intermediate [Els79, Kil81, SC95, SW83]. Internal [BN78, FL85, NR92, BF78, KHPS61, Tho60, Woo70d, Woo70c]. International [Ano75a, Gat65]. Internet [TH98]. Interpolating [Hos70, Win69]. Interpolation [Du96, HKA72, Hos74, MKB87, Mau73, McL76, Sal70b, Spä69b, TI75, Vit75, Bic68, Bic69, For75, For72, Hal68, HK72, Hun68, Lar67, Pow67, Sal72, Sal73, Sco69, Spä69a]. Interpretation [Car70, Kin71, Wu97, Ber87b, CFS89, Kin69a]. Interpreter [MA89, Pag76, Fos65, FE69, Rey71b]. Interpreter-Oriented [Pag76]. interpreters [FL89]. Interpreting [PT90]. interpretive [Bem74, Daw73, Hor60, KB62, Mac74, Rob59, SS63b]. Interprocess [PB86, Mar82]. Interrogating [Ran74]. Interrupted [Wil83b]. interruption [Wet66]. Intersection [BKMT96, BKMT97, KPP^{+}95, NW93, Sab76, YG93]. Interval [FvLMS98, GH73, Han70, KKR96, LPS94, RL75, TL98, VT87, vW93, Dew71, Dew72a, Dew74, OS69, Ruž91]. intervals [MKP94]. intra [RV93]. intra-media [RV93]. Intrinsic [Wal98, Coo72b, Ecc71, McC70b]. Introducing [HK80]. Introduction [BPTV92, Bro74b, DBCRW92, EMCO92, EMO92, Eme80, HR73, IL74, Pau80, RL73, Ull73, WS73, WB99, Bai88, Bry59, Dic91, For65, Pol60, Sla60, Tho61, TL85c, Woo60, Cor79]. introductory [Maz78]. Invariant [Sco70, McC71b, SM71]. inventory [For66, FGR67, Lew65a]. Inverse [Fle68, Pow69, Ull88, AL72, Boh68, May66]. Inverses [PW70, Tew68]. Inversion [DC90a, Dur74, Pie82, NE67, Pie71, Stu84, ZL73]. Inverted [Ing74, JW82, McD77, NS77]. Invertibility [Hie96, Hie97]. inverting [Goo75]. investigation [All66a, Dea76, Mil81b, Pak67, Phi74, Tal84]. Investment [Wil90, AB84]. Investments [Coc97]. invocation [ESK95]. invoicing [Jen61, Pal62]. involving [LW66b, MP65a, MP66, McC78, Tay70b]. IPSEs [Le 88]. Irregular [Cra72, BJ80, JKW94, LM68]. IRS [AM92]. Irving [PCSG63]. Irwin [PCSG63]. ISAM [MKP87]. Island [CDZ74]. ISO [ATL93, Nau79]. isolation [JBJ84]. Issue [GV99b, Ham93, Hon93b, Hut93, KP97, Kum99, Mel95, Mof97, NF94, SK93, Mur98]. Issues [KU99, Mir98, Nel79, VP80]. items [Ada76, Woo70d]. Iteration [BBP94, Jar68, MH94, Mid76, Car66, CJ70, DM64, Eva63, Eva64, FZ94, Gor79, Jar66b, Lew71b, Pri68, Sea71, Tee63a, Wol69, Wri63]. Iterative [Ahm72, AM98, BE82, BW71, CM79, Er84c, Gow58, Jar67a, Lew65a, MT61, Osb63, Phi70, Roh87, Wen59, van88, BW69b, DMS^{+}89, Erd65, EF63a, Eve63, Ged96, Had71, Hun61, Jam87, Jam89, JN65, Jar66a, Kha63, Lar64, Lu89, Maj94, Nor64, Pow62, PM72, Tew69b, Tha62]. IV [Ehr72, Mee72, Art75, CPR75, Fin72, HV74, Tay76].

J [BB80a, BL81a, CF81, FJA80, Fox87, GS73, HR73, IL74, JR81, Kil81, Llo74, NLW74, NQ80, TCB71, WY71, WD80, Wre71]. Jackson [Poo91, PL94b, SW91b]. Jacobi [Cha62, Loi72, Sea69, Wri63]. James [WY71]. JANUS [Lav70]. Japanese [Tak59]. Jape [BS99]. Jardine [CW70, Roh75]. Jardine-Sibson [CW70, Roh75]. jet [Geo63]. jigsaw [HM98]. JK [Kai72]. Job [Bar74, Fra76, HR73, Li98, Moo79, NP68]. job-shop [NP68]. Jobs [GP81, MS74, Pag62]. jobstream [HPLG79]. John [Wil70]. Johnson [VR81b]. Johnson-Trotter [VR81b]. Join [Bra86b, Bra88, She95, Mul92]. Joins [HR97, Seg87]. joint [Ada76, Har80]. Josef [YFM71]. journal [BB89, Gil73, Pag73, Sha82, SPMD83, Daw74, KTW^{+}73]. journals [Chu68]. Jr [NLW74]. JSD [GGS89]. Judicial [Cla88]. July [Ano75a]. Jumping [ASG79, Atk84, Hil80, MR82, Rob80, Tay80]. Junction [Sav69, Sav70]. June [Gol59]. Just [NN97]. Justification [Pri81].

KALDAS [DP66]. Karney [Wil70]. Karp [Akm91]. Karp-Luby [Akm91]. KATE [Tes77]. Katzan [NLW74]. KDF [Dun63]. KDF9 [Bar66, Dun62a, Sam66, Bar65b, BHJV66, Dun62b, PL68, WR67, WHM71]. KDF9ALGOL [Mul69]. Kendall [Sla71]. Kennedy [CF81]. Kent [Nee63]. kernel [UC90]. Key [BD82, BK79b, CC87, SB85, BDG99, CLC92, Col63b, Hun85, TH82, Wic87]. Keys [SD78, Pan70, SS96a]. keyword [Bon71]. Khalil [BL81a]. kind [AW71, BFMW64, Lew72, Lin69, RD80]. Kinds [Sol99, Hut89]. Kit [BS87, AFWH88]. Klint [Kil81]. km [Fen90]. Knapsack [SM88, AO80b]. knitting [CZW94]. knots [Hos70, HKA72, Hos74]. know [Hig63]. Knowledge [BDB92, Fro85, KC88, PK86, BMNS87, BKM78, Kni93, MMH87, Rus86, Sar96, Tay90]. knowledge-based [BMNS87]. Knuth [Dic91, Thi86, War94]. Kochenburger [GS73]. Koffmann [NPCA81]. Kolmogorov [Ris99b, GV99a, GV99b, She99, WD99a]. Kriegspiel [BW75b, WBB72]. Kron [Bow90]. Kuo [WS73]. Kutta [Whi72, CL78, Day66, Eng69, Kar66, Mar58, O’R70, Scr64a, Whe59].

L [BL81a, Kit74, NPCA81, RL73, Ris99b, TCB71, Wil67b, Wil70]. Label [KKR96]. Labeling [BGO94]. Labelling [SK85, UHS85, UVS87, HW90]. labels [Pag82a]. Laboratory [Var67, FSWL82]. Lady [SHH81]. Lagrangian [Sal73, Sal72]. lags [Mea73]. Laguerre [GZ72]. Lamé [Wil65b]. LAN [Par88]. Lance [Sib71]. Lanczos [Wri66a, Kiz66, Wil58b, Wra66]. Language [ACDP91, Atk76, BBG^{+}76, BLR94, Bla92, BHW78, DB90, Ear75, ER80, ER86a, Els79, Fro83, Gri90, Har84a, HS83, Hig76a, Ili69, JJ67, JK80, Kil81, KP79, Mac91, MM84, Mil70, Moo79, NF76, NF80, NPCA81, PDY74, Pie74, Pou92, PJ80, Rey90, Sar90, SC95, Sme92, Sta77, SMY95, TL85a, WD80, Wol78, YCC92, YP86, BBG^{+}63, Bag61a, Bar72a, BPG87, BLY70, BHY89, BS73a, BG79, Bow73, Bra63, BFH61, BM60, Bro73c, Bur67, BKK78, BL62, Cam69, CMRW72, CCL72, Cle66, CW62, Cor65, CP62, Dak75, DM60, EB89, Elc65, ELR78, ESK95, FS66, FL89, GCS89, God96, Gri66, Han94, Hig76b, Hol78, JB94, KL77, KM74, Kel98, Ken80, Ker63, KM65, LWS68, Lay85, Lee72, LMP79, LP73a, Luk67, MP95, Mar91a, Mol71, Nap67, Pal73, Par75d]. language [Per76, Rey71b, Ser81a, Ser81b, Smi68, Smi70b, Sta71, Ste74, SB81, UC90, UK74, Wan80, WP75, Wil68a, Wil82b, WN81, YAB92, Bag62, Cor62, HR73, HR74]. language-compiler-computer [Cor65]. language-oriented [UK74]. Languages [Ano75a, ABM88, Bar74, Bur88, CTW94, Cow86, FD88, Har82, HK88, HH80, Hou86, LP92, Pag76, Pag81, PW90, SS92, TL85b, WD74, WF89, Bac71, Bar72b, BML94, Bra72, BMRS90, Dan92, ER71, Gha72, Goo75, Hil72, HSP66, KL89, Kom74, Pai65, Pal72, PC78, PRO80a, Pey89, SEP89, Tal84, Tes77, TP91, TL85c, Wal82, Wal71, Wat91c, Woo69b, Woo70b, Woo71a, Yel89, FSWL82, Tre82]. Lano [BB80a]. LANs [TH98]. Laplace [AL72, BW71, Dur74, Had68, Had71, JM68, LM68, Pie71, Pie82, Ran68b, ZL73]. Large [Chr73b, FM83, KA98, PCSG63, RKA92, TM95, Wha90, BP79, BPSS72, Cha63, CDHL93, Che64, CRS^{+}93, Cla77b, Com75a, Com75b, Cut59, Dou65, GK96, GF91, Har88, Kec87, Kha63, Lew65b, Liv60, MMH87, Rig64, Rit89, Sch76, SWZ93, SK76, SWA97, Sla60, Smi66b, Sut60, TH82, Tem72, Wil59, Woo69c, Wri60]. Large-Capacity [PCSG63]. Large-Scale [KA98, BPSS72, Cut59, GK96, Har88, Tem72]. late [Kol83]. Latency [Sam81]. Latent [WB98, DP64, Han62]. lattice [Pal88]. Lattices [Sar96]. Laurent [BS73b]. Law [BM94b]. lawyers [HK67]. Layered [Mar92]. layout [KK94]. lazy [FL89, AJ89]. Lazy-ML [AJ89]. LCF [Age95, Pau85]. leading [Chu68]. LEAPS [Rya63]. Learned [Pau85]. Learning [Mic91, WHF83, MG85, Mic63, Mic82, OF93, Ski96, FSWL82]. Least [Fle68, Mir98, PW70, TI75, ABBS83, Cad61, Del77, JM68, Jon70, Kha63, LR72, Pow67, PM72, Rei67, Ros60, Tew69b, ZL73]. least-cost [ABBS83]. least-square [Kha63]. Least-Squares [Mir98, TI75, ZL73]. Leathrum [BL81a]. Leaves [OW83b]. Lebow [PCSG63]. Left [Vel82, Kom74, Tak96, Woo69b, Woo70b]. left-to-right [Tak96]. Legal [F^{+}73, Lei86, McO66, Bel70a, BB71]. Legendre [Cha68, Ede65, GL64]. Legibility [IB96]. LEGOL [Sta77]. Lehman [Cor79]. Lehmer [Ful76]. Length [Arg91, CT97, GW73, Goo85b, Ham69d, Ris99b, TM95, WD99a, Wel72, Art75, CCL72, HW67, Hon71, ID73, Lle85, RM89, Sam72a, Tay76, Tur79]. Length-limited [TM95]. Lengths [Bra85, Bra86a, KCP99].

Lenient [GP90]. LEO [Ano60, Cam97, Gos60, Lew63, Pin61]. Lessons [Les83b, Pau85, Lic89]. let [Mus66a]. Letter [CL86, Par67, RW62, Str65b, Gla73, Ros63b]. Letter-oriented [CL86]. Letters [KTW^{+}73]. Level [Dea79, DBH85, Dob84, FR87, Geh83, Ger70, GKS79, Tre82, TL85b, Wil85b, Won99, Aki88, Bac71, BH68, BPG87, BLY70, Bra72, Bro60, DN64, EB89, ER71, FH77, Hav95, Hil72, Ken80, LW67c, Pag82a, Pal72, Per76, RW61, Wal82, Wal71]. Levels [Sch77a, Duk66, Jon66, Rac65]. Levy [BL81a]. Lewin [CF81]. Lexical [AM97, HV78]. Lexicographic [Er87b, Er88b, Irv84, Boo67i, Er87a, Spo85, Wil83b]. lexicographical [Roh90, Roh91]. Lexicographically [Er85, Er83b]. Librarian [CGRJ92]. Library [HW92, Not70, Teu87]. Liebmann [MST64]. Life [Sca85, Gil64, Ros82, Sut60]. light [Bry59]. Like [Har84b]. limit [GlM68, May64]. Limited [Aki77, Car70, Kin69a, Kin71, TM95]. Limiting [GH75]. LINC [SPMD83]. Line [Bre82, Ear80, Jon71b, Par75a, BBG^{+}71, Bar69, CW71, DP76, Deu69, Eva62, Eva65a, Fra70, HP71, MW68, MMG91, Phi68a, PG82, PL68, RW67, Sab72, Spä68, WM69]. line-syntax [HP71]. line-thinning [Fra70]. LINEAL [KU74]. Linear [Abe85, BE82, Cha93b, CM93, Dob91, DD87, DD88b, Fen91, Fle75, Fle68, Had90, HY81, Jak85, Kra71, KKR96, KS97, Lar85, LW93, Lyn72, OT91, PS96, Pow65, Tew69c, Wil85a, YFM71, Aha65, Akk73, AGNS91, AHM74, Bak64, Bar65a, BSD60, BCH91, BE73, Boy60, Bur67, BP69, Cas75, CF89, Dei95, DS67, DA77, Del77, EF63b, Eva64, EA70, Eva71, Eva72, EH76, FB75, FGH71, GV91, Gol78b, GE71, GlM68, Has61, Hit78, HM77, Imb93, JS67, Jen66, Jon70, KU74, Kro70b, Kro70c, Kro70d, Kro70e, Law69, LC90, LS78, Lit70, Lyn64, MT61, MB66, Nor64, O’B59b, Ord64, OW67, OW68, Øst79, Rei67, SWZ93, SS63a, Scr65, SK76, SG78, Sne70, Tew72, Tho68b, Til69, Tza73, UK74, WA91, Wit80, Zam74]. Linear-Phase [Lyn72]. linearisation [Tho72]. Lines [Ber76, CP87, DF90, Ric71, Sal70a]. Link [BW75a, Com75a, Com75b, Mar94, Def77, Pit65, Sib73]. linkage [MCC^{+}95]. linked [NO92b, ON94, Tho72]. linked-list [ON94]. Links [MH98]. Liouville [GM65]. Lisp [d’I69, Jor73, Sip88, Der79, PBF91]. List [Atk76, Ing74, Mil70, RC70, Bar65b, BCH91, Boo67e, Bro79a, Col63b, CW62, Fos65, Gor79, Hou69b, NO92b, ON94, Tru66, Wis64]. list-processing [Bar65b, CW62, Fos65, Hou69b, Wis64]. listed [Woo69d, Woo70f]. Listing [Er87b]. Lists [Wil65a, Elc65, WJ61]. literacy [Osb93]. Literal [Miy90]. Literate [Knu84, Thi86]. literature [DGM68, Sco82]. LITHP [Ham66a, Tru66]. live [Mil95]. livelock [Pac88]. livelock-free [Pac88]. liver [WY62]. LKBs [MMH87]. LL [BS73a]. Load [BM93, EM81b, GIB90, HLK94]. Load-balancing [HLK94]. loading [Cle67, Sha58]. Lobatto [JS67]. Local [CD69, GP88, IEH^{+}85, MHS85, Dee80, Mal59, McC71b, SM71]. Locality [SM97, BM84, Leu83b]. Locally [Sib78, WW76, ABBS83, CC93, TB86]. Locating [Jar67a, Bae62]. Location [CCL98, DS99, EHH86, ZSP99, vSHBT98]. lock [Sch76]. Locking [BJ98, Wet81, BS84, Che95c]. Lockout [Lor76]. locus [Mea73]. Lof [Ire93, BK87]. logarithmic [Cha67, WA91]. Logbooks [vBFvdW96]. Logic [AM98, CC92, CG92, DRT96, Edw78, Eme80, EH78, Gal92a, Gal92b, Gol82, HR73, Kha91, Lei86, MK96, MK85, Mar91b, Mel95, NL90, PL74, SB99, Sco74, War86, YCC92, van86, von95, AA68, Ben72, BVH74, BKK78, Cho95, Cla90, CRS^{+}93, CFS89, DPT94, DW77, DZW75, Ede63, Edw72, Gib88, GRB93, Hur79, KL77, LEN63, Lew71a, LQ96, MS97, Sal87, VH72, Wis64, FSWL82, PCSG63]. Logic-Based [MK85, Mar91b]. Logical [JCCJ99, Meg96, Tay90, Wel80, JL80, Ree80b, Tal84, Wil68b]. London [Wil67b, CGH^{+}68, Dol66]. long [LC95]. long-duration [LC95]. Longest [Van91]. Longo [HR73]. Look [Wil77a]. Look-Up [Wil77a]. lookup [Mad80, OW66]. Loop [DFO79, KP97, NN97, SM97, Tak86]. Loops [LF97, PD96b, JC97, KPT96, MT83, SP82]. Loss [Sch83, Kni93]. Lossless [MW97, Mof97, SH97, TWMG93]. lossy [WBM^{+}94]. lost [YG93]. Louis [RL73]. Lovelace [SHH81]. Low [FH77, Bac71, ER71, Hil72, Wal71]. lower [Smi66a]. Lsix [Hou69a]. Ltd [Wil70]. Luby [Akm91]. Ludgate [WR72, Ran71]. Lump [SNH76]. lumped [Ree84b]. lumped-state [Ree84b]. Lyapunov [Man71a].

M [BL81a, BS73b, Den80, FSWL82, HR73, NLW74, NQ80, Sla71, Wil67b, CDCL97, Smy91]. M-way [Smy91]. Machine [AD76, BC73, GH75, Hie97, Kac91, KLB91, KRK73, Mic91, PF86, Tre82, WD80, Ahm73, Bem74, BG79, CD77, CCJ91, Cur95, Daw73, EAL66, GN88, Hie98, IWT73, Mac74, NDD96, Nau93, PGM84, Ran71, Ric58, RB86, Sal74, UC90, WR72, Clo87, Ili69, TL92]. machine-checked [Cur95]. machine-independent [CCJ91]. machine-readable [RB86]. Machines [Nar96, Shi85, Ski91, BB94, Fox61, Tza73, PCSG63]. Macro [Bro84b, TL85b, Bro69b, Cam78, Dan92, Lay85, Nap67, Tes77]. macro-processor [Tes77]. macrogenerator [Str65a]. macroprocessor [Mac71, Nud77]. Macros [Mid76, Bow73]. Maddison [Bre81]. magic [AGNS91]. Magnetic [Edw77, ID73, Wat75b, Wri83, BSD60, Cha62, DGM68, EAL66, Gil66, Goo64, Gri59, Liv59, Wat71]. magnetic-tape [BSD60, Cha62, Goo64]. magnitudes [Boo67g, Hil69, Woo70g]. mail [Mit89b]. Main [BK75, HH76, LE86, BBH^{+}63, BRC66, Gla69]. Mainframe [CK95, Hea81, Wat82]. Mainly [Mau73]. Maintainable [Hut86b]. Maintaining [Lev96]. Maintenance [BIY88, CCD86, JSKCG94, Orm91, PSK96, BCH91, FW83, HK82, Pag60, Ros66]. Majority [CM91, Gri92, Wat77b, HW74, Par74, Van73]. make [BJ79]. Makes [Sha86]. Mäkinen [Coh01, TAC01]. Making [WCH78, Ben70, Coo89b, TB82]. Mal’cev [TCB71]. man [Bri67, Mus66b, Mus66a, Mus67]. man-computer [Bri67]. Management [BY79, BK75, BL90, CO76, CCL98, Coa74, Dea74b, Dea79, FB99, Gor78, GL94, H. 68, Har82, HM70, KU99, KSD77, NW74, OD91, PDY74, PSD^{+}81, VCB93, Wil84, ZSP99, Ber87b, Dea74a, Dou59a, GG59, KMR94, Lew65b, Lic89, MBC^{+}86, Noo84, PS74, RSS87, SD73, TU96, NLW74, NPCA81, VP80]. manager [Fis76, Hek89, Dix85]. Managerial [Nob91, PRLW93]. Managing [LB92, ZG99]. Manchester [Bro58a, Bro58b, HPS61, KHPS61]. Manipulating [Mar91b]. Manipulation [FS87, PDY74, BMT95, Luk67, MMG91, Obr68, Ros67, Uck85]. manipulative [Bar67, BF72]. Manipulator [d’I69]. Mano [HR73]. manufacturers [Hea81, Wat82]. manufacturing [CZW94, Rus63]. Many [FR98, YCC92, Ber69, Jar75, Sal68]. Many-Sorted [YCC92]. map [BKK87, Lun71]. map-folding [Lun71]. Mapping [BHW78, FL85, She92, Sla64b]. mappings [Gri80]. maps [Sau72]. Mark [HV92, Bro83, Wil86]. mark-scaling [Bro83, Wil86]. Market [Coo60a, Gos60]. Marking [Gro68]. Markov [CH87, Hav95, Leu83a, LW83, Lle87]. Markovian [HH95, Pfl84, Ret95]. marks [Var80b]. Marquardt [TI75]. Marshall [PCSG63]. Martin [Bro74b, Wre71, BK87, Ire93]. Martin-Lof [Ire93, BK87]. Marwick [VP80]. MASCOT [SJ79]. Mass [Fra69, Hav95, SRS94]. Massachusetts [Ano74]. Masses [PS93]. Masson [Wil62a, Wil63a]. Match [DS91, Wu97, Zve80]. Matching [Eke95, How97, Pou93, SH97, Wri61b, APD99, BLPL92, Coo89c, Hex67, Hor93, PK85, Sza73, Tak96]. Mate [Pro74, Man71b]. Mate-In-Two [Pro74]. Mathematical [BT63, CDZ74, Cor79, Hou63, JR81, LW85c, LM88, PCSG63, Sla71, Ull73, FK68, Tee67]. Mathematics [BT63, Coo92b, DG58, McG77]. Matrices [Loi72, OK63, RL73, SC81, Wil58c, Wil70, Cay61, CJ70, CHPS85, Edw72, FJ66, Gou70, GRB68, GZ72, Hat76, Hur73, Kha63, LoD70, NE67, Par67, Sea71, Sea69, Tew67, Tew69a, Wil65d, Win59b, Zam70]. Matrix [Boo67b, Coo75, Gra71, Pow69, Wil62b, Wil67b, AM65, Bra74, CL83, DP64, DW73, Eva75, Fet65, Han62, Kai72, KK94, Man71a, MT61, May66, ME88a, Meg90, O’B67, RW61, SE82, SF61, Wil59, Wri63]. Matrizen [Alb71]. matter [Hut84]. Matters [Hug89, Ral61]. Maturing [Wil91a]. Maurer [BS73b]. MAX [Nud77]. Maximum [Smi71, Wet81, Wil78c, Phi68b, Pow67]. May [Var67]. McNeary [WS73]. McPOPLOG [Bru92]. MDL [Ris99a, WD99b]. Mead [Aki77]. Meadow [IL74]. Mean [Bra85, Bra86a, Deo83, GE71, Lit70]. Meaning [BMD92, BS73b]. means [Cox71, ILPP98, Sow87]. Measure [BW73a, BW75a, Pra84, WB68, Yok97, Hun75, WD79]. Measurement [CCD86, RS88, Sou76, BM84, JC88]. Measurements [DBCRW92, Fen90, Muc82, Jac71]. Measures [Lev99, RS87, Tiz60].

Measuring

[BR90, Jan63, Lem80, JC86, GT63]. meat [Mus66b, Mus66a, Mus67]. Mechanical [Cho95, HK92, Lan64]. mechanically [HM95b]. Mechanics [Pau80]. Mechanism [KS82, BP82]. Mechanisms [RKA92, O’S88, ZH93]. mechanization [FR63, Mic63, Mic82]. Mechanized [Nee63]. Mechanizing [Tre92, Wri60]. Media [LS93b, RV93, TB94]. median [Erk84]. median-of-three [Erk84]. medical [JO74]. medicine [Ler69]. Medium [CDCL97, GP88, MD84]. medium-sized [MD84]. Meek [FJA80, NPCA81, HWW87]. Meet [Mum85]. melting [ND59]. member [HW74, Par74, Van73]. MEMBERS [BC73]. membranes [O’C69]. Memorial [Ano73b]. Memories [Dob84, RKA92, Par96a]. Memory [AM99, BJ98, Bur87a, DH97, HH76, KS97, LE86, Lop94, NMM91, PCSG63, Pfl84, Ree87, Ver87, Wri83, BBKR89, CWH93, Dav82, DO72, KVKP95, LBC^{+}94, NDD96, Stu71, TF96, TU96, WA91, WR67]. memory-its [Stu71]. Memoryless [Fen74]. Merbach [YFM71]. Mercury [And66, Bro58b, FdVR58, Gib61, PM62]. merge [Woo70d, Woo70c]. mergeable [OOW91]. Merging [DD87, DĎ88a, DD88b, JD59, Sym95, Wri82, DD91, HL92, Win59a]. merits [LOKM99]. Mersenne [Hil79a]. Mersey [Hon74]. Mesh [Cra72, Eva62]. Meshes [Har76, PTV97, Bro95, Lep95, PHS96]. Message [Gro73, Ris99b, Rob88, WD99a, YS91, MR94]. Messages [Gee91, KC97, Ram99]. Messaging [MRW90, GGH^{+}87]. Meta [YS91]. Meta-Message [YS91]. metal [ILPP98]. Metalanguage [Pag79]. Method [Aki77, Bro69d, CC87, CW70, Du96, Dur74, FR98, Fro85, Gal92b, Gow58, Had90, HD72, HH94b, Hun70, Jar67a, JLV99, LV79, Loi72, Mal77, McL85, Med70, Miy90, Nas75, NM65, Oka98, Pow64, Pow65, Pro74, SL78, TUW98, TL97, Wat67, Wil58b, Wil60, Wil62b, Abb70, Akm91, AGNS91, All59, All66b, Bak64, Bar65a, BWZ71, Bow90, Box65, Boy63, Bur66, Bye67, Cam71, CC88, CLC92, Cha62, Chr73a, Cla77b, Col63b, CR63, Day65, Day66, Def77, DF90, Dun79, Ell63, Eva63, Eva64, Eva65b, Eva66, FM64, FM66, Fik75, FP63, For61, FN93, Ful76, GJM68, Gen69, Gib60, Gil58a, Gou68, GRB68, Gui68, HB70, Had68, Han62, HV74, Hay69, Hil75, HWW87, Hun61, Hun68, HSW92, Jar67b, JM68, Jon66, Kai72, Kar66, Kec87, Kha63, Kiz66, KL71, Lei87]. method [Lil70, Lil71, LP73a, MP95, Man71a, Mar74a, MP67, ML65, O’B68b, Oli72, Osb64b, Osb64c, Pec70, Poo91, PL94b, Pow62, PM72, Pri68, Rei66, Riv68, Rob79, RW61, Ros60, Sab76, Scr70, Sea69, SK76, Sib73, Smi66a, Sne66, Sow87, Sub85, Tay70a, Tay70b, Tew68, Tew69b, Tha62, Tho58b, War74, Whe59, Wri66a, Wri63, Var79]. Methodologies [CD81, FSW85, Ham91b, Mic91, Nob91, SWB87]. Methodology [BR86, HS83, Wat74a, Wat74b, ELR78, Mar76, TAP89]. Methodology-Directed [HS83]. Methods [Alb71, BW69b, Bro79b, Coo92a, EMCO92, Fle68, GWWB98, Hou63, HOH91, JR81, Laz81, LCP81, MWHC96, MM70, NV77, Nud69, O’R70, OK63, Phi70, Rob90, Sta94, Tew69c, WS73, vFV92, All66a, AW71, BM92, BS75, Box65, Box66a, Bri79, CW61, CL78, Coo92b, Del77, Eck74, EHW89, Eva77, Fox62, Goo61, Hal67, HB66, HO93, JC86, Jar70, JS71, Jar66a, Ler69, Lew72, LB76, LC93, Lin69, Mak68b, Mar58, MT61, MD71, MM71, Moj77, Mul92, MS70, OS65, Osb64a, Osb65, OW68, Pak67, Pal69, Pea69, Phi74, Ral61, Ram70, Rv88, RDPS95, Spo85, Tho61, Wat68, Whi72, Wie95, Wil59, Wis64, WP68, Wra66, Wri64, Wri66b, Zaf71, WS73]. Metric [DW88, Fle70, Pea69]. Metrication [FW86]. metrics [Nav87]. Metzger [BT63].

Michaelson [Wil67b]. micro [MR85]. micro-computer [MR85]. Microcode [Dea79, Fos86]. Microcoded [BGL^{+}78]. Microcomputer [Dea82, EM81a, Ham81b]. Microcomputers [CS85, Mar79]. Microelectronics [Jam81]. Microprocessor [ES81, WW83, MMB87]. Microprocessor-Based [ES81]. microprocessors [Hea81, Wat82, NQ80]. Microprogram [LCP81]. Microprogrammable [NR79]. Microprogrammed [BC73]. Microprogramming [Bro74a, WS74]. Microsoft [BL90]. Microtones [O’B70]. Mikhailov [Mea73]. mill [Mus67]. Millar [MHF93]. Milne [Smi66a]. Milton [Fry71]. MIMD [ZE88]. mincing [Rea70]. Mini [Gil69]. Mini-COBOL [Gil69]. Minicomputer [Gor78, Ken80, KK79]. minicomputers [Hea81, Wat82]. Minimal [CL86, CM93, LF97, SM88, Col66a, WC94]. minimax [CP66b, CO66, CP66a, MR76a, OW69, Tem69, Tew72]. Minimaxing [Mic81]. Minimisation [Aki77, Hos71, Hur68, LoD71, MS70, Pea69, ZD73]. Minimising [LT79, Pec70]. Minimization [FR64, Fle65, NM65, Dav68, FP63]. Minimize [Deo83]. Minimizing [Miy90, MB96, Pow65, Zan67]. MINIMOP [MT70]. Minimum [FB75, MF88, Pow64, WD99a, KL71, Lil70, Lil71, Roh73, Ris99b]. Ministry [Ano60]. Mintree [Boo78, HO78, Obr78a, Obr78b, Pit78, RT78, WT79]. MINX [Boo79b, PHJ67]. MIS [EDS86]. miss [DSB97]. missing [Kol83]. Mitchell [Alb71]. Mix [FG77, MS74]. Mixed [Dak65, DL97, MM70, NC76, BP94, DP73, LW71, LWS68, MM71]. mixing [Bem74, Mac74]. ML [AJ89, GM95, MA89, O’N92]. MLS [LWS68]. MML [WD99b]. Mobile [BGS99, CCL98, CL99a, CTH99, DS99, FLW99, FB99, KU99, Kum99, Wai70]. Mode [Pri78, ZSP99, Col66b, VW96].

Model [Aky92, AD76, BPR74, CL99a, Cor79, DS99, DH97, Dea74b, DRT96, FM86b, FS81, FR98, FH98, FH97, Gom81, Goo85b, GKS79, GIB90, JW98, KA86, Lal99, Mac81, MVL92, NL90, Orm91, PW81a, Ree87, Rob88, Sar90, Sla71, SK95, WB98, Wor94, Zah81, BCH91, BRP93, CPV93, CM89, DC84, Dea74a, Dee80, Fun84, KGC85, KM94, Law69, Leu82b, Leu83b, LL90, Lle73, MKP87, Mic63, Per80, RL89, Ree84b, Rv88, Sha58, TR93, Tur78, VM93, WO75, ZR88, vdWHvB98]. Model-Based [FR98]. Modeling [Cor79]. Modelling [AA99, AHL92b, AHL92a, Coh01, CH87, Du96, GH95, HJ99, Huz79, LB92, LO99, LM88, Mäk01, Meg96, Par82, Sch77a, TAC98, TAC01, vBFvdW96, BWM88, EM71, Fox78, FN93, God96, KH87, PL94b, Rob79, SC87, TLF96, Var79, Wie95, ZR88]. Models [Bas86, CNTR97, CC92, Flo76, FFP99, FRS^{+}87, Fuh92, LW83, LW85b, LW85c, Lle83, Low95, OM88, PW90, RW74, TC92, WJ74, ZE88, BS75, BS79, Coo89c, Dav84, Don78, Fur89, HTW89, HH95, Hav95, Jar70, Min88, Roj95, Ros82, WBM^{+}94]. modern [BHY89, HS69]. modest [Tho80]. modification [Ash68, Gui68, MR76a, SW91a, Weg93]. Modifications [Hit78, LS78, Pow69, SG78]. Modified [BBG^{+}76, DHW78, MM84, Eva63, Eva64, KL71, Lil70, Lil71, MD71, Sea69, SE82, SAA88, Tho58b]. MODULA [Wan80, Bai88]. Modula-2 [Bai88]. Modular [DC90b, Eme80, Wat91c]. Module [GM95]. Modules [DAEH81, MG88]. Modulo [CC87, LT79, CC88, CLC92]. molecular [Bar75, BJ78, CRS^{+}93, CR63, FR63]. molecules [Duk66, Jon66, Rac65]. Moler [Hit78, LS78, SG78]. Monitor [CR83]. Monitoring [CT94, Dea78, Huz79, PAW86, Wet66]. Monitors [Wie79]. monothetic [LW65].

Monte

[Akm91, CR63, KB62, KP79, Mus59]. MONTECODE [KB62]. Moore [WY71, Tak96]. Moring [HR73]. Morphic [FD88]. Morris [Eme80, HR73]. most [Chu68]. Motion [GLR93, Abb64]. Motivated [Bun97]. mouse [MMG91]. move [Man71b]. Movements [BK79b]. moves [BJ79, Man71b]. Moving [PW81a, LW66b, Rit89]. MP [Mac71]. MP/1 [Mac71]. MP1 [JI93]. Mu [Smy91]. Mu-balancing [Smy91]. MU5 [Eme80, CMRW72, Fra76, Ibb72, IH77, MDFS72]. MU6P [WW83]. Müller [YFM71]. Müller-Merbach [YFM71]. Multi [BT71, BJ98, CC87, GKS79, Ham69d, Ing74, Kac91, KKR96, Lun71, Mit89b, Par88, TL92, TL97, Wat77b, WW73, All59, BH68, BM93, CWH93, Che90, Com75a, Com75b, Fra68, HW74, JO92, Ken85, Lam63, LW67c, MT70, O’B59b, OKMS96, PM90, Par74, Par78, TB94, Van73, WV70, You60]. Multi-Access [WW73, BT71, Fra68, MT70]. multi-attribute [JO92]. Multi-Combinator [TL92]. multi-computer [Com75a, Com75b]. multi-criteria [Par78]. Multi-destination [Mit89b]. Multi-dimensional [Lun71, Ken85]. Multi-Disc [CC87]. Multi-Label [KKR96]. Multi-LAN [Par88]. Multi-Length [Ham69d]. Multi-Level [GKS79, BH68, LW67c]. Multi-List [Ing74]. multi-media [TB94]. multi-member [HW74, Par74, Van73]. multi-order [O’B59b]. Multi-Partition [TL97]. Multi-Processing [Wat77b]. multi-project [Lam63]. multi-stage [OKMS96, PM90, You60]. Multi-Transputer [Kac91]. multi-user [BM93]. multi-valued [CWH93, Che90, WV70]. multi-variable [All59]. Multi-View [BJ98]. multiattribute [MHLS93]. Multicast [CTH99, JCCJ99]. Multicomponent [Sto73]. Multicomputers [AM99, BK92]. Multics [Jar75]. Multidimensional [Lun75, RV78, CP68, JC88, LM63b, Smi70a]. Multikey [KP92]. Multilanguage [Bru92]. multilevel [ZH93]. multilist [CC93]. Multimedia [CW96a, HS93, HJ99, Hut93, NH89, OD91, RLD94, HGL93, RV93, VM93]. Multioutput [EH78]. Multipath [CH99, TGE96]. Multiple [BHK85, CC87, CL66, GP88, How67b, Jon72, Li98, LC98, Wu97, BSH70, CLC92, Col64b, Hea63a, Lo84, MMB87, MKP87, NSW87, PHS96, SS63b, WA91, WWN91, Wil67a, WN68, Wil65d]. multiple-access [Wil67a, WN68]. multiple-precision [SS63b]. multiplexing [MCL89]. Multiplexor [ES81]. multiplication [Joh61]. Multiplicative [DR67]. Multiplier [PB76]. multipliers [BVH74, Kro70c, Kro70d, Kro70e, VH72]. Multipoint [Jar66a]. MultiPolynomial [Man93]. Multiprocessor [Aky92, BWW81, BGL^{+}78, Bur87a, DT96, MS92, NR79, SMR98, Wil91b, MCL89, SP91, WJW95]. Multiprocessors [KS97, PS90]. Multiprogramming [LY79, HLN68, Sym71]. multiscale [JKW94]. Multisection [BES83]. Multithreading [Sin97]. Multivalued [Jaj86]. Multiway [TW77, Gow62]. Music [O’B70, Gil63, Woo66]. Mutch [Ano73b]. Mutual [HPR88, TK99, Mad98, PJ89, YJ91, YJ94]. Myers [FJA80]. Mystical [Sta85]. Myths [Tho98].

N [Eme80, Kil81, CD69, MLB63, Hal68]. N. [FSWL82]. Naive [Eas76]. name [IH77]. names [Bra63, Elc65, Ker63, Tho64]. NAND [LT79, RDT80, TD77, ZD71]. NAND-NOR [TD77]. NAND-NOR-XOR [LT79]. NAND/NOR [RDT80]. narrowing [Pad91]. Nashelsky [RL73]. national [Dru63, Noo84, Ano60, Coo60b, Gra67, Ker76, Liv59, Var67]. National-Elliott [Coo60b, Liv59]. Natural [Pau85, PJ80, Sme92, Wol78, CCL72, FL89, MP95, Mar91a, YAB92]. natural-language [Mar91a]. Nature [Tul85, Wil77b, Bri67, Sca89]. Naur [Roh68, CCH73]. Navier [Bye65]. Navigation [AM92]. Near [Luk99]. NEBULA [BFH61, Rou62]. Necessary [CP66a, Thi95, BS73a]. necessity [Dun82]. need [BS89]. Needed [GG86]. Needs [Mum85, Cor65]. Neighbour [MZ96]. Neighbour-Greeting [MZ96]. Nelder [Aki77]. Nested [LF97, dHZ97, KPT96, LL90, Mad97, MT83, SP82]. Nesting [BBP94, DK74, CD85]. Net [Tur92, Che95c, Gal87, CZW94]. nets [BBG95, Rib95, Wil68b]. Network [CZL97, DS99, EM81b, FRS^{+}87, FH97, Gee91, HLK94, HPR88, IEH^{+}85, KLB91, LT79, MHS85, RLD94, Ste77b, AO87, Boo67e, Boo67c, Boo67d, Car66, CF89, Eth71, Fun85, JI93, Jon79, JKW94, Ken80, KH81, Nic66, PPRB89, PPB84c, Sol60, Wak93, WC91, Bar78b]. Networks [AM91, CCL98, CH99, EM81a, GP88, Har84b, HCH96a, Jen99, KA98, LC98, Phi67a, Pri78, SK85, She92, TK99, Wil91b, ZSP99, van96, BP79, Bro95, Bur66, HH94a, Lip80, Mar69, Mar74a, Nud69, OKMS96, PM90, SP91, Cla92, Eme80]. Neural [Cla92, FRS^{+}87, Tur92, WC91]. neurons [CWH93]. Neville [Hun68]. Newton [MD71, Med70, Sac62]. next [Cut59]. NF [Wil62a, Wil63a]. Nicholson [Dav67, KM66, dD66]. Nimbus [Cro74]. No [Wu97]. Nodal [Cra72]. node [Boo67e, Boo67c, Boo67d, GP96a, GP96b]. node-to-node [GP96a, GP96b]. Nodes [LCH99, Boo67e, Boo67d, For75, HW90]. noisy [Cla94, DF90]. Nomenclature [Wil83a]. Non [Bay76, BKMT96, CC74, Er88b, Fle68, Goo80, HH80, JW74, LP92, Lea77, Pow65, SD78, SS92, Wen59, Wey82, WF89, ZS96, van86, Aha65, Bar65a, BL72, Chr71b, Dei95, DS67, FB75, Gol94, Goo85a, GlM68, Gri66, Has61, Hat84, HB67, JW72, JS68, Jon70, JC97, Kni68, Mic64, Nor64, Smi68, Smi70b, Tza73, WP68, Wol69, Woo71c]. Non-Analytical [Wen59]. non-binary [Chr71b]. Non-Classical [van86]. Non-Constant [BKMT96]. Non-Determinism [SS92]. Non-Deterministic [CC74]. Non-Executable [LP92]. non-hierarchic [JS68]. Non-Linear [Fle68, Pow65, Aha65, Bar65a, Dei95, DS67, FB75, GlM68, Has61, Jon70, Nor64, Tza73]. non-numerical [BL72]. Non-Overflow [Goo80]. non-paged [Kni68]. Non-Parametric [JW74]. Non-Procedural [Lea77]. non-programmer [Smi68, Smi70b]. non-programming [Gri66]. Non-Recursive [Bay76, Hat84, Woo71c]. Non-Regular [Er88b]. Non-repudiation [ZS96]. non-scientist [Mic64]. non-singular [WP68, Wol69]. non-standard [HB67]. non-statistical [JW72]. Non-Strict [WF89, Gol94]. Non-Testable [Wey82]. Non-Underflow [Goo80, Goo85a]. non-uniform [JC97]. Non-Unique [SD78]. Nondeterminism [Bur88]. nondeterministic [HRW95, Hie98]. Nonlinear [Box66b, Bro69d, GK73, Gre70, Med70, Mot79, Mot85, Myl66, Sou76, SL78, TI75, TS75, Abb80, AJ70, BRH70, Bro69c, CN63, Fie70, HB70, IZP81, MR76a, MM78, Mot80, OW69, OW71, PM72, PD79, Tem69]. nonlinearly [OM64]. Nonmorphological [Hay79, Tur78]. Nonrecursive [Bur75b, Sou77]. Noot [Kil81]. NOR [LT79, RDT80, Sal87, TD77, Wil68b, ZD71]. NOR-gate [Sal87]. norm [SKY73]. Normal [DP88, Fry71, JW74, Loi72, Ada69, But61, Woo69a, Woo70a]. normalarea [HJBH67]. Normalization [AC84]. Notable [LT71]. Notation [Bol97, DBH85, Hul86, DZW75, Ham62]. Note [And90, Ano60, Bae62, Ber69, BB71, Che64, Che65, Cob69, CND66, Col63a, Col63b, Day67, Dew72a, Dew74, Ede65, Elc65, Eva62, EF63b, FFMR66, Fet65, Fie70, FN78, GE71, Gow58, GL64, Ham69d, HJBH67, HP67, Hil69, HW74, Hos74, Hun60, Jon66, KL71, Kro71, LW67c, LW68, Lil71, Mat63, McC72, Mic81, Mil60, Mux70, O’B59b, OW66, OW68, Ove67c, Par74, PHJ67, Pit78, PP69, Ran68a, Red71, Rud67b, Ryd64, SM71, Sal70c, Sal70b, Sco69, SS63a, Tem69, WH87, Whe59, Wic87, Wic92, Win69, Woo70f, Woo70g, vR70b, And66, BY66, Bar69, Bea65, BM89, BS75, BN76, BS71, Boo67l, Boo67m, Boo67n, Boo68b, Cam71, Cha67, CPR75, Chu68, Cla77b, Dav82, Dav84, Don78, Er82a, Er83b, FP82, Ham66a, Hay77, HM77, Hou69b, HW90, Hys72, Jam87]. note [Jar67b, Jor73, Kom74, LW71, LL70, Low77, Maj94, NO92b, O’N92, Pag58, Pag63, Pez76, Pro83, Ran68b, Rei67, Rei71, RM69, Roh68, RL72, Ruž91, Sou77, Wat59, Wis68, Wod78, Woo71a, Wyn71, vdV70]. Notes [Cla60, Akm91, Ano78, BLPL92, Cha92, Fur92, JO92, NO92b, Pap68, Per65, PR98, Val74]. Notice [Ham66b, Ano75b]. Novel [VVW96, WL87, Tee63b]. November [Ano73c]. NRDC [The91a]. Nucleic [CCL87]. Null [HY81]. Number [AM98, Ano73a, BP92, BT63, Cha87, Chr71a, Hol75, McI82, Mit76, PW81a, Par92, PB76, PB78, SD82, SNH76, Smi71, And74, BP94, BW69a, Bea74, BF78, Cob69, CH60, DR67, Hon71, Hun60, MT83, MP68, PP69, Pre68, Sam72a, SP82, Str59, WP67]. Numbers [Er83a, Col67, Coo68, Edm60, Ful76, Sam72b, Tho58b]. numeral [Ros63b]. Numerical [Abb70, BFMW64, BW70a, Boh68, CM79, Dur74, Gre69, Hun72, JS67, Lin69, MLB63, OP73a, Phi67c, Phi67b, Phi70, RV78, Sco70, SS72b, SL78, WS73, ZL73, Abb64, Aha65, All66a, AW71, BY66, BL72, BW70b, Cas75, CP71, Day65, Day66, Day67, Del68, Ell63, FCHK96, Gre67, Hai69, Hen78, HH79a, Hun61, KU74, Lew71b, Lew72, LB76, LW66b, Mak68b, Mil60, MP62, MP67, New68, Osb64a, OM64, Osb69, Pak67, Pie71, Ral61, RM69, Ros63a, SS63a, Scr64b, Tom67, Twi75, UK74, Val71, WCL71, Wol69, Wol71, Wra66]. Numerik [Alb71]. Numériques [Wil62a, Wil63a].

O [Llo74, DD88b, LV79, MK82, Per80, Ros63b, Sil79]. O.-J [Llo74]. O.C.R. [Pai69]. Obituary [Fox87]. OBJ [GCS89]. Object [CNTR97, CTW94, Coo89a, CT92, DE86, FA88, GL94, HW89, HS93, Mar92, PBF91, SCC^{+}96, SMY92, vBFvdW96, CG99, Coo89c, Cro80, DPT94, DO72, ESK95, Fur89, KL89, LSV89, LQ96, MR96, OM91, PL94b, RDPS95, SZNM94, TCCH94, TCM94, VM93, Yel89, ZR88]. object-based [Fur89]. Object-Oriented [CTW94, Coo89a, CT92, HW89, PBF91, SCC^{+}96, SMY92, vBFvdW96, GL94, DPT94, ESK95, KL89, LSV89, LQ96, PL94b, RDPS95, SZNM94, TCCH94, TCM94, VM93, Yel89, ZR88]. Objectives [Par75c]. Objects [BKMT96, CR97, DMW91, Hop81, NW93, NF94, Cho99, CFL84, DPSW89, Far78]. Observations [PW74b, LR72, Sib71]. Obtain [Wil78c]. Obtaining [Bye65]. Occam [B– D96]. Occupancy [BW73b]. occurring [GE71, Lit70]. OCR [WP69]. octrees [MHF93]. Octtree [BM90c]. ODA [ATL93, KMK95]. odd [HKA72, Hos74]. Odds [van96]. Oettinger [RP63]. off [Pag58, Wal64, YFM71]. Office [Geh83, Nob91, Ano60, Sha94, Sut60, Tho58a, Tyl88, JBD73a, JBD73b]. Offices [AHL92a, AHL92b]. offs [SGR96]. offset [Pri68]. Offshore [She85]. Oggetto [Mar92]. on-ground [Tur75]. On-Line [Par75a, CW71, BBG^{+}71, Bar69, PL68, WM69]. One [BG94, Bro74c, CCD86, Lyo85, Mus66b, Mus66a, Mus67, Nas75, Pow69, Var67, BRH70, BP69, Chr71b, Col63a, Com85, Day65, Sim69]. One-bit [BG94]. One-day [Var67]. One-In [CCD86]. One-In/One-Out [CCD86]. One-Out [CCD86]. One-sided [Nas75]. one-step [Day65]. one-track [Sim69]. Only [CD82, Sal87]. Ontario [Lio67]. OO [BB94]. opaque [Gri73]. Open [Bra85, Bra86a, RBCM97, AS63, ATL93, TCM94]. open-shop [AS63]. Operating [BJ98, Das91, Dij62, Fra76, Gle62, Hop81, Kin62, Pri78, SH72, Var76a, BMNS87, Clo65, Cor62, EMM72, Fel72, FS66, HPS61, JS89, KHPS61, MT70, MDFS72, MT86, SS72c, SS72d, Tur68, VG73, WHM71]. Operation [Dea74b, Str72, Cle67, Rig64]. Operational [Mil63, FGR67, Ano73c]. Operations [Abe85, Bla82, GN83, Ham69d, LPS94, OW83a, OW83b, TCB71, Ull88, Wet81, Wre71, Ara94, CMW98, God75, Ham67a, ND59, O’B67, Rey83]. Operator [BBK87, Wil79, BW71, HL76a, Tee63b]. operators [BML94, ZD71]. OPLA [LQ96]. optical [Pai66]. optically [MCL89]. Optimal [Akl87, BGR86, DO72, GH94, GP96b, HR97, HK85b, Kle97, KW84, KPT96, Luk99, MH94, Pag82b, Pap82, SRS94, Ste67, Sto72, Sur86, Sym95, WFD98, BP94, Bra80, Cam71, Cha97, CD85, Dun73, EL82, GV91, HW90, KR89, Lep95, Leu82a, Wym73, YJ94]. optimal-fit [Cam71, Leu82a]. Optimality [CLC92, dS77]. Optimally [CCMS97, Sib73]. Optimisation [Den80, DFO79, HH76, LW85a, SMY92, TD77, BW69b, Bro69a, Hys72, MD71, OP75, PW74a, Pal69, Phi74, Pri77, Sav69, Sav70, Tay90, Tho72]. Optimising [BK79b, PG90, Seg87, Tim87, AGNS91, BHY89, SA73]. Optimization [FG90, Fir61, Gea64, HZ95, HD82, KCP99, YFM71, Box65, Box66a, Fin68, Gui68, HB70, Ram70, Tal84]. Optimized [Sin97]. Optimizing [KP92, Li98, EF63a, MW68]. Optimum [Had71, Joh61, MS74, Akk73, Car61, HK82, LM68, Ran68b, RW67, Rei66]. option [AB78]. Options [Tri80, Wak93]. orbitals [FR63]. Ord [Roh91]. Ord-Smith [Roh91]. Order [BD81, Boo67g, DW91, Er83a, Er88b, Fry71, Har92, Hil69, Irv84, Mel95, O’R70, Ram99, SB85, Woo70g, YCC92, von95, Ada76, All66a, All66b, Boo67i, Cam73, CWH93, Cho95, Chr73a, Dav67, Er87a, Eva65b, Eva66, Fyf70, Hen88, Hop71, HKA72, Hos74, Kar66, Mac73, Mak68b, Man71a, Mil60, O’B59b, Oli71, RM69, Scr64b, Twi75, Whi72, Wil83b, Wol71]. Order- [Er83a]. Order-sorted [DW91]. Ordered [AK74, DY80, Er85, Atk93, Bro79a, Col63b, Er89a, Eva64, Far78, Fos68a, Lyn64, Pag71, BPG87]. Ordering [Pro74, SN75, CND66, DM64, Jen61, Spo85]. Ordinary [Nor76, WY71, BWZ71, BS71, Cay61, Chr73a, CN63, Eng69, For61, Fox62, May64, Nor64, Oli69, Osb64a, Osb64b, OW68, RM69, Tay70a, Whe59, Wil76a, Wol71, Wri64]. ordinates [OSS69]. Organisation [Bar78b, CS77, HK85b, Moo79, Sca85, Sur86, BE79]. organisational [Le 88]. organisations [EH88]. Organising [Dea74b, Fen91, Dea74a, PS74, SD73]. Organization [Bag61b, Cha81, Buc60, Cas63, DC84, Dav84, KHPS61, MN61, Pit65, Pla61]. organizational [PRLW93]. Organizations [CN82]. organizing [NO92b]. orientated [Gro73, KL77]. Oriented [AA99, Bag62, CTW94, Coo89a, CT92, HW89, KB87, LCL92, Low95, Mar92, PBF91, Pag76, SCC^{+}96, SMY92, XTU97, vBFvdW96, Alp71, Bag61a, Bai88, BML94, Bol97, CL86, CG99, DPT94, ELR78, ESK95, For66, FGR67, GL94, JC97, KL89, KM74, KM65, LSV89, LQ96, ML65, PL94b, RDPS95, SZNM94, TCCH94, TCM94, TAP89, UK74, UPT89, VM93, Yel89, ZR88]. origin [Mul75]. Origins [Cle92]. Orion [LT71, TH64, Tay68]. Orthogonal [Pow68, SZNM94, And69, CW61, Hay69, Mer90]. orthogonalising [Pal69]. OS6 [SS72c, SS72d]. Osborne [MR76a]. oscillating [Low77]. Oscillatory [AEH75, AEH76]. OSSAD [AHL92a]. Other [NF94, ABN61, Boo67d, Box65, CD77]. OUFDM [Lei87]. our [Mus67]. outcomes [Par77]. outer [Bro73b, Wil68a]. Outline [Atk76]. outlook [Wil61a]. Output [WCH78, BS75, Dun63, HPB84, Hoa63, Sal74, SS72d, Tay76, WH87, Zar69]. over-relaxation [Car61, Eva62, EF63a, Eva65a, LM68, Rei66, Tee63a]. Overcoming [Ver91]. Overflow [Bra85, Bra86a, Goo77, Goo80, Sam76, MKP87, Sri91]. overhead [Abb70]. Overlap [Hie96]. Overlapping [BHK85, CW70, Wil91b, BKMK90]. Overlay [AA99, SAM88, BKK87]. overrelaxation [Ran68b]. Overview [BCV93, CG99]. own [Hig76b, O’B67].

P [Bur80, Cor79, Kil81, NQ80, TCB71, WY71, AF88]. P-Compressed [AF88]. PACE [PM62]. Package [PBR^{+}90, BLY70, GP80, Wil64b]. Packages [SLH83]. packet [SP91]. Packetized [Wak93]. Packing [Ben82, Coo72a]. Packs [Cha75]. Pad [SKW85]. Page [BGS99]. paged [Kni68]. pages [Wil67b, Wil70]. Paging [Hig67, Jos70, LWS69b, PB69, BCR72, Hoa73, LKK71, OL68, Wil73]. paid [Pol60]. Paige [Roh91]. paper [DMR88, Ris99b, Sib71]. Papers [Cla99, Daw99, Ano75a, Ano80a, Bre81, Chu68]. Parabolas [Smi71]. Parabolic [MM70, SB70, Eva65b, Eva66, KS71, LW66b, MM71, PC64]. Paradigm [YG93, HKT95]. paradigms [KH87]. Parallel [Akl87, Arm78, BE82, BES83, BCV93, CT94, Cla90, Clo87, CCL87, DM85, DS90, DFO79, Er83c, EW80, FRS^{+}87, Gil58b, GB83, HM93, Har92, Hel93, JM96, JJ86, Jel79, Kac91, KP92, Lea78, Lea86, Lis74, NCR97, Par96b, Pey89, RW74, SK93, She92, She95, Ski91, Tho93, WI84, Woo77a, BE79, BL72, BLPL92, BBKR89, BVH74, Cha90, CRS^{+}93, Cle67, CFS89, ES77, GK96, GHR^{+}93, HH94a, Han94, JI93, Lai94, Mil63, PGM84, Par96a, TCCH94, UC90, VH72, WWN91, War74, ZM95, NL90]. Parallelism [Bar85, MSS93, NN97, Sha95, SM97, DL76]. Parallelization [KP97, JC97]. Parameter [Cra72, Jok90, Abb80, BRH70, DS67, Jon70, OM64, Ran68b, Rei66, Tom67]. parameterised [Dan92]. Parameters [Bir79, PRS96, Hur73, KCP99, Lew71b, Mal61, Mic63]. Parametric [BMT95, GM83, Gro71, JW74, HK72, Kro70e, McC71b, SM71]. Parametrized [SM97]. parenthesis [Er83b]. Paris [Wil62a, Wil63a]. Parity [War86, Day77]. park [Lei62]. Parser [Sto86, Bar75]. parsers [Ter73]. Parsing [Wil79, Bar75, BJ78, HL76a, SEP89]. Part [BJ78, Com75b, EMCO92, EMO92, Rey71b, Rey71a, Fra61, Fra62, HPS61, KHPS61, Mus66b, Mus66a, Mus67, OS70, PL68, WN68, All66a, For66, FGR67, HS69, JBD73a, JBD73b, Mic63, OS71, Ree80a, SS72c, SS72d, Woo69b, Woo70b]. Partial [Alb71, DS91, Fox63, Goo80, NS87, Sch77b, SB70, SN75, Zve80, Cay61, Day66, DW73, Eva65b, Eva66, Hop71, KS71, Osb63, Osb65, PC64, TW64, Twi75, Zam74]. Partial-Match [DS91, Zve80]. partiality [MS97]. partially [JB94]. Participative [Mum84]. Partition [CY96, TL97, AO80b]. Partitioned [Sch83, Seg87]. Partitioning [AD95, Bel70b, FM83, HR97, LF97, Car66, CG99, OM91, Stu71]. Partitions [GH94, DMS^{+}89, Er88a, FL80]. partly [Col63b]. Pascal [NPCA81, BKK87, Sil79, Tan78, Wic82, AK81, DB90, DE86, Jan88, KS82, WC85, WC86]. Pascal-FC [DB90]. passing [Coo72b, Ecc71, McC70b]. password [APW86, ZH93]. password-capability [APW86]. Patches [McL85]. PATCOSY [Ker76]. Path [ADD86, Arg91, DE86, GLR93, Kho93, Kie88, Van91, Boo67e, Boo67c, Boo67d, DE86]. Paths [HK80, Jen99]. patient [JO74]. Pattern [Bra78, BKM78, Cla94, Hal68, How97, Pou93, Ull73, PK85, Tra82, Tra84, Ull65, Ull67]. Pattern-based [BKM78]. Pattern-Matching [Pou93]. Patterns [LW85b, ZBS81]. Paulson [Ash68]. Pavlidis [Bra78]. Pay [Sla60, TS64, Cas63, Moo63]. payroll [WF78]. PCTE [NO92a]. PDP [WD80]. PDP-11 [WD80]. Peaceman [Gou68, ML65]. Peak [EM81b]. Pebble [GIB90]. Pebble-crunching [GIB90]. pedagogical [Gra87]. peg [Lu89]. Pegasus [BM63, CF59, Gea61, Gib61, Gri59, Hor60, Ryd64, Wil59]. Penalty [DSB97]. pensions [Pol60, Ano60]. PEPA [Hol95]. Perfect [CL86, CHK85, CM93, MWHC96, CCJ91, Cze98]. Performability [PRS96]. Performance [Aky92, BPR74, BK92, Bri85, CC87, CCL98, CL99a, Cla77a, Dea78, DBCRW92, Fen90, FS81, GH95, IEH^{+}85, IP86, JW82, KSD77, Lar82, Lar85, Lep95, Mah84, MZ96, ME98, MK82, Muc82, Mul85, NMM91, NPCA81, TL97, Ver87, WJW95, Wil91b, dHZ97, CPV93, CC88, CMW98, Don78, Fox78, Gos63, GN88, GE85, HPLG79, JC88, Leu82b, LC95, LOKM99, MMG91, MB90, Rey83, WM84, Yat62, Dak73]. Performance-Based [BK92]. Performances [ES81, Cro75]. period [Eck74, Ful76]. Periodic [Che95a, For75, HK72, MC95, BE73, HKA72, Hos74]. peripheral [Cle67]. Perlog [MG88]. PERM [Boo79a]. Permanent [SH72]. Permutation [AM97, Atk80, Boo67j, Irv84, MF82, Roy78, VR81b, Boo67k, Boo67h, Er87a, Erk84, Fik75, LoD70, NP68, OS70, OS71, Roh90, Roh91, Tew67, Woo77b]. Permutations [Akl87, Boo67i, Boo67b, GB83, Hea63b, ON90, Roh76, Roh78, Top82, CF89, Er89a, Spo85, Var80b, Wil83b]. Perseus [Hun59]. Persistence [MG88, SZNM94]. Persistent [ABC^{+}83, CNTR97, PC64, DB88, DPSW89, MCC^{+}95]. Personal [Wil84, Ano76, Ham80a, Nau93]. personnel [BG71]. Perspective [LC99, TM94, TLF96]. PERT [How67a, Rob63]. Pertaining [Wri83]. perturbation [Sch94]. Peter [FJA80, PR98]. Peterson [HV92]. Petri [CZW94, Che95c, Gal87, Rib95]. petroleum [Gal59]. Ph.D. [HL76b, Lau75]. Phase [Lyn72]. Phenomenon [FE80]. Philosophy [Bla92, Cap86]. Phrase [Has74, Hou71, BM60]. Physical [ADK^{+}93, Sch77a, Var67, BMT85]. Physics [BT63]. Pictorial [DP88]. Picture [Ede63, PW81b]. pictures [SHCP91, Kil81]. Pieces [Ben82, RHS89]. Piecewise [Bic68, Bic69, Phi68a]. Pigs [Pon91]. pipe [Meg90]. pipeline [CS95, Ibb72, ME88a, TK95]. Pipelined [AM90, JM96]. Pipelines [dHM93]. Pipelining [BM90b, Kau92, NN97]. pis [Cr081]. pitfalls [Pai69, WP69]. pivoting [Zam74]. PL [FJA80, Cuf72, DO72, Mer70, TS68, Val74, Var67]. PL/1 [FJA80, DO72, Mer70, TS68, Var67]. PL/I [Cuf72, Val74]. place [Hig76b, Str61]. Placement [Bel84, CW82, Pag84, Dee85, LoD71]. Plagiarism [Jan88, VW96]. Planar [LR78, Wil85a, WM84]. Plane [Col66a, GS79, Che65, GS78, Lar64, Rei71]. Planning [GLR93, Osb93, WJ74, Gal59, Lam63, MHLS93, Mus59]. plans [War60]. plant [Cam69, You60]. plants [KM74]. Platform [CBH^{+}94, MSS93]. Play [O’B70, Bel70a, BB71]. played [Fin60]. playing [Bra80]. plot [Roh74]. plotter [Gri73, LN79, McC71a, McC72, Pit67]. Plotters [Ear80, Mau73, PW81b]. Plotting [Cra72, Che65, Le 69, Mea73]. PMS [DBH85]. pocket [BL77]. Poet [KBTB97]. Point [Bla82, BKK84, EHH86, RS88, SA79, Wic92, AH69, Bic68, Bic69, Fyf69, Has61, Kiz66, McC71b, Pri68, Rya66, SM71, Smi64, SS63b, Wic89, Wri66a]. Point-In-Polygon [BKK84]. Pointer [van88]. Points [BJ80, Cra72, GS79, Jar67a, Kio81, McL74, Sch82a, Smi71, dS80, Coo72b, Ecc71, HK82, MKP94, McC70b, McC71b, Nic66, Ord64, SM71, Sal72, Sal73, Sne66]. Poisson [Lew71b]. Policies [CW96a, Poo91, Web72]. Policy [Ham67f, She73a, She73b, Bar73, Day78, DL76, GT72, Lew65a, MT72, She72b, She74, She79]. Polish [Col63a, Ham62, Rei81]. Polygon [Bra88, BKK84, SA79]. polygonal [Par89]. polygons [Kio81, KVW89, dS80]. polyhedra [Col66a]. Polyhedral [FA88]. Polyhedron [MHF93]. Polymorphic [Cow86, Har84a]. Polymorphism [Ram99]. Polynomial [GH73, GH75, OSS69, Riv68, Wat67, BJS74, Bea68, Bra75, GJM68, Ham67a, KK94, McC93, MR76b, O’B68b, OP75, Phi68b, Pow67, Var77, Wes80b, Win69, War94]. polynomial-time [KK94]. Polynomials [Han70, OP73a, Wil65b, CD69, Ede65, For72, Gou68, Hay69, vdV70]. Polyominoes [Lun75]. polyphase [Cam73, EL82, Mac73]. Polytechnics [Fur92]. polythetic [LW66a]. polytopes [Wat81]. popularisation [Pen77]. population [TB63]. Populations [Wha90]. Port [Sil81]. Portability [NQ80, Tri78]. Portable [HH79b, MSS93, Wal82]. Porting [Cro80]. Portions [Pag81]. Ports [BPG87]. Position [Spr92]. positive [Jen71, MT61, Zam70]. Possible [Ric88, Hea81, Wat82]. Post [JBD73a, JBD73b]. Postage [Mit89a, Lun69, Sel72]. postfix [DZW75, Rei81]. postman [Bar78a, Gol78a]. Postmodern [RHHR98]. Power [BY79, Bru92, HD72, JJ86, Nor76, BS73a, Gor79, RW67, SD62, Wol90]. Powerful [FKT90]. pp [Wil62a, Wil63a]. PPM [Bun97, CT97]. PPMB [WJW95]. Practical [CHK85, Hut86b, Muc82, ON79, Paa91, Par82, Ski91, ST82, TM95, Bro73b, Dac63, GK96, HRS93]. Practice [Sta94, AWH91, EL82, Gil59, HN88, Tay81, Tho60]. Practitioner [DBCRW92]. pragmatic [Mar91b]. PRAM [Lep95]. PRAMs [ADK^{+}93]. PRAXIS [AB90]. PRBS [Jac71]. Precedence [Has74, PS90, Wat76a, Wil79, Er82a, God75, HL76a, LL70, Wat77a]. Precedence-Constrained [PS90]. PRECI [BD82, DNV81, DAODT85, WS86]. PRECI* [DAT87]. Precise [PP98]. Precision [Aki77, GH75, CL66, Hon71, How67b, Hun75, Sam72a, SS63b]. precompilation [Ber87b]. Precondition [CW98]. Predicates [WC98]. Predicting [Maz78, YA61, Per80]. Prediction [DS99, Won99, MC95, Wit80]. Predictive [MW97]. predictor [Ano80b, For61, Hal67, Ral61, Sne80, Wes80a]. predictor-corrector [Ano80b, For61, Hal67, Ral61, Sne80, Wes80a]. preferences [MMG91]. preferential [HW74, Par74, Van73]. preliminary [CO72, Phi74]. Premature [Hun81]. Preparation [Pai66, Col64a, Dav63, Smi86, Wal82]. preprocessing [Tak96]. preprocessor [Bur79, HB67]. Prescheduling [FG77]. prescribed [Atk93, CL83]. presence [Hio67]. Present [Wil61c, All66b]. presentation [RSS87]. Presentations [HJ99, VM93]. Preserving [LO75, Wil72]. Presidential [Mic64]. Press [Wil67b]. Presumed [PRLW93]. pretend [Mus66a]. pretreatment [vdV70]. Prevention [TF87]. Previously [Ham67b, Ham67c, Ham67d, Ham67e, Ham68a, Ham68b, Ham68c, Ham69a, Ham69b, Ham69c, Ham70a, Ham70b, Ham70c, Ham70d, Ham71a, Ham71b, She71, She72a]. price [Sla64b]. Pricing [BN78, EM81b, GK84, SZ74, Wil90, Zmu75a, Zmu75b]. primal [Kro70b, Kro70c, Kro70d]. primality [Ara94]. Primary [HK80]. Prime [CH60, Hil79a, HSW92, Iga80]. primitive [Sil79]. principal [Del68]. Principle [Ris99a, VG99]. Principles [Bag62, BBKR89, BB81, Clo87, FSWL82, Hut86b, JR81, Kit74, TL85b, Bel58, Fen96, SS72c, Tho62]. printed [YS61]. Prior [FG77]. Prioritizing [SS96b]. Priority [BBK87, FK89, GP88, MG95, NMM91, NT79, Sil82, Che95d, OOW91, TB94]. Privacy [Gol73]. Prize [Ano73b, Ano81b]. prizes [Gil65]. Probabilistic [Fuh92, JW74, Li98, Ret95, Sol99, van92, CY95, Hav95]. Probabilities [Fry71]. Probability [SR76, Hav95, McC78]. probable [Par77]. Probe [OT91]. Probes [Lyo85]. Probing [DJ95, Dob91, Kra71]. Problem [AD95, BES83, Bar84, BGR86, BR86, GC87, Jon71b, KL77, KPP^{+}95, Mit89a, NPCA81, PW70, Sel88, Wil83a, AO80a, Bag61a, Bai88, Bar78a, BLPL92, BN76, Boh68, BP79, Bro69a, BP69, Car64, CK93, Day66, Er82b, Er88c, ELR78, FM66, Gol78a, Hay77, Hop61, Hop71, HM80b, How68, Hut84, KH81, Lar64, Law69, Lig73, Lun69, ME90, Mot80, Mus59, Obr68, OM91, PM72, PW67, Sch76, Sel72, TPvdW96, TAP89, Wod78, Wra66, YJ91, Zak77]. problem-oriented [Bag61a, Bai88, TAP89]. Problem-Solving [BR86]. Problems [ADD86, Bel84, BW70a, BWW81, BKMT97, BS73b, Dak65, Gil64, KV84, Lor76, Mal59, Mur85, dP60, Pro74, Sco70, Sla60, SM88, Tho60, UHS85, Ver91, Zve75, tHO99, Aha65, AO80b, AH69, All66a, AJ70, ABN61, Bag62, BJS74, Bic68, Bic69, Box66b, Box66a, Bra75, BKK87, Cas75, Chr73b, CRS^{+}93, CP64, Col67, Eva77, Fir61, For66, FGR67, Fos64, Fun85, Fyf69, Fyf70, Gil73, Gil65, Gil58c, Gil58d, GM65, Gre67, Gre69, JW72, LW66b, LT61, MP65a, MP66, McC70a, MM78, Myl66, NPS89, Oom93, Osb64c, OM64, Pag63, PW74a, PJ89, Pak67, Pap68, Per65, Pri68, Rus63, Smi66b, Sol60, Str60, Tay70a, WP67, Woo69c]. Procedural [HH80, Lea77, CD85]. Procedure [BB80b, GE71, Lit70, ME80, Weg74, ZG99, GE69, HW67, HW88, HR96, Kai72, Le 69, Leu82a, LM63b, McC71a, McC71b, McC72, NP68, Pal69, PJ86, PFW74, Pri77, PD79, Roh91, SM71, TS68, Tho62, Woo70c, dD66]. Procedures [Dew71, Dew72a, Dew74, Gun70, Ham69d, RS88, BY66, CY95, Jon72, KJ75, Osb63, Wat63]. Proceedings [RP63]. Process [Aky92, BBK87, BBP94, BPR74, BKLL95, Ger70, GH95, HR74, SJ79, TM94, TF87, tHO99, Bar72b, BGMW96, Boo74a, Boo74b, Com75a, Com75b, Erd65, EF63a, FZ94, FH77, GK96, HH95, HM95a, Hos71, MST64, Rea70, Ret95, Rib95, Ser95, Shr75, Sla64a, Tee63a]. Processable [EW80, War74]. Processe [Sil82]. Processed [MS74, Wat71]. Processes [CR97, Cla78, Lis74, NW74, RW74, Wen59, BM94a, BM59, Bye65, Hai69, Nat86, RW62, Ros63a, Scr64a, Win59a]. Processing [AM90, Atk76, BR91, CCL87, Dea82, DFO79, FG77, GWWB98, GS73, HM84, KP92, Lea86, Lea77, MLR90, Mir98, MK82, MB96, OBLP75, RC70, SK93, SR76, Sme92, Ste77a, Wat77b, YS91, AF76, Alp71, Bar65b, BL72, BLY70, BFH61, CBB^{+}76, CW62, DM60, ET88, ES77, Fos65, Fri65, Gat65, Gil64, HN79, Ham81b, Hat76, Hin60, Hop61, Hou69b, Hun59, LR67, Lee78a, Lic89, MC95, Mal59, MMH87, MP65b, Mer70, Mil63, Nau79, Par77, Par78, Rus63, SI80, Sla60, Sow87, TCM94, Thi81, WKRJ60, Wis64, YS61]. Processor [BTD99, Jel79, Lea78, Sin97, Bro69b, Dan92, DW77, Fos71, GB69, LLR73, Mac70b, NSW87, TCL89, Tes77, Tru66, WH68, WN81]. Processors [Bro84b, Lea86, MF82, PS96, TL85b, War86, CP79, Hor93, Lay85]. procurement [Web72]. Produce [BPTV92]. produced [Ged96]. producing [ABN61]. Product [Har76, CC88, ME88a, Meg90, Ser95]. Production [Lei83, BB89, Bry59, HK59, Hol95, Hug61, LP73b, MP95, Mus59, War60]. Productivity [Laz81]. Products [LT79, EH88, GL64, Tyl88]. Professional [HL76b]. profile [BW69b]. profitability [Jan63]. Proforma [AB90]. Program [AB90, Bas86, BPTV92, BK79a, BKL74, CCD86, CR83, Cle92, CCS87, DC90a, Eme80, FW86, Gol82, HW77, Jos70, LO75, LB86, Mae78, Mot79, Mot85, Pai99, RS87, SHH81, Tse87a, Tse87b, War96, WB99, Wha90, Woo86, And66, AMKM66, BH63, Bar69, BHS65, Bel70a, BD72, BB71, Bir89, BW70b, BSPI65, BM80, BM60, BM61, Bro74d, BW75b, BSH70, Bur69a, BKK87, Cad61, Col66b, CHH73, Dol66, FCHK96, FH71, Fos68b, FK86, GCS89, Gib60, HK73, HLOS65, Hoa72, How67a, Hut86a, Jon79, KB62, KV63, Kro70b, Kro70c, Kro70d, Kro70e, LW68, LW71, LLR73, LB76, Lio67, LMS73, Lum64, LP73b, MP89, MWC70, MM78, Mot80, Nau93, New66, Pan70, SAHS71, Smi72, Sol60, Sow87, Str65b, SS63b, Tow77, Und63, Wet66, WD79, Wig75, YS60, YG93]. programme [Ano74, Sha82]. programmed [Lei87]. Programmer [Bel76, FR87, GN83, Wil64a, Smi68, Smi70b]. Programmer-Defined [GN83]. programmers [Fis76, Gea59]. programmes [Mar79]. Programming [All83, All85, Ano75a, Arm78, ABC^{+}83, ABM88, Bel76, BGR86, BS73b, Bur88, CNTR97, CG92, Cra66a, Dak65, DRT96, DE86, Ear75, FSWL82, Fle75, Ger70, Gri88, Gri90, HM93, Hea63a, HS83, Hel84, HS93, Hug89, Kha91, Knu84, Lee82, Lei86, LM88, MK85, Mot79, Mot85, NF76, NPCA81, OBLP75, PBF91, Pag76, Pag81, Par75c, PW90, RP63, Roh76, Ski91, Thi86, Tri80, VP80, Wai70, Wat91b, WB96, WD74, WD80, WF89, Ake66, AJ70, Bai88, BJS74, BPG87, BRP93, Box66b, Boy60, Bra75, Bra76, BM80, BFH61, Bro73c, Bur67, Bur87b, BP69, Cas63, Chr71b, Cla90, CRS^{+}93, DI78, EB89, Er86b, FB75, Fra67, Fur92, Gil58b, Gil59, Gri66, GHR^{+}93, HKT95, Hig76b, Hol78, Hop78, How59, Ing76, JS89, KM65, Law69, Lay85, LQ96, LP73a, LK76, Maz78]. programming [MW68, Mol71, MM78, Mot80, Myl66, OL78, Pey89, Phi67a, Pon91, PW74b, Sla64a, TCCH94, Wal82, Wan80, Wat91c, Wil68a, Wil82b, Wil89, Wym73, ZM95, d’A62, BT63, Bro74b, Coo89a, Cor79, NPCA81, TCB71, YFM71, Llo74]. Programs [Axf90, Bar74, Bir74, Bur69b, CT94, Cha75, CGK80, Fro92, Ham77, HW89, JP76, Jan88, KC97, Lew84, LL75, Lop94, NL90, O’B70, Oul82, Ros93, She92, SS75, Sto86, SAM88, Top82, TGE96, Wey82, AGNS91, BF72, Ber66, Bra80, Bra74, Bro58a, BM84, Cam69, CFS89, DN64, DO72, Dun78, Dun82, Gib61, GRB93, Gol94, HV78, HM80b, JP69, LW65, LW66a, LM77a, McO66, Mul69, Nie65, NI70, O’N92, Osb71, Rob59, RT76, SAHS71, Tim87, War74, WF78, WC85, Woo59]. Progress [Sme92, Pol60, Rou62, WY71]. progressive [LM63b]. Project [KMK95, Bro73a, Fis74a, Hon74, Lam63, Nud69, Par77, SPMD83, GGH^{+}87]. Projection [Tew69c]. Projections [Ull88, Col66a, Oom93]. Prolog [Foo86, KB87, Kac91, MC95, MG88, Paa91, PG90, Ros93, Shi90, TGE96, WC86]. Prolog-Based [WC86]. prolonged [CM89]. PROM [PW93]. promise [Yat62]. Proof [BG94, BS99, Der79, FH71, Hoa72, Lew84, Smi72, BM86, Cur95, PJ86, Woo69a]. Proofs [Pau85, von95, Bra77]. Propagation [KS85]. Properties [Bur69b, CCMS97, CDCL97, FD88, Kra71, Spr92, V’y99, Bar61, CLC92, Con66, Hil79a, Lyn65, MP89, Pal88, SW78]. property [Bra94, HM98]. PROPLAN [Pen77].

Proportional [CCMS97, LC99]. Proposal [Ste77a, Coo89b, Hou83, Tho80]. Proposals [HKT75, BM61, Lau75, Nap67]. Proposed [BR90, CP62, DCHR76, Tri83]. propositional [Gib88]. prospects [LT61]. prosthesis [LBC^{+}94]. protected [FKLM76]. Protecting [BGR98]. Protection [Cha87, NW74, Sho81, Wil84, CF96, McO66]. Protein [CCL87]. Proteus [BGL^{+}78]. Protocol [BG94, CDCL97, HH94b, JCCJ99, KA98, NH89, Vau95, vFV92, BS84, HCH96b, TF96]. Protocols [DSB97, MZ96, Nai97, DJ95, Pac88]. Prototype [SS81, Sta77, DAT87, Gre69]. Prototypes [DM83]. prototyping [Har88, HR96, MD87, UPT89]. Provably [VH99]. provide [McC71b, SM71]. Proving [Bur69b, MP89, Mel95, SN75, Mel66]. provision [Aga96]. Proxies [LC94]. Proximity [HAM98, JW98]. PS [PA90]. PS-Algol [PA90]. Pseudo [PW70, DR67, Edm60, Ful76, Hil79a, Hun60, May66, MP68, Roh91, Tho58b, Var77]. pseudo-inverse [May66]. Pseudo-Inverses [PW70]. pseudo-lexicographical [Roh91]. pseudo-random [DR67, Edm60, Ful76, Hil79a, Hun60, MP68, Tho58b]. pseudo-remainder [Var77]. psychology [Pla64]. publication [Ham66b]. publications [Smi86]. published [Dun78, Ham67b, Ham67c, Ham67d, Ham67e, Ham68a, Ham68b, Ham68c, Ham69a, Ham69b, Ham69c, Ham70a, Ham70b, Ham70c, Ham70d, Ham71a, Ham71b, She71, She72a]. publishing [BB89, Dun82]. punched [cGR61]. purchase [AB78]. purely [Ara94]. Purpose [ASG79, Bri79, Hil80, Jon65, LMP79, MR82, PC78, Rob80, Str65a, Tay80, Und63]. purposes [Gri62, WO75]. put [Bro73c]. puzzle [Maj94, RG86]. Puzzles [Smi83, Bur69a].

Q [Dun79]. QMC [Rob88]. Quad [Woo82, Woo84, BK83]. Quadratic [Bur75a, HD72, Loi72, ST81, Eck74, Hon93a, Jam87, Ord64, Tho72]. quadratically [MS70]. Quadrature [AEH75, Bar61, Cha68, Haz78, JS67, LM63a, LK77, MLB63, OS68, Oli72, Pie71, Sac62, Ste67]. Quadratures [Cha70, Tha62]. quadric [Sab76]. Quadtree [OW83a, OW83b, Suf90, TUW98, Wil88, HW90, JM96]. Quadtree-Based [TUW98]. Quadtrees [Abe85, Bre93, FM86a, Ste86, UVS87, BKK87, GV91]. Quadtress [AF88]. Quality [RS87, THHS93, WCH78, All71, Jon71a, MN61]. quantification [NPS89]. Quantifier [Dub93, Hon93a, Hon93b, LW93, Can93b, Cuc93, LS93a]. quantitative [Hun75]. quantities [Ada76]. Quasi [Cze98, YW82, Cay61, Coo72b, Ecc71, HH95, McC70b]. quasi-diagonal [Cay61]. Quasi-Equifrequent [YW82]. quasi-intrinsic [Coo72b, Ecc71, McC70b]. Quasi-perfect [Cze98]. quasi-reversible [HH95]. Quaternions [BM90c]. ‘Queensway’ [Hon74]. Queries [BR91, LG86, NW93, Pap82, BJ78, ON94, PG90, Tay90]. Query [BKMT97, CTW94, Mac91, MLR90, MB96, OM88, Sar90, SMY92, TPvdW96, BML94, Dan92, ET88, PPRB89, Ser81a, Ser81b, Sv83, Tal84, TL85c]. questions [BSH70]. Queue [MG95, NMM91, NT79, NDD96, OOW91, Tur79]. Queueing [Par75a, Lip80]. Queues [Har84b, Che95d]. Queuing [Sil82, Str60]. Quickshunt [McC82]. quicksort [Erk84, FH71, Hoa62a, Ver87]. Quindiagonal [SC81, BE73, Eva75, SE82]. quintic [Hos70, Sal70b, Spä69b]. Quorum [TK99]. Quorum-Based [TK99]. Quotient [Bur75a].

R [Alb71, BB80a, Bro74b, BS73b, Bur80, Cor79, FSWL82, GS73, Llo74, NLW74, RL73, RP63, Wil67b, Wre71]. R. [FSWL82].

Rachford [ML65, FM64, Gou68]. Racing [KC97]. radar [JP69]. Rademacher [Hur73]. Rademacher-Walsh [Hur73]. Radford [NLW74]. radiational [MP65a, MP66]. Radix [Dav92, Par92, BP94, Boo74a, Boo74b, Ham74, Woo69d, Woo70f]. RAID [TL97]. railway [Gil58c, Gil58d]. railways [Abb70]. Rainbow [SKW85, WSG^{+}84]. Ramacher [Cla92]. Ramon [WY71]. RAMPS [Lam63]. RANCOMB [Pik79]. Random [But61, FvLMS98, GH94, IH86, Mah86, MH94, McL76, NW87, ON90, Rea70, SW84, WN86, BW69a, Bea74, Bra94, CDHL93, DR67, Edm60, EN82, Ful76, Hil79a, Hun60, Jac73, KV63, Kol78, MP68, Pit74, Pri77, Ree79c, Ree80a, Ree83, Tho58b]. Randomized [HM96]. randomness [SL73]. Range [KP92, LG86, She95, HP77]. Range-Join [She95]. Rank [Fry71, Pow69, SB85, Boo67f, Hil69, Woo70g]. Ranking [Er87b, Pal86]. ranks [Woo70d]. rapid [Cra66b, IW86, OW66]. Rapidly [Jam89, FP63, PM72]. RAPIDWRITE [Ayr64, Hum62]. Raster [Col87, GV91, Par89]. raster-scanned [GV91]. rat [MMG91]. rates [BF78, Eva65a]. Ratio [GOW83, Ash68, BB69]. Rational [Vit75, Cob69, CO66, How59, JN65, Jar66b, Lar67, PP69, Pre68, Tha66]. rationale [WCL71]. Ratios [Wat76b, Pez76]. Raulefs [FJA80]. Ray [SMR98]. RCC [NF80]. Re [BW70a, BGS99, FS75, GP90, PD96b, WI84, AW85, Ada76, NO92b, YG93]. Re-Authoring [BGS99]. Re-Circulating [WI84]. Re-engineering [PD96b]. Re-Entrant [BW70a, GP90]. re-evaluation [YG93]. re-examination [AW85]. re-order [Ada76]. re-organizing [NO92b]. Re-Set [FS75]. reactions [Bae64]. reactor [McC70a]. Read [DSB97, BJ79]. Read-miss [DSB97]. readability [Hum62]. readable [RB86]. Readers [KRR82, BN76].

Reading [Art75, DMR88, HV74, MP85]. Readings [FSWL82]. Ready [BBK87, Deo83]. Ready-Trace [BBK87]. Real [Bar72b, BC73, BKL74, ER86a, Ham91a, Han70, HOH91, JP86, KU99, LPSC99, Low95, Oka98, Sha72a, SMY95, Wil77c, Wu96, BD72, CL83, CS95, Che95b, CL66, CJ70, CHPS85, Cob69, CK93, DPT94, DO72, FJ66, GRB68, Gri62, Hop61, Kai72, Lut80, Par67, Phi67a, PP69, Pre68, Sea71, Sto72, TB94, Tom82, Wil65d]. Real-Time [BKL74, HOH91, JP86, KU99, Low95, SMY95, BC73, LPSC99, BD72, CS95, Che95b, DPT94, Phi67a, Sto72, TB94]. Real-Valued [Ham91a, Wu96]. realisation [WS86]. realistic [Ben72]. realization [BM61, Wat63]. really [Bae64]. reals [HRS93]. Rearrange [Boo67h]. rearranging [Win60]. Reasoning [Gru96, Lal99, BGMW96, PL94a]. Receipt [ZS96]. Recipes [Coo92b]. Recoding [Wol78]. Recognition [KVW89, SCC^{+}96, VVW96, ZBS81, Bau65, Cla94, CP61, GB61, Hin61, New61, Pai66, Sha65, Sha94, Str61, Tra82, Tra84, Ull65, War74, Bra78, Ull73]. Recognizing [Jaj86]. recommendations [CF96]. recomposition [JM96]. Recompression [Kle97]. Reconciling [BB94]. Reconfigurable [PTV97, Hor93]. reconfiguration [BP82]. reconsideration [Lee78b]. Reconstruction [Kho93]. Reconstructions [CTH99]. Record [Bae75, LCL92, Tho64, JO74, KCP99, YAH80]. Record-Classes [Bae75]. Record-Oriented [LCL92]. recorded [Fai58]. Recording [PAW86, Dou58c, Dou59c, Dru63, Jen61, MTPL88, Pol60]. recordings [Bir81]. Records [Bra85, Bra86a, CN82, SD78, Art75, Gil66, Tri74]. Recovery [BDS82, Gun70, Har81, KS85, MS92, Mul81, ABBS83, Lee78b, Mad97, RM89]. Recreating [Bay76]. Rectangle [KPP^{+}95]. Rectangular [Ben82, EHH86, Har76, AO80a, BW73b, Phi67c]. recurrence [MB66, Not70, Pag71]. Recursion [Bir79, WB99, Bra77, WTW91]. Recursive [Bay76, Bla82, BGR86, Lyn72, TCB71, CC88, Eva75, FH71, Hat84, PJ86, PS92, Woo71c, Woo71b]. Recursively [Has75]. Red [BKMT96]. Red-Blue [BKMT96]. Redistribution [MSNW99]. reduce [Aus71, Har79]. Reduced [HK85a, BS89, Zam74]. Reducer [KLB91]. Reducible [KV84]. Reducing [Bas86, DSB97, Mal77, Wil62b, AM65, OS65, YJ94]. Reduction [Bur87a, CL99b, JW74, Pad91, Tho72, All59, BBKR89, Boy63, LC93, SF61, Tew69a]. reductions [Rei67]. redundancy [Bro74d, HK73]. Redundant [AM98, SD82, OS65]. Reed [PCSG63]. referee [BW75b]. refereeing [MP85]. Reference [Axf90, MS91, Dou59c]. referencing [Leu83b]. Referential [Bur88]. Refined [CCMS97]. Refinement [Cle92, LZ95, Woo92]. Refinements [LO75, WD99b, van88]. Reflections [The91a, Lav70]. regarding [Gil66]. Regenerative [Aki77]. region [Phi67c]. Regional [Bar78b]. Regions [Kam77, LR78, LM68, Par89, Rei71]. Registers [GK73]. registration [Ano82a]. Regression [Gom81, Sla62, Hea63a, Per80]. Regular [Er88b, Gol93, Sal74, SY72]. rehashing [Mad80]. Reification [Cle92]. Reified [MK96]. rejection [PFW74]. Rejoinder [Ris99c, WD99c]. Related [BMSA88, Mur85, OW83b, Spr92, BKK87, Eck74, FL80, Wil65c]. relating [Ral61]. Relation [Jaj86, LL90]. Relational [AK81, Bra86b, Bra88, CO76, CS85, Fro82, Mac81, Ull88, BM90a, BMT85, CAC82, Dan92, ET88, GE85, LB91, SS96a, Tal84, TL85c, Uck85, WS86]. Relations [HR97, Has74, Lev96, LPS94, Sha81, MP79, Pag71]. Relationship [Bra88, DP88, Par82, Wil78d]. Relationships [RG95, Azm90, HO93, Kin69a]. Relative [Wil78c, KM94, Loi68, SGR96]. Relaxation [HM83, Bye65, Car61, Eva62, EF63a, Eva65a, LM68, Rei66, Tee63a]. Relevance [Mid87, BM92]. Relevant [The91a, Tul85]. Reliability [BL81a, BK79a, LCH99, ZE88, Fun85, LC93]. Reliable [FJA80, KA98, Nai97]. Relles [FSWL82]. reloading [Aus71]. remainder [Var77]. Remark [Obr78b, Sut76b]. Remarks [HO78, Var76b, Fin60, SS72a]. Remote [FC87, ZG99, AS63, BBG^{+}71, Har70, HR96, MB66, Teu87]. Removal [WB99, Ric71, Tom82]. Removal/Introduction [WB99]. removes [Par75d]. Removing [PD96a]. Renal [Gor78]. renaming [Mel66]. Rendering [Mor89]. rendezvous [SSDS89]. Reorganization [Che95a, Fun84]. reorganize [ON94]. Repair [CW96b]. repeating [Hun60]. repetition [HN79]. Replacement [CO76, Dob84, Jak85, TW77]. replenishment [Ada76]. Replicated [CW96b, LW85a, WB96, ZG99, Cho99, MP89, YGS63]. replicating [Akm86, Tim87]. Replication [Ben85, Cav83, FD88, KA98]. Reply [Coh01, TAC01]. Report [BBG^{+}76, Cho82, Hoa62b, Mid87, BBG^{+}63, DHW76, Pol60, Rou62, Tre82, Aga96, Gol73]. Reporting [BKMT96]. represent [Ken85, OV92]. Representation [Bay76, BS73b, Er84b, Fro85, HAM98, Mae78, ME98, PK86, SW83, Bar64, BKM78, BKMK90, CCH73, Er82b, FL80, GS63, KVW89, Par89, RB86, Sab72, Sam63, Tay64]. Representations [Flo76, SD82, Kni93]. Representing [BT81, Lal99, Poo91, von95]. Reprinting [Ano78]. Reproducing [AB90]. reproduction [Dun78]. repudiation [ZS96]. Requirements [Mum85, NH89, PDY74, TCCH94, Boy63, HO93, Sam68, TP91, Zam74, dSA92]. Research [Ano73c, Har80, JJ67, TCB71, WY71, AS63, AWH91, BMPS71, Cra66a, EDS86, Fox61, Gos60, JBD73a, JBD73b, Las68, YR58, Yat62, Wre71]. reservations [Geo63]. Reservoir [TW77]. reshuffling [Er88c]. Residue [AM90, BP94]. residues [Hil79a]. resilience [Dea76]. resistance [HN88]. Resolution [Par88, FS94, Mel66, Stu84]. Resolving [LPSC99]. Resource [Lam63, NN97, Lic89, Wan80]. Resource-Directed [NN97]. Resources [GG86, SZ74, Zmu75a, Zmu75b]. respect [Con66]. Response [JP86, Li98, PAW86, TDW^{+}99, GE71, Lit70, Per80]. Restart [SH72]. restricted [BDG99, TF96]. Restructuring [WC85]. Resultant [Man93]. resultants [Hon93a]. resulting [Ano73c]. Results [BKMT97, SR76, Sto73, Bra76, EDS86, FM66, Mel66, MM60]. Retained [Hop81]. Retrieval [Bla92, CC92, Fuh92, HW92, Har80, HR74, HM70, Mac81, Mal77, Meg96, NBL95, OD91, OT91, Sme92, SW79, TC92, WK77, WB98, Zve80, van86, van92, All82, CH60, EHW89, Fai58, FE69, Gob75, JO92, Llo64, OF93, ON94, RSS87, Sv83, Smi70b, TPvdW96, van71, vdWHvB98, Nee63]. Retrievals [BK79b]. Retrieving [Lal99, Mac91]. retroactive [Kol83]. return [BF78]. Reuse [FFP99, RL89]. Reusing [ZHS94]. Reve [Maj94, RG86]. reversal [Ull77, YY89]. Reverse [HCH96b, War94, Col63a]. reversible [HH95]. Reversing [Roh90]. Review [BB80a, Fle65, Gol73, RP63, Wil62a, Wil63a, Wil67b, Wil70, BF72, BL71, Bra78, Bri79, Bur80, Cha73, Cla92, Den80, Fry71, GRB93, GT63, Her61, Hou63, JC86, Kil81, Kit74, Llo74, Nee63, Pau80, RDPS95, Sla71, Ull73, Web70, Wre71]. Reviews [Ano96a, Ham91b, MdlIDS99, RLT99, San99, Alb71, Ano96b, BL81a, BT63, Bro74b, BS73b, CF81, Cor79, Eme80, FSWL82, FE80, FJA80, GS73, HV92, HR73, IL74, JR81, NLW74, NQ80, NPCA81, PCSG63, RL73, RP63, TCB71, VP80, WS73, WY71, WD80, YFM71]. Revised [BBG^{+}63, DHW76, MKP94]. Revision [CGRJ92]. Revisited [Bar84, NR92, Wat08, van92, DĎ88a, OW80]. rewrite [AGNS91]. rewriting [DW91, HKK91, Nip91]. Rhodesia [Che62a]. Richard [Eme80]. Right [Coc97, Vel82, Tak96]. Right-To-Left [Vel82]. rigid [Duk66, Jon66, Rac65]. Rigorous [Wil61b, HKT95]. Rina [WD80]. Ring [JCCJ99, BLY70, Bro95, WH68]. Ringing [Tre70, Pap60]. rings [Pac88]. ripple [Sal87]. RISC [BTD99, TK95]. RISC-Style [BTD99]. Rissanen [Cla99, Daw99]. RMIT [AE90a]. RNS [HAAZ99]. Road [Bei72, VV94, VVW96, Bri79]. Roadmaps [Can93a]. Robert [BL81a, BT63, FE80, Wil70]. Robot [Arm78]. robots [BMPS71]. Robust [Lev99]. Rogers [Her61]. Role [Cle92, EH88, MV88, SC87]. Rolland [Eme80]. Romberg [Fox67, Sha72b]. ROMs [Cla88]. Ronald [HR73]. Root [Gow58, JBJ84, Hon71, Pro83, Sam72a, Str59, Wes80b]. rooted [Atk93]. Roots [Han70, DP64, Eve63, Han62, Jam89, Mal61]. Rose [RL73]. Rosenbrock [MR65, Pal69]. rotating [Mar82]. Rotation [van88, AO87, Pal88]. Rotational [Mak91a, Duk66, Jon66, Rac65]. Rotational-Admissible [Mak91a]. rotations [Er89b]. Roth [BH68]. Round [YFM71, Cha71, Cha72, Dew72b, Fla72, Pag58, Wal64]. Round-off [YFM71, Pag58, Wal64]. Route [CTH99, Moo59, Nic66, Pap68, Per65]. routine [BB59, FdVR58, Pag60, PM62]. Routines [Dix85, GP90, HCC74, BMR62, GL70, Lyn69, LK77, RD80, Sol68, Teu87]. Routing [FvLMS98, KKR96, SK85, TL98, VT87, BP79, GP96a, GP96b, Ruž91, SP91]. Row [Tew67]. Row-column [Tew67]. Rows [Boo67b]. Roy [Fry71]. Royal [Dav63, Sla60, TS64]. RSA [Hun85, Wic87]. Rt [ON79]. Ruckert [Cla92]. Rudiments [Woo92]. Rule [Har95, Mic82, AGNS91, BK87, MV88, PJ86, PW74b, Wym73]. Rule-based [Mic82, MV88]. Rules [HY81, HK80, Hou86, Lei83, LO99, CND66, Moj77]. Run [CT94, NC76, PG82]. Run-Time [CT94]. Runge [Whi72, CL78, Day66, Eng69, Kar66, Mar58, O’R70, Scr64a, Whe59]. Running [Gib61, Wei91]. runs [Dow69, KV63]. Rushby [FSWL82]. Russian [GT63]. Rutishauser [Alb71].

S [BL81a, Cor79, NPCA81, PCSG63, Ris99b, WS73, Wil67b]. S.O.R. [LM68, O’C69]. S.S.O.R. [DM64]. SABER [Hop61]. Safe [CY96, DB88, Coo89b]. Safety [BMD92, BM92, Sil79, WM95, dSA92]. safety-critical [BM92, dSA92]. Sale [SW91a]. sales [Mui58]. salesman [Obr68]. Samhita [PSD^{+}81]. Sample [Fry71, LMP79]. Sampling [CCMS97, Jak85, LC99, But61, EN82, PFW74]. SAMPO [AHL92b, AHL92a]. SAS [Geo63]. Satellite [Pea63, Tur75]. satisfaction [vdWHvB98]. satisfiability [FG90]. satisfying [Mil60]. Saving [ET75, RDT80, Zak84]. Scalable [Hel93, HCH96a, PTV97, NDD96, TF96]. Scale [KA98, BPSS72, Cut59, GK96, Har88, HM95a, Sla60, Tem72, Woo69c, You60]. scaling [Bra94, Bro83, PM62, Wil86]. Scan [Col87, Ste86]. scanned [GV91]. scanner [GB61, MN61]. Scanning [Gri90, PW81b, QY89]. Scatter [Kra71, Mal77]. Scattered [Sch82a, Sne66]. scenes [Ric71]. Schade [WY71]. schedule [KV63]. Scheduling [AM99, CZL97, Deo83, DT96, FG77, Low95, PS90, VH99, Wal77, Wil90, WFD98, ABN61, BMRS90, CS95, Chr73b, FG90, HJW62, Kni68, KPT96, Lam63, Lio67, LR72, NP68, Pag62, PW67, TB94]. Scheduling-Oriented [Low95]. Schema [FL85, Fro83, Sha81, Dee80, SS96a, Uck85]. SCHEMAL [Fro83]. Schemas [Jaj86, Oul87]. Schemata [PG81, Wil83a, Wil82a]. Schematics [MTW88]. Scheme [AA99, BP81, CL86, Cha87, CCC92, FKA^{+}93, HLK94, LW83, Pag84, Pik81, ZSP99, Ada76, AGNS91, Bar65b, Bar67, BM89, Ber59, BDG99, CCJ91, CC93, Coo72b, DM64, Ecc71, Fra70, Hor60, IJ62, Jam87, Jen66, Lee78b, McC70b, Nat86, Pol60, RW61, SR74]. Schemes [BDB92, CCD86, CM79, GP88, KKR96, RL82, TK99, ZHS94, dHZ97, Ahm72, And74, Bak94, Had71, LS93a, Mak89, MB90, Win69]. school [ABN61, Aus76, Law69, Lio67]. Schultz [Bro74b]. Schwarz [Alb71]. Science [Che86, Cla77a, FE80, Fis74a, GG86, Cap86, GGH^{+}87, Lau75, Mar79, Bel74, BL81a, Bro73a, FJA80, HL76b, Mic64]. Scientific [ET75, Mar91b, NF76, Sta85, Smi86]. scientist [Mic64]. Scope [Hou86]. Scoping [Mer90]. Scotland [Ano75a]. SCR [Cro74]. Screen [BL90, Dix85, DMR88, Dix85]. Seamed [ST81]. Search [Bra85, Bra86a, Dak65, DS91, FK93, JW98, KV95, Les83b, Mah86, MKB87, SD78, Smy91, Wei93, AO80b, Boo79c, Chr71b, CCB95, CM89, Eck74, FP82, Hav95, Hil79b, HM96, Mic71, Ove67c, Pal69, Par96a, PHJ67, PP79, Pri77, RHS89, Wil78d, dS77]. searcher [vdWHvB98]. Searches [Fen91, KP92, Lyo85]. Searching [Bur75a, CCL87, MKH86, Sur86, CLW98, Day69, Kol78, PH73, She89]. Seasonal [Hio67]. secant [Bar65a]. Second [BD81, Sla71, All66b, Hug61, Mil60, RD80, Scr64b, Twi75, Wil58a, All66a]. second-order [All66b, Scr64b]. Secondary [BK79b, HK85b, SB85, Aki88, BMT85]. Secret [MSNW99, ZHS94]. section [Boo67g, Boo67h, Boo67f, Hil69, Hil79b, Woo70g, dP60]. Secure [MRW90, TH98, WFD98, ZSM99, Mit89b]. Securing [RG95]. Security [AE90b, BMD92, GKS79, MRW90, Wil84, O’S94, NLW74]. Seek [Bas86, SR76, Kol78]. seeking [Sne66]. Seeks [Wat75b]. Segmentation [Hig67, Las68, DN64]. Segments [Pag82b, Sab72]. segregation [WR86]. SEL [Mol71]. seldom [Gil73]. Selected [FSWL82, Kiz66, Wri66a]. selecting [Wym73]. Selection [Bel76, Dob84, HK85a, HK85b, Kie88, MLR90, PHS96, Ris99a, TW77, AB84, BMT85, Gea59, Lew71b, MHLS93, Par96a]. Selective [PBH86, Mat63]. Self [Aki77, Dea74b, Fen91, LC94, PS74, She92, SD73, Wat75a, WL87, Akm86, Bra94, Dea74a, Hut86a, Kro70e, LLR73, Mol71, Tim87]. Self-adjusting [She92]. self-affine [Bra94]. self-augmenting [Hut86a]. self-dual [Kro70e]. self-extensible [Mol71]. Self-Indexing [Wat75a]. Self-Organising [Fen91, PS74, SD73, Dea74a]. self-replicating [Akm86, Tim87]. Self-Test [WL87, LLR73]. selfmodifying [Nie65]. Selling [Bro69a]. Semantic [ABD84, AA99, AM92, Col97, Das91, Der79, Fro85, Gol78b, OD91, Pag82a, Wil78a, WBM^{+}94, WB98, Bak94, BD80, FW83, God96, Lo84]. Semantically [Bun97, SEP89]. Semantico [Mar91b]. Semantico-pragmatic [Mar91b]. Semantics [All83, BBK87, CW98, MA89, NPS89, Pag79, Tre92, vtv92, All85, BBG95, MR94, Ser81a, Ser81b, Tak86, Yel89]. semaphore [YJ91]. Semaphores [KRR79, KRR82, FK89]. Semi [Can93a, Coo75, HAAZ99, LW83, Nar96, CL78, Tee63a]. Semi- [Nar96]. Semi-Algebraic [Can93a]. Semi-Custom [HAAZ99]. Semi-Hybrid [Coo75]. semi-implicit [CL78]. semi-iteration [Tee63a]. Semi-Markov [LW83]. Sensitive [AMN81, Ran74, Gha72]. sentences [YAB92]. sentential [Tur78]. separable [Ake66, Osb63, Osb65]. Separated [PW81a, Osb69]. Separating [NM90]. separation [BM59]. Sequence [CCL87, Hie96, Tre92, CLW98, MR94, MB66]. Sequences [Hie97, VR92, V’y99, YT79, Cla94, Far78, Hil79a, OS70, OS71, Pag71, Woo77b]. sequencing [LN79]. Sequential [BZN98, Huz79, KW84, MKH86, NV77, PL74, YT79, Zve80, HD71, KJ75, Kol78, LW66b, Sal74, Tza73]. Sequentially [MS74, Wat71]. serial [Sow87]. Series [Boo68a, GM65, Hun70, IS81, Jar68, Nor76, Sla71, Wri66b, AEH76, BBG^{+}71, BWZ71, Bea65, BS71, Cha67, CN63, Ede65, Ell63, Gib60, KS71, Mat63, Nor64, Oli69, Pin61, Scr65, Scr70, Sne70, Wat59, Wil76a, Wol69]. Server [LS93b, MHS85, PRS96, TU96]. Servers [CW96a, Lip80]. Service [AF76, FB99, THHS93, Wie79, Bag61b, Kin62, MR96, Pit65, vSHBT98, Ker76]. Services [BN78, GK84, Wil90, HG71]. Servomechanism [Coo75]. Session [Var62]. Sessions [CR97]. Set [DAEH81, FS75, GS79, HK85a, HD82, KB87, Ree84a, She95, WW82, AGNS91, Atk80, Bae84, Ben72, Con66, Coo72b, DMS^{+}89, Ecc71, Er88a, Far78, HTW89, McC70b, O’K85, Ros64a, WC94, dSA92]. Set-Oriented [KB87]. set-theoretic [HTW89]. Sets [Can93a, Ham77, KRR79, Nor76, BS89, Gri80, JC97, Wil59]. Seven [Nob91]. Several [Pow64, Box66a, GRB93, Pow62]. SGML [HV92, Smi86]. Shaded [PW81b]. Shaded-Picture [PW81b]. Shan [WS73]. Shared [Bur87a, DMW91, KS97, Bae66, KVKP95, LWS69a, MR96, NDD96, Pen66, TF96]. Shared-Memory [Bur87a, KS97]. Shares [MSNW99, ZHS94]. Sharing [LCL92, ZHS94, Coo60b, DL76, HLK94, Lew63, Lip80, MR96, Mil63]. Sharp [Bro74b]. Sharpe [Web70]. Shell [GR75].

Shellsort [Wei91]. Shift

[Fry71, GK73, GZ72]. Shifts [PB76, PB78]. shop [AS63, ND59, NP68, Sha58]. Short [Akm91, BLPL92, Cha92, Fur92, JO92, NO92b, O’N92, PM90, FP82]. shortcut [PG82]. Shortest [Boo67c, Boo67d, Kho93, YT79, Boo67e, Nic66, Pap68, Per65]. Shortest-Path [Kho93]. should [Hig63, Hil67]. SHRDLU [Kli75]. Sibson [CW70, Roh75]. sided [BJ80, Nas75]. sieve [Hoa72]. sign [Can93b]. signature [BMT95, KCP99]. significance [Tri76, Tri83]. significant [Ber69, Sal68]. significantly [Weg93]. Siklossy [BS73b]. silent [CT74, Fis74b, Gea73]. SIMD [PBR^{+}90]. similar [BLY70, BKMK90, Web72]. Similarity [GWWB98, Wha90, Yok97, LW66a, Ull67]. Simon [BS73b]. SIMPL [Pal73]. SIMPL/I [Pal73]. Simple [Bar85, Er88b, HM83, NS87, BBB68, Cam73, CS75, Col66b, Fos65, HM95a, Kec87, Leu82b, Mac73, Sub85, Weg93, FKT90]. Simplex [Aki77, NM65, Kro70b, Kro70c, Kro70d, Kro70e, PD79]. Simplicity [Wol90]. simplification [Fit73, Jar70, Wil77d]. simplified [Ber59]. Simplifying [Dix85]. SIMULA [Hel84, KM74, PGM84, IP86]. SIMULA-based [KM74]. simulate [WY62]. simulated [Gri62]. Simulating [MES97, RLD94, WA91, WC91]. Simulation [Bei72, BPR74, BCR72, BKL74, CD81, Cor79, DBH85, ER80, GS73, Gom81, ILPP98, KP79, Nel79, PBR^{+}90, Pri78, Sch82b, Sto73, Tur92, You60, Bae66, Bea68, Bra74, Bri79, BPSS72, BL62, Cam69, Che64, Cle66, Cro75, DP66, GP73, Har88, Har59, HG67, HB66, KM74, KK79, LEN63, Lei62, Lep95, Lew65a, MES97, Mus59, ND59, Pal73, PC78, Pen66, Per76, PRO80a, RWT64, SWZ93, SB81, Tem72, UPT89]. Simulation/Regression [Gom81]. simulations [Bux66, CM89, Hen88, KB62]. simulator [BBS66, Dem60, LB91, Wil64b].

Simulators [Kor94]. Simultaneous [Boy60, Bro69d, OP73b, BSD60, Bro69c, Chr73a, CJ70, Fie70, IZP81, Jen66, Jen71, LC90, PD79, Sea71, Wri63]. Singh [Den80]. Single [BW75a, HWW87, JCCJ99, Lar85, OT91, SL78, Tse87b, Wri82, Col63a, Hon71, Kro70d, Mar73, PL68, Sam72a, Sib73]. single-address [Col63a]. Single-File [Lar85]. single-link [Sib73]. Single-Probe [OT91]. Singular [Nas75, NS87, DP64, Fox67, May66, Sha72b, Tay70a, Tew69b, WP68, Wol69]. Singularities [RV78]. singularity [Cha67]. Sint [Kil81]. Site [MLR90]. sites [MW68]. siting [Lev65]. Situ [SW84]. Size [BKMT96, HH76, HD72, LV79, O’R70, Sam76, Eva62, Iga80, She89, Whi72, WCL71]. Sized [Pag82b, MD84]. Sizes [Edw77, Sto72]. skeleton [Lei87]. Skies’ [GGH^{+}87]. skip [NI70, Osb71]. SL5 [Han78]. Slices [Skl81]. Sliding [BG94]. SLINK [Sib73]. slip [MP67]. slope [Hon93a, WC94]. slopes [McC71b, SM71]. slots [Hut89]. Slotted [BBP98, Pac88]. Slotted-FIFO [BBP98]. SLPL [Atk76]. Small [Bei72, Pie74, BHRS69, BHRS70, Che62b, CW71, Col64a, Coo60a, Eva62, GP73, Hil77, HRS68, Kec87, Lic89, MP65b, Mul87, SS72c, SS72d, WH68]. small-store [Col64a]. Smalltalk [HW89]. Smalltalk-80 [HW89]. Smart [BDB92]. Smith [WY71, Roh91]. smooth [Coo72b, Ecc71, Er88c, McC70b]. Smoothing [MZ86, Med70, CD69, Hun72, dW64]. SNAP [Nap67]. SOAP [SAHS71]. Social [Mum80]. Society [Ano73c]. SODA [Tur68]. sofa [How68]. Soft [How97, WW82]. Soft-Edged [WW82]. Software [BB80a, BL81a, BPTV92, Bro79b, BM94b, Che86, DH97, Dea82, EMCO92, Fal81, FW86, FJA80, JK80, Kha91, Laz81, LE86, LB86, MP85, Mic91, MSS93, Mul69, NQ80, PSK96, PBR^{+}90, Pra84, Rin75, SWA97, SWB87, TM94, Tre91, Ver91, BM94a, Bro69b, CF96, GW88, HKT95, Hen78, HH79a, HP77, IW86, Jac79, LP73a, MBC^{+}86, MR77, Mar76, MES97, MCC^{+}95, Nef65, RL89, RHHR98, Sam68, ST89, Wil67c]. soldier [Cas63, Moo63]. solid [HTW89]. solids [Gri73]. Solution [BE82, BSD60, BGR86, Bro69c, CM79, Cay61, DS67, DAEH81, Fie70, Fle68, GH73, GH75, Jen71, Pro74, Sch85, Sco70, Sel88, SS72b, SL78, Tho68b, Wat67, Wil60, Wil59, Abb80, Aha65, AO80a, AHM74, And75, BFMW64, Bak64, BJS74, Bar78a, BWZ71, BW71, BE73, BLPL92, Bir68, Boh68, Bra75, Car64, Cas75, Cla77b, CN63, CK93, Day66, Day67, DA77, Eg69, Ell63, EF63b, Eva65b, Eva66, EA70, Eva71, Eva72, EH76, Fir61, For61, Fyf69, Fyf70, Ged96, Gil58c, Gil58d, Gol78a, GM65, Had71, Hai69, Has61, Hop71, HM77, JS67, JN65, Jen66, Kar66, KS71, Lew71b, LW66b, LM68, Loi68, Lu89, Mak68b, Man71a, McC70a, ME90, MP62, Nor64, Oli69, Osb64a, Osb64c, OM64, Osb65, Osb69, Øst79, PW74a, PJ89, PM72, PD79, Ran68a, Ros63a, SS63a]. solution [Scr64b, Scr65, Ser95, SK76, Sne70, Twi75, WP68, Wol69, Wol71, Wra66, YJ91]. Solutions [BD81, Bla82, Nor76, Wil62a, Wil63a, Zam74, AH69, Bye65, Eng69, OM91, Rei67, Tew72]. solvable [YJ94]. Solve [LG86]. solver [Del77]. Solving [BR86, BWW81, Bro69d, Chr73a, CRS^{+}93, McC93, NPCA81, PS96, Phi70, SWZ93, SB70, Tew69c, TS75, UHS85, All66a, All66b, AJ70, Bar65a, Bur69a, Dav67, Er88c, Eva63, Eva77, FM64, Fun85, GJM68, Had68, Hit78, IZP81, Jam87, Jar66a, Jar66b, Jar67b, Kha63, KH81, LC90, LS78, MP67, MM78, O’B68b, Osb63, OW68, Pak67, Pri68, RD80, SG78, Wil76a, Wri63]. Some [Abe85, All88, ASG79, Bay73, Bon69, Bro60, BM61, Bro74e, CW61, CCMS97, Dem60, Dru63, Edw78, FM66, FR80, Fin60, Gil66, The91a, Har59, HK80, Hut86b, Kra71, Lar67, LR67, Les83b, LM77b, LR72, Mid87, Nap67, New61, Pal88, Pap68, Per65, Pie71, PR98, Ral61, Roh87, Ros63a, Sib71, SR76, Sla64b, SW78, TL85b, Var76b, VR81b, WP68, Wri83, Zaf71, tHO99, Abb64, All66a, AL72, Bra76, Chu68, CR63, EDS86, FK68, Fun85, Gea61, Goo61, Hil80, Kin67b, Lic89, Mel66, MR82, Pak67, Rob80, Sab72, Sal72, Sha82, Tay80, Wol90, Cut59]. Sons [Wil70]. SOR [Lyn64]. Sort [Boo67f, Dav92, Dob91, Had90, Hil69, NR92, TW77, Woo70g, Yok97, Cam73, Cra66b, EL82, Low77, Mac73, Woo69d, Woo70d, Woo70f, Woo71c, Woo71b]. Sort-Based [Yok97]. Sorted [YCC92, DW91]. Sorter [MG95, WI84]. Sorting [AM97, CM91, CMS80, DS90, Dro87, Er83c, Fos68a, Gri92, Had90, HR97, JD59, KV84, McC82, SW84, WI84, Bel58, Col63b, Fen96, GR75, HL92, LW67a, LW67b, Lau91, SAA88, TH82, VR81a, Woo70c]. sound [Bir81]. Source [FH97, Lle87, Pie74, Won99, Elc65, PW93, Per76]. source-language [Elc65]. Sources [GV99a, DGM68]. SP [Mac70b, SM94]. SP/1 [Mac70b]. SPA [HRW95]. Space [BD82, Col87, DW88, DD87, DD88b, FA88, Har81, Har76, MM70, MCL89, Sch77a, SGR96, BKMK90, Had71, HL92, KR89, KPT96, MM71, Tay70b, WA91, Zak84]. Space-Based [DW88]. Space-Filling [Col87]. space-optimal [KR89]. Space-time [SGR96]. spaced [GM72, OP71, Twi75]. Spanning [Obr68, ADD90, Cha90, Roh73]. Sparse [RL73, Tew69c, Bra74, Cla77b, FCHK96, Hat76, Kha63, NE67, SWZ93, SK76, Tew67, Tew69a]. sparsely [Dav82]. Spatial [BWW81, Bra86b, JW98, NW93, NF94, Wor94, Bre93, JKW94, Mil81b, RB86]. Spatially [Wal98]. Special [Ano74, Ano75b, Boo68b, GV99b, Ham93, Hon93b, Hut93, KP97, Kum99, Mel95, Mof97, NF94, SK93, Tay96, Coo68, GB61, Jon65, LMP79, Mot80, Mur98]. special-purpose [Jon65]. specialisation [Clo86]. specialised [HT78]. specialized [YGS63]. Specification [BPTV92, Dro87, DF87, EMCO92, EMO92, JK80, LP92, Mai85, MMR^{+}98, MVL92, Rob88, SC95, Sil82, ATL93, Bra94, Dun79, Hol95, HO93, Kel98, Lei87, O’S94, Pag82a, SFD92, Shi89, TK95, TP91, TL85c, Wic89, WS86, Wri74]. Specifications [CCS87, Dix85, EMO92, Kuh92, Les83a, Sta74, VM94, Wat79, Wil85b, tHO99, FCHK96, FM85, GCS89, MP95, O’N92]. specified [Cra72, Hav95]. Specifying [Cla78, PPB84c, Gri66]. Specimen [Cho82]. SPECOL [Smi68, Smi70b]. spectra [SD62]. spectral [Hur79]. Speech [Sha94, LR67, Woo66]. Speed [Dea79, HH76, PB76, PB78, Ahm72, Ahm73, IWT73, Mar58]. Speeding [Bir74, Erd65, Sow87]. speeds [Fos71]. spelling [Lee74, Mar74b]. SPG [OBLP75]. Spiral [CK94, Mul85, Jon70]. Spline [BT81, Bro74b, Man74, San75, Sco69, Spä69a, AH69, BY66, For75, HK72]. Splines [MZ86, Cox71, Fyf69, Fyf70, Hos70, HKA72, HK72, Hos74, Sal70b, Spä69b, Pha90]. Split [Sel88]. Splitting [PBH86, Sri91]. spot [GB61]. Spotting [Oka98]. Spurious [Ram99]. SQL [NPS89, WWK^{+}90]. square [Eve63, GE71, Hon71, Jam89, Kha63, Lit70, Pro83, Sam72a, Str59]. squared [PFW74]. Squares [Fle68, Mir98, PW70, Pow65, TI75, Cad61, Del77, JM68, Jon70, Pec70, Pow67, PM72, Rei67, Tew69b, ZL73]. SQUIRREL [WWK^{+}90]. SSOR [Lyn64]. SST [Tow77]. Stabbing [NW93]. Stability [JW72, Kar66, BCH91, Car86, Hal67, LS93a, MP67, Tay70b, Wat68, dD66]. Stable [DD87, Mak68b, Rod70, Sym95, Eva65b, Eva66, HL92, Zaf71]. Stack [Goo77, Goo80, Goo85b, Lop94, Goo85a]. Stack-Length [Goo85b]. staff [YA61]. stage [FN93, OKMS96, PM90, You60].

Stamp [Mit89a, Lun69, Sel72]. Standard [BR90, Mil81a, MN90, Ros70, VP80, ATL93, Clo86, DCHR76, HB67, MRW75, Tri76, BP81, MA89, O’N92, Tri83]. Standardisation [NPCA81]. Standards [Bro89, LP92, THHS93, Wie79, HL76b]. start [Boo67e, Boo67c, Boo67d]. Starting [Eve63]. starvation [YJ91]. starvation-free [YJ91]. State [FS75, Gol93, Hie97, Lle83, Tza73, Woo77a, Atk84, Cla60, CW71, Day70, Dou59b, Gol59, Hie98, Lan71, Lle73, Ree84b, Sla66, TR93, Tow77]. Statement [Day78, Ham67f, She72b, She73a, She73b, She74, She79, Gal73]. Statements [PD96a, DI78, Hop78, Ing76, MP65b, Pag82a, Sal71b, WC85]. States [GG59, McI82]. Static [KS82, Mah84, Wil78a, Won99, Wie95]. Stationary [Ord64, Pow62]. statistic [LW68]. Statistical [CF81, FSWL82, Gea58, Sch83, She99, SLH83, DM60, JW72]. statistics [Gri62, O’B67, YR58]. STATLILB [FSWL82]. steady [Abb64]. steelworks [Mas63]. Stefan [Wra66]. Steiner [Col82]. Step [O’R70, Whi72, Chr73a, Day65, DHW78, FW83, GRB68]. step-until [DHW78]. Steps [Sha81, Yel89]. stereo [Ort71]. stereographic [Col66a]. Stiefel [Alb71]. stiff [Cas75]. still [PR98]. Stochastic [BKLL95, Che95c, FH97, OM91, Pri95, V’y99, BBG95, HM95a, Rib95, Ser95, Tza73]. stock [Ada76, AO80a, Dou58c, Pal62]. stock-control [Pal62]. Stoer [YFM71]. Stokes [Bye65]. stopping [Moj77]. Storage [Bas86, BK75, Bro79b, CO76, Fra69, Fro82, Gom81, Har81, Har82, Kra71, LS97, LS93b, Mal77, McI82, Mul85, OT91, Rob77, Sam81, Sch77a, Ste77b, Ste86, Wol78, Aki88, All82, Boy63, Bro60, Cam71, CY95, CPR75, Fos71, HW67, IJ62, Jen66, Kni68, Lan65, Leu82a, Leu82b, Pez76, Ree84b, SRS94, SR74, TH82, Win59a, Zam74]. Store [DK74, Mar92, BSD60, BJ79, Cha62, Col64a, DN64, IH77, Ree79c, Ree80a, Ree83, RW61]. Storing [AF88, Hay69]. Stove [RWT64]. straight [CP87, Phi68a, Sab72, Sal70a, Spä68]. Strategies [CS77, Cha81, CY96, CCL98, CL99a, CD82, DS99, Rob77, AW85, BCH91, BM59, Bra80, Bre93, BM93, LW67a, LW67b, Noo84, WLDC71]. Strategy [CCMS97, CL99b, JES98, LC99, BWM88, Lan76, NE67]. Stratified [BV92]. stratifying [Bag61a]. stream [BS89]. Streams [All90, Cla78]. Strema [Cla78]. Strict [WF89, Gol94, McC93]. String [Gre70, Gri90, Hou75, Lea86, Mil68, Rey77, BLPL92, Mac70b, Tak96]. Strings [MJ75, CCL72, Er83b, OV92]. strongly [DB88]. Structural [Bra78, Bur69b, CGK80, Car86, CP87, Cuc93, Liv60, VG73]. Structure [AM92, DC90a, GWWB98, Has74, Lah91, Mil70, SH72, Woo82, Wri83, dRC74, BBH70, Ben70, Bre93, Bri67, BM60, BK83, CS75, CC93, Dak75, FKLM76, FS94, Fur89, HLN68, Hut86a, Ire93, KM74, Mar73, MP79, MT83, MDFS72, PJ86, SP82, SS72c, Stu84, Tri74, Wil78b, Wis68, WH68]. structure-directed [PJ86]. Structured [ADD86, ADD90, Das91, DI78, Eme80, Gol82, GKS79, Hop78, Hou86, Ing76, Lal99, Lei83, LL75, Llo74, Lop94, MMR^{+}98, Mil81a, MTW88, NPCA81, Tri80, Wil77b, WO78, BB80b, Bro89, CCB95, Han94, Hoa72, Hoa73, Kol78, Lem80, LK76, ME80, OL78, Oul87, Pin77, SFD92, Wat77a, WP75]. Structures [BIY88, BV92, BHK85, DL97, Fro82, Gol85, HC78, Ing74, IS81, Mac91, Mae78, MSNW99, NMM91, SKW85, Sur86, Wat91a, Bir81, Eth71, FCHK96, FK86, HH95, Hou71, OV92, SR74, She85]. Structuring [Bay73, ER86a, FM86b, Mir98, Ran86, Rom99, Wil85b, NS77, van71]. Student [Jan88, FG90, Lem80, Maz78, TS68]. Students [WS73, GS73]. Studies [BY79, KP79, Pri78, Coo83, Gre69, GW88, Har59, Lee78a]. Study [BSH90, FKA^{+}93, Goo85b, HZ95, Hou63, IP86, Laz81, Lle83, LZ95, MB96, Mum85, Nel79, Ole91, SW79, VR92, Wei91, Aki88, Car64, CO72, GR75, HN79, HK59, KV63, Lan65, Lip80, Mad97, Osb93, Sha82, Tri78, WWN91, dSA92]. Studying [GM95]. Sturm [GM65]. Style [BTD99, Lee82, Pag81, Nau93]. Styles [PW90]. Sub [Sar96]. Sub-knowledge [Sar96]. subarrays [Smi70a]. subdivision [OS69]. Subject [Ano97b, Sta85, AJ70, HB70, LP73b, OSS69, Osb69]. Sublinear [DD87]. Sublist [NT79]. subroutine [CL83, CHPS85, Coo68, IZP81, SHCP91]. Subroutines [BL90, Ham91a, Dun73]. subscript [Rya66]. subscripted [HV74]. subscripts [Boo67g, Boo67h, Hil69, Woo70g]. Subset [Pai99, SM88, Var80b]. Subset-Generating [SM88]. Subsets [CHM94, Wil78c]. Substitution [HM83, Ull77]. Substitutions [AG98]. Substring [WK77]. subsystem [Bro82, MRW75]. Subsystems [Bru92, Fen74, Kie88]. Subtypes [BL81b]. success [EDS86, Lic89, Maz78]. successful [HM96]. successive [Boy63, Car61, EF63a, PL94a, Ran68b, Rei66, Tee63a]. such [PC78]. SUD [Ree79a]. sufficient [BS73a]. Suggested [Ehr72, Fin72, Mee72]. Suitable [CCL87, FG77, CF89, Kha63]. Suite [CL99b, LP73b]. Sum [LT79, Pow65, Bur69a, Pec70]. SUMFAC [Boo67a]. Summation [Edw78, Mat63]. summer [Ano74]. SUMT [Vit75]. SuperPascal [Han94]. Superposed [FS87]. superscalar [MC95]. supervision [Wan80]. Supervisor [Cav83, HLN68, LWS69a]. Supplement [GT63, Day78, DHW76, Ham67b, Ham67c, Ham67d, Ham67e, Ham68a, Ham68b, Ham68c, Ham69a, Ham69b, Ham69c, Ham70a, Ham70b, Ham70c, Ham70d, Ham71a, Ham71b, She71, She72a, She72b, She73a, She73b, She74, She79, Var68, Ham66b]. supplier [EH88]. supply [Bur66, Dav63]. Support [BJ98, PSK96, Tur92, Bur79, ST89]. Supporting [BM94a, BBF^{+}94, HS93, FH77, KM94]. supports [Sha94]. Surface [BS99, FR80, Oxl85, Wil77c, Cad61, CW61, Hay69, Thr69, Tom82]. Surfaces [Mac70a, McL85, NW87, MP67, Sab76]. Survey [LCP81, Mur85, Pau85, PCSG63, Tre91, And66, CHH73, FNK94, Fur92, Lon61, Mäk91b, Mur83, Nav87, Red60, Sco82, UK74]. surveys [Coo60a, YS60, YS61]. Suspected [VW96]. Swann [Hos71, Pal69]. Switching [Ram85, TD77, TH98, WV70, ZD71]. SWN [Roj95]. symbol [Day70, Lan71, Luk67, Tow77]. symbol-state [Day70, Lan71]. Symbolic [CCC92, Har79, Gol94]. Symbols [Mil81a, KGC85]. Symmetric [HAM98, OK63, SC81, ZOG85, Abb64, CL83, CJ70, CHPS85, DD91, EF63a, Eva71, FJ66, GZ72, Jen66, Kai72, MT61, Par67, PW69, RM89, Ran68a, RW61, Sea71, Sea69, SE82, Wil65d, Zam70]. Symmetrical [Has74, AM65]. Symmetries [Sha95]. Symmetrischer [Alb71]. Symposium [Gea59, Gri59, Liv59, RP63, Har80, Var67]. Synchronisation [HS93, SJ79, CF88, Nat86, RM89, Shr75]. synchronization [RV93]. Synchronizing [KRR82]. Synchronous [SS72b, MP89]. synonym [GF87]. Syntactic [Bar75, BJ78, Pag81, Cam78, Lo84]. Syntax [Bur72, HR70, Mux73, Alp71, Bar75, Bro73b, CND66, Fos68b, FE69, FK68, HP71, Ree67, Sch65, Ser81a, Ser81b, Sim69, Wil68a, Wil76a, Zar69]. syntax-analyzers [CND66]. Syntax-Directed [Bur72, Mux73, Ree67, Sch65]. syntax-driven [FE69]. Synthesis [Cle92, CGK80, Mar69, BJ78, CZW94, CD80, Dol66, Hur73, Hur79, Lau91, Wil68b, Win75, Woo66]. synthesized [ON94].

Synthetic [GP81, CW76, HPLG79]. System [Aky92, BT81, Bas86, BSH90, BY79, BPR74, BW78, BTD99, Bro74a, Bru92, BGL^{+}78, CD81, Cav83, CGRJ92, CDZ74, Cra72, Dea74b, Dea78, DM83, DNV81, DAODT85, DM85, DMW91, EMO92, Fal81, Fra76, Fra69, Fro85, Gom81, Gor78, GM95, Har79, HH80, Hop81, JP86, Ker76, KSD77, KBTB97, LY79, LV79, Laz81, LE86, Mac81, MS92, MM84, McI82, Mic97, Mum85, Nob91, Ole91, OBLP75, Par82, Par88, PS96, PAW86, PSD^{+}81, Ran86, Rob88, Sha72a, Shi90, SH72, Sta77, Ste77a, Teu87, Var76a, Wai70, Wil64a, WC86, Zmu75a, Zmu75b, Abb70, AB84, APW86, Bae66, BBG^{+}71, BWM88, BM86, Bar66, BBB68, BBF70, BBH70, BL72, BMNS87, BCR72, Bra94, BG71, BPSS72, Bry59, BHJV66, Cha92, CS95, Chr73a, CRS^{+}93, Cla77b, Cle67, Col70, Cra66b, Cra86, DP76, Dav63]. system [Dea74a, DTIR88, Dow71, Eth71, FKLM76, For66, Fox78, FS66, Fra68, Gob75, Gri62, GHR^{+}93, Hic60, Hoa63, Hoa73, HPS61, HJW62, HS82, HGL93, Hun59, Jan63, Jen61, KU74, KHPS61, KP60, KK79, Lan76, LWS69a, LWS68, Lar95, LN79, LMP79, LQ96, MBC^{+}86, MMB87, Mal61, MP85, MT70, MHLS93, MR67, MWC70, MDFS72, MT86, Nav87, NPW68, O’B67, Ort71, Pal62, Pen66, Per80, PR77, Pin77, Poo91, Rai87, RSS87, Rei67, RB85, Ros66, Sam66, Sha82, Shi89, Sma88, Sv83, SS72c, SS72d, SW91b, TH64, Tem72, Tur79, Tur68, UPT89, VG73, WWK^{+}90, WHM71, Wil76a, WM69, Wis64, Woo68, WC91, BW69a, Bea74, Eme80, WW73]. Systematic [BSH90, CR97, JSKCG94, Pag71, ST82]. Systematics [BBdD66, Gri66, Gri68, Kin67b, Ser81a, Ser81b]. Systems [Abe85, AM98, BB80a, BE82, BP92, Bar84, Bar85, BGR98, BS87, BJ98, BL91, CC87, CCC92, Cla77a, Col82, CT92, Das91, Dav81, DH97, Dea82, DM83, Den80, DBH85, F^{+}73, FE80, FSW85, Fle68, FS75, Geh83, GM83, GP90, Gun70, Ham91b, HCC74, Hel93, Her91, Hil76, HCH96a, HM84, Hul86, HOH91, Hut86b, Hut93, IP86, KU99, Kin69b, LH95, LPSC99, Li98, LCH99, LB92, Lis74, Lom88, LZ95, Low95, Mah84, MS91, MMR^{+}98, Mil81a, MH98, MK82, NF80, NLW74, NR79, NH89, NCR97, Nob91, OD91, Par75a, PDY74, PB86, PBH86, RBCM97, RG95, Rob90, Rom99, Sch82b, Sip88, Smi66b, SBRS94, Ste77a, SMY95, SMY92, Tew69c, Tho93, TCB71, TS75, TF87, Vau95, VW96, VS87, WJ74, WL96, Wat74a, Wat74b, WB96, Wic92, WHF83]. Systems [Zah81, ZE88, AHM74, Ano63, AFWH88, AWH91, Bak64, BBP98, BM92, BWZ71, BM81, BS61, BL71, BE73, Ber87b, BP82, Bri67, CPV93, Cay61, CZW94, Che95b, CY95, Che95c, Clo65, Com75a, Com75b, Coo83, Cor65, Cra66a, DS67, ET88, EMM72, Eng69, EF63b, Eva64, EA70, Eva71, Eva72, EH76, Fel72, For61, Gea80, GF91, Gha72, Goo64, GE69, GE71, GlM68, Gri66, H. 68, Ham80a, Ham81b, HM80a, HKK91, HN88, Hit78, HM77, HG67, Imb93, IZP81, JC86, JS89, Kar66, Kec87, Kel98, KMR94, LW67a, LW67b, Lan76, Lav64, LC93, Lip80, LS78, Lit70, Liv60, LQ96, Lyn64, MV88, Mak68b, Mar82, Mea73, MCL89, Mic80, Mil95, MES97, MB90, Mum80, Mum84, O’S94, Øst79, Par78, Pen76, Pen77, PRLW93, PL94b, PS74, RV93, Rig64, RM69]. systems [Ros82, Ros66, Rus63, SI80, Seg87, SRS94, SK76, SWA97, SG78, Sne70, Spu88, SD73, Sto72, Str61, TCCH94, TCM94, TF96, TB82, Tho72, Tom82, TU96, UK74, Var80a, Vin64, WJW95, WS74, Wil67a, WN68, WM95, WN81, WHF82, Zam74, dSA92, BB80a, BC73, CF81, GS73, Hou63, IL74, JR81, LM88, PCSG63]. Systolic [Par92, WI84, LC90, ME88a, ME88b, Meg90, ME90].

T [BB80a, Bra78, Her61, IL74, Kil81, Wil70]. Table [Bur75a, HD72, Lew84, Lyo85, Sam76, Sam81, Wil77a, Car70, Day77, Fro81, Kin69a, Kin71, Mil60, OW66, PH73]. Tables [AK74, Bay73, Mae78, SD78, Sch85, SS75, Van91, Col64a, Day70, Gow62, Gri68, Kin67a, KJ75, Lan71, Mad80, MV88, Sha72b, Tow77, WL72, YS61]. tabulate [Cra66b]. Tagged [MG95]. tailoring [FCHK96]. taking [Str59]. Talbot [YFM71]. Talk [Fok96]. TANE [HKPT99]. TANT [Mar69]. Tape [Edw77, BSD60, Cha62, DGM68, EAL66, Gil66, Goo64, Gri59, Liv59]. tapes [Gri59, cGR61]. Target [FL85, KBTB97, CMRW72, CP62]. Target-System [KBTB97]. Task [Sta74, VH99, BM80]. Tasks [PS90, Sme92, BM86]. Tattoo [MR85]. tax [God75]. taxonometric [WLDC71]. Taxonomic [Coo83]. taxonomy [Fur89, WHF82]. Taylor [BWZ71, Gib60, Wil76a]. TC8 [VS87]. teach [HS82]. Teaching [AD76, DB90, McG77, WHF83, Bow73, Cap86, Fox61, Fur92, GGH^{+}87, GRB93, Lee78a, Spu88]. team [Bri67]. tearing [Bow90]. Tears [MH77, Lin72a]. tech [Lic89]. Technical [PT90, KC88]. Technique [BB80a, Bay76, Dou59c, Fen91, Kra71, Kuh92, Mul81, Wil77c, CP87, CZW94, Che65, CP61, Cox70, Dav67, Dee85, Eva65a, Fin68, FB75, Fos64, GR75, Gil63, HM95a, How59, LK77, Mea73, MW68, New68, NI70, OB68a, Oli69, ON94, Osb71, RHS89, RB86, Sav69, Sav70, SC71, Sri91, Sza73, Ull77, Wet66, Woo69c]. Techniques [ABN61, BH63, Bay73, Col87, Coo75, DF87, ES81, GL94, HK88, IB96, JW82, MZ86, Mal77, MSNW99, MW97, NPCA81, PCSG63, SB70, Ull73, Wal76, Woo86, AL72, Bra74, Bro60, Cha93a, CO72, Hon71, How67b, Hur68, Hus61, Hys72, Jac71, Lar67, PH73, RV93, Rib95, Sam72a, SFD92, Str60, TK95, Wym73, ZH93, BS73b, Cha73, Kit74, PCSG63]. technology [Cap86, Cla88, Ham80b, Mor89, Sha94, Ano74, Pri64, RL73]. TELCOMP [Mul69]. tele [Hop61]. tele-processing [Hop61]. telecommunication [Geo63]. telecommunications [Ano73c, DTIR88]. teleprinter [GS63, Sam63]. Teleprocessing [BB80a]. television [Bro69a]. template [Coo89c, HW88, Hor93]. Temporal [Bar87, HJ99, Lom88, MK94, MK96, Wor94, Kni93, KM94]. Ten [oH59, The91b, Kil81]. Tennent [FSWL82]. Tensor [Har76]. term [DW91, EA70, Eva71, GB69, HKK91, Rya66]. Terminal [Ste77b, PKK87, Sac62]. termination [Er86b, Hun81]. terminology [GT65]. terms [Gat65, MB66]. Ternary [BVH74, GK73, VH72, Hur68, Hur79, SW78]. Terrain [Du96]. tessellation [Wat81]. tessellations [Bow81, GS78]. Tesseral [BM90c]. Test [Cha75, CL99b, DW88, Hie96, Nai97, RS88, WL87, Woo59, YT79, Ben72, Box66b, Dow69, Hun60, Inc87, LLR73, Myl66]. Test-Data [DW88]. Testable [Wey82]. testbed [Hen78, HH79a]. Testing [BTD99, CY96, Ham77, Hie97, IH86, JES98, Kha91, KC97, Ris99a, Sam76, WL87, Wey82, Wil70, Ara94, BM90a, BW69a, Day77, Hie98, KJ75, LLR73, LMS73, Tay61]. Tests [BZN98, BKK84, Gro68]. Text [BS86, BHK85, FH98, How97, Lea78, LJ98, Mac91, Pik81, TC92, Zve75, CCB95, Col70, Fen96, Gob75, KR89, Mar91a, Mar91b, TR93, Thi81, WBM^{+}94]. Textbook [Nee63]. texts [KC88]. textual [Mul87]. textural [She89]. Their [DM83, Hou75, OP73a, RP63, Sel88, ABN61, BF72, Elc65, Fis76, GH95, Jam89, MA89, NE67, RL73, Ros82, You63]. them [Cam58]. Themes [VS87]. Themselves [Fok96]. Theorem [Mel95, Mel66, Pow69, SN75, Goo68, Woo69a, Woo70a, dS77]. Theorem-Proving [SN75, Mel66]. Theoretic [CCS87, Fos64, HTW89, MHLS93]. Theoretical [BK92, Pag60, Bae66, HRS93, Pen66, Ral61].

Theories [Pra88, SN75, Tul85, Pad91]. Theory [CF81, Cha87, Coo92b, GV99a, Jel79, Kor94, Lan85, MK94, Mic81, Par75a, PCSG63, Pra84, Rze85, SS96b, SR76, Sta85, TCB71, WY71, BK87, BWM88, Bra76, BMRS90, CAC82, EL82, Gil59, HRS93, Hig76b, HN88, JW72, LW67a, LW67b, LL90, Lin72b, MKP94, SM94, Tho60, Woo69b, Woo70b, Ire93, YFM71]. theory-based [BWM88]. thesis [Lau75, PR98]. thetas [Cr081]. thinning [Deu69, Fra70, SHCP91]. Thomas [PCSG63]. Thread [Sin97]. Threaded [BF81]. Three [Bro74c, Wat77b, Che65, Erk84, EA70, Eva71, Had71, KK94, McC71b, MM71, Phi67b, Rya63, SM71, Sco69, Spä69a]. three-dimensional [Che65]. three-point [McC71b, SM71]. three-track [KK94]. Threshold [BR91, CW96a, Edw78, EH78, Par96b, Win75]. Threshold-based [CW96a]. Threshold-Logic [EH78]. throughs [Lem80]. Tight [KP98]. Time [BP94, Bur72, BKL74, CT94, CZL97, Coo60b, Coo75, CM93, Deo83, DT96, DD87, DD88b, ER86a, HY81, HOH91, Jak85, JP86, KU99, LH95, Lew63, Li98, LB92, Low95, Med70, Mux73, NC76, Sha72a, SMY95, Tak99, TDW^{+}99, Wei91, Wil77c, Coc97, Bae66, Bar72b, BML94, BD72, BC73, Bro69a, BKMK90, CWH93, CS95, Che95b, Col64a, Dei95, DPT94, ET75, GE69, Hex67, HM95a, Hop61, Hor93, Joh61, KK94, KM94, Kol78, KPT96, LPSC99, LWS69a, Lip80, LR72, Lut80, MKP94, Mil63, Pen66, Per80, Phi67a, RW68, SGR96, Sto72, Tho68a, TB94, Tom82, Tur79, WA91, Win59a, YY89, BM93]. Time-constrained [LH95]. Time-Domain [Coo75]. time-oriented [BML94]. time-shared [Bae66, LWS69a, Pen66]. Time-sharing [Coo60b]. time-tables [Col64a]. Times [Deo83, Har84b, JP86, Mal77, PAW86, SD78, SR76, Nud69, Wil78d]. timesharing [HG71]. timestamps [Mad98]. Timetable [SS80, Sel88, Akk73, Alm69, BF71, DD70]. timetables [Alm66, ABN61, Bar65c, FL68, Woo68, Yul68]. timetabling [Aus76, Col67, CK93, Law69, WP67, Woo69c]. Timing [CDCL97, Lor76, Che64]. Timothy [HV92]. Titan [LWS68]. titles [Bon71]. Titli [Den80]. Token [JCCJ99]. Tokyo [BMPS71]. Tokyo-Edinburgh [BMPS71]. Tolerance [JP76]. Tolerant [BENM98, MS91, MS92, AE90b, GP96b, GIB90]. Tompkins [Roh91]. Tompkins-Paige [Roh91]. Tool [BS87, BS86, Sch77b, AFWH88, BWM88, KK79, Lam63, PL94a, ZM95]. toolkit [HR96]. Tools [CD81, CG92, FKT90, KMK95, PB86, GGS89, GG59, HKPW89, IW86]. Toolset [IHH95, KMR94]. Top [ABD84, Lau91, Duk66, HW88, Jon66, Rac65, Woo71a]. Top-Down [ABD84, Lau91, HW88, Woo71a]. topographic [VV94]. Topological [BM90a, Er83c, KV84, VR81a]. Topology [Gee91, TK99, BP82, Las68, VV94]. Torquay [Ano73c]. Torso [vO78]. Total [SN75, MS97, PJ86]. Tower [Er82b]. Towers [Ged96, Wod78, MW68, RW67, Er84a, Er84b, Er84c, GC87, Hay77, Lu89, Roh87]. Trace [BBK87]. Traces [Foo86]. Tracing [SMR98, Vin59]. track [KK94, Sim69]. Tracking [Cur95, Ear80, GLR93]. Tracks [PW81a]. tracts [PR98]. trade [SGR96]. trade-offs [SGR96]. Traffic [Bei72, Bri79, TB94]. train [dSA92]. training [Gea59, GGH^{+}87]. trains [CMS80, KV63]. trajectory [Vin59]. Transaction [BJ98, KU99, MK82, OBLP75, TDW^{+}99, dHZ97, Che95c, Mad97, MB90, TCM94]. Transaction-Processing [MK82]. Transactions [LH95, ZG99, dHZ97, LC95]. transcendental [JBJ84]. Transcription [Bro74e]. Transfer [MS92, ZSP99, Bea68]. transferable [HWW87]. Transferral [BHW78]. Transfers [Huz79, MB96]. Transform [Ham91a, MH77, PTV97, AL72, Pie71, Pie82, Fen96]. Transformation [DY80, HK88, HK92, Nas75, War96, WB99, FCHK96, Fra61, Fra62, Ged96, Oul87, PR77, Scr72, War94, Wyn71]. transformation-directed [PR77]. Transformational [Gru96]. Transformations [BT81, BPTV92, KS97, Mid87, Box66a, Hen88, YY89]. Transforming [Oom93, LL70]. Transforms [Dur74, Bel63, Hur79, ZL73]. Transitions [Bro84a]. translate [Mul69]. Translating [WC86]. Translation [Ham62, OM88, BB59, Cle59, Dak75, O’N92, Ric58]. translator [DH60, Hoa62b, McE67, Pul64, Ree67, WP75]. translators [Goo75, LP73a]. Transmission [Jok90, Wol78, Ano63, Dac63, Dav63, Fen87, LT61, Lon61, MW68, RW67, Rus63, Smi63, Str61, Wil61a, Wil61c, Wri61a, NLW74]. transmitting [Wil63b]. Transparency [Bur88]. TRANSPLINE [BT81]. Transport [NH89, PW67]. Transportable [Moo79]. transportation [HV78, Lig73]. Transposing [Win59b]. Transposition [Bro74e, GN94]. Transpositions [AG98]. transputer [AO87, SP91, B– D96, Bow90, Kac91, KLB91, She92]. transputer-based [SP91]. Trap [Bra88]. travel [Dei95]. travelling [Obr68]. Traversal [Bri85, Tak99, van88, Hat84, Sou77]. Traversals [And90, Bur75b, Azm90, Er89a]. Treating [MS97]. Treatment [BW70a, TI75, DPSW89]. Tree [And90, Arg91, BHK85, CS77, Che95a, Dak65, Har84b, JJ86, Jen99, ME98, Nar96, Pad73, SW83, Sur86, Tak99, Wei93, Woo82, ZOG85, van88, AO80b, BK83, Cha92, Chr71b, Day76, FL80, Hat84, Mäk91b, Obr68, OV92, Roh73, TB86, Woo71c, Woo71b]. Tree-Based [Nar96]. Tree-Like [Har84b]. tree-search [Chr71b]. Trees [BD84, BF81, Bri85, BMR62, Bur75b, Cha91, DS91, DY80, Er85, Er87b, Er88b, EA83, FK93, Gol93, GOW83, KV95, LG86, Mah86, Mak91a, Mul81, Pal86, PP98, Sip88, Smy91, Spr92, WW76, Win60, Woo84, XTU97, ADD90, Atk93, Bae70, Bro95, BKMK90, Cha90, CM89, Er89a, Er89b, ES77, FP82, Lyn65, Obr67, OW80, Pal88, SGR96, Sou77, Sri91, SW78, Zak84]. trend [Hio67]. Trends [VS87]. Tri [Wil62b, Bak64, EF63b]. Tri-Diagonal [Wil62b, Bak64, EF63b]. Trials [Sha72a, BS61]. Triangulated [JW98, JKW94]. Triangulating [vO78]. Triangulation [LR78, Oxl85]. triangulations [Sib78]. Tricky [PS93]. Tridiagonal [OK63, AHM74, Eva72, FJ66, GZ72, HM77, Øst79, Par67]. TRIES [NS77]. Trigonometric [OP71, Hun68, OP73b, Wat59]. Triodic [Gol82]. Triple [Col82, Mar92]. Trojans [Coh01, Mäk01, TAC01, TAC98]. trolley [Abb70]. Trotter [VR81b]. truncation [FNK94, OP75, Scr64a, Smi66a]. TSAS [LWS69a]. Tube [Ste77b]. Tuning [Wie79]. tunnel [HB58a, HB58b, Hon74]. tuple [Dee85]. Tuples [BMSA88, YAB92]. turbines [Lan65]. Turchin [FE80]. Turing [BB94, CD77, Nau93]. Turning [Jar67a]. Tutorial [Lan82, PP98]. twenty [Chu68]. twenty-five [Chu68]. Two [Bro74c, Cha93a, CCC92, Cra72, DF87, EB89, ES81, Fry71, GH75, Ham91a, JW82, KV84, Lyo85, MM70, McL76, Pac88, PW81a, PB86, Pro74, Sol99, Str60, Suf90, Wat77b, Wil85b, Woo77a, Zmu75a, Zmu75b, AH69, Bic68, Bic69, Bro60, DN64, Dei95, FN93, Fyf69, Has61, Mak89, Nic66, Obr67, OW68, Pag82a, PRO80a, Phi67b, Pri68, RS69, RW61, Sab76, Seg87, SPMD83, Woo70c, HP67]. Two-Dimensional [Ham91a, McL76]. Two-Disc [CCC92]. two-headed [Dei95]. Two-Level [Wil85b, EB89, Bro60, DN64, Pag82a, RW61]. Two-Parameter [Cra72]. two-point [AH69, Bic68, Bic69, Fyf69, Has61, Pri68]. two-stage [FN93]. two-way [Seg87, Woo70c]. TWOBYTWO [HP79]. Type [ABM88, Cow86, Hig76a, Les83a, Sho81, Coo89b, Eng69, FM64, Hen88, Hex67, Hou83, LL90, LM68, Mid77, Mor89, PC64, Sac62, SRS94, WS86]. type-matching [Hex67]. type-safe [Coo89b]. typed [DB88]. Typeless [Hig76a]. Types [BL81b, Har84a, HK92, Mai85, OD91, YP86, BK87, Obr67, Ire93]. Typesetting [CW82, Gol69].

U.K. [Fur92, Wil67c]. U.S.S.R [BT63]. U.S.S.R. [Cla60]. UET [BMRS90]. UFL [YP86]. UK [The91a]. Ullman [JR81]. Ulrich [Cla92]. UMIST [GCS89]. UML [Pai99]. unbiased [EN82]. Unbounded [CT97]. uncertain [Qui83]. Uncertainties [Sou76]. uncertainty [AB84]. Unconditional [Wig81, Tho80]. unconstrained [MD71]. Uncoupling [Fos71]. Undefined [SB99, Cze88]. Underflow [Goo77, Goo80, Goo85a]. undergraduates [Fur92]. underground [Lan65]. Underlying [Rob88, GK96]. Understandability [Var76a]. Understanding [Eas76, HC78, Nau93, Cla88]. unequally [GM72, OP71, Twi75]. unfoldable [JB94]. Unification [Ver91, ME88b]. Unified [Wor94]. Uniform [Atk93, Mal77, JC97]. Unifying [Lo84, Wol90]. union [Atk80]. Unique [JC97, SD78, NF94]. Uniqueness [vtv92]. Uniquest [vtv92]. Unit [BK75, LKK71, JJ67]. unitary [Fra61]. United [GG59]. units [Gla69, Wil68b]. Unity [Ver91]. Universal [Zah81, Jor73, Nau93, NPS89, Bag62]. Universities [Fur92, Aga96, Dou59b, Sam68, Ski96, VW95, Web72, You63]. University [Ano75a, Bro58a, DD70, Dol66, GG86, HPS61, KHPS61, Muc82, Wil67b, Akk73, Buc60, Buc65, WKRJ60, Woo68, Alm66, Alm69, Col70, Yul68]. UNIX [Hek89, MBC^{+}86]. UNIX-based [MBC^{+}86]. unnoted [Sal72]. unpredictability [BRP93]. Unranking [Pal86, SGR96]. UNRAVEL [Bro73c]. Unravelling [Oul82]. Unstable [DD88b, Sco70]. Unstructured [CCD86, Oul82, WO78]. Unstructuredness [Tse87a, Tse87b, Wil77b]. unsuccessful [HM96]. until [DHW78]. unwanted [Par75d]. Up/Down [MG95]. Update [LPS94]. updates [CM89]. Updating [JW82, Bro79a, Cha90]. Upon [Cha87, YY89]. upper [WP67]. URPC [HR96]. Usability [Sha86, Tyl88]. Usage [EM81b, LE86, Bro73b, Leh73]. Use [AS63, BBS66, Bra85, Bra86a, Cha63, DK74, Gol82, Jar68, Lee78a, Mid76, OP73a, Sur86, Ver87, Wic92, AEH76, AL72, Bae70, BH68, BM92, BS61, BLY70, Box66a, Bra74, Cla77b, Clo65, Col66b, Col70, Cra66a, Dac63, Day70, DO72, Dou65, Er86b, Eva77, FT73, Fyf69, Fyf70, Gal59, Gea61, GB69, Gri59, Gri68, Gri80, HPLG79, Har88, Hil67, Hoc62, JN65, KK79, LM63a, Lan71, Liv59, McC71a, McC72, MHF93, OP75, OV92, WNT73, Wra66, YR58]. used [Gat65, Hay69]. Useful [AC84, Wil65a]. User [AM92, Ayr64, CGRJ92, Eas76, Fra68, Gol64, Har84a, NF76, NF80, NM90, PJ80, Ric64, SLH83, Bae64, BM93, CT74, Cor65, Fis74b, Gea59, Gea73, HN88, HPS61, JC86, Jam80, JBD73a, JBD73b, KM65, Lav64, O’S88, RB85, Smi64, Wak93]. User-Defined [Har84a]. User-Extensible [NF80]. User-Friendly [SLH83]. Users [BGS99, Bac71, ER71, Hil72, Wal71, Web72]. uses [Mus66b]. Using [Aki77, BT81, BGR86, Bro69b, BHK85, Cha91, CTH99, Cla78, Col87, Coo75, CH87, CT92, DE86, Fro85, GW73, IP86, JCCJ99, JL80, Leu83a, LW85b, LG86, Lom88, MZ86, Mil81a, NMW97, NC76, PB76, PB78, Pik81, Ram99, Ree80b, SD78, Sar96, Sch82b, Sho81, Ski96, SW83, Thi86, TH98, UVS87, Wat67, Wel80, Wel72, WK77, ZBS81, vFV92, van88, Ara94, Art75, BSD60, BHS65, Bec58, BS71, Bow73, Bri79, CC93, Chr73a, Chr71b, CRS^{+}93, Cla94, CMW98, Col64a, Ede65, Er89b, EM71, EH76, FK68, GS63, Gib60, Gob75, GRB68, Gri73, GB61, HV74, HW74, Hon71, HW89, Hop63, HKA72, HK72, Hos74, How97, HSW92, JM96, JS67, Jon70, JES98, KC88, KL71, Kro70c, Kro70d, Kro70e, Lil70, Lil71, Mad98, Mar69, Mar74a, McC93, ME88a]. using [Mul87, Oli71, Oom93, ON94, Pag82a, Pag71, Pal62, Par74, PC78, PG90, PRO80a, Red60, Rv88, Roh73, Sam63, Sam72a, Sch65, Sea69, Tho64, Van73, VV94, WC94, Wil59, Wil89, Woo70c, Yok97]. USSOR [Lyn64]. Utilisation [LY79, OT91, Gla69]. Utility [Wil64a]. utilizing [All82].

V [Eme80, FE80, Hou63]. Validating [Fos86, Jac79]. Validation [FS81, Lis74, LCP81, Pad73, O’S94, Smi72, VV94]. Value [BW70a, Nas75, NS87, Sco70, AH69, Bon69, Car64, Cas75, Del68, Eva77, Fyf69, Fyf70, Hop71, Par96a, Pri68, Ros60, Sch76, Tay70a]. Valued [Ham91a, Wu96, CWH93, Che90, WV70]. Values [BR91, Sho81, Pow62, Tee67]. Vandermonde [TS75]. Variability [Rin75]. Variable [Bir79, Cha93b, GW73, Imb93, PB76, PB78, Pea69, Sam76, Wel72, All59, Art75, CCL72, Fle70, HW67, Jen71, Lle85, MT83, RM89, SP82, Tay76, Tho64, Tri74, War74]. variable-length [CCL72, HW67, Lle85, Tay76]. Variables [Pow64, Sou76, Suf90, Cze88, Had71, Hig76b, Pow62]. Variance [Dav68, Ash68, BB69, PFW74]. Variant [Thi86, Weg93, WBB72]. Variants [Bun97]. variate [Col64b]. variation [DS67]. Variations [Woo77a, All66a, Pak67]. variety [Hop63, Pai65]. Various [Fro85, Mit76]. varying [KCP99]. VDM [Gib88, O’N92]. VDU [Bro79a]. Vector [Ben85, Kei79, McL78, MF82, Aha65, Boo67k, Boo67i, Boo67j, Par89, Pri68]. Vectors [Pow68, And69, DP64, Pal69]. Veen [Kil81]. Vegas [EL89]. Verdicts [Wat77b]. Verification [HH94b, Kha91, LCP81, tHO99, Cho95, Gra87, HKT95, HM95b, HCH96b, LB76, O’S94, TK95]. verified [HM95b]. Version [Lar85]. Versions [Roh87]. Versus [BD84, HH76, Wri82, DMR88, MZ96, Nie65, PRLW93]. Vertex [Sel88]. vertical [CG99]. Vertices [Sel88]. Very [Tre82]. VI [HS69, HL70, Hea68, Wel70]. Via [Eke95, FR98, BHW78, Els79, FG90, Hon93a, Li98, Ull88]. Viable [Ear75]. Vibration [Wil67b]. Vibrations [Gre70]. Vickery [Kit74]. Video [THHS93, Wak93]. View [BJ98, EMO92, Ano76, Lav64, Mar73, Mor89, Shr75, Smi64]. Viewdata [MM84]. viewing [Ort71]. Viewpoint [Mal77]. views [Dan92, Smi63]. Virtual [BBF^{+}94, Bro79b, Gom81, JCCJ99, NMM91, RKA92, Ver87, Cro80, Las68]. Virus [Coh01, Mäk01, TAC98, TAC01]. viscous [Abb64]. Visibility [FA88]. Vision [Arm78]. Visputer [ZM95]. visualization [ZM95]. Visualizations [IB96, KBTB97]. VLSI [Cla92, AM90, ACDP91, CF89, DS90, HAAZ99, Lea86, Sta94, Tre82, Wil91a]. VM [Don78]. VM/ [Don78]. vocabulary [Gat65, Nau79]. Vol [WY71]. Volterra [Lew72, AW71, JS67, Lin69, WP68]. Volume [Bur80, RP63, Wil63a]. volumes [Oxl89, Rit89]. Voronoı [Wat81]. vote [HWW87]. votes [HW74, Par74, Van73]. Voting [Par96b]. Vs [MKH86, Wym73]. VULCAN [MM84].

W [BT63, Bro74b, CF81, FE80, GT63, Her61, JR81, Kil81, Llo74, NPCA81, PCSG63, WY71]. W. [FSWL82]. Waiting [Arg91, Tur79]. walk [Bra94, Lem80]. walk-throughs [Lem80]. Wallace [Ris99b, Cla99, Daw99]. Walsh [Hur73]. Wang [PCSG63]. want [Bae64]. Warp [BM93]. Wasserman [VP80]. Watson [MR76a]. Watt [Gen69]. wave [Bea68]. waveguides [MCL89]. Way [AD95, Bar87, Ken85, Seg87, Woo70c, Smy91]. Ways [Nob91]. Weakest [CW98]. Web [BGS99, Thi86]. Weber [VP80, LL70]. weekend [Ano76]. weekly [Sha58]. Wegner [HV92, PR98]. Weighted [Arg91, Ric88, Sch85, BMT95, PD79, Wil86, ZL73]. Weighting [Arg91]. Welcome [Gea80, van93]. Well [CCD86, Er83b]. Well-Formed [CCD86, Er83b]. Westreich [Bre81]. Wheeler [Fen96]. Where [Cut59]. Which [FR98, O’B70, Fin60, OM64, PJ89, Par75d, Sal74, SAHS71]. while [BK87, Tak86]. while-rule [BK87]. White [Man71b]. Wide [KA98, vSHBT98]. wide-area [vSHBT98]. widely [Tay70a]. Wiley [Wil70]. Wilkes [Ano81b, HT81]. Wilkinson [Fox87]. William [Eme80, MM69, Web70]. Williams [BS73b, Cor79, Sib71]. Willoughby [RL73]. wind [HB58a, HB58b]. Window [BG94, Hek89, Mic82, PKK87]. Windscale [WM69]. Winston [Bur80]. wire [Abb70]. Wireless [FLW99, ZSP99]. Wirth [LL70]. Wirth-Weber [LL70]. Within [AD76, BHW78, Cha81, Ada76, Gri68]. Without [Arg91, Pow65, Fle65, Jak85, Pec70, Pow64, Wie79, Zan67]. Witzgall [YFM71]. Wombat [Cla88]. wonderland [Fel72]. Wooloms [IL74]. word [Bur69a, FT73, Mul75, Thi81, WNT73]. words [Coo68, Ull77]. Work [BBF^{+}94, The91a, Che62b, Cut59, GT65, HL76b, Lep95, Tho58a]. work-optimal [Lep95]. Working [Ham69d]. Workloads [DL97]. Workshop [FJA80, Tre82]. Workstation [SKW85, WSG^{+}84]. Workstations [CZL97, BM93, SWZ93]. World [Oka98, HS69]. Worst [Joh74, Lar82, Ric88, Rob77, Erk84].

Worst-Case [Joh74, Lar82].

Worst-Possible [Ric88]. writer [BN76]. Writers [KRR82]. Writing [Akm86, BW78, Bux66, HW89, Paa91, Pag81, All71, For65, HV74, Jon71a]. Written [BP81]. WSI [Lea86]. WWW [HJ99].

x [Loi68, Loo59]. X-1 [Loo59]. X.400 [MRW90]. X.400-1988 [MRW90]. XENO [Smi83]. XOR [LT79]. Xunrang [Weg93]. XVII [TCB71].

Yarmish [WD80]. year

[Bar58, Hug61, Sut60, War60, YR58, Tho98]. Years [The91b, oH59, Rya63, SPMD83, Tho58a]. Yen [Pau80]. Younger [Gol73]. yourself [Nef65]. Yovits [PCSG63]. Yuzhang [Weg93].

Zero [Bre71, BDB92, Chr71b, BP69, Cox70, Ros63b, SC71]. Zero-Knowledge [BDB92]. Zero-one [Chr71b, BP69]. zeros [Che65, Cr081, Lar64]. Zipf [Pra88]. zones [HTW89]. Zubov [Hou63].

References

Aleksander:1968:ALC

[AA68] I. Aleksander and R. C. Albrow. Adaptive logic circuits. The Computer Journal, 11(1):65–71, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                                    //www3.oup.co.uk/computer_
                                                    journal/hdb/Volume_11/Issue_
                                                    01/110065.sgm.abs.html;
                                                    http://www3.oup.co.uk/computer_
                                                    journal/hdb/Volume_11/Issue_
                                                    01/tiff/65.tif;   http://
                                                    www3.oup.co.uk/computer_journal/

hdb/Volume_11/Issue_01/tiff/
66.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_11/Issue_01/tiff/67.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/68.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/69.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_01/tiff/
70.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_11/Issue_01/tiff/71.
tif.

Agius:1999:CCO

[AA99] Harry W. Agius and Marios C. Angelides. COSMOScontent oriented semantic modelling overlay scheme. The Computer Journal, 42(3):153–176, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
03/420153.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
03/pdf/420153.pdf.

Ahituv:1978:APO

[AB78] N. Ahituv and I. Borovits.

Analysis of the purchase option of computers. The Computer Journal, 21(2):105–109, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/105.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        02/tiff/106.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        02/tiff/107.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        02/tiff/108.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        02/tiff/109.tif.

Agmon:1984:DUC

[AB84] T. Agmon and I. Borovits. Dynamic and uncertainty considerations in the selection and the acquisition of an information system: an adoption of a break-even analysis to investment decisions. The Computer Journal, 27(1):87–89, February 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        01/tiff/87.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_27/Issue_01/tiff/
                                        88.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_27/Issue_01/tiff/89.
                                        tif.

Alcock:1990:PPR

[AB90] D. Alcock and D. Brown. PRAXIS: A program for reproducing proforma design calculations. The Computer Journal, 33(4):337–343, August 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/337.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/338.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/339.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/340.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/341.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/342.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/343.tif.

Abbott:1964:ASS

[Abb64] M. R. Abbott. Axially symmetric steady motion of a viscous incompressible fluid: some numerical experiments. The Computer Journal, 7(1):47–53, April 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
01/070047.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
01/tiff/47.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_01/tiff/
48.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_07/Issue_01/tiff/49.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_

                                        07/Issue_01/tiff/50.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        01/tiff/51.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_07/Issue_01/tiff/
                                        52.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_07/Issue_01/tiff/53.
                                        tif.

Abbott:1970:NMC

[Abb70]  M. R. Abbott.  Numeri-

cal method for calculating the dynamic behaviour of a trolley wire overhead contact system for electric railways. The Computer Journal, 13(4):363–368, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/130363.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/363.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/364.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/365.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/366.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/367.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/368.tif.

Abbott:1980:ACS

[Abb80]  J. P. Abbott.  Algorithm

110: Computing solution arcs of nonlinear equations with a parameter. The Computer Journal, 23(1):85–89, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
01/tiff/85.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/
86.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_23/Issue_01/tiff/87.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
23/Issue_01/tiff/88.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
01/tiff/89.tif.

Anderson:1983:ALL

[ABBS83] S. O. Anderson, R. C. Backhouse, E. H. Bugge, and C. P. Stirling. An assessment of locally least-cost error recovery. The Computer Journal, 26(1):15–24, February 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
01/tiff/15.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_01/tiff/
16.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_26/Issue_01/tiff/17.
tif;   http://www3.oup.co.

uk/computer_journal/hdb/Volume_
26/Issue_01/tiff/18.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
01/tiff/19.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_01/tiff/
20.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_26/Issue_01/tiff/21.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
26/Issue_01/tiff/22.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
01/tiff/23.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_01/tiff/
24.tif.

Atkinson:1983:APP

[ABC^{+}83] M. P. Atkinson, P. J. Bailey, K. J. Chisholm, P. W. Cockshott, and R. Morrison. An approach to persistent programming. The Computer Journal, 26(4):360–365, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/360.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/361.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/362.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/363.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_

04/tiff/364.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/365.tif. Also pub-

lished in/as: In “Readings in Object-Oriented Database Systems” edited by S. Zdonik and D. Maier, Morgan Kaufman, 1990.

Adorni:1984:TSA

[ABD84] G. Adorni, A. Boccalatte, and M. Di Manzo. Top-down semantic analysis. The Computer Journal, 27(3):233–237, August 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/233.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/234.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/235.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/236.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/237.tif.

Abel:1985:SEO

[Abe85] D. J. Abel. Some elemental operations on linear quadtrees for geographic information systems. The Computer Journal, 28(1):73–77, February 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_

journal/hdb/Volume_28/Issue_
01/tiff/73.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_01/tiff/
74.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_28/Issue_01/tiff/75.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
28/Issue_01/tiff/76.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
01/tiff/77.tif.

Atkinson:1988:BTC

[ABM88] M. Atkinson, P. Buneman, and R. Morrison. Binding and type checking in database programming languages. The Computer Journal, 31(2):99–109, April 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/100.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/101.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/102.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/103.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/104.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/105.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/106.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/107.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/108.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/109.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/99.tif.

Appleby:1961:TPS

[ABN61] J. S. Appleby, D. V. Blake, and E. A. Newman. Techniques for producing school timetables on a computer and their application to other scheduling problems. The Computer Journal, 3(4):237–245, January 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/030237.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/237.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/238.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/239.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/240.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/241.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_

                                        04/tiff/242.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        04/tiff/243.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        04/tiff/244.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        04/tiff/245.tif.

Arora:1984:NCB

[AC84] Adarsh K. Arora and C. R. Carlson. Normalization could be useful. The Computer Journal, 27(1):57–61, February 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        01/tiff/57.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_27/Issue_01/tiff/
                                        58.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_27/Issue_01/tiff/59.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        27/Issue_01/tiff/60.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        01/tiff/61.tif.

Ancona:1991:HLH

[ACDP91] M. Ancona, A. Clematis, L. De Floriani, and E. Puppo. HIDEL: A language for hierarchical VLSI design. The Computer Journal, 34(3):195–206, June 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_

journal/hdb/Volume_34/Issue_
03/tiff/195.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/196.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/197.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/198.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/199.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/200.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/201.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/202.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/203.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/204.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/205.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/206.tif.

Atkinson:1976:MCG

[AD76] L. V. Atkinson and G. J. Duffus. Machine code generation within a compiler — a teaching model. The Computer Journal, 19(2):160–165, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/160.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/161.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/162.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/163.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/164.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/165.tif.

Ahmad:1995:WGP

[AD95] Imtiaz Ahmad and Muhammad K. Dhodhi. On the mway graph partitioning problem. The Computer Journal, 38(3):237–244, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_38/Issue_03/
                                        Vol38_03.body.html#AbstractAhmad.

Adams:1969:AAU

[Ada69]  A. G. Adams.  Algorithm

39: Areas under the normal curve. The Computer Journal, 12(2):197–198, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        02/tiff/197.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/198.tif.

Adamczewski:1976:FER

[Ada76] B. F. M. Adamczewski. A formula for establishing re-order quantities for individual items within a joint stock replenishment scheme. The Computer Journal, 19(3):194–197, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/194.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/195.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/196.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/197.tif.

Ancona:1986:PPS

[ADD86] M. Ancona, L. De Floriani, and J. S. Deogun. Path problems in structured graphs. The Computer Journal, 29 (6):553–563, December 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/553.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/554.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_

06/tiff/555.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/556.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/557.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/558.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/559.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/560.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/561.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/562.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/563.tif.

Ancona:1990:SST

[ADD90] M. Ancona, L. De Floriani, and J. S. Deogun. Structured spanning trees. The Computer Journal, 33(4):344–355, August 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/344.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/345.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/346.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_33/Issue_
04/tiff/347.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/348.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/349.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/350.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/351.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/352.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/353.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/354.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/355.tif.

Abolhassan:1993:PDP

[ADK^{+}93] F. Abolhassan, R. Drefenstedt, J. Keller, W. J. Paul, and D. Scheerer. On the physical design of PRAMs. The Computer Journal, 36(8):756–762, ???? 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_08/
Vol36_08.body.html#AbstractAbolhassan.

Abramson:1990:RDF

[AE90a] D. Abramson and G. Egan. The RMIT data flow computer: A hybrid architecture. The Computer Journal, 33(3):230–240, June 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/230.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/231.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/232.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/233.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/234.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/235.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/236.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/237.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/238.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/239.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/240.tif.

Agrawal:1990:ISF

[AE90b] D. Agrawal and A. El Abbadi. Integrating security with fault-tolerant distributed databases. The Computer Journal, 33(1):71–78, February 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
01/tiff/71.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_01/tiff/
72.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_33/Issue_01/tiff/73.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/74.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
01/tiff/75.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_01/tiff/
76.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_33/Issue_01/tiff/77.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/78.tif.

Alaylioglu:1975:AGQ

[AEH75] Ayse Alaylioglu, G. A. Evans, and J. Hyslop. Automatic generation of quadrature formulae for oscillatory integrals. The Computer Journal, 18(2):173–176, May 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/173.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/174.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_18/Issue_
                                        02/tiff/175.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        02/tiff/176.tif.

Alaylioglu:1976:UCS

[AEH76] A. Alaylioglu, G. A. Evans, and J. Hyslop. The use of Chebyshev series for the evaluation of oscillatory integrals. The Computer Journal, 19(3):258–267, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/258.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/259.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/260.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/261.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/262.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/263.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/264.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/265.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/266.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_19/Issue_
03/tiff/267.tif.

Alewine:1976:SBI

[AF76] T. Alewine and R. A. Fleck, Jr. Service bureau or in-house data processing. The Computer Journal, 19(3):198–201, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/198.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/199.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/200.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/201.tif.

Anedda:1988:PQI

[AF88] C. Anedda and L. Felician. Pcompressed quadtress for image storing. The Computer Journal, 31(4):353–357, August 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/353.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/354.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/355.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/356.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        04/tiff/357.tif.

Avison:1988:ISD

[AFWH88] D. E. Avison, G. Fitzgerald, and A. T. Wood-Harper. Information systems development: a tool kit is not enough. The Computer Journal, 31(4):379–380, August 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        04/tiff/379.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        04/tiff/380.tif.

Abdel-Ghaffar:1998:DST

[AG98] Khaled A. A. Abdel-Ghaffar. Detecting substitutions and transpositions of characters. The Computer Journal, 41 (4):270–277, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_41/Issue_04/
                                        Vol41_04.body.html#AbstractAbdel-
                                        Ghaffar.

Agar:1996:HCF

[Aga96] Jon Agar. History of computing feature: The provision of digital computers to British universities up to the Flowers Report (1966). The Computer Journal, 39(7):630–642, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www.oup.co.uk/jnls/
list/comjnl/hdb/Volume_39/
Issue_07/390630.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_07/
Vol39_07.body.html#AbstractAgar.

Agerholm:1995:LEH

[Age95] Sten Agerholm. LCF examples in HOL. The Computer Journal, 38(2):121–130, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_38/Issue_02/
Vol38_02.body.html#AbstractAgerholm.

Alagar:1991:IMS

[AGNS91] V. S. Alagar, P. Goyal, P. S.

Nair, and F. Sadri. Integrated magic set method: a rule rewrite scheme for optimising linear datalog programs. The Computer Journal, 34(6):A121– A129, December 1991. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Albasiny:1969:CSS

[AH69] E. L. Albasiny and W. D.

Hoskins. Cubic spline solutions to two-point boundary value problems. The Computer Journal, 12(2):151–153, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/120151.sgm.abs.html;
http://www3.oup.co.uk/computer_

journal/hdb/Volume_12/Issue_
02/tiff/151.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/152.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/153.tif.

Ahamed:1965:ACN

[Aha65] S. V. Ahamed. Accelerated convergence of numerical solution of linear and non-linear vector field problems. The Computer Journal, 8(1):73–76, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
01/080073.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
01/tiff/73.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_01/tiff/
74.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_08/Issue_01/tiff/75.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
08/Issue_01/tiff/76.tif.

Auramaki:1992:MOTb

[AHL92a] E. Auramäki, R. Hirschheim, and K. Lyytinen. Modelling offices through discourse analysis: A comparison and evaluation of SAMPO with OS-SAD and ICN. The Computer Journal, 35(5):492–500, October 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_35/Issue_05/
Vol35_05.body.html#AbstractAuramaki.

Auramaki:1992:MOTa

[AHL92b] E. Auramäki, R. Hjrschheimi, and K. Lyytinen. Modelling offices through discourse analysis: the SAMPO approach. The Computer Journal, 35(4):342–352, August 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_35/Issue_04/
Vol35_04.body.html#AbstractAuramakil.

Ahmad:1972:ISH

[Ahm72] M. Ahmad. Iterative schemes for high speed division. The Computer Journal, 15(4):333– 336, November 1972. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
15/Issue_04/150333.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
15/Issue_04/tiff/333.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/334.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/335.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/336.tif.

Ahmad:1973:CSN

[Ahm73]  M. Ahmad.  The comput-

ing speed of a new machine. The Computer Journal, 16(2):161–164, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/160161.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/161.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/162.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/163.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/164.tif.

Andres:1974:ACA

[AHM74] T. Andres, W. D. Hoskins, and G. E. McMaster. Algorithm 84: A coupled algorithm for the solution of certain tridiagonal systems of linear equations. The Computer Journal, 17(4):378–379, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/378.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/379.tif.

Allran:1970:ASN

[AJ70] R. R. Allran and S. E. J. Johnsen. An algorithm for solving nonlinear programming problems subject to nonlinear inequality constraints. The Computer Journal, 13(2):171–177, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/130171.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/171.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/172.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/173.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/174.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/175.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/176.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/177.tif.

Augustsson:1989:CLC

[AJ89] L. Augustsson and T. Johnsson. The Chalmers Lazy-ML compiler. The Computer Journal, 32(2):127–141, April 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/127.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_32/Issue_
                                        02/tiff/128.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        02/tiff/129.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        02/tiff/130.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        02/tiff/131.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        02/tiff/132.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        02/tiff/133.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        02/tiff/134.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        02/tiff/135.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        02/tiff/136.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        02/tiff/137.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        02/tiff/138.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        02/tiff/139.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        02/tiff/140.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        02/tiff/141.tif.

Amble:1974:OHT

[AK74] O. Amble and D. E. Knuth. Ordered hash tables. The Computer Journal, 17(2):135–142, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/135.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/136.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/137.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/138.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/139.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/140.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/141.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/142.tif.

Alagic:1981:RPD

[AK81] S. Alagic and A. Kulenovic. Relational Pascal data base interface. The Computer Journal, 24(2):112–117, May 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/112.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/113.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        02/tiff/114.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        02/tiff/115.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        02/tiff/116.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        02/tiff/117.tif.

Akeroyd:1966:ASP

[Ake66] A. J. Akeroyd. An application of separable programming. The Computer Journal, 8(4):344–346, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        04/080344.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        04/tiff/344.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        04/tiff/345.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        04/tiff/346.tif.

Akitt:1977:FMU

[Aki77] J. W. Akitt. Function minimisation using the Nelder and Mead simplex method with limited arithmetic precision: the self regenerative simplex. The Computer Journal, 20(1):84–85, February 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
01/tiff/84.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_01/tiff/
85.tif.

Akinyokun:1988:ESC

[Aki88] O. C. Akinyokun. The experimental study of CODA-SYL database/administration at the secondary storage level. The Computer Journal, 31 (4):344–352, August 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/344.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/345.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/346.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/347.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/348.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/349.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/350.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/351.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_31/Issue_
                                        04/tiff/352.tif.

Akkoyunlu:1973:LAC

[Akk73]  E. A. Akkoyunlu.  A lin-

ear algorithm for computing the optimum university timetable. The Computer Journal, 16(4):347–350, November 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        04/160347.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        04/tiff/347.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        04/tiff/348.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        04/tiff/349.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        04/tiff/350.tif.

Akl:1987:AOP

[Akl87] S. G. Akl. Adaptive and optimal parallel algorithms for enumerating permutations and combinations. The Computer Journal, 30(5):433–436, October 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        05/tiff/433.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        05/tiff/434.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_30/Issue_
05/tiff/435.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/436.tif.

Akman:1986:WSC

[Akm86]  V. Akman.  Writing self-

replicating code. The Computer Journal, 29(6):573–574, December 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/573.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/574.tif.

Akman:1991:SNI

[Akm91]  V. Akman.  Short notes:

Implementation of Karp-Luby Monte Carlo method: an exercise in approximate counting. The Computer Journal, 34(3):279–282, June 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/279.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/280.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/281.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/282.tif.

Akyildiz:1992:PAM

[Aky92] I. F. Akyildiz. Performance analysis of a multiprocessor system model with process communication. The Computer Journal, 35(1):52–61, February 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_35/Issue_01/
                                        Vol35_01.body.html#AbstractAkyildiz.

Atkinson:1972:CSI

[AL72] M. P. Atkinson and S. R. Lang. A comparison of some inverse Laplace transform techniques for use in circuit design. The Computer Journal, 15(2):138–139, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        02/150138.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        02/tiff/138.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        02/tiff/139.tif.

Albaseny:1971:BRN

[Alb71] E. L. Albaseny. Book reviews: Numerik Symmetrischer Matrizen, by H. Rutishauser, E. Stiefel and H. Schwarz, 1968; Computational Methods in Partial Differential Equations, by A. R. Mitchell, 1969. The Computer Journal, 14(4):374, November 1971.

CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/374.tif.

Allen:1959:MRE

[All59]  C. D. Allen.  A method

for the reduction of empirical multi-variable functions. The Computer Journal, 1 (4):196–200, January 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/010196.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/tiff/196.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/tiff/197.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/tiff/198.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/tiff/199.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/tiff/200.tif.

Allen:1966:IDN

[All66a]  B. T. Allen.  An investi-

gation into direct numerical methods for solving some calculus of variations problems. Part 1 — Second order methods. The Computer Journal, 9(2):205–210, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/090205.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/205.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/206.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/207.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/208.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/209.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/210.tif. See corre-

spondence [Pak67].

Allen:1966:NMS

[All66b] B. T. Allen. A new method of solving second-order differential equations when the first derivative is present. The Computer Journal, 8(4):392–394, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/080392.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/392.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/393.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/394.tif.

Allen:1971:CQW

[All71] R. L. Allen. Correspondence:

On quality of writing. The Computer Journal, 14(1):86, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/140086.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/tiff/86.tif. See [Jon71a].

Allen:1975:IC

[All75] K. J. Allen. Insuring computers. The Computer Journal, 18(3):194–197, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/194.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/195.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/196.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/197.tif.

Allen:1982:FID

[All82] F. W. Allen. A file index for document storage and retrieval utilizing descriptor fragments. The Computer Journal, 25(1):2–6, February 1982.

CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/2.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/
3.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_25/Issue_01/tiff/4.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/5.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/6.tif.

Allison:1983:PDS

[All83] Lloyd Allison. Programming denotational semantics. The Computer Journal, 26(2):164–174, May 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/164.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/165.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/166.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/167.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/168.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/169.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_26/Issue_
02/tiff/170.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/171.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/172.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/173.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/174.tif.

Allison:1985:PDS

[All85]  Lloyd Allison.  Program-

ming denotational semantics. II. The Computer Journal, 28(5):480–486, November 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/480.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/481.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/482.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/483.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/484.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/485.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/486.tif.

Allison:1988:SAC

[All88] L. Allison. Some applications of continuations. The Computer Journal, 31(1):9–11, February 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        01/tiff/10.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_31/Issue_01/tiff/
                                        11.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_31/Issue_01/tiff/9.
                                        tif.

Allison:1990:CIG

[All90] L. Allison. Continuations implement generators and streams. The Computer Journal, 33 (5):460–465, October 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        05/tiff/460.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        05/tiff/461.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        05/tiff/462.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        05/tiff/463.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        05/tiff/464.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        05/tiff/465.tif.

Almond:1966:ACU

[Alm66] M. Almond. An algorithm for constructing University timetables. The Computer Journal, 8(4):331–340, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/080331.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/331.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/332.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/333.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/334.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/335.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/336.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/337.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/338.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/339.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/340.tif.

Almond:1969:UFT

[Alm69] M. Almond. A University faculty timetable. The Computer Journal, 12(3):215–217, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        03/120215.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        03/tiff/215.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        03/tiff/216.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        03/tiff/217.tif. See corre-

spondence [DD70].

Alpiar:1971:DSO

[Alp71]  R. Alpiar.  Double syntax

oriented processing. The Computer Journal, 14(1):25– 37, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        14/Issue_01/140025.sgm.abs.
                                        html;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        14/Issue_01/tiff/25.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        01/tiff/26.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_14/Issue_01/tiff/
                                        27.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_14/Issue_01/tiff/28.
                                        tif;   http://www3.oup.co.

uk/computer_journal/hdb/Volume_
14/Issue_01/tiff/29.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/tiff/30.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_01/tiff/
31.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_14/Issue_01/tiff/32.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
14/Issue_01/tiff/33.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/tiff/34.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_01/tiff/
35.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_14/Issue_01/tiff/36.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
14/Issue_01/tiff/37.tif.

Alway:1971:GFA

[Alw71] G. G. Alway. A general factorising algorithm. The Computer Journal, 14(2):166–168, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/140166.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/166.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/167.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/168.tif.

Alway:1965:ARB

[AM65] G. G. Alway and D. W. Martin. An algorithm for reducing the bandwidth of a matrix of symmetrical configuration. The Computer Journal, 8(3):264–272, October 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        03/080264.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        03/tiff/264.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        03/tiff/265.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        03/tiff/266.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        03/tiff/267.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        03/tiff/268.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        03/tiff/269.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        03/tiff/270.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        03/tiff/271.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        03/tiff/272.tif.

Alia:1990:VBC

[AM90] G. Alia and E. Martinelli.

VLSI binary-residue converters for pipelined processing. The Computer Journal, 33 (5):473–475, October 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/473.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/474.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/475.tif.

Agrawal:1991:EDI

[AM91] D. Agrawal and A. Malpani.

Efficient dissemination of information in computer networks. The Computer Journal, 34(6):534–541, December 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/534.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/535.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/536.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/537.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/538.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/539.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        06/tiff/540.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        06/tiff/541.tif.

Agosti:1992:UNI

[AM92] Maristella Agosti and Pier Giorgio Marchetti. User navigation in the IRS conceptual structure through a semantic association function. The Computer Journal, 35(3):194–199, June 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_35/Issue_03/
                                        Vol35_03.body.html#AbstractAgosti.

Arnavut:1997:LPS

[AM97] Z. Arnavut and S. Magliveras. Lexical permutation sorting algorithm. The Computer Journal, 40(5):292–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www.oup.co.uk/computer_
                                        journal/Volume_40/Issue_05/
                                        Vol40_05.body.html#AbstractArnavut;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_40/Issue_05/
                                        Vol40_05.body.html#AbstractArnavut.

Alia:1998:LDF

[AM98] Giuseppe Alia and Enrico Martinelli. Logic design of a fast circuit for iterative additions in redundant hybrid number systems. The Computer Journal, 41(1):45–51, ????

1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_41/Issue_01/
Vol41_01.body.html#AbstractAlia.

Al-Mouhamed:1999:EBS

[AM99] Mayez Al-Mouhamed. Evolutionbased scheduling of computations and communications on distributed-memory multicomputers. The Computer Journal, 42(5):373–390, 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
05/420373.sgm.abs.html.

ApSimon:1966:CIP

[AMKM66] H. G. ApSimon, W. D. Maurer, P. J. H. King, and C. H. R. Morris. Correspondence: An impossible program. The Computer Journal, 8(4):329–330, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/329.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/330.tif.

Atkinson:1981:CSE

[AMN81] L. V. Atkinson, J. J. Mc-Gregor, and S. D. North. Context sensitive editing as an approach to incremental compilation. The Computer Journal, 24(3):222–229, August 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/222.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/223.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/224.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/225.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/226.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/227.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/228.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/229.tif.

Anderson:1966:NCG

[And66] A. J. B. Anderson. A note on the construction of a general survey program in Extended Mercury Autocode. The Computer Journal, 8 (4):312–314, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/080312.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_

04/tiff/312.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/313.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/314.tif.

Andrew:1969:CCO

[And69] A. M. Andrew. Correspondence:

The calculation of orthogonal vectors. The Computer Journal, 12(4):411, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/411.tif. See [Pow68].

Andrew:1974:CDN

[And74] A. M. Andrew. Correspondence: Decimal number checking schemes. The Computer Journal, 17(4):382–383, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/382.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/383.tif. See cor-

respondence [And75, How75, ES75].

Andrew:1975:CDE

[And75] A. M. Andrew. Correspondence: Decimal error-correction — a solution. The Computer Journal, 18(3):284–285, August 1975. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/284.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/285.tif. See [And74,

ES75, How75].

Andersson:1990:NEB

[And90] A. Andersson. A note on the expected behaviour of binary tree traversals. The Computer Journal, 33(5):471–472, October 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/471.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/472.tif.

Anon:1960:NCL

[Ano60] Anon. Note on commissioning of LEO automatic office at the Ministry of Pensions and National Insurance. The Computer Journal, 2(4):198, January 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/198.tif.

Anonymous:1963:DFS

[Ano63] Anonymous. Discussion forum — the systems approach to data transmission. The Computer Journal, 6(3):228–231, October 1963. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/228.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/229.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/230.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/231.tif.

Anonymous:1973:CIB

[Ano73a]  Anonymous.  COBOL in-

formation bulletin number

16. The Computer Journal, 16(1):81–92, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
01/tiff/81.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_01/tiff/
82.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_16/Issue_01/tiff/83.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
16/Issue_01/tiff/84.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
01/tiff/85.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_01/tiff/
86.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_16/Issue_01/tiff/87.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_

                                        16/Issue_01/tiff/88.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        01/tiff/89.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_16/Issue_01/tiff/
                                        90.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_16/Issue_01/tiff/91.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        16/Issue_01/tiff/92.tif.

Anonymous:1973:EMM

[Ano73b]  Anonymous.  Eric Mutch

memorial prize. The Computer Journal, 16(3):244, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        03/tiff/244.tif.

Anonymous:1973:ORS

[Ano73c] Anonymous. Operational Research Society annual conference: Torquay November 13–16 1973 — Changes resulting from developments in telecommunications. The Computer Journal, 16(3):280, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        03/tiff/280.tif.

Anonymous:1974:SSP

[Ano74] Anonymous. Special summer programme at Massachusetts Institute of Technology. The Computer Journal, 17(2):192, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/192.tif.

Anonymous:1975:CPI

[Ano75a]  Anonymous.  Call for pa-

pers: 3rd International Conference on Automata, Languages and Programming, Edinburgh University, Scotland, July 20–23, 1976. The Computer Journal, 18(3):287, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/287.tif.

Anonymous:1975:SN

[Ano75b]  Anonymous.  Special no-

tice. The Computer Journal, 18(3):287, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/287.tif.

Anonymous:1976:EPV

[Ano76]  Anonymous.  Editorial: A

personal view of the ‘Beaumont’ weekend. The Computer Journal, 19(1):2, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
01/tiff/2.tif.

Anonymous:1978:ENR

[Ano78] Anonymous. Editorial notes: Reprinting of algorithms. The Computer Journal, 21(3):276, August 1978. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        03/tiff/276.tif.

Anonymous:1979:E

[Ano79a] Anonymous. Editorial. The Computer Journal, 22(1):2, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        22/Issue_01/tiff/2.tif.

Anonymous:1979:EA

[Ano79b] Anonymous. Erratum: Algorithm 105. The Computer Journal, 22(1):90, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        01/tiff/90.tif. Refers to

November 1978 issue, but no such entry yet exists.

Anonymous:1980:CP

[Ano80a] Anonymous. Calls for papers. The Computer Journal, 23(1):33, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_23/Issue_
01/tiff/33.tif.

Anonymous:1980:EEP

[Ano80b] Anonymous. Erratum: “An efficient predictor-corrector algorithm”. The Computer Journal, 23(4):383, November 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/383.tif. See [Wes80a,

Sne80].

Anonymous:1981:IA

[Ano81a] Anonymous. Information for authors. The Computer Journal, 24(1):2, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
01/tiff/2.tif.

Anonymous:1981:WP

[Ano81b]  Anonymous.  The Wilkes

Prize. The Computer Journal, 24(1):1, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
01/tiff/1.tif.

Anonymous:1982:CR

[Ano82a] Anonymous. Call for registration. The Computer Journal, 25(2):286, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/286.tif.

Anonymous:1982:IA

[Ano82b] Anonymous. Information for authors. The Computer Journal, 25(2):288, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/288.tif.

Anonymous:1996:BRa

[Ano96a] Anonymous. Book reviews. The Computer Journal, 39(7):??, ???? 1996. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_39/Issue_07/
                                        Vol39_07.body.html#AbstractReviews397.

Anonymous:1996:BRb

[Ano96b] Anonymous. Book reviews. The Computer Journal, 39(8):723– 730, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://www.oup.co.uk/jnls/
                                        list/comjnl/hdb/Volume_39/
                                        Issue_08/390723.sgm.abs.html.

Anonymous:1997:AI

[Ano97a] Anonymous. Author index for 1997. The Computer Journal, 40(10):??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_40/Issue_10/
Vol40_10.body.html#AbstractAuthors.

Anonymous:1997:SI

[Ano97b] Anonymous. Subject index for 1997. The Computer Journal, 40(10):??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_40/Issue_10/
Vol40_10.body.html#AbstractSubjects.

Albano:1980:HSR

[AO80a] A. Albano and R. Orsini. A heuristic solution of the rectangular cutting stock problem. The Computer Journal, 23(4):338–343, November 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/338.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/339.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/340.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/341.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/342.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/343.tif.

Albano:1980:TSA

[AO80b] A. Albano and R. Orsini. A tree search approach to the M-partition and Knapsack problems. The Computer Journal, 23(3):256–261, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/256.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/257.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/258.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/259.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/260.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/261.tif.

Arabnia:1987:TNA

[AO87] H. R. Arabnia and M. A. Oliver. A transputer network for the arbitrary rotation of digitised images. The Computer Journal, 30(5):425–432, October 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/425.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/426.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/427.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/428.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/429.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/430.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/431.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/432.tif.

Allison:1999:CAM

[APD99] L. Allison, D. Powell, and T. I. Dix. Compression and approximate matching. The Computer Journal, 42(1):1–10, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
01/420001.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
01/pdf/420001.pdf.

Anderson:1986:PS

[APW86] M. Anderson, R. D. Pose, and C. S. Wallace. A passwordcapability system. The Computer Journal, 29(1):1–8, February 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_

                                        01/tiff/1.tif;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_29/Issue_01/tiff/
                                        2.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_29/Issue_01/tiff/3.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        29/Issue_01/tiff/4.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        01/tiff/5.tif;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_29/Issue_01/tiff/
                                        6.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_29/Issue_01/tiff/7.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        29/Issue_01/tiff/8.tif.

Arazi:1994:PTU

[Ara94] Benjamin Arazi. On primality testing using purely divisionless operations. The Computer Journal, 37(3):219–222, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_37/Issue_03/
                                        Vol37_03.body.html#AbstractArazi.

Argo:1991:WWW

[Arg91] G. Argo. Weighting without waiting: the weighted path length tree. The Computer Journal, 34(5):444–449, October 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        05/tiff/444.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/445.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/446.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/447.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/448.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/449.tif.

Armstrong:1978:PPC

[Arm78] J. L. Armstrong. Programming a parallel computer for robot vision. The Computer Journal, 21(3):215–218, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/215.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/216.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/217.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/218.tif.

Arthur:1975:CRV

[Art75] D. F. Arthur. Correspondence:

Reading variable length records using FORTRAN IV. The Computer Journal, 18(3):283, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        02/tiff/283.tif. See [HV74,

ET75].

Anstey:1963:URD

[AS63] T. H. Anstey and K. W. Smillie. Use of a remote digital computer on an open-shop basis in agricultural research. The Computer Journal, 6(2):118–120, July 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/060118.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/tiff/118.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/tiff/119.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/tiff/120.tif.

Arblaster:1979:JSP

[ASG79] A. T. Arblaster, M. E. Sime, and T. R. G. Green. Jumping to some purpose. The Computer Journal, 22(2):105–109, May 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/105.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/106.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_

02/tiff/107.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/108.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/109.tif. See cor-

respondence [Hil80, Rob80, Tay80, MR82].

Ashby:1968:MPA

[Ash68]  T. Ashby.  A modification

to Paulson’s approximation to the variance ratio distribution. The Computer Journal, 11(2):209–210, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/110209.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/209.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/210.tif.

Allen:1970:CWA

[AT70] R. L. Allen and The Editor. Correspondence: What is an analyst? The Computer Journal, 13(1):119–120, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/tiff/119.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/tiff/120.tif. See [Fel69,

Fel70].

Atkinson:1976:LPL

[Atk76] L. V. Atkinson. List processing language, SLPL — an informal outline. The Computer Journal, 19(1):32–39, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        01/tiff/32.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_19/Issue_01/tiff/
                                        33.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_19/Issue_01/tiff/34.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        19/Issue_01/tiff/35.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        01/tiff/36.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_19/Issue_01/tiff/
                                        37.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_19/Issue_01/tiff/38.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        19/Issue_01/tiff/39.tif.

Atkinson:1980:PGS

[Atk80] M. D. Atkinson. Permutation groups and set union algorithms. The Computer Journal, 23(2):187, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        02/tiff/187.tif.

Atkinson:1984:JAG

[Atk84]  L. V. Atkinson.  Jumping

about and getting into a state. The Computer Journal, 27(1):42–46, February 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
01/tiff/42.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_01/tiff/
43.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_27/Issue_01/tiff/44.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
27/Issue_01/tiff/45.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
01/tiff/46.tif.

Atkinson:1993:UGR

[Atk93]  M. D. Atkinson.  Uniform

generation of rooted ordered trees with prescribed degrees. The Computer Journal, 36 (6):593–594, December 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_36/Issue_06/
Vol36_06.body.html#AbstractAtkinson.

Appelt:1993:FSI

[ATL93] W. Appelt and N. Tetteh-Lartey. The formal specification of the ISO open document architecture (ODA) standard. The Computer Journal, 36(3):269–279, June 1993.

CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_36/Issue_03/
Vol36_03.body.html#AbstractAppelt.

Austin:1971:DDA

[Aus71] B. J. Austin. A dynamic disc allocation algorithm designed to reduce fragmentation during file reloading. The Computer Journal, 14(4):378–381, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/140378.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/378.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/379.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/380.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/381.tif.

Aust:1976:IAS

[Aus76] R. J. Aust. An improvement algorithm for school timetabling. The Computer Journal, 19(4):339–343, November 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/339.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_19/Issue_
04/tiff/340.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/341.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/342.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/343.tif.

Anderssen:1971:INM

[AW71] A. S. Anderssen and E. T.

White. Improved numerical methods for Volterra integral equations of the first kind. The Computer Journal, 14(4):442–443, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/140442.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/442.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/443.tif. See correc-

tions [Lew72].

Abel:1985:RFC

[AW85] D. J. Abel and W. T.

Williams. A re-examination of four classificatory fusion strategies. The Computer Journal, 28(4):439–443, August 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_

04/tiff/439.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/440.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/441.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/442.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/443.tif.

Avison:1991:ISD

[AWH91] D. E. Avison and A. T. Wood-Harper. Information systems development research: an exploration of ideas in practice. The Computer Journal, 34(2):98–112, April 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/100.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/101.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/102.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/103.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/104.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/105.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_

02/tiff/106.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/107.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/108.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/109.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/110.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/111.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/112.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/98.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_02/tiff/
99.tif.

Axford:1990:RCC

[Axf90] T. H. Axford. Reference counting of cyclic graphs for functional programs. The Computer Journal, 33(5):466–470, October 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/466.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/467.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/468.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_33/Issue_
                                        05/tiff/469.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        05/tiff/470.tif.

Ayre:1964:UER

[Ayr64] T. H. Ayre. User’s experience of RAPIDWRITE. The Computer Journal, 7(2):101–102, July 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        02/070101.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        02/tiff/101.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        02/tiff/102.tif.

Azmoodeh:1990:BDI

[Azm90]  M. Azmoodeh.  BRMQ: a

database interface facility based on graph traversals and extended relationships on groups of entities. The Computer Journal, 33(1):31–39, February 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        01/tiff/31.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_33/Issue_01/tiff/
                                        32.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_33/Issue_01/tiff/33.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        33/Issue_01/tiff/34.tif;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
01/tiff/35.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_01/tiff/
36.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_33/Issue_01/tiff/37.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/38.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
01/tiff/39.tif.

Borger:1996:CCO

[B– D96] Egon. Börger and Igor – Dur– danovic.

Correctness of compiling Occam to Transputer code. The Computer Journal, 39(1):52– 92, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www.oup.co.uk/jnls/
list/comjnl/hdb/Volume_39/
Issue_01/390052.sgm.abs.html;
 http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_01/
Vol39_01.body.html#AbstractEgon.

Bjorn-Andersen:1988:HFH

[BA88] N. Bjorn-Andersen. Are ‘human factors’ human? The Computer Journal, 31(5):386–390, October 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/386.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/387.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_31/Issue_
                                        05/tiff/388.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        05/tiff/389.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        05/tiff/390.tif.

Backhouse:1971:CHL

[Bac71] R. C. Backhouse. Correspondence: On “High level languages for low level users”. The Computer Journal, 14(4): 365, November 1971. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        14/Issue_04/tiff/365.tif.

See [ER71, Wal71, Fel72, Hil72].

Baer:1962:NEL

[Bae62]  R. B. Baer.  Note on

an extremum locating algorithm. The Computer Journal, 5(3):193, October 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        03/050193.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        03/tiff/193.tif.

Baecker:1964:IUR

[Bae64] H. D. Baecker. Initial users’ reactions: what do they really want? The Computer Journal, 7(2):108–109, July 1964. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/070108.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/108.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/109.tif.

Baecker:1966:CAB

[Bae66] H. D. Baecker. Correspondence: On “An analysis, both theoretical and by simulation, of a time-shared computer system”. The Computer Journal, 9(2):128, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/090128.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/128.tif. See [Pen66].

Baecker:1970:UAT

[Bae70]  H. D. Baecker.  The use

of ALGOL 68 for trees. The Computer Journal, 13 (1):25–27, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/130025.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/tiff/25.tif;   http://
www3.oup.co.uk/computer_journal/

                                        hdb/Volume_13/Issue_01/tiff/
                                        26.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_13/Issue_01/tiff/27.
                                        tif.

Baecker:1975:AR

[Bae75] H. D. Baecker. Areas and record-classes. The Computer Journal, 18(3):223–226, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/223.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/224.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/225.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/226.tif.

Baecker:1984:ISE

[Bae84] H. D. Baecker. On instruction set encoding. The Computer Journal, 27(3):283, August 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        03/tiff/283.tif.

Bagley:1961:IPO

[Bag61a]  P. R. Bagley.  Improving

problem-oriented language by stratifying it. The Computer Journal, 4(3):217–221, October 1961. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/040217.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/217.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/218.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/219.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/220.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/221.tif.

Bagshaw:1961:OCS

[Bag61b] A. R. Bagshaw. Organization of a computing service for industry and commerce. The Computer Journal, 4(2):181–184, July 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/040181.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/181.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/182.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/183.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/184.tif.

Bagley:1962:PPU

[Bag62]  P. R. Bagley.  Principles

and problems of a Universal Computer-Oriented Language. The Computer Journal, 4(4):305–312, January 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        04/040305.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        04/tiff/305.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        04/tiff/306.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        04/tiff/307.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        04/tiff/308.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        04/tiff/309.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        04/tiff/310.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        04/tiff/311.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        04/tiff/312.tif.

Bailes:1988:API

[Bai88] P. A. Bailes. An advanced problem-oriented introduction to imperative programming with Modula-2. The Computer Journal, 31(6):531–539, December 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/531.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/532.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/533.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/534.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/535.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/536.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/537.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/538.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/539.tif.

Bakes:1964:AMS

[Bak64] M. D. Bakes. An alternative method of solution of certain tri-diagonal systems of linear equations. The Computer Journal, 7(2):135–136, July 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/070135.sgm.abs.html;
http://www3.oup.co.uk/computer_

                                        journal/hdb/Volume_07/Issue_
                                        02/tiff/135.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        02/tiff/136.tif.

Bakker:1994:SAE

[Bak94]  J. A. Bakker.  A seman-

tic approach to enforce correctness of data distribution schemes. The Computer Journal, 37(7):561–575, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_37/Issue_07/
                                        Vol37_07.body.html#AbstractBakker.

Barnard:1958:FYB

[Bar58] A. J. Barnard. The first year with a business computer. The Computer Journal, 1(1):29–36, April 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        01/010029.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        01/tiff/29.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_01/Issue_01/tiff/
                                        30.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_01/Issue_01/tiff/31.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        01/Issue_01/tiff/32.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        01/tiff/33.tif;   http://
                                        www3.oup.co.uk/computer_journal/

hdb/Volume_01/Issue_01/tiff/
34.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_01/Issue_01/tiff/35.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
01/Issue_01/tiff/36.tif.

Barrett:1961:CPG

[Bar61]  W. Barrett.  Convergence

properties of Gaussian quadrature formulae. The Computer Journal, 3(4):272–277, January 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/030272.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/272.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/273.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/274.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/275.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/276.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/277.tif.

Barron:1964:CRA

[Bar64] D. W. Barron. Correspondence: On the representation of ALGOL. The Computer Journal, 6(4):335, January 1964. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        04/tiff/335.tif. See [Sam63,

Tay64].

Barnes:1965:ASN

[Bar65a] J. G. P. Barnes. An algorithm for solving non-linear equations based on the secant method. The Computer Journal, 8(1):66–72, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        01/080066.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        01/tiff/66.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_08/Issue_01/tiff/
                                        67.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_08/Issue_01/tiff/68.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        08/Issue_01/tiff/69.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        01/tiff/70.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_08/Issue_01/tiff/
                                        71.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_08/Issue_01/tiff/72.
                                        tif.

Barnes:1965:KAL

[Bar65b] J. G. P. Barnes. A KDF9 ALGOL list-processing scheme. The Computer Journal, 8(2):

113–119, July 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
08/Issue_02/080113.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
08/Issue_02/tiff/113.tif;
 http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/114.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/115.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/116.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/117.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/118.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/119.tif.

Barraclough:1965:ADC

[Bar65c] E. D. Barraclough. The application of a digital computer to the construction of timetables. The Computer Journal, 8(2):136–146, July 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/080136.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/136.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_08/Issue_
                                        02/tiff/137.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        02/tiff/138.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        02/tiff/139.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        02/tiff/140.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        02/tiff/141.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        02/tiff/142.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        02/tiff/143.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        02/tiff/144.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        02/tiff/145.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        02/tiff/146.tif.

Barron:1966:CES

[Bar66] D. W. Barron. Correspondence: On “The Egdon system for the KDF9”. The Computer Journal, 9(3):247, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        03/090247.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_

03/tiff/247.tif. See [BHJV66]

and correspondence [Sam66].

Barton:1967:SMA

[Bar67] D. Barton. A scheme for manipulative algebra on a computer. The Computer Journal, 9(4):340–344, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/090340.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/340.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/341.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/342.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/343.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/344.tif.

Barron:1969:DCN

[Bar69] D. W. Barron. Discussion and correspondence: a note on program debugging in an online environment. The Computer Journal, 12(1):104, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
01/120104.sgm.abs.html;
http://www3.oup.co.uk/computer_

                                        journal/hdb/Volume_12/Issue_
                                        01/tiff/104.tif.

Barron:1971:ACF

[Bar71] D. W. Barron. Approaches to conversational FORTRAN. The Computer Journal, 14(2): 123–127, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        14/Issue_02/140123.sgm.abs.
                                        html;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        14/Issue_02/tiff/123.tif;
                                         http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/124.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/125.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/126.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/127.tif.

Bares:1972:CFD

[Bar72a] P. Bareš. Correspondence: On “The formal definition of the BASIC language”. The Computer Journal, 15(3):276, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/150276.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/276.tif. See [Lee72].

Barnes:1972:RTL

[Bar72b] J. G. P. Barnes. Real time languages for process control. The Computer Journal, 15(1):15–17, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/150015.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/tiff/15.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_01/tiff/
16.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_15/Issue_01/tiff/17.
tif.

Barron:1973:CEP

[Bar73] D. W. Barron. Correspondence:

On editorial policy. The Computer Journal, 16(3):284, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/284.tif. See [GT72].

Barron:1974:JCL

[Bar74]  D. W. Barron.  Job con-

trol languages and job control programs. The Computer Journal, 17(3):282–286, August 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/282.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/283.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/284.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/285.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/286.tif.

Barker:1975:SDP

[Bar75] P. G. Barker. Syntactic definition and parsing of molecular formulae. I. initial syntax definition and parser implementation. The Computer Journal, 18(4):355–359, November 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/355.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/356.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/357.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/358.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/359.tif.

Barlow:1978:ASP

[Bar78a] J. F. Barlow. An algorithm for the solution of the postman’s problem. The Computer Journal, 21(1):86–89, February 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
01/tiff/86.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_01/tiff/
87.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_21/Issue_01/tiff/88.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
21/Issue_01/tiff/89.tif. See

correspondence [Gol78a].

Barry:1978:DCR

[Bar78b] P. T. Barry. Discussion and correspondence: the Regional Computing Organisation Network. The Computer Journal, 21(2):184–187, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/184.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/185.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/186.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/187.tif.

Barton:1984:PAS

[Bar84]  D. Barton.  Problem with

algebra systems — revisited. The Computer Journal, 27(2):159–164, May 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/159.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/160.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/161.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/162.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/163.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/164.tif.

Barton:1985:PSA

[Bar85] David Barton. Parallelism in simple algebra systems. The Computer Journal, 28(2):142– 147, May 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
28/Issue_02/tiff/142.tif;
 http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/143.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/144.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/145.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/146.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_28/Issue_
02/tiff/147.tif.

Barringer:1987:TW

[Bar87] H. Barringer. Up and down the temporal way. The Computer Journal, 30(2):134–148, April 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/134.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/135.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/136.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/137.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/138.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/139.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/140.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/141.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/142.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/143.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/144.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_30/Issue_
                                        02/tiff/145.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        02/tiff/146.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        02/tiff/147.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        02/tiff/148.tif.

Bassiouni:1986:SPM

[Bas86]  M. A. Bassiouni.  System

and program models of storage allocation for reducing seek delay. The Computer Journal, 29(1):47–51, February 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        01/tiff/47.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_29/Issue_01/tiff/
                                        48.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_29/Issue_01/tiff/49.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        29/Issue_01/tiff/50.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        01/tiff/51.tif.

Bauldreay:1965:CCR

[Bau65]  J. Bauldreay.  Correspon-

dence: character recognition. The Computer Journal, 8(3):207, October 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/207.tif.

Bays:1973:STS

[Bay73] C. Bays. Some techniques for structuring chained hash tables. The Computer Journal, 16(2):126–131, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/160126.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/126.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/127.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/128.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/129.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/130.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/131.tif.

Bays:1976:NTR

[Bay76] C. Bays. Non-recursive technique for recreating a digraph from its K-formula representation. The Computer Journal, 19(4):326–328, November 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

journal/hdb/Volume_19/Issue_
04/tiff/326.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/327.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/328.tif.

Brigham:1959:TRD

[BB59] R. C. Brigham and C. G. Bell. A translation routine for the DEUCE computer. The Computer Journal, 2(2):76–84, July 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
02/020076.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
02/tiff/76.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_02/tiff/
77.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_02/Issue_02/tiff/78.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
02/Issue_02/tiff/79.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
02/tiff/80.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_02/tiff/
81.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_02/Issue_02/tiff/82.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
02/Issue_02/tiff/83.tif;
http://www3.oup.co.uk/computer_

journal/hdb/Volume_02/Issue_
02/tiff/84.tif.

Box:1969:CVR

[BB69] M. J. Box and R. M. Box. Computation of the variance ratio distribution. The Computer Journal, 12(3):277–278, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/120277.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/277.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/278.tif.

Berry:1971:NAH

[BB71] J. L. Berry and A. G. Bell.

Note on Algorithm 50: How to program a computer to play legal chess. The Computer Journal, 14(1):106–107, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/tiff/106.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/tiff/107.tif. See [Bel70a].

Barber:1980:BRD

[BB80a] D. L. A. Barber and J. M.

Brady. Book review: Data Communications and Teleprocessing Systems, by T. Horsley, 1979; A Technique for Software and Systems Design, by R. J. Lano, 1979. The Computer Journal, 23(2):141, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/141.tif.

Bishop:1980:PCS

[BB80b] J. M. Bishop and D. W. Barron. Procedure calling and structured architecture. The Computer Journal, 23(2):115–123, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/115.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/116.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/117.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/118.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/119.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/120.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/121.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/122.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_

02/tiff/123.tif. See corre-

spondence [ME80].

Bishop:1981:PD

[BB81] J. M. Bishop and D. W. Barron. Principles of descriptors. The Computer Journal, 24(3):210–221, August 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/210.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/211.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/212.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/213.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/214.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/215.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/216.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/217.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/218.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/219.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_

03/tiff/220.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/221.tif.

Brailsford:1989:EPJ

[BB89] D. F. Brailsford and R. J. Beach. Electronic publishing — a journal and its production. The Computer Journal, 32(6):482–493, December 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/482.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/483.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/484.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/485.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/486.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/487.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/488.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/489.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/490.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/491.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/492.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/493.tif.

Berrisford:1994:ROT

[BB94] G. Berrisford and M. Burrows. Reconciling OO with Turing machines. The Computer Journal, 37(10):888–906, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_37/Issue_10/
Vol37_10.body.html#AbstractBerrisford.

Barton:1968:SAS

[BBB68] D. Barton, S. R. Bourne, and C. J. Burgess. A simple algebra system. The Computer Journal, 11(3):293–298, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/110293.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/293.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/294.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/295.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/296.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_

                                        03/tiff/297.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        03/tiff/298.tif.

Burbridge:1966:CS

[BBdD66] D. A. Burbridge, A. Robert Brown, A. d’Agapeyeff, and F. G. Duncan. Correspondence: Systematics. The Computer Journal, 9(3):229– 229, 237–237, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        03/tiff/229.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        03/tiff/237.tif. See [Gri66].

Barton:1970:AS

[BBF70] D. Barton, S. R. Bourne, and J. P. Fitch. An algebra system. The Computer Journal, 13(1):32–39, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        01/130032.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        01/tiff/32.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_13/Issue_01/tiff/
                                        33.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_13/Issue_01/tiff/34.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        13/Issue_01/tiff/35.tif;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/tiff/36.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_01/tiff/
37.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/38.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
13/Issue_01/tiff/39.tif.

Benford:1994:SCW

[BBF^{+}94] Steve Benford, John Bowers, Lennart E. Fahlen, John Mariani, and Tom Rodden. Supporting cooperative work in virtual environments. The Computer Journal, 37(8):653–668, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_37/Issue_08/
Vol37_08.body.html#AbstractBenford.

Backus:1963:RRA

[BBG^{+}63] J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. Mc-Carthy, P. Naur, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Revised report on the algorithmic language ALGOL 60. The Computer Journal, 5(4):349–367, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/050349.sgm.abs.html;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/349.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/350.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/351.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/352.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/353.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/354.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/355.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/356.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/357.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/358.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/359.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/360.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/361.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/362.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_

04/tiff/363.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/364.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/365.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/366.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/367.tif.

Ball:1971:FRA

[BBG^{+}71] D. Ball, P. M. Blackall, V. Gerard, G. R. Macleod, P. J. Marcer, and E. M. Palandri. FOCUS — a remote access file handling system on-line to a CDC 6000 series computer. The Computer Journal, 14(2):113–118, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/140113.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/113.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/114.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/115.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/116.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/117.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/118.tif.

Backus:1976:MRA

[BBG^{+}76] J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. Mc-Carthy, P. Naur, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, M. Woodger, R. M. De Morgan, I. D. Hill, and B. A. Wichmann. Modified report on the algorithmic language ALGOL-60. The Computer Journal, 19(4):364–379, November 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        04/tiff/364.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        04/tiff/365.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        04/tiff/366.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        04/tiff/367.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        04/tiff/368.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        04/tiff/369.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        04/tiff/370.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        04/tiff/371.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/372.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/373.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/374.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/375.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/376.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/377.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/378.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/379.tif.

Bernardo:1995:DSE

[BBG95] Marco Bernardo, Nadia Busi, and Roberto Gorrieri. A distributed semantics for EMPA based on stochastic contextual nets. The Computer Journal, 38(7):492–509, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_38/Issue_07/
Vol38_07.body.html#AbstractBernardo.

Barron:1963:MFC

[BBH^{+}63] D. W. Barron, J. N. Buxton, D. F. Hartley, E. Nixon, and C. Strachey. The main features of CPL. The Computer Journal, 6(2):134–143, July 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/060134.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/134.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/135.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/136.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/137.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/138.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/139.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/140.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/141.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/142.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/143.tif.

Barton:1970:SCA

[BBH70] D. Barton, S. R. Bourne, and J. R. Horton. The structure of the Cambridge algebra system. The Computer Journal, 13(3):243–247, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/130243.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/243.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/244.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/245.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/246.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/247.tif.

Baeten:1987:RSC

[BBK87] J. C. M. Baeten, J. A. Bergstra, and J. W. Klop. Readytrace semantics for concrete process algebra with the priority operator. The Computer Journal, 30(6):498–506, December 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/498.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/499.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/500.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_

                                        06/tiff/501.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        06/tiff/502.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        06/tiff/503.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        06/tiff/504.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        06/tiff/505.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        06/tiff/506.tif.

Bevan:1989:PDD

[BBKR89] D. I. Bevan, G. L. Burn, R. J. Karia, and J. D. Robson. Principles for the design of a distributed memory architecture for parallel graph reduction. The Computer Journal, 32(5):461–469, October 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        05/tiff/461.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        05/tiff/462.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        05/tiff/463.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        05/tiff/464.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        05/tiff/465.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_32/Issue_
05/tiff/466.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/467.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/468.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/469.tif.

Bergstra:1994:PAI

[BBP94] Jan A. Bergstra, Inge Bethke, and Alban Ponse. Process algebra with iteration and nesting. The Computer Journal, 37(4):243–258, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_37/Issue_04/
Vol37_04.body.html#AbstractBergstra.

Baldoni:1998:SFC

[BBP98] Roberto Baldoni, Roberto Beraldi, and Ravi Prakash. Slotted-FIFO communication for asynchronous distributed systems. The Computer Journal, 41(5):337–348, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_41/Issue_05/
Vol41_05.body.html#AbstractBaldoni.

Baum:1966:UDA

[BBS66] M. M. Baum, R. G. Blake, and R. J. Smale. Use of digital analogue simulator (DAS). The Computer Journal, 9(2):175–180, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/090175.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/175.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/176.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/177.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/178.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/179.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/180.tif.

Brooker:1967:IDC

[BC67] R. A. Brooker and S. R. Clark. An index directed compiler. The Computer Journal, 10(1):60–63, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/100060.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/60.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_01/tiff/
61.tif;   http://www3.oup.
co.uk/computer_journal/hdb/

Volume_10/Issue_01/tiff/62.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/63.tif.

Broadbent:1973:MME

[BC73] J. K. Broadbent and G. F.

Coulouris. MEMBERS — a Microprogrammed Experimental Machine with a Basic Executive for Real-time Systems. The Computer Journal, 16(3):205–208, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/160205.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/205.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/206.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/207.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/208.tif.

Bastani:1991:MSA

[BCH91] F. B. Bastani, I.-R. Chen, and W. Hilal. A model for the stability analysis of maintenance strategies for linear list. The Computer Journal, 34(1):80–87, February 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_

01/tiff/80.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_01/tiff/
81.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_34/Issue_01/tiff/82.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
34/Issue_01/tiff/83.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
01/tiff/84.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_01/tiff/
85.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_34/Issue_01/tiff/86.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
34/Issue_01/tiff/87.tif.

Boote:1972:SPC

[BCR72] W. P. Boote, S. R. Clark, and T. A. Rourke. Simulation of a paging computer system. The Computer Journal, 15(1):51–57, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/150051.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/tiff/51.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_01/tiff/
52.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_15/Issue_01/tiff/53.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
15/Issue_01/tiff/54.tif;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/tiff/55.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_01/tiff/
56.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_15/Issue_01/tiff/57.
tif.

Bergsten:1993:OPA

[BCV93] B. Bergsten, M. Couprie, and P. Valduriez. Overview of parallel architectures for databases. The Computer Journal, 36(8):734–740, ???? 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_08/
Vol36_08.body.html#AbstractBergsten.

Bennett:1961:CAC

[BD61] J. M. Bennett and R. J.

Dakin. Computers as an aid in computer design assessment. The Computer Journal, 3(4):253–256, January 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/030253.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/253.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/254.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/255.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        04/tiff/256.tif.

Bennett:1972:FRT

[BD72] D. G. Bennett and R. A. Davenport. A facility for real-time program development. The Computer Journal, 15(1):5–7, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        01/150005.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        01/tiff/5.tif;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_15/Issue_01/tiff/
                                        6.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_15/Issue_01/tiff/7.
                                        tif.

Boccalatte:1980:ADS

[BD80] A. Boccalatte and M. Di Manzo. An approach to the detection of semantic errors. The Computer Journal, 23(4):317–323, November 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        04/tiff/317.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        04/tiff/318.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        04/tiff/319.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_23/Issue_
04/tiff/320.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/321.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/322.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/323.tif.

Bateman:1981:CAS

[BD81] S. O. Bateman, Jr. and S. D. Danielopoulos. Computerised analytic solutions of second order differential equations. The Computer Journal, 24(2):180–183, May 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/180.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/181.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/182.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/183.tif.

Bell:1982:KSC

[BD82] D. A. Bell and S. M. Deen. Key space compression and hashing in PRECI. The Computer Journal, 25(4):486–492, November 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

journal/hdb/Volume_25/Issue_
04/tiff/486.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/487.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/488.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/489.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/490.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/491.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/492.tif.

Bell:1984:HTV

[BD84] D. A. Bell and S. M. Deen. Hash trees versus B-trees. The Computer Journal, 27 (3):218–224, August 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/218.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/219.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/220.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/221.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/222.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/223.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/224.tif.

Burmester:1992:EZK

[BDB92] M. V. D. Burmester, Y. G.

Desmedt, and T. Beth. Efficient zero-knowledge identification schemes for smart cards. The Computer Journal, 35(1):21–29, February 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_35/Issue_01/
Vol35_01.body.html#AbstractBurmester.

Blundo:1999:RKA

[BDG99] Carlo Blundo, P. D’Arco, and A. G. Gaggia. A τ-restricted key agreement scheme. The Computer Journal, 42(1):51–61, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
01/420051.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
01/pdf/420051.pdf.

Boccalatte:1982:ERA

[BDS82] A. Boccalatte, M. Di Manzo, and D. Sciarra. Error recovery with attribute grammars. The Computer Journal, 25(3):331–337, August 1982. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/331.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/332.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/333.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/334.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/335.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/336.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/337.tif.

Benson:1973:AAS

[BE73] A. Benson and D. J. Evans. Algorithm 80: An algorithm for the solution of periodic quindiagonal systems of linear equations. The Computer Journal, 16(3):278–279, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/278.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/279.tif.

Barlow:1979:POB

[BE79] R. H. Barlow and D. J. Evans. A parallel organisation of the bisection algorithm. The Computer Journal, 22(3):267–269, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/267.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/268.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/269.tif.

Barlow:1982:PAI

[BE82] R. H. Barlow and D. J.

Evans. Parallel algorithms for the iterative solution to linear systems. The Computer Journal, 25(1):56–60, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/56.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/
57.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_25/Issue_01/tiff/58.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/59.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/60.tif.

Beasley:1965:NAC

[Bea65] J. D. Beasley. A note on the arrangement of Chebyshev series. The Computer Journal, 8(3):278–279, October 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        03/080278.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        03/tiff/278.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        03/tiff/279.tif.

Beauchamp:1968:SWF

[Bea68] K. G. Beauchamp. The simulation of wave filters having polynomial transfer functions on an analogue computer. The Computer Journal, 10(4):352–359, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/100352.sgm.abs.html.

Beasley:1974:CSR

[Bea74] J. D. Beasley. Correspondence: System 4 random number. The Computer Journal, 17(4):381, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        04/tiff/381.tif.

Beck:1958:HAU

[Bec58]  F. Beck.  Harmonic anal-

ysis using a digital computer. The Computer Journal, 1(3):117, October 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/010117.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/117.tif.

Beilby:1972:RTS

[Bei72] M. H. Beilby. Road traffic simulation on a small computer. The Computer Journal, 15(2):134–137, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/150134.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/134.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/135.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/136.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/137.tif.

Bell:1958:PS

[Bel58]  D. A. Bell.  The princi-

ples of sorting. The Computer Journal, 1(2):71–77, July 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

journal/hdb/Volume_01/Issue_
02/010071.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
02/tiff/71.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_02/tiff/
72.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_01/Issue_02/tiff/73.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
01/Issue_02/tiff/74.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
02/tiff/75.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_02/tiff/
76.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_01/Issue_02/tiff/77.
tif.

Bell:1963:AFT

[Bel63]  D. A. Bell.  Approxima-

tions in Fourier transforms. The Computer Journal, 6 (3):244–247, October 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/060244.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/244.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/245.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/246.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_06/Issue_
03/tiff/247.tif.

Bell:1970:AHP

[Bel70a] A. G. Bell. Algorithm 50:

How to program a computer to play legal chess. The Computer Journal, 13(2):208–219, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/208.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/209.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/210.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/211.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/212.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/213.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/214.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/215.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/216.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/217.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/218.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/tiff/219.tif. See note

[BB71].

Bell:1970:PID

[Bel70b] A. G. Bell. Partitioning integers in N dimensions. The Computer Journal, 13(3):278–283, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/130278.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/278.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/279.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/280.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/281.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/282.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/283.tif.

Bell:1973:H

[Bel73]  A. G. Bell.  Himmelbett.

The Computer Journal, 16 (3):269–272, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        03/160269.sgm.abs.html;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/269.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/270.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/271.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/272.tif. See corre-

spondence [Fre74, Str74].

Bell:1974:CCS

[Bel74] D. A. Bell. Correspondence:

Computing Science. The Computer Journal, 17(4):383, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
17/Issue_04/tiff/383.tif.

See [Bro73a, Fis74a].

Bell:1976:PSP

[Bel76]  D. Bell.  Programmer se-

lection and programming errors. The Computer Journal, 19(3):202–206, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/202.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/203.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/204.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/205.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/206.tif.

Bell:1984:DDP

[Bel84] David A. Bell. Difficult data placement problems. The Computer Journal, 27(4):315– 320, November 1984. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        27/Issue_04/tiff/315.tif;
                                         http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        04/tiff/316.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        04/tiff/317.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        04/tiff/318.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        04/tiff/319.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        04/tiff/320.tif.

Bemer:1974:CMI

[Bem74] R. W. Bemer. Correspondence: On mixing interpretive and machine code. The Computer Journal, 17(2):188–189, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        02/tiff/188.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/189.tif. See [Daw73,

DP73, Mac74].

Bennett:1970:DCC

[Ben70] R. D. Bennett. Discussion and correspondence: computers and the business information structure: new horizons in decision making effectiveness. The Computer Journal, 13(3):327–329, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/130327.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/327.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/328.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/329.tif.

Bennetts:1972:RAD

[Ben72] R. G. Bennetts. A realistic approach to detection test set generation for combinational logic circuits. The Computer Journal, 15(3):238–246, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/150238.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/238.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/239.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/240.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/241.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/242.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/243.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/244.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/245.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/246.tif.

Bengtsson:1982:PRP

[Ben82] Bengt Erik Bengtsson. Packing rectangular pieces — a heuristic approach. The Computer Journal, 25(3):353–357, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/353.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/354.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/355.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_

03/tiff/356.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/357.tif.

Benyon:1985:EVC

[Ben85] Peter R. Benyon. Exploiting vector computers by replication. The Computer Journal, 28(2):138–141, May 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/138.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/139.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/140.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/141.tif.

Bai:1998:FTB

[BENM98] L. Q. Bai, H. Ebara, H. Nakano, and H. Maeda. Fault-tolerant broadcasting on the arrangement graph. The Computer Journal, 41(3):171–184, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_41/Issue_03/
Vol41_03.body.html#AbstractBai.

Berry:1959:ISC

[Ber59] F. J. Berry. Intercode, a simplified coding scheme for AMOS. The Computer Journal, 2(2):55–58, July 1959.

CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
02/020055.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
02/tiff/55.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_02/tiff/
56.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_02/Issue_02/tiff/57.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
02/Issue_02/tiff/58.tif.

Berry:1966:GP

[Ber66] R. E. Berry. Grader programs. The Computer Journal, 9(3):252–256, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/090252.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/252.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/253.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/254.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/255.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/256.tif.

Bergson:1969:CNC

[Ber69] A. Bergson. Correspondence:

Note on the calculation of e to many significant digits. The Computer Journal, 12(1):105, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
01/tiff/105.tif.

Berry:1976:ADD

[Ber76]  G. Berry.  Algorithm 92:

The drawing of dashed lines. The Computer Journal, 19 (4):361–363, November 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/361.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/362.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/363.tif.

Berghel:1987:CCH

[Ber87a] H. Berghel. Crossword compilation with Horn clauses. The Computer Journal, 30(2):183– 188, April 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
30/Issue_02/tiff/183.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_

02/tiff/184.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/185.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/186.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/187.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/188.tif.

Bertino:1987:EPI

[Ber87b]  Lisa Bertino.  An evalua-

tion of precompilation and interpretation in distributed database management systems. The Computer Journal, 30 (6):519–528, December 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/519.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/520.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/521.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/522.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/523.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/524.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_

06/tiff/525.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/526.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/527.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/528.tif.

Barlow:1983:PMA

[BES83] R. H. Barlow, D. J. Evans, and J. Shanehchi. Parallel multisection applied to the eigenvalue problem. The Computer Journal, 26(1):6–9, February 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
01/tiff/6.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_26/Issue_01/tiff/
7.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_26/Issue_01/tiff/8.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
26/Issue_01/tiff/9.tif.

Brittan:1971:CTC

[BF71] J. N. G. Brittan and F. J. M.

Farley. College timetable construction by computer. The Computer Journal, 14 (4):361–365, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/140361.sgm.abs.html;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/361.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/362.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/363.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/364.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/365.tif.

Barton:1972:RAM

[BF72] D. Barton and J. P. Fitch. A review of algebraic manipulative programs and their application. The Computer Journal, 15(4):362–381, November 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/150362.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/362.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/363.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/364.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/365.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/366.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_15/Issue_
04/tiff/367.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/368.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/369.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/370.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/371.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/372.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/373.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/374.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/375.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/376.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/377.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/378.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/379.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/380.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/381.tif.

Buckley:1978:ADN

[BF78] James J. Buckley and Robert A. Fleck, Jr. Algorithm 105: Determining the number of internal rates of return. The Computer Journal, 21(4):373– 377, November 1978. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        21/Issue_04/tiff/373.tif;
                                         http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        04/tiff/374.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        04/tiff/375.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        04/tiff/376.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        04/tiff/377.tif.

Brinck:1981:AAT

[BF81] K. Brinck and N. Y. Foo. Analysis of algorithms on threaded trees. The Computer Journal, 24(2):148–155, May 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        02/tiff/148.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        02/tiff/149.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        02/tiff/150.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_24/Issue_
02/tiff/151.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/152.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/153.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/154.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/155.tif.

Braunholtz:1961:NPL

[BFH61] T. G. H. Braunholtz, A. G.

Fraser, and P. M. Hunt. NEB-ULA: a programming language for data processing. The Computer Journal, 4 (3):197–211, October 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/040197.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/197.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/198.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/199.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/200.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/201.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_04/Issue_
                                        03/tiff/202.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        03/tiff/203.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        03/tiff/204.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        03/tiff/205.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        03/tiff/206.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        03/tiff/207.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        03/tiff/208.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        03/tiff/209.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        03/tiff/210.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        03/tiff/211.tif.

Baker:1964:NSF

[BFMW64] C. T. H. Baker, L. Fox, D. F. Mayers, and K. Wright. Numerical solution of Fredholm integral equations of first kind. The Computer Journal, 7(2):141–148, July 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        02/070141.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_

journal/hdb/Volume_07/Issue_
02/tiff/141.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/142.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/143.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/144.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/145.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/146.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/147.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/148.tif.

Bridle:1971:CPI

[BG71] J. W. Bridle and R. J.

Gregersen. CMSR — a personnel information system. The Computer Journal, 14 (4):338–343, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/140338.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/338.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/339.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_

                                        04/tiff/340.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/341.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/342.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/343.tif.

Boulton:1979:MDL

[BG79] P. I. P. Boulton and J. R. Goguen. A machine description language. The Computer Journal, 22(2):132–135, May 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/132.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/133.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/134.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/135.tif.

Bezem:1994:CPO

[BG94] M. A. Bezem and J. F. Groote. A correctness proof of a onebit sliding window protocol in μCRL. The Computer Journal, 37(4):289–307, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_37/Issue_04/
                                        Vol37_04.body.html#AbstractBezem.

Bull:1978:PMM

[BGL^{+}78] G. M. Bull, S. Gilbey, Kathleen Levine, A. G. Lippiatt, and C. H. C. Machin. Proteus — a microcoded multiprocessor system. The Computer Journal, 21(2):154–160, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/154.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/155.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/156.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/157.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/158.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/159.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/160.tif.

Barringer:1996:PAF

[BGMW96] Howard Barringer, Graham Gough, Brian Monahan, and Alan Williams. A process algebra foundation for reasoning about Core ELLA. The Computer Journal, 39(4):303– 324, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www.oup.co.uk/jnls/

                                        list/comjnl/hdb/Volume_39/
                                        Issue_04/390303.sgm.abs.html;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_39/Issue_04/
                                        Vol39_04.body.html#AbstractBarringer.

Bhagavathi:1994:GHL

[BGO94] D. Bhagavathi, C. E. Grosch, and S. Olariu. A greedy hypercube-labeling algorithm. The Computer Journal, 37 (2):124–128, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_37/Issue_02/
                                        Vol37_02.body.html#AbstractBhagavathi.

Boardman:1986:RAO

[BGR86] J. T. Boardman, C. Garrett, and G. C. A. Robson. A recursive algorithm for the optimal solution of a complex allocation problem using a dynamic programming formulation. The Computer Journal, 29(2):182–186, April 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        02/tiff/182.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        02/tiff/183.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        02/tiff/184.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        02/tiff/185.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/186.tif.

Benjamin:1998:FAP

[BGR98] R. Benjamin, B. Gladman, and B. Randell. Protecting IT systems from Cyber-Crime. The Computer Journal, 41(7):429–443, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_41/Issue_07/
Vol41_07.body.html#AbstractBenjamin.

Bickmore:1999:WPF

[BGS99] Timothy Bickmore, Andreas Girgensohn, and Joseph W. Sullivan. Web page filtering and reauthoring for mobile users. The Computer Journal, 42(6):534– 546, 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
06/420534.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
06/pdf/420534.pdf.

Barron:1963:TPE

[BH63] D. W. Barron and D. F. Hartley. Techniques for program error diagnosis on EDSAC 2. The Computer Journal, 6(1):44–49, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_

01/060044.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/tiff/44.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/
45.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_06/Issue_01/tiff/46.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
06/Issue_01/tiff/47.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/tiff/48.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/
49.tif.

Barnard:1968:URD

[BH68] D. F. Barnard and D. F. Holman. The use of Roth’s decomposition algorithm in multi-level design of circuits. The Computer Journal, 11 (3):269–276, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/110269.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/269.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/270.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/271.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/272.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/273.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/274.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/275.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/276.tif.

Bruyendonck:1969:CCD

[BH69] D. Bruyendonck and D. Hirschberg.

Correspondence: Convex differentiable curves. The Computer Journal, 12(2):200, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/200.tif. See [Phi68a].

Burns:1966:ESK

[BHJV66] D. Burns, E. N. Hawkins, D. R. Judd, and J. L. Venn. The Egdon system for the KDF9. The Computer Journal, 8(4):297–302, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/080297.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/297.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/298.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_08/Issue_
                                        04/tiff/299.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        04/tiff/300.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        04/tiff/301.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        04/tiff/302.tif. See corre-

spondence [Bar66, Sam66].

Burton:1985:MGT

[BHK85] F. Warren Burton, Matthew M. Huntbach, and J. (Yiannis) G. Kollias. Multiple generation text files using overlapping tree structures. The Computer Journal, 28(4):414–416, August 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        04/tiff/414.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        04/tiff/415.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        04/tiff/416.tif.

Buckle:1969:CSF

[BHRS69] D. G. R. Buckle, C. H. Hunt, Maureen M. Ryder, and S. H. Storey. Computers and the small firm. 2. The Computer Journal, 12(1):10–14, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_

01/120010.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
01/tiff/10.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_01/tiff/
11.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_12/Issue_01/tiff/12.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/13.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
01/tiff/14.tif.

Buckle:1970:CSF

[BHRS70] D. G. R. Buckle, C. H. Hunt, M. M. Ryder, and S. H. Storey. Computers and the small firm: 3. The Computer Journal, 13(1):7–9, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/130007.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/tiff/7.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_13/Issue_01/tiff/
8.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/9.
tif.

Beale:1965:EUD

[BHS65] E. M. L. Beale, P. A. B.

Hughes, and R. E. Small. Experiences in using a decomposition program. The Computer Journal, 8(1):13–18, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
01/080013.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
01/tiff/13.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_01/tiff/
14.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_08/Issue_01/tiff/15.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
08/Issue_01/tiff/16.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
01/tiff/17.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_01/tiff/
18.tif.

Bonczek:1978:ITW

[BHW78] R. H. Bonczek, C. W. Holsapple, and A. B. Whinston. Information transferral within a distributed data base via a generalised mapping language. The Computer Journal, 21(2):110–116, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/110.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/111.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/112.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/113.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/114.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/115.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/116.tif.

Bloss:1989:OCM

[BHY89] Adrienne Bloss, P. R. Hudak, and J. H. Young. An optimising compiler for a modern functional language. The Computer Journal, 32(2):152–161, April 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/152.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/153.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/154.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/155.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/156.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/157.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/158.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_32/Issue_
                                        02/tiff/159.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        02/tiff/160.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        02/tiff/161.tif.

Bickley:1968:PCI

[Bic68]  W. G. Bickley.  Piece-

wise cubic interpolation and two-point boundary problems. The Computer Journal, 11 (2):206–208, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        02/110206.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        02/tiff/206.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        02/tiff/207.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        02/tiff/208.tif.

Bickley:1969:CPC

[Bic69] W. G. Bickley. Correspondence: Piecewise cubic interpolation and two-point boundary problems. The Computer Journal, 12(1):105, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        01/tiff/105.tif.

Birtwistle:1968:ESE

[Bir68] G. M. Birtwistle. The explicit solution of the equation of heat conduction. The Computer Journal, 11(3):317–323, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/110317.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/317.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/318.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/319.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/320.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/321.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/322.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/323.tif.

Bird:1974:SP

[Bir74] R. Bird. Speeding up programs. The Computer Journal, 17(4):337–339, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/337.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        04/tiff/338.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        04/tiff/339.tif.

Bird:1979:REV

[Bir79] R. S. Bird. Recursion elimination with variable parameters. The Computer Journal, 22(2):151–154, May 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/151.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/152.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/153.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/154.tif.

Bird:1981:DCD

[Bir81] P. R. Bird. Discussion and correspondence: Data structures for catalogues of sound recordings. The Computer Journal, 24(1):92–94, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        01/tiff/92.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_24/Issue_01/tiff/
                                        93.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/

Volume_24/Issue_01/tiff/94.
tif.

Bird:1989:AIP

[Bir89] R. S. Bird. Algebraic identities for program calculation. The Computer Journal, 32(2):122–126, April 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/122.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/123.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/124.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/125.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/126.tif.

Bastani:1988:CMD

[BIY88] F. B. Bastani, S. S. Iyengar, and I-Ling Yen. Concurrent maintenance of data structures in a distributed environment. The Computer Journal, 31(2):165–174, April 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/165.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/166.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_31/Issue_
02/tiff/167.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/168.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/169.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/170.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/171.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/172.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/173.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/174.tif.

Barker:1978:SDP

[BJ78] P. G. Barker and P. S. Jones. Syntactic definition and parsing of molecular formulae. part 2: Graphical synthesis of molecular formulae for data base queries. The Computer Journal, 21(3):224–233, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/224.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/225.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/226.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/227.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/228.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/229.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/230.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/231.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/232.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/233.tif.

Bell:1979:HRM

[BJ79] A. G. Bell and N. Jacobi.

How to read, make and store chess moves. The Computer Journal, 22(1):71–76, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
01/tiff/71.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_01/tiff/
72.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_22/Issue_01/tiff/73.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
22/Issue_01/tiff/74.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
01/tiff/75.tif;   http://

                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_22/Issue_01/tiff/
                                        76.tif.

Bryn-Jones:1980:CPS

[BJ80]  Mervyn Bryn-Jones.  Cor-

respondence: Points and nsided irregular figures. The Computer Journal, 23(2):190, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        02/tiff/190.tif.

Bodorik:1998:MVM

[BJ98] P. Bodorik and D. Jutla. Multiview memory support to operating systems in locking for transaction and database systems. The Computer Journal, 41(2):84–97, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_41/Issue_02/
                                        Vol41_02.body.html#AbstractBodorik.

Ballard:1974:ASC

[BJS74] Dana H. Ballard, C. O. Jelinek, and R. Schinzinger. An algorithm for the solution of constrained generalised polynomial programming problems. The Computer Journal, 17(3):261–266, August 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        03/tiff/261.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/262.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/263.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/264.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/265.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/266.tif. See corre-

spondence [Bra75].

Berkling:1975:CDH

[BK75] K. Berkling and W. Kluge.

Conceptual design of a hardwired main storage management unit. The Computer Journal, 18(4):360–365, November 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/360.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/361.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/362.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/363.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/364.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_

                                        04/tiff/365.tif.

Boulton:1979:EPR

[BK79a] P. I. P. Boulton and M. A. R. Kittler. Estimating program reliability. The Computer Journal, 22(4):328–331, November 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        04/tiff/328.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        04/tiff/329.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        04/tiff/330.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        04/tiff/331.tif.

Burton:1979:ODH

[BK79b] F. W. Burton and J. Kollias. Optimising disc head movements in secondary key retrievals. The Computer Journal, 22(3):206–208, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        03/tiff/206.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        03/tiff/207.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        03/tiff/208.tif.

Burton:1983:CEQ

[BK83] F. Warren Burton and J. G.

Kollias. Comment on “The explicit quad tree as a structure for computer graphics”. The Computer Journal, 26(2):188, May 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/188.tif.

Backhouse:1987:WRM

[BK87] R. C. Backhouse and A. A.

Khamiss. A while-rule in Martin-Löf’s theory of types. The Computer Journal, 30 (1):27–36, February 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
01/tiff/27.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_01/tiff/
28.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_30/Issue_01/tiff/29.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
30/Issue_01/tiff/30.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
01/tiff/31.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_01/tiff/
32.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_30/Issue_01/tiff/33.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
30/Issue_01/tiff/34.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_

                                        01/tiff/35.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_30/Issue_01/tiff/
                                        36.tif.

Born:1992:TPB

[BK92] R. G. Börn and J. R. Kenevan. Theoretical performance-based cost-effectiveness of multicomputers. The Computer Journal, 35(1):62–70, February 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_35/Issue_01/
                                        Vol35_01.body.html#AbstractBorn.

Burston:1978:DLA

[BKK78] A. K. Burston, D. J. Kinniment, and Hilary Kahn. A design language for asynchronous logic. The Computer Journal, 21(4):347–354, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        04/tiff/347.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        04/tiff/348.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        04/tiff/349.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        04/tiff/350.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        04/tiff/351.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_

04/tiff/352.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/353.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/354.tif.

Burton:1984:CPT

[BKK84] F. W. Burton, V. J. Kollias, and J. G. Kollias. Consistency in point-in-polygon tests. The Computer Journal, 27(4):375–376, November 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/375.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/376.tif.

Burton:1987:GPP

[BKK87] F. W. Burton, V. J. Kollias, and J. G. Kollias. A general PASCAL program for map overlay of quadtrees and related problems. The Computer Journal, 30(4):355–361, August 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/355.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/356.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/357.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_30/Issue_
                                        04/tiff/358.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        04/tiff/359.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        04/tiff/360.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        04/tiff/361.tif.

Burnett:1974:SRP

[BKL74] P. Burnett, P. A. Kidd, and A. M. Lister. Simulation of real-time program faults. The Computer Journal, 17(1):25–27, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        01/tiff/25.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_17/Issue_01/tiff/
                                        26.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_17/Issue_01/tiff/27.
                                        tif.

Brinksma:1995:SCB

[BKLL95] Ed Brinksma, Joost-Pieter Katoen, Rom Langerak, and Diego Latella. A stochastic causality-based process algebra. The Computer Journal, 38(7):552–565, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_38/Issue_07/
                                        Vol38_07.body.html#AbstractBrinksma.

Bratko:1978:PRC

[BKM78] I. Bratko, D. Kopec, and D. Michie. Pattern-based representation of chess endgame knowledge. The Computer Journal, 21(2):149–153, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/149.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/150.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/151.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/152.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/153.tif.

Burton:1990:IOB

[BKMK90] F. W. Burton, J. G. Kollias, D. G. Matsakis, and V. G. Kollias. Implementation of overlapping B-trees for time and space efficient representation of collections of similar files. The Computer Journal, 33(3):279–280, June 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/279.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/280.tif.

Bozanis:1996:RBI

[BKMT96] Panayiotis Bozanis, Nectarios Kitsios, Christos Makris, and Athanasios Tsakalidis. Red-blue intersection reporting for objects of non-constant size. The Computer Journal, 39(6):541– 546, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://www.oup.co.uk/jnls/
                                        list/comjnl/hdb/Volume_39/
                                        Issue_06/390541.sgm.abs.html;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_39/Issue_06/
                                        Vol39_06.body.html#AbstractBozanis.

Bozanis:1997:NRI

[BKMT97] P. Bozanis, N. Kitsios, C. Makris, and A. Tsakalidis. New results on intersection query problems. The Computer Journal, 40(1):22–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www.oup.co.uk/computer_
                                        journal/Volume_40/Issue_01/
                                        Vol40_01.body.html#AbstractBozanis;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_40/Issue_01/
                                        Vol40_01.body.html#AbstractBozanis.

Buxton:1962:CSL

[BL62] J. N. Buxton and J. G. Laski. Control and simulation language. The Computer Journal, 5(3):194–199, October 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_

03/050194.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/194.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/195.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/196.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/197.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/198.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/199.tif.

Bennetts:1971:FDD

[BL71] R. G. Bennetts and D. W.

Lewin. Fault diagnosis of digital systems — a review. The Computer Journal, 14(2):199–206, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/140199.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/199.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/200.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/201.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/202.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/203.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/204.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/205.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/206.tif.

Beaven:1972:APP

[BL72] P. A. Beaven and D. W. Lewin. An associative parallel processing system for non-numerical computation. The Computer Journal, 15(4):343–349, November 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/150343.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/tiff/343.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/tiff/344.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/tiff/345.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/tiff/346.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/tiff/347.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/tiff/348.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_15/Issue_
04/tiff/349.tif.

Blazie:1977:CCP

[BL77] Deane B. Blazie and L. S. Levy.

A cross compiler for pocket calculators. The Computer Journal, 20(3):213–221, August 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/213.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/214.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/215.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/216.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/217.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/218.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/219.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/220.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/221.tif.

Bell:1981:BRS

[BL81a] A. G. Bell and R. Lovett. Book reviews: Software Reliability Guidebook, by Robert L. Glass, 1979; Foundations of Computer Science, by M. S. Carberry, H. M. Khalil and J. F. Leathrum and L. S. Levy, 1979. The Computer Journal, 24(1):91, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        01/tiff/91.tif.

Burton:1981:ADT

[BL81b] F. Warren Burton and Brian J. Lings. Abstract data types, subtypes and data independence. The Computer Journal, 24(4):308–311, November 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        04/tiff/308.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        04/tiff/309.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        04/tiff/310.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        04/tiff/311.tif.

Bilodeau:1990:DSM

[BL90] M. Bilodeau and D. Laguitton. Development of screen management subroutines compatible with Microsoft FOR-TRAN compiler. The Computer Journal, 33(5):444–452, October 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_

journal/hdb/Volume_33/Issue_
05/tiff/444.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/445.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/446.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/447.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/448.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/449.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/450.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/451.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/452.tif.

Burns:1991:FBD

[BL91] A. Burns and A. M. Lister.

A framework for building dependable systems. The Computer Journal, 34(2):173–181, April 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/173.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/174.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/175.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        02/tiff/176.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        02/tiff/177.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        02/tiff/178.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        02/tiff/179.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        02/tiff/180.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        02/tiff/181.tif.

Blake:1982:FPS

[Bla82] A. Blake. Fixed point solutions of recursive operations on Boolean arrays. The Computer Journal, 25(2):231–234, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/231.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/232.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/233.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/234.tif.

Blair:1992:IRP

[Bla92]  D. C. Blair.  Information

retrieval and the philosophy of language. The Computer Journal, 35(3):200–207, June 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_35/Issue_03/
Vol35_03.body.html#AbstractBlair.

Bertossi:1992:SNP

[BLPL92] A. A. Bertossi, F. Luccio, L. Pagli, and E. Lodi. Short notes: A parallel solution to the approximate string matching problem. The Computer Journal, 35(5):524–526, October 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_35/Issue_05/
350524.sgm.abs.html;   http:
//www3.oup.co.uk/computer_
journal/Volume_35/Issue_05/
tiff/524.tif;   http://www3.
oup.co.uk/computer_journal/
Volume_35/Issue_05/tiff/525.
tif;   http://www3.oup.co.
uk/computer_journal/Volume_
35/Issue_05/tiff/526.tif.

Balter:1994:GL

[BLR94] Roland Balter, Serge Lacourte, and Michel Riveill. The Guide language. The Computer Journal, 37(6):519–530, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_37/Issue_06/
Vol37_06.body.html#AbstractBalter.

Blake:1970:RPP

[BLY70] L. F. Blake, R. E. Lawson, and I. M. Yuille. A ring processing package for use with FORTRAN or a similar highlevel language. The Computer Journal, 13(1):40–47, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/130040.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/tiff/40.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_01/tiff/
41.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/42.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
13/Issue_01/tiff/43.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/tiff/44.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_01/tiff/
45.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/46.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
13/Issue_01/tiff/47.tif.

Bernard:1959:GSC

[BM59] E. E. Bernard and P. D. A. Mole. Generating strategies for continuous separation processes. The Computer Journal, 2(2):87–89, July 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
02/020087.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
02/tiff/87.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_02/tiff/
88.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_02/Issue_02/tiff/89.
tif.

Brooker:1960:APP

[BM60] R. A. Brooker and D. Morris. An assembly program for a phrase structure language. The Computer Journal, 3(3):168–174, October 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/030168.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/168.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/169.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/170.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/171.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/172.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_

                                        03/tiff/173.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        03/tiff/174.tif.

Brooker:1961:SPR

[BM61] R. A. Brooker and D. Morris. Some proposals for the realization of a certain assembly program. The Computer Journal, 3(4):220–231, January 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        04/030220.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        04/tiff/220.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        04/tiff/221.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        04/tiff/222.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        04/tiff/223.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        04/tiff/224.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        04/tiff/225.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        04/tiff/226.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        04/tiff/227.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        04/tiff/228.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/229.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/230.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/231.tif.

Barrett:1963:EAP

[BM63] W. Barrett and A. J. Mitchell.

An extended autocode for Pegasus. The Computer Journal, 6(3):237–240, October 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/060237.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/237.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/238.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/239.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/240.tif.

Bratko:1980:APC

[BM80] I. Bratko and D. Michie. An advice program for a complex chess programming task. The Computer Journal, 23 (4):353–359, November 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_23/Issue_
                                        04/tiff/353.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        04/tiff/354.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        04/tiff/355.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        04/tiff/356.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        04/tiff/357.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        04/tiff/358.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        04/tiff/359.tif.

Bell:1981:CES

[BM81] D. Bell and Donald Michie. Correspondence: Expert systems. The Computer Journal, 24(2):191, May 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        02/tiff/191.tif. See [Mic80].

Bunt:1984:MLB

[BM84] Richard B. Bunt and Jennifer M. Murphy. The measurement of locality and the behaviour of programs. The Computer Journal, 27 (3):238–245, August 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_27/Issue_
03/tiff/238.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/239.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/240.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/241.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/242.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/243.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/244.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/245.tif.

Barringer:1986:PSA

[BM86] H. Barringer and I. Mearns.

A proof system for Ada tasks. The Computer Journal, 29 (5):404–415, October 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/404.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/405.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/406.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/407.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        05/tiff/408.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        05/tiff/409.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        05/tiff/410.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        05/tiff/411.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        05/tiff/412.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        05/tiff/413.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        05/tiff/414.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        05/tiff/415.tif.

Benyon:1988:EAI

[BM88] D. Benyon and Dianne Murray. Experience with adaptive interfaces. The Computer Journal, 31(5):465–473, October 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        05/tiff/465.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        05/tiff/466.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        05/tiff/467.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_

05/tiff/468.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/469.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/470.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/471.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/472.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/473.tif.

Bell:1989:NDD

[BM89] T. Bell and A. Moffat. A note on the DMC data compression scheme. The Computer Journal, 32(1):16–20, February 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
01/tiff/16.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_01/tiff/
17.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_32/Issue_01/tiff/18.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
32/Issue_01/tiff/19.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
01/tiff/20.tif.

Baik:1990:TAT

[BM90a] K. H. Baik and L. L. Miller.

Topological approach for testing equivalence in heterogeneous relational databases. The Computer Journal, 33(1):2– 10, February 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/10.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
01/tiff/2.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_33/Issue_01/tiff/
3.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_33/Issue_01/tiff/4.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/5.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
01/tiff/6.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_33/Issue_01/tiff/
7.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_33/Issue_01/tiff/8.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/9.tif.

Bailey:1990:PDC

[BM90b] R. L. Bailey and R. Mukkalama. Pipelining data compression algorithms. The Computer Journal, 33(4):308–313, August 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/308.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_

04/tiff/309.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/310.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/311.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/312.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/313.tif.

Bell:1990:TQO

[BM90c] S. B. M. Bell and D. C. Mason. Tesseral quaternions for the octtree. The Computer Journal, 33(5):386–397, October 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/386.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/387.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/388.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/389.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/390.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/391.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/392.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_33/Issue_
                                        05/tiff/393.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        05/tiff/394.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        05/tiff/395.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        05/tiff/396.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        05/tiff/397.tif.

Barroca:1992:FMU

[BM92] L. M. Barroca and J. A. Mc-Dermid. Formal methods: use and relevance for the development of safety-critical systems. The Computer Journal, 35(6):579–599, December 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_36/Issue_06/
                                        Vol35_06.body.html#AbstractBarroca.

See editorial [Coo92a].

Burdorf:1993:LBS

[BM93] C. Burdorf and J. Marti. Load balancing strategies for Time Warp on multi-user workstations. The Computer Journal, 36(2):168–176, April 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_36/Issue_02/
                                        Vol36_02.body.html#AbstractBurdorf.

Belkhatir:1994:SSD

[BM94a] Noureddine Belkhatir and Walcelio L. Melo. Supporting software development processes in Adele 2. The Computer Journal, 37(7):621–628, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_37/Issue_07/
Vol37_07.body.html#AbstractBelkhatir.

Bundy:1994:NSC

[BM94b] Alan Bundy and Hector Mac-Queen. The new software copyright law. The Computer Journal, 37(2):79–82, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_37/Issue_02/
Vol37_02.body.html#AbstractBundy.

Burns:1992:MSS

[BMD92] A. Burns, J. McDermid, and J. Dobson. On the meaning of safety and security. The Computer Journal, 35(1):3–15, February 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_35/Issue_01/
Vol35_01.body.html#AbstractBurns.

Bassiouni:1994:DIE

[BML94] Mostafa A. Bassiouni, Amar Mukherjee, and Mark J. Llewellyn. Design and implementation of extended Boolean and comparison operators for time-oriented query languages. The Computer Journal, 37 (7):576–587, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_37/Issue_07/
Vol37_07.body.html#AbstractBassiouni.

Blair:1987:KOS

[BMNS87] G. S. Blair, J. A. Mariani, J. R. Nicol, and D. Shepherd. A knowledge-based operating system. The Computer Journal, 30(3):193–200, June 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/193.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/194.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/195.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/196.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/197.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/198.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/199.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_

03/tiff/200.tif.

Barrow:1971:TED

[BMPS71] H. G. Barrow, D. Michie, R. J.

Popplestone, and S. H. Salter. Tokyo-Edinburgh dialogue on robots in artificial intelligence research. The Computer Journal, 14(1):91–95, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/140091.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/tiff/91.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_01/tiff/
92.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_14/Issue_01/tiff/93.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
14/Issue_01/tiff/94.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/tiff/95.tif.

Brooker:1962:TR

[BMR62] R. A. Brooker, D. Morris, and J. S. Rohl. Trees and routines. The Computer Journal, 5(1):33–47, April 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
01/050033.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
01/tiff/33.tif;   http://

www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_01/tiff/
34.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_05/Issue_01/tiff/35.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
05/Issue_01/tiff/36.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
01/tiff/37.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_01/tiff/
38.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_05/Issue_01/tiff/39.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
05/Issue_01/tiff/40.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
01/tiff/41.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_01/tiff/
42.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_05/Issue_01/tiff/43.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
05/Issue_01/tiff/44.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
01/tiff/45.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_01/tiff/
46.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_05/Issue_01/tiff/47.
tif.

Brooker:1967:CCF

[BMR67a] R. A. Brooker, D. Morris, and J. S. Rohl. Compiler Compiler facilities in Atlas Autocode. The Computer Journal, 9(4):350–352, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/090350.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/350.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/351.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/352.tif.

Brooker:1967:ECC

[BMR67b] R. A. Brooker, D. Morris, and J. S. Rohl. Experience with the compiler compiler. The Computer Journal, 9(4):345–349, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/090345.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/345.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/346.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/347.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/348.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_09/Issue_
                                        04/tiff/349.tif.

Burton:1990:AUS

[BMRS90] F. W. Burton, G. P. McKeown, and V. J. Rayward-Smith. Applications of UET scheduling theory to the implementation of declarative languages. The Computer Journal, 33(4):330–336, August 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        04/tiff/330.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        04/tiff/331.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        04/tiff/332.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        04/tiff/333.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        04/tiff/334.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        04/tiff/335.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        04/tiff/336.tif.

Bell:1988:CRT

[BMSA88] D. A. Bell, F. J. McErlean, P. M. Stewart, and W. Arbuckle. Clustering related tuples in databases. The Computer Journal, 31(3):253–257, June 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/253.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/254.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/255.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/256.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/257.tif.

Bonanno:1985:AAS

[BMT85] R. Bonanno, D. Maio, and P. Tiberio. An approximation algorithm for secondary index selection in relational database physical design. The Computer Journal, 28(4):398–405, August 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/398.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/399.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/400.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/401.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/402.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_

                                        04/tiff/403.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        04/tiff/404.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        04/tiff/405.tif.

Bozanis:1995:PWF

[BMT95] Panayiotis Bozanis, Christos Makris, and Athanasios Tsakalidis. Parametric weighted filter: an efficient dynamic manipulation of signature files. The Computer Journal, 38(6):478–488, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_38/Issue_06/
                                        Vol38_06.body.html#AbstractBozanis.

Boari:1976:NDR

[BN76] M. Boari and G. Neri. A note on deadlock in the ‘readers and writers’ problem. The Computer Journal, 19(3):275, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/275.tif.

Borovits:1978:IPC

[BN78] I. Borovits and S. Neumann. Internal pricing for computer services. The Computer Journal, 21(3):199–204, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_

journal/hdb/Volume_21/Issue_
03/tiff/199.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/200.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/201.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/202.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/203.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/204.tif.

Bohte:1968:NSI

[Boh68] Z. Bohte. Numerical solution of the inverse algebraic eigenvalue problem. The Computer Journal, 10(4):385–388, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/100385.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/385.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/386.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/387.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/388.tif.

Bolognesi:1997:EFC

[Bol97] T. Bolognesi. Expressive flexibility of a constraint-oriented notation. The Computer Journal, 40(5):259–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www.oup.co.uk/computer_
                                        journal/Volume_40/Issue_05/
                                        Vol40_05.body.html#AbstractBolognesi;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_40/Issue_05/
                                        Vol40_05.body.html#AbstractBolognesi.

Bonney:1969:SCC

[Bon69] M. C. Bonney. Some considerations of the cost and value of information. The Computer Journal, 12(2):118–123, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        02/120118.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        02/tiff/118.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        02/tiff/119.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        02/tiff/120.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        02/tiff/121.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        02/tiff/122.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        02/tiff/123.tif.

Bonney:1971:KIF

[Bon71] R. W. Bonney. A keyword index file of British book titles. The Computer Journal, 14(4):333–337, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/140333.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/333.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/334.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/335.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/336.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/337.tif.

Boothroyd:1967:AS

[Boo67a]  J. Boothroyd.  Algorithm

18: SUMFAC. The Computer Journal, 9(4):416, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/416.tif.

Boothroyd:1967:APR

[Boo67b]  J. Boothroyd.  Algorithm

20: Permutations of the rows or columns of a matrix. The Computer Journal, 10(2):206–207, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/206.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/207.tif.

Boothroyd:1967:ASPa

[Boo67c] J. Boothroyd. Algorithm 22: Shortest path between start node and end node of a network. The Computer Journal, 10(3):306–307, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/306.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/307.tif.

Boothroyd:1967:ASPb

[Boo67d] J. Boothroyd. Algorithm 23: Shortest path between start node and all other nodes of a network. The Computer Journal, 10(3):307–308, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/307.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/308.tif.

Boothroyd:1967:ALN

[Boo67e]  J. Boothroyd.  Algorithm

24: The list of nodes on the shortest path from start node to end node of a network. The Computer Journal, 10(3):308, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/308.tif. See note

[Ham69d].

Boothroyd:1967:ASS

[Boo67f] J. Boothroyd. Algorithm 25:

Sort a section of the elements of an array by determining the rank of each element. The Computer Journal, 10(3):308–309, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/308.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/309.tif. See note

[Hil69, Woo70g].

Boothroyd:1967:AOS

[Boo67g]  J. Boothroyd.  Algorithm

26: Order the subscripts of an array section according to the magnitudes of the elements. The Computer Journal, 10(3):309–310, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        03/tiff/309.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        03/tiff/310.tif. See note

[Hil69, Woo70g].

Boothroyd:1967:ARE

[Boo67h]  J. Boothroyd.  Algorithm

27: Rearrange the elements of an array section according to a permutation of the subscripts. The Computer Journal, 10(3):310, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        03/tiff/310.tif.

Boothroyd:1967:APEa

[Boo67i]  J. Boothroyd.  Algorithm

28: Permutations of the elements of a vector in lexicographic order. The Computer Journal, 10(3):311, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        03/tiff/311.tif.

Boothroyd:1967:APEb

[Boo67j] J. Boothroyd. Algorithm 29: Permutation of the elements of a vector. The Computer Journal, 10(3):311, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/311.tif.

Boothroyd:1967:AFP

[Boo67k]  J. Boothroyd.  Algorithm

30: Fast permutation of the elements of a vector. The Computer Journal, 10 (3):311–312, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/311.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/312.tif.

Boothroyd:1967:AANa

[Boo67l]  J. Boothroyd.  Algorithms:

Author’s note on algorithms 22, 23, 24. The Computer Journal, 10(3):306, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/306.tif.

Boothroyd:1967:AANb

[Boo67m]  J. Boothroyd.  Algorithms:

Author’s note on algorithms 25, 26, 27. The Computer Journal, 10(3):308, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/308.tif.

Boothroyd:1967:AANc

[Boo67n]  J. Boothroyd.  Algorithms:

Author’s note on algorithms 28, 29, 30. The Computer Journal, 10(3):310, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        03/tiff/310.tif.

Boothroyd:1968:ACF

[Boo68a]  J. Boothroyd.  Algorithm

31: Complex Fourier series. The Computer Journal, 10 (4):414–416, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/tiff/414.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/tiff/415.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/tiff/416.tif.

Boothroyd:1968:SNA

[Boo68b] J. Boothroyd. Special note on Algorithm 31. complex Fourier analysis. The Computer Journal, 11(1):115, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        01/tiff/115.tif.

Boothroyd:1974:CRC

[Boo74a]  J. Boothroyd.  Correspon-

dence: On a radix conversion process. The Computer Journal, 17(1):95, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
01/tiff/95.tif. See erratum

[Boo74b] and correspondence [Ham74].

Boothroyd:1974:ERC

[Boo74b]  J. Boothroyd.  Erratum:

“On a radix conversion process”. The Computer Journal, 17(4):381, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/381.tif. See [Boo74a,

Ham74].

Boothroyd:1978:CAM

[Boo78] J. Boothroyd. Certification of Algorithm 1: Mintree. The Computer Journal, 21(3): 277, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
21/Issue_03/tiff/277.tif.

See [Obr78a].

Boothroyd:1979:AP

[Boo79a]  J. Boothroyd.  Algorithm

6: PERM. The Computer Journal, 22(1):88–89, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
01/tiff/88.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_01/tiff/

89.tif. Reprinted from The Computer Bulletin, September 1964, p. 67.

Boothroyd:1979:AM

[Boo79b] J. Boothroyd. Algorithm 7: MINX. The Computer Journal, 22(1):89, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
01/tiff/89.tif. Reprinted

from The Computer Bulletin, September 1964, p. 67.

Boothroyd:1979:CAF

[Boo79c] J. Boothroyd. Certification of Algorithm 2: Fibonacci search. The Computer Journal, 22(1):86, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
01/tiff/86.tif. Reprinted

from The Computer Bulletin. See [PP79].

Boothroyd:1979:CAE

[Boo79d] J. Boothroyd. Certification of Algorithm 3: Easter. The Computer Journal, 22(1):87, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
01/tiff/87.tif. See [Hil79c].

Bowring:1973:DCU

[Bow73] C. A. Bowring. Discussion and correspondence: using macros to aid assembly-language teaching. The Computer Journal, 16(3):281–283, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/160281.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/281.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/282.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/283.tif.

Bowyer:1981:CDT

[Bow81] A. Bowyer. Computing Dirichlet tessellations. The Computer Journal, 24(2):162–166, May 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/162.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/163.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_24/Issue_
                                        02/tiff/164.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        02/tiff/165.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        02/tiff/166.tif.

Bowers:1984:DAA

[Bow84] David S. Bowers. A database architecture for aggregateincomplete data. The Computer Journal, 27(4):294–300, November 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        04/tiff/294.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        04/tiff/295.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        04/tiff/296.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        04/tiff/297.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        04/tiff/298.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        04/tiff/299.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        04/tiff/300.tif.

Bowden:1990:KMT

[Bow90] K. Bowden. Kron’s method of tearing on a Transputer array. The Computer Journal, 33(5):453–459, October 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/453.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/454.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/455.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/456.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/457.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/458.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/459.tif.

Box:1965:NMC

[Box65] M. J. Box. A new method of constrained optimization and a comparison with other methods. The Computer Journal, 8(1):42–52, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
01/080042.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
01/tiff/42.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_01/tiff/
43.tif;   http://www3.oup.
co.uk/computer_journal/hdb/

                                        Volume_08/Issue_01/tiff/44.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        08/Issue_01/tiff/45.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        01/tiff/46.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_08/Issue_01/tiff/
                                        47.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_08/Issue_01/tiff/48.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        08/Issue_01/tiff/49.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        01/tiff/50.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_08/Issue_01/tiff/
                                        51.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_08/Issue_01/tiff/52.
                                        tif.

Box:1966:CSC

[Box66a] M. J. Box. A comparison of several current optimization methods, and the use of transformations in constrained problems. The Computer Journal, 9(1):67–77, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        01/090067.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        01/tiff/67.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_09/Issue_01/tiff/
                                        68.tif;   http://www3.oup.

co.uk/computer_journal/hdb/
Volume_09/Issue_01/tiff/69.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
09/Issue_01/tiff/70.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/71.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_01/tiff/
72.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_09/Issue_01/tiff/73.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
09/Issue_01/tiff/74.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/75.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_01/tiff/
76.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_09/Issue_01/tiff/77.
tif.

Box:1966:CNP

[Box66b] M. J. Box. Correspondence:

On “Nonlinear programming test problems”. The Computer Journal, 9(1):115–116, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/115.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/116.tif. See [Myl66].

Boyd:1960:SEL

[Boy60] K. T. Boyd. Simultaneous equations and linear programming. The Computer Journal, 3(1):45–46, April 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        01/030045.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        01/tiff/45.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_03/Issue_01/tiff/
                                        46.tif.

Boyell:1963:MSG

[Boy63] R. L. Boyell. The method of successive grids for reduction of function storage requirements. The Computer Journal, 5(4):320–321, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        04/050320.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        04/tiff/320.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        04/tiff/321.tif.

Byrne:1969:IGI

[BP69] J. L. Byrne and L. G. Proll. Initialising Geoffrion’s implicit enumeration algorithm for the zero-one linear programming problem. The Computer Journal, 12(4):381–384, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/120381.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/381.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/382.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/383.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/384.tif.

Bozyigit:1979:FRP

[BP79] M. Bozyigit and Y. Paker.

A fixed routing problem in large and high connectivity networks. The Computer Journal, 22(3):246–250, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/246.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/247.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/248.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/249.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/250.tif.

Butler:1981:FDI

[BP81] G. F. Butler and J. Pike. Freeformat data input scheme written in Standard FORTRAN (ANSI 66). The Computer Journal, 24(3):256–257, August 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        03/tiff/256.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        03/tiff/257.tif.

Bozyigit:1982:TRM

[BP82] M. Bozyigit and Y. Paker. A topology reconfiguration mechanism for distributed computer systems. The Computer Journal, 25(1):87–92, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/87.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_25/Issue_01/tiff/
                                        88.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_25/Issue_01/tiff/89.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        25/Issue_01/tiff/90.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/91.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_25/Issue_01/tiff/
                                        92.tif.

Barsi:1992:AFH

[BP92] F. Barsi and M. C. Pinotti. Adding flexibility to hybrid number systems. The Computer Journal, 35(6):630–635, December 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_06/
Vol35_06.body.html#AbstractBarsi.

Barsi:1994:TOM

[BP94] Ferruccio Barsi and M. Cristina Pinotti. Time optimal mixed radix conversion for residue number applications. The Computer Journal, 37(10):907–916, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_37/Issue_10/
Vol37_10.body.html#AbstractBarsi.

Basu:1987:OPL

[BPG87] Julie Basu, L. M. Patnaik, and A. K. Goswami. Ordered Ports — a language concept for high-level distributed programming. The Computer Journal, 30(6):487–497, December 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/487.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/488.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/489.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/490.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/491.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/492.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/493.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/494.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/495.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/496.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/497.tif.

Boari:1974:PEP

[BPR74] M. Boari, P. Pellizzardi, and R. Rossi. Performance evaluation of a process control system by simulation model. The Computer Journal, 17 (4):365–370, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/365.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/366.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/367.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/368.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/369.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/370.tif.

Brown:1972:DCD

[BPSS72] A. E. Brown, C. E. Phillips, J. S. Scandale, and D. P. Sparrow. A data collection and display system for a largescale simulation. The Computer Journal, 15(2):105–108, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/150105.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/105.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/106.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/107.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/108.tif. See [Tem72].

Boiten:1992:HPC

[BPTV92] E. A. Boiten, H. A. Partsch, D. Tuijnman, and N. Volker. How to produce correct software — an introduction to formal specification and program development by transformations. The Computer Journal, 35(6):547–554, December 1992. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_06/
Vol35_06.body.html#AbstractBoiten.

Boardman:1986:TPM

[BR86] J. T. Boardman and G. C. A. Robson. Towards a problemsolving methodology for coping with increasing complexity: an engineering approach. The Computer Journal, 29(2):161–166, April 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/161.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/162.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/163.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/164.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/165.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/166.tif.

Berghel:1990:PSM

[BR90] H. Berghel and R. Rankin. A proposed standard for measuring crossword compilation efficiency. The Computer Journal, 33(2):181–184, April 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/181.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/182.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/183.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/184.tif.

Bodorik:1991:TVP

[BR91] P. Bodorik and J. S. Riordon. Threshold values for processing distributed queries. The Computer Journal, 34 (6):551–558, December 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/551.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/552.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/553.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/554.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/555.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/556.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/557.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        06/tiff/558.tif.

Brace:1963:DCE

[Bra63] D. A. Brace. Direct coding of English language names. The Computer Journal, 6(2): 113–117, July 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        06/Issue_02/060113.sgm.abs.
                                        html;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        06/Issue_02/tiff/113.tif;
                                         http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/tiff/114.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/tiff/115.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/tiff/116.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/tiff/117.tif. See corre-

spondence [Ker63].

Bradly:1972:CHL

[Bra72] H. W. Bradly. Correspondence: On high level languages. The Computer Journal, 15(3):258, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/258.tif. See [Pal72].

Brandon:1974:IUS

[Bra74] D. M. Brandon, Jr. The implementation and use of sparse matrix techniques in general simulation programs. The Computer Journal, 17(2):165–171, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/165.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/166.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/167.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/168.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/169.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/170.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/171.tif.

Bradley:1975:CAS

[Bra75] John Bradley. Correspondence:

On “An algorithm for the solution of constrained generalised polynomial programming problems”. The Computer Journal, 18(3):285–286, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_

                                        03/tiff/285.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/286.tif. See [BJS74].

Brady:1976:PAS

[Bra76] J. M. Brady. A programming approach to some concepts and results in the theory of computation. The Computer Journal, 19(3):234–237, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/234.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/235.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/236.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/237.tif.

Brady:1977:HPR

[Bra77]  J. M. Brady.  Hints on

proofs by recursion induction. The Computer Journal, 20(4):353–355, November 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        04/tiff/353.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        04/tiff/354.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        04/tiff/355.tif.

Brady:1978:BRS

[Bra78] J. M. Brady. Book review:

Structural Pattern Recognition, by T. Pavlidis, 1977. The Computer Journal, 21(4):384, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
21/Issue_04/tiff/384.tif.

Bramer:1980:COS

[Bra80] M. A. Bramer. Correct and optimal strategies in game playing programs. The Computer Journal, 23(4):347–352, November 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/347.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/348.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/349.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/350.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/351.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/352.tif.

Bradley:1985:UMD

[Bra85] James Bradley. Use of mean distance between overflow records to compute average search lengths in hash files with open addressing. The Computer Journal, ??(??):??, August 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Bradley:1986:UMD

[Bra86a] J. Bradley. Use of mean distance between overflow records to compute average search lengths in hash files with open addressing. The Computer Journal, 29(2):167–170, April 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/167.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/168.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/169.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/170.tif.

Bradley:1986:JDR

[Bra86b] James Bradley. Join dependencies in relational databases and the geometry of spatial grids. The Computer Journal, 29(4):378–380, August 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/378.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_

04/tiff/379.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/380.tif.

Bradley:1988:PJD

[Bra88] J. Bradley. Polygon join dependencies, closed co-relationship chains and the connection trap in relational databases. The Computer Journal, 31(2):141– 146, April 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
31/Issue_02/tiff/141.tif;
 http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/142.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/143.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/144.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/145.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/146.tif.

Bradley:1994:CSS

[Bra94] James Bradley. A complete Lsystem specification for generating an exact self-affine growthdecay function with a random walk scaling property. The Computer Journal, 37(1):58–66, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

                                        journal/Volume_37/Issue_01/
                                        Vol37_01.body.html#AbstractBradley.

Brooker:1966:MFA

[BRC66] R. A. Brooker, J. S. Rohl, and S. R. Clark. The main features of Atlas Autocode. The Computer Journal, 8(4):303–310, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        04/080303.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        04/tiff/303.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        04/tiff/304.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        04/tiff/305.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        04/tiff/306.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        04/tiff/307.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        04/tiff/308.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        04/tiff/309.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        04/tiff/310.tif. See com-

ment [CL66].

Brent:1971:AGC

[Bre71] R. P. Brent. An algorithm with guaranteed convergence for finding a zero of a function. The Computer Journal, 14(4):422–425, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/140422.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/422.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/423.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/424.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/425.tif.

Brent:1981:CCP

[Bre81]  R. P. Brent.  Correspon-

dence: Comments on papers by Maddison and Westreich. The Computer Journal, 24(1):95–96, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
01/tiff/95.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_01/tiff/

96.tif. See [Mad80, Wes80b].

Bresenham:1982:ILC

[Bre82]  Jack E. Bresenham.  In-

cremental line compaction. The Computer Journal, 25 (1):116–120, February 1982.

CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/116.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/117.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/118.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/119.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/120.tif.

Breene:1993:QHG

[Bre93]  L. A. Breene.  Quadtrees

and hypercubes: grid embedding strategies based on spatial data structure addressing. The Computer Journal, 36(6):562–569, December 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_36/Issue_06/
                                        Vol36_06.body.html#AbstractBreene.

Barrodale:1970:CBA

[BRH70] I. Barrodale, F. D. K. Roberts, and C. R. Hunt. Computing best ℓ_{p} approximations by functions nonlinear in one parameter. The Computer Journal, 13(4):382–386, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_

journal/hdb/Volume_13/Issue_
04/130382.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/382.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/383.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/384.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/385.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/386.tif.

Brittan:1967:NSM

[Bri67] J. N. G. Brittan. The nature and structure of man-computer communication in team control systems. The Computer Journal, 10(3):219–226, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/100219.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/219.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/220.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/221.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/222.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_

                                        03/tiff/223.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        03/tiff/224.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        03/tiff/225.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        03/tiff/226.tif.

Briggs:1979:RRT

[Bri79] T. Briggs. A review of road traffic simulation methods using a general purpose digital computer. The Computer Journal, 22(1):11–16, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        01/tiff/11.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_22/Issue_01/tiff/
                                        12.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_22/Issue_01/tiff/13.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        22/Issue_01/tiff/14.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        01/tiff/15.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_22/Issue_01/tiff/
                                        16.tif.

Brittain:1980:DCE

[Bri80]  J. N. G. Brittain.  De-

sign for a changing environment. The Computer Journal, 23(1):13–22, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
01/tiff/13.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/
14.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_23/Issue_01/tiff/15.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
23/Issue_01/tiff/16.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
01/tiff/17.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/
18.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_23/Issue_01/tiff/19.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
23/Issue_01/tiff/20.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
01/tiff/21.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/
22.tif.

Brinck:1985:EPT

[Bri85] Keith Brinck. The expected performance of traversal algorithms in binary trees. The Computer Journal, 28(4):426–432, August 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/426.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_

                                        04/tiff/427.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        04/tiff/428.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        04/tiff/429.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        04/tiff/430.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        04/tiff/431.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        04/tiff/432.tif.

Brooker:1958:APD

[Bro58a] R. A. Brooker. The autocode programs developed for the Manchester University Computers. The Computer Journal, 1(1):15–21, April 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        01/010015.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        01/tiff/15.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_01/Issue_01/tiff/
                                        16.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_01/Issue_01/tiff/17.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        01/Issue_01/tiff/18.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        01/tiff/19.tif;   http://
                                        www3.oup.co.uk/computer_journal/

hdb/Volume_01/Issue_01/tiff/
20.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_01/Issue_01/tiff/21.
tif.

Brooker:1958:FAF

[Bro58b]  R. A. Brooker.  Further

autocode facilities for the Manchester (mercury) computer. The Computer Journal, 1(3):124–127, October 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/010124.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/124.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/125.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/126.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/127.tif.

Brooker:1960:STD

[Bro60] R. A. Brooker. Some techniques for dealing with twolevel storage. The Computer Journal, 2(4):189–194, January 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/020189.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_

04/tiff/189.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/190.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/191.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/192.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/193.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/194.tif.

Brown:1969:STT

[Bro69a] A. R. Brown. Selling television time: an optimisation problem. The Computer Journal, 12(3):201–207, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/120201.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/201.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/202.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/203.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/204.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/205.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_12/Issue_
03/tiff/206.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/207.tif.

Brown:1969:UMP

[Bro69b] P. J. Brown. Using a macro processor to aid software implementation. The Computer Journal, 12(4):328–331, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/120328.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/328.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/329.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/330.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/331.tif.

Broyden:1969:ASN

[Bro69c] C. G. Broyden. Algorithm 44:

Solution of nonlinear simultaneous equations. The Computer Journal, 12(4):406–408, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/406.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/407.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/408.tif. See note

[Fie70, Red71].

Broyden:1969:NMS

[Bro69d] C. G. Broyden. A new method of solving nonlinear simultaneous equations. The Computer Journal, 12(1):94–99, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        01/120094.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        01/tiff/94.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_12/Issue_01/tiff/
                                        95.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_12/Issue_01/tiff/96.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        12/Issue_01/tiff/97.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        01/tiff/98.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_12/Issue_01/tiff/
                                        99.tif.

Brown:1973:DCP

[Bro73a] D. C. Brown. Discussion and correspondence: the project, and the future of Computing Science courses. The Computer Journal, 16(4):380–382, November 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_

journal/hdb/Volume_16/Issue_
04/160380.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/380.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/381.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/382.tif. See corre-

spondence [Fis74a, Bel74].

Brown:1973:PUO

[Bro73b] P. J. Brown. A practical usage of outer and inner syntax. The Computer Journal, 16(2):122–123, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/160122.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/122.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/123.tif.

Brown:1973:UPL

[Bro73c] P. J. Brown. UNRAVEL — a programming language to put intelligence into dumps. The Computer Journal, 16 (1):10–12, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
01/160010.sgm.abs.html;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
01/tiff/10.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_01/tiff/
11.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_16/Issue_01/tiff/12.
tif.

Broadbent:1974:MSA

[Bro74a] J. K. Broadbent. Microprogramming and system architecture. The Computer Journal, 17(1):2–8, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
01/tiff/2.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/
3.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_17/Issue_01/tiff/4.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
17/Issue_01/tiff/5.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
01/tiff/6.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/
7.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_17/Issue_01/tiff/8.

tif. See correspondence [WS74].

Broderick:1974:BRS

[Bro74b] W. R. Broderick. Book reviews: Spline Analysis, by Martin H. Schultz, 1972; Introduction to Programming, by A. Sharp; Algol Programming, by R. W. Sharp. The Computer Journal, 17(1):92, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
01/tiff/92.tif.

Brown:1974:BDE

[Bro74c] D. A. H. Brown. Biquinary decimal error detection codes with one, two and three check digits. The Computer Journal, 17(3):201–204, August 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/201.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/202.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/203.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/204.tif.

Brown:1974:CCR

[Bro74d]  D. A. H. Brown.  Cor-

respondence: Cyclic redundancy checking by program. The Computer Journal, 17 (1):94–95, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_

                                        01/tiff/94.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_17/Issue_01/tiff/
                                        95.tif.

Brown:1974:SEC

[Bro74e] D. A. H. Brown. Some error correcting codes for certain transposition and transcription errors in decimal integers. The Computer Journal, 17(1):9–12, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        01/tiff/10.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_17/Issue_01/tiff/
                                        11.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_17/Issue_01/tiff/12.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        17/Issue_01/tiff/9.tif.

Brown:1979:AAU

[Bro79a] A. Brown. Algorithm 106: An algorithm for updating an ordered list on a VDU. The Computer Journal, 22(2):188–189, May 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/188.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/189.tif.

Brown:1979:SMV

[Bro79b] P. J. Brown. Software methods for virtual storage of executable code. The Computer Journal, 22(1):50–52, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
01/tiff/50.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_01/tiff/
51.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_22/Issue_01/tiff/52.
tif.

Brooker:1982:SB

[Bro82] R. A. Brooker. A ‘database’ subsystem for BCPL. The Computer Journal, 25(4):448– 464, November 1982. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
25/Issue_04/tiff/448.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/449.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/450.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/451.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/452.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/453.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/454.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/455.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/456.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/457.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/458.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/459.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/460.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/461.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/462.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/463.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/464.tif.

Broyden:1983:MA

[Bro83] C. G. Broyden. A markscaling algorithm. The Computer Journal, 26(2):109–112, May 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/109.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/110.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/111.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/112.tif.

Brown:1984:TAI

[Bro84a]  D. H. J. Brown.  Transi-

tions to Ada: an incremental approach. The Computer Journal, 27(1):37–41, February 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
01/tiff/37.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_01/tiff/
38.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_27/Issue_01/tiff/39.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
27/Issue_01/tiff/40.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
01/tiff/41.tif.

Brown:1984:EMP

[Bro84b] P. J. Brown. Embedded macro processors. The Computer Journal, 27(4):348–353, November 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/348.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/349.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_27/Issue_
04/tiff/350.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/351.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/352.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/353.tif.

Brown:1989:SSD

[Bro89]  H. Brown.  Standards for

structured documents. The Computer Journal, 32(6):505– 514, December 1989. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
32/Issue_06/tiff/505.tif;
 http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/506.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/507.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/508.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/509.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/510.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/511.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/512.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_32/Issue_
06/tiff/513.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/514.tif.

Browne:1995:EMT

[Bro95] R. F. Browne. The embedding of meshes and trees into degree four chordal ring networks. The Computer Journal, 38(1):71–77, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_38/Issue_01/
Vol38_01.body.html#AbstractBrowne.

Berrington:1993:GUP

[BRP93] Neil Berrington, David De Roure, and Julian Padget. Guaranteeing unpredictability (programming model). The Computer Journal, 36(8):723– 733, ???? 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/Volume_
36/Issue_08/Vol36_08.body.
html#AbstractBerrington.

Brink:1991:AB

[BRS91] C. Brink, I. M. Rewitzky, and R. A. Schmidt. Autodescriptivity: beware! The Computer Journal, 34(4):380–381, August 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/380.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/381.tif.

Bruha:1992:AMS

[Bru92] I. Bruha. AI multilanguage system McPOPLOG: The power of communication between its subsystems. The Computer Journal, 35(6):625–629, December 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_06/
Vol35_06.body.html#AbstractBruha.

Bryen:1959:IES

[Bry59] J. F. A. Bryen. The introduction and establishment of a system of computer production control in a light engineering factory. The Computer Journal, 2(3):115–118, October 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/020115.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/115.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/116.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/117.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/118.tif.

Bennett:1961:ATC

[BS61] R. O. Bennett and J. B.

Stringer. Acceptance trials of computer systems for government use. The Computer Journal, 4(3):185–196, October 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/040185.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/185.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/186.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/187.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/188.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/189.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/190.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/191.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/192.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/193.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/194.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_04/Issue_
                                        03/tiff/195.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        03/tiff/196.tif.

Boggs:1971:NIO

[BS71] R. A. C. Boggs and F. J. Smith. A note on the integration of ordinary differential equations using Chebyshev series. The Computer Journal, 14(3):270–271, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/140270.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/tiff/270.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/tiff/271.tif.

Bordier:1973:NSC

[BS73a] J. Bordier and H. Saya. A necessary and sufficient condition for a power language to be LL(k). The Computer Journal, 16(4):351–356, November 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        04/160351.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        04/tiff/351.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        04/tiff/352.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/353.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/354.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/355.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/356.tif.

Burstall:1973:BRR

[BS73b] R. M. Burstall and P. A. Samet.

Book reviews: Representation and Meaning, by Herbert A. Simon and Laurent Siklossy, 1972; A Collection of Programming Problems and Techniques, by H. A. Maurer and M. R. Williams, 1972. The Computer Journal, 16(3):253, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/253.tif.

Bezdek:1975:NCM

[BS75] R. H. Bezdek and A. H. Sameh.

A note on computational methods for input-output econometric models. The Computer Journal, 18(3):248–249, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/248.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_18/Issue_
03/tiff/249.tif.

Burdekin:1979:CCE

[BS79] R. Burdekin and R. S. Scowen. Computers and the construction of economic models. The Computer Journal, 22(1):3– 10, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
22/Issue_01/tiff/10.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
01/tiff/3.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_22/Issue_01/tiff/
4.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_22/Issue_01/tiff/5.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
22/Issue_01/tiff/6.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
01/tiff/7.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_22/Issue_01/tiff/
8.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_22/Issue_01/tiff/9.
tif.

Buckley:1984:HGL

[BS84] G. N. Buckley and A. Silberschatz. On the heterogeneous guard locking protocol. The Computer Journal, 27(1):86–87, February 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
01/tiff/86.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_01/tiff/
87.tif.

Blair:1986:DTD

[BS86] A. Blair and I. Sommerville.

DSA — a tool for descriptive text analysis. The Computer Journal, 29(5):460–466, October 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/460.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/461.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/462.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/463.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/464.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/465.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/466.tif.

Benyon:1987:TTK

[BS87] D. Benyon and S. Skidmore.

Towards a tool kit for the systems analyst. The Computer Journal, 30(1):2–7, February 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        01/tiff/2.tif;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_30/Issue_01/tiff/
                                        3.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_30/Issue_01/tiff/4.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        30/Issue_01/tiff/5.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        01/tiff/6.tif;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_30/Issue_01/tiff/
                                        7.tif.

Bennett:1989:NRB

[BS89] J. P. Bennett and G. C. Smith. The need for reduced byte stream instruction sets. The Computer Journal, 32(4):370–373, August 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        04/tiff/370.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        04/tiff/371.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        04/tiff/372.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        04/tiff/373.tif.

Bornat:1999:AFP

[BS99] Richard Bornat and Bernard Sufrin. Animating formal proof at the surface: The Jape proof calculator. The Computer Journal, 42(3):177–192, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
03/420177.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
03/pdf/420177.pdf.

Barron:1960:SSL

[BSD60] D. W. Barron and H. P. F.

Swinnerton-Dyer. Solution of simultaneous linear equations using a magnetic-tape store. The Computer Journal, 3(1):28–33, April 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
01/030028.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
01/tiff/28.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_01/tiff/
29.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_03/Issue_01/tiff/30.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
03/Issue_01/tiff/31.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
01/tiff/32.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_01/tiff/
33.tif.

Buckley-Sharp:1970:CPB

[BSH70] M. D. Buckley-Sharp and F. T. C. Harris. A computer program for banking multiple choice questions. The Computer Journal, 13(3):230–236, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/130230.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/230.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/231.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/232.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/233.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/234.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/235.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/236.tif.

Bell:1990:IDD

[BSH90] D. A. Bell, J. Shao, and M. Elizabeth C. Hull. Integrated deductive database system implementation. A systematic study. The Computer Journal, 33(1):40–48, February 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
01/tiff/40.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_01/tiff/
41.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_33/Issue_01/tiff/42.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/43.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
01/tiff/44.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_01/tiff/
45.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_33/Issue_01/tiff/46.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/47.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
01/tiff/48.tif.

Boutel:1965:CIP

[BSPI65] B. E. Boutel, C. Strachey, J. H. G. Phillips, and Michael Irish. Correspondence: An impossible program. The Computer Journal, 8(3):215, October 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/215.tif.

Brace:1963:BRE

[BT63] D. A. Brace and G. J.

Tee. Book reviews: Elementary Mathematical Programming, by Robert W. Metzger, 1963; U.S.S.R. Computational Mathematics and Mathematical Physics, Number 1, 1962. The Computer Journal, 6(3):276–276, 286–286, October 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        03/tiff/276.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        03/tiff/286.tif.

Baldwin:1971:MAC

[BT71] J. C. Baldwin and R. E. Thomas. Multi-access on the Chilton Atlas. The Computer Journal, 14(2):119–122, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/140119.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/119.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/120.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/121.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/122.tif.

Barsky:1981:TSR

[BT81] Brian A. Barsky and Spencer W. Thomas. TRANSPLINE — a system for representing curves using transformations among four spline formulations. The Computer Journal, 24(3):271–277, August 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/271.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/272.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/273.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/274.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/275.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/276.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/277.tif.

Bojovic:1999:IDT

[BTD99] Miroslav Bojovic, Milo Tomasevic, and Jovan Djordjevic. The interactive development and testing system for a RISCstyle processor. The Computer Journal, 42(5):422–429, 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
05/420422.sgm.abs.html.

Buckingham:1960:OUC

[Buc60] R. A. Buckingham. The organization of a university computer centre. The Computer Journal, 3(3):131–135, October 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        03/030131.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        03/tiff/131.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        03/tiff/132.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        03/tiff/133.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        03/tiff/134.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        03/tiff/135.tif.

Buckingham:1965:CU

[Buc65] R. A. Buckingham. The computer in the university. The Computer Journal, 8(1):1–7, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        01/080001.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        01/tiff/1.tif;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_08/Issue_01/tiff/
                                        2.tif;   http://www3.oup.

co.uk/computer_journal/hdb/
Volume_08/Issue_01/tiff/3.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
08/Issue_01/tiff/4.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
01/tiff/5.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_08/Issue_01/tiff/
6.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_08/Issue_01/tiff/7.
tif.

Bull:1972:DDB

[Bul72] G. M. Bull. Dynamic debugging in BASIC. The Computer Journal, 15(1):21–24, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/150021.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/tiff/21.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_01/tiff/
22.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_15/Issue_01/tiff/23.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
15/Issue_01/tiff/24.tif. See

correspondence [Sil72].

Bunton:1997:SMI

[Bun97]  S. Bunton.  Semantically

motivated improvements for PPM variants. The Computer Journal, 40(2/3):76–??, ???? 1997. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www.oup.co.uk/computer_
                                        journal/Volume_40/Issue_02/
                                        Vol40_02.body.html#AbstractBunton;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_40/Issue_02/
                                        Vol40_02.body.html#AbstractBunton;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_40/Issue_02/
                                        Vol40_03.body.html#AbstractBunton.

Burstall:1966:CDE

[Bur66] R. M. Burstall. Computer design of electricity supply networks by a heuristic method. The Computer Journal, 9 (3):263–274, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        03/090263.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        03/tiff/263.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        03/tiff/264.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        03/tiff/265.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        03/tiff/266.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        03/tiff/267.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        03/tiff/268.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_09/Issue_
03/tiff/269.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/270.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/271.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/272.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/273.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/274.tif.

Burley:1967:PLL

[Bur67] H. T. Burley. A programming language for linear algebra. The Computer Journal, 10(1):69–73, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/100069.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/69.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_01/tiff/
70.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_10/Issue_01/tiff/71.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/72.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/73.tif.

Burstall:1969:PSW

[Bur69a] R. M. Burstall. A program for solving word sum puzzles. The Computer Journal, 12(1):48–51, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        01/120048.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        01/tiff/48.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_12/Issue_01/tiff/
                                        49.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_12/Issue_01/tiff/50.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        12/Issue_01/tiff/51.tif.

Burstall:1969:PPP

[Bur69b] Rod M. Burstall. Proving properties of programs by structural induction. The Computer Journal, 12(1):41–48, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        01/120041.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        01/tiff/41.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_12/Issue_01/tiff/
                                        42.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_12/Issue_01/tiff/43.
                                        tif;   http://www3.oup.co.

uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/44.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
01/tiff/45.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_01/tiff/
46.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_12/Issue_01/tiff/47.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/48.tif.

Burgess:1972:CTE

[Bur72] C. J. Burgess. Compile-time error diagnostics in syntaxdirected compilers. The Computer Journal, 15(4):302–307, November 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
15/Issue_04/150302.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
15/Issue_04/tiff/302.tif;
 http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/303.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/304.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/305.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/306.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/307.tif. See corre-

spondence [Mux73].

Burkhard:1975:FTQ

[Bur75a] W. A. Burkhard. Full table quadratic quotient searching. The Computer Journal, 18(2):161–163, May 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        02/tiff/161.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        02/tiff/162.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        02/tiff/163.tif.

Burkhard:1975:NTT

[Bur75b]  W. A. Burkhard.  Nonre-

cursive traversals of trees. The Computer Journal, 18 (3):227–230, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/227.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/228.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/229.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/230.tif.

Burton:1979:FPS

[Bur79] W. Burton. A FORTRAN preprocessor to support encapsulated data abstraction definitions. The Computer Journal, 22(4):307–312, November 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/307.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/308.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/309.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/310.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/311.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/312.tif.

Burton:1980:BRA

[Bur80]  C. T. Burton.  Book re-

view: Artificial Intelligence, Volume 1 by P. H. Winston and R. H. Brown. The Computer Journal, 23(2):106, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/106.tif.

Burkimsher:1987:CRS

[Bur87a] P. C. Burkimsher. Combinator reduction in a shared-memory multiprocessor. The Computer Journal, 30(3):214–222, June 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/214.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/215.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/216.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/217.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/218.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/219.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/220.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/221.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/222.tif.

Burton:1987:FPC

[Bur87b] F. W. Burton. Functional programming for concurrent and distributed computing. The Computer Journal, 30 (5):437–450, October 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/437.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/438.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/439.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/440.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/441.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/442.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/443.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/444.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/445.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/446.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/447.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/448.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/449.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/450.tif.

Burton:1988:NRT

[Bur88]  F. W. Burton.  Nondeter-

minism with referential transparency in functional programming languages. The Computer Journal, 31(3):243–247, June 1988. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        03/tiff/243.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        03/tiff/244.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        03/tiff/245.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        03/tiff/246.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        03/tiff/247.tif.

Butcher:1961:RSN

[But61] J. C. Butcher. Random sampling from the normal distribution. The Computer Journal, 3(4):251–253, January 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        04/030251.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        04/tiff/251.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        04/tiff/252.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        04/tiff/253.tif.

Buxton:1966:WSC

[Bux66] J. N. Buxton. Writing simulations in CSL. The Computer Journal, 9(2):137–143, August 1966. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/090137.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/137.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/138.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/139.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/140.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/141.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/142.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/143.tif.

Bruza:1992:SHS

[BV92] P. D. Bruza and T. P. Van Der Weide. Stratified hypermedia structures for information disclosure. The Computer Journal, 35(3):208–220, June 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_35/Issue_03/
Vol35_03.body.html#AbstractBruza.

Bromley:1974:CTL

[BVH74] A. G. Bromley, Z. G. Vranesic, and V. C. Hamacher. Correspondence: Ternary logic in parallel multipliers. The Computer Journal, 17(2):189–190, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        02/tiff/189.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        02/tiff/190.tif. See [VH72].

Beasley:1969:DTS

[BW69a] J. D. Beasley and K. Wilson. Design and testing of the System 4 random number generator. The Computer Journal, 12(4):368–372, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/120368.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/368.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/369.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/370.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/371.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/372.tif.

Bellamy:1969:MPO

[BW69b] N. W. Bellamy and M. J. West. Methods of profile optimisation by iterative analogue computation. The Computer Journal, 12(2):132–138, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/120132.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/132.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/133.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/134.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/135.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/136.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/137.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/138.tif.

Bernal:1970:NTB

[BW70a] M. J. M. Bernal and J. R.

Whiteman. Numerical treatment of biharmonic boundary value problems with re-entrant boundaries. The Computer Journal, 13(1):87–91, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/130087.sgm.abs.html;
http://www3.oup.co.uk/computer_

                                        journal/hdb/Volume_13/Issue_
                                        01/tiff/87.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_13/Issue_01/tiff/
                                        88.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_13/Issue_01/tiff/89.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        13/Issue_01/tiff/90.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        01/tiff/91.tif.

Boulton:1970:PNC

[BW70b] D. M. Boulton and C. S. Wallace. A program for numerical classification. The Computer Journal, 13(1):63–69, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        01/130063.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        01/tiff/63.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_13/Issue_01/tiff/
                                        64.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_13/Issue_01/tiff/65.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        13/Issue_01/tiff/66.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        01/tiff/67.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_13/Issue_01/tiff/
                                        68.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/

Volume_13/Issue_01/tiff/69.
tif.

Beaubien:1971:IFD

[BW71] M. J. Beaubien and A. Wexler.

Iterative, finite difference solution of interior eigenvalues and eigenfunctions of Laplace’s operator. The Computer Journal, 14(3):263–269, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/140263.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/263.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/264.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/265.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/266.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/267.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/268.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/269.tif.

Boulton:1973:IMH

[BW73a] D. M. Boulton and C. S. Wallace. An information measure for hierarchic classification. The Computer Journal, 16(3):254–261, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/160254.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/254.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/255.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/256.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/257.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/258.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/259.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/260.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/261.tif.

Boulton:1973:ORA

[BW73b] D. M. Boulton and C. S. Wallace. Occupancy of a rectangular array. The Computer Journal, 16(1):57–63, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
01/160057.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_

01/tiff/57.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_01/tiff/
58.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_16/Issue_01/tiff/59.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
16/Issue_01/tiff/60.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
01/tiff/61.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_01/tiff/
62.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_16/Issue_01/tiff/63.
tif.

Boulton:1975:IMS

[BW75a] D. M. Boulton and C. S. Wallace. An information measure for single link classification. The Computer Journal, 18(3):236–238, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/236.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/237.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/238.tif.

Buckholtz:1975:APR

[BW75b] T. J. Buckholtz and C. S.

Wetherell. Algorithm 89: A program to referee Kriegspiel and chess. The Computer Journal, 18(2):177–183, May 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/177.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/178.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/179.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/180.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/181.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/182.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/183.tif.

Bochmann:1978:CWS

[BW78] Gregor V. Bochmann and P. Ward. Compiler writing system for attribute grammars. The Computer Journal, 21(2):144–148, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/144.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/145.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/146.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_21/Issue_
02/tiff/147.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/148.tif. See also:

publication 199, Dépt. d’Informatique, University de Montréal (July 1975).

Barnard:1988:AMC

[BWM88] P. Barnard, M. Wilson, and A. Maclean. Approximate modelling of cognitive activity with an expert system: a theory-based strategy for developing an interactive design tool. The Computer Journal, 31(5):445–456, October 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/445.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/446.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/447.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/448.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/449.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/450.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/451.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_31/Issue_
                                        05/tiff/452.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        05/tiff/453.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        05/tiff/454.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        05/tiff/455.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        05/tiff/456.tif.

Bowyer:1981:MAS

[BWW81] Adrian Bowyer, Philip J. Willis, and John R. Woodwark. A multiprocessor architecture for solving spatial problems. The Computer Journal, 24(4):353–357, November 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        04/tiff/353.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        04/tiff/354.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        04/tiff/355.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        04/tiff/356.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        04/tiff/357.tif.

Barton:1971:ASS

[BWZ71] David Barton, I. M. Willers, and R. V. M. Zahar. The automatic solution of systems of ordinary differential equations by the method of Taylor series. The Computer Journal, 14(3):243–248, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/140243.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/243.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/244.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/245.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/246.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/247.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/248.tif.

Barrodale:1966:NNP

[BY66] I. Barrodale and A. Young.

A note on numerical procedures for approximation by spline functions. The Computer Journal, 9(3):318–320, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/090318.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_

                                        03/tiff/318.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        03/tiff/319.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        03/tiff/320.tif.

Ben-Yaacov:1979:ICD

[BY79]  G. Z. Ben-Yaacov.  Inter-

active computation and data management for power system studies. The Computer Journal, 22(1):76–79, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        01/tiff/76.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_22/Issue_01/tiff/
                                        77.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_22/Issue_01/tiff/78.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        22/Issue_01/tiff/79.tif.

Berghel:1989:CC

[BY89] H. Berghel and C. Yi. Crossword compiler-compilation. The Computer Journal, 32(3):276– 280, June 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        32/Issue_03/tiff/276.tif;
                                         http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        03/tiff/277.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_

03/tiff/278.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/279.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/280.tif.

Bye:1965:OSN

[Bye65] J. A. T. Bye. Obtaining solutions of the Navier–Stokes equation by relaxation processes. The Computer Journal, 8(1):53–56, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
01/080053.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
01/tiff/53.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_01/tiff/
54.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_08/Issue_01/tiff/55.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
08/Issue_01/tiff/56.tif.

Bye:1967:NMD

[Bye67] J. A. T. Bye. A new method of determining eigenvalues and eigenfunctions. The Computer Journal, 10(1):108–111, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/100108.sgm.abs.html;
http://www3.oup.co.uk/computer_

                                        journal/hdb/Volume_10/Issue_
                                        01/tiff/108.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        01/tiff/109.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        01/tiff/110.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        01/tiff/111.tif.

Biasizzo:1998:SDA

[BZN98] Anton Biasizzo, Alenka Zuzek, and Franc Novak. Sequential diagnosis with asymmetrical tests. The Computer Journal, 41(3):163–170, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_41/Issue_03/
                                        Vol41_03.body.html#AbstractBiasizzo.

Carlson:1982:AFD

[CAC82] C. Robert Carlson, Adarsh K. Arora, and Miroslava Milosavljevic Carlson. The application of functional dependency theory to relational databases. The Computer Journal, 25(1):68–73, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/68.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_25/Issue_01/tiff/
                                        69.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_25/Issue_01/tiff/70.

tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/71.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/72.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/
73.tif.

Cadwell:1961:LSS

[Cad61]  J. H. Cadwell.  A least

squares surface fitting program. The Computer Journal, 3(4):266–269, January 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/030266.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/266.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/267.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/268.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/269.tif.

Caminer:1958:HAT

[Cam58] D. T. Caminer. And how to avoid them. The Computer Journal, 1(1):11–14, April 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
01/010011a.sgm.abs.html;

                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        01/tiff/11.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_01/Issue_01/tiff/
                                        12.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_01/Issue_01/tiff/13.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        01/Issue_01/tiff/14.tif.

Cameron:1969:LCP

[Cam69] P. T. Cameron. A language for chemical plant design and simulation programs. The Computer Journal, 12(1):24–28, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        01/120024.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        01/tiff/24.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_12/Issue_01/tiff/
                                        25.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_12/Issue_01/tiff/26.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        12/Issue_01/tiff/27.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        01/tiff/28.tif.

Campbell:1971:NOF

[Cam71] J. A. Campbell. A note on an optimal-fit method for dynamic allocation of storage. The Computer Journal, 14(1):7–9, February 1971. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/140007.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/tiff/7.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_14/Issue_01/tiff/
8.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_14/Issue_01/tiff/9.
tif.

Campbell:1973:ESA

[Cam73] D. V. A. Campbell. Errata: “A simple analysis of the n^{th} order polyphase sort”. The Computer Journal, 16(3):244, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/244.tif. See [Mac73].

Campbell:1978:CDF

[Cam78] W. R. Campbell. A compiler definition facility based on the syntactic macro. The Computer Journal, 21(1):35–41, February 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
01/tiff/35.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_01/tiff/
36.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_21/Issue_01/tiff/37.
tif;   http://www3.oup.co.

                                        uk/computer_journal/hdb/Volume_
                                        21/Issue_01/tiff/38.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        01/tiff/39.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_21/Issue_01/tiff/
                                        40.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_21/Issue_01/tiff/41.
                                        tif.

Caminer:1997:LAB

[Cam97] David Tresman Caminer, OBE. LEO and its applications: the beginning of business computing. The Computer Journal, 40(10):585–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_40/Issue_10/
                                        Vol40_10.body.html#AbstractCaminer.

Canny:1993:CRG

[Can93a] John Canny. Computing roadmaps of general semi-algebraic sets. The Computer Journal, 36(5):504–514, October 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_36/Issue_05/
                                        Vol36_05.body.html#AbstractCanny2.

Canny:1993:IAS

[Can93b] John Canny. Improved algorithms for sign determination and existential quantifier elimination. The Computer Journal, 36(5):409–418, October 1993. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_05/
Vol36_05.body.html#AbstractCanny.

Capper:1986:PTC

[Cap86] L. Capper. A philosophy for the teaching of computer science and information technology. The Computer Journal, 29(1):83–89, February 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
01/tiff/83.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_01/tiff/
84.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_29/Issue_01/tiff/85.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
29/Issue_01/tiff/86.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
01/tiff/87.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_01/tiff/
88.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_29/Issue_01/tiff/89.
tif.

Carre:1961:DOA

[Car61] B. A. Carre. The determination of the optimum accelerating factor for successive over-relaxation. The Computer Journal, 4(1):73–79, April 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/040073.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/tiff/73.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_01/tiff/
74.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_04/Issue_01/tiff/75.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
04/Issue_01/tiff/76.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/tiff/77.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_01/tiff/
78.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_04/Issue_01/tiff/79.
tif.

Carling:1964:SSI

[Car64] L. N. Carling. A study of the solution of an initial-value problem with a hybrid computer. The Computer Journal, 7(1):40–46, April 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
01/070040.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
01/tiff/40.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_01/tiff/
41.tif;   http://www3.oup.
co.uk/computer_journal/hdb/

Volume_07/Issue_01/tiff/42.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
07/Issue_01/tiff/43.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
01/tiff/44.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_01/tiff/
45.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_07/Issue_01/tiff/46.
tif.

Carre:1966:PNE

[Car66] B. A. Carre. The partitioning of network equations for block iteration. The Computer Journal, 9(1):84–97, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/090084.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/84.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_01/tiff/
85.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_09/Issue_01/tiff/86.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
09/Issue_01/tiff/87.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/88.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_01/tiff/
89.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_09/Issue_01/tiff/90.

                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        09/Issue_01/tiff/91.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        01/tiff/92.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_09/Issue_01/tiff/
                                        93.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_09/Issue_01/tiff/94.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        09/Issue_01/tiff/95.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        01/tiff/96.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_09/Issue_01/tiff/
                                        97.tif.

Carrick:1970:CIL

[Car70] D. A. Carrick. Correspondence: Interpretation of limited entry decision table format. The Computer Journal, 13(2):220–221, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/tiff/220.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/tiff/221.tif. See [Kin69a,

Kin71].

Carden:1986:SSC

[Car86] J. Carden. The structural stability of corporate and devolved databases. The Computer Journal, 29(4):361–367, August 1986. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/361.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/362.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/363.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/364.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/365.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/366.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/367.tif.

Caskey:1963:ASP

[Cas63] W. S. Caskey. Accounting for the soldier’s pay — organization of programming. The Computer Journal, 5(4):258–263, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/050258.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/258.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/259.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/260.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        04/tiff/261.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        04/tiff/262.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        04/tiff/263.tif.

Cash:1975:NSL

[Cas75] J. R. Cash. The numerical solution of linear stiff boundary value problems. The Computer Journal, 18(4):371–372, November 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        04/tiff/371.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        04/tiff/372.tif.

Cavouras:1983:CSE

[Cav83] John C. Cavouras. Computer system evaluation through supervisor replication. The Computer Journal, 26(2):134–141, May 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        02/tiff/134.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        02/tiff/135.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        02/tiff/136.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_

02/tiff/137.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/138.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/139.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/140.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/141.tif.

Cayless:1961:SSO

[Cay61] M. A. Cayless. Solution of systems of ordinary and partial differential equations by quasi-diagonal matrices. The Computer Journal, 4(1):54–61, April 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/040054.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/tiff/54.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_01/tiff/
55.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_04/Issue_01/tiff/56.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
04/Issue_01/tiff/57.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/tiff/58.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_01/tiff/
59.tif;   http://www3.oup.
co.uk/computer_journal/hdb/

                                        Volume_04/Issue_01/tiff/60.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        04/Issue_01/tiff/61.tif.

Coles:1976:ACP

[CBB^{+}76] E. C. Coles, L. J. Beilin, C. J. Bulpitt, C. T. Dollery, B. F. Johnson, A. D. Munro-Faure, and S. C. Turner. An approach to computer processing of clinical data. The Computer Journal, 19(2):104–109, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/104.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/105.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/106.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/107.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/108.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/109.tif.

Cahill:1994:CDA

[CBH^{+}94] Vinny Cahill, Roland Balter, David Harper, Neville Harris, Xavier Rousset De Pina, and Pedro Sousa. The Comandos distributed application platform. The Computer Journal, 37(6):477–486, ???? 1994. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_37/Issue_06/
Vol37_06.body.html#AbstractCahill.

Cohen:1974:NF

[CC74] J. Cohen and Eileen Carton. Non-deterministic Fortran. The Computer Journal, 17(1):44–51, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
01/tiff/44.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/
45.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_17/Issue_01/tiff/46.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
17/Issue_01/tiff/47.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
01/tiff/48.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/
49.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_17/Issue_01/tiff/50.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
17/Issue_01/tiff/51.tif.

Chang:1987:PAG

[CC87] C. C. Chang and C. Y. Chen.

Performance analysis of the generalised disc modulo allocation method for multiple key hashing files on multi-disc systems. The Computer Journal, 30(6):535–540, December 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        06/tiff/535.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        06/tiff/536.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        06/tiff/537.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        06/tiff/538.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        06/tiff/539.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        06/tiff/540.tif.

Chang:1988:RPF

[CC88] C. C. Chang and C. Y. Cheng. A recursive performance formula of the disc modulo allocation method for binary Cartesian product files. The Computer Journal, 31 (6):565–566, December 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        06/tiff/565.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        06/tiff/566.tif.

Chiaramella:1992:ARM

[CC92] Y. Chiaramella and J. P. Chevallet. About retrieval models and logic. The Computer Journal, 35(3):233–242, June 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_35/Issue_03/
Vol35_03.body.html#AbstractChiaramella.

Chang:1993:NLA

[CC93] Henry Ker-Chang Chang and Shing Hong Chen. A new locally adaptive data compression scheme using multilist structure. The Computer Journal, 36(6):570–578, December 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_06/
Vol36_06.body.html#AbstractChang.

Clarke:1995:AST

[CCB95] Charles L. A. Clarke, G. V.

Cormack, and F. J. Burkowski. An algebra for structured text search and a framework for its implementation. The Computer Journal, 38(1):43–56, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_38/Issue_01/
Vol38_01.body.html#AbstractClarke.

Chang:1992:SGC

[CCC92] C. C. Chang, H. Y. Chen, and C. Y. Chen. Symbolic Gray code as a data allocation scheme for two-disc systems. The Computer Journal, 35(3):299–305, June 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_35/Issue_03/
Vol35_03.body.html#AbstractChang.

Cantone:1986:WCU

[CCD86] G. Cantone, A. Cimitile, and U. De Carlini. Well-formed conversion of unstructured one-in/ one-out schemes for complexity measurement and program maintenance. The Computer Journal, 29(4):322–329, August 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/322.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/323.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/324.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/325.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/326.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/327.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/328.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/329.tif.

Chaplin:1973:GRB

[CCH73] R. I. Chaplin, R. E. Crosbie, and J. L. Hay. A graphical representation of the Backus-Naur form. The Computer Journal, 16(1):28–29, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
01/160028.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
01/tiff/28.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_01/tiff/
29.tif.

Chang:1991:DMP

[CCJ91] C. C. Chang, C. Y. Chen, and J. K. Jan. On the design of a machine-independent perfect hashing scheme. The Computer Journal, 34(5):469–474, October 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/469.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/470.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/471.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/472.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/473.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        05/tiff/474.tif.

Clare:1972:IVL

[CCL72] A. C. Clare, E. M. Cook, and M. F. Lynch. The identification of variable-length, equifrequent character strings in a natural language data base. The Computer Journal, 15(3):259–262, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/150259.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/259.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/260.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/261.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/262.tif.

Coulson:1987:PNA

[CCL87] A. F. W. Coulson, J. F. Collins, and A. Lyall. Protein and nucleic acid sequence database searching: a suitable case for parallel processing. The Computer Journal, 30(5):420–424, October 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_

05/tiff/420.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/421.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/422.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/423.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/424.tif.

Chen:1998:PEF

[CCL98] Ing-Ray Chen, Tsong-Min Chen, and Chiang Lee. Performance evaluation of forwarding strategies for location management in mobile networks. The Computer Journal, 41(4):243–??, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_41/Issue_04/
Vol41_04.body.html#AbstractChen.

Chan:1997:SPO

[CCMS97] Fun Ting Chan, Tsong Yueh Chen, Ieng Kei Mark, and Shir Ming Shen. On some properties of the optimally refined proportional sampling strategy. The Computer Journal, 40(4):194–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www.oup.co.uk/computer_
journal/Volume_40/Issue_03/
Vol40_04.body.html#AbstractChan;
http://www3.oup.co.uk/computer_

journal/Volume_40/Issue_04/
Vol40_04.body.html#AbstractChan.

Coulter:1987:ICP

[CCS87] N. S. Coulter, R. B. Cooper, and M. K. Solomon. Informationtheoretic complexity of program specifications. The Computer Journal, 30(3):223–227, June 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/223.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/224.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/225.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/226.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/227.tif.

Cole:1969:LSP

[CD69] A. J. Cole and A. J. T. Davie. Local smoothing by polynomials in n dimensions. The Computer Journal, 12(1):72–76, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
01/120072.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
01/tiff/72.tif;   http://
www3.oup.co.uk/computer_journal/

hdb/Volume_12/Issue_01/tiff/
73.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_12/Issue_01/tiff/74.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/75.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
01/tiff/76.tif.

Carpenter:1977:OTM

[CD77] B. E. Carpenter and R. W. Doran. The other Turing machine. The Computer Journal, 20(3):269–279, August 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/269.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/270.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/271.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/272.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/273.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/274.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/275.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/276.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_20/Issue_
                                        03/tiff/277.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        03/tiff/278.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        03/tiff/279.tif. Hodges page

318 note 6.1.

Clark:1980:ACT

[CD80] K. L. Clark and J. Darlington. Algorithm classification through synthesis. The Computer Journal, 23 (1):61–65, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        01/tiff/61.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_23/Issue_01/tiff/
                                        62.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_23/Issue_01/tiff/63.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        23/Issue_01/tiff/64.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        01/tiff/65.tif.

Cavouras:1981:STC

[CD81] J. C. Cavouras and R. H. Davis. Simulation tools in computer system design methodologies. The Computer Journal, 24(1):25–28, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_24/Issue_
01/tiff/25.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_01/tiff/
26.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_24/Issue_01/tiff/27.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
24/Issue_01/tiff/28.tif.

Cocu:1982:CAS

[CD82] J. J. Cocu and R. E. Devillers. On a class of allocation strategies inducing bounded delays only. The Computer Journal, 25(1):52–55, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/52.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/
53.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_25/Issue_01/tiff/54.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/55.tif.

Crookes:1985:AOP

[CD85] D. Crookes and C. J. Devlin.

An algorithm for optimal procedural nesting. The Computer Journal, 28(1):17–21, February 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
01/tiff/17.tif;   http://
www3.oup.co.uk/computer_journal/

                                        hdb/Volume_28/Issue_01/tiff/
                                        18.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_28/Issue_01/tiff/19.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        28/Issue_01/tiff/20.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        01/tiff/21.tif.

Chan:1997:TPF

[CDCL97] E. Chan, C. Daoxu, J. Cao, and C.-H Lee. Timing properties of the FDDI-M medium access protocol. The Computer Journal, 40(1):43–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www.oup.co.uk/computer_
                                        journal/Volume_40/Issue_01/
                                        Vol40_01.body.html#AbstractChan;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_40/Issue_01/
                                        Vol40_01.body.html#AbstractChan.

Charlton:1993:AGR

[CDHL93] C. C. Charlton, P. E. Dunne, K. Halewood, and P. H. Leng. An algorithm to generate random large combinational circuits. The Computer Journal, 36(2):200–206, April 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_36/Issue_02/
                                        Vol36_02.body.html#AbstractCharlton.

Chau:1974:IIG

[CDZ74] A. Y. C. Chau, B. W. Davies, and B. Zacharov. Island — an interactive graphics system for mathematical analysis. The Computer Journal, 17(2):104–112, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/104.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/105.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/106.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/107.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/108.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/109.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/110.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/111.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/112.tif.

Coulouris:1972:TCA

[CEM72] G. F. Coulouris, J. M. Evans, and R. W. Mitchell. Towards content-addressing in data bases. The Computer Journal, 15(2):95–98, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        02/150095.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        02/tiff/95.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_15/Issue_02/tiff/
                                        96.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_15/Issue_02/tiff/97.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        15/Issue_02/tiff/98.tif.

Clarke:1959:PA

[CF59] B. Clarke and G. E. Felton. The Pegasus autocode. The Computer Journal, 1 (4):192–195, January 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        04/010192.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        04/tiff/192.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        04/tiff/193.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        04/tiff/194.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        04/tiff/195.tif.

Cairns:1981:BRT

[CF81] D. J. Cairns and P. R. Freeman. Book reviews: Theory and Design of Digital Computer Systems, 2nd ed., by Douglas Lewin, 1980; Statistical Computing, by W. J. Kennedy and J. E. Gentle, 1980. The Computer Journal, 24(1):96, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
01/tiff/96.tif. See [Mad80,

Wes80b].

Cole:1988:ECS

[CF88] R. Cole and Clare Foxcroft.

An experiment in clock synchronisation. The Computer Journal, 31(6):496–502, December 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/496.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/497.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/498.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/499.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/500.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/501.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/502.tif.

Cosnard:1989:GPV

[CF89] M. Cosnard and A. G. Ferreira. Generating permutations on a VLSI suitable linear network. The Computer Journal, 32(6):571–573, December 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        06/tiff/571.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        06/tiff/572.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        06/tiff/573.tif.

Cifuentes:1996:ARC

[CF96] Cristina Cifuentes and Anne Fitzgerald. Australian recommendations on computer software protection. The Computer Journal, 39(7):566–576, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://www.oup.co.uk/jnls/
                                        list/comjnl/hdb/Volume_39/
                                        Issue_07/390566.sgm.abs.html;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_39/Issue_07/
                                        Vol39_07.body.html#AbstractCifuentes.

Corsini:1984:IAO

[CFL84] P. Corsini, G. Frosini, and L. Lopriore. The implementation of abstract objects in a capability based addressing architecture. The Computer Journal, 27(2):127–134, May 1984. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/127.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/128.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/129.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/130.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/131.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/132.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/133.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/134.tif.

Corsini:1989:PIL

[CFS89] P. Corsini, G. Frosini, and G. Speranza. The parallel interpretation of logic programs in distributed architectures. The Computer Journal, 32(1):29–35, February 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
01/tiff/29.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_01/tiff/
30.tif;   http://www3.oup.
co.uk/computer_journal/hdb/

                                        Volume_32/Issue_01/tiff/31.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        32/Issue_01/tiff/32.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        01/tiff/33.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_32/Issue_01/tiff/
                                        34.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_32/Issue_01/tiff/35.
                                        tif.

Cooper:1965:CID

[CG65] G. J. Cooper and E. Gal. Correspondence: Integration of differential equations. The Computer Journal, 8(2):176, July 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        02/tiff/176.tif.

Conlon:1992:DTC

[CG92] T. Conlon and S. Gregory. Debugging tools for concurrent logic programming. The Computer Journal, 35(2):157–169, April 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_35/Issue_02/
                                        Vol35_02.body.html#AbstractConlon.

Chinchwadkar:1999:OVP

[CG99] G. S. Chinchwadkar and A. Goh. An overview of vertical partitioning in object oriented databases. The Computer Journal, 42(1):39–50, ???? 1999.

CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
01/420039.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
01/pdf/420039.pdf.

Coulouris:1968:LCC

[CGH^{+}68] G. F. Coulouris, T. J. Goodey, R. W. Hill, R. W. Keeling, and D. Levin. The London CPL1 compiler. The Computer Journal, 11(1):26–30, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/110026.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/26.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_01/tiff/
27.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_11/Issue_01/tiff/28.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/29.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/30.tif.

Cowell:1980:SSA

[CGK80] D. F. Cowell, D. F. Gillies, and A. A. Kaposi. Synthesis and structural analysis of abstract programs. The Computer Journal, 23(3):243–247, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/243.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/244.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/245.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/246.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/247.tif.

Ross:1961:CCC

[cGR61] H. M. c. G. Ross. Considerations in choosing a character code for computers and punched tapes. The Computer Journal, 3(4):202–210, January 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/030202.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/202.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/203.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/204.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/205.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_03/Issue_
04/tiff/206.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/207.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/208.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/209.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/210.tif.

Cawsey:1992:ALB

[CGRJ92] Alison Cawsey, Julia Galliers, Steven Reece, and Karen Sparck Jones. Automating the librarian: Belief revision as a base for system action and communication with the user. The Computer Journal, 35(3):221–232, June 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_35/Issue_03/
Vol35_03.body.html#AbstractCawsey.

Cockayne:1960:PNC

[CH60] A. H. Cockayne and E. Hyde.

Prime number code for information retrieval. The Computer Journal, 3(1):21–22, April 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
01/030021.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
01/tiff/21.tif;   http://

www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_01/tiff/
22.tif.

Cormack:1987:DCU

[CH87] G. V. Cormack and R. N. S. Horspool. Data compression using dynamic Markov modelling. The Computer Journal, 30(6):541–550, December 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/541.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/542.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/543.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/544.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/545.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/546.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/547.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/548.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/549.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/550.tif.

Chuang:1999:EFD

[CH99] Po-Jen Chuang and Huey-Ming Hsu. Effective fault diagnosis approaches for multipath networks. The Computer Journal, 42(5):409–421, 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
05/420409.sgm.abs.html.

Chartres:1962:AJM

[Cha62] B. A. Chartres. Adaptation of the Jacobi method for a computer with magnetic-tape backing store. The Computer Journal, 5(1):51–60, April 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
01/050051.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
01/tiff/51.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_01/tiff/
52.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_05/Issue_01/tiff/53.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
05/Issue_01/tiff/54.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
01/tiff/55.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_01/tiff/
56.tif;   http://www3.oup.
co.uk/computer_journal/hdb/

                                        Volume_05/Issue_01/tiff/57.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        05/Issue_01/tiff/58.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        01/tiff/59.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_05/Issue_01/tiff/
                                        60.tif.

Chapman:1963:ULC

[Cha63]  F. G. Chapman.  Use of

large computers at a distance. The Computer Journal, 6(3):214–218, October 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        03/060214.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        03/tiff/214.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        03/tiff/215.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        03/tiff/216.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        03/tiff/217.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        03/tiff/218.tif.

Chawla:1967:NEC

[Cha67] M. M. Chawla. A note on the estimation of the coefficients in the Chebyshev series expansion of a function having a logarithmic singularity. The Computer Journal, 9(4):413, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/090413.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/413.tif.

Chawla:1968:AEE

[Cha68]  M. M. Chawla.  Asymp-

totic estimates for the error of the Gauss–Legendre quadrature formula. The Computer Journal, 11(3):339–340, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/110339.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/339.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/340.tif.

Chawla:1970:EEG

[Cha70] M. M. Chawla. Estimation of errors of Gauss–Chebyshev quadratures. The Computer Journal, 13(1):107–109, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/130107.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/tiff/107.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        01/tiff/108.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        01/tiff/109.tif.

Chambers:1971:ARF

[Cha71] J. M. Chambers. Another round of FORTRAN. The Computer Journal, 14(3):312–314, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/140312.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/tiff/312.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/tiff/313.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/tiff/314.tif. See cor-

respondence [Dew72b, Fla72, Cha72].

Chambers:1972:CAR

[Cha72] J. M. Chambers. Correspondence: On “Another round of FORTRAN”. The Computer Journal, 15(4):381, November 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/150381.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/tiff/381.tif. See [Cha71,

Dew72b, Fla72].

Chance:1973:BRF

[Cha73] E. M. Chance. Book review:

FORTRAN Techniques, by A. Colin Day, 1972. The Computer Journal, 16(3):265, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/265.tif.

Chambers:1975:ATP

[Cha75] A. D. Chambers. Audit test packs and computer audit programs. The Computer Journal, 18(2):98–101, May 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/100.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/101.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/98.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_02/tiff/
99.tif.

Chambers:1978:CFA

[Cha78] A. D. Chambers. Computer fraud and abuse. The Computer Journal, 21(3):194–198, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/194.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/195.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/196.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/197.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/198.tif.

Chambers:1981:CSC

[Cha81] A. D. Chambers. Current strategies for computer auditing within an organization. The Computer Journal, 24(4):290–294, November 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/290.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/291.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/292.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/293.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/294.tif.

Chang:1987:IPS

[Cha87] C. C. Chang. An information protection scheme based upon number theory. The Computer Journal, 30(3):249–253, June 1987. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/249.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/250.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/251.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/252.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/253.tif.

Chaudhuri:1990:FUD

[Cha90] P. Chaudhuri. Finding and updating depth-first spanning trees of acyclic digraphs in parallel. The Computer Journal, 33(3):247–251, June 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/247.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/248.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/249.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/250.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/251.tif.

Challab:1991:IFA

[Cha91] D. J. Challab. Implementation of flexible arrays using balanced trees. The Computer Journal, 34(5):386–396, October 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        05/tiff/386.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        05/tiff/387.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        05/tiff/388.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        05/tiff/389.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        05/tiff/390.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        05/tiff/391.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        05/tiff/392.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        05/tiff/393.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        05/tiff/394.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        05/tiff/395.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        05/tiff/396.tif.

Challab:1992:SNF

[Cha92] D. J. Challab. Short notes:

Flexible arrays: evaluation of the B-tree based system. The Computer Journal, 35 (5):527–529, October 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_35/Issue_05/
350527.sgm.abs.html;   http:
//www3.oup.co.uk/computer_
journal/Volume_35/Issue_05/
tiff/527.tif;   http://www3.
oup.co.uk/computer_journal/
Volume_35/Issue_05/tiff/528.
tif;   http://www3.oup.co.
uk/computer_journal/Volume_
35/Issue_05/tiff/529.tif.

Challis:1993:TNT

[Cha93a]  M. F. Challis.  Two new

techniques for computing extremal h-bases A_{k}. The Computer Journal, 36(2):117–126, April 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_02/
Vol36_02.body.html#AbstractChallis.

Chandru:1993:VEL

[Cha93b] Vijay Chandru. Variable elimination in linear constraints. The Computer Journal, 36 (5):463–472, October 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_

journal/Volume_36/Issue_05/
Vol36_05.body.html#AbstractChandru.

Chaudhuri:1997:ODA

[Cha97] Pranay Chaudhuri. An optimal distributed algorithm for computing bridge- connected components. The Computer Journal, 40(4):200–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www.oup.co.uk/computer_
journal/Volume_40/Issue_03/
Vol40_04.body.html#AbstractChaudhuri;
 http://www3.oup.co.uk/computer_
journal/Volume_40/Issue_04/
Vol40_04.body.html#AbstractChaudhuri.

Checksfield:1962:FCR

[Che62a] A. E. Checksfield. The first computer in Rhodesia. The Computer Journal, 5(2):79–87, July 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/050079.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/79.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_02/tiff/
80.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_05/Issue_02/tiff/81.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
05/Issue_02/tiff/82.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/83.tif;   http://

www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_02/tiff/
84.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_05/Issue_02/tiff/85.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
05/Issue_02/tiff/86.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/87.tif.

Cheesman:1962:SBC

[Che62b] D. V. Cheesman. A small business computer at work. The Computer Journal, 5(1):1–6, April 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
01/050001.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
01/tiff/1.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_05/Issue_01/tiff/
2.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_05/Issue_01/tiff/3.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
05/Issue_01/tiff/4.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
01/tiff/5.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_05/Issue_01/tiff/
6.tif.

Chew:1964:NTS

[Che64] R. L. Chew. Note on timing simulation of a large asynchronous computer. The Computer Journal, 7(2):122–123, July 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        02/070122.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        02/tiff/122.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        02/tiff/123.tif.

Chezem:1965:NTD

[Che65] C. G. Chezem. Note on ‘threedimensional’ plotting as a technique for finding the zeros of functions in the complex plane. The Computer Journal, 8(3):288, October 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        03/080288.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        03/tiff/288.tif.

Chen:1986:DAT

[Che86] T. Y. Chen. Dynamic approach towards software science. The Computer Journal, 29(2):189, April 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        02/tiff/189.tif.

Chen:1990:CMF

[Che90] T. Y. Chen. On the consistency of multi-valued functions. The Computer Journal, 33(6):570–572, December 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/570.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/571.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/572.tif.

Chen:1995:DBT

[Che95a] Ing Ray Chen. A degradable B^{link}-tree with periodic data reorganization. The Computer Journal, 38(3):245–256, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_38/Issue_03/
Vol38_03.body.html#AbstractChen.

Chen:1995:AIC

[Che95b] Ing-Ray Chen. On applying imprecise computation to realtime AI systems. The Computer Journal, 38(6):434–442, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_38/Issue_06/
Vol38_06.body.html#AbstractChen.

Chen:1995:SPN

[Che95c] Ing-Ray Chen. Stochastic Petri net analysis of deadlock detection algorithms in transaction database systems with dynamic locking. The Computer Journal, 38(9):717–733, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_38/Issue_09/
Vol38_09.body.html#AbstractChen.

Chen:1995:ECA

[Che95d] Jingsen Chen. An efficient construction algorithm for a class of implicit double-ended priority queues. The Computer Journal, 38(10):818–821, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_38/Issue_10/
Vol38_10.body.html#AbstractChen.

Coulson:1973:ASP

[CHH73] V. G. Coulson, I. D. Hill, and V. F. Hillier. ALGOL survey program. The Computer Journal, 16(1):13–15, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
01/160013.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
01/tiff/13.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_01/tiff/
14.tif;   http://www3.oup.
co.uk/computer_journal/hdb/

Volume_16/Issue_01/tiff/15.
tif.

Cormack:1985:PPH

[CHK85] G. V. Cormack, R. N. S.

Horspool, and M. Kaiserswerth. Practical perfect hashing. The Computer Journal, 28(1):54–58, February 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
01/tiff/54.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_01/tiff/
55.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_28/Issue_01/tiff/56.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
28/Issue_01/tiff/57.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
01/tiff/58.tif.

Christensen:1994:DSC

[CHM94] Søren Christensen, Yoram Hirshfeld, and Faron Moller. Decidable subsets of CCS. The Computer Journal, 37(4):233–242, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_37/Issue_04/
Vol37_04.body.html#AbstractChristensen.

Cho:1982:ARF

[Cho82] J. R. Lin Sin Cho. Automatic report formatting from a report specimen. The Computer Journal, 25(2):242–247, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/242.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/243.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/244.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/245.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/246.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/247.tif.

Chou:1995:MVD

[Cho95] Ching-Tsun Chou. Mechanical verification of distributed algorithms in higher-order logic. The Computer Journal, 38 (2):152–161, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_38/Issue_02/
Vol38_02.body.html#AbstractChou.

Choy:1999:CIR

[Cho99] M. Choy. Consistent implementations of replicated objects. The Computer Journal, 42(1):24–38, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_

journal/hdb/Volume_42/Issue_
01/420024.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
01/pdf/420024.pdf.

Clint:1985:ADF

[CHPS85] M. Clint, C. Holt, R. H. Perrott, and A. Stewart. Algorithm 120: A DAP FORTRAN subroutine for the eigensolution of real symmetric matrices. The Computer Journal, 28(3):340–342, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/340.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/341.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/342.tif.

Christofides:1971:ACN

[Chr71a]  N. Christofides.  An algo-

rithm for the chromatic number of a graph. The Computer Journal, 14(1):38–39, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/140038.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/tiff/38.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_01/tiff/
39.tif.

Christofides:1971:ZOP

[Chr71b] N. Christofides. Zero-one programming using non-binary tree-search. The Computer Journal, 14(4):418–421, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/140418.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/418.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/419.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/420.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/421.tif.

Christiansen:1973:ASS

[Chr73a] Jan Christiansen. Algorithm 77: Solving a system of simultaneous ordinary differential equations of the first order using a method for automatic step change. The Computer Journal, 16(2):187–188, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/187.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/188.tif.

Christofides:1973:LSP

[Chr73b] N. Christofides. Large scheduling problems with bivalent costs. The Computer Journal, 16(3):262–264, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/160262.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/262.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/263.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/264.tif.

Churchhouse:1968:DCN

[Chu68] R. F. Churchhouse. Discussion and correspondence: note on the twenty-five most cited papers in some leading journals. The Computer Journal, 11(1):116–120, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/110116.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/116.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/117.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/118.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        01/tiff/119.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        01/tiff/120.tif.

Clint:1970:EEE

[CJ70] M. Clint and A. Jennings. The evaluation of eigenvalues and eigenvectors of real symmetric matrices by simultaneous iteration. The Computer Journal, 13(1):76–80, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        01/130076.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        01/tiff/76.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_13/Issue_01/tiff/
                                        77.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_13/Issue_01/tiff/78.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        13/Issue_01/tiff/79.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        01/tiff/80.tif. See erratum

[Sea71].

Cooper:1993:SRI

[CK93] Tim B. Cooper and Jeffrey H. Kingston. The solution of real instances of the timetabling problem. The Computer Journal, 36(7):645–653, ???? 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_36/Issue_07/
Vol36_07.body.html#AbstractCooper.

Chu:1994:ASH

[CK94] Jiang-Hsing Chu and Gary D.

Knott. An analysis of spiral hashing. The Computer Journal, 37(8):715–719, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_37/Issue_08/
Vol37_08.body.html#AbstractChu.

Campbell-Kelly:1995:IEB

[CK95] M. Campbell-Kelly. ICL and the evolution of the British mainframe. The Computer Journal, 38(5):400–412, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_38/Issue_05/
Vol38_05.body.html#AbstractKelly.

Clark:1966:CMP

[CL66] S. R. Clark and W. F. Lunnon. Correspondence: Multiple precision arithmetic (real and complex). The Computer Journal, 9(2):174, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/174.tif. See [BRC66].

Cash:1978:CAS

[CL78] J. R. Cash and C. B. Liem. On the computational aspects of semi-implicit Runge–Kutta methods. The Computer Journal, 21(4):363–365, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/363.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/364.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/365.tif.

Chan:1983:AFS

[CL83] N. N. Chan and Kim-Hung Li. Algorithm 115: A FOR-TRAN subroutine for finding a real symmetric matrix with prescribed diagonal elements and eigenvalues. The Computer Journal, 26(2):184–186, May 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/184.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/185.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/186.tif.

Chang:1986:LOM

[CL86] C. C. Chang and R. C. T. Lee. A letter-oriented minimal perfect hashing scheme. The Computer Journal, 29(3):277–281, June 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/277.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/278.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/279.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/280.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/281.tif.

Chen:1999:AMP

[CL99a] Guan-Chi Chen and Suh-Yin Lee. An analytic model for performance analysis of concurrency control strategies in mobile environments. The Computer Journal, 42(6):511– 521, 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
06/420511.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
06/pdf/420511.pdf.

Chen:1999:CTS

[CL99b] Tsong Yueh Chen and Man Fai Lau. On the completeness of a test suite reduction strategy. The Computer Journal, 42(5):430–440, 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        05/420430.sgm.abs.html.

Clarke:1960:NSD

[Cla60]  L. Clarke.  Notes on the

state of digital computing in the U.S.S.R. The Computer Journal, 3(3):164–167, October 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        03/030164.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        03/tiff/164.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        03/tiff/165.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        03/tiff/166.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        03/tiff/167.tif.

Clapson:1977:TPS

[Cla77a]  P. J. Clapson.  Toward a

performance science: a comparative analysis of computing systems. The Computer Journal, 20(4):308–315, November 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        04/tiff/308.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_20/Issue_
04/tiff/309.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/310.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/311.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/312.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/313.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/314.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/315.tif.

Clasen:1977:NUC

[Cla77b] R. J. Clasen. A note on the use of the conjugate gradient method in the solution of a large system of sparse equations. The Computer Journal, 20(2):185–186, May 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/185.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/186.tif.

Clark:1978:SSA

[Cla78] I. A. Clark. Strema: Specifying application processes using streams. The Computer Journal, 21(1):25–30, February 1978. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
01/tiff/25.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_01/tiff/
26.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_21/Issue_01/tiff/27.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
21/Issue_01/tiff/28.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
01/tiff/29.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_01/tiff/
30.tif.

Clarke:1988:JUI

[Cla88] R. A. Clarke. Judicial understanding of information technology: the case of the Wombat ROMs. The Computer Journal, 31(1):25–33, February 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
01/tiff/25.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_01/tiff/
26.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_31/Issue_01/tiff/27.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
31/Issue_01/tiff/28.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
01/tiff/29.tif;   http://
www3.oup.co.uk/computer_journal/

hdb/Volume_31/Issue_01/tiff/
30.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_31/Issue_01/tiff/31.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
31/Issue_01/tiff/32.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
01/tiff/33.tif.

Clark:1990:PLP

[Cla90] K. L. Clark. Parallel logic programming. The Computer Journal, 33(6):482–493, December 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/482.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/483.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/484.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/485.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/486.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/487.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/488.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/489.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_

                                        06/tiff/490.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/491.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/492.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/493.tif.

Clarkson:1992:BRV

[Cla92] Trevor Clarkson. Book review: VLSI Design of Neural Networks, by Ulrich Ramacher and Ulrich Ruckert. The Computer Journal, 35(6):624, December 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_35/Issue_
                                        06/tiff/624.tif.

Clark:1994:PRN

[Cla94] A. N. Clark. Pattern recognition of noisy sequences of behavioural events using functional combinators. The Computer Journal, 37(5):385–398, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_37/Issue_05/
                                        Vol37_05.body.html#AbstractClark.

Clarke:1999:DPR

[Cla99] B. Clarke. Discussion of the papers by Rissanen, and by Wallace and Dowe. The Computer Journal, 42(4):338–339, ???? 1999. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
04/420338.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
04/pdf/420338.pdf.

Chang:1992:OPD

[CLC92] C. C. Chang, H. F. Lin, and C. Y. Chen. Optimality properties of the disc modulo allocation method for k-ary multiple key hashing files. The Computer Journal, 35(4):390–394, August 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_35/Issue_04/
Vol35_04.body.html#AbstractChang.

Cleave:1959:AFT

[Cle59]  J. P. Cleave.  Algorithms

for formula translation. The Computer Journal, 2(2):53–54, July 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
02/020053.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
02/tiff/53.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_02/tiff/
54.tif.

Clenshaw:1960:CFD

[Cle60]  C. W. Clenshaw.  Curve

fitting with a digital computer. The Computer Journal, 2(4):170–173, January 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/020170.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/170.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/171.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/172.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/173.tif.

Clementson:1966:ECS

[Cle66] A. T. Clementson. Extended control and simulation language. The Computer Journal, 9(3):215–220, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/090215.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/215.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/216.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/217.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/218.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_09/Issue_
03/tiff/219.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/220.tif.

Clementson:1967:ALS

[Cle67] A. T. Clementson. An assembly and loading system for computers with parallel peripheral operation. The Computer Journal, 9(4):370–372, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/090370.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/370.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/371.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/372.tif.

Clement:1992:RDR

[Cle92]  T. Clement.  The role of

data reification in program refinement: Origins, synthesis and appraisal. The Computer Journal, 35(5):451–459, October 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_35/Issue_05/
Vol35_05.body.html#AbstractClement.

Clippinger:1962:C

[Cli62a] R. F. Clippinger. COBOL.

The Computer Journal, 5 (3):177–180, October 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/050177.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/177.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/178.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/179.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/180.tif.

Clippinger:1962:F

[Cli62b]  R. F. Clippinger.  FACT.

The Computer Journal, 5(2): 112–125, July 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
05/Issue_02/050112.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
05/Issue_02/tiff/112.tif;
 http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/113.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/114.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/115.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/116.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/117.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/118.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/119.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/120.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/121.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/122.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/123.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/124.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/125.tif.

Clippinger:1962:IA

[Cli62c] R. F. Clippinger. Information algebra. The Computer Journal, 5(3):180–192, October 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/050180.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/180.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/181.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        03/tiff/182.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        03/tiff/183.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        03/tiff/184.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        03/tiff/185.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        03/tiff/186.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        03/tiff/187.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        03/tiff/188.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        03/tiff/189.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        03/tiff/190.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        03/tiff/191.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        03/tiff/192.tif.

Cloot:1965:WUO

[Clo65] P. L. Cloot. What is the use of operating systems? The Computer Journal, 7 (4):249–254, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_

04/070249.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/249.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/250.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/251.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/252.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/253.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/254.tif.

Clocksin:1986:ASS

[Clo86] W. F. Clocksin. Automatic specialisation of standard designs. The Computer Journal, 29(6):495–499, December 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/495.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/496.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/497.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/498.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/499.tif.

Clocksin:1987:PDP

[Clo87] W. F. Clocksin. Principles of the Delphi Parallel Inference Machine. The Computer Journal, 30(5):386–392, October 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        05/tiff/386.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        05/tiff/387.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        05/tiff/388.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        05/tiff/389.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        05/tiff/390.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        05/tiff/391.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        05/tiff/392.tif.

Chu:1998:EHB

[CLW98] K. W. Chu, S. K. Lam, and M. H. Wong. An efficient hashbased algorithm for sequence data searching. The Computer Journal, 41(6):402–415, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_41/Issue_06/
                                        Vol41_06.body.html#AbstractChu.

Cash:1979:IAN

[CM79] J. R. Cash and J. C. P. Miller.

On an iterative approach to the numerical solution of difference schemes. The Computer Journal, 22(2):184–187, May 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/184.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/185.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/186.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/187.tif.

Culberson:1989:EBB

[CM89] J. Culberson and J. I. Munro.

Explaining the behaviour of binary search trees under prolonged updates: a model and simulations. The Computer Journal, 32(1):68–75, February 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
01/tiff/68.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_01/tiff/
69.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_32/Issue_01/tiff/70.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
32/Issue_01/tiff/71.tif;

                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        01/tiff/72.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_32/Issue_01/tiff/
                                        73.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_32/Issue_01/tiff/74.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        32/Issue_01/tiff/75.tif.

Campbell:1991:FMW

[CM91] D. Campbell and T. McNeill. Finding a majority when sorting is not available. The Computer Journal, 34(2):186, April 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        02/tiff/186.tif. See corre-

spondence [Gri92].

Czech:1993:LTA

[CM93] Zbigniew J. Czech and Bohdan S. Majewski. A linear time algorithm for finding minimal perfect hash functions. The Computer Journal, 36(6):579–587, December 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_36/Issue_06/
                                        Vol36_06.body.html#AbstractCzech.

Capon:1972:MCT

[CMRW72] P. C. Capon, D. Morris, J. S. Rohl, and I. R. Wilson. The MU5 compiler target language and autocode. The Computer Journal, 15(2):109–112, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/150109.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/109.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/110.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/111.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/112.tif.

Colin:1980:ST

[CMS80] A. J. T. Colin, A. D. McGettrick, and P. D. Smith. Sorting trains. The Computer Journal, 23(3):270–273, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/270.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/271.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/272.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/273.tif.

Cockshott:1998:HPO

[CMW98] W. Paul Cockshott, Douglas McGregor, and John Wilson. High performance operations using a compressed database architecture. The Computer Journal, 41(5):283–296, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_41/Issue_05/
                                        Vol41_05.body.html#AbstractCockshott.

Clenshaw:1963:SNO

[CN63] C. W. Clenshaw and H. J. Norton. The solution of nonlinear ordinary differential equations in Chebyshev series. The Computer Journal, 6(1):88–92, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        01/060088.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        01/tiff/88.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_06/Issue_01/tiff/
                                        89.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_06/Issue_01/tiff/90.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        06/Issue_01/tiff/91.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        01/tiff/92.tif.

Chan:1982:ECA

[CN82] Arvola Chan and Bahram Niamir. On estimating the cost of accessing records in blocked database organizations. The Computer Journal, 25(3):368–374, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/368.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/369.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/370.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/371.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/372.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/373.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/374.tif.

Cohen:1966:NOG

[CND66] J. Cohen and X. Nguyen-Dinh. Note on ordering of grammar rules in syntaxanalyzers. The Computer Journal, 9(3):250–251, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/090250.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/250.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_09/Issue_
                                        03/tiff/251.tif.

Cahill:1997:OMD

[CNTR97] Vinny Cahill, Paddy Nixon, Brendan Tangney, and Fethi Rabhi. Object models for distributed or persistent programming. The Computer Journal, 40(8):513–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_40/Issue_08/
                                        Vol40_08.body.html#AbstractCahill.

Curtis:1966:CMR

[CO66] A. Curtis and M. R. Osborne. The construction of minimax rational approximations to functions. The Computer Journal, 9(3):286–293, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        03/090286.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        03/tiff/286.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        03/tiff/287.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        03/tiff/288.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        03/tiff/289.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        03/tiff/290.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/291.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/292.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/293.tif.

Cunningham:1972:EHG

[CO72] K. M. Cunningham and J. C.

Ogilvie. Evaluation of hierarchical grouping techniques; a preliminary study. The Computer Journal, 15(3):209–213, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/150209.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/209.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/210.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/211.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/212.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/213.tif.

Casey:1976:RAS

[CO76] R. G. Casey and I. M. Osman. Replacement algorithms for storage management in relational data bases. The Computer Journal, 19 (4):306–314, November 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/306.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/307.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/308.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/309.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/310.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/311.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/312.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/313.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/314.tif.

Coates:1974:MCB

[Coa74]  H. N. Coates.  Manage-

ment in the computer business. The Computer Journal, 17(3):279–282, August 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/279.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_17/Issue_
03/tiff/280.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/281.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/282.tif.

Cobb:1969:NA

[Cob69] S. M. Cobb. Note on Algorithm 35: Best rational approximation to a real number. The Computer Journal, 12(3):293–295, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/293.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/294.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/295.tif. See [Pre68,

PP69].

Cochrane:1997:CGR

[Coc97] P. Cochrane. From copper to glass: the right idea, decisions and investments at the right time. … The Computer Journal, 40(1):1–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www.oup.co.uk/computer_
journal/Volume_40/Issue_01/
Vol40_01.body.html#AbstractCochrane;
 http://www3.oup.co.uk/computer_
journal/Volume_40/Issue_01/
Vol40_01.body.html#AbstractCochrane.

Cohen:2001:RCS

[Coh01] Fred Cohen. Reply to “Comment on ‘A Framework for Modelling Trojans and Computer Virus Infection’ by E. Mäkinen”. The Computer Journal, 44(4):326–327, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_44/Issue_
                                        04/440326.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_44/Issue_
                                        04/pdf/440326.pdf.  See

[TAC98, Mäk01, TAC01].

Colin:1963:NCR

[Col63a] A. J. T. Colin. Note on coding reverse Polish expressions for single-address computers with one accumulator. The Computer Journal, 6(1):67–68, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        01/060067.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        01/tiff/67.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_06/Issue_01/tiff/
                                        68.tif.

Collison:1963:NMF

[Col63b] D. M. Collison. Note on a method of forming a sorting key for a partly ordered list, and an application. The Computer Journal, 6(1):74, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/060074.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/tiff/74.tif.

Cole:1964:PET

[Col64a]  A. J. Cole.  The prepa-

ration of examination timetables using a small-store computer. The Computer Journal, 7(2):117–121, July 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/070117.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/117.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/118.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/119.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/120.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/121.tif.

Colin:1964:MVC

[Col64b] A. Colin. The multiple variate counter. The Computer Journal, 6(4):339–347, January 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/060339.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/339.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/340.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/341.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/342.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/343.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/344.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/345.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/346.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/347.tif.

Cole:1966:PSP

[Col66a] A. J. Cole. Plane and stereographic projections of convex polyhedra from minimal information. The Computer Journal, 9(1):27–31, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/090027.sgm.abs.html;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/27.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_01/tiff/
28.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_09/Issue_01/tiff/29.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
09/Issue_01/tiff/30.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/31.tif.

Colin:1966:SPU

[Col66b] A. J. T. Colin. A simple program for use in the ‘conversational mode’. The Computer Journal, 9(3):238–241, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/090238.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/238.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/239.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/240.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/241.tif.

Cole:1967:CCN

[Col67]  A. J. Cole.  Correspon-

dence: Chromatic numbers and timetabling problems. The Computer Journal, 10(2):

217, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        10/Issue_02/tiff/217.tif.

See [WP67].

Colin:1970:DTF

[Col70] A. J. T. Colin. DOLPHIN — a text filing system for University use. The Computer Journal, 13(2):136–141, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/130136.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/tiff/136.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/tiff/137.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/tiff/138.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/tiff/139.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/tiff/140.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/tiff/141.tif.

Colbourn:1982:CCI

[Col82] Charles J. Colbourn. Computing the chromatic index of Steiner triple systems. The Computer Journal, 25 (3):338–339, August 1982.

CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/338.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/339.tif.

Cole:1987:CTR

[Col87] A. J. Cole. Compaction techniques for raster scan graphics using space-filling curves. The Computer Journal, 30 (1):87–92, February 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
01/tiff/87.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_01/tiff/
88.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_30/Issue_01/tiff/89.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
30/Issue_01/tiff/90.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
01/tiff/91.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_01/tiff/
92.tif.

Colomb:1997:ISH

[Col97] R. M. Colomb. Impact of semantic heterogeneity on federating databases. The Computer Journal, 40(5):235–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www.oup.co.uk/computer_
                                        journal/Volume_40/Issue_05/
                                        Vol40_05.body.html#AbstractColomb;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_40/Issue_05/
                                        Vol40_05.body.html#AbstractColomb.

Comerford:1975:LSMa

[Com75a] P. J. Comerford. Link systems for multi-computer control of a large process. I. The Computer Journal, 18(2):144–147, May 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        02/tiff/144.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        02/tiff/145.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        02/tiff/146.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        02/tiff/147.tif.

Comerford:1975:LSMb

[Com75b] P. J. Comerford. Link systems for multi-computer control of a large process: Part 2: Fault detection. The Computer Journal, 18(3):220–222, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/220.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/221.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/222.tif.

Compton:1985:AED

[Com85] A. J. Compton. An algorithm for the even distribution of entities in one dimension. The Computer Journal, 28(5):530–537, November 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/530.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/531.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/532.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/533.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/534.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/535.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/536.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/537.tif.

Constantinescu:1966:CSE

[Con66] P. Constantinescu. The classification of a set of elements with respect to a set of properties. The Computer Journal, 8(4):352–357, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/080352.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/352.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/353.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/354.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/355.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/356.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/357.tif.

Cook:1960:MSS

[Coo60a] R. L. Cook. Market surveys with a small computer. The Computer Journal, 3 (3):140–142, October 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/030140.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/140.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/141.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_03/Issue_
03/tiff/142.tif.

Cook:1960:TSN

[Coo60b] R. L. Cook. Time-sharing on the National-Elliott 802. The Computer Journal, 2 (4):185–188, January 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/020185.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/185.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/186.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/187.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/188.tif.

Cooper:1966:ECC

[Coo66] D. C. Cooper. The equivalence of a certain computations. The Computer Journal, 9(1):45–52, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/090045.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/45.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_01/tiff/
46.tif;   http://www3.oup.
co.uk/computer_journal/hdb/

Volume_09/Issue_01/tiff/47.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
09/Issue_01/tiff/48.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/49.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_01/tiff/
50.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_09/Issue_01/tiff/51.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
09/Issue_01/tiff/52.tif.

Cooper:1967:ECC

[Coo67] G. J. Cooper. The evaluation of the coefficients in the Chebyshev expansion. The Computer Journal, 10(1):94–100, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/100094.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/100.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/94.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_01/tiff/
95.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_10/Issue_01/tiff/96.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/97.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/98.tif;   http://

www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_01/tiff/
99.tif.

Cooper:1968:BSI

[Coo68] B. E. Cooper. Basic subroutine for the input of numbers, words and special characters. The Computer Journal, 11(2):157–159, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/110157.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/157.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/158.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/159.tif.

Cook:1972:CPF

[Coo72a]  D. B. Cook.  Correspon-

dence: ‘Packing’ in FOR-TRAN. The Computer Journal, 15(4):361, November 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/361.tif.

Cooper:1972:CQI

[Coo72b] W. Cooper. Correspondence:

On “A quasi-intrinsic scheme for passing a smooth curve through a discrete set of points”. The Computer Journal, 15(3):246, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/150246.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/246.tif. See [McC70b,

Ecc71, Smi71, Sab72].

Cooper:1975:TME

[Coo75] B. Cooper. Time-domain matrix evaluation of a digital servomechanism using semihybrid computational techniques. The Computer Journal, 18(2):169–172, May 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/169.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/170.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/171.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/172.tif.

Cookson:1983:TSC

[Coo83] M. J. Cookson. Taxonomic studies on current approaches to systems analysis. The Computer Journal, 26(3):283–284, August 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/283.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/284.tif.

Cook:1989:ECO

[Coo89a] W. R. Cook, editor. European Conference of Object-Oriented Programming (ECOOP ’89), volume 32(4) of The Computer Journal. British Computer Society, London, UK, August 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Cook:1989:PME

[Coo89b] W. R. Cook. A proposal for making Eiffel type-safe. The Computer Journal, 32 (4):305–311, August 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/305.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/306.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/307.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/308.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/309.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/310.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/311.tif.

Cooper:1989:FHO

[Coo89c] M. C. Cooper. Formal hierarchical object models for fast template matching. The Computer Journal, 32(4):351–361, August 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/351.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/352.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/353.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/354.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/355.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/356.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/357.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/358.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/359.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/360.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_32/Issue_
04/tiff/361.tif.

Cooke:1992:EFM

[Coo92a] J. Cooke. Editorial – formal methods: What? why? and when? The Computer Journal, 35(5):417–418, October 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_35/Issue_
05/tiff/417.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_35/Issue_
05/tiff/418.tif.

Cooke:1992:FMM

[Coo92b] J. Cooke. Formal methods — mathematics, theory, recipes or what? The Computer Journal, 35(5):419–423, October 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_35/Issue_05/
Vol35_05.body.html#AbstractCooke.

See editorial [Coo92a].

Cormack:1962:EOE

[Cor62] A. S. Cormack. Early operating experience with Language H. The Computer Journal, 5(3):158–161, October 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/050158.sgm.abs.html;
http://www3.oup.co.uk/computer_

journal/hdb/Volume_05/Issue_
03/tiff/158.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/159.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/160.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/161.tif.

Cormack:1965:IBU

[Cor65]  A. S. Cormack.  Interac-

tion between user’s needs and language-compiler-computer systems. The Computer Journal, 8(1):8–12, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
01/080008.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
01/tiff/10.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_01/tiff/
11.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_08/Issue_01/tiff/12.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
08/Issue_01/tiff/8.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
01/tiff/9.tif.

Corner:1979:BRM

[Cor79]  Lewis Corner.  Book re-

views: Model Building in Mathematical Programming, by H. P. Williams, 1978; Computer Simulation and Modeling: An Introduction, by R. S. Lehman. The Computer Journal, 22(1):2, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
01/tiff/2.tif.

Cowling:1986:TCP

[Cow86] A. J. Cowling. Type checking in polymorphic languages. The Computer Journal, 29 (6):538–544, December 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/538.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/539.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/540.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/541.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/542.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/543.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/544.tif.

Cox:1970:BTC

[Cox70] M. G. Cox. A bracketing technique for computing a zero of a function. The Computer Journal, 13(1):101–102, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        01/130101.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        01/tiff/101.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        01/tiff/102.tif. See note

[SC71].

Cox:1971:AAC

[Cox71] M. G. Cox. An algorithm for approximating convex functions by means of first degree splines. The Computer Journal, 14(3):272–275, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/140272.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/tiff/272.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/tiff/273.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/tiff/274.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/tiff/275.tif.

Clowes:1961:NTA

[CP61] M. B. Clowes and J. R. Parks. A new technique in automatic character recognition. The Computer Journal, 4(2):121–128, July 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/040121.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/121.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/122.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/123.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/124.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/125.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/126.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/127.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/128.tif.

Curtis:1962:PTL

[CP62] A. R. Curtis and I. C. Pyle.

A proposed target language for compilers on Atlas. The Computer Journal, 5(2):100–106, July 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_

                                        02/050100.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        02/tiff/100.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        02/tiff/101.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        02/tiff/102.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        02/tiff/103.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        02/tiff/104.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        02/tiff/105.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        02/tiff/106.tif.

Clowes:1964:AP

[CP64] J. S. Clowes and E. S. Page. Assignment problems. The Computer Journal, 6(4):304–307, January 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        04/060304.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        04/tiff/304.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        04/tiff/305.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        04/tiff/306.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_06/Issue_
04/tiff/307.tif.

Curtis:1966:NCM

[CP66a] A. R. Curtis and M. J. D.

Powell. Necessary conditions for a minimax approximation. The Computer Journal, 8(4):358–361, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/080358.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/358.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/359.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/360.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/361.tif.

Curtis:1966:CEA

[CP66b] A. R. Curtis and M. J. D.

Powell. On the convergence of exchange algorithms for calculating minimax approximations. The Computer Journal, 9(1):78–80, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/78.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_01/tiff/
79.tif;   http://www3.oup.

co.uk/computer_journal/hdb/
Volume_09/Issue_01/tiff/80.
tif.

Cranley:1968:EMI

[CP68] R. Cranley and T. N. L. Patterson. The evaluation of multidimensional integrals. The Computer Journal, 11(1):102–110, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/110102.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/102.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/103.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/104.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/105.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/106.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/107.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/108.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/109.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/110.tif.

Cranley:1971:ANE

[CP71] R. Cranley and T. N. L. Patterson. On the automatic numerical evaluation of definite integrals. The Computer Journal, 14(2):189–198, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/140189.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/189.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/190.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/191.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/192.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/193.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/194.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/195.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/196.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/197.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/198.tif. See corre-

spondence [Val71].

Chattergy:1979:DFC

[CP79] R. Chattergy and U. W. Pooch. A distributed function computer with dedicated processors. The Computer Journal, 22(1):37–40, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        01/tiff/37.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_22/Issue_01/tiff/
                                        38.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_22/Issue_01/tiff/39.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        22/Issue_01/tiff/40.tif.

Castle:1987:EST

[CP87] C. M. A. Castle and M. L. V. Pitteway. An efficient structural technique for encoding ‘bestfit’ straight lines. The Computer Journal, 30(2):168–175, April 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        02/tiff/168.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        02/tiff/169.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        02/tiff/170.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        02/tiff/171.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_30/Issue_
02/tiff/172.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/173.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/174.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/175.tif.

Chung-Phillips:1975:NDD

[CPR75] Alice Chung-Phillips and R. W.

Rosen. A note on dynamic data storage in FORTRAN IV. The Computer Journal, 18(4):342–343, November 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/342.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/343.tif.

Catania:1993:MPE

[CPV93] V. Catania, A. Puliafito, and L. Vita. A model for performance evaluation of gracefully degrading systems. The Computer Journal, 36(2):177–185, April 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_02/
Vol36_02.body.html#AbstractCatania.

Cr0xf0rd:1981:CZT

[Cr081] A. A. Cr0xf0rd. Correspondence: Os, zeros, thetas, and pis. The Computer Journal, 24(1):95, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
01/tiff/95.tif. The use of

the digit zero, rather than the letter oh, in the author’s name is intentional.

Cowdrey:1963:AMC

[CR63] D. R. Cowdrey and C. M. Reeves. An application of the Monte Carlo method to the evaluation of some molecular integrals. The Computer Journal, 6(3):277–286, October 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/060277.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/277.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/278.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/279.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/280.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/281.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/282.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_06/Issue_
03/tiff/283.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/284.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/285.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/286.tif.

Clark:1983:GIP

[CR83] B. E. J. Clark and S. K.

Robinson. A graphically interacting program monitor. The Computer Journal, 26 (3):235–238, August 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/235.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/236.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/237.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/238.tif.

Challab:1987:BA

[CR87] D. J. Challab and J. D.

Roberts. Buddy algorithms. The Computer Journal, 30 (4):308–315, August 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_

                                        04/tiff/308.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        04/tiff/309.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        04/tiff/310.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        04/tiff/311.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        04/tiff/312.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        04/tiff/313.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        04/tiff/314.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        04/tiff/315.tif.

Chandy:1997:SCD

[CR97] K. Mandy Chandy and Adam Rikfin. Systematic composition of distributed objects: Processes and sessions. The Computer Journal, 40(8):465–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_40/Issue_08/
                                        Vol40_08.body.html#AbstractChandy.

Craddock:1966:PSU

[Cra66a] J. M. Craddock. Programming systems for use in research computation. The Computer Journal, 8(4):315–318, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/080315.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/315.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/316.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/317.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/318.tif.

Craig:1966:CCR

[Cra66b] T. K. Craig. CRESTS — Courtauld’s rapid extract, sort and tabulate system. The Computer Journal, 9(1):3–10, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/090003.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/10.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_01/tiff/
3.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_09/Issue_01/tiff/4.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
09/Issue_01/tiff/5.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/6.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_09/Issue_01/tiff/
7.tif;   http://www3.oup.

                                        co.uk/computer_journal/hdb/
                                        Volume_09/Issue_01/tiff/8.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        09/Issue_01/tiff/9.tif.

Craddock:1967:CAC

[Cra67] I. D. Craddock. Correspondence: On analogue computers. The Computer Journal, 10(2):217, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        02/tiff/217.tif.

Crane:1972:ACP

[Cra72] C. M. Crane. Algorithm 75: Contour plotting for functions specified at nodal points of an irregular mesh based on an arbitrary two-parameter coordinate system. The Computer Journal, 15(4):382–384, November 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/tiff/382.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/tiff/383.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/tiff/384.tif.

Craig:1986:ABS

[Cra86] I. D. Craig. The Ariadne-1 blackboard system. The Computer Journal, 29(3):235–240, June 1986. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/235.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/236.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/237.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/238.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/239.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/240.tif.

Cross:1974:NSD

[Cro74] M. J. Cross. Nimbus 4 SCR data handling. The Computer Journal, 17(2):172–180, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/172.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/173.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/174.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/175.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/176.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_17/Issue_
                                        02/tiff/177.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        02/tiff/178.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        02/tiff/179.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        02/tiff/180.tif.

Crombie:1975:CCH

[Cro75] D. B. Crombie. Correspondence: Comparison of hybrid and digital computation performances for distillation simulation. The Computer Journal, 18(3):285, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/285.tif.

Crowe:1980:PVO

[Cro80] M. K. Crowe. Porting virtual object files. The Computer Journal, 23(3):287–288, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        03/tiff/287.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        03/tiff/288.tif.

Clark:1993:SLC

[CRS^{+}93] D. A. Clark, C. J. Rawlings, J. Shirazi, L.-L. Li, K. Schuerman, M. Reeve, and A. Veron. Solving large combinatorial problems in molecular biology using the ElipSys parallel constraint logic programming system. The Computer Journal, 36(8):690–701, ???? 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_08/
Vol36_08.body.html#AbstractClark.

Cardenas:1975:SDS

[CS75] A. F. Cardenas and R. W.

Seeley. A simple data structure for interactive graphic design/drafting. The Computer Journal, 18(1):30–33, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
01/tiff/30.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_01/tiff/
31.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_18/Issue_01/tiff/32.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
18/Issue_01/tiff/33.tif.

Cardenas:1977:DTD

[CS77] Alfonso F. Cardenas and James P. Sagamang. Doublychained tree data base organisation — analysis and design strategies. The Computer Journal, 20(1):15–26, February 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
01/tiff/15.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_01/tiff/
16.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_20/Issue_01/tiff/17.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
20/Issue_01/tiff/18.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
01/tiff/19.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_01/tiff/
20.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_20/Issue_01/tiff/21.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
20/Issue_01/tiff/22.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
01/tiff/23.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_01/tiff/
24.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_20/Issue_01/tiff/25.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
20/Issue_01/tiff/26.tif.

Cesarini:1985:DIR

[CS85] F. Cesarini and G. Soda. Design and implementation of a relational DBMS for microcomputers. The Computer Journal, 28(4):391–397, August 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

journal/hdb/Volume_28/Issue_
04/tiff/391.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/392.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/393.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/394.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/395.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/396.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/397.tif.

Chatterjee:1995:DPS

[CS95] Saurav Chatterjee and Jay Strosnider. Distributed pipeline scheduling: a framework for distributed, heterogeneous realtime system design. The Computer Journal, 38(4):271–285, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_38/Issue_04/
Vol38_04.body.html#AbstractChatterjee.

Coates:1974:CSB

[CT74] H. N. Coates and The Editor. Correspondence: The silent business user. The Computer Journal, 17(1):96, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

                                        journal/hdb/Volume_17/Issue_
                                        01/tiff/96.tif. See [TT73].

Czejdo:1992:IIS

[CT92] B. Czejdo and M. C. Taylor. Integration of information systems using an object-oriented approach. The Computer Journal, 35(5):501–513, October 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_35/Issue_05/
                                        Vol35_05.body.html#AbstractCzejdo.

Cai:1994:ART

[CT94] Wentong Cai and Stephen J. Turner. An approach to the run-time monitoring of parallel programs. The Computer Journal, 37(4):333–345, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_37/Issue_04/
                                        Vol37_04.body.html#AbstractCai.

Cleary:1997:ULC

[CT97] J. G. Cleary and W. J. Teahan. Unbounded length contexts for PPM. The Computer Journal, 40(2/3):67–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www.oup.co.uk/computer_
                                        journal/Volume_40/Issue_02/
                                        Vol40_02.body.html#AbstractCleary;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_40/Issue_02/
                                        Vol40_02.body.html#AbstractCleary.

Chi:1999:FMM

[CTH99] Kuang-Hwei Chi, Chien-Chao Tseng, and Ting-Lu Huang. A framework for mobile multicast using dynamic route reconstructions. The Computer Journal, 42(6):522–533, 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
06/420522.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
06/pdf/420522.pdf.

Chan:1994:EOO

[CTW94] D. K. C. Chan, P. W. Trinder, and R. C. Welland. Evaluating object-oriented query languages. The Computer Journal, 37(10):858–872, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_
10/tiff/858.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_
10/tiff/859.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_
10/tiff/860.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_
10/tiff/861.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_
10/tiff/862.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_

10/tiff/863.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_
10/tiff/864.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_
10/tiff/865.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_
10/tiff/866.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_
10/tiff/867.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_
10/tiff/868.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_
10/tiff/869.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_
10/tiff/870.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_
10/tiff/871.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_
10/tiff/872.tif;   http:/
/www3.oup.co.uk/computer_
journal/Volume_37/Issue_10/
Vol37_10.body.html#AbstractChan.

Cucker:1993:CQE

[Cuc93] Felipe Cucker. On the complexity of quantifier elimination: the structural approach. The Computer Journal, 36(5):400–408, October 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_05/

Vol36_05.body.html#AbstractCucker.

Cuff:1972:CCF

[Cuf72]  R. N. Cuff.  A conver-

sational compiler for full PL/I. The Computer Journal, 15(2):99–104, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/150099.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/100.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/101.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/102.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/103.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/104.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/99.tif.

Curzon:1995:TDC

[Cur95] Paul Curzon. Tracking design changes with formal machinechecked proof. The Computer Journal, 38(2):91–100, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_38/Issue_02/
Vol38_02.body.html#AbstractCurzon.

Cuttle:1959:WNS

[Cut59] G. Cuttle. Where next? Some conjectures on the future of the large-scale computer in integrated commercial work. The Computer Journal, 2(2):85–86, July 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
02/020085.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
02/tiff/85.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_02/tiff/
86.tif.

Chen:1986:DAC

[CV86] W. C. Chen and J. S. Vitter. Deletion algorithms for coalesced hashing. The Computer Journal, 29(5):436–450, October 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/436.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/437.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/438.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/439.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/440.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_29/Issue_
05/tiff/441.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/442.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/443.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/444.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/445.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/446.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/447.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/448.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/449.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/450.tif.

Cadwell:1961:SOM

[CW61] J. H. Cadwell and D. E.

Williams. Some orthogonal methods of curve and surface fitting. The Computer Journal, 4(3):260–264, October 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/040260.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_

                                        03/tiff/260.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        03/tiff/261.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        03/tiff/262.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        03/tiff/263.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        03/tiff/264.tif.

Cooper:1962:AAL

[CW62] D. C. Cooper and H. Whitfield. ALP: an autocode listprocessing language. The Computer Journal, 5(1):28–32, April 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        01/050028.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        01/tiff/28.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_05/Issue_01/tiff/
                                        29.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_05/Issue_01/tiff/30.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        05/Issue_01/tiff/31.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        01/tiff/32.tif.

Cole:1970:IAJ

[CW70] A. J. Cole and D. Wishart. An improved algorithm for the Jardine-Sibson method of generating overlapping clusters. The Computer Journal, 13(2):156–163, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/130156.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/156.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/157.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/158.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/159.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/160.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/161.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/162.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/163.tif.

Coggan:1971:LSE

[CW71] G. C. Coggan and J. A. Wilson.

On-line state estimation with a small computer. The Computer Journal, 14(1):61–64, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_

                                        01/140061.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        01/tiff/61.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_14/Issue_01/tiff/
                                        62.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_14/Issue_01/tiff/63.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        14/Issue_01/tiff/64.tif.

Curnow:1976:SB

[CW76] H. J. Curnow and B. A. Wichmann. A synthetic benchmark. The Computer Journal, 19(1):43–49, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        01/tiff/43.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_19/Issue_01/tiff/
                                        44.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_19/Issue_01/tiff/45.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        19/Issue_01/tiff/46.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        01/tiff/47.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_19/Issue_01/tiff/
                                        48.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_19/Issue_01/tiff/49.
                                        tif.

Corbett:1982:IPD

[CW82] Chris Corbett and Ian H. Witten. On the inclusion and placement of documentation graphics in computer typesetting. The Computer Journal, 25(2):272–277, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/272.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/273.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/274.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/275.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/276.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/277.tif.

Chen:1996:TBA

[CW96a] I.-R. Chen and C.-M. Wang.

Threshold-based admission control policies for multimedia servers. The Computer Journal, 39(9):757–??, ???? 1996. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www.oup.co.uk/computer_
journal/Volume_39/Issue_09/
Vol39_09.body.html#AbstractChen1;
 http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_09/
Vol39_09.body.html#AbstractChen1.

Chen:1996:ARD

[CW96b] I.-R. Chen and D.-C Wang. Analysis of replicated data with repair dependency. The Computer Journal, 39(9):767–??, ???? 1996. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www.oup.co.uk/computer_
                                        journal/Volume_39/Issue_09/
                                        Vol39_09.body.html#AbstractChen2;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_39/Issue_09/
                                        Vol39_09.body.html#AbstractChen2.

Cavalcanti:1998:WPS

[CW98] Ana Cavalcanti and Jim Woodcock. A weakest precondition semantics for Z. The Computer Journal, 41(1):1–15, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_41/Issue_01/
                                        Vol41_01.body.html#AbstractCavalcanti.

Chao:1993:CTM

[CWH93] Daniel Y. Chao, David T. Wang, and Douglas D. C. Hung. Convergence time and memory capacity of higherorder Hopfield associative memory with multi-valued neurons. The Computer Journal, 36(6):554–561, December 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_36/Issue_06/
                                        Vol36_06.body.html#AbstractChao.

Chen:1995:APE

[CY95] Ing-Ray Chen and I.-Ling Yen.

Analysis of probabilistic error checking procedures on storage systems. The Computer Journal, 38(5):348–354, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_38/Issue_05/
Vol38_05.body.html#AbstractChen.

Chen:1996:CSP

[CY96] Tsong Yueh Chen and Yuen Tak Yu. Constraints for safe partition testing strategies. The Computer Journal, 39(7):619– 625, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www.oup.co.uk/jnls/
list/comjnl/hdb/Volume_39/
Issue_07/390619.sgm.abs.html;
 http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_07/
Vol39_07.body.html#AbstractChen.

Czech:1988:EID

[Cze88] Z. J. Czech. Efficient implementation of detection of undefined variables. The Computer Journal, 31(6):545–549, December 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/545.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/546.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_31/Issue_
                                        06/tiff/547.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        06/tiff/548.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        06/tiff/549.tif.

Czech:1998:QPH

[Cze98] Zbigniew J. Czech. Quasiperfect hashing. The Computer Journal, 41(6):416–421, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_41/Issue_06/
                                        Vol41_06.body.html#AbstractCzech.

Cierniak:1997:CTS

[CZL97] Michal Cierniak, Mohammed Javeed Zaki, and Wei Li. Compiletime scheduling algorithms for a heterogeneous network of workstations. The Computer Journal, 40(6):356–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www.oup.co.uk/computer_
                                        journal/Volume_40/Issue_06/
                                        Vol40_06.body.html#AbstractCierniak;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_40/Issue_06/
                                        Vol40_06.body.html#AbstractCierniak.

Chao:1994:EKT

[CZW94] Daniel Y. Chao, Mengchu Zhou, and David T. Wang. Extending the knitting technique to Petri Net synthesis of automated manufacturing systems. The Computer Journal, 37(1):67–76, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_37/Issue_01/
Vol37_01.body.html#AbstractChao.

dAgapeyeff:1962:CDC

[d’A62] A. d’Agapeyeff. Current developments in commercial automatic programming. The Computer Journal, 5(2):107–111, July 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/050107.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/107.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/108.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/109.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/110.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/111.tif.

Delves:1977:AFG

[DA77] L. M. Delves and L. F. Abd.

El Al. Algorithm 97: The fast Galerkin algorithm for the solution of linear Fredholm equations. The Computer Journal, 20(4):374–376, November 1977. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/374.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/375.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/376.tif.

Dace:1963:EPU

[Dac63] D. J. Dace. Experience in the practical use of data transmission. The Computer Journal, 6(1):17–23, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/060017.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/tiff/17.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/
18.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_06/Issue_01/tiff/19.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
06/Issue_01/tiff/20.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/tiff/21.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/
22.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_06/Issue_01/tiff/23.
tif.

Delves:1981:SMS

[DAEH81] L. M. Delves, L. F. Abd-Elal, and J. A. Hendry. Set of modules for the solution of integral equations. The Computer Journal, 24(2):184–190, May 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/184.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/185.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/186.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/187.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/188.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/189.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/190.tif.

Dakin:1965:TSA

[Dak65] R. J. Dakin. A tree search algorithm for mixed integer programming problems. The Computer Journal, 8 (3):250–255, October 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/080250.sgm.abs.html;

                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        03/tiff/250.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        03/tiff/251.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        03/tiff/252.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        03/tiff/253.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        03/tiff/254.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        03/tiff/255.tif.

Dakin:1973:CIF

[Dak73]  R. J. Dakin.  Correspon-

dence: Impressions following the BCS Conference on Computer Performance. The Computer Journal, 16(1):18, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        01/tiff/18.tif. See corre-

spondence [Gla73].

Dakin:1975:GCL

[Dak75] R. J. Dakin. A general control language: language structure and translation. The Computer Journal, 18(4):324–332, November 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        04/tiff/324.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/325.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/326.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/327.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/328.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/329.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/330.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/331.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/332.tif.

Dandeker:1992:IPV

[Dan92]  K. N. Dandeker.  Imple-

mentation of parameterised views in relational query languages with a macro processor. The Computer Journal, 35(5):A425–??, October 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_35/Issue_05/
35000A425.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/Volume_35/Issue_05/
tiff/A425.tif.

Deen:1985:AGD

[DAODT85] S. M. Deen, R. R. Amin, G. O.

Ofori-Dwumfuo, and M. C. Taylor. Architecture of a generalised distributed database system — PRECI. The Computer Journal, 28(3):282–290, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/282.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/283.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/284.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/285.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/286.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/287.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/288.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/289.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/290.tif.

Daszczuk:1991:SSD

[Das91] W. B. Daszczuk. A structured semantic design of distributed operating systems. The Computer Journal, 34 (6):482–492, December 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/482.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/483.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/484.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/485.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/486.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/487.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/488.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/489.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/490.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/491.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/492.tif.

Deen:1987:IPP

[DAT87] S. M. Deen, R. R. Amin, and M. C. Taylor. Implementation of a prototype for PRECI*. The Computer Journal, 30(2):157–162, April 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        02/tiff/157.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        02/tiff/158.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        02/tiff/159.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        02/tiff/160.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        02/tiff/161.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        02/tiff/162.tif.

Davey:1963:DPT

[Dav63] L. B. Davey. Data preparation and transmission in the Royal Air Force integrated supply system. The Computer Journal, 6(3):219–222, October 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        03/060219.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        03/tiff/219.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        03/tiff/220.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        03/tiff/221.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        03/tiff/222.tif.

Davison:1967:HOC

[Dav67]  E. J. Davison.  A high-

order Crank-Nicholson technique for solving differential equations. The Computer Journal, 10(2):195–197, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/100195.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/195.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/196.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/197.tif.

Davidon:1968:VAM

[Dav68]  W. C. Davidon.  Variance

algorithm for minimization. The Computer Journal, 10 (4):406–410, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/100406.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/406.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/407.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/408.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_10/Issue_
                                        04/tiff/409.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/tiff/410.tif.

Davenport:1976:DI

[Dav76] R. A. Davenport. Database integrity. The Computer Journal, 19(2):110–116, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/110.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/111.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/112.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/113.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/114.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/115.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/116.tif.

Davenport:1978:DCD

[Dav78]  R. A. Davenport.  Dis-

tributed or centralised data base. The Computer Journal, 21(1):7–14, February 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_21/Issue_
01/tiff/10.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_01/tiff/
11.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_21/Issue_01/tiff/12.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
21/Issue_01/tiff/13.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
01/tiff/14.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_01/tiff/
7.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_21/Issue_01/tiff/8.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
21/Issue_01/tiff/9.tif.

Davenport:1981:DDD

[Dav81]  R. A. Davenport.  Design

of distributed data base systems. The Computer Journal, 24(1):31–41, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
01/tiff/31.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_01/tiff/
32.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_24/Issue_01/tiff/33.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
24/Issue_01/tiff/34.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
01/tiff/35.tif;   http://

                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_24/Issue_01/tiff/
                                        36.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_24/Issue_01/tiff/37.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        24/Issue_01/tiff/38.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        01/tiff/39.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_24/Issue_01/tiff/
                                        40.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_24/Issue_01/tiff/41.
                                        tif.

Davies:1982:NSF

[Dav82] D. J. M. Davies. A note on sparsely filled dynamically allocated memory. The Computer Journal, 25(1):159, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/159.tif.

Davis:1984:NDF

[Dav84] R. H. Davis. A note on dynamic file organization models. The Computer Journal, 27(3):284, August 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        03/tiff/284.tif.

Davis:1992:FRS

[Dav92] I. J. Davis. A fast radix sort. The Computer Journal, 35(6):636–642, December 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_06/
Vol35_06.body.html#AbstractDavis.

Dawson:1973:CIC

[Daw73]  J. L. Dawson.  Combining

interpretive code with machine code. The Computer Journal, 16(3):216–219, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/160216.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/216.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/217.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/218.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/219.tif. See corre-

spondence [Mac74, Bem74].

Dawkins:1974:CCJ

[Daw74] B. W. Dawkins. Correspondence: The Computer Journal. The Computer Journal, 17(2):190, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_17/Issue_
                                        02/tiff/190.tif.

Dawid:1999:DPR

[Daw99] A. P. Dawid. Discussion of the papers by Rissanen and by Wallace and Dowe. The Computer Journal, 42(4):323–326, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        04/420323.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        04/pdf/420323.pdf.

Day:1965:OSM

[Day65] J. T. Day. A one-step method for the numerical integration of the differential equation y′ = f(x)y + g(x). The Computer Journal, 7(4):314–317, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        04/070314.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        04/tiff/314.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        04/tiff/315.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        04/tiff/316.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        04/tiff/317.tif.

Day:1966:RKM

[Day66] J. T. Day. A Runge–Kutta method for the numerical solution of the Goursat problem in hyperbolic partial differential equations. The Computer Journal, 9(1):81–83, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/090081.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/81.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_01/tiff/
82.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_09/Issue_01/tiff/83.
tif.

Day:1967:NNS

[Day67] J. T. Day. Note on the numerical solution of integrodifferential equations. The Computer Journal, 9(4):394– 395, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
09/Issue_04/090394.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
09/Issue_04/tiff/394.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/395.tif.

Day:1969:ASA

[Day69]  A. C. Day.  Algorithm

38: A searching algorithm. The Computer Journal, 12 (1):101–102, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
01/tiff/101.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
01/tiff/102.tif.

Day:1970:USS

[Day70] A. C. Day. The use of symbolstate tables. The Computer Journal, 13(4):332–339, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/130332.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/332.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/333.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/334.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/335.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/336.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/337.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/338.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/339.tif. See note

[Lan71].

Day:1976:ABB

[Day76]  A. C. Day.  Algorithm

91: Balancing a binary tree. The Computer Journal, 19 (4):360–361, November 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/360.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/361.tif.

Day:1977:AGP

[Day77] A. C. Day. Algorithm 95:

Generating a parity testing table. The Computer Journal, 20(3):280, August 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/280.tif.

Day:1978:ASS

[Day78] A. C. Day. Algorithms supplement: Statement of policy. The Computer Journal, 21(4):377, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/377.tif.

Dutton:1981:NGC

[DB81] R. D. Dutton and R. C. Brigham. A new graph colouring algorithm. The Computer Journal, 24(1):85–86, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
01/tiff/85.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_01/tiff/
86.tif.

Dearle:1988:SBS

[DB88] A. Dearle and A. L. Brown. Safe browsing in a strongly typed persistent environment. The Computer Journal, 31 (6):540–544, December 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/540.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/541.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/542.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/543.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/544.tif.

Davies:1990:TLP

[DB90] G. L. Davies and A. Burns. The teaching language Pascal-FC. The Computer Journal, 33(2):147–154, April 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/147.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/148.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/149.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/150.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/151.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/152.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/153.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/154.tif.

Dietrich:1992:PID

[DBCRW92] S. W. Dietrich, M. Brown, E. Cortes-Rello, and S. Wunderlin. A practitioner’s introduction to database performance benchmarks and measurements. The Computer Journal, 35(4):322–331, August 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_35/Issue_04/
Vol35_04.body.html#AbstractDietrich.

Djordjevic:1985:PLN

[DBH85] Jovan Djordjevic, Mario Barbacci, and Brad Hosler. A PMS level notation for the description and simulation of digital systems. The Computer Journal, 28(4):357–365, August 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        04/tiff/357.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        04/tiff/358.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        04/tiff/359.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        04/tiff/360.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        04/tiff/361.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        04/tiff/362.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        04/tiff/363.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        04/tiff/364.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        04/tiff/365.tif. PMS.

Davis:1984:DFO

[DC84] R. H. Davis and P. Coumpas. A dynamic file organization model. The Computer Journal, 27(2):143–150, May 1984.

CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/143.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/144.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/145.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/146.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/147.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/148.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/149.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/150.tif.

Dromey:1990:SCA

[DC90a] R. G. Dromey and T. A. Chorvat. Structure clashes — an alternative to program inversion. The Computer Journal, 33(2):126–132, April 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/126.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/127.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_33/Issue_
                                        02/tiff/128.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        02/tiff/129.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        02/tiff/130.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        02/tiff/131.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        02/tiff/132.tif.

Dueck:1990:MAG

[DC90b] Gerald D. P. Dueck and Gordon V. Cormack. Modular attribute grammars. The Computer Journal, 33(2):164–172, April 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        02/tiff/164.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        02/tiff/165.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        02/tiff/166.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        02/tiff/167.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        02/tiff/168.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        02/tiff/169.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        02/tiff/170.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/171.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/172.tif. See also: re-

search report CS-88-19, University of Waterloo (May 1988).

Day:1976:PNS

[DCHR76] A. C. Day, P. A. Clarke, D. Hill, and J. K. Reid. The proposed new standard for FORTRAN: a critical examination. The Computer Journal, 19(3):268–271, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/268.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/269.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/270.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/271.tif.

dAgapeyeff:1966:CSC

[dD66] A. d’Agapeyeff and F. G. Duncan. Correspondence: On the stability of Crank and Nicholson’s procedure. The Computer Journal, 9(3):320– 320, 324–324, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_

                                        03/090320.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        03/tiff/320.tif. See [PC64,

KM66].

Dawkins:1970:CUF

[DD70] B. W. Dawkins and H. C. Dawkins. Correspondence: On “A University faculty timetable”. The Computer Journal, 13(1):121–122, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        01/tiff/121.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        01/tiff/122.tif. See [Alm69].

Dvorak:1987:SLT

[DD87] S. Dvorak and B. Durian. Stable linear time sublinear space merging. The Computer Journal, 30(4):372–375, August 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        04/tiff/372.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        04/tiff/373.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        04/tiff/374.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        04/tiff/375.tif.

Dvorak:1988:MDR

[DĎ88a] S. Dvořák and B. Ďurian. Merging by decomposition revisited. The Computer Journal, 31(6):553–556, December 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/553.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/554.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/555.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/556.tif.

Dvorak:1988:ULT

[DD88b] S. Dvorak and B. Durian. Unstable linear time O(1) space merging. The Computer Journal, 31(3):279–282, June 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/279.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/280.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/281.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/282.tif.

Dvorak:1991:DMS

[DD91] S. Dvorak and B. Durian. Decomposition merging with symmetric exchanges. The Computer Journal, 34(6):A029– A033, December 1991. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

DePrisco:1998:DEH

[DdS98] Roberto De Prisco and Alfredo de Santis. On the data expansion of the Huffman compression algorithm. The Computer Journal, 41(3):137–144, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_41/Issue_03/
                                        Vol41_03.body.html#AbstractDePrisco.

Dowsing:1986:PBB

[DE86] R. D. Dowsing and R. Elliott. Programming a bounded buffer using the object and path expression constructs of Path Pascal. The Computer Journal, 29(5):423–429, October 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        05/tiff/423.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        05/tiff/424.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        05/tiff/425.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_

05/tiff/426.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/427.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/428.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/429.tif.

Dearnley:1974:MSD

[Dea74a] P. A. Dearnley. A model of a self-organising data management system. The Computer Journal, 17(1):13–16, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
01/tiff/13.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/
14.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_17/Issue_01/tiff/15.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
17/Issue_01/tiff/16.tif.

Dearnley:1974:OMS

[Dea74b] P. A. Dearnley. The operation of a model self organising data management system. The Computer Journal, 17(3):205–210, August 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/205.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_17/Issue_
                                        03/tiff/206.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        03/tiff/207.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        03/tiff/208.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        03/tiff/209.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        03/tiff/210.tif.

Dearnley:1976:IDR

[Dea76]  P. A. Dearnley.  An in-

vestigation into database resilience. The Computer Journal, 19(2):117–121, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/117.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/118.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/119.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/120.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/121.tif.

Dearnley:1978:MDS

[Dea78]  P. Dearnley.  Monitoring

database system performance. The Computer Journal, 21 (1):15–19, February 1978.

CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
01/tiff/15.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_01/tiff/
16.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_21/Issue_01/tiff/17.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
21/Issue_01/tiff/18.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
01/tiff/19.tif.

Dearnley:1979:ALM

[Dea79]  P. A. Dearnley.  Appli-

cation level microcode to speed data base management. The Computer Journal, 22(3):200–202, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/200.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/201.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/202.tif.

Dearnley:1982:SDM

[Dea82] P. A. Dearnley. Software development for microcomputer data processing systems. The Computer Journal, 25(2):253–256, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/253.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/254.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/255.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/256.tif.

Deen:1980:CSG

[Dee80]  S. M. Deen.  A canoni-

cal schema for a generalised data model with local interfaces. The Computer Journal, 23(3):201–206, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        03/tiff/201.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        03/tiff/202.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        03/tiff/203.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        03/tiff/204.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        03/tiff/205.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        03/tiff/206.tif.

Deen:1985:TPT

[Dee85]  S. Misbah Deen.  A tu-

ple placement technique for databases. The Computer Journal, 28(3):291–297, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/291.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/292.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/293.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/294.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/295.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/296.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/297.tif.

Defays:1977:EAC

[Def77] D. Defays. An efficient algorithm for a complete link method. The Computer Journal, 20(4):364–366, November 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/364.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/365.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_20/Issue_
                                        04/tiff/366.tif.

Deighton:1995:ATT

[Dei95] Ken Deighton. Average time to travel on two-headed, nonlinear discs. The Computer Journal, 38(10):811–817, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_38/Issue_10/
                                        Vol38_10.body.html#AbstractDeighton.

Delves:1968:NEP

[Del68] L. M. Delves. The numerical evaluation of principal value integrals. The Computer Journal, 10(4):389–391, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/100389.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/tiff/389.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/tiff/390.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/tiff/391.tif.

Delves:1977:LES

[Del77]  L. M. Delves.  Algorithm

96: A linear equation solver for Galerkin and least squares methods. The Computer Journal, 20(4):371–374, November 1977. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/371.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/372.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/373.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/374.tif.

Dempsey:1960:SAS

[Dem60] J. M. Dempsey. Some aspects of simulator design. The Computer Journal, 3(3):158–161, October 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/030158.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/158.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/159.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/160.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/161.tif.

Denham:1980:BRS

[Den80] M. J. Denham. Book review:

Systems Decomposition, Optimisation and Control, by M. G. Singh and A. Titli. The Computer Journal, 23(1):89, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
01/tiff/89.tif.

Deogun:1983:SRT

[Deo83] J. S. Deogun. On scheduling with ready times to minimize mean flow time. The Computer Journal, 26(4):320–328, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/320.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/321.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/322.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/323.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/324.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/325.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/326.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/327.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/328.tif.

Deransart:1979:PSA

[Der79] Pierre Deransart. Proof by semantic attributes of a Lisp compiler. The Computer Journal, 22(3):240–245, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/240.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/241.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/242.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/243.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/244.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/245.tif.

Deutsch:1969:CCL

[Deu69] E. S. Deutsch. Correspondence:

Comments on a line thinning algorithm. The Computer Journal, 12(4):412, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/412.tif. See corre-

spondence [Fra70].

Dewar:1971:API

[Dew71] J. K. S. Dewar. Algorithm 71: Procedures for interval arithmetic. The Computer Journal, 14(4):447–450, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/447.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/448.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/449.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/450.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/451.tif. See note

[Dew72a, Dew74].

Dewar:1972:NAP

[Dew72a] J. K. S. Dewar. Note on Algorithm 71: Procedures for interval arithmetic. The Computer Journal, 15(3):285, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/285.tif. See [Dew71,

Dew74].

Dewar:1972:CAR

[Dew72b] R. B. K. Dewar. Correspondence: On “Another round of FORTRAN”. The Computer Journal, 15(1):7, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/150007.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/tiff/7.tif. See [Cha71,

Fla72, Cha72].

Dewar:1974:NAP

[Dew74] J. K. S. Dewar. Note on Algorithm 71: Procedures for interval arithmetic. The Computer Journal, 17(4):379–380, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/379.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/380.tif. See [Dew71,

Dew72a].

Ferranti:1966:HB

[dF66]  B. de Ferranti.  The hu-

man brain. The Computer Journal, 9(2):117–123, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/090117.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/117.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/118.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/119.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/120.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/121.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/122.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/123.tif.

Duce:1987:FSC

[DF87] D. A. Duce and E. V. C. Fielding. Formal specification — a comparison of two techniques. The Computer Journal, 30(4):316–327, August 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/316.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/317.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/318.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/319.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/320.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/321.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/322.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_30/Issue_
04/tiff/323.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/324.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/325.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/326.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/327.tif.

Dehne:1990:ECG

[DF90] F. Dehne and L. Ficocelli.

An efficient computational geometry method for detecting dotted lines in noisy images. The Computer Journal, 33(5):424–428, October 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/424.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/425.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/426.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/427.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/428.tif.

DiManzo:1979:LOP

[DFO79] M. Di Manzo, A. L. Frisiani, and G. Olimpo. Loop optimisation for parallel processing. The Computer Journal, 22(3):234–239, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/234.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/235.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/236.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/237.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/238.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/239.tif.

Dowse:1958:MB

[DG58] R. G. Dowse and H. W. Gearing. Mathematics in business. The Computer Journal, 1(1):22–24, April 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
01/010022.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
01/tiff/22.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_01/tiff/
23.tif;   http://www3.oup.
co.uk/computer_journal/hdb/

Volume_01/Issue_01/tiff/24.
tif.

Davison:1968:AMT

[DGM68] P. S. Davison, P. Giles, and D. A. R. Matthews. Ageing of magnetic tape: A critical bibliography and comparison of literature sources. The Computer Journal, 11(3):241–246, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/110241.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/241.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/242.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/243.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/244.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/245.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/246.tif.

Duncan:1960:DAT

[DH60] F. G. Duncan and D. H. R.

Huxtable. The DEUCE Alphacode translator. The Computer Journal, 3(2):98–107, July 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

journal/hdb/Volume_03/Issue_
02/030098.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
02/tiff/100.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
02/tiff/101.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
02/tiff/102.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
02/tiff/103.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
02/tiff/104.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
02/tiff/105.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
02/tiff/106.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
02/tiff/107.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
02/tiff/98.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_02/tiff/
99.tif.

Dearden:1997:SEM

[DH97] Andrew M. Dearden and Michael D. Harrison. A software engineering model for case memory systems. The Computer Journal, 40(4):167–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www.oup.co.uk/computer_
journal/Volume_40/Issue_03/

Vol40_04.body.html#AbstractDearden;
http://www3.oup.co.uk/computer_
journal/Volume_40/Issue_04/
Vol40_04.body.html#AbstractDearden.

deGuzman:1993:PDC

[dHM93] I. P. de Guzmán, P. G. Harrison, and E. Medina. Pipelines for divide-and-conquer functions. The Computer Journal, 36(3):254–268, June 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_36/Issue_03/
Vol36_03.body.html#AbstractGuzman.

DeMorgan:1976:SAR

[DHW76] R. M. De Morgan, I. D. Hill, and B. A. Wichmann. A supplement to the ALGOL 60 revised report. The Computer Journal, 19(3):276–288, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/276.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/277.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/278.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/279.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/280.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_19/Issue_
03/tiff/281.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/282.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/283.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/284.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/285.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/286.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/287.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/288.tif.

DeMorgan:1978:MAS

[DHW78] R. M. De Morgan, I. D. Hill, and B. A. Wichmann. Modified ALGOL 60 and the stepuntil element. The Computer Journal, 21(3):282, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/282.tif.

deFerreiraRezende:1997:TIN

[dHZ97] F. de Ferreira Rezende, T. Härder, and J. Zielinski. Transaction identifiers in nested transactions: Implementation schemes and performance. The Computer Journal, 40(5):245–258, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www.oup.co.uk/computer_
journal/Volume_40/Issue_05/
Vol40_05.body.html#AbstractRezende;
http://www3.oup.co.uk/computer_
journal/Volume_40/Issue_05/
Vol40_05.body.html#AbstractRezende.

dInverno:1969:AAL

[d’I69] R. A. d’Inverno. ALAM — Atlas Lisp Algebraic Manipulator. The Computer Journal, 12(2):124–127, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/120124.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/124.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/125.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/126.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/127.tif.

Dwyer:1978:CSP

[DI78] B. Dwyer and J. Inglis. Correspondence: On “Structured programming and input statements”. The Computer Journal, 21(2):190–191, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_

journal/hdb/Volume_21/Issue_
02/tiff/190.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/191.tif. See [Ing76,

Hop78].

Dick:1991:IKC

[Dic91] A. J. J. Dick. An introduction to Knuth–Bendix completion. The Computer Journal, 34(1):2–15, February 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
01/tiff/10.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_01/tiff/
11.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_34/Issue_01/tiff/12.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
34/Issue_01/tiff/13.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
01/tiff/14.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_01/tiff/
15.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_34/Issue_01/tiff/2.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
34/Issue_01/tiff/3.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
01/tiff/4.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_34/Issue_01/tiff/
5.tif;   http://www3.oup.

co.uk/computer_journal/hdb/
Volume_34/Issue_01/tiff/6.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
34/Issue_01/tiff/7.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
01/tiff/8.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_34/Issue_01/tiff/
9.tif.

Dijkstra:1962:OEA

[Dij62]  E. W. Dijkstra.  Operat-

ing experience with ALGOL

60. The Computer Journal, 5(2):125–127, July 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/050125.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/125.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/126.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/127.tif.

Dixon:1985:SSS

[Dix85]  F. J. Dixon.  Simplifying

screen specifications — the ‘Full Screen Manager’ interface and ‘Screen Form’ generating routines. The Computer Journal, 28(2):117–127, May 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

journal/hdb/Volume_28/Issue_
02/tiff/117.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/118.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/119.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/120.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/121.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/122.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/123.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/124.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/125.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/126.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/127.tif.

Dawson:1995:PFI

[DJ95] Scott Dawson and Farham Jahanian. Probing and fault injection of dependable distributed protocols. The Computer Journal, 38(4):286–300, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

journal/Volume_38/Issue_04/
Vol38_04.body.html#AbstractDawson.

Dvorak:1974:DGG

[DK74] S. Dvorak and B. Kropac. Decomposition of the Gozinto’s graph with the use of a nesting store. The Computer Journal, 17(3):245–248, August 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/245.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/246.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/247.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/248.tif.

Devillers:1976:IPF

[DL76] R. Devillers and G. Louchard.

Improvement of parallelism in a finite buffer sharing policy. The Computer Journal, 19(3):238–245, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/238.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/239.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/240.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/241.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/242.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/243.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/244.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/245.tif.

Delis:1997:CAS

[DL97] Alex Delis and Quang LeViet. Contemporary access structures under mixed workloads. The Computer Journal, 40 (4):183–193, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www.oup.co.uk/computer_
                                        journal/Volume_40/Issue_03/
                                        Vol40_04.body.html#AbstractDelis;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_40/Issue_04/
                                        Vol40_04.body.html#AbstractDelis.

Douglas:1960:ALS

[DM60] A. S. Douglas and A. J. Mitchell. AUTOSTAT: a language for statistical data processing. The Computer Journal, 3(2):61–66, July 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        02/030061.sgm.abs.html;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
02/tiff/61.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_02/tiff/
62.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_03/Issue_02/tiff/63.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
03/Issue_02/tiff/64.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
02/tiff/65.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_02/tiff/
66.tif.

DSylva:1964:RIS

[DM64] E. D’Sylva and G. A. Miles.

The S.S.O.R. iteration scheme for equations with σ_{1} ordering. The Computer Journal, 6(4):366–367, January 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/060366.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/366.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/367.tif.

Dearnley:1983:FSP

[DM83] P. A. Dearnley and P. J. Mayhew. In favour of system prototypes and their integration into the systems development cycle. The Computer Journal, 26(1):36–42, February 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
01/tiff/36.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_01/tiff/
37.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_26/Issue_01/tiff/38.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
26/Issue_01/tiff/39.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
01/tiff/40.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_01/tiff/
41.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_26/Issue_01/tiff/42.
tif.

Delves:1985:DDS

[DM85] L. M. Delves and S. C. Mawdsley. DAP-Algol: a development system for parallel algorithms. The Computer Journal, 28(2):148–153, May 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/148.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/149.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/150.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_

02/tiff/151.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/152.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/153.tif.

Dillon:1988:RPV

[DMR88] A. Dillon, C. McKnight, and J. Richardson. Reading from paper versus reading from screen. The Computer Journal, 31(5):457–464, October 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/457.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/458.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/459.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/460.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/461.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/462.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/463.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/464.tif.

Djokic:1989:FIA

[DMS^{+}89] B. Djokic, M. Miyakawa, S. Sekiguchi, I. Semba, and I. Stojmenovic. A fast iterative algorithm for generating set partitions. The Computer Journal, 32(3):281–282, June 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/281.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/282.tif.

Dollimore:1991:DSD

[DMW91] J. Dollimore, E. Miranda, and Xu Wang. The design of a system for distributing shared objects. The Computer Journal, 34(6):514–521, December 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/514.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/515.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/516.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/517.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/518.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/519.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_

06/tiff/520.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/521.tif.

Dearnley:1964:ASP

[DN64] F. H. Dearnley and G. B.

Newell. Automatic segmentation of programs for a two-level store computer. The Computer Journal, 7(3):185–187, October 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/070185.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/185.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/186.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/187.tif.

Deen:1981:DCD

[DNV81] S. M. Deen, D. Nikodem, and A. Vashishta. Design of a canonical database system (PRECI). The Computer Journal, 24(3):200–209, August 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/200.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/201.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_

03/tiff/202.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/203.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/204.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/205.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/206.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/207.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/208.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/209.tif.

Dos:1972:ODU

[DO72] K. Dos and H. Otto. Optimal dynamic use of memory for PL/1 object programs in a real memory environment. The Computer Journal, 15(1):18–20, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/150018.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/tiff/18.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_01/tiff/
19.tif;   http://www3.oup.
co.uk/computer_journal/hdb/

Volume_15/Issue_01/tiff/20.
tif.

Dobosiewicz:1984:RSM

[Dob84] Wlodzimierz Dobosiewicz. Replacement selection in 3-level memories. The Computer Journal, 27(4):334–339, November 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/334.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/335.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/336.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/337.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/338.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/339.tif.

Dobosiewicz:1991:ALP

[Dob91] W. Dobosiewicz. Algorithm 126: Linear probing sort. The Computer Journal, 34 (4):370–373, August 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/370.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/371.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        04/tiff/372.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        04/tiff/373.tif.

Dollimore:1966:GFS

[Dol66]  J. Dollimore.  A general

Fourier synthesis program for the London University Atlas computer. The Computer Journal, 8(4):347–351, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        04/080347.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        04/tiff/347.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        04/tiff/348.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        04/tiff/349.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        04/tiff/350.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        04/tiff/351.tif.

Donovan:1978:NPV

[Don78] J. J. Donovan. A note on performance of VM/ 370 in the integration of models and data bases. The Computer Journal, 21(1):20–24, February 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
01/tiff/20.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_01/tiff/
21.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_21/Issue_01/tiff/22.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
21/Issue_01/tiff/23.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
01/tiff/24.tif.

Douglas:1958:CCb

[Dou58a] A. S. Douglas. Computer and commerce: 2. The Computer Journal, 1(3):132–137, October 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/010132.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/132.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/133.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/134.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/135.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/136.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/137.tif.

Douglas:1958:CCa

[Dou58b]  A. S. Douglas.  Comput-

ers and commerce: 1. The Computer Journal, 1(2):69–70, July 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        02/010069.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        02/tiff/69.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_01/Issue_02/tiff/
                                        70.tif.

Douglas:1958:CCS

[Dou58c] A. S. Douglas. Computers and commerce: 3-stock recording and control. The Computer Journal, 1(3):137–141, October 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        03/010137.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        03/tiff/137.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        03/tiff/138.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        03/tiff/139.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        03/tiff/140.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        03/tiff/141.tif.

Douglas:1959:CCM

[Dou59a] A. S. Douglas. Computer and commerce: 4 — management and control. The Computer Journal, 1(4):168–171, January 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/010168.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/tiff/168.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/tiff/169.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/tiff/170.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/tiff/171.tif.

Douglas:1959:SAB

[Dou59b] A. S. Douglas. The state of the art — (b) Computers in British universities. The Computer Journal, 2(3):100–102, October 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/020100.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/100.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/101.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_02/Issue_
03/tiff/102.tif.

Douglas:1959:TRR

[Dou59c] A. S. Douglas. Technique for the recording or, and reference to data in a computer. The Computer Journal, 2(1):1–9, April 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
01/020001.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
01/tiff/1.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_02/Issue_01/tiff/
2.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_02/Issue_01/tiff/3.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
02/Issue_01/tiff/4.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
01/tiff/5.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_02/Issue_01/tiff/
6.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_02/Issue_01/tiff/7.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
02/Issue_01/tiff/8.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
01/tiff/9.tif.

Douglas:1961:ICD

[Dou61] A. S. Douglas. The impact of computers on documentation. The Computer Journal, 4(2):145–149, July 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/040145.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/145.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/146.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/147.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/148.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/149.tif.

Douglas:1965:ULC

[Dou65] A. S. Douglas. The use of a large computer on a bureau basis. The Computer Journal, 7(4):262–263, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/070262.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/262.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/263.tif.

Downham:1969:RT

[Dow69] D. Y. Downham. The runs up and down test. The Computer Journal, 12(4):373–376, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/120373.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/373.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/374.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/375.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/376.tif.

Downing:1971:CCC

[Dow71] J. Downing. A computer controlled circulation system. The Computer Journal, 14(1):2– 6, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
14/Issue_01/140002.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
14/Issue_01/tiff/2.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/tiff/3.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_14/Issue_01/tiff/
4.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_14/Issue_01/tiff/5.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
14/Issue_01/tiff/6.tif.

Paula:1960:PAC

[dP60] F. C. de Paula. Problems of auditing computer data. Section 2: The external auditor and computers. The Computer Journal, 3(1):11–14, April 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
01/030011.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
01/tiff/11.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_01/tiff/
12.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_03/Issue_01/tiff/13.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
03/Issue_01/tiff/14.tif.

Doust:1964:LRV

[DP64] A. Doust and V. E. Price. The latent roots and vectors of a singular matrix. The Computer Journal, 7(3):222–227, October 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/070222.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/222.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/223.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_

03/tiff/224.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/225.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/226.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/227.tif.

Dineley:1966:KAB

[DP66] J. L. Dineley and C. Preece. KALDAS, an algorithmically based digital simulation of analogue computation. The Computer Journal, 9(2):181– 187, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
09/Issue_02/090181.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
09/Issue_02/tiff/181.tif;
 http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/182.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/183.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/184.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/185.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/186.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/187.tif.

Dakin:1973:MCA

[DP73] R. J. Dakin and P. C.

Poole. A mixed code approach. The Computer Journal, 16(3):219–222, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/160219.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/219.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/220.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/221.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/222.tif. See corre-

spondence [Mac74, Bem74].

Dance:1976:ALD

[DP76] D. L. Dance and U. W. Pooch.

An adaptive on line data compression system. The Computer Journal, 19(3):216–224, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/216.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/217.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/218.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_19/Issue_
03/tiff/219.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/220.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/221.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/222.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/223.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/224.tif.

Dawson:1988:EDF

[DP88] Kathryn S. Dawson and Lorraine M. Purgailis Parker. From entity-relationship diagrams to fourth normal form: a pictorial aid to analysis. The Computer Journal, 31(3):258–268, June 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/258.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/259.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/260.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/261.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/262.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_31/Issue_
03/tiff/263.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/264.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/265.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/266.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/267.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/268.tif.

Dixon:1989:TPO

[DPSW89] G. N. Dixon, G. D. Parrington, S. K. Shrivastava, and S. M. Wheater. The treatment of persistent objects in Arjuna. The Computer Journal, 32(4):323–332, August 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/323.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/324.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/325.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/326.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/327.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_32/Issue_
                                        04/tiff/328.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        04/tiff/329.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        04/tiff/330.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        04/tiff/331.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        04/tiff/332.tif.

Diaz:1994:DDR

[DPT94] M. Diaz, E. Pimentel, and J. M. Troya. DROL: a distributed and real-time object-oriented logic environment. The Computer Journal, 37(5):407–421, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_37/Issue_05/
                                        Vol37_05.body.html#AbstractDiaz.

Downham:1967:MCP

[DR67] D. Y. Downham and F. D. K. Roberts. Multiplicative congruential pseudo-random number generators. The Computer Journal, 10(1):74–77, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        01/100074.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        01/tiff/74.tif;   http://
                                        www3.oup.co.uk/computer_journal/

hdb/Volume_10/Issue_01/tiff/
75.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_10/Issue_01/tiff/76.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/77.tif.

dInverno:1974:CFS

[dRC74] R. A. d’Inverno and R. A.

Russell-Clark. CLAM — its function, structure and implementation. The Computer Journal, 17(3):229–233, August 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/229.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/230.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/231.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/232.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/233.tif.

Dromey:1987:DSA

[Dro87] R. G. Dromey. Derivation of sorting algorithms from a specification. The Computer Journal, 30(6):512–518, December 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/512.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        06/tiff/513.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        06/tiff/514.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        06/tiff/515.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        06/tiff/516.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        06/tiff/517.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        06/tiff/518.tif.

Diaz:1996:DPL

[DRT96] M. Diaz, B. Rubio, and J. M. Troya. Distributed programming with a logic channel based coordination model. The Computer Journal, 39(10):876–??, ???? 1996. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www.oup.co.uk/computer_
                                        journal/Volume_39/Issue_10/
                                        Vol39_10.body.html#AbstractDiaz;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_39/Issue_10/
                                        Vol39_10.body.html#AbstractDiaz.

Drummond:1963:SAR

[Dru63] J. Drummond. Some aspects of recording graduated national insurance contributions. The Computer Journal, 6(1):1–4, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_

journal/hdb/Volume_06/Issue_
01/060001.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/tiff/1.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/
2.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_06/Issue_01/tiff/3.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
06/Issue_01/tiff/4.tif.

Deist:1967:SSN

[DS67] F. H. Deist and L. Sefor. Solution of systems of non-linear equations by parameter variation. The Computer Journal, 10(1):78–82, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/100078.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/78.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_01/tiff/
79.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_10/Issue_01/tiff/80.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/81.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/82.tif.

deChampeaux:1977:OTB

[dS77] D. de Champeaux and L. Sint.

An optimality theorem for a bi-directional heuristic search algorithm. The Computer Journal, 20(2):148–150, May 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        02/tiff/148.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        02/tiff/149.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        02/tiff/150.tif.

deSmith:1980:PPA

[dS80] Michael de Smith. Points, polygons, and areas. The Computer Journal, 23(2):189, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        02/tiff/189.tif. See corre-

spondence [Kio81].

Dey:1990:NPS

[DS90] S. Dey and P. K. Srimani. A new parallel sorting algorithm and its efficient VLSI implementation. The Computer Journal, 33(3):241–246, June 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        03/tiff/241.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        03/tiff/242.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_33/Issue_
03/tiff/243.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/244.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/245.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/246.tif.

Dandamudi:1991:IPS

[DS91] S. P. Dandamudi and P. G.

Sorenson. Improved partialmatch search algorithms for BD trees. The Computer Journal, 34(5):415–422, October 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/415.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/416.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/417.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/418.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/419.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/420.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/421.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_34/Issue_
                                        05/tiff/422.tif.

Das:1999:ALP

[DS99] Sajal K. Das and Sanjoy K. Sen. Adaptive location prediction strategies based on a hierarchical network model in a cellular mobile environment. The Computer Journal, 42(6):473– 486, 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        06/420473.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        06/pdf/420473.pdf.

deLemos:1992:TSC

[dSA92] R. de Lemos, A. Saeed, and T. Anderson. A train set as a case study for the requirements analysis of safetycritical systems. The Computer Journal, 35(1):30–40, February 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_35/Issue_01/
                                        Vol35_01.body.html#AbstractLemos.

Dahlgren:1997:RRM

[DSB97] Fredrick H. Dahlgren, Per Stenström, and Marten Björkman. Reducing the read-miss penalty for flat COMA protocols. The Computer Journal, 40(4):208– ??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www.oup.co.uk/
computer_journal/Volume_40/
Issue_03/Vol40_04.body.html#
AbstractDahlgren;   http:
//www3.oup.co.uk/computer_
journal/Volume_40/Issue_04/
Vol40_04.body.html#AbstractDahlgren.

Djordjevic:1996:CTS

[DT96] Goran Lj. Djordjevic and Milorad B. Tosic. A compile-time scheduling heuristic for multiprocessor architectures. The Computer Journal, 39(8):663– ??, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www.oup.co.uk/jnls/
list/comjnl/hdb/Volume_39/
Issue_08/390663.sgm.abs.html;
 http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_08/
Vol39_08.body.html#AbstractDjord.

Deen:1988:DDD

[DTIR88] S. M. Deen, M. C. Taylor, P. A. Ingram, and K. W. Rayner. A distributed directory database system for telecommunications. The Computer Journal, 31(2):175–181, April 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/175.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/176.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/177.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        02/tiff/178.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        02/tiff/179.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        02/tiff/180.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        02/tiff/181.tif.

Du:1996:IMG

[Du96]  C. Du.  An interpolation

method for grid-based terrain modelling. The Computer Journal, 39(10):837–??, ???? 1996. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www.oup.co.uk/computer_
                                        journal/Volume_39/Issue_10/
                                        Vol39_10.body.html#AbstractDu;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_39/Issue_10/
                                        Vol39_10.body.html#AbstractDu.

Dubhashi:1993:QEA

[Dub93] Devdatt P. Dubhashi. Quantifier elimination in p-adic fields. The Computer Journal, 36(5):419–426, October 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_36/Issue_05/
                                        Vol36_05.body.html#AbstractDubhashi.

Duke:1966:CCR

[Duk66] B. J. Duke. Correspondence: On the computation of rotational levels of rigid asymmetric top molecules. The Computer Journal, 9(2):210, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/090210.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/210.tif. See note

[Rac65] and correspondence [Jon66].

Duncan:1962:EIA

[Dun62a]  F. G. Duncan.  Errata:

“Implementation of ALGOL 60 for the English Electric KDF9”. The Computer Journal, 5(3):176, October 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/176.tif. See [Dun62b].

Duncan:1962:IAE

[Dun62b] F. G. Duncan. Implementation of ALGOL 60 for the English Electric KDF9. The Computer Journal, 5(2):130–131, July 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/050130.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/130.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_05/Issue_
                                        02/tiff/131.tif. See errata

[Dun62a].

Duncan:1963:IOA

[Dun63] F. G. Duncan. Input and output for ALGOL 60 on KDF

9. The Computer Journal, 5(4):341–344, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        04/050341.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        04/tiff/341.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        04/tiff/342.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        04/tiff/343.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        04/tiff/344.tif.

Dunham:1973:WAA

[Dun73] C. B. Dunham. When absolute approximations are optimal in function subroutines. The Computer Journal, 16(3):265, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        03/160265.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        03/tiff/265.tif.

Dunham:1978:CCR

[Dun78]  C. B. Dunham.  Corre-

spondence: Checking reproduction of published programs. The Computer Journal, 21(4):383, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/383.tif. See corre-

spondence [Jac79].

Duncan:1979:QCM

[Dun79]  G. Duncan.  Q charts —

a method of specification. The Computer Journal, 22 (4):380–381, November 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/380.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/381.tif.

Dunham:1982:NPP

[Dun82]  C. B. Dunham.  The ne-

cessity of publishing programs. The Computer Journal, 25(1):61–62, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/61.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/
62.tif.

Durbin:1974:NIL

[Dur74]  F. Durbin.  Numerical in-

version of Laplace transforms: an efficient improvement to Dubner and Abate’s method. The Computer Journal, 17 (4):371–376, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/371.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/372.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/373.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/374.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/375.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/376.tif.

Witte:1964:EES

[dW64] L. de Witte. Error estimates for smoothing and extrapolation formulae. The Computer Journal, 7(3):220–221, October 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/070220.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/220.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/221.tif.

Dunham:1973:CDM

[DW73] C. B. Dunham and G. A. Watson. Correspondence: Degeneracy in the matrix of partial derivatives. The Computer Journal, 16(3):284, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/284.tif. See [OW71].

Duff:1977:CLA

[DW77] M. J. B. Duff and D. M. Watson. The cellular logic array image processor. The Computer Journal, 20(1):68–72, February 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
01/tiff/68.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_01/tiff/
69.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_20/Issue_01/tiff/70.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
20/Issue_01/tiff/71.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
01/tiff/72.tif.

Davis:1988:MST

[DW88] Martin Davis and Elaine Weyuker. Metric spacebased test-data adequacy criteria. The Computer Journal, 31(1):17–24, February 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
01/tiff/17.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_01/tiff/
18.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_31/Issue_01/tiff/19.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
31/Issue_01/tiff/20.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
01/tiff/21.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_01/tiff/
22.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_31/Issue_01/tiff/23.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
31/Issue_01/tiff/24.tif.

Dick:1991:OTR

[DW91] A. J. J. Dick and P. Watson. Order-sorted term rewriting. The Computer Journal, 34(1):16–19, February 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
01/tiff/16.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_01/tiff/
17.tif;   http://www3.oup.
co.uk/computer_journal/hdb/

Volume_34/Issue_01/tiff/18.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
34/Issue_01/tiff/19.tif.

Dasarathy:1980:TOT

[DY80] B. Dasarathy and Cheng Yang.

A transformation on ordered trees. The Computer Journal, 23(2):161–164, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/161.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/162.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/163.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/164.tif.

Duncan:1975:PNL

[DZW75] F. G. Duncan, D. Zissos, and Maureen Walls. A postfix notation for logic circuits. The Computer Journal, 18(1):63–69, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
01/tiff/63.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_01/tiff/
64.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_18/Issue_01/tiff/65.
tif;   http://www3.oup.co.

                                        uk/computer_journal/hdb/Volume_
                                        18/Issue_01/tiff/66.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        01/tiff/67.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_18/Issue_01/tiff/
                                        68.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_18/Issue_01/tiff/69.
                                        tif.

Evans:1970:AAS

[EA70] D. J. Evans and L. V. Atkinson. Algorithm 51: An algorithm for the solution of general three term linear systems. The Computer Journal, 13(3):323–324, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/323.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/324.tif.

Evans:1983:ECB

[EA83] D. J. Evans and F. Abdollahzadeh. Efficient construction of balanced binary trees. The Computer Journal, 26(3):193–195, August 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        03/tiff/193.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        03/tiff/194.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/195.tif.

Edwards:1966:AMT

[EAL66] D. B. G. Edwards, D. Aspinal, and R. Lane. Addressing magnetic tape on the Atlas 1 computing machine. The Computer Journal, 8(4):323–329, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/080323.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/323.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/324.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/325.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/326.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/327.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/328.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/329.tif.

Earnshaw:1975:AVP

[Ear75] R. A. Earnshaw. Is APL a viable programming language? The Computer Journal, 18(4):318–323, November 1975. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/318.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/319.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/320.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/321.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/322.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/323.tif.

Earnshaw:1980:LTI

[Ear80]  R. A. Earnshaw.  Line

tracking for incremental plotters. The Computer Journal, 23(1):46–52, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
01/tiff/46.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/
47.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_23/Issue_01/tiff/48.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
23/Issue_01/tiff/49.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
01/tiff/50.tif;   http://
www3.oup.co.uk/computer_journal/

hdb/Volume_23/Issue_01/tiff/
51.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_23/Issue_01/tiff/52.
tif.

Eason:1976:UNC

[Eas76] K. D. Eason. Understanding the naive computer user. The Computer Journal, 19(1):3–7, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
01/tiff/3.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_19/Issue_01/tiff/
4.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_19/Issue_01/tiff/5.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
19/Issue_01/tiff/6.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
01/tiff/7.tif.

Edupuganty:1989:TGF

[EB89] B. Edupuganty and B. R.

Bryant. Two-level grammar as a functional programming language. The Computer Journal, 32(1):36–44, February 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
01/tiff/36.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_01/tiff/
37.tif;   http://www3.oup.
co.uk/computer_journal/hdb/

                                        Volume_32/Issue_01/tiff/38.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        32/Issue_01/tiff/39.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        01/tiff/40.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_32/Issue_01/tiff/
                                        41.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_32/Issue_01/tiff/42.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        32/Issue_01/tiff/43.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        01/tiff/44.tif.

Ecclestone:1971:CQI

[Ecc71] M. J. Ecclestone. Correspondence: On “A quasi-intrinsic scheme for passing a smooth curve through a discrete set of points”. The Computer Journal, 14(4):351, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/140351.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/351.tif. See [McC70b,

Smi71, Sab72, Coo72b].

Ecker:1974:PSQ

[Eck74] A. Ecker. The period of search for the quadratic and related hash methods. The Computer Journal, 17(4):340–343, November 1974. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/340.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/341.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/342.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/343.tif.

Edelstein:1963:PLB

[Ede63]  L. A. Edelstein.  ‘picture

logic’ for ‘Bacchus’ a fourthgeneration computer. The Computer Journal, 6(2):144–153, July 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/060144.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/144.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/145.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/146.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/147.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/148.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/149.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/tiff/150.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/tiff/151.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/tiff/152.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/tiff/153.tif.

Edelstein:1965:NAA

[Ede65] L. A. Edelstein. Note on analytic angular integrations using series in Legendre polynomials. The Computer Journal, 8(2):157–158, July 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        02/tiff/157.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        02/tiff/158.tif.

Edmonds:1960:GPR

[Edm60] A. R. Edmonds. The generation of pseudo-random numbers on electronic digital computers. The Computer Journal, 2(4):181–185, January 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_02/Issue_
                                        04/020181.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_02/Issue_
                                        04/tiff/181.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/182.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/183.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/184.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/185.tif.

Ein-Dor:1986:AAS

[EDS86] P. Ein-Dor and E. Segev. Attitudes, association and success of MIS: some empirical results from research in the context of a business game. The Computer Journal, 29(3):212–221, June 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/212.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/213.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/214.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/215.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/216.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/217.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/218.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        03/tiff/219.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        03/tiff/220.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        03/tiff/221.tif.

Edwards:1972:LBM

[Edw72] C. R. Edwards. The logic of Boolean matrices. The Computer Journal, 15(3):247– 253, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        15/Issue_03/150247.sgm.abs.
                                        html;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        15/Issue_03/tiff/247.tif;
                                         http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/248.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/249.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/250.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/251.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/252.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/253.tif.

Edwards:1977:CBS

[Edw77] B. J. Edwards. Choice of block sizes for magnetic tape files. The Computer Journal, 20(1):10–14, February 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
01/tiff/10.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_01/tiff/
11.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_20/Issue_01/tiff/12.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
20/Issue_01/tiff/13.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
01/tiff/14.tif.

Edwards:1978:SID

[Edw78] C. R. Edwards. Some improved designs for the Digital Summation Threshold Logic (DSTL) gate. The Computer Journal, 21(1):73–78, February 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
01/tiff/73.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_01/tiff/
74.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_21/Issue_01/tiff/75.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
21/Issue_01/tiff/76.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
01/tiff/77.tif;   http://

                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_21/Issue_01/tiff/
                                        78.tif.

Evans:1963:IPO

[EF63a] D. J. Evans and C. V. D. Forrington. An iterative process for optimizing symmetric successive over-relaxation. The Computer Journal, 6 (3):271–273, October 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        03/060271.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        03/tiff/271.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        03/tiff/272.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        03/tiff/273.tif.

Evans:1963:NSC

[EF63b] D. J. Evans and C. V. D. Forrington. Note on the solution of certain tri-diagonal systems of linear equations. The Computer Journal, 5(4):327–328, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        04/050327.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        04/tiff/327.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_05/Issue_
04/tiff/328.tif.

El-gendi:1969:CSD

[Eg69] S. E. El-gendi. Chebyshev solution of differential, integral and integro differential equations. The Computer Journal, 12(3):282–287, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/120282.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/282.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/283.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/284.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/285.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/286.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/287.tif.

Evans:1976:SCB

[EH76] D. J. Evans and M. Hatzopoulos. The solution of certain banded systems of linear equations using the folding algorithm. The Computer Journal, 19(2):184–187, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/184.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/185.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/186.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/187.tif.

Eves:1978:IMT

[EH78] J. M. Eves and S. L. Hurst. Improvements in multioutput threshold-logic gates. The Computer Journal, 21(1):79– 85, February 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        21/Issue_01/tiff/79.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        01/tiff/80.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_21/Issue_01/tiff/
                                        81.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_21/Issue_01/tiff/82.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        21/Issue_01/tiff/83.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        01/tiff/84.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_21/Issue_01/tiff/
                                        85.tif.

Eason:1988:SRD

[EH88] K. D. Eason and Susan Harker.

The supplier’s role in the design of products for organisations. The Computer Journal, 31(5):426–430, October 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/426.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/427.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/428.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/429.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/430.tif.

Edelsbrunner:1986:RPL

[EHH86] H. Edelsbrunner, G. Haring, and David Hilbert. Rectangular point location in D dimensions with applications. The Computer Journal, 29(1):76–82, February 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
01/tiff/76.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_01/tiff/
77.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_29/Issue_01/tiff/78.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
29/Issue_01/tiff/79.tif;

                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        01/tiff/80.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_29/Issue_01/tiff/
                                        81.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_29/Issue_01/tiff/82.
                                        tif.

Ehrman:1972:CSE

[Ehr72] J. R. Ehrman. Correspondence: On “Suggested Extensions to FORTRAN IV”. The Computer Journal, 15(4):332, November 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/150332.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/tiff/332.tif. See [Fin72].

El-Hamdouchi:1989:CHA

[EHW89] A. El-Hamdouchi and P. Willett. Comparison of hierarchic agglomerative clustering methods for document retrieval. The Computer Journal, 32(3):220–227, June 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        03/tiff/220.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        03/tiff/221.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        03/tiff/222.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/223.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/224.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/225.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/226.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/227.tif.

Eker:1995:ACM

[Eke95]  S. M. Eker.  Associative-

commutative matching via bipartite graph matching. The Computer Journal, 38(5):381– 399, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/Volume_
38/Issue_05/Vol38_05.body.
html#AbstractEker.

Er:1982:TPC

[EL82] M. C. Er and B. G. T. Lowden. The theory and practice of constructing an optimal polyphase sort. The Computer Journal, 25(1):93–101, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/100.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/101.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/93.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_25/Issue_01/tiff/
                                        94.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_25/Issue_01/tiff/95.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        25/Issue_01/tiff/96.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/97.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_25/Issue_01/tiff/
                                        98.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_25/Issue_01/tiff/99.
                                        tif.

Ellis:1989:VGC

[EL89] J. A. Ellis and P. M. Lepolesa. A Las Vegas graph colouring algorithm. The Computer Journal, 32(5):474–476, October 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        05/tiff/474.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        05/tiff/475.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        05/tiff/476.tif.

Elcock:1965:NAL

[Elc65] E. W. Elcock. Note on the addressing of lists by their sourcelanguage names. The Computer Journal, 8(3):242–243, October 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/080242.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/242.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/243.tif.

Elliott:1963:CSM

[Ell63] D. Elliott. A Chebyshev series method for the numerical solution of Fredholm integral equations. The Computer Journal, 6(1):102–112, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/060102.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/tiff/102.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/tiff/103.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/tiff/104.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/tiff/105.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/tiff/106.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/tiff/107.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_06/Issue_
                                        01/tiff/108.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        01/tiff/109.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        01/tiff/110.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        01/tiff/111.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        01/tiff/112.tif.

Evans:1978:CCM

[ELR78] R. V. Evans, G. S. Lockington, and T. N. Reid. A compiler compiler and methodology for problem oriented language compiler implementors. The Computer Journal, 21(2):117–121, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        02/tiff/117.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        02/tiff/118.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        02/tiff/119.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        02/tiff/120.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        02/tiff/121.tif.

Elsworth:1979:CIL

[Els79]  E. F. Elsworth.  Compila-

tion via an intermediate language. The Computer Journal, 22(3):226–233, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/226.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/227.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/228.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/229.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/230.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/231.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/232.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/233.tif.

Evangelisti:1971:GMU

[EM71] C. J. Evangelisti and S. P.

Morse. Graphical modelling using contextually implied functions. The Computer Journal, 14(4):382–389, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/140382.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_

                                        04/tiff/382.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/383.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/384.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/385.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/386.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/387.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/388.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/389.tif.

Ekanadham:1981:MN

[EM81a] K. Ekanadham and A. Mahjoub. Microcomputer networks. The Computer Journal, 24(1):17– 24, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        24/Issue_01/tiff/17.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        01/tiff/18.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_24/Issue_01/tiff/
                                        19.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_24/Issue_01/tiff/20.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        24/Issue_01/tiff/21.tif;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
01/tiff/22.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_01/tiff/
23.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_24/Issue_01/tiff/24.
tif.

Ewusi-Mensah:1981:CNU

[EM81b]  K. Ewusi-Mensah.  Com-

puter network usage control through peak load pricing. The Computer Journal, 24(1):71–77, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
01/tiff/71.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_01/tiff/
72.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_24/Issue_01/tiff/73.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
24/Issue_01/tiff/74.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
01/tiff/75.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_01/tiff/
76.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_24/Issue_01/tiff/77.
tif.

Ehrig:1992:IASa

[EMCO92] H. Ehrig, B. Mahr, I. Classen, and F. Orejas. Introduction to algebraic specification. part 1: Formal methods for software development. The Computer Journal, 35(5):460–467, October 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_35/Issue_05/
                                        Vol35_05.body.html#AbstractEhrig1.

Emery:1980:BRA

[Eme80] G. Emery. Book reviews: Associative networks, edited by N. V. Findler, 1979; The MU5 Computer System, by Derrick Morris and Rolland N. Ibbett, 1979; Cobol: An Introduction to Structured Logic and Modular Program Design, by William Davis and Richard Fisher, 1979. The Computer Journal, 23(2):191, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        02/tiff/191.tif.

Elcock:1972:DDC

[EMM72] E. W. Elcock, J. J. McGregor, and A. M. Murray. Data directed control and operating systems. The Computer Journal, 15(2):125–129, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        02/150125.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        02/tiff/125.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/126.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/127.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/128.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/129.tif.

Ehrig:1992:IASb

[EMO92] H. Ehrig, B. Mahr, and F. Orejas. Introduction to algebraic specification. part 2: From classical view to foundations of system specifications. The Computer Journal, 35(5):468–477, October 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_35/Issue_05/
Vol35_05.body.html#AbstractEhrig2.

Ernvall:1982:AUR

[EN82] Jarmo Ernvall and Olli Nevalainen.

An algorithm for unbiased random sampling. The Computer Journal, 25(1):45–47, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/45.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/
46.tif;   http://www3.oup.
co.uk/computer_journal/hdb/

Volume_25/Issue_01/tiff/47.
tif.

England:1969:EER

[Eng69] R. England. Error estimates for Runge–Kutta type solutions to systems of ordinary differential equations. The Computer Journal, 12(2):166–170, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/120166.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/166.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/167.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/168.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/169.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/170.tif.

Evershed:1971:HLL

[ER71] D. G. Evershed and G. E. Rippon. High level languages for low level users. The Computer Journal, 14(1):87–90, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/140087.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_

01/tiff/87.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_01/tiff/
88.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_14/Issue_01/tiff/89.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
14/Issue_01/tiff/90.tif. See

correspondence [Bac71, Wal71, Fel72, Hil72].

Eidelson:1980:IBI

[ER80] A. F. Eidelson and I. J. Robinson. Implementation of BED-SOCS: an interactive simulation language. The Computer Journal, 23(3):233–236, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/233.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/234.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/235.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/236.tif. See corre-

spondence [SB81].

Er:1982:NCP

[Er82a]  M. C. Er.  A note on

computing precedence functions. The Computer Journal, 25(3):397–398, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/397.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/398.tif.

Er:1982:RAT

[Er82b]  M. C. Er.  A representa-

tion approach to the Tower of Hanoi problem. The Computer Journal, 25(4):442–447, November 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/442.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/443.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/444.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/445.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/446.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/447.tif.

Er:1983:FAC

[Er83a] M. C. Er. Fast algorithm for computing order-K Fibonacci numbers. The Computer Journal, 26(3):224–227, August 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_

03/tiff/224.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/225.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/226.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/227.tif.

Er:1983:NGW

[Er83b] M. C. Er. A note on generating well-formed parenthesis strings lexicographically. The Computer Journal, 26 (3):205–207, August 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/205.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/206.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/207.tif.

Er:1983:PCA

[Er83c] M. C. Er. A parallel computation approach to topological sorting. The Computer Journal, 26(4):293–295, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/293.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/294.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        04/tiff/295.tif.

Er:1984:CTHa

[Er84a] M. C. Er. The colour Towers of Hanoi: a generalization. The Computer Journal, 27(1):80–82, February 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        01/tiff/80.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_27/Issue_01/tiff/
                                        81.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_27/Issue_01/tiff/82.
                                        tif.

Er:1984:CTHb

[Er84b] M. C. Er. The cyclic Towers of Hanoi: a representation approach. The Computer Journal, 27(2):171–175, May 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        02/tiff/171.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        02/tiff/172.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        02/tiff/173.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        02/tiff/174.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_27/Issue_
02/tiff/175.tif.

Er:1984:GCT

[Er84c] M. C. Er. The generalized colour Towers of Hanoi: an iterative algorithm. The Computer Journal, 27(3):278–282, August 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/278.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/279.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/280.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/281.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/282.tif.

Er:1985:EOT

[Er85] M. C. Er. Enumerating ordered trees lexicographically. The Computer Journal, 28 (5):538–542, November 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/538.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/539.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/540.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        05/tiff/541.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        05/tiff/542.tif.

Elloy:1986:ELC

[ER86a] J. P. Elloy and O. Roux. Electre: a language for control structuring in real time. The Computer Journal, 29(3):229– 234, June 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        29/Issue_03/tiff/229.tif;
                                         http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        03/tiff/230.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        03/tiff/231.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        03/tiff/232.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        03/tiff/233.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        03/tiff/234.tif.

Er:1986:UTI

[Er86b] M. C. Er. The use of termination indicators in computer programming. The Computer Journal, 29(5):430–433, October 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_

05/tiff/430.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/431.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/432.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/433.tif.

Er:1987:EIP

[Er87a] M. C. Er. An efficient implementation of permutation backtracking in lexicographic order. The Computer Journal, 30(3):282, June 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/282.tif.

Er:1987:LLR

[Er87b]  M. C. Er.  Lexicographic

listing and ranking of T -ary trees. The Computer Journal, 30(6):569–572, December 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/569.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/570.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/571.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/572.tif.

Er:1988:FAG

[Er88a]  M. C. Er.  A fast algo-

rithm for generating set partitions. The Computer Journal, 31(3):283–284, June 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        03/tiff/283.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        03/tiff/284.tif.

Er:1988:SAG

[Er88b]  M. C. Er.  A simple al-

gorithm for generating nonregular trees in lexicographic order. The Computer Journal, 31(1):61–64, February 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        01/tiff/61.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_31/Issue_01/tiff/
                                        62.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_31/Issue_01/tiff/63.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        31/Issue_01/tiff/64.tif.

Er:1988:SRA

[Er88c] M. C. Er. A smooth reshuffling algorithm for solving the Bulterman’s problem. The Computer Journal, 31(6):557–560, December 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/557.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/558.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/559.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/560.tif.

Er:1989:CAP

[Er89a] M. C. Er. Classes of admissible permutations that are generatable by depth-first traversals of ordered trees. The Computer Journal, 32(1):76–85, February 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
01/tiff/76.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_01/tiff/
77.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_32/Issue_01/tiff/78.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
32/Issue_01/tiff/79.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
01/tiff/80.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_01/tiff/
81.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_32/Issue_01/tiff/82.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_

                                        32/Issue_01/tiff/83.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        01/tiff/84.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_32/Issue_01/tiff/
                                        85.tif.

Er:1989:NAG

[Er89b] M. C. Er. A new algorithm for generating binary trees using rotations. The Computer Journal, 32(5):470–473, October 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        05/tiff/470.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        05/tiff/471.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        05/tiff/472.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        05/tiff/473.tif.

Erdelyi:1965:SCI

[Erd65] I. Erdelyi. On speeding convergence of an iterative eigenvalue process. The Computer Journal, 8(2):159–165, July 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        02/080159.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        02/tiff/159.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_08/Issue_
02/tiff/160.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/161.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/162.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/163.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/164.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/165.tif.

Eriksen:1974:DBC

[Eri74] Sven Eriksen. The data base concept. The Computer Journal, 8(1):??, ???? 1974. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Erkio:1984:WCP

[Erk84]  Hannu Erkio.  The worst

case permutation for median-of-three quicksort. The Computer Journal, 27(3):276–277, August 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/276.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/277.tif.

Ewen-Smith:1975:CCD

[ES75] B. M. Ewen-Smith. Correspondence: Check digits and error correction. The Computer Journal, 18(3):284, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/284.tif. See [And74,

And75, How75].

Evans:1977:ACB

[ES77] D. J. Evans and S. A. Smith. Algorithm 99: On the construction of balanced binary trees for parallel processing. The Computer Journal, 20 (4):378–379, November 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/378.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/379.tif.

Elliman:1981:PTM

[ES81] A. D. Elliman and A. R. Sudhindra. Performances of two microprocessor-based multiplexor techniques. The Computer Journal, 24(4):343–346, November 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
24/Issue_04/tiff/343.tif;
 http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/344.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_

04/tiff/345.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/346.tif.

Evered:1995:FOI

[ESK95] Mark Evered, Axel Schmolitzky, and Michael Kölling. A flexible object invocation language based on object-oriented language definition. The Computer Journal, 38(3):181–191, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_38/Issue_03/
Vol38_03.body.html#AbstractEvered.

Edwards:1975:CSC

[ET75] J. P. N. Edwards and The Editor. Correspondence: Saving CPU time in scientific computation. The Computer Journal, 18(3):285, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/285.tif. See [HV74,

Art75, Hat76].

Egyhazy:1988:QPA

[ET88] C. Egyhazy and K. Triantis. A query processing algorithm for distributed relational database systems. The Computer Journal, 31(1):34–40, February 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

journal/hdb/Volume_31/Issue_
01/tiff/34.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_01/tiff/
35.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_31/Issue_01/tiff/36.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
31/Issue_01/tiff/37.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
01/tiff/38.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_01/tiff/
39.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_31/Issue_01/tiff/40.
tif.

Etherton:1971:DSN

[Eth71]  M. Etherton.  Data struc-

tures for a network design system. The Computer Journal, 14(4):366–374, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/140366.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/366.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/367.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/368.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/369.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_14/Issue_
04/tiff/370.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/371.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/372.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/373.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/374.tif.

Evans:1962:NLR

[Eva62] D. J. Evans. Note on the line over-relaxation factor for small mesh size. The Computer Journal, 5(1):48–50, April 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
01/050048.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
01/tiff/48.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_01/tiff/
49.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_05/Issue_01/tiff/50.
tif.

Evans:1963:EMA

[Eva63] D. J. Evans. The extrapolated modified Aitken iteration method for solving elliptic difference equations. The Computer Journal, 6(2):193–201, July 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/060193.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/193.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/194.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/195.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/196.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/197.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/198.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/199.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/200.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/201.tif.

Evans:1964:EMA

[Eva64] D. J. Evans. The extrapolated modified Aitken iteration method applied to σ_{1}- ordered systems of linear equations. The Computer Journal, 7(2):137–140, July 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_

02/070137.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/137.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/138.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/139.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/140.tif.

Evans:1965:ELR

[Eva65a] D. J. Evans. Estimation of the line over-relaxation factor and convergence rates of an alternating direction line over-relaxation technique. The Computer Journal, 7 (4):318–321, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/070318.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/318.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/319.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/320.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/321.tif.

Evans:1965:SEM

[Eva65b] D. J. Evans. A stable explicit method for the finitedifference solution of a fourthorder parabolic partial differential equation. The Computer Journal, 8(3):280–287, October 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/080280.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/280.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/281.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/282.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/283.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/284.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/285.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/286.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/287.tif. See com-

ments [Eva66].

Evans:1966:CSE

[Eva66] D. J. Evans. Correspondence: On “A stable explicit method for the finite-difference solution of a fourth-order parabolic partial differential equation”. The Computer Journal, 9(1):115, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/115.tif. See [Eva65b].

Evans:1971:AAS

[Eva71] D. J. Evans. Algorithm 69: An algorithm for the solution of symmetric general three term linear systems. The Computer Journal, 14(4):444–445, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/444.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/445.tif.

Evans:1972:ASC

[Eva72] D. J. Evans. An algorithm for the solution of certain tridiagonal systems of linear equations. The Computer Journal, 15(4):356–359, November 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/150356.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/356.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/357.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/358.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/tiff/359.tif.

Evans:1975:RAD

[Eva75] D. J. Evans. A recursive algorithm for determining the eigenvalues of a quindiagonal matrix. The Computer Journal, 18(1):70–73, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        01/tiff/70.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_18/Issue_01/tiff/
                                        71.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_18/Issue_01/tiff/72.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        18/Issue_01/tiff/73.tif.

Evans:1977:UFM

[Eva77] D. J. Evans. On the use of fast methods for solving boundary value problems. The Computer Journal, 20(2):181–184, May 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        02/tiff/181.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        02/tiff/182.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        02/tiff/183.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_20/Issue_
02/tiff/184.tif.

Eve:1963:SAI

[Eve63] J. Eve. Starting approximation for the iterative calculation of square roots. The Computer Journal, 6(3):274–276, October 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/060274.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/274.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/275.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/276.tif.

Evans:1980:ADP

[EW80] D. J. Evans and Shirley A.

Williams. Analysis and detection of parallel processable code. The Computer Journal, 23(1):66–72, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
01/tiff/66.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/
67.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_23/Issue_01/tiff/68.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
23/Issue_01/tiff/69.tif;

                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        01/tiff/70.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_23/Issue_01/tiff/
                                        71.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_23/Issue_01/tiff/72.
                                        tif.

Episkopou:1986:TFC

[EWH86] D. M. Episkopou and A. T. Wood-Harper. Towards a framework to choose appropriate IS approaches. The Computer Journal, 29(3):222–228, June 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        03/tiff/222.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        03/tiff/223.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        03/tiff/224.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        03/tiff/225.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        03/tiff/226.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        03/tiff/227.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        03/tiff/228.tif.

Farmakides:1973:LAC

[F^{+}73] John B. Farmakides et al. Legal aspects of computerized information systems. The Computer Journal, 7(1):??, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Franklin:1988:AGP

[FA88] W. R. Franklin and V. Akman.

Adaptive grid for polyhedral visibility in object space: an implementation. The Computer Journal, 31(1):56–60, February 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
01/tiff/56.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_01/tiff/
57.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_31/Issue_01/tiff/58.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
31/Issue_01/tiff/59.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
01/tiff/60.tif.

Fairthorne:1958:ARR

[Fai58] R. A. Fairthorne. Automatic retrieval of recorded information. The Computer Journal, 1(1):36–41, April 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
01/010036.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
01/tiff/36.tif;   http://

www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_01/tiff/
37.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_01/Issue_01/tiff/38.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
01/Issue_01/tiff/39.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
01/tiff/40.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_01/tiff/
41.tif.

Falla:1981:GSE

[Fal81] M. E. Falla. Gamma software engineering system. The Computer Journal, 24(3):235– 242, August 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
24/Issue_03/tiff/235.tif;
 http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/236.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/237.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/238.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/239.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/240.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/241.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_24/Issue_
03/tiff/242.tif.

Farr:1978:AGA

[Far78] G. L. Farr. Algorithm 104: Generation of the alphabetically ordered sequences from a set of N distinct objects. The Computer Journal, 21(4):373, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/373.tif.

Fleck:1975:AMN

[FB75] R. Fleck and J. Bailey. Algorithm 87: Minimum of a nonlinear function by the application of the geometric programming technique. The Computer Journal, 18(1):86–89, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
01/tiff/86.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_01/tiff/
87.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_18/Issue_01/tiff/88.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
18/Issue_01/tiff/89.tif.

Froese:1999:CMM

[FB99] Kevin W. Froese and Richard B.

Bunt. Cache management for mobile file service. The Computer Journal, 42(6):442– 454, 1999. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
06/420442.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
06/pdf/420442.pdf.

Foster:1987:RC

[FC87] J. M. Foster and I. F. Currie. Remote capabilities. The Computer Journal, 30 (5):451–457, October 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/451.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/452.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/453.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/454.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/455.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/456.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/457.tif.

Fitzpatrick:1996:TAF

[FCHK96] Stephen Fitzpatrick, M. Clint, T. J. Harmer, and P. Kilpatrick. The tailoring of abstract functional specifications of numerical algorithms for sparse data structures through automated program derivation and transformation. The Computer Journal, 39(2):145–168, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www.oup.co.uk/jnls/
list/comjnl/hdb/Volume_39/
Issue_02/390145.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_02/
Vol39_02.body.html#AbstractFitzpatrick.

Fang:1988:CPC

[FD88] Z. Fang and J. S. Deogun.

Closure properties of certain classes of languages under generalized morphic replication. The Computer Journal, 31(4):325–329, August 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/325.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/326.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/327.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/328.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/329.tif.

Fotheringham:1958:IRF

[FdVR58] J. A. Fotheringham and M. de V. Roberts.

An input routine for the Ferranti Mercury computer. The Computer Journal, 1 (3):128–131, October 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/010128.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/128.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/129.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/130.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/131.tif.

Fox:1969:ISD

[FE69] A. J. Fox and P. W. Edwards. Implementation of a syntax-driven interpreter for data retrieval. The Computer Journal, 12(3):225–232, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/120225.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/225.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/226.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/227.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_12/Issue_
03/tiff/228.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/229.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/230.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/231.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/232.tif.

Fellget:1980:BRP

[FE80] P. B. Fellget and C. D. Easteal. Book reviews: The Phenomenon of Science: A Cybernetic Approach to Human Evolution, by V. F. Turchin, 1977; Business Systems Handbook, by Robert W. Gilmour, 1979. The Computer Journal, 23(2):114, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/114.tif.

Fellgett:1969:CWA

[Fel69] Peter Fellgett. Correspondence:

What is an analyst? The Computer Journal, 12(1):104, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
01/tiff/104.tif. See corre-

spondence [Fel70, AT70].

Fellgett:1970:CWA

[Fel70] P. B. Fellgett. Correspondence: What is an analyst? The Computer Journal, 13(1):118–119, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        01/tiff/118.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        01/tiff/119.tif. See [Fel69,

AT70].

Fellgett:1972:CWO

[Fel72] P. B. Fellgett. Correspondence: The wonderland of operating systems. The Computer Journal, 15(2):131, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        02/tiff/131.tif. See note

[Bac71, ER71, Wal71, Hil72].

Fenton:1974:MS

[Fen74]  J. S. Fenton.  Memoryless

subsystems. The Computer Journal, 17(2):143–147, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        02/tiff/143.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        02/tiff/144.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_17/Issue_
02/tiff/145.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/146.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/147.tif.

Fenwick:1987:FAC

[Fen87] P. M. Fenwick. A fast-carry adder with CMOS transmission gates. The Computer Journal, 30(1):77–79, February 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
01/tiff/77.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_01/tiff/
78.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_30/Issue_01/tiff/79.
tif.

Fenwick:1990:PMK

[Fen90] P. Fenwick. Performance measurements on a 39 km extended Ethernet. The Computer Journal, 33(4):314–322, August 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/314.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/315.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/316.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/317.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/318.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/319.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/320.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/321.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/322.tif.

Fenwick:1991:NTS

[Fen91] P. M. Fenwick. A new technique for self-organising linear searches. The Computer Journal, 34(5):450–454, October 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/450.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/451.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/452.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/453.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/454.tif.

Fenwick:1996:BWT

[Fen96] P. M. Fenwick. The Burrows– Wheeler Transform for block sorting text compression: principles and improvements. The Computer Journal, 39(9):731– ??, ???? 1996. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www.oup.co.uk/
computer_journal/Volume_39/
Issue_09/Vol39_09.body.html#
AbstractFenwick;   http:/
/www3.oup.co.uk/computer_
journal/Volume_39/Issue_09/
Vol39_09.body.html#AbstractFenwick.

Fettis:1965:NME

[Fet65] H. E. Fettis. Note on the matrix equation Ax = λBx. The Computer Journal, 8(3):279, October 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/080279.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/279.tif.

Fair:1966:NCA

[FFMR66] G. R. Fair, A. D. J. Flowerdew, W. G. Munro, and D. Rowley. Note on the computer as an aid to the architect. The Computer Journal, 9(1):16–20, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/090016.sgm.abs.html;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/16.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_01/tiff/
17.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_09/Issue_01/tiff/18.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
09/Issue_01/tiff/19.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/20.tif.

Fontana:1999:CMI

[FFP99] Francesca Arcelli Fontana, Ferrante Formato, and Remo Pareschi. Computational models for information reuse. The Computer Journal, 42(7):582– 591, 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
07/420582.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
07/pdf/420582.pdf.

Forbes:1977:PAS

[FG77] K. Forbes and A. W. Goldsworthy. Prescheduling algorithm — scheduling a suitable mix prior to processing. The Computer Journal, 20(1):27–29, February 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
01/tiff/27.tif;   http://
www3.oup.co.uk/computer_journal/

hdb/Volume_20/Issue_01/tiff/
28.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_20/Issue_01/tiff/29.
tif.

Feldman:1990:OAS

[FG90] R. Feldman and M. C. Golumbic.

Optimization algorithms for student scheduling via constraint satisfiability. The Computer Journal, 33(4):356– 364, August 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
33/Issue_04/tiff/356.tif;
 http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/357.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/358.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/359.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/360.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/361.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/362.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/363.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/364.tif.

Fletcher:1971:CLB

[FGH71] R. Fletcher, J. A. Grant, and M. D. Hebden. The calculation of linear best L_{p} approximations. The Computer Journal, 14(3):276–279, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/140276.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/tiff/276.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/tiff/277.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/tiff/278.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/tiff/279.tif.

Forrington:1967:AOC

[FGR67] C. V. D. Forrington, S. M. Grey, and M. M. Reid. An application-oriented computer approach to inventory control problems. Part 2: Implementation and operational experience. The Computer Journal, 9(4):325–329, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        04/090325.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        04/tiff/325.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/326.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/327.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/328.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/329.tif.

Foley:1971:PRP

[FH71] M. Foley and C. A. R. Hoare.

Proof of a recursive program: Quicksort. The Computer Journal, 14(4):391–395, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/140391.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/391.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/392.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/393.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/394.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/395.tif.

Ford:1977:LLA

[FH77] W. S. Ford and V. C. Hamacher.

Low level architecture features for supporting process communication. The Computer Journal, 20(2):156–162, May 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/156.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/157.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/158.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/159.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/160.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/161.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/162.tif.

Frey:1997:ESS

[FH97] B. J. Frey and G. E. Hinton. Efficient stochastic source coding and an application to a Bayesian network source model. The Computer Journal, 40(2/3):157–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www.oup.co.uk/computer_
journal/Volume_40/Issue_02/
Vol40_02.body.html#AbstractFrey;
 http://www3.oup.co.uk/computer_
journal/Volume_40/Issue_02/
Vol40_03.body.html#AbstractFrey.

Franti:1998:CMA

[FH98] Pasi Franti and Timo Hatakka.

Context model automata for text compression. The Computer Journal, 41(7):474–485, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_41/Issue_07/
Vol41_07.body.html#AbstractFranti.

Fielding:1970:NAS

[Fie70] K. Fielding. Note on Algorithm 44: Solution of nonlinear simultaneous equations. The Computer Journal, 13(2):219, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/219.tif. See [Bro69c,

Red71].

Fike:1975:PGM

[Fik75] C. T. Fike. A permutation generation method. The Computer Journal, 18(1):21–22, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
01/tiff/21.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_01/tiff/
22.tif.

Findler:1960:SRG

[Fin60]  N. V. Findler.  Some re-

marks on the game ‘dama’ which can be played on a digital computer. The Computer Journal, 3(1):40–44, April 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
01/030040.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
01/tiff/40.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_01/tiff/
41.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_03/Issue_01/tiff/42.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
03/Issue_01/tiff/43.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
01/tiff/44.tif.

Finkelstein:1968:COT

[Fin68] M. Finkelstein. A compiler optimization technique. The Computer Journal, 11(1):22–25, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/110022.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/22.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_01/tiff/
23.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_11/Issue_01/tiff/24.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/25.tif.

Finn:1972:CSE

[Fin72] A. J. Finn. Correspondence:

Suggested extensions to FOR-TRAN IV. The Computer Journal, 15(1):12, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/150012.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/tiff/12.tif.  See cor-

respondence [Mee72, Hal72, Ehr72].

Finkelstein:1993:ECC

[Fin93] Anthony Finkelstein. European computing curricula: a guide and comparative analysis. The Computer Journal, 36(4):299–319, August 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_04/
Vol36_04.body.html#AbstractFinklestein1.

Firth:1961:OPS

[Fir61] A. W. O. Firth. Optimization problems: solution by an analogue computer. The Computer Journal, 4(1):68–72, April 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/040068.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_

                                        01/tiff/68.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_04/Issue_01/tiff/
                                        69.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_04/Issue_01/tiff/70.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        04/Issue_01/tiff/71.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        01/tiff/72.tif.

Fisher:1974:CPF

[Fis74a]  D. L. Fisher.  Correspon-

dence: On “The project, and the future of Computing Science courses”. The Computer Journal, 17(2):190, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        02/tiff/190.tif. See [Bro73a,

Bel74].

Fisher:1974:CES

[Fis74b] G. A. Fisher. Correspondence: On “Editorial: the silent business user”. The Computer Journal, 17(1):95–96, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        01/tiff/95.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_17/Issue_01/tiff/

96.tif. See [Gea73, TT73].

Fisher:1976:BPM

[Fis76]  D. L. Fisher.  Blind pro-

grammers — their manager’s experience. The Computer Journal, 19(2):189–190, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/189.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/190.tif.

Fitch:1973:AS

[Fit73] J. P. Fitch. On algebraic simplification. The Computer Journal, 16(1):23–27, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
01/160023.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
01/tiff/23.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_01/tiff/
24.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_16/Issue_01/tiff/25.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
16/Issue_01/tiff/26.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
01/tiff/27.tif.

Fox:1966:FER

[FJ66] A. J. Fox and F. A. Johnson. On finding the eigenvalues of real symmetric tridiagonal matrices. The Computer Journal, 9(1):98–105, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/090098.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/100.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/101.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/102.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/103.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/104.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/105.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/98.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_01/tiff/
99.tif.

Ford:1980:BRW

[FJA80] Brian Ford, David Jacobs, and A. M. Addyman. Book reviews: Workshop on Reliable Software: Applied Computer Science, by Peter Raulefs, 1979; Reliable Software Through Composite Design, by G. J. Myers, 1975; Fortran, PL/1 and the Algols, by Brian Meek, 1978. The Computer Journal, 23(2):146, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/146.tif.

Foxley:1968:ISA

[FK68] E. Foxley and P. King. The implementation of syntax analysis using ALGOL, and some mathematical applications. The Computer Journal, 10(4):325– 335, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
10/Issue_04/100325.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
10/Issue_04/tiff/325.tif;
 http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/326.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/327.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/328.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/329.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/330.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/331.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/332.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/tiff/333.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/tiff/334.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/tiff/335.tif.

Fucik:1986:HPC

[FK86] J. Fucik and J. Kral. The hierarchy of program control structures. The Computer Journal, 29(1):24–32, February 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        01/tiff/24.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_29/Issue_01/tiff/
                                        25.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_29/Issue_01/tiff/26.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        29/Issue_01/tiff/27.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        01/tiff/28.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_29/Issue_01/tiff/
                                        29.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_29/Issue_01/tiff/30.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        29/Issue_01/tiff/31.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        01/tiff/32.tif.

Freisleben:1989:PS

[FK89] B. Freisleben and J. L. Keedy.

Priority semaphores. The Computer Journal, 32(1):24– 28, February 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
32/Issue_01/tiff/24.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
01/tiff/25.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_01/tiff/
26.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_32/Issue_01/tiff/27.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
32/Issue_01/tiff/28.tif.

Fraenkel:1990:BHC

[FK90] A. S. Fraenkel and S. T.

Klein. Bidirectional Huffman coding. The Computer Journal, 33(4):296–307, August 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/296.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/297.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/298.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/299.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_33/Issue_
                                        04/tiff/300.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        04/tiff/301.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        04/tiff/302.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        04/tiff/303.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        04/tiff/304.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        04/tiff/305.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        04/tiff/306.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        04/tiff/307.tif.

Fraenkel:1993:BDS

[FK93] A. S. Fraenkel and S. T. Klein. Bounding the depth of search trees. The Computer Journal, 36(7):668–678, ???? 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_36/Issue_07/
                                        Vol36_07.body.html#AbstractFraenkel.

Finkelstein:1993:IES

[FKA^{+}93] Anthony Finkelstein, Jeff Kramer, Samson Abramsky, Krysia Broda, Sophia Drossopoulou, and Susan Eisenbach. An integrated engineering study scheme in computing. The Computer Journal, 36(4):320– 334, August 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/Volume_
36/Issue_04/Vol36_04.body.
html#AbstractFinklestein2.

Ferrie:1976:ESP

[FKLM76] J. Ferrie, C. Kaiser, D. Lanciaux, and B. Martin. An extensible structure for protected systems’ design. The Computer Journal, 19(4):315–321, November 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/315.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/316.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/317.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/318.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/319.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/320.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/321.tif.

Foster:1990:CSC

[FKT90] I. Foster, C. Kesselman, and S. Taylor. Concurrency: Simple concepts and powerful tools. The Computer Journal, 33(6):501–507, December 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/501.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/502.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/503.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/504.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/505.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/506.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/507.tif.

Foxley:1968:CET

[FL68] E. Foxley and K. Lockyer. The construction of examination timetables by computer. The Computer Journal, 11(3):264–268, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/110264.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/264.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_

03/tiff/265.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/266.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/267.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/268.tif.

Fenner:1980:BTR

[FL80] T. I. Fenner and G. Loizou.

A binary tree representation and related algorithms for generating integer partitions. The Computer Journal, 23 (4):332–337, November 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/332.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/333.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/334.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/335.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/336.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/337.tif.

Flynn:1985:MCS

[FL85] D. J. Flynn and A. H. F. Laender. Mapping from a conceptual schema to a target internal schema. The Computer Journal, 28(5):508–517, November 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/508.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/509.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/510.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/511.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/512.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/513.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/514.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/515.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/516.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/517.tif.

Frost:1989:CNL

[FL89] R. Frost and J. Launchbury. Constructing natural language interpreters in a lazy functional language. The Computer Journal, 32(2):108–121, April 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/108.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/109.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/110.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/111.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/112.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/113.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/114.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/115.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/116.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/117.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/118.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/119.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/120.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/121.tif.

Flavell:1972:CAR

[Fla72]  A. J. Flavell.  Correspon-

dence: On “Another round of FORTRAN”. The Computer Journal, 15(1):92, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        01/150092.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        01/tiff/92.tif. See [Cha71,

Dew72b, Hal72, Cha72].

Fletcher:1965:FME

[Fle65] R. Fletcher. Function minimization without evaluating derivatives — a review. The Computer Journal, 8(1):33–41, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        01/080033.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        01/tiff/33.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_08/Issue_01/tiff/
                                        34.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_08/Issue_01/tiff/35.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        08/Issue_01/tiff/36.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        01/tiff/37.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_08/Issue_01/tiff/

38.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_08/Issue_01/tiff/39.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
08/Issue_01/tiff/40.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
01/tiff/41.tif.

Fletcher:1968:GIM

[Fle68] R. Fletcher. Generalized inverse methods for the best least squares solution of systems of non-linear equations. The Computer Journal, 10 (4):392–399, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/100392.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/392.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/393.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/394.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/395.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/396.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/397.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/398.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/tiff/399.tif.

Fletcher:1970:NAV

[Fle70]  R. Fletcher.  A new ap-

proach to variable metric algorithms. The Computer Journal, 13(3):317–322, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/130317.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/317.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/318.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/319.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/320.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/321.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/322.tif.

Fleck:1975:CEL

[Fle75] Robert A. Fleck, Jr. Computer efficiency and linear programming. The Computer Journal, 18(2):115–116, May 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_

02/tiff/115.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/116.tif.

Florentin:1974:CAD

[Flo74]  J. J. Florentin.  Consis-

tency auditing of databases. The Computer Journal, 17 (1):52–58, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
01/tiff/52.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/
53.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_17/Issue_01/tiff/54.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
17/Issue_01/tiff/55.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
01/tiff/56.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/
57.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_17/Issue_01/tiff/58.
tif.

Florentin:1976:DBR

[Flo76] J. J. Florentin. Data base representations of application models. The Computer Journal, 19(1):13–16, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_

                                        01/tiff/13.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_19/Issue_01/tiff/
                                        14.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_19/Issue_01/tiff/15.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        19/Issue_01/tiff/16.tif.

Fong:1999:DCB

[FLW99] Cedric C. F. Fong, John C. S. Lui, and M. H. Wong. Distributed caching and broadcast in a wireless mobile computing environment. The Computer Journal, 42(6):455–472, 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        06/420455.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        06/pdf/420455.pdf.

Fairweather:1964:GAD

[FM64] G. Fairweather and A. R. Mitchell. A generalized alternating direction method of Douglas-Rachford type for solving the biharmonic equation. The Computer Journal, 7(3):242–245, October 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        03/070242.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        03/tiff/242.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/243.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/244.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/245.tif.

Fairweather:1966:SCR

[FM66] G. Fairweather and A. R.

Mitchell. Some computational results of an improved A.D.I. method for the Dirichlet problem. The Computer Journal, 9(3):298–303, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/090298.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/298.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/299.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/300.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/301.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/302.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/303.tif.

Fraenkel:1983:CCP

[FM83] Aviezri S. Fraenkel and Moshe Mor. Combinatorial compression and partitioning of large dictionaries. The Computer Journal, 26(4):336–343, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/336.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/337.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/338.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/339.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/340.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/341.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/342.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/343.tif.

Furtado:1985:IAF

[FM85] A. L. Furtado and T. S. E. Maibaum. An informal approach to formal (algebraic) specifications. The Computer Journal, 28(1):59–67, February 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
01/tiff/59.tif;   http://

www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_01/tiff/
60.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_28/Issue_01/tiff/61.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
28/Issue_01/tiff/62.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
01/tiff/63.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_01/tiff/
64.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_28/Issue_01/tiff/65.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
28/Issue_01/tiff/66.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
01/tiff/67.tif.

Fabbrini:1986:AQ

[FM86a] F. Fabbrini and C. Montani. Autumnal quadtrees. The Computer Journal, 29 (5):472–474, October 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/472.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/473.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/474.tif.

Feldman:1986:EMC

[FM86b] P. Feldman and D. Miller. Entity model clustering: Structuring a data model by abstraction. The Computer Journal, 29(4):348–360, August 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/348.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/349.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/350.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/351.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/352.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/353.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/354.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/355.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/356.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/357.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/358.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/359.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_

04/tiff/360.tif.

Fitch:1978:NCG

[FN78] J. P. Fitch and A. C. Norman.

A note on compacting garbage collection. The Computer Journal, 21(1):31–34, February 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
01/tiff/31.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_01/tiff/
32.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_21/Issue_01/tiff/33.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
21/Issue_01/tiff/34.tif.

Franti:1993:TSM

[FN93] Pasi Fränti and Olli Nevalainen.

A two-stage modelling method for compressing binary images by arithmetic coding. The Computer Journal, 36(7):615–622, ???? 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_07/
Vol36_07.body.html#AbstractFranti.

Franti:1994:CDI

[FNK94] Pasi Fränti, Olli Nevalainen, and Timo Kaukoranta. Compression of digital images by block truncation coding: a survey. The Computer Journal, 37(4):308–332, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_37/Issue_04/
                                        Vol37_04.body.html#AbstractFranti.

Fokkinga:1996:ETA

[Fok96] Maarten Fokkinga. Expressions that talk about themselves. The Computer Journal, 39(5):408– 412, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://www.oup.co.uk/jnls/
                                        list/comjnl/hdb/Volume_39/
                                        Issue_05/390408.sgm.abs.html;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_39/Issue_05/
                                        Vol39_05.body.html#AbstractFokkinga.

Foo:1986:DIP

[Foo86] N. Y. Foo. Dewey indexing of Prolog traces. The Computer Journal, 29(1):17–19, February 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        01/tiff/17.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_29/Issue_01/tiff/
                                        18.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_29/Issue_01/tiff/19.
                                        tif.

Forrington:1961:EPC

[For61] C. V. D. Forrington. Extensions of the predictor-corrector method for the solution of systems of ordinary differential equations. The Computer Journal, 4(1):80–85, April 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/040080.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/tiff/80.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_01/tiff/
81.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_04/Issue_01/tiff/82.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
04/Issue_01/tiff/83.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/tiff/84.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_01/tiff/
85.tif.

Forbes:1965:ICW

[For65] J. M. Forbes. An introduction to compiler writing. The Computer Journal, 8(2):98–102, July 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/080098.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/100.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/101.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/102.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_08/Issue_
                                        02/tiff/98.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_08/Issue_02/tiff/
                                        99.tif.

Forrington:1966:AOC

[For66] C. V. D. Forrington. An application-oriented computer approach to inventory control problems. Part 1: Facilities of the system. The Computer Journal, 9(2):129–136, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        02/090129.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        02/tiff/129.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        02/tiff/130.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        02/tiff/131.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        02/tiff/132.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        02/tiff/133.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        02/tiff/134.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        02/tiff/135.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        02/tiff/136.tif.

Forrest:1972:IIA

[For72]  A. R. Forrest.  Interac-

tive interpolation and approximation by Bézier polynomials. The Computer Journal, 15(1):71–79, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/150071.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/tiff/71.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_01/tiff/
72.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_15/Issue_01/tiff/73.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
15/Issue_01/tiff/74.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/tiff/75.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_01/tiff/
76.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_15/Issue_01/tiff/77.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
15/Issue_01/tiff/78.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/tiff/79.tif.

Ford:1975:APC

[For75]  W. S. Ford.  Algorithm

90: Periodic cubic spline interpolation with equidistant nodes. The Computer Journal, 18(2):183–184, May 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/183.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/184.tif.

Foster:1964:CTG

[Fos64] F. G. Foster. A computer technique for game-theoretic problems I: chemin-de-fer analyzed. The Computer Journal, 7(2):124–130, July 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/070124.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/124.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/125.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/126.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/127.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/128.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/129.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/130.tif.

Foster:1965:SLP

[Fos65] D. M. Foster. A simple listprocessing interpreter. The Computer Journal, 8(2):120– 129, July 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
08/Issue_02/080120.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
08/Issue_02/tiff/120.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/121.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/122.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/123.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/124.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/125.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/126.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/127.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/128.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/129.tif.

Foster:1968:SAO

[Fos68a] C. C. Foster. Sorting almost ordered arrays. The Computer Journal, 11(2):134–137, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/110134.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/134.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/135.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/136.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/137.tif.

Foster:1968:SIP

[Fos68b] J. M. Foster. A syntax improving program. The Computer Journal, 11(1):31–34, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/110031.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/31.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_01/tiff/
32.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_11/Issue_01/tiff/33.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/34.tif.

Foster:1971:UCP

[Fos71] C. C. Foster. Uncoupling central processor and storage device speeds. The Computer Journal, 14(1):45–48, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/140045.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/tiff/45.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_01/tiff/
46.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_14/Issue_01/tiff/47.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
14/Issue_01/tiff/48.tif.

Foster:1986:VMA

[Fos86]  J. M. Foster.  Validat-

ing microcode algebraically. The Computer Journal, 29 (5):416–422, October 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/416.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/417.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/418.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/419.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        05/tiff/420.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        05/tiff/421.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        05/tiff/422.tif.

Fox:1961:CMT

[Fox61]  L. Fox.  Computing ma-

chines for teaching and research. The Computer Journal, 4(3):212–216, October 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        03/040212.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        03/tiff/212.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        03/tiff/213.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        03/tiff/214.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        03/tiff/215.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        03/tiff/216.tif.

Fox:1962:CMO

[Fox62]  L. Fox.  Chebyshev meth-

ods for ordinary differential equations. The Computer Journal, 4(4):318–331, January 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/040318.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/318.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/319.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/320.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/321.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/322.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/323.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/324.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/325.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/326.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/327.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/328.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/329.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/330.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_04/Issue_
                                        04/tiff/331.tif.

Fox:1963:PDE

[Fox63] L. Fox. Partial differential equations. The Computer Journal, 6(1):69–74, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        01/060069.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        01/tiff/69.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_06/Issue_01/tiff/
                                        70.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_06/Issue_01/tiff/71.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        06/Issue_01/tiff/72.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        01/tiff/73.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_06/Issue_01/tiff/
                                        74.tif.

Fox:1967:RIC

[Fox67] L. Fox. Romberg integration for a class of singular integrands. The Computer Journal, 10(1):87–93, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        01/100087.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_

01/tiff/87.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_01/tiff/
88.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_10/Issue_01/tiff/89.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/90.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/91.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_01/tiff/
92.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_10/Issue_01/tiff/93.
tif.

Foxley:1978:HCP

[Fox78] Eric Foxley. A hybrid computer performance modelling system. The Computer Journal, 21(3):205–209, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/205.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/206.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/207.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/208.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/209.tif.

Fox:1987:ODJ

[Fox87] Leslie Fox. Obituary: Dr. J. H. Wilkinson, F.R.S. The Computer Journal, 30(1):1, 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        01/tiff/1.tif.

Fletcher:1963:RCD

[FP63] R. Fletcher and M. J. D. Powell. A rapidly convergent descent method for minimization. The Computer Journal, 6(2):163–168, July 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/060163.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/tiff/163.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/tiff/164.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/tiff/165.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/tiff/166.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/tiff/167.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/tiff/168.tif.

Frost:1982:SNB

[FP82] R. A. Frost and M. M. Peterson. A short note on binary search trees. The Computer Journal, 25(1):158, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/158.tif.

Fletcher:1963:MAD

[FR63] R. Fletcher and C. M. Reeves.

A mechanization of algebraic differentiation and the automatic generation of formulae for molecular integral of Gaussian orbitals. The Computer Journal, 6(3):287–292, October 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/060287.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/287.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/288.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/289.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/290.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/291.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_06/Issue_
03/tiff/292.tif.

Fletcher:1964:FMC

[FR64] R. Fletcher and C. M. Reeves. Function minimization by conjugate gradients. The Computer Journal, 7(2):149–154, July 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/070149.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/149.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/150.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/151.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/152.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/153.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/154.tif.

Feng:1980:SNS

[FR80] D. Y. Feng and R. F. Riesenfeld. Some new surface forms for computer aided geometric design. The Computer Journal, 23(4):324–331, November 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_

04/tiff/324.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/325.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/326.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/327.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/328.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/329.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/330.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/331.tif.

Faidhi:1987:PEL

[FR87] J. A. W. Faidhi and S. K.

Robinson. Programmer experience level indicators. The Computer Journal, 30(1):52– 62, February 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
30/Issue_01/tiff/52.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
01/tiff/53.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_01/tiff/
54.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_30/Issue_01/tiff/55.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_

                                        30/Issue_01/tiff/56.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        01/tiff/57.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_30/Issue_01/tiff/
                                        58.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_30/Issue_01/tiff/59.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        30/Issue_01/tiff/60.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        01/tiff/61.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_30/Issue_01/tiff/
                                        62.tif.

Fraley:1998:HMC

[FR98] Chris Fraley and Adrian E. Raftery. How many clusters? which clustering method? answers via model-based cluster analysis. The Computer Journal, 41(8):578–588, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_41/Issue_08/
                                        Vol41_08.body.html#AbstractFraley.

Francis:1961:TUA

[Fra61] J. G. F. Francis. The QR transformation a unitary analogue to the LR transformation — part 1. The Computer Journal, 4(3):265–271, October 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_

03/040265.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/265.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/266.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/267.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/268.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/269.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/270.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/271.tif. See also

[Fra62, Wat08, Wat11].

Francis:1962:TP

[Fra62]  J. G. F. Francis.  The

QR transformation — part

2. The Computer Journal, 4(4):332–345, January 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/040332.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/332.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/333.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_

                                        04/tiff/334.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        04/tiff/335.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        04/tiff/336.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        04/tiff/337.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        04/tiff/338.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        04/tiff/339.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        04/tiff/340.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        04/tiff/341.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        04/tiff/342.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        04/tiff/343.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        04/tiff/344.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        04/tiff/345.tif. See also

[Fra61, Wat08, Wat11].

Fraser:1967:DCA

[Fra67] A. G. Fraser. Data compression and automatic programming. The Computer Journal, 10(2):165–167, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/100165.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/165.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/166.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/167.tif.

Fraser:1968:UCM

[Fra68] A. G. Fraser. User control in a multi-access system. The Computer Journal, 11(1):12–16, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/110012.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/12.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_01/tiff/
13.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_11/Issue_01/tiff/14.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/15.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/16.tif.

Fraser:1969:IMS

[Fra69] A. G. Fraser. Integrity of a mass storage filing system. The Computer Journal, 12(1):1–5, February 1969. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        01/120001.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        01/tiff/1.tif;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_12/Issue_01/tiff/
                                        2.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_12/Issue_01/tiff/3.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        12/Issue_01/tiff/4.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        01/tiff/5.tif.

Fraser:1970:CFC

[Fra70] J. G. Fraser. Correspondence: Further comments on a linethinning scheme. The Computer Journal, 13(2):221–222, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/tiff/221.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/tiff/222.tif. See corre-

spondence [Deu69].

Frank:1976:JCM

[Fra76] G. R. Frank. Job control in the MU5 operating system. The Computer Journal, 19(2):139–143, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/139.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/140.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/141.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/142.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/143.tif.

Freeman:1974:CH

[Fre74] K. Freeman. Correspondence:

On “Himmelbett”. The Computer Journal, 17(2):189, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/189.tif. See [Bel73,

Str74].

Frielink:1965:AED

[Fri65] A. B. Frielink. The auditor and electronic data processing. The Computer Journal, 8(2):95–97, July 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/080095.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/95.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_02/tiff/

                                        96.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_08/Issue_02/tiff/97.
                                        tif.

Frost:1981:ADI

[Fro81]  R. A. Frost.  Algorithm

112: Dumping the index of a dynamic hash table. The Computer Journal, 24 (4):383–384, November 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        04/tiff/383.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        04/tiff/384.tif.

Frost:1982:BSS

[Fro82] R. A. Frost. Binary-relational storage structures. The Computer Journal, 25(3):358– 367, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        25/Issue_03/tiff/358.tif;
                                         http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        03/tiff/359.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        03/tiff/360.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        03/tiff/361.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        03/tiff/362.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_25/Issue_
03/tiff/363.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/364.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/365.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/366.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/367.tif. Discusses

the dynamic hashing scheme used by ASDAS, under development at Strathclyde University.

Frost:1983:SYA

[Fro83] R. A. Frost. SCHEMAL: Yet another conceptual schema definition language. The Computer Journal, 26(3):228–234, August 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/228.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/229.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/230.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/231.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/232.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_

                                        03/tiff/233.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        03/tiff/234.tif.

Frost:1985:USC

[Fro85] R. A. Frost. Using semantic concepts to characterise various knowledge representation formalisms: a method of facilitating the interface of knowledge base system components. The Computer Journal, 28(2):112–116, May 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        02/tiff/112.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        02/tiff/113.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        02/tiff/114.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        02/tiff/115.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        02/tiff/116.tif.

Frost:1992:CPE

[Fro92] R. A. Frost. Constructing programs as executable attribute grammars. The Computer Journal, 35(4):376–389, August 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_35/Issue_04/
                                        Vol35_04.body.html#AbstractFrost.

Forrest:1987:INN

[FRS^{+}87] B. M. Forrest, D. Roweth, N. Stroud, D. J. Wallace, and G. V. Wilson. Implementing neural network models on parallel computers. The Computer Journal, 30(5):413–419, October 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/413.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/414.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/415.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/416.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/417.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/418.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/419.tif.

Fryer:1971:BRR

[Fry71] J. G. Fryer. Book review:

Rank Order Probabilities: Two Sample Normal Shift Alternatives, by Roy C. Milton, 1970. The Computer Journal, 14(1):107, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_

journal/hdb/Volume_14/Issue_
01/tiff/107.tif.

Fraser:1966:CLO

[FS66] A. G. Fraser and J. D. Smart. The COMPL language and operating system. The Computer Journal, 9(2):144–156, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/090144.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/144.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/145.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/146.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/147.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/148.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/149.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/150.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/151.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/152.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/153.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/154.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/155.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/156.tif.

Florentin:1975:SSR

[FS75] J. J. Florentin and A. J.

Sammes. Systems with state re-set. The Computer Journal, 18(2):135–139, May 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/135.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/136.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/137.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/138.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/139.tif.

Foxley:1981:VAM

[FS81] E. Foxley and O. Salman.

Validation of an analytic model of computer performance. The Computer Journal, 24(4):347–352, November 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

journal/hdb/Volume_24/Issue_
04/tiff/347.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/348.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/349.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/350.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/351.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/352.tif.

Firby:1987:CMS

[FS87] P. A. Firby and D. J. Stone. Colour manipulation of superposed families of curves. The Computer Journal, 30 (4):349–354, August 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/349.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/350.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/351.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/352.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/353.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_30/Issue_
04/tiff/354.tif.

Firby:1994:ASR

[FS94] P. A. Firby and D. J. Stone.

Analysis of the structure and resolution of a graphics device. The Computer Journal, 37(5):367–376, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_37/Issue_05/
Vol37_05.body.html#AbstractFirby.

Fitzgerald:1985:FAC

[FSW85] G. Fitzgerald, N. Stokes, and J. R. G. Wood. Feature analysis of contemporary information systems methodologies. The Computer Journal, 28(3):223–230, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/223.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/224.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/225.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/226.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/227.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/228.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        03/tiff/229.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        03/tiff/230.tif.

Farlie:1982:BRW

[FSWL82] D. J. G. Farlie, Mildred L. G. Shaw, B. A. Wichmann, and S. H. Lavington. Book reviews: W. M. Brelsford and R. A. Relles, STATLILB: A Statistical Computing Laboratory, 1981; N. Rushby (ed.), Selected Readings in Computer-Based Learning, 1981; R. D. Tennent, Principles of Programming Languages, 1981; D. Hutchison, Fundamentals of Computer Logic, 1981. The Computer Journal, 25(2):287, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/287.tif.

Fairthorne:1973:CWU

[FT73] Robert A. Fairthorne and The Editor. Correspondence: On word use. The Computer Journal, 16(3):286, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        03/tiff/286.tif. See [WNT73].

Fuhr:1992:PMI

[Fuh92]  Norbert Fuhr.  Probabilis-

tic models in information retrieval. The Computer Journal, 35(3):243–255, June 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_35/Issue_03/
Vol35_03.body.html#AbstractFuhr.

Fuller:1976:PPN

[Ful76] A. T. Fuller. The period of pseudo-random numbers generated by Lehmer’s congruential method. The Computer Journal, 19(2):173–177, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/173.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/174.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/175.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/176.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/177.tif.

Fung:1984:RMB

[Fun84] K. T. Fung. A reorganization model based on the database entropy concept. The Computer Journal, 27(1):67–71, February 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_

                                        01/tiff/67.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_27/Issue_01/tiff/
                                        68.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_27/Issue_01/tiff/69.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        27/Issue_01/tiff/70.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        01/tiff/71.tif.

Fung:1985:EAS

[Fun85]  K. T. Fung.  An entropy

approach to solving some network reliability problems. The Computer Journal, 28 (4):353–356, August 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        04/tiff/353.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        04/tiff/354.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        04/tiff/355.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        04/tiff/356.tif.

Furuta:1989:OTA

[Fur89] R. Furuta. An object-based taxonomy for abstract structure in document models. The Computer Journal, 32 (6):494–504, December 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/494.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/495.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/496.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/497.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/498.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/499.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/500.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/501.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/502.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/503.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/504.tif.

Furber:1992:SNS

[Fur92] D. Furber. Short notes: A survey of the teaching of programming to computing undergraduates in U.K. Universities and Polytechnics. The Computer Journal, 35(5):530–533, October 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_35/Issue_05/
                                        350530.sgm.abs.html;   http:
                                        //www3.oup.co.uk/computer_
                                        journal/Volume_35/Issue_05/
                                        tiff/530.tif;   http://www3.
                                        oup.co.uk/computer_journal/
                                        Volume_35/Issue_05/tiff/531.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/Volume_
                                        35/Issue_05/tiff/532.tif;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_35/Issue_05/
                                        tiff/533.tif.

Flammini:1998:CIR

[FvLMS98] Michele Flammini, Jan van Leeuwen, and Alberto Marchetti-Spaccamela. The complexity of interval routing on random graphs. The Computer Journal, 41(1):16–25, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_41/Issue_01/
                                        Vol41_01.body.html#AbstractFlammini.

Frost:1983:STA

[FW83] R. A. Frost and S. Whittaker. A step towards the automatic maintenance of the semantic integrity of databases. The Computer Journal, 26(2):124–133, May 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        02/tiff/124.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        02/tiff/125.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/126.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/127.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/128.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/129.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/130.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/131.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/132.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/133.tif.

Fenton:1986:AAS

[FW86] N. E. Fenton and R. W. Whitty.

Axiomatic approach to software metrication through program decomposition. The Computer Journal, 29(4):330–339, August 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/330.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/331.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/332.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_29/Issue_
                                        04/tiff/333.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        04/tiff/334.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        04/tiff/335.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        04/tiff/336.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        04/tiff/337.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        04/tiff/338.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        04/tiff/339.tif.

Fyfe:1969:UCS

[Fyf69] D. J. Fyfe. The use of cubic splines in the solution of two-point boundary value problems. The Computer Journal, 12(2):188–192, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        02/120188.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        02/tiff/188.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        02/tiff/189.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        02/tiff/190.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_

02/tiff/191.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/192.tif.

Fyfe:1970:UCS

[Fyf70] D. J. Fyfe. The use of cubic splines in the solution of certain fourth order boundary value problems. The Computer Journal, 13(2):204–205, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/130204.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/204.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/205.tif.

Fokkink:1994:BPA

[FZ94] Wan Fokkink and Hans Zantema. Basic process algebra with iteration: completeness of its equational axioms. The Computer Journal, 37(4):259–267, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_37/Issue_04/
Vol37_04.body.html#AbstractFokkink.

Galer:1959:UCE

[Gal59] G. S. Galer. The use of computers for economic planning in the petroleum chemical industry. The Computer Journal, 2(3):145–150, October 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/020145.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/145.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/146.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/147.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/148.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/149.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/150.tif.

Gale:1971:CFE

[Gal71] J. W. Gale. Correspondence: On the February 1971 editorial. The Computer Journal, 14(4):360, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/140360.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/360.tif.

Gallaher:1973:CDG

[Gal73] L. J. Gallaher. Correspondence: On Dijkstra and the GOTO statement. The Computer Journal, 16(3):284–285, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/284.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/285.tif.

Gallard:1987:EDP

[Gal87] R. H. Gallard. An extension in the definition of a Petri net execution. The Computer Journal, 30(1):16–19, February 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
01/tiff/16.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_01/tiff/
17.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_30/Issue_01/tiff/18.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
30/Issue_01/tiff/19.tif.

Galton:1992:CLC

[Gal92a] A. Galton. Classical logic:

a crash course for beginners. The Computer Journal, 35(5):424–430, October 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_35/Issue_05/
Vol35_05.body.html#AbstractGalton1.

Galton:1992:LFM

[Gal92b] A. Galton. Logic as a formal method. The Computer Journal, 35(5):431–440, October 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_35/Issue_05/
                                        Vol35_05.body.html#AbstractGalton2.

Garg:1978:ACB

[Gar78] V. K. Garg. Algorithm 103: Computing the Bessel functions Y_{n}(x + iy) and K_{n}(x + iy). The Computer Journal, 21(3):272–276, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        03/tiff/272.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        03/tiff/273.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        03/tiff/274.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        03/tiff/275.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        03/tiff/276.tif.

Gatehouse:1965:EDP

[Gat65] J. S. Gatehouse. Electronic data processing for the international vocabulary of terms used in information processing. The Computer Journal, 7(4):271–274, January 1965.

CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/070271.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/271.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/272.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/273.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/274.tif.

Grimsdale:1961:CRD

[GB61] R. L. Grimsdale and J. M.

Bullingham. Character recognition by digital computer using a special flying-spot scanner. The Computer Journal, 4(2):129–136, July 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/040129.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/129.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/130.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/131.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/132.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        02/tiff/133.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        02/tiff/134.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        02/tiff/135.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        02/tiff/136.tif.

Gould:1969:CUT

[GB69] I. H. Gould and R. W. Bemer. Correspondence: On the use of the term ‘processor’. The Computer Journal, 12(4):411–412, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/411.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/412.tif. See [LWS68,

PB69, LWS69b].

Gupta:1983:PGP

[GB83] Phalguni Gupta and G. P. Bhattacharjee. Parallel generation of permutations. The Computer Journal, 26(2):97–105, May 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        02/tiff/100.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        02/tiff/101.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/102.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/103.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/104.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/105.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/97.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_02/tiff/
98.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_26/Issue_02/tiff/99.
tif.

Gault:1987:FAT

[GC87] D. Gault and M. Clint. A fast algorithm for the Towers of Hanoi problem. The Computer Journal, 30(4):376–378, August 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/376.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/377.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/378.tif.

Gallimore:1989:UOL

[GCS89] R. M. Gallimore, D. Coleman, and V. Stavridou. UMIST OBJ: a language for executable program specifications. The Computer Journal, 32 (5):413–421, October 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/413.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/414.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/415.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/416.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/417.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/418.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/419.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/420.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/421.tif.

Gough:1969:GPE

[GE69] N. E. Gough and J. B. A. Epton. A general procedure for evaluating the controllability of time delay feedback control systems. The Computer Journal, 12(3):273–276, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/120273.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/273.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/274.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/275.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/276.tif.

Gough:1971:NAP

[GE71] N. E. Gough and J. B. A.

Epton. Note on Algorithm 48: Procedure for the evaluation of an integral occurring in the mean square response analysis of linear systems. The Computer Journal, 14(2):215, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/215.tif. See [Lit70].

Gray:1985:PRI

[GE85] P. M. D. Gray and P. Esslemont. The performance of a relational interface to a Codasyl database. The Computer Journal, 28(5):501–507, November 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/501.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_28/Issue_
05/tiff/502.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/503.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/504.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/505.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/506.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/507.tif.

Gearing:1958:SFB

[Gea58] H. W. Gearing. Statistical foundations for business forecasts. The Computer Journal, 1(2):59–63, July 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
02/010059.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
02/tiff/59.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_02/tiff/
60.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_01/Issue_02/tiff/61.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
01/Issue_02/tiff/62.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
02/tiff/63.tif.

Gearing:1959:SST

[Gea59] H. W. Gearing. Symposium on the selection and training of programmers — 1: a business user’s approach. The Computer Journal, 2(3):107–109, October 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/020107.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/107.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/108.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/109.tif.

Gearing:1961:UPA

[Gea61]  H. W. Gearing.  The use

of Pegasus Autocode in some experimental business applications of computers. The Computer Journal, 4(1):30–34, April 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/040030.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/tiff/30.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_01/tiff/
31.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_04/Issue_01/tiff/32.
tif;   http://www3.oup.co.

                                        uk/computer_journal/hdb/Volume_
                                        04/Issue_01/tiff/33.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        01/tiff/34.tif.

Gear:1964:OAF

[Gea64] C. W. Gear. Optimization of the address field compilation in the ILLIAC 2 assembler. The Computer Journal, 6(4):332–335, January 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        04/tiff/332.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        04/tiff/333.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        04/tiff/334.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        04/tiff/335.tif.

Gearing:1973:ESB

[Gea73]  H. W. G. Gearing.  Ed-

itorial: the silent business user. The Computer Journal, 16(3):194–195, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        03/160194.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        03/tiff/194.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_

03/tiff/195.tif. See corre-

spondence [Fis74b].

Gearing:1980:NAS

[Gea80] H. W. G. Gearing. New approaches to systems analysis design. welcome. The Computer Journal, 23(1):3–5, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
01/tiff/3.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/
4.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_23/Issue_01/tiff/5.
tif.

Gedeon:1996:CTH

[Ged96] T. D. Gedeon. The Cyclic Towers of Hanoi: an iterative solution produced by transformation. The Computer Journal, 39(4):353–356, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www.oup.co.uk/jnls/
list/comjnl/hdb/Volume_39/
Issue_04/390353.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_04/
Vol39_04.body.html#AbstractGedeon.

Geesink:1991:CDA

[Gee91] L. H. Geesink. The coordination of distributed active messages in a dynamic network topology. The Computer Journal, 34(6):542–550, December 1991. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/542.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/543.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/544.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/545.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/546.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/547.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/548.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/549.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/550.tif.

Gehani:1983:HLF

[Geh83] N. H. Gehani. High level form definition in office information systems. The Computer Journal, 26(1):52–59, February 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
01/tiff/52.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_01/tiff/
53.tif;   http://www3.oup.

co.uk/computer_journal/hdb/
Volume_26/Issue_01/tiff/54.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
26/Issue_01/tiff/55.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
01/tiff/56.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_01/tiff/
57.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_26/Issue_01/tiff/58.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
26/Issue_01/tiff/59.tif.

Gentleman:1969:EAG

[Gen69] W. M. Gentleman. An error analysis of Goertzel’s (watt’s) method for computing Fourier coefficients. The Computer Journal, 12(2):160–164, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/120160.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/160.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/161.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/162.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/163.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/164.tif.

George:1963:SAJ

[Geo63] A. F. George. SAS aids for the jet age: telecommunication for electronic reservations. The Computer Journal, 6(1):14–17, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        01/060014.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        01/tiff/14.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_06/Issue_01/tiff/
                                        15.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_06/Issue_01/tiff/16.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        06/Issue_01/tiff/17.tif.

George:1966:CAI

[Geo66] F. H. George. Computer assisted instruction. The Computer Journal, 9(1):32–34, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        01/090032.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        01/tiff/32.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_09/Issue_01/tiff/
                                        33.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_09/Issue_01/tiff/34.
                                        tif.

Gertler:1970:HLP

[Ger70] J. Gertler. High-level programming for process control. The Computer Journal, 13(1):70–75, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/130070.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/tiff/70.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_01/tiff/
71.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/72.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
13/Issue_01/tiff/73.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/tiff/74.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_01/tiff/
75.tif.

Gwei:1987:FSI

[GF87] G. M. Gwei and E. Foxley. A flexible synonym interface with application examples in CAL and help environments. The Computer Journal, 30(6):551–557, December 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/551.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_30/Issue_
                                        06/tiff/552.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        06/tiff/553.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        06/tiff/554.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        06/tiff/555.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        06/tiff/556.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        06/tiff/557.tif.

Ghafoor:1991:DCC

[GF91] A. Ghafoor and F. Y. Farhat. Dynamic concurrency control algorithms for large distributed database systems. The Computer Journal, 34(6):A095– A104, December 1991. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Gregory:1959:ECT

[GG59] R. H. Gregory and H. W. Gearing. Electronic computers as tools for management in the United States of America: 1956. The Computer Journal, 1(4):179–191, January 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        04/010179.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        04/tiff/179.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/tiff/180.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/tiff/181.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/tiff/182.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/tiff/183.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/tiff/184.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/tiff/185.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/tiff/186.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/tiff/187.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/tiff/188.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/tiff/189.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/tiff/190.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/tiff/191.tif.

Georgeff:1986:RNU

[GG86] M. P. Georgeff and G. K.

Gupta. Resources needed for a university computer science department. The Computer Journal, 29(1):90–95, February 1986. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        01/tiff/90.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_29/Issue_01/tiff/
                                        91.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_29/Issue_01/tiff/92.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        29/Issue_01/tiff/93.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        01/tiff/94.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_29/Issue_01/tiff/
                                        95.tif.

Godfrey:1987:PCS

[GGH^{+}87] D. Godfrey, S. Gong, R. Hart, N. Koorland, and S. Smit. Project ‘Clear Skies’: teaching computer science by computer based training and electronic messaging in China. The Computer Journal, 30 (5):469–474, October 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        05/tiff/469.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        05/tiff/470.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        05/tiff/471.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        05/tiff/472.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/473.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/474.tif.

Godwin:1989:CJD

[GGS89] A. N. Godwin, M. B. Gore, and D. W. Salt. A comparison of JSD and DFD as descriptive tools. The Computer Journal, 32(3):202–211, June 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/202.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/203.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/204.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/205.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/206.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/207.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/208.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/209.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/210.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_32/Issue_
                                        03/tiff/211.tif.

Grant:1973:SPE

[GH73] J. A. Grant and G. D. Hitchins. The solution of polynomial equations in interval arithmetic. The Computer Journal, 16(1):69–72, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        01/160069.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        01/tiff/69.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_16/Issue_01/tiff/
                                        70.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_16/Issue_01/tiff/71.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        16/Issue_01/tiff/72.tif.

Grant:1975:TAS

[GH75] J. A. Grant and G. D. Hitchins. Two algorithms for the solution of polynomial equations to limiting machine precision. The Computer Journal, 18(3):258–264, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/258.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/259.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_18/Issue_
03/tiff/260.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/261.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/262.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/263.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/264.tif.

Ganley:1994:ORP

[GH94] Joseph L. Ganley and Lenwood S.

Heath. Optimal and random partitions of random graphs. The Computer Journal, 37(7):641–643, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_37/Issue_07/
Vol37_07.body.html#AbstractGanley.

Gilmore:1995:PAT

[GH95] Stephen Gilmore and Jane Hillston. Process algebras and their application to performance modelling. The Computer Journal, 38(7):489–491, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_38/Issue_07/
Vol38_07.body.html#AbstractGilmore.

Proceedings of the Third Workshop on Process Algebra and Performance Modelling, Edinburgh, Scotland.

Ghandour:1972:FSA

[Gha72] Z. J. Ghandour. Formal systems and analysis of context sensitive languages. The Computer Journal, 15(3):229–237, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/150229.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/229.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/230.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/231.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/232.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/233.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/234.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/235.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/236.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/237.tif.

Guder:1993:GPP

[GHR^{+}93] U. Geuder, M. Hardtner, A. Reuter, B. Worner, and R. Zink. GRIDS — a parallel programming system for gridbased algorithms. The Computer Journal, 36(8):702–711, ???? 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_08/
Vol36_08.body.html#AbstractGeuder.

Gibbons:1960:PAI

[Gib60]  A. Gibbons.  A program

for the automatic integration of differential equations using the method of Taylor series. The Computer Journal, 3(2):108–111, July 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
02/030108.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
02/tiff/108.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
02/tiff/109.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
02/tiff/110.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
02/tiff/111.tif.

Gibbons:1961:RPA

[Gib61] A. Gibbons. Running Pegasus autocode programs on Mercury. The Computer Journal, 3(4):232–236, January 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/030232.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/232.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/233.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/234.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/235.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/236.tif.

Gibbons:1968:KA

[Gib68] A. Gibbons. K Autocode. The Computer Journal, 11 (2):160–168, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/110160.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/160.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/161.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/162.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/163.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_

02/tiff/164.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/165.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/166.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/167.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/168.tif.

Gibbins:1988:VAP

[Gib88] P. F. Gibbins. VDM: axiomatising its propositional logic. The Computer Journal, 31(6):510–516, December 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/510.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/511.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/512.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/513.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/514.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/515.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/516.tif.

Gulati:1990:PMF

[GIB90] S. Gulati, S. S. Iyengar, and J. Barhen. The pebblecrunching model for faulttolerant load balancing in hypercube ensembles. The Computer Journal, 33(3):204–214, June 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/204.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/205.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/206.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/207.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/208.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/209.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/210.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/211.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/212.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/213.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/214.tif.

Gill:1958:BFH

[Gil58a] S. Gill. A binary form of Horner’s method. The Computer Journal, 1(2):84–86, July 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
02/010084.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
02/tiff/84.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_02/tiff/
85.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_01/Issue_02/tiff/86.
tif.

Gill:1958:PP

[Gil58b]  S. Gill.  Parallel program-

ming. The Computer Journal, 1(1):2–10, April 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
01/010002.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
01/tiff/10.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_01/tiff/
2.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_01/Issue_01/tiff/3.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
01/Issue_01/tiff/4.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_

                                        01/tiff/5.tif;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_01/Issue_01/tiff/
                                        6.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_01/Issue_01/tiff/7.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        01/Issue_01/tiff/8.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        01/tiff/9.tif.

Gilmour:1958:SRPa

[Gil58c] A. Gilmour. The solution of railway problems on a digital computer: 1. The Computer Journal, 1(1):25–28, April 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        01/010025.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        01/tiff/25.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_01/Issue_01/tiff/
                                        26.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_01/Issue_01/tiff/27.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        01/Issue_01/tiff/28.tif.

Gilmour:1958:SRPb

[Gil58d] A. Gilmour. The solution of railway problems on a digital computer: 2. The Computer Journal, 1(2):78–83, July 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
02/010078.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
02/tiff/78.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_02/tiff/
79.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_01/Issue_02/tiff/80.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
01/Issue_02/tiff/81.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
02/tiff/82.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_02/tiff/
83.tif.

Gill:1959:CTP

[Gil59] S. Gill. Current theory and practice of automatic programming. The Computer Journal, 2(3):110–114, October 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/020110.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/110.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/111.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/112.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/113.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_02/Issue_
                                        03/tiff/114.tif.

Gill:1963:TCM

[Gil63] S. Gill. A technique for the composition of music in a computer. The Computer Journal, 6(2):129–133, July 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/060129.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/tiff/129.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/tiff/130.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/tiff/131.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/tiff/132.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/tiff/133.tif.

Giles:1964:PLA

[Gil64] P. Giles. Problems of life assurance data processing. The Computer Journal, 7(2):95–98, July 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        02/070095.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        02/tiff/95.tif;   http://

www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_02/tiff/
96.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_07/Issue_02/tiff/97.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
07/Issue_02/tiff/98.tif.

Gill:1965:ACP

[Gil65]  S. Gill.  Automatic com-

puting: its problems and prizes. The Computer Journal, 8(3):177–189, October 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/080177.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/177.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/178.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/179.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/180.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/181.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/182.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/183.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/184.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        03/tiff/185.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        03/tiff/186.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        03/tiff/187.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        03/tiff/188.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        03/tiff/189.tif.

Giles:1966:SCR

[Gil66]  P. Giles.  Some considera-

tions regarding arrangement of records on magnetic tape. The Computer Journal, 8 (4):319–322, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        04/080319.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        04/tiff/319.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        04/tiff/320.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        04/tiff/321.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        04/tiff/322.tif.

Giles:1969:MC

[Gil69]  P. Giles.  Mini-COBOL.

The Computer Journal, 12 (3):208–214, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/120208.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/208.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/209.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/210.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/211.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/212.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/213.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/214.tif.

Giles:1973:IPW

[Gil73] P. Giles. Implementation problems — why are they seldom aired in the journal? The Computer Journal, 16(3):283, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/283.tif.

Garside:1968:NMS

[GJM68] G. R. Garside, P. Jarratt, and C. Mack. A new method for solving polynomial equations. The Computer Journal, 11(1):87–90, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/110087.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/87.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_01/tiff/
88.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_11/Issue_01/tiff/89.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/90.tif. See

correspondence [O’B68b].

Green:1973:NTF

[GK73] D. H. Green and R. G. Kelsch. Nonlinear ternary feedback shift registers. The Computer Journal, 16(4):360–367, November 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/160360.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/360.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/361.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/362.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_16/Issue_
04/tiff/363.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/364.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/365.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/366.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/367.tif.

Gale:1984:PIC

[GK84] William A. Gale and Roger Koenker. Pricing interactive computer services. The Computer Journal, 27(1):8– 17, February 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
27/Issue_01/tiff/10.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
01/tiff/11.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_01/tiff/
12.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_27/Issue_01/tiff/13.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
27/Issue_01/tiff/14.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
01/tiff/15.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_01/tiff/
16.tif;   http://www3.oup.
co.uk/computer_journal/hdb/

                                        Volume_27/Issue_01/tiff/17.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        27/Issue_01/tiff/8.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        01/tiff/9.tif.

Geschiere:1996:PDP

[GK96] J. P. Geschiere and W. H. F. J. Körver. A practical development process for parallel largescale applications and its underlying formal framework. The Computer Journal, 39(5):386– 407, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://www.oup.co.uk/jnls/
                                        list/comjnl/hdb/Volume_39/
                                        Issue_05/390386.sgm.abs.html;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_39/Issue_05/
                                        Vol39_05.body.html#AbstractGeschiere.

Gudes:1979:SMS

[GKS79] E. Gudes, H. S. Koch, and F. A. Stahl. Security in a multi-level structured model of a data base. The Computer Journal, 22(4):303–306, November 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        04/tiff/303.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        04/tiff/304.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        04/tiff/305.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_22/Issue_
04/tiff/306.tif.

Grant:1964:NIP

[GL64] J. A. Grant and O. G. Ludwig. Note on the integrals of products of associated Legendre functions. The Computer Journal, 6(4):356–357, January 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/060356.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/356.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/357.tif.

Garwick:1970:CEI

[GL70] Jan V. Garwick and J. N. Lyness. Correspondence: On “The effect of inadequate convergence criteria in automatic routines”. The Computer Journal, 13(1):122, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/tiff/122.tif. See [Lyn69].

Gunther:1994:OOT

[GL94] Oliver Günther and Johannes Lamberts. Object-oriented techniques for the management of geographic and environmental data. The Computer Journal, 37(1):16–25, ????

1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_37/Issue_01/
Vol37_01.body.html#AbstractGunther.

Gladwin:1969:UGD

[Gla69] B. J. Gladwin. The utilisation of graphic display units as the main form of computer input. The Computer Journal, 12(2):114–117, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/120114.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/114.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/115.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/116.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/117.tif.

Glading:1973:CDL

[Gla73] D. G. Glading. Correspondence: On Dakin’s letter. The Computer Journal, 16(3):287, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/287.tif. See [Dak73].

Glennie:1962:OEF

[Gle62]  A. E. Glennie.  Operat-

ing experience with FOR-TRAN. The Computer Journal, 5(2):132–139, July 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/050132.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/132.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/133.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/134.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/135.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/136.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/137.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/138.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/139.tif.

Gourlay:1968:DAL

[GlM68] A. R. Gourlay and J. L. l. Morris. Deferred approach to the limit in non-linear hyperbolic systems. The Computer Journal, 11(1):95–101, May 1968. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        01/110095.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        01/tiff/100.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        01/tiff/101.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        01/tiff/95.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_11/Issue_01/tiff/
                                        96.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_11/Issue_01/tiff/97.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        11/Issue_01/tiff/98.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        01/tiff/99.tif.

Gonzalez-Lopez:1993:PTM

[GLR93] M. J. González-López and T. Recio. Path tracking in motion planning. The Computer Journal, 36(5):515–524, October 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_36/Issue_05/
                                        Vol36_05.body.html#AbstractLopez.

Green:1965:SSC

[GM65] D. J. Green and S. Michaelson. Series solution of certain Sturm–Liouville eigenvalue problems. The Computer Journal, 7(4):322–336, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/070322.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/322.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/323.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/324.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/325.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/326.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/327.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/328.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/329.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/330.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/331.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/332.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/333.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_07/Issue_
                                        04/tiff/334.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        04/tiff/335.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        04/tiff/336.tif.

Gill:1972:AIU

[GM72] P. E. Gill and G. F. Miller. An algorithm for the integration of unequally spaced data. The Computer Journal, 15(1):80–83, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        01/150080.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        01/tiff/80.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_15/Issue_01/tiff/
                                        81.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_15/Issue_01/tiff/82.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        15/Issue_01/tiff/83.tif.

Ghosh:1983:PCG

[GM83] P. K. Ghosh and S. P. Mudur. Parametric curves for graphic design systems. The Computer Journal, 26(4):312–319, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        04/tiff/312.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/313.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/314.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/315.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/316.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/317.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/318.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/319.tif.

Gunter:1995:SMM

[GM95] Elsa Gunter and Savi Maharaj. Studying the ML module system in HOL. The Computer Journal, 38(2):142–151, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_38/Issue_02/
Vol38_02.body.html#AbstractGunter.

Griswold:1983:PCO

[GN83] R. E. Griswold and Michael Novak. Programmer-defined control operations. The Computer Journal, 26(2):175–183, May 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_

02/tiff/175.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/176.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/177.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/178.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/179.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/180.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/181.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/182.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/183.tif.

Goyal:1988:DMI

[GN88] P. Goyal and T. S. Narayanan. Dictionary machine with improved performance. The Computer Journal, 31(6):490– 495, December 1988. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
31/Issue_06/tiff/490.tif;
 http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/491.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/492.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_31/Issue_
06/tiff/493.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/494.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/495.tif.

Giddy:1994:ACT

[GN94] J. P. Giddy and R. Safavi Naini. Automated cryptanalysis of transposition ciphers. The Computer Journal, 37(5):429–436, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_37/Issue_05/
Vol37_05.body.html#AbstractGiddy.

Goble:1975:FRS

[Gob75]  C. E. Goble.  A free-text

retrieval system using hash codes. The Computer Journal, 18(1):18–20, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
01/tiff/18.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_01/tiff/
19.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_18/Issue_01/tiff/20.
tif.

Goddu:1975:APO

[God75] F. J. B. Goddu. The application of precedence operations to interactive Canadian income tax calculations. The Computer Journal, 18(4):344–348, November 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/344.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/345.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/346.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/347.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/348.tif.

Godwin:1996:BSD

[God96] Anthony N. Godwin. Building a semantic description for a data modelling language. The Computer Journal, 39(4):275– 290, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www.oup.co.uk/jnls/
list/comjnl/hdb/Volume_39/
Issue_04/390275.sgm.abs.html;
 http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_04/
Vol39_04.body.html#AbstractGodwin.

Goldsmith:1959:SAC

[Gol59] J. A. Goldsmith. The state of the art — (a) Commercial computers in Britain, June 1959. The Computer Journal, 2(3):97–99, October 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/020097.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/97.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_03/tiff/
98.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_02/Issue_03/tiff/99.
tif.

Golds:1964:UEC

[Gol64] I. M. Golds. User’s experience of COBOL. The Computer Journal, 7(2):99–101, July 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/070099.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/100.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/101.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/99.tif.

Goldman:1969:CAT

[Gol69] M. S. Goldman. Computeraided typesetting with a CRT. The Computer Journal, 12 (4):357–364, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_

                                        04/120357.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/357.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/358.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/359.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/360.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/361.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/362.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/363.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/364.tif.

Goldsworthy:1973:CPA

[Gol73] A. W. Goldsworthy. Computers and privacy — A review of the Younger Comm. Report. The Computer Journal, 5(1):??, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Golden:1978:CAS

[Gol78a] Bruce L. Golden. Correspondence: On “An algorithm for the solution of the postman’s problem”. The Computer Journal, 21(4):383, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/383.tif. See [Bar78a].

Goldsack:1978:CSD

[Gol78b] S. J. Goldsack. Correspondence: Semantic definition of linear control algorithms. The Computer Journal, 21 (4):381–382, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/381.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/382.tif.

Goller:1982:TLU

[Gol82] N. E. Goller. Triodic logic and its use in structured program design. The Computer Journal, 25(2):218–226, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/218.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/219.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/220.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/221.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/222.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_25/Issue_
02/tiff/223.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/224.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/225.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/226.tif.

Goller:1985:HDS

[Gol85] N. E. Goller. Hybrid data structures defined by indirection. The Computer Journal, 28(1):44–53, February 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
01/tiff/44.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_01/tiff/
45.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_28/Issue_01/tiff/46.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
28/Issue_01/tiff/47.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
01/tiff/48.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_01/tiff/
49.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_28/Issue_01/tiff/50.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
28/Issue_01/tiff/51.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_

01/tiff/52.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_01/tiff/
53.tif.

Goldberg:1993:FSA

[Gol93] Robert R. Goldberg. Finite state automata from regular expression trees. The Computer Journal, 36(7):623–630, ???? 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_07/
Vol36_07.body.html#AbstractGoldberg.

Goldson:1994:SCN

[Gol94] Doug Goldson. A symbolic calculator for non-strict functional programs. The Computer Journal, 37(3):177–187, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_37/Issue_03/
Vol37_03.body.html#AbstractGoldson.

Gomaa:1981:HSM

[Gom81] H. Gomaa. Hybrid simulation/ regression model of a virtual storage system. The Computer Journal, 24(4):358–366, November 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/358.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/359.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/360.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/361.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/362.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/363.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/364.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/365.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/366.tif.

Good:1961:CSM

[Goo61] I. J. Good. A comparison of some methods of calculating covariance functions on an electronic computer. The Computer Journal, 3(4):262–265, January 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/030262.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/262.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/263.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/264.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_03/Issue_
04/tiff/265.tif.

Goodman:1964:AMT

[Goo64] J. V. Goodman. Auditing magnetic-tape systems. The Computer Journal, 7(1):4–7, April 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
01/070004.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
01/tiff/4.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_07/Issue_01/tiff/
5.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_07/Issue_01/tiff/6.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
07/Issue_01/tiff/7.tif.

Good:1968:CGT

[Goo68] I. J. Good. Correspondence:

Gödel’s theorem. The Computer Journal, 10(4):417, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/417.tif.

Goodwin:1975:AIC

[Goo75]  D. T. Goodwin.  An al-

gorithm for inverting certain translators of context-free languages. The Computer Journal, 18(4):349–354, November 1975. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/349.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/350.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/351.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/352.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/353.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/354.tif.

Goodwin:1977:CUO

[Goo77] D. T. Goodwin. Conditions for underflow and overflow of an arithmetic stack. The Computer Journal, 20(1):56–62, February 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
01/tiff/56.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_01/tiff/
57.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_20/Issue_01/tiff/58.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
20/Issue_01/tiff/59.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
01/tiff/60.tif;   http://
www3.oup.co.uk/computer_journal/

hdb/Volume_20/Issue_01/tiff/
61.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_20/Issue_01/tiff/62.
tif.

Goodwin:1980:PNN

[Goo80] D. T. Goodwin. Partial nonunderflow and non-overflow of an arithmetic stack. The Computer Journal, 23(2):153–160, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/153.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/154.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/155.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/156.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/157.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/158.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/159.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/160.tif.

Goodwin:1985:CSN

[Goo85a] D. T. Goodwin. The computability of stack non-underflow. The Computer Journal, 28(3): 257–263, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/257.tif;
 http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/258.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/259.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/260.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/261.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/262.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/263.tif.

Goodwin:1985:FSS

[Goo85b] D. T. Goodwin. Further study of a stack-length model. The Computer Journal, 28(1):9– 16, February 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
28/Issue_01/tiff/10.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
01/tiff/11.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_01/tiff/
12.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_28/Issue_01/tiff/13.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_

28/Issue_01/tiff/14.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
01/tiff/15.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_01/tiff/
16.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_28/Issue_01/tiff/9.
tif.

Gordon:1978:IGM

[Gor78]  M. Gordon.  An interac-

tive graphical minicomputer system for the management and exploration of clinical renal data. The Computer Journal, 21(1):91–93, February 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
01/tiff/91.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_01/tiff/
92.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_21/Issue_01/tiff/93.
tif.

Gordon:1979:PLI

[Gor79] M. J. C. Gordon. On the power of list iteration. The Computer Journal, 22(4):376–379, November 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/376.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/377.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        04/tiff/378.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        04/tiff/379.tif.

Gosden:1960:MRA

[Gos60] J. A. Gosden. Market research applications on LEO. The Computer Journal, 3 (3):142–143, October 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        03/030142.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        03/tiff/142.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        03/tiff/143.tif.

Gosden:1963:ECP

[Gos63]  J. A. Gosden.  Estimat-

ing computer performance. The Computer Journal, 5 (4):276–283, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        04/050276.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        04/tiff/276.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        04/tiff/277.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_

04/tiff/278.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/279.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/280.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/281.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/282.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/283.tif.

Gourlay:1968:APR

[Gou68]  A. R. Gourlay.  The ac-

celeration of the Peaceman– Rachford method by Chebyshev polynomials. The Computer Journal, 10(4):378–382, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/100378.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/378.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/379.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/380.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/381.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/382.tif.

Gourlay:1970:GEH

[Gou70] A. R. Gourlay. Generalisation of elementary Hermitian matrices. The Computer Journal, 13(4):411–412, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/130411.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/411.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/412.tif.

Gower:1958:NIM

[Gow58] J. C. Gower. A note on an iterative method for root extraction. The Computer Journal, 1(3):142–143, October 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        03/010142.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        03/tiff/142.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        03/tiff/143.tif.

Gower:1962:HMT

[Gow62]  J. C. Gower.  The han-

dling of multiway tables on computers. The Computer Journal, 4(4):280–286, January 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/040280.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/280.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/281.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/282.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/283.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/284.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/285.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/286.tif.

Gonnet:1983:HT

[GOW83] Gaston H. Gonnet, Henk J.

Olivié, and Derick Wood. Height-ratio-balanced trees. The Computer Journal, 26(2): 106–108, May 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
26/Issue_02/tiff/106.tif;
 http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/107.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/108.tif.

Gay:1973:IDS

[GP73] B. Gay and S. G. Payne. Interactive digital simulation on a small computer. The Computer Journal, 16(2):118–121, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/160118.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/118.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/119.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/120.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/121.tif.

Garside:1980:API

[GP80] G. R. Garside and P. E. Pintelas. An ALGOL 68 package for implementing graph algorithms. The Computer Journal, 23(3):237–242, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        03/tiff/237.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        03/tiff/238.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        03/tiff/239.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_23/Issue_
03/tiff/240.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/241.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/242.tif.

Graybeal:1981:DCE

[GP81] W. T. Graybeal and U. W.

Pooch. Design of calibration experiments for synthetic jobs. The Computer Journal, 24(2):132–138, May 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/132.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/133.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/134.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/135.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/136.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/137.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/138.tif.

Gopalakrishnan:1988:MAC

[GP88] M. Gopalakrishnan and L. M.

Patnaik. Medium access control schemes for local area networks with multiple priority function. The Computer Journal, 31(3):209–219, June 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/209.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/210.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/211.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/212.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/213.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/214.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/215.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/216.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/217.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/218.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/219.tif.

Govindarajan:1990:LEC

[GP90] R. Govindarajan and L. M. Patnaik. Lenient execution and concurrent execution of reentrant routines: Efficient implementation in data flow systems. The Computer Journal, 33(2):185–187, April 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/185.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/186.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/187.tif.

Gu:1996:EAN

[GP96a] Qian-Ping Gu and Shietung Peng. An efficient algorithm for node-to-node routing in hypercubes with faulty clusters. The Computer Journal, 39(1): 14–19, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www.oup.co.uk/jnls/
list/comjnl/hdb/Volume_39/
Issue_01/390014.sgm.abs.html;
 http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_01/
Vol39_01.body.html#AbstractGu.

Gu:1996:OAN

[GP96b] Qian-Ping Gu and Shietung Peng. Optimal algorithms for node-to-node fault tolerant routing in hypercubes. The Computer Journal, 39(7):626– 629, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www.oup.co.uk/jnls/
list/comjnl/hdb/Volume_39/

                                        Issue_07/390626.sgm.abs.html;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_39/Issue_07/
                                        Vol39_07.body.html#AbstractGu.

Ghoshdastidar:1975:SES

[GR75] D. Ghoshdastidar and M. K. Roy. A study on the evaluation of Shell’s sorting technique. The Computer Journal, 18(3):234–235, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/234.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/235.tif.

Grady:1967:NG

[Gra67]  J. W. Grady.  The Na-

tional Giro. The Computer Journal, 10(2):121–125, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        02/100121.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        02/tiff/121.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        02/tiff/122.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        02/tiff/123.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        02/tiff/124.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_10/Issue_
02/tiff/125.tif.

Grad:1971:MB

[Gra71]  J. Grad.  Matrix balanc-

ing. The Computer Journal, 14(3):280–284, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/140280.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/280.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/281.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/282.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/283.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/284.tif.

Gray:1987:PVC

[Gra87]  D. Gray.  A pedagogical

verification condition generator. The Computer Journal, 30(3):239–248, June 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/239.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/240.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_30/Issue_
03/tiff/241.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/242.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/243.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/244.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/245.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/246.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/247.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/248.tif.

Grad:1968:ACE

[GRB68] J. Grad, K. A. Redish, and M. A. Brebner. Algorithm 32: Calculation of eigenvalues of real matrices by the QR method using double QR step. The Computer Journal, 11(1):112–114, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/112.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/113.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/114.tif.

Goldson:1993:RSP

[GRB93] Doug Goldson, Steve Reeves, and Richard Bornat. A review of several programs for the teaching of logic. The Computer Journal, 36(4):373–386, August 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_04/
Vol36_04.body.html#AbstractGoldson.

Greenspan:1967:NAB

[Gre67]  D. Greenspan.  A numeri-

cal approach to biharmonic problems. The Computer Journal, 10(2):198–201, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/100198.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/198.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/199.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/200.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/201.tif.

Greenspan:1969:NSP

[Gre69] D. Greenspan. Numerical studies of prototype cavity flow problems. The Computer Journal, 12(1):88–93, February 1969. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
01/120088.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
01/tiff/88.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_01/tiff/
89.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_12/Issue_01/tiff/90.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/91.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
01/tiff/92.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_01/tiff/
93.tif.

Greenspan:1970:DNS

[Gre70] D. Greenspan. Discrete, nonlinear string vibrations. The Computer Journal, 13(2):195–201, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/130195.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/195.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/196.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/197.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_

02/tiff/198.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/199.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/200.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/201.tif.

Griffiths:1959:SEU

[Gri59] G. B. Griffiths. Symposium on experiences with the use of magnetic tape — 1: magnetic tapes on a Ferranti Pegasus. The Computer Journal, 2(3):118–119, October 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/020118.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/118.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/119.tif.

Grimmond:1962:ARS

[Gri62] R. Grimmond. An analysis of real and simulated statistics for system design purposes. The Computer Journal, 5(2):94–99, July 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/050094.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_

02/tiff/94.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_02/tiff/
95.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_05/Issue_02/tiff/96.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
05/Issue_02/tiff/97.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/98.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_02/tiff/
99.tif.

Grindley:1966:SNP

[Gri66] C. B. B. Grindley. Systematics — a non-programming language for designing and specifying commercial systems for computers. The Computer Journal, 9(2):124–128, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/090124.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/124.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/125.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/126.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/127.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_

02/tiff/128.tif. See com-

ments [BBdD66].

Grindley:1968:UDT

[Gri68]  C. B. B. Grindley.  The

use of decision tables within systematics. The Computer Journal, 11(2):128–133, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/110128.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/128.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/129.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/130.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/131.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/132.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/133.tif.

Griffiths:1973:DOS

[Gri73]  J. G. Griffiths.  Drawing

opaque solids using an incremental plotter. The Computer Journal, 16(1):35–38, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
01/160035.sgm.abs.html;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
01/tiff/35.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_01/tiff/
36.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_16/Issue_01/tiff/37.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
16/Issue_01/tiff/38.tif.

Griswold:1980:UCS

[Gri80] R. E. Griswold. The use of character sets and character mappings in Icon. The Computer Journal, 23(2):107–114, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/107.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/108.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/109.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/110.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/111.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/112.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/113.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/114.tif.

Griswold:1988:PG

[Gri88] R. E. Griswold. Programming with generators. The Computer Journal, 31(3):220–228, June 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/220.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/221.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/222.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/223.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/224.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/225.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/226.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/227.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/228.tif.

Griswold:1990:SSI

[Gri90] R. E. Griswold. String scanning in the Icon programming language. The Computer Journal, 33(2):98–107, April 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/100.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/101.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/102.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/103.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/104.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/105.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/106.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/107.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/98.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_02/tiff/
99.tif.

Gries:1992:CFM

[Gri92] David Gries. Correspondence: On “Finding a Majority When Sorting is Not Available”. The Computer Journal, 35(3):298, June 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_35/Issue_
03/tiff/298.tif. See [CM91].

Points out that the supposedly new algorithm was previously published by Boyer and Moore (1982), Misra and Gries (1982), and Gries (1990).

Groves:1968:MEC

[Gro68] P. D. Groves. Marking and evaluating class tests and examinations by computer. The Computer Journal, 10 (4):365–367, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/100365.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/365.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/366.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/367.tif.

Grossman:1971:PCF

[Gro71]  M. Grossman.  Parametric

curve fitting. The Computer Journal, 14(2):169–172, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/140169.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/169.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/170.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_

                                        02/tiff/171.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/172.tif.

Grover:1973:MOI

[Gro73]  D. J. Grover.  Message

orientated interactive graphics. The Computer Journal, 16(1):39–45, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        01/160039.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        01/tiff/39.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_16/Issue_01/tiff/
                                        40.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_16/Issue_01/tiff/41.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        16/Issue_01/tiff/42.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        01/tiff/43.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_16/Issue_01/tiff/
                                        44.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_16/Issue_01/tiff/45.
                                        tif.

Grundy:1996:THR

[Gru96] Jim Grundy. Transformational hierarchical reasoning. The Computer Journal, 39(4):291– 302, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www.oup.co.uk/jnls/
list/comjnl/hdb/Volume_39/
Issue_04/390291.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_04/
Vol39_04.body.html#AbstractGrundy.

Gerard:1963:HRA

[GS63] J. M. Gerard and A. Sambles. A hardware representation for AL-GOL 60 using Creed teleprinter equipment. The Computer Journal, 5(4):338–340, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/050338.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/338.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/339.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/340.tif.

Girling:1973:BRC

[GS73] B. Girling and D. A. Sturt.

Book reviews: Computer Simulation of Dynamic Systems, by R. J. Kochenburger, 1972; Electronic Data Processing and Computers for Commercial Students, by E. A. Bird, 1972. The Computer Journal, 16(3):215, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_

journal/hdb/Volume_16/Issue_
03/tiff/215.tif.

Green:1978:CDT

[GS78] P. J. Green and R. Sibson. Computing Dirichlet tessellations in the plane. The Computer Journal, 21(2):168–173, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/168.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/169.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/170.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/171.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/172.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/173.tif.

Green:1979:CCH

[GS79] P. J. Green and B. W. Silverman. Constructing the convex hull of a set of points in the plane. The Computer Journal, 22(3):262–266, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/262.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_

03/tiff/263.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/264.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/265.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/266.tif.

Goyer:1963:BRR

[GT63] M. Goyer and G. C. Tootill.

Boor review: Russian Supplement to Elsevier’s Dictionary of Automation, Computers, Control and Measuring, by W. E. Clason, 1962. The Computer Journal, 6(3):273, October 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/273.tif.

Gould:1965:TWI

[GT65] I. H. Gould and G. C. Tootill.

The terminology work of IFIP and ICC. The Computer Journal, 7(4):264–270, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/070264.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/264.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/265.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_07/Issue_
                                        04/tiff/266.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        04/tiff/267.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        04/tiff/268.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        04/tiff/269.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        04/tiff/270.tif.

Gayler:1972:CEP

[GT72] D. R. Gayler and The Editor. Correspondence: On editorial policy. The Computer Journal, 15(4):307, November 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/150307.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/tiff/307.tif. See [Wal71,

MT72, Bar73, KTW^{+}73].

Guin:1968:DCM

[Gui68]  J. A. Guin.  Discussion

and correspondence: modification of the complex method of constrained optimization. The Computer Journal, 10 (4):416–417, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/100416.sgm.abs.html;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/416.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/417.tif.

Gunton:1970:RPD

[Gun70] A. Gunton. Recovery procedures for direct access commercial systems. The Computer Journal, 13(2):123–126, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/130123.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/123.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/124.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/125.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/126.tif.

Goel:1991:OFC

[GV91] V. Goel and Y. V. Venkatesh.

On an optimal and faster construction of linear quadtrees from raster-scanned images. The Computer Journal, 34(6): A073–A083, December 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Gammerman:1999:KCS

[GV99a] Alexander Gammerman and Vladimir Vovk. Kolmogorov complexity: Sources, theory and applications. The Computer Journal, 42(4):252–255, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
04/420252.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
04/pdf/420252.pdf.

Gammerman:1999:SIE

[GV99b] Alexander Gammerman and Vladimir Vovk. Special issue editorial: Kolmogorov complexity. The Computer Journal, 42(4):251, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
04/420251.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
04/pdf/420251.pdf.

Gibbons:1973:CFC

[GW73] Charles J. Gibbons and M. Wells. Correspondence: On “File Compression Using Variable Length Encodings”. The Computer Journal, 16(3):287, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/287.tif. See [Wel72].

Guariso:1988:SBE

[GW88] G. Guariso and H. Werthner.

A software base for environmental studies. The Computer Journal, 31(6):550–552, December 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/550.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/551.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/552.tif.

Gillet:1998:SDM

[GWWB98] Valerie J. Gillet, David J.

Wild, Peter Willett, and John Bradshaw. Similarity and dissimilarity methods for processing chemical structure databases. The Computer Journal, 41(8):548–558, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_41/Issue_08/
Vol41_08.body.html#AbstractGillet.

Grad:1972:ALS

[GZ72] J. Grad and E. Zakrajsek.

LR algorithm with Laguerre shift for symmetric tridiagonal matrices. The Computer Journal, 15(3):268–270, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

                                        journal/hdb/Volume_15/Issue_
                                        03/150268.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/268.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/269.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/270.tif.

HWG:1968:MIS

[H. 68] H. W. G. Management information systems. The Computer Journal, 10(4):418, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/tiff/418.tif.

Hiasat:1999:SCV

[HAAZ99] Ahmad A. Hiasat and Hoda Abdel-Aty-Zohdy. Semi-custom VLSI design and implementation of a new efficient RNS division algorithm. The Computer Journal, 42(3):232–240, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        03/420232.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        03/pdf/420232.pdf.

Hadjidimos:1968:GAD

[Had68] A. Hadjidimos. On a generalised alternating direction implicit method for solving Laplace’s equation. The Computer Journal, 11(3):324–328, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/110324.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/324.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/325.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/326.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/327.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/328.tif.

Hadjidimos:1971:OEA

[Had71] A. Hadjidimos. Optimum extrapolated ADI iterative difference schemes for the solution of Laplace’s equation in three space variables. The Computer Journal, 14(2):179–183, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/140179.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/179.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/180.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/181.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/182.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/183.tif.

Haddon:1990:CLS

[Had90]  B. K. Haddon.  Cycle-

sort: A linear sorting method. The Computer Journal, 33 (4):365–367, August 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        04/tiff/365.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        04/tiff/366.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        04/tiff/367.tif.

Haines:1969:IIP

[Hai69] C. F. Haines. Implicit integration processes with error estimate for the numerical solution of differential equations. The Computer Journal, 12(2):183–187, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        02/120183.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        02/tiff/183.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/184.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/185.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/186.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/187.tif.

Hall:1967:SPC

[Hal67]  G. Hall.  The stability of

predictor-corrector methods. The Computer Journal, 9 (4):410–412, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/090410.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/410.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/411.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/412.tif.

Hall:1968:PCI

[Hal68] P. A. V. Hall. Pattern classification as interpolation in N dimensions. The Computer Journal, 11(3):287–292, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_

                                        03/110287.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        03/tiff/287.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        03/tiff/288.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        03/tiff/289.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        03/tiff/290.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        03/tiff/291.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        03/tiff/292.tif.

Hall:1972:CFC

[Hal72]  D. T. Hall.  Correspon-

dence: On FORTRAN complainants. The Computer Journal, 15(3):273, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/150273.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/273.tif. See [Hil72,

Fin72, Fla72, Pal72].

Hamblin:1962:TPN

[Ham62]  C. L. Hamblin.  Transla-

tion to and from Polish notation. The Computer Journal, 5(3):210–213, October 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/050210.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/210.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/211.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/212.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/213.tif.

Hammersley:1966:NIL

[Ham66a] P. Hammersley. A note on the implementation of LITHP on the I.C.T. 1905. The Computer Journal, 9(2):173–174, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/090173.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/173.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/174.tif.

Hammersley:1966:NFP

[Ham66b] P. Hammersley. Notice: Future publication of Algorithms Supplement. The Computer Journal, 9(1):80, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_09/Issue_
                                        01/tiff/80.tif.

Hamblin:1967:APO

[Ham67a] C. L. Hamblin. An algorithm for polynomial operations. The Computer Journal, 10(2):168–171, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        02/100168.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        02/tiff/168.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        02/tiff/169.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        02/tiff/170.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        02/tiff/171.tif.

Hammersley:1967:ASPa

[Ham67b] P. Hammersley. Algorithms supplement: Previously published algorithms. The Computer Journal, 9(4):414, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        04/tiff/414.tif.

Hammersley:1967:ASPb

[Ham67c] P. Hammersley. Algorithms supplement: Previously published algorithms. The Computer Journal, 10(1):112, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/112.tif.

Hammersley:1967:ASPc

[Ham67d] P. Hammersley. Algorithms supplement: Previously published algorithms. The Computer Journal, 10(2):206, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/206.tif.

Hammersley:1967:ASPd

[Ham67e] P. Hammersley. Algorithms supplement: Previously published algorithms. The Computer Journal, 10(3):306, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/306.tif.

Hammersley:1967:SP

[Ham67f]  P. Hammersley.  Statement

of policy. The Computer Journal, 10(2):209–210, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/209.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/210.tif.

Hammersley:1968:ASPa

[Ham68a] P. Hammersley. Algorithms supplement: Previously published algorithms. The Computer Journal, 10(4):414, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/tiff/414.tif.

Hammersley:1968:ASPb

[Ham68b] P. Hammersley. Algorithms supplement: Previously published algorithms. The Computer Journal, 11(1):112, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        01/tiff/112.tif.

Hammersley:1968:ASPc

[Ham68c] P. Hammersley. Algorithms supplement: Previously published algorithms. The Computer Journal, 11(3):347, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        03/tiff/347.tif.

Hammersley:1969:ASPa

[Ham69a] P. Hammersley. Algorithms supplement: Previously published algorithms. The Computer Journal, 12(2):197, May 1969. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/197.tif.

Hammersley:1969:ASPb

[Ham69b] P. Hammersley. Algorithms supplement: Previously published algorithms. The Computer Journal, 12(3):291, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/291.tif.

Hammersley:1969:ASPc

[Ham69c] P. Hammersley. Algorithms supplement: Previously published algorithms. The Computer Journal, 12(4):405–406, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
12/Issue_04/tiff/405.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/406.tif.

Hammersley:1969:NAP

[Ham69d] P. Hammersley. Note on Algorithm 34: Procedures for the basic arithmetical operations in multi-length working. The Computer Journal, 12(1):102–103, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

journal/hdb/Volume_12/Issue_
01/tiff/102.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
01/tiff/103.tif. See [Boo67e].

Hammersley:1970:ASPa

[Ham70a] P. Hammersley. Algorithms supplement: Previously published algorithms. The Computer Journal, 13(1):110, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/tiff/110.tif.

Hammersley:1970:ASPb

[Ham70b] P. Hammersley. Algorithms supplement: Previously published algorithms. The Computer Journal, 13(2):206–207, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/206.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/207.tif.

Hammersley:1970:ASPc

[Ham70c] P. Hammersley. Algorithms supplement: Previously published algorithms. The Computer Journal, 13(3):323, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/323.tif.

Hammersley:1970:ASPd

[Ham70d] P. Hammersley. Algorithms supplement: Previously published algorithms. The Computer Journal, 13(4):421, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/421.tif.

Hammersley:1971:ASPa

[Ham71a] P. Hammersley. Algorithms supplement: Previously published algorithms. The Computer Journal, 14(1):103, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/tiff/103.tif.

Hammersley:1971:ASPb

[Ham71b] P. Hammersley. Algorithms supplement: Previously published algorithms. The Computer Journal, 14(2):207, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/207.tif.

Hammersley:1973:E

[Ham73]  P. Hammersley.  Editorial.

The Computer Journal, 16(1): 2–3, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
16/Issue_01/tiff/2.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_

01/tiff/3.tif. See correspondence [TT73].

HamiltonKelly:1974:CRC

[Ham74] B. Hamilton Kelly. Correspondence: On radix conversion. The Computer Journal, 17(4):382, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/382.tif. See [Boo74a,

Boo74b].

Hamlet:1977:TPF

[Ham77] R. G. Hamlet. Testing programs with finite sets of data. The Computer Journal, 20(3):232–237, August 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/232.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/233.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/234.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/235.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_

03/tiff/236.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/237.tif.

Hammersley:1980:CSA

[Ham80a]  P. Hammersley.  Corporate

systems analysis — a personal editorial. The Computer Journal, 23(1):2, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
01/tiff/2.tif.

Hammersley:1980:ICT

[Ham80b] P. Hammersley. The impact of changing technology. The Computer Journal, 23(1):7–9, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
01/tiff/7.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/
8.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_23/Issue_01/tiff/9.
tif.

Hammersley:1981:E

[Ham81a]  P. Hammersley.  Editorial.

The Computer Journal, 24(1): 1, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
24/Issue_01/tiff/1.tif.

Hammersley:1981:IMS

[Ham81b] P. Hammersley. The impact of microcomputer systems on commercial data processing. The Computer Journal, 24(1):14–16, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
01/tiff/14.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_01/tiff/
15.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_24/Issue_01/tiff/16.
tif.

Hammersley:1984:E

[Ham84]  P. Hammersley.  Editorial.

The Computer Journal, 27(1): 1–2, February 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
27/Issue_01/tiff/1.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
01/tiff/2.tif.

Hammersley:1991:AET

[Ham91a] A. P. Hammersley. Algorithm 125: Efficient two-dimensional fast Fourier transform subroutines for real-valued or Hermitian data. The Computer Journal, 34(4):362–370, August 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

journal/hdb/Volume_34/Issue_
04/tiff/362.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/363.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/364.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/365.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/366.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/367.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/368.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/369.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/370.tif.

Hammersley:1991:ISD

[Ham91b] P. Hammersley. Information systems design methodologies — book reviews. The Computer Journal, 34(2):182–185, April 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/182.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/183.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/184.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        02/tiff/185.tif.

Hammersley:1993:SIE

[Ham93] Peter Hammersley. Special issue editorial: Computing curricula. The Computer Journal, 36(4):297–??, August 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_36/Issue_04/
                                        Vol36_04.body.html#AbstractHammersley.

Hubert:1998:RSP

[HAM98] Lawrence Hubert, Phipps Arabie, and Jacqueline Meulman. The representation of symmetric proximity data: Dimensions and classifications. The Computer Journal, 41(8):566–577, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_41/Issue_08/
                                        Vol41_08.body.html#AbstractHubert.

Handscomb:1962:CLR

[Han62]  D. C. Handscomb.  Com-

putation of the latent roots of a Hessenberg matrix by Bairstow’s method. The Computer Journal, 5(2):139–141, July 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        02/050139.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_

journal/hdb/Volume_05/Issue_
02/tiff/139.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/140.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/141.tif.

Hanson:1970:AEB

[Han70] Richard J. Hanson. Automatic error bounds for real roots of polynomials having interval coefficients. The Computer Journal, 13(3):284–288, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/130284.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/284.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/285.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/286.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/287.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/288.tif.

Hanson:1978:FS

[Han78]  D. R. Hanson.  Filters in

SL5. The Computer Journal, 21(2):134–143, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/134.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/135.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/136.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/137.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/138.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/139.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/140.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/141.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/142.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/143.tif.

Hansen:1994:ICS

[Han94] Per Brinch Hansen. Interference control in SuperPascal — a block-structured parallel language. The Computer Journal, 37(5):399–406, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_37/Issue_05/
Vol37_05.body.html#AbstractHansen.

Harrison:1959:SHS

[Har59] J. M. Harrison. Some helicopter simulation studies. The Computer Journal, 2(1):10–23, April 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
01/020010.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
01/tiff/10.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_01/tiff/
11.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_02/Issue_01/tiff/12.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
02/Issue_01/tiff/13.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
01/tiff/14.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_01/tiff/
15.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_02/Issue_01/tiff/16.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
02/Issue_01/tiff/17.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
01/tiff/18.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_01/tiff/
19.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_02/Issue_01/tiff/20.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
02/Issue_01/tiff/21.tif;

                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_02/Issue_
                                        01/tiff/22.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_02/Issue_01/tiff/
                                        23.tif.

Harrop:1970:BAR

[Har70] M. Harrop. A basic approach to remote access. The Computer Journal, 13(2):131–135, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/130131.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/tiff/131.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/tiff/132.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/tiff/133.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/tiff/134.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/tiff/135.tif.

Hartley:1976:TPA

[Har76] P. J. Hartley. Tensor product approximations to data defined on rectangular meshes in N-space. The Computer Journal, 19(4):348–352, November 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_

04/tiff/348.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/349.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/350.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/351.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/352.tif.

Harrington:1979:NSI

[Har79]  Steven J. Harrington.  A

new symbolic integration system in REDUCE. The Computer Journal, 22(2):127–131, May 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/127.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/128.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/129.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/130.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/131.tif.

Hart:1980:JBA

[Har80] Deborah Hart. Joint BCS ACM symposium on Research and Development in Information Retrieval . The Computer Journal, 23(2):131, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        02/tiff/131.tif.

Harrington:1981:SEC

[Har81] Steven J. Harrington. Space efficient copying storage recovery. The Computer Journal, 24(4):316–319, November 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        04/tiff/316.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        04/tiff/317.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        04/tiff/318.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        04/tiff/319.tif.

Harrison:1982:ESM

[Har82]  P. G. Harrison.  Efficient

storage management for functional languages. The Computer Journal, 25(2):264–271, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/264.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/265.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_

02/tiff/266.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/267.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/268.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/269.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/270.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/271.tif.

Harland:1984:UTP

[Har84a]  David M. Harland.  User-

defined types in a polymorphic language. The Computer Journal, 27(1):47–56, February 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
01/tiff/47.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_01/tiff/
48.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_27/Issue_01/tiff/49.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
27/Issue_01/tiff/50.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
01/tiff/51.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_01/tiff/
52.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_27/Issue_01/tiff/53.

                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        27/Issue_01/tiff/54.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        01/tiff/55.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_27/Issue_01/tiff/
                                        56.tif.

Harrison:1984:DCT

[Har84b] P. G. Harrison. The distribution of cycle times in tree-like networks of queues. The Computer Journal, 27(1):27–36, February 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        01/tiff/27.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_27/Issue_01/tiff/
                                        28.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_27/Issue_01/tiff/29.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        27/Issue_01/tiff/30.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        01/tiff/31.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_27/Issue_01/tiff/
                                        32.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_27/Issue_01/tiff/33.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        27/Issue_01/tiff/34.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        01/tiff/35.tif;   http://
                                        www3.oup.co.uk/computer_journal/

hdb/Volume_27/Issue_01/tiff/
36.tif.

Harker:1988:UPS

[Har88] Susan Harker. The use of prototyping and simulation in the development of large-scale applications. The Computer Journal, 31(5):420–425, October 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/420.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/421.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/422.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/423.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/424.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/425.tif.

Harrison:1992:HOA

[Har92] P. G. Harrison. A higherorder approach to parallel algorithms. The Computer Journal, 35(6):555–566, December 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_06/
Vol35_06.body.html#AbstractHarrison.

Harrison:1995:BDD

[Har95] John Harrison. Binary decision diagrams as a HOL derived rule. The Computer Journal, 38(2):162–170, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_38/Issue_02/
                                        Vol38_02.body.html#AbstractHarrison.

Haselgrove:1961:SNL

[Has61] C. B. Haselgrove. The solution of non-linear equations and of differential equations with two-point boundary conditions. The Computer Journal, 4(3):255–259, October 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        03/040255.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        03/tiff/255.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        03/tiff/256.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        03/tiff/257.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        03/tiff/258.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        03/tiff/259.tif.

Haskell:1974:SPR

[Has74]  R. Haskell.  Symmetrical

precedence relations on general phrase structure grammars. The Computer Journal, 17(3):234–241, August 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/234.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/235.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/236.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/237.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/238.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/239.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/240.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/241.tif.

Haskell:1975:EIC

[Has75]  R. Haskell.  Efficient im-

plementation of a class of recursively defined functions. The Computer Journal, 18 (1):23–29, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
01/tiff/23.tif;   http://

                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_18/Issue_01/tiff/
                                        24.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_18/Issue_01/tiff/25.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        18/Issue_01/tiff/26.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        01/tiff/27.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_18/Issue_01/tiff/
                                        28.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_18/Issue_01/tiff/29.
                                        tif.

Hatter:1976:PSM

[Hat76] D. J. Hatter. On processing sparse matrices. The Computer Journal, 19(3):273–274, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/273.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/274.tif. See [ET75].

Hatter:1984:AGN

[Hat84] D. J. Hatter. Algorithm 116: A generalized non-recursive binary tree traversal algorithm. The Computer Journal, 27(2):178–179, May 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        02/tiff/178.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/179.tif.

Haverkort:1995:SPM

[Hav95] Boudewijn R. Haverkort. In search of probability mass: probabilistic evaluation of highlevel specified Markov models. The Computer Journal, 38(7):521–529, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_38/Issue_07/
Vol38_07.body.html#AbstractHaverkort.

Hayes:1969:MSO

[Hay69] D. G. Hayes. A method of storing the orthogonal polynomials used for curve and surface fitting. The Computer Journal, 12(2):148–150, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/120148.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/148.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/149.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/150.tif.

Hayes:1977:NTH

[Hay77]  P. J. Hayes.  A note on

the Towers of Hanoi problem. The Computer Journal, 20(3):282–285, August 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/282.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/283.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/284.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/285.tif. See corre-

spondence [Wod78].

Hayden:1979:CND

[Hay79]  Jane Hayden.  Correspon-

dence: Nonmorphological data collection. The Computer Journal, 22(3):286–287, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/286.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/287.tif. See [Tur78].

Hazlewood:1978:AAC

[Haz78] L. Hazlewood. Algorithm 101: An algorithm for cautious adaptive quadrature. The Computer Journal, 21(2):180–183, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

journal/hdb/Volume_21/Issue_
02/tiff/180.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/181.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/182.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/183.tif.

Hollingdale:1958:ACWa

[HB58a] S. H. Hollingdale and M. M.

Barritt. An application of a computer to wind tunnel design: 1. The Computer Journal, 1(1):42–47, April 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
01/010042.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
01/tiff/42.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_01/tiff/
43.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_01/Issue_01/tiff/44.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
01/Issue_01/tiff/45.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
01/tiff/46.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_01/tiff/
47.tif.

Hollingdale:1958:ACWb

[HB58b] S. H. Hollingdale and M. M.

Barritt. An application of a computer to wind tunnel design: 2. The Computer Journal, 1(2):64–68, July 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
02/010064.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
02/tiff/64.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_02/tiff/
65.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_01/Issue_02/tiff/66.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
01/Issue_02/tiff/67.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
02/tiff/68.tif.

Hughes:1966:DSA

[HB66] F. M. Hughes and A. Brameller. Digital simulation of analogue methods. The Computer Journal, 9(1):35–44, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/090035.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/35.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_01/tiff/
36.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_09/Issue_01/tiff/37.

tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
09/Issue_01/tiff/38.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/39.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_01/tiff/
40.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_09/Issue_01/tiff/41.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
09/Issue_01/tiff/42.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/43.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_01/tiff/
44.tif.

Hopgood:1967:AAP

[HB67] F. R. A. Hopgood and A. G.

Bell. The Atlas ALGOL preprocessor for non-standard dialects. The Computer Journal, 9(4):360–364, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/090360.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/360.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/361.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/362.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_

                                        04/tiff/363.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        04/tiff/364.tif.

Haarhoff:1970:NMO

[HB70] P. C. Haarhoff and J. D. Buys. A new method for the optimization of a nonlinear function subject to nonlinear constraints. The Computer Journal, 13(2):178–184, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/130178.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/tiff/178.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/tiff/179.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/tiff/180.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/tiff/181.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/tiff/182.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/tiff/183.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/tiff/184.tif.

Honig:1978:TUA

[HC78] William L. Honig and C. Robert Carlson. Toward an understanding of (actual) data structures. The Computer Journal, 21(2):98–104, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/100.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/101.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/102.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/103.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/104.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/98.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_02/tiff/
99.tif.

Hay:1974:IRS

[HCC74] J. L. Hay, R. E. Crosbie, and R. I. Chaplin. Integration routines for systems with discontinuities. The Computer Journal, 17(3):275–278, August 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/275.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/276.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_

                                        03/tiff/277.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        03/tiff/278.tif.

Hsu:1996:GNS

[HCH96a] Wen-Jing Hsu, Moon Jung Chung, and Zejun Hu. Gaussian networks for scalable distributed systems. The Computer Journal, 39(5):417–426, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://www.oup.co.uk/jnls/
                                        list/comjnl/hdb/Volume_39/
                                        Issue_05/390417.sgm.abs.html;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_39/Issue_05/
                                        Vol39_05.body.html#AbstractHsu.

Huang:1996:RPV

[HCH96b] Chung-Ming Huang, Ye-In Chang, and Duen-Tay Huang. Reverse protocol verification: concept, algorithm and application. The Computer Journal, 39(6):511–524, ???? 1996. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://www.oup.co.uk/jnls/
                                        list/comjnl/hdb/Volume_39/
                                        Issue_06/390511.sgm.abs.html;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_39/Issue_06/
                                        Vol39_06.body.html#AbstractHuang1.

Hanson:1971:CAS

[HD71] R. J. Hanson and P. Dyer. A computational algorithm for sequential estimation. The Computer Journal, 14(3):285– 290, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
14/Issue_03/140285.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
14/Issue_03/tiff/285.tif;
 http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/286.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/287.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/288.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/289.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/290.tif.

Hopgood:1972:QHM

[HD72] F. R. A. Hopgood and J. Davenport. The quadratic hash method when the table size is a power of 2. The Computer Journal, 15(4):314–315, November 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/150314.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/314.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/315.tif. See corre-

spondence [PH73].

Horspool:1982:AOS

[HD82] R. Nigel Horspool and Laurence W. Dunkelman. Analysis and optimization of set expressions. The Computer Journal, 25(3):340–346, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        03/tiff/340.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        03/tiff/341.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        03/tiff/342.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        03/tiff/343.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        03/tiff/344.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        03/tiff/345.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        03/tiff/346.tif.

Healy:1963:PMR

[Hea63a] M. J. R. Healy. Programming multiple regression. The Computer Journal, 6(1):57–61, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        01/060057.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_

01/tiff/57.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/
58.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_06/Issue_01/tiff/59.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
06/Issue_01/tiff/60.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/tiff/61.tif.

Heap:1963:PI

[Hea63b] B. R. Heap. Permutations by interchanges. The Computer Journal, 6(3):293–294, October 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/060293.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/293.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/294.tif.

Healy:1968:TFV

[Hea68] M. J. R. Healy. Towards FOR-TRAN VI? The Computer Journal, 11(2):169–172, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/110169.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/169.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_11/Issue_
                                        02/tiff/170.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        02/tiff/171.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        02/tiff/172.tif.

Healey:1981:PIM

[Hea81] M. Healey. The possible impact of minicomputers and microprocessors on mainframe computer manufacturers. The Computer Journal, 24(1):3–8, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        01/tiff/3.tif;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_24/Issue_01/tiff/
                                        4.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_24/Issue_01/tiff/5.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        24/Issue_01/tiff/6.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        01/tiff/7.tif;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_24/Issue_01/tiff/

8.tif. See correspondence [Wat82].

Hekmatpour:1989:WMU

[Hek89] S. Hekmatpour. A window manager for UNIX. The Computer Journal, 32(1):21–23, February 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
01/tiff/21.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_01/tiff/
22.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_32/Issue_01/tiff/23.
tif.

Helsgaun:1984:BPS

[Hel84]  Keld Helsgaun.  Backtrack

programming with SIMULA. The Computer Journal, 27(2): 151–158, May 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
27/Issue_02/tiff/151.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/152.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/153.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/154.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/155.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/156.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/157.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/158.tif.

Hellwagner:1993:DCS

[Hel93] Hermann Hellwagner. Design considerations for scalable parallel file systems. The Computer Journal, 36(8):741–755, ???? 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_08/
Vol36_08.body.html#AbstractHellwagner.

Hennell:1978:ETN

[Hen78] M. A. Hennell. An experimental testbed for numerical software. The Computer Journal, 21(4):333–336, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/333.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/334.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/335.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/336.tif.

Henson:1988:HTT

[Hen88] M. C. Henson. Higher-order transformations and type simulations. The Computer Journal, 31(6):517–524, December 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/517.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_

06/tiff/518.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/519.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/520.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/521.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/522.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/523.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/524.tif.

Herring:1961:BRA

[Her61] G. J. Herring. Book review:

Analogue Computation in Engineering Design, by A. E. Rogers and T. W. Connolly, 1960. The Computer Journal, 4(1):41, April 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/tiff/41.tif.

Herbert:1991:EDS

[Her91] Andrew Herbert. Editorial:

Distributed systems. The Computer Journal, 34(6):481, December 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
34/Issue_06/tiff/481.tif.

Hext:1967:CTT

[Hex67]  J. B. Hext.  Compile-time

type-matching. The Computer Journal, 9(4):365–369, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/090365.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/365.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/366.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/367.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/368.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/369.tif.

Huesmann:1967:ECS

[HG67] L. Rowell Huesmann and R. P. Goldberg. Evaluating computer systems through simulation. The Computer Journal, 10(2):150–156, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/100150.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/150.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_

02/tiff/151.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/152.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/153.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/154.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/155.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/156.tif.

Houston:1971:IBT

[HG71] G. Houston and R. Gillespie.

Integration of batch and timesharing services. The Computer Journal, 14(4):348–351, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/140348.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/348.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/349.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/350.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/351.tif.

Hui:1993:MIS

[HGL93] S. C. Hui, A. Goh, and L. H.

Lau. A multimedia information system for IC failure analysis. The Computer Journal, 36(1):97–105, February 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_36/Issue_01/
                                        Vol36_01.body.html#AbstractHui.

HughesJones:1975:AGC

[HGMM75] R. Hughes Jones, P. R. Graves-Morris, and G. J. Makinson. Algorithm 85: Generation of Chisholm approximants. The Computer Journal, 18(1):81–83, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        01/tiff/81.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_18/Issue_01/tiff/
                                        82.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_18/Issue_01/tiff/83.
                                        tif.

Hittos:1976:OMM

[HH76] R. L. Hittos and D. S. Henderson. Optimisation of main memory size versus speed. The Computer Journal, 19(1):40–42, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        01/tiff/40.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_19/Issue_01/tiff/

41.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_19/Issue_01/tiff/42.
tif.

Hennell:1979:ETN

[HH79a] M. A. Hennell and D. Hedley. An experimental testbed for numerical software. II. AL-GOL 68. The Computer Journal, 22(1):53–56, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
01/tiff/53.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_01/tiff/
54.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_22/Issue_01/tiff/55.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
22/Issue_01/tiff/56.tif.

Howkins:1979:TMP

[HH79b] T. J. Howkins and M. T. Harandi. Towards more portable Cobol. The Computer Journal, 22(4):290–295, November 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/290.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/291.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/292.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_22/Issue_
04/tiff/293.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/294.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/295.tif.

Haskell:1980:SCN

[HH80] R. Haskell and P. G. Harrison. System conventions for non procedural languages. The Computer Journal, 23(2):132–141, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/132.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/133.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/134.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/135.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/136.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/137.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/138.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/139.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/140.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/141.tif.

Hamdi:1994:ECI

[HH94a] M. Hamdi and R. W. Hall.

An efficient class of interconnection networks for parallel computations. The Computer Journal, 37(3):206–218, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_37/Issue_03/
Vol37_03.body.html#AbstractHamdi.

Huang:1994:IPV

[HH94b] Chung-Ming Huang and Jenq-Muh Hsu. An incremental protocol verification method. The Computer Journal, 37 (8):698–710, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_37/Issue_08/
Vol37_08.body.html#AbstractHuang.

Harrison:1995:EQR

[HH95] Peter Harrison and Jane Hillston. Exploiting quasi-reversible structures in Markovian process algebra models. The Computer Journal, 38(7):510–520, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_38/Issue_07/
Vol38_07.body.html#AbstractHarrison.

Hickman:1960:EED

[Hic60] T. C. Hickman. Early experiences with an E.D.P. system. The Computer Journal, 2(4):152–163, January 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/020152.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/152.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/153.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/154.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/155.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/156.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/157.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/158.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/159.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/160.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/161.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_

04/tiff/162.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/163.tif.

Hierons:1996:ETS

[Hie96]  R. M. Hierons.  Extend-

ing test sequence overlap by invertibility. The Computer Journal, 39(4):325–330, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www.oup.co.uk/jnls/
list/comjnl/hdb/Volume_39/
Issue_04/390325.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_04/
Vol39_04.body.html#AbstractHierons.

Hierons:1997:TFS

[Hie97]  Robert M. Hierons.  Test-

ing from a finite-state machine: Extending invertibility to sequences. The Computer Journal, 40(4):220–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www.oup.co.uk/computer_
journal/Volume_40/Issue_03/
Vol40_04.body.html#AbstractHierons;
http://www3.oup.co.uk/computer_
journal/Volume_40/Issue_04/
Vol40_04.body.html#AbstractHierons.

Hierons:1998:ATD

[Hie98] R. M. Hierons. Adaptive testing of a deterministic implementation against a nondeterministic finite state machine. The Computer Journal, 41(5):349–355, ???? 1998. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_41/Issue_05/
Vol41_05.body.html#AbstractHierons.

Higman:1963:WES

[Hig63] B. Higman. What EVERY-BODY should know about AL-GOL. The Computer Journal, 6(1):50–56, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/060050.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/tiff/50.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/
51.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_06/Issue_01/tiff/52.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
06/Issue_01/tiff/53.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/tiff/54.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/
55.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_06/Issue_01/tiff/56.
tif.

Higman:1967:CPS

[Hig67] Bryan Higman. Correspondence: Paging and segmentation. The Computer Journal, 10(2):216, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/216.tif.

Higley:1976:TCT

[Hig76a] C. J. Higley. Type checking in a typeless language. The Computer Journal, 19(2):166–169, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/166.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/167.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/168.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/169.tif.

Higman:1976:POV

[Hig76b] B. Higman. The place of own variables in programming language theory. The Computer Journal, 19(3):225–228, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/225.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/226.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/227.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/228.tif.

Hill:1967:CEC

[Hil67] I. D. Hill. Correspondence: On “Efficiency of computer use should not be a consideration”. The Computer Journal, 10(2):209, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        02/tiff/209.tif.

Hill:1969:NAS

[Hil69] I. D. Hill. Note on Algorithms 25, 26. 25: Sort a section of the elements of an array by determining the rank of each element. 26: Order the subscripts of an array section according to the magnitudes of the elements. The Computer Journal, 12(4):409, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/409.tif. See [Boo67f,

Boo67g, Woo70g].

Hill:1971:FFA

[Hil71] I. D. Hill. Faults in functions, in ALGOL and FOR-TRAN. The Computer Journal, 14(3):315–316, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_

journal/hdb/Volume_14/Issue_
03/140315.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/315.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/316.tif. See notes

[SS72a].

Hill:1972:CHL

[Hil72] I. D. Hill. Correspondence:

On “High level languages for low level users”. The Computer Journal, 15(1):17, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/150017.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/tiff/17.tif. See [Bac71,

ER71, Wal71, Fel72, Hal72].

Hill:1975:AEM

[Hil75] I. D. Hill. Algorithm 88:

An electoral method. The Computer Journal, 18(1):89– 90, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
18/Issue_01/tiff/89.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
01/tiff/90.tif.

Hilden:1976:GDI

[Hil76]  J. Hilden.  Guidelines for

the design of interactive systems. The Computer Journal, 19(2):144–150, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/144.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/145.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/146.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/147.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/148.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/149.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/150.tif.

Hill:1977:AHD

[Hil77] I. D. Hill. Algorithm 94: Handling dates on small computers. The Computer Journal, 20(3):280, August 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/280.tif.

Hill:1979:CPP

[Hil79a]  G. W. Hill.  Cyclic prop-

erties of pseudo-random sequences of Mersenne prime residues. The Computer Journal, 22(1):80–85, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
01/tiff/80.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_01/tiff/
81.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_22/Issue_01/tiff/82.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
22/Issue_01/tiff/83.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
01/tiff/84.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_01/tiff/
85.tif.

Hill:1979:AGS

[Hil79b] I. D. Hill. Algorithm 109:

Golden section search. The Computer Journal, 22(4):383– 384, November 1979. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
22/Issue_03/tiff/383.tif;
 http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/384.tif.

Hill:1979:AE

[Hil79c]  I. D. Hill.  Algorithm 3:

Easter. The Computer Journal, 22(1):86–87, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_22/Issue_
                                        01/tiff/86.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_22/Issue_01/tiff/

87.tif. Reprinted from The Computer Bulletin. See [Boo79d, Ran79].

Hill:1980:CJS

[Hil80]  I. D. Hill.  Correspon-

dence: Jumping to some purpose. The Computer Journal, 23(1):94, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        01/tiff/94.tif. See [ASG79,

Rob80, Tay80, MR82].

Hindle:1960:BAA

[Hin60] R. Hindle. A bank adopts automatic data processing. The Computer Journal, 3 (3):127–130, October 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        03/030127.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        03/tiff/127.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        03/tiff/128.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        03/tiff/129.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_03/Issue_
03/tiff/130.tif.

Hindle:1961:CRD

[Hin61] R. Hindle. Character recognition and document handling in banks. The Computer Journal, 4(2):157–160, July 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/040157.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/157.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/158.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/159.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/160.tif.

Hiorns:1967:SAF

[Hio67]  R. H. Hiorns.  Seasonal

adjustment and forecasting in the presence of trend. The Computer Journal, 10 (2):143–149, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/100143.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/143.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_

                                        02/tiff/144.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        02/tiff/145.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        02/tiff/146.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        02/tiff/147.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        02/tiff/148.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        02/tiff/149.tif.

Hitchin:1978:CMF

[Hit78] D. Hitchin. Correspondence: On “Modifications to the Forsythe-Moler algorithm for solving linear algebraic systems”. The Computer Journal, 21(4):384, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        04/tiff/384.tif. See [SG78,

LS78].

Huang:1999:ITB

[HJ99] Chung-Ming Huang and Ming-Yuhe Jang. Interactive temporal behaviours and modelling for multimedia presentations in the WWW environment. The Computer Journal, 42(2):112–128, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_

02/420112.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
02/pdf/420112.pdf.

Hill:1967:NAN

[HJBH67] I. D. Hill, S. A. Joyce, A. Bergson, and P. Hammersley. Note on Algorithm 13: normalarea. The Computer Journal, 10(1):113–115, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/113.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/114.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/115.tif.

Howarth:1962:ASS

[HJW62] D. J. Howarth, P. D. Jones, and M. T. Wyld. The Atlas scheduling system. The Computer Journal, 5(3):238–246, October 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/050238.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/238.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/239.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/240.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/241.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/242.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/243.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/244.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/245.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/246.tif.

Hemy:1959:SAC

[HK59] D. C. Hemy and W. J. Kease. The study of the application of a computer to production control. The Computer Journal, 2(1):24–38, April 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
01/020024.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
01/tiff/24.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_01/tiff/
25.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_02/Issue_01/tiff/26.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
02/Issue_01/tiff/27.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_

01/tiff/28.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_01/tiff/
29.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_02/Issue_01/tiff/30.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
02/Issue_01/tiff/31.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
01/tiff/32.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_01/tiff/
33.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_02/Issue_01/tiff/34.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
02/Issue_01/tiff/35.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
01/tiff/36.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_01/tiff/
37.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_02/Issue_01/tiff/38.
tif.

Harris:1967:CAL

[HK67] D. J. Harris and A. K. Kent.

The computer as an aid to lawyers. The Computer Journal, 10(1):22–28, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/100022.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/22.tif;   http://

                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_10/Issue_01/tiff/
                                        23.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_10/Issue_01/tiff/24.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        10/Issue_01/tiff/25.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        01/tiff/26.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_10/Issue_01/tiff/
                                        27.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_10/Issue_01/tiff/28.
                                        tif.

Hoskins:1972:APC

[HK72] W. D. Hoskins and P. R. King. Algorithm 73: Periodic cubic spline interpolation using parametric splines. The Computer Journal, 15(3):282–283, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/282.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/283.tif.

Higginson:1973:CCR

[HK73] P. L. Higginson and P. T. Kirstein. On the computation of cyclic redundancy checks by program. The Computer Journal, 16(1):19–22, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_

journal/hdb/Volume_16/Issue_
01/160019.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
01/tiff/19.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_01/tiff/
20.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_16/Issue_01/tiff/21.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
16/Issue_01/tiff/22.tif.

Hatzopoulos:1980:SRI

[HK80] M. Hatzopoulos and J. G. Kollias. Some rules for introducing indexing paths in a primary file. The Computer Journal, 23(3):207–211, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/207.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/208.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/209.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/210.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/211.tif.

Hatzopoulos:1982:DOD

[HK82] M. Hatzopoulos and J. G. Kollias. The determination of the optimum database maintenance points. The Computer Journal, 25(1):126–129, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/126.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/127.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/128.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/129.tif.

Hatzopoulos:1985:SRS

[HK85a] Michael Hatzopoulos and J. (Yiannis) Kollias. On the selection of a reduced set of indexes. The Computer Journal, 28(4):406–408, August 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        04/tiff/406.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        04/tiff/407.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        04/tiff/408.tif.

Hatzopoulos:1985:TOS

[HK85b] Michalis Hatzopoulos and J. G. Kollias. Towards the optimal secondary index organisation and secondary index selection. The Computer Journal, 28(5):524–529, November 1985. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/524.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/525.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/526.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/527.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/528.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/529.tif.

Harrison:1988:ATT

[HK88] P. G. Harrison and H. Khoshnevisan. Algebraic transformation techniques for functional languages. The Computer Journal, 31(3):229–242, June 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/229.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/230.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/231.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/232.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_

                                        03/tiff/233.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        03/tiff/234.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        03/tiff/235.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        03/tiff/236.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        03/tiff/237.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        03/tiff/238.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        03/tiff/239.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        03/tiff/240.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        03/tiff/241.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        03/tiff/242.tif.

Harrison:1992:MTD

[HK92] P. G. Harrison and H. Khoshnevisan. The mechanical transformation of data types. The Computer Journal, 35(2):138– 147, April 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/Volume_
                                        35/Issue_02/Vol35_02.body.
                                        html#AbstractHarrison.

Hoskins:1972:AIU

[HKA72] W. D. Hoskins, P. R. King, and T. H. Andres. Algorithm 74: Interpolation using periodic splines of odd order with equidistant knots. The Computer Journal, 15(3):283–285, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/283.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/284.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/285.tif. See note

[Hos74].

Hermann:1991:ITR

[HKK91] M. Hermann, C. Kirchner, and H. Kirchner. Implementations of term rewriting systems. The Computer Journal, 34(1):20–33, February 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
01/tiff/20.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_01/tiff/
21.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_34/Issue_01/tiff/22.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
34/Issue_01/tiff/23.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
01/tiff/24.tif;   http://
www3.oup.co.uk/computer_journal/

                                        hdb/Volume_34/Issue_01/tiff/
                                        25.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_34/Issue_01/tiff/26.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        34/Issue_01/tiff/27.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        01/tiff/28.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_34/Issue_01/tiff/
                                        29.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_34/Issue_01/tiff/30.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        34/Issue_01/tiff/31.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        01/tiff/32.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_34/Issue_01/tiff/
                                        33.tif.

Huhtala:1999:TEA

[HKPT99] Ykä Huhtala, Juha Kärkkäinen, Pasi Porkka, and Hannu Toivonen. TANE: An efficient algorithm for discovering functional and approximate dependencies. The Computer Journal, 42(2):100–111, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        02/420100.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        02/pdf/420100.pdf.

Heuring:1989:CDF

[HKPW89] V. P. Heuring, U. Kastens, R. G. Plummer, and W. M. Waite. COMAR: a data format for integration of CFG tools. The Computer Journal, 32(5):445–452, October 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/445.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/446.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/447.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/448.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/449.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/450.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/451.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/452.tif.

Hawley:1975:DCC

[HKT75] D. A. Hawley, J. S. Knowles, and E. E. Tozer. Database consistency and the CO-DASYL DBTG proposals. The Computer Journal, 18 (3):206–212, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/206.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/207.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/208.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/209.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/210.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/211.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/212.tif.

Halang:1995:EGP

[HKT95] Wolfgang A. Halang, Bernd Krämer, and Leszek Trybus. Exploiting a graphical programming paradigm to facilitate rigorous verification of embedded software. The Computer Journal, 38(4):301–309, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_38/Issue_04/
                                        Vol38_04.body.html#AbstractHalang.

Hammersley:1970:CTF

[HL70] P. Hammersley and J. Larmouth. Correspondence: Towards Fortran VI? The Computer Journal, 13(2):220, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/220.tif. See [Red71,

Wel70].

Henderson:1976:EOP

[HL76a] D. S. Henderson and M. R.

Levy. An extended operator precedence parsing algorithm. The Computer Journal, 19(3):229–233, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/229.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/230.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/231.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/232.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/233.tif.

Henshell:1976:PDW

[HL76b] R. D. Henshell and H. C.

Lauer. Ph.D. work in Computer Science and Professional Standards. The Computer Journal, 19(3):274–275, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_

                                        02/tiff/274.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/275.tif. See [Lau75].

Huang:1992:FSM

[HL92] B.-C. Huang and M. A. Langston. Fast stable merging and sorting in constant extra space. The Computer Journal, 35(6):643–650, December 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_36/Issue_06/
                                        Vol35_06.body.html#AbstractHuang.

Hao:1994:ASL

[HLK94] Ying Hao, Jyh-Charn S. Liu, and Junguk L. Kim. An allsharing load-balancing scheme on the COMA/CD network and its analysis. The Computer Journal, 37(9):779–794, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_37/Issue_09/
                                        Vol37_09.body.html#AbstractHao.

Hartley:1968:SMS

[HLN68] D. F. Hartley, B. Landy, and R. M. Needham. The structure of a multiprogramming supervisor. The Computer Journal, 11(3):247–255, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_

03/110247.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/247.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/248.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/249.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/250.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/251.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/252.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/253.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/254.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/255.tif.

Higman:1965:CIP

[HLOS65] Bryan Higman, W. F. Lunnon, C. F. J. Outred, and C. Strachey. Correspondence: An impossible program. The Computer Journal, 8(2):175–176, July 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/175.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/176.tif.

Humphrey:1970:MIR

[HM70] A. L. Humphrey and W. G. Munro. Management information retrieval. The Computer Journal, 13(2):127–130, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/130127.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/tiff/127.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/tiff/128.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/tiff/129.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/tiff/130.tif.

Hoskins:1977:ANS

[HM77] W. D. Hoskins and G. E. Mc-Master. Algorithm 98: A note on the solution of certain tridiagonal systems of linear equations. The Computer Journal, 20(4):376–378, November 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        04/tiff/376.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        04/tiff/377.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        04/tiff/378.tif.

Hansen:1980:CAA

[HM80a] J. V. Hansen and Lynn J.

McKell. A computer aid for the analysis of complex systems. The Computer Journal, 23(2):98–106, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/100.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/101.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/102.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/103.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/104.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/105.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/106.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/98.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_02/tiff/
99.tif.

Horspool:1980:APD

[HM80b] R. N. Horspool and N. Marovac.

An approach to the problem of detranslation of computer programs. The Computer Journal, 23(3):223–229, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/223.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/224.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/225.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/226.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/227.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/228.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/229.tif.

Hunter:1983:ERA

[HM83] D. G. N. Hunter and A. R. McKenzie. Experiments with relaxation algorithms for breaking simple substitution ciphers. The Computer Journal, 26(1):68–71, February 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
01/tiff/68.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_01/tiff/
69.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_26/Issue_01/tiff/70.
tif;   http://www3.oup.co.

uk/computer_journal/hdb/Volume_
26/Issue_01/tiff/71.tif.

Hull:1984:CDD

[HM84] M. Elizabeth C. Hull and R. M.

McKeag. Concurrency in the design of data processing systems. The Computer Journal, 27(4):289–293, November 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/289.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/290.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/291.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/292.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/293.tif.

Hains:1993:PFP

[HM93] G. Hains and L. M. R. Mullin.

Parallel functional programming with arrays. The Computer Journal, 36(3):238–245, June 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_03/
Vol36_03.body.html#AbstractHains.

Hillston:1995:STS

[HM95a] Jane Hillston and Vassilis Mertsiotakis. A simple time scale decomposition technique for stochastic process algebras. The Computer Journal, 38(7):566–577, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_38/Issue_07/
                                        Vol38_07.body.html#AbstractHillston.

Homeier:1995:MVV

[HM95b] Peter V. Homeier and David F. Martin. A mechanically verified verification condition generator. The Computer Journal, 38(2):131–141, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_38/Issue_02/
                                        Vol38_02.body.html#AbstractHomeier.

Hui:1996:RCA

[HM96] Lucas Chi Kwong Hui and Charles U. Martel. Randomized competitive algorithms for successful and unsuccessful search. The Computer Journal, 39(5):427–438, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://www.oup.co.uk/jnls/
                                        list/comjnl/hdb/Volume_39/
                                        Issue_05/390427.sgm.abs.html;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_39/Issue_05/
                                        Vol39_05.body.html#AbstractHui.

Hoggar:1998:FCJ

[HM98] Stuart Hoggar and Lindsey Menzies. Fractal compression and the jigsaw property I. The Computer Journal, 41(5):319–336, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_41/Issue_05/
Vol41_05.body.html#AbstractHoggar.

Haley:1979:SPR

[HN79] J. R. Haley and A. S. Noetzell. A study of processing repetition. The Computer Journal, 22(4):352–356, November 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/352.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/353.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/354.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/355.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/356.tif.

Hirschheim:1988:ISU

[HN88] R. Hirschheim and M. Newman. Information systems and user resistance: theory and practice. The Computer Journal, 31(5):398–408, October 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/398.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/399.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/400.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/401.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/402.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/403.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/404.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/405.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/406.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/407.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/408.tif.

Hill:1978:RAM

[HO78] I. D. Hill and Alex K. Obruča. Remarks on Algorithm 1: Mintree. The Computer Journal, 21(3):277, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_

journal/hdb/Volume_21/Issue_
03/tiff/277.tif. See [Obr78a].

Hull:1993:FRB

[HO93] M. Elizabeth C. Hull and P. G. O’Donoghue. Family relationships between requirements and design specification methods. The Computer Journal, 36(2):153–167, April 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_02/
Vol36_02.body.html#AbstractHull.

Hoare:1962:Q

[Hoa62a] C. A. R. Hoare. Quicksort. The Computer Journal, 5(1):10–16, April 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
01/050010.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
01/tiff/10.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_01/tiff/
11.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_05/Issue_01/tiff/12.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
05/Issue_01/tiff/13.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
01/tiff/14.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_01/tiff/
15.tif;   http://www3.oup.

co.uk/computer_journal/hdb/
Volume_05/Issue_01/tiff/16.
tif.

Hoare:1962:REA

[Hoa62b] C. A. R. Hoare. Report on the Elliott ALGOL translator. The Computer Journal, 5(2):127–129, July 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/050127.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/127.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/128.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/129.tif.

Hoare:1963:EAI

[Hoa63] C. A. R. Hoare. The Elliott ALGOL input/output system. The Computer Journal, 5(4):345–348, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/050345.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/345.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/346.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_

04/tiff/347.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/348.tif.

Hoare:1972:PSP

[Hoa72] C. A. R. Hoare. Proof of a structured program: The sieve of Eratosthenes’. The Computer Journal, 15(4):321–325, November 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/150321.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/321.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/322.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/323.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/324.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/325.tif.

Hoare:1973:SPS

[Hoa73] C. A. R. Hoare. A structured paging system. The Computer Journal, 16(3):209–215, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/160209.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_

03/tiff/209.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/210.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/211.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/212.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/213.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/214.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/215.tif.

Hockney:1962:AAE

[Hoc62] R. W. Hockney. ABS12 AL-GOL: an extension to Algol 60 for industrial use. The Computer Journal, 4 (4):292–300, January 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/040292.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/292.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/293.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/294.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/295.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/296.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/297.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/298.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/299.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/300.tif.

Hull:1991:DMR

[HOH91] M. Elizabeth C. Hull, P. G.

O’Donoghue, and B. J. Hagan. Development methods for real-time systems. The Computer Journal, 34(2):164–172, April 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/164.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/165.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/166.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/167.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/168.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/169.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_34/Issue_
                                        02/tiff/170.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        02/tiff/171.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        02/tiff/172.tif.

Holmes:1975:IND

[Hol75] W. N. Holmes. Identification number design. The Computer Journal, 18(2):102–107, May 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        02/tiff/102.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        02/tiff/103.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        02/tiff/104.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        02/tiff/105.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        02/tiff/106.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        02/tiff/107.tif.

Holmes:1978:APL

[Hol78] W. N. Holmes. Is APL a programming language? The Computer Journal, 21(2):128–131, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_

02/tiff/128.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/129.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/130.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/131.tif.

Holton:1995:PSI

[Hol95] D. R. W. Holton. A PEPA specification of an industrial production cell. The Computer Journal, 38(7):542–551, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_38/Issue_07/
Vol38_07.body.html#AbstractHolton.

Honey:1971:CCD

[Hon71] D. W. Honey. Correspondence: Calculation of a doublelength square root from a double length number using single precision techniques. The Computer Journal, 14(4):443, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/140443.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/443.tif.

Honey:1974:MTE

[Hon74] D. W. Honey. The ‘Queensway’ Mersey tunnel entrance control project. The Computer Journal, 17(4):290–294, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/290.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/291.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/292.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/293.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/294.tif.

Hong:1993:QEF

[Hon93a] Hoon Hong. Quantifier elimination for formulas constrained by quadratic equations via slope resultants. The Computer Journal, 36(5):439–449, October 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_05/
Vol36_05.body.html#AbstractHong.

Hong:1993:SIE

[Hon93b] Hoon Hong. Special issue editorial: Computational quantifier elimination. The Computer Journal, 36(5):399–??, October 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

journal/Volume_36/Issue_05/
Vol36_05.body.html#AbstractHong.

Hope:1961:SRT

[Hop61]  K. S. Hope.  SABER: a

real time problem in teleprocessing. The Computer Journal, 4(2):109–113, July 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/040109.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/109.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/110.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/111.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/112.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/113.tif.

Hopkinson:1963:IAU

[Hop63] J. R. Hopkinson. Integrated accounting using a variety of equipment. The Computer Journal, 6(1):5–6, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/060005.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/tiff/5.tif;   http://www3.

                                        oup.co.uk/computer_journal/
                                        hdb/Volume_06/Issue_01/tiff/
                                        6.tif.

Hope:1971:HOS

[Hop71]  E. L. V. Hope.  A high-

order solution to the characteristic initial value problem in hyperbolic partial differential equations. The Computer Journal, 14(1):65–72, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        01/140065.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        01/tiff/65.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_14/Issue_01/tiff/
                                        66.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_14/Issue_01/tiff/67.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        14/Issue_01/tiff/68.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        01/tiff/69.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_14/Issue_01/tiff/
                                        70.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_14/Issue_01/tiff/71.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        14/Issue_01/tiff/72.tif.

Hope:1978:CSP

[Hop78]  Eric L. V. Hope.  Corre-

spondence: Structured programming input statements.

The Computer Journal, 21(4): 382, November 1978. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
21/Issue_04/tiff/382.tif.

See [Ing76, DI78].

Hopper:1981:ROO

[Hop81] K. Hopper. Retained objects and operating system interfaces. The Computer Journal, 24(4):331–338, November 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/331.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/332.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/333.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/334.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/335.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/336.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/337.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/338.tif.

Hornsby:1960:FIS

[Hor60] J. S. Hornsby. A function interpretive scheme for Pegasus. The Computer Journal, 2(4):174–180, January 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/020174.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/174.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/175.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/176.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/177.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/178.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/179.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/180.tif.

Horng:1993:CTA

[Hor93] S. J. Horng. Constant time algorithm for template matching on a reconfigurable array of processors. The Computer Journal, 36(3):246–253, June 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_03/
Vol36_03.body.html#AbstractHorng.

Hoskins:1970:AIQ

[Hos70]  W. D. Hoskins.  Algo-

rithm 62: Interpolating quintic splines on equidistant knots. The Computer Journal, 13 (4):437–438, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/437.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/438.tif.

Hoshino:1971:DSC

[Hos71] S. Hoshino. On Davies, Swann and Campey minimisation process. The Computer Journal, 14(4):426–427, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/140426.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/426.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/427.tif.

Hoskins:1974:NAI

[Hos74]  W. D. Hoskins.  Note on

Algorithm 74: Interpolation using periodic splines of odd order with equi-distant knots. The Computer Journal, 17(1):91, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
01/tiff/91.tif. See [HKA72].

Houstoun:1963:BRM

[Hou63] J. Houstoun. Book review: Mathematical Methods for the Study of Automatic Control Systems, by V. I. Zubov, 1962. The Computer Journal, 6(3):247, October 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/247.tif.

Housden:1969:DIL

[Hou69a] R. J. W. Housden. The definition and implementation of Lsix in BCL. The Computer Journal, 12(1):15–23, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
01/120015.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
01/tiff/15.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_01/tiff/
16.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_12/Issue_01/tiff/17.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/18.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
01/tiff/19.tif;   http://

www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_01/tiff/
20.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_12/Issue_01/tiff/21.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/22.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
01/tiff/23.tif.

Housden:1969:NLP

[Hou69b] R. J. W. Housden. A note on list-processing in BCL. The Computer Journal, 12 (4):332–341, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/120332.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/332.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/333.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/334.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/335.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/336.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/337.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/338.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/339.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/340.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/341.tif.

Housden:1971:PSF

[Hou71] R. J. W. Housden. Phrase structures in FORTRAN. The Computer Journal, 14(3):224– 228, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        14/Issue_03/140224.sgm.abs.
                                        html;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        14/Issue_03/tiff/224.tif;
                                         http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/tiff/225.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/tiff/226.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/tiff/227.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/tiff/228.tif.

Housden:1975:SCT

[Hou75] R. J. W. Housden. On string concepts and their implementation. The Computer Journal, 18(2):150–156, May 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/150.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/151.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/152.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/153.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/154.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/155.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/156.tif.

House:1983:PEF

[Hou83] R. T. House. A proposal for an extended form of type checking of expressions. The Computer Journal, 26(4):366–374, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/366.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/367.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/368.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/369.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_

                                        04/tiff/370.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        04/tiff/371.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        04/tiff/372.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        04/tiff/373.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        04/tiff/374.tif.

House:1986:ASR

[Hou86]  R. T. House.  Alternative

scope rules for block-structured languages. The Computer Journal, 29(3):253–260, June 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        03/tiff/253.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        03/tiff/254.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        03/tiff/255.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        03/tiff/256.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        03/tiff/257.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        03/tiff/258.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        03/tiff/259.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_29/Issue_
03/tiff/260.tif.

Howell:1959:NPT

[How59] K. M. Howell. A new programming technique for rational fractions. The Computer Journal, 1(4):176–178, January 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/010176.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/tiff/176.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/tiff/177.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/tiff/178.tif.

Howden:1967:PCP

[How67a] W. E. Howden. A program for the construction of PERT flow charts. The Computer Journal, 10(3):278–281, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/100278.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/278.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/279.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/280.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        03/tiff/281.tif.

Howell:1967:MPA

[How67b] K. M. Howell. Multiple precision arithmetic techniques. The Computer Journal, 9 (4):383–387, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        04/090383.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        04/tiff/383.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        04/tiff/384.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        04/tiff/385.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        04/tiff/386.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        04/tiff/387.tif.

Howden:1968:SP

[How68] W. E. Howden. The sofa problem. The Computer Journal, 11(3):299–301, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        03/110299.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        03/tiff/299.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/300.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/301.tif.

Howell:1975:CCD

[How75] K. M. Howell. Correspondence: Check digits and error correction. The Computer Journal, 18(3):283–284, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/283.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/284.tif. See [And74,

And75, ES75].

Howard:1997:TIC

[How97] P. G. Howard. Text image compression using soft pattern matching. The Computer Journal, 40(2/3):146–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www.oup.co.uk/computer_
journal/Volume_40/Issue_02/
Vol40_02.body.html#AbstractHoward;
http://www3.oup.co.uk/computer_
journal/Volume_40/Issue_02/
Vol40_03.body.html#AbstractHoward.

Hill:1967:NAT

[HP67] I. D. Hill and M. C. Pike.

Note on algorithm 4: TWO BY TWO. The Computer Journal, 9(4):416, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/416.tif. See [HP79].

Haddon:1971:ALS

[HP71] E. W. Haddon and L. G. Proll. An ALGOL linesyntax checker. The Computer Journal, 14(2):128–132, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/140128.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/128.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/129.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/130.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/131.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/132.tif.

Huxtable:1977:HSI

[HP77] D. H. R. Huxtable and J. M. M. Pinkerton. The hardware/ software interface of the ICL 2900 range of computers. The Computer Journal, 20 (4):290–295, November 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/290.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/291.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/292.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/293.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/294.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/295.tif.

Hill:1979:AT

[HP79] I. D. Hill and M. C. Pike.

Algorithm 4: TWOBYTWO. The Computer Journal, 22 (1):87–88, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
01/tiff/87.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_01/tiff/

88.tif. Reprinted from The Computer Bulletin, September 1964, p. 67. See note [HP67].

Henstridge:1984:ABO

[HPB84] J. D. Henstridge, R. W. Payne, and R. J. Baker. Algorithm 117: Buffered output in FORTRAN. The Computer Journal, 27(2):179–184, May 1984. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/179.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/180.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/181.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/182.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/183.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/184.tif. See note

[WH87].

Haring:1979:USJ

[HPLG79] G. Haring, R. Posch, C. Leonhardt, and G. Gell. The use of a synthetic jobstream in performance evaluation. The Computer Journal, 22 (3):209–219, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/209.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/210.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/211.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_

03/tiff/212.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/213.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/214.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/215.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/216.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/217.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/218.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/219.tif.

Helary:1988:DAM

[HPR88] J.-M. Hélary, N. Plouzeau, and M. Raynal. A distributed algorithm for mutual exclusion in an arbitrary network. The Computer Journal, 31(4):289–295, August 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/289.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/290.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/291.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_

                                        04/tiff/292.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        04/tiff/293.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        04/tiff/294.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        04/tiff/295.tif.

Howarth:1961:MUA

[HPS61] D. J. Howarth, R. B. Payne, and F. H. Sumner. The Manchester University Atlas operating system part II: users’ description. The Computer Journal, 4(3):226–229, October 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        03/040226.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        03/tiff/226.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        03/tiff/227.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        03/tiff/228.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        03/tiff/229.tif.

Hext:1970:SAD

[HR70] J. B. Hext and P. S. Roberts. Syntax analysis by Domolki’s algorithm. The Computer Journal, 13(3):263–271, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/130263.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/263.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/264.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/265.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/266.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/267.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/268.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/269.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/270.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/271.tif.

Heath:1973:BRC

[HR73] F. G. Heath and C. M. Reeves.

Book reviews: Computer Logic Design, by M. Morris Mano, 1972; Logic and Logic Design, By B. Girling and H. G. Moring; An Introduction to Job Control Language, by Ronald J. DeMasi and Francis J. Longo, 1972. The Computer Journal, 16(3):204, August 1973. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        03/tiff/204.tif.

Heindel:1974:ARP

[HR74] L. E. Heindel and J. T. Roberto. ARPL — A Retrieval Process Language. The Computer Journal, 17(2):113–116, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        02/tiff/113.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        02/tiff/114.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        02/tiff/115.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        02/tiff/116.tif.

Huang:1996:UTP

[HR96] Yen-Min Huang and Chinya V. Ravishankar. URPC: a toolkit for prototyping remote procedure calls. The Computer Journal, 39(6):525–540, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://www.oup.co.uk/jnls/
                                        list/comjnl/hdb/Volume_39/
                                        Issue_06/390525.sgm.abs.html;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_39/Issue_06/
                                        Vol39_06.body.html#AbstractHuang2.

Harris:1997:OCR

[HR97] Evan P. Harris and Kotagiri Ramamohanarao. Optimal clustering of relations to improve sorting and partitioning for joins. The Computer Journal, 40(7):416–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_40/Issue_07/
Vol40_07.body.html#AbstractHarris.

Hunt:1968:CSF

[HRS68] C. H. Hunt, M. M. Ryder, and S. H. Storey. Computers and the small firm: 1. The Computer Journal, 11(1):1–4, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/110001.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/1.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_11/Issue_01/tiff/
2.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_11/Issue_01/tiff/3.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/4.tif.

Heintz:1993:TPC

[HRS93] Joos Heintz, Marie-Françoise Roy, and Pablo Solernó. On the theoretical and practical complexity of the existential theory of reals. The Computer Journal, 36(5):427–431, October 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_05/
Vol36_05.body.html#AbstractHeintz.

Hermanns:1995:FCI

[HRW95] Holger Hermanns, Michael Rettelbach, and Thorsten Weiss. Formal characterisation of immediate actions in SPA with nondeterministic branching. The Computer Journal, 38(7):530–541, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_38/Issue_07/
Vol38_07.body.html#AbstractHermanns.

Hendry:1969:TFV

[HS69] D. F. Hendry and P. A. Samet. Towards FORTRAN VI? Part

2. FORTRAN in the modern world. The Computer Journal, 12(3):218–220, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/120218.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/218.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/219.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_12/Issue_
03/tiff/220.tif.

Higgins:1971:DAA

[HS71] L. D. Higgins and Francis J.

Smith. Disc access algorithms. The Computer Journal, 14(3):249–253, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/140249.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/249.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/250.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/251.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/252.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/253.tif.

Hudson:1982:DST

[HS82] R. W. A. Hudson and J. A.

Self. A dialogue system to teach database concepts. The Computer Journal, 25 (1):135–139, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/135.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_

                                        01/tiff/136.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/137.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/138.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/139.tif.

Hehner:1983:PGE

[HS83] Eric C. R. Hehner and Brad A. Silverberg. Programming with grammars: an exercise in methodology-directed language design. The Computer Journal, 26(3):277–281, August 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        03/tiff/277.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        03/tiff/278.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        03/tiff/279.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        03/tiff/280.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        03/tiff/281.tif.

Horn:1993:PSM

[HS93] F. Horn and J. B. Stefani. On programming and supporting multimedia object synchronisation. The Computer Journal, 36(1):4–18, February 1993. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_01/
Vol36_01.body.html#AbstractHorn.

Humby:1966:CCE

[HSP66] E. Humby, P. Shackleton, and R. M. Paine. Correspondence: Commercial English languages. The Computer Journal, 8(4):310–311, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/310.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/311.tif.

Hwang:1992:NAC

[HSW92] J.-J. Hwang, B.-M. Shao, and P. C. Wang. A new access control method using prime factorisation. The Computer Journal, 35(1):16–20, February 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_35/Issue_01/
Vol35_01.body.html#AbstractHwang.

Haskell:1978:CDS

[HT78] R. Haskell and The Editor.

Correspondence: Dissemination of specialised information. The Computer Journal, 21(4):382, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/382.tif.

Hartley:1981:CWA

[HT81] D. F. Hartley and The Editor. Correspondence: The Wilkes Award. The Computer Journal, 24(2):192, May 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/192.tif.

Halbert:1989:GAZ

[HTW89] A. R. Halbert, S. J. P. Todd, and J. R. Woodwark. Generalizing active zones for set-theoretic solid models. The Computer Journal, 32(1):86–89, February 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
01/tiff/86.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_01/tiff/
87.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_32/Issue_01/tiff/88.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
32/Issue_01/tiff/89.tif.

Hughes:1961:CPC

[Hug61] D. J. L. Hughes. Computer production control — the second year. The Computer Journal, 3(4):198–201, January 1961. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/030198.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/198.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/199.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/200.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/201.tif.

Hughes:1989:WFP

[Hug89]  R. J. M. Hughes.  Why

functional programming matters. The Computer Journal, 32(2):98–107, April 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/100.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/101.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/102.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/103.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/104.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/105.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        02/tiff/106.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        02/tiff/107.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        02/tiff/98.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_32/Issue_02/tiff/
                                        99.tif.

Hull:1986:ICN

[Hul86] M. Elizabeth C. Hull. Implementations of the CSP notation for concurrent systems. The Computer Journal, 29 (6):500–505, December 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        06/tiff/500.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        06/tiff/501.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        06/tiff/502.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        06/tiff/503.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        06/tiff/504.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        06/tiff/505.tif.

Humby:1962:RNA

[Hum62] E. Humby. RAPIDWRITE — a new approach to COBOL readability. The Computer Journal, 4(4):301–304, January 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/040301.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/301.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/302.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/303.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/304.tif.

Hunt:1959:FPD

[Hun59] P. M. Hunt. The Ferranti Perseus data-processing system. The Computer Journal, 2(2):68–75, July 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
02/020068.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
02/tiff/68.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_02/tiff/
69.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_02/Issue_02/tiff/70.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
02/Issue_02/tiff/71.tif;
http://www3.oup.co.uk/computer_

                                        journal/hdb/Volume_02/Issue_
                                        02/tiff/72.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_02/Issue_02/tiff/
                                        73.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_02/Issue_02/tiff/74.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        02/Issue_02/tiff/75.tif.

Hunter:1960:NTR

[Hun60] D. G. N. Hunter. Note on a test for repeating cycles in a pseudorandom number generator. The Computer Journal, 3(1):9, April 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        01/030009.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        01/tiff/9.tif.

Hunter:1961:IMN

[Hun61]  D. B. Hunter.  An itera-

tive method of numerical differentiation. The Computer Journal, 3(4):270–271, January 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        04/030270.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        04/tiff/270.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        04/tiff/271.tif.

Hunter:1968:NMT

[Hun68] D. B. Hunter. Neville’s method for trigonometric interpolation. The Computer Journal, 11(3):311–313, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/110311.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/311.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/312.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/313.tif.

Hunter:1970:CME

[Hun70] D. B. Hunter. Clenshaw’s method for evaluating certain finite series. The Computer Journal, 13(4):378–381, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/130378.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/378.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/379.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/380.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_13/Issue_
04/tiff/381.tif.

Hunt:1972:NSF

[Hun72] L. B. Hunt. Numerical smoothing and filtering in N dimensions. The Computer Journal, 15(1):58–65, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/150058.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/tiff/58.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_01/tiff/
59.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_15/Issue_01/tiff/60.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
15/Issue_01/tiff/61.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/tiff/62.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_01/tiff/
63.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_15/Issue_01/tiff/64.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
15/Issue_01/tiff/65.tif.

Hunter:1975:QMP

[Hun75] G. Hunter. A quantitative measure of precision. The Computer Journal, 18(3):231–233, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/231.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/232.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/233.tif.

Hunt:1981:CPT

[Hun81]  J. G. Hunt.  Correspon-

dence: Premature termination. The Computer Journal, 24(1):95, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
01/tiff/95.tif. See [Zar80].

Hunter:1985:ARK

[Hun85]  D. G. N. Hunter.  Algo-

rithm 121: RSA key calculation in Ada. The Computer Journal, 28(3):343–348, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/343.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/344.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/345.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/346.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_

03/tiff/347.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/348.tif. See note

[Wic87].

Hurst:1968:EBM

[Hur68]  S. L. Hurst.  An exten-

sion of binary minimisation techniques to ternary equations. The Computer Journal, 11(3):277–286, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/110277.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/277.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/278.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/279.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/280.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/281.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/282.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/283.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/284.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_11/Issue_
03/tiff/285.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/286.tif.

Hurst:1973:ACP

[Hur73]  S. L. Hurst.  The appli-

cation of Chow parameters and Rademacher-Walsh matrices in the synthesis of binary functions. The Computer Journal, 16(2):165–173, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/160165.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/165.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/166.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/167.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/168.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/169.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/170.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/171.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/172.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/173.tif.

Hurst:1979:ECS

[Hur79] S. L. Hurst. An engineering consideration of spectral transforms for ternary logic synthesis. The Computer Journal, 22(2):173–183, May 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/173.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/174.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/175.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/176.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/177.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/178.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/179.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/180.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/181.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/182.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_22/Issue_
02/tiff/183.tif.

Huskey:1961:CTA

[Hus61] H. D. Huskey. Compiling techniques for algebraic expressions. The Computer Journal, 4(1):10–19, April 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/040010.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/tiff/10.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_01/tiff/
11.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_04/Issue_01/tiff/12.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
04/Issue_01/tiff/13.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/tiff/14.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_01/tiff/
15.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_04/Issue_01/tiff/16.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
04/Issue_01/tiff/17.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/tiff/18.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_01/tiff/
19.tif.

Hutchinson:1984:HPD

[Hut84] Alan Hutchinson. The halting problem does not matter. The Computer Journal, 27(4):376, November 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        04/tiff/376.tif.

Hutchinson:1986:DSA

[Hut86a] A. Hutchinson. A data structure and algorithm for a self-augmenting heuristic program. The Computer Journal, 29(2):135–150, April 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        02/tiff/135.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        02/tiff/136.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        02/tiff/137.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        02/tiff/138.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        02/tiff/139.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        02/tiff/140.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        02/tiff/141.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_29/Issue_
02/tiff/142.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/143.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/144.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/145.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/146.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/147.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/148.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/149.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/150.tif.

Hutchinson:1986:SPP

[Hut86b] A. Hutchinson. Some practical principles for design of maintainable systems. The Computer Journal, 29(1):20– 23, February 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
29/Issue_01/tiff/20.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
01/tiff/21.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_01/tiff/
22.tif;   http://www3.oup.

                                        co.uk/computer_journal/hdb/
                                        Volume_29/Issue_01/tiff/23.
                                        tif.

Hutchinson:1989:IKS

[Hut89] A. Hutchinson. Inheritance and kinds of slots. The Computer Journal, 32(1):63–67, February 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        01/tiff/63.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_32/Issue_01/tiff/
                                        64.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_32/Issue_01/tiff/65.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        32/Issue_01/tiff/66.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        01/tiff/67.tif.

Hutchison:1993:SIE

[Hut93] David Hutchison. Special issue editorial: Distributed multimedia systems. The Computer Journal, 36(1):2–??, February 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_36/Issue_01/
                                        Vol36_01.body.html#AbstractHuthchison.

Huzan:1979:MIS

[Huz79] Eva Huzan. Modelling of indexed sequential files: Monitoring disc transfers. The Computer Journal, 22(1):22– 27, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
22/Issue_01/tiff/22.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
01/tiff/23.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_01/tiff/
24.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_22/Issue_01/tiff/25.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
22/Issue_01/tiff/26.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
01/tiff/27.tif.

Hathaway:1974:CEM

[HV74] P. J. Hathaway and D. Van Vliet. Correspondence: An efficient method for reading and writing subscripted arrays using FORTRAN IV. The Computer Journal, 17(2):190–191, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/190.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/191.tif. See corre-

spondence [Art75, ET75].

Haentjens:1978:LAT

[HV78] R. Haentjens and E. Van de Vel.

Lexical analysis and the transportation of programs. The Computer Journal, 21(2):190, May 1978. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        02/tiff/190.tif.

Hankins:1992:BRS

[HV92] Christopher Hankins and John Vince. Book reviews: The SGML Handbook, by Charles Goldfarb, 1990; Fractal Creations, by Timothy Wegner and Mark Peterson, 1991. The Computer Journal, 35(1):40, February 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_35/Issue_
                                        01/tiff/40.tif.

Haddon:1967:CPV

[HW67] B. K. Haddon and W. M. Waite. A compaction procedure for variable-length storage elements. The Computer Journal, 10(2):162–165, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        02/100162.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        02/tiff/162.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        02/tiff/163.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        02/tiff/164.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_10/Issue_
02/tiff/165.tif.

Hill:1974:NAC

[HW74] I. D. Hill and R. W. M.

Wedderburn. Note on Algorithm 78: Counting preferential votes in multi-member constituencies using absolute majority criteria. The Computer Journal, 17(4):380, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/380.tif. See [Van73,

Par74].

Haseman:1977:AAP

[HW77] W. D. Haseman and A. B.

Whinston. Automatic application program interface to a data base. The Computer Journal, 20(3):222–226, August 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/222.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/223.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/224.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/225.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/226.tif.

Henschen:1988:TIP

[HW88] L. J. Henschen and B. P. Weems. A top-down inference procedure for template dependencies. The Computer Journal, 31(1):51–55, February 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
01/tiff/51.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_01/tiff/
52.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_31/Issue_01/tiff/53.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
31/Issue_01/tiff/54.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
01/tiff/55.tif.

Hopkins:1989:WCO

[HW89] T. P. Hopkins and M. I. Wolczko. Writing concurrent object-oriented programs using Smalltalk-80. The Computer Journal, 32(4):341–350, August 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/341.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/342.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/343.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_32/Issue_
04/tiff/344.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/345.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/346.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/347.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/348.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/349.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/350.tif.

Hunter:1990:NOL

[HW90] A. Hunter and P. J. Willis. A note on the optimal labelling of quadtree nodes. The Computer Journal, 33(5):398–401, October 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/398.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/399.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/400.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/401.tif.

Harper:1992:EEC

[HW92] David J. Harper and Andrew D. M. Walker. ECLAIR: an extensible class library for information retrieval. The Computer Journal, 35(3):256–267, June 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_35/Issue_03/
                                        Vol35_03.body.html#AbstractHarper.

Hill:1987:STV

[HWW87] I. D. Hill, B. A. Wichmann, and D. R. Woodall. Algorithm 123: Single transferable vote by Meek’s method. The Computer Journal, 30(3):277–281, June 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        03/tiff/277.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        03/tiff/278.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        03/tiff/279.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        03/tiff/280.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        03/tiff/281.tif.

Harrison:1981:ENR

[HY81] M. A. Harrison and A. Yehudai. Eliminating null rules in linear time. The Computer Journal, 24(2):156–161, May 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/156.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/157.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/158.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/159.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/160.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/161.tif.

Hyslop:1972:NAO

[Hys72] J. Hyslop. A note on the accuracy of optimisation techniques. The Computer Journal, 15(2):140, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/150140.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/140.tif.

He:1995:CSO

[HZ95] Weidong He and Chaochen Zhou. A case study of optimization. The Computer Journal, 38(9):734–746, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

                                        journal/Volume_38/Issue_09/
                                        Vol38_09.body.html#AbstractHe.

Ingram:1996:ALT

[IB96] R. Ingram and S. Benford. The application of legibility techniques to enhance information visualizations. The Computer Journal, 39(10):819–??, ???? 1996. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www.oup.co.uk/computer_
                                        journal/Volume_39/Issue_10/
                                        Vol39_10.body.html#AbstractIngram;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_39/Issue_10/
                                        Vol39_10.body.html#AbstractIngram.

Ibbett:1972:MIP

[Ibb72] R. N. Ibbett. The MU5 instruction pipeline. The Computer Journal, 15(1):42–50, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        01/150042.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        01/tiff/42.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_15/Issue_01/tiff/
                                        43.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_15/Issue_01/tiff/44.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        15/Issue_01/tiff/45.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        01/tiff/46.tif;   http://

www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_01/tiff/
47.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_15/Issue_01/tiff/48.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
15/Issue_01/tiff/49.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/tiff/50.tif.

Inglis:1973:FBL

[ID73] J. Inglis and E. G. Dee. Flexibility of block-length for magnetic files. The Computer Journal, 16(4):303–307, November 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/160303.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/303.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/304.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/305.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/306.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/307.tif.

Ibbett:1985:CHP

[IEH^{+}85] R. N. Ibbett, D. A. Edwards, T. P. Hopkins, C. K. Cadogan, and D. A. Train. Centrenet — a high performance local area network. The Computer Journal, 28(3):231–242, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/231.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/232.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/233.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/234.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/235.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/236.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/237.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/238.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/239.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/240.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/241.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/242.tif.

Igarashi:1980:SAP

[Iga80] Y. Igarashi. The size of arrays for a prime implicant generating algorithm. The Computer Journal, 23(1):73–77, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
01/tiff/73.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/
74.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_23/Issue_01/tiff/75.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
23/Issue_01/tiff/76.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
01/tiff/77.tif.

Ibbett:1977:MNS

[IH77] R. N. Ibbett and M. A.

Husband. The MU5 name store. The Computer Journal, 20(3):227–231, August 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/227.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/228.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/229.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/230.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/231.tif.

Ince:1986:EER

[IH86] D. C. Ince and S. Hekmatpour. Empirical evaluation of random testing. The Computer Journal, 29(4):380, August 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        04/tiff/380.tif.

Ibbett:1995:HFT

[IHH95] R. N. Ibbett, P. E. Heywood, and F. W. Howell. HASE: A flexible toolset for computer architects. The Computer Journal, 38(10):755–764, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_38/Issue_10/
                                        Vol38_10.body.html#AbstractIbbett.

Iliffe:1962:DSA

[IJ62] J. K. Iliffe and J. G. Jodeit. A dynamic storage allocation scheme. The Computer Journal, 5(3):200–209, October 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        03/050200.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        03/tiff/200.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        03/tiff/201.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_05/Issue_
03/tiff/202.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/203.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/204.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/205.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/206.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/207.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/208.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/209.tif.

Inglis:1974:BRA

[IL74] J. Inglis and S. H. Lavington.

Book reviews: The Analysis of Information Systems, by C. T. Meadow, 1973; Introduction to Digital Computer Design, by D. J. Wooloms, 1973. The Computer Journal, 17(2):192, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/192.tif.

Iliffe:1969:EBB

[Ili69]  J. K. Iliffe.  Elements of

BLM (Basic Language Machine). The Computer Journal, 12(3):251–258, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/120251.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/251.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/252.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/253.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/254.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/255.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/256.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/257.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/258.tif.

Ierotheou:1998:SDM

[ILPP98] C. Ierotheou, C. H. Lai, C. J. Palansuriya, and K. A. Pericleous. Simulation of 2-D metal cutting by means of a distributed algorithm. The Computer Journal, 41(1):57–63, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_41/Issue_01/
Vol41_01.body.html#AbstractIerotheou.

Imbert:1993:VED

[Imb93]  Jean-Louis Imbert.  Vari-

able elimination for disequations in generalized linear constraint systems. The Computer Journal, 36(5):473–484, October 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_05/
Vol36_05.body.html#AbstractImbert.

Ince:1987:AGT

[Inc87]  D. C. Ince.  The auto-

matic generation of test data. The Computer Journal, 30 (1):63–69, February 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
01/tiff/63.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_01/tiff/
64.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_30/Issue_01/tiff/65.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
30/Issue_01/tiff/66.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
01/tiff/67.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_01/tiff/
68.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_30/Issue_01/tiff/69.
tif.

Inglis:1974:IIM

[Ing74] J. Inglis. Inverted indexes and multi-list structures. The Computer Journal, 17(1):59– 63, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        17/Issue_01/tiff/59.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        01/tiff/60.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_17/Issue_01/tiff/
                                        61.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_17/Issue_01/tiff/62.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        17/Issue_01/tiff/63.tif.

Inglis:1976:SPI

[Ing76]  J. Inglis.  Structured pro-

gramming and input statements. The Computer Journal, 19(2):188–189, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/188.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/189.tif. See corre-

spondence [DI78, Hop78].

Iyer:1986:PSC

[IP86] K. Viswanathan Iyer and L. M. Patnaik. Performance study of a centralized concurrency control algorithm for distributed database systems using Simula. The Computer Journal, 29(2):118–126, April 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/118.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/119.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/120.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/121.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/122.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/123.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/124.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/125.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/126.tif.

Ireland:1993:ESM

[Ire93] Andrew Ireland. On exploiting the structure of Martin-Lof’s Theory of Types. The Computer Journal, 36(4):387–398, August 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

                                        journal/Volume_36/Issue_04/
                                        Vol36_04.body.html#AbstractIreland.

Irving:1984:PBL

[Irv84] Robert W. Irving. Permutation backtracking in lexicographic order. The Computer Journal, 27(4):373–375, November 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        04/tiff/373.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        04/tiff/374.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        04/tiff/375.tif.

Izatt:1981:DSD

[IS81] W. T. Izatt and E. A. Schmitz. Data structures and descriptors in the ICL 2900 series and beyond. The Computer Journal, 24(4):301–307, November 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        04/tiff/301.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        04/tiff/302.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        04/tiff/303.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        04/tiff/304.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_

04/tiff/305.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/306.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/307.tif.

Ince:1986:RGC

[IW86] D. Ince and M. Woodman.

The rapid generation of a class of software tools. The Computer Journal, 29(2):151–160, April 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/151.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/152.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/153.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/154.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/155.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/156.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/157.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/158.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/159.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_29/Issue_
                                        02/tiff/160.tif.

Ibbett:1973:CCS

[IWT73] R. N. Ibbett, E. T. Warburton, and The Editor. Correspondence: The computing speed of a new machine. The Computer Journal, 16(3):287, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        03/tiff/287.tif.

Incerti:1981:AFS

[IZP81] S. Incerti, F. Zirilli, and V. Parisi. Algorithm 111: A FORTRAN subroutine for solving systems of nonlinear simultaneous equations. The Computer Journal, 24 (1):87–91, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        01/tiff/87.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_24/Issue_01/tiff/
                                        88.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_24/Issue_01/tiff/89.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        24/Issue_01/tiff/90.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        01/tiff/91.tif.

Jackson:1971:PCC

[Jac71] P. A. Jackson. PRBS crosscorrelation measurements by hybrid computational techniques. The Computer Journal, 14(1):49–54, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/140049.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/tiff/49.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_01/tiff/
50.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_14/Issue_01/tiff/51.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
14/Issue_01/tiff/52.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/tiff/53.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_01/tiff/
54.tif.

Jackson:1973:CFA

[Jac73]  D. M. Jackson.  Closed

form approximations for random errors in distance functions. The Computer Journal, 16(1):52–56, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
01/160052.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
01/tiff/52.tif;   http://
www3.oup.co.uk/computer_journal/

                                        hdb/Volume_16/Issue_01/tiff/
                                        53.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_16/Issue_01/tiff/54.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        16/Issue_01/tiff/55.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        01/tiff/56.tif.

Jacobs:1979:CVS

[Jac79] David Jacobs. Correspondence: Validating software. The Computer Journal, 22(3):287, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        03/tiff/287.tif. See [Dun78].

Jajodia:1986:RMD

[Jaj86] S. Jajodia. Recognizing multivalued dependencies in relation schemas. The Computer Journal, 29(5):458–459, October 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        05/tiff/458.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        05/tiff/459.tif.

Jakobsson:1985:SRL

[Jak85] Matti Jakobsson. Sampling without replacement in linear time. The Computer Journal, 28(4):412–413, August 1985. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/412.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/413.tif.

James:1980:UI

[Jam80] E. B. James. The user interface. The Computer Journal, 23(1):25–28, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
01/tiff/25.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/
26.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_23/Issue_01/tiff/28.
tif.

James:1981:MIC

[Jam81]  E. B. James.  Microelec-

tronics: the impact on computer centres. The Computer Journal, 24(1):9–13, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
01/tiff/10.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_01/tiff/
11.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_24/Issue_01/tiff/12.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_

                                        24/Issue_01/tiff/13.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        01/tiff/9.tif.

Jamieson:1987:NCI

[Jam87]  M. J. Jamieson.  A note

on the convergence of an iterative scheme for solving a quadratic equation. The Computer Journal, 30(2):189–190, April 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        02/tiff/189.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        02/tiff/190.tif.

Jamieson:1989:RCI

[Jam89]  M. J. Jamieson.  Rapidly

converging iterative formulae for finding square roots and their computational efficiencies. The Computer Journal, 32(1):93–94, February 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        01/tiff/93.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_32/Issue_01/tiff/
                                        94.tif.

Janes:1963:MPC

[Jan63] J. D. W. Janes. Measuring the profitability of a computer system. The Computer Journal, 5(4):284–293, January 1963. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/050284.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/284.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/285.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/286.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/287.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/288.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/289.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/290.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/291.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/292.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/293.tif.

Jankowitz:1988:DPS

[Jan88]  H. T. Jankowitz.  Detect-

ing plagiarism in student Pascal programs. The Computer Journal, 31(1):1–8, February 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        01/tiff/1.tif;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_31/Issue_01/tiff/
                                        2.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_31/Issue_01/tiff/3.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        31/Issue_01/tiff/4.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        01/tiff/5.tif;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_31/Issue_01/tiff/
                                        6.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_31/Issue_01/tiff/7.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        31/Issue_01/tiff/8.tif.

Jarratt:1966:MIM

[Jar66a] P. Jarratt. Multipoint iterative methods for solving certain equations. The Computer Journal, 8(4):398–400, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        04/080398.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        04/tiff/398.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        04/tiff/399.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        04/tiff/400.tif.

Jarratt:1966:RIF

[Jar66b] P. Jarratt. A rational iteration function for solving equations. The Computer Journal, 9(3):304–307, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/090304.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/304.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/305.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/306.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/307.tif.

Jarratt:1967:IML

[Jar67a] P. Jarratt. An iterative method for locating turning points. The Computer Journal, 10(1):82–84, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/100082.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/82.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_01/tiff/
83.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_10/Issue_01/tiff/84.
tif.

Jarratt:1967:NAE

[Jar67b] P. Jarratt. A note on the asymptotic error constant of a certain method for solving equations. The Computer Journal, 9(4):408–409, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        04/090408.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        04/tiff/408.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        04/tiff/409.tif.

Jarratt:1968:UCS

[Jar68] P. Jarratt. The use of comparison series in analysing iteration functions. The Computer Journal, 11(3):314–316, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        03/110314.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        03/tiff/314.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        03/tiff/315.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        03/tiff/316.tif.

Jardine:1970:DCA

[Jar70] N. Jardine. Discussion and correspondence: Algorithms, methods and models in the simplification of complex data. The Computer Journal, 13 (1):116–117, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/130116.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/tiff/116.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/tiff/117.tif.

Jarvis:1975:MFM

[Jar75]  J. E. Jarvis.  The many

faces of Multics. The Computer Journal, 18(1):2–6, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
01/tiff/2.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_18/Issue_01/tiff/
3.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_18/Issue_01/tiff/4.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
18/Issue_01/tiff/5.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
01/tiff/6.tif.

Jackson:1994:DUP

[JB94] W. Ken Jackson and F. Warren Burton. A definite and unfoldable, partially deterministic language. The Computer Journal, 37(8):711–714, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_37/Issue_08/
Vol37_08.body.html#AbstractJackson.

Jorre:1973:DCUa

[JBD73a] R. Jorre, G. Boyfield, and J. Darkin. Data communication user research in the Post Office: Part 1. The Computer Journal, 16(3):196–204, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/160196.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/196.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/197.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/198.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/199.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/200.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/201.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/202.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_

03/tiff/203.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/204.tif.

Jorre:1973:DCUb

[JBD73b] R. Jorre, G. Boyfield, and J. Darkin. Data communication user research in the Post Office: Part 2. The Computer Journal, 16(4):292–302, November 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/160292.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/292.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/293.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/294.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/295.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/296.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/297.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/298.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/299.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/300.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        04/tiff/301.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        04/tiff/302.tif.

Jones:1984:ARI

[JBJ84] B. Jones, M. Banerjee, and L. Jones. Algorithm 118: Root isolation for transcendental equations. The Computer Journal, 27(2):184–187, May 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        02/tiff/184.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        02/tiff/185.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        02/tiff/186.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        02/tiff/187.tif.

Jagodzinski:1986:RMM

[JC86] A. P. Jagodzinski and D. D. Clarke. A review of methods for measuring and describing users’ attitudes as an essential constituent of systems analysis and design. The Computer Journal, 29(2):97–102, April 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        02/tiff/100.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_29/Issue_
02/tiff/101.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/102.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/97.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_02/tiff/
98.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_29/Issue_02/tiff/99.
tif.

Jagodzinski:1988:MAM

[JC88] A. P. Jagodzinski and D. D.

Clarke. A multidimensional approach to the measurement of human-computer performance. The Computer Journal, 31(5):409–419, October 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/409.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/410.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/411.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/412.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/413.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/414.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_31/Issue_
                                        05/tiff/415.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        05/tiff/416.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        05/tiff/417.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        05/tiff/418.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        05/tiff/419.tif.

Ju:1997:USO

[JC97] Jialin Ju and Vipin Chaudhary. Unique sets oriented parallelization of loops with non-uniform dependences. The Computer Journal, 40(6):322–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www.oup.co.uk/computer_
                                        journal/Volume_40/Issue_06/
                                        Vol40_06.body.html#AbstractJu;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_40/Issue_06/
                                        Vol40_06.body.html#AbstractJu.

Jia:1999:MPB

[JCCJ99] Weijia Jia, Jiannong Cao, To-Yat Cheung, and Xiaohua Jia. A multicast protocol based on a single logical ring using a virtual token and logical clocks. The Computer Journal, 42(3):202–220, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_

03/420202.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
03/pdf/420202.pdf.

Jones:1959:CSM

[JD59] R. W. Jones and A. S. Douglas. Correspondence: Sorting by merging. The Computer Journal, 2(2):95–96, July 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
02/tiff/95.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_02/tiff/
96.tif.

Jelinek:1979:ATP

[Jel79] J. Jelínek. An algebraic theory for parallel processor design. The Computer Journal, 22(4):363–375, November 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/363.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/364.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/365.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/366.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/367.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_22/Issue_
04/tiff/368.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/369.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/370.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/371.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/372.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/373.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/374.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/375.tif.

Jennings:1961:DOR

[Jen61] G. Jennings. A direct ordering, recording and invoicing system. The Computer Journal, 4(2):150–156, July 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/040150.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/150.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/151.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/152.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/153.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/154.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/155.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/156.tif.

Jennings:1966:CSS

[Jen66] A. Jennings. A compact storage scheme for the solution of symmetric linear simultaneous equations. The Computer Journal, 9(3):281–285, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/090281.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/281.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/282.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/283.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/284.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/285.tif.

Jennings:1971:ASV

[Jen71] A. Jennings. Algorithm 70:

Solution of variable bandwidth positive definite simultaneous equations. The Computer Journal, 14(4):446, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/446.tif.

Jennings:1999:DAF

[Jen99] Esther Jennings. Distributed algorithms for finding central paths in tree networks. The Computer Journal, 42(7):609– 612, 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        07/420609.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        07/pdf/420609.pdf.

Jones:1998:SUG

[JES98] B. F. Jones, D. E. Eyres, and H. H. Sthamer. A strategy for using genetic algorithms to automate branch and fault-based testing. The Computer Journal, 41(2):98–107, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_41/Issue_02/
                                        Vol41_02.body.html#AbstractJones2.

Jesshope:1993:MNC

[JI93] C. R. Jesshope and C. Izu. The MP1 network chip and its application to parallel computers. The Computer Journal, 36(8):763–777, ???? 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_36/Issue_08/
Vol36_08.body.html#AbstractJesshope.

Jones:1967:CAC

[JJ67] K. Sparck Jones and D. Jackson. Current approaches to classification and clump-finding at the Cambridge Language Research Unit. The Computer Journal, 10(1):29–37, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/100029.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/29.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_01/tiff/
30.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_10/Issue_01/tiff/31.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/32.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/33.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_01/tiff/
34.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_10/Issue_01/tiff/35.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_

                                        10/Issue_01/tiff/36.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        01/tiff/37.tif.

Jarosz:1986:CTP

[JJ86] J. Jarosz and J. R. Jaworowski. Computer tree — the power of parallel computations. The Computer Journal, 29(2):103–108, April 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        02/tiff/103.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        02/tiff/104.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        02/tiff/105.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        02/tiff/106.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        02/tiff/107.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        02/tiff/108.tif.

Jones:1980:ASD

[JK80] W. T. Jones and S. A. Kirk. APL as a software design specification language. The Computer Journal, 23(3):230–232, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_

03/tiff/230.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/231.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/232.tif.

Jones:1994:ITI

[JKW94] Christopher B. Jones, David B.

Kidner, and J. Mark Ware. The implicit triangulated irregular network and multiscale spatial databases. The Computer Journal, 37(1):43–57, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_37/Issue_01/
Vol37_01.body.html#AbstractJones.

Johnson:1980:CUD

[JL80] S. N. Johnson and Jay Louise.

Correspondence: Using data abstraction for logical design. The Computer Journal, 23(4):380, November 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/380.tif. See [Wel80,

Ree80b].

Jiang:1999:NAI

[JLV99] Tao Jiang, Ming Li, and Paul Vitányi. New applications of the incompressibility method. The Computer Journal, 42(4):287–293, ???? 1999. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
04/420287.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
04/pdf/420287.pdf.

Jarratt:1968:LSM

[JM68] P. Jarratt and C. Mack. A least squares method for Laplace’s equation with Dirichlet boundary conditions. The Computer Journal, 11(1):83–87, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/110083.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/83.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_01/tiff/
84.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_11/Issue_01/tiff/85.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/86.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/87.tif.

Jackson:1996:PPF

[JM96] David Jeff Jackson and Wagdy Mahmoud. Parallel pipelined fractal image compression using quadtree recomposition. The Computer Journal, 39(1):1–13, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www.oup.co.uk/jnls/
list/comjnl/hdb/Volume_39/
Issue_01/390001.sgm.abs.html;
 http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_01/
Vol39_01.body.html#AbstractJackson.

Jarratt:1965:URF

[JN65] P. Jarratt and D. Nudds. The use of rational functions in the iterative solution of equations on a digital computer. The Computer Journal, 8(1):62–65, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
01/080062.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
01/tiff/62.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_01/tiff/
63.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_08/Issue_01/tiff/64.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
08/Issue_01/tiff/65.tif.

Jones:1974:PMR

[JO74] B. E. Jones and M. A. Ould.

The patient medical record as a database. The Computer Journal, 17(4):295–301, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/295.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_17/Issue_
                                        04/tiff/296.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        04/tiff/297.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        04/tiff/298.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        04/tiff/299.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        04/tiff/300.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        04/tiff/301.tif.

Jamison:1992:SNI

[JO92] B. Jamison and S. Olariu. Short notes: Indexing for multi-attribute retrieval. The Computer Journal, 35(5):535– 537, October 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/Volume_
                                        35/Issue_05/350535.sgm.abs.
                                        html;   http://www3.oup.co.
                                        uk/computer_journal/Volume_
                                        35/Issue_05/tiff/535.tif;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_35/Issue_05/
                                        tiff/536.tif;   http://www3.
                                        oup.co.uk/computer_journal/
                                        Volume_35/Issue_05/tiff/537.
                                        tif.

Johnson:1961:OTM

[Joh61] H. H. Johnson. Optimum time for multiplication on a digital computer. The Computer Journal, 3(4):256–261, January 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/030256.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/256.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/257.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/258.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/259.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/260.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/261.tif.

Johnson:1974:PSC

[Joh74] David S. Johnson. Worstcase behavior of graph coloring algorithms. In Proceedings 5th Southeastern Conference on Combinatorics, Graph Theory, and Computing, pages 513–

527. Utilitas Mathematica Publishing, Winnipeg, Manitoba, Canada, 1974. See [Woo69c].

Jokinen:1990:PTA

[Jok90] M. O. Jokinen. Parameter transmission abstractions. The Computer Journal, 33(2):133– 139, April 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
33/Issue_02/tiff/133.tif;
 http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/134.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/135.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/136.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/137.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/138.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/139.tif.

Jones:1965:SPC

[Jon65]  C. B. Jones.  A special-

purpose compiler. The Computer Journal, 8(2):109–112, July 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/080109.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/109.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/110.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/111.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/112.tif.

Jones:1966:NAM

[Jon66] E. L. Jones. Note on an alternate method for the computation of rotational energy levels of rigid asymmetric top molecules. The Computer Journal, 9(1):65–66, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/090065.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/65.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_01/tiff/

66.tif. See [Rac65] and correspondence [Duk66].

Jones:1970:SNA

[Jon70] A. Jones. Spiral — A new algorithm for non-linear parameter estimation using least squares. The Computer Journal, 13(3):301–308, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/130301.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/301.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/302.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/303.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/304.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/305.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/306.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/307.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/308.tif.

Jones:1971:CQW

[Jon71a]  A. J. Jones.  Correspon-

dence: On quality of writing. The Computer Journal, 14(4):337, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/140337.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/337.tif. See [All71].

Jones:1971:NAH

[Jon71b] C. B. Jones. A new approach to the ‘hidden line’ problem. The Computer Journal, 14(3):232–237, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/140232.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/tiff/232.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/233.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/234.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/235.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/236.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/237.tif.

Jones:1972:AMI

[Jon72]  H. S. P. Jones.  Algo-

rithm 72: Multiple integer arithmetic procedures in AL-GOL. The Computer Journal, 15(3):281–282, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/281.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/282.tif.

Jones:1979:NDP

[Jon79] Susan Jones. A network display program. The Computer Journal, 22(2):98–104, May 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/100.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_

                                        02/tiff/101.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/102.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/103.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/104.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/98.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_22/Issue_02/tiff/
                                        99.tif.

Jordan:1973:NLU

[Jor73] C. R. Jordan. A note on LISP universal S-functions. The Computer Journal, 16(2):124– 125, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        16/Issue_02/160124.sgm.abs.
                                        html;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        16/Issue_02/tiff/124.tif;
                                         http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/125.tif.

Joseph:1970:APP

[Jos70]  M. Joseph.  An analysis

of paging and program behaviour. The Computer Journal, 13(1):48–54, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_

01/130048.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/tiff/48.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_01/tiff/
49.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/50.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
13/Issue_01/tiff/51.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/tiff/52.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_01/tiff/
53.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/54.
tif.

Jackson:1969:DAC

[JP69] K. Jackson and J. R. Prior.

Debugging and assessment of control programs for an automatic radar. The Computer Journal, 12(4):303–306, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/120303.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/303.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/304.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/305.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/306.tif.

James:1976:TIC

[JP76] E. B. James and D. P. Partridge. Tolerance to inaccuracy in computer programs. The Computer Journal, 19(3):207–212, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/207.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/208.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/209.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/210.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/211.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/212.tif.

Joseph:1986:FRT

[JP86] M. Joseph and P. Pandya. Finding response times in a realtime system. The Computer Journal, 29(5):390–395, October 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        05/tiff/390.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_

05/tiff/391.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/392.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/393.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/394.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/395.tif.

Jones:1981:BRP

[JR81] S. Jones and C. M. Reeves.

Book reviews: Principles of Database Systems, by J. D. Ullman, 1980; Mathematical Methods in Computer Graphics and Design, edited by K. W. Brodie, 1980. The Computer Journal, 24(2):190, May 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/190.tif.

Jain:1967:NSL

[JS67] M. K. Jain and K. D. Sharma.

Numerical solution of linear differential equations and Volterra’s integral equation using Lobatto quadrature formula. The Computer Journal, 10(1):101–107, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/100101.sgm.abs.html;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/101.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/102.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/103.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/104.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/105.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/106.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/107.tif.

Jardine:1968:CHN

[JS68] N. Jardine and R. Sibson. The construction of hierarchic and non-hierarchic classifications. The Computer Journal, 11(2):177–184, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/110177.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/177.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/178.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/179.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_11/Issue_
02/tiff/180.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/181.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/182.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/183.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/184.tif.

Jardine:1971:CMA

[JS71] N. Jardine and R. Sibson.

Choice of methods for automatic classification. The Computer Journal, 14(4):404– 406, November 1971. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
14/Issue_04/140404.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
14/Issue_04/tiff/404.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/405.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/406.tif.

Jones:1989:FPO

[JS89] S. B. Jones and A. F. Sinclair.

Functional programming and operating systems. The Computer Journal, 32(2):162–174, April 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/162.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/163.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/164.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/165.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/166.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/167.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/168.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/169.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/170.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/171.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/172.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/173.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/174.tif.

Jambor-Sadeghi:1994:SAC

[JSKCG94] Kamyar Jambor-Sadeghi, Mohammad A. Ketabchi, Junjie Chue, and M. Ghiassi. A systematic approach to corrective maintenance. The Computer Journal, 37(9):764–778, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_37/Issue_09/
Vol37_09.body.html#AbstractSadeghi.

Jackson:1972:SPN

[JW72] D. M. Jackson and L. J.

White. Stability problems in non-statistical classification theory. The Computer Journal, 15(3):214–221, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/150214.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/214.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/215.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/216.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/217.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/218.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/219.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/220.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/221.tif.

Jackson:1974:NFR

[JW74] D. M. Jackson and W. W. Wadge. Normal form reduction of probabilistic computations in non-parametric classification. The Computer Journal, 17(3):249–252, August 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        03/tiff/249.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        03/tiff/250.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        03/tiff/251.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        03/tiff/252.tif.

Johnson:1982:UII

[JW82] J. S. Johnson and D. B. Webster. Updating an inverted index — a performance comparison of two techniques. The Computer Journal, 25(2):169–175, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/169.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/170.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_25/Issue_
02/tiff/171.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/172.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/173.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/174.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/175.tif.

Jones:1998:PST

[JW98] Christopher B. Jones and J. Mark Ware. Proximity search with a triangulated spatial model. The Computer Journal, 41(2):71–83, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_41/Issue_02/
Vol41_02.body.html#AbstractJones.

Kulkarni:1986:EEF

[KA86] K. G. Kulkarni and M. P.

Atkinson. EFDM: Extended functional data model. The Computer Journal, 29(1):38– 46, February 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
29/Issue_01/tiff/38.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
01/tiff/39.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_01/tiff/

                                        40.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_29/Issue_01/tiff/41.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        29/Issue_01/tiff/42.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        01/tiff/43.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_29/Issue_01/tiff/
                                        44.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_29/Issue_01/tiff/45.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        29/Issue_01/tiff/46.tif.

Kumar:1998:HRC

[KA98] Akhil Kumar and Noha Adly. HPP: A reliable causal broadcast protocol for large-scale replication in wide area networks. The Computer Journal, 41(2):108–124, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_41/Issue_02/
                                        Vol41_02.body.html#AbstractKumar.

Kacsuk:1991:PPA

[Kac91] P. Kacsuk. A parallel Prolog abstract machine and its multi-Transputer implementation. The Computer Journal, 34(1):52–63, February 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        01/tiff/52.tif;   http://

www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_01/tiff/
53.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_34/Issue_01/tiff/54.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
34/Issue_01/tiff/55.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
01/tiff/56.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_01/tiff/
57.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_34/Issue_01/tiff/58.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
34/Issue_01/tiff/59.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
01/tiff/60.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_01/tiff/
61.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_34/Issue_01/tiff/62.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
34/Issue_01/tiff/63.tif.

Kaiser:1972:JMP

[Kai72]  H. F. Kaiser.  The JK

method: a procedure for finding the eigenvectors and eigenvalues of a real symmetric matrix. The Computer Journal, 15(3):271–273, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/150271.sgm.abs.html;

                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/271.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/272.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/273.tif.

Kammerer:1977:ER

[Kam77] P. Kammerer. Excluding regions. The Computer Journal, 20(2):128–131, May 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        02/tiff/128.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        02/tiff/129.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        02/tiff/130.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        02/tiff/131.tif.

Karim:1966:SFO

[Kar66] A. I. Abdel Karim. Stability of the fourth order Runge– Kutta method for the solution of systems of differential equations. The Computer Journal, 9(3):308–311, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        03/090308.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_

journal/hdb/Volume_09/Issue_
03/tiff/308.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/309.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/310.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/311.tif.

Kaur:1992:HHP

[Kau92] D. Kaur. Hierarchical hypernet for pipelining. The Computer Journal, 35(1):81–87, February 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_35/Issue_01/
Vol35_01.body.html#AbstractKaur.

Kelly:1962:MIP

[KB62] D. H. Kelly and J. N. Buxton. MONTECODE — an interpretive program for Monte Carlo simulations. The Computer Journal, 5(2):88–93, July 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/050088.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/88.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_02/tiff/
89.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_05/Issue_02/tiff/90.
tif;   http://www3.oup.co.

uk/computer_journal/hdb/Volume_
05/Issue_02/tiff/91.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/92.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_02/tiff/
93.tif.

Kacsuk:1987:DPS

[KB87] P. Kacsuk and A. Bale. DAP Prolog: a set-oriented approach to Prolog. The Computer Journal, 30(5):393–403, October 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/393.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/394.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/395.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/396.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/397.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/398.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/399.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/400.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/401.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/402.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/403.tif.

Kunz:1997:PTS

[KBTB97] T. Kunz, J. P. Black, D. J.

Taylor, and T. Basten. Poet: Target-system independent visualizations of complex distributedapplication executions. The Computer Journal, 40(8):499– ??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/Volume_
40/Issue_08/Vol40_08.body.
html#AbstractKunz.

Kontos:1988:KAT

[KC88] John Kontos and J. C. Cavouras.

Knowledge acquisition from technical texts using attribute grammars. The Computer Journal, 31(6):525–530, December 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/525.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/526.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/527.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/528.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_31/Issue_
                                        06/tiff/529.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        06/tiff/530.tif.

Kilgore:1997:TDP

[KC97] Richard B. Kilgore and Craig M. Chase. Testing distributed programs containing racing messages. The Computer Journal, 40(8):489–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_40/Issue_08/
                                        Vol40_08.body.html#AbstractKilgore.

Kocberber:1999:OSF

[KCP99] S. Kocberber, F. Can, and J. M. Patton. Optimization of signature file parameters for databases with varying record lengths. The Computer Journal, 42(1):11–23, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        01/420011.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        01/pdf/420011.pdf.

Kecskes:1987:SMB

[Kec87] Cs. Kecskes. A simple method for building up large computer systems from small computers. The Computer Journal, 30(4):295–297, August 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/295.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/296.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/297.tif.

Keir:1979:CVA

[Kei79]  K. M. Keir.  Correspon-

dence: Vector approximation to curves. The Computer Journal, 22(4):381, November 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/381.tif.

Kelly:1998:ASL

[Kel98] Declan P. Kelly. ARC: A specification language for communicating systems. The Computer Journal, 41(6):386–401, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_41/Issue_06/
Vol41_06.body.html#AbstractKelly.

Kent:1980:MBH

[Ken80]  S. A. Kent.  A minicom-

puter based high level network control language. The Computer Journal, 23(2):124–131, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/124.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/125.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/126.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/127.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/128.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/129.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/130.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/131.tif.

Kent:1985:ENW

[Ken85] Philippe Kent. An efficient new way to represent multidimensional data. The Computer Journal, 28(2):184–190, May 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/184.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/185.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/186.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_28/Issue_
02/tiff/187.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/188.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/189.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/190.tif.

Kermode:1963:CDC

[Ker63] E. R. Kermode. Correspondence: On “Direct coding of English language names”. The Computer Journal, 6(3): 236, October 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
06/Issue_03/tiff/236.tif.

See [Bra63].

Kerridge:1976:PDS

[Ker76] J. M. Kerridge. PATCOSY — a database system for the National Health Service. The Computer Journal, 19(2):98–103, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/100.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/101.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/102.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_

02/tiff/103.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/98.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_02/tiff/
99.tif.

Kimura:1985:DDM

[KGC85] T. D. Kimura, W. D. Gillett, and J. R. Cox, Jr. A design of a data model based on abstraction of symbols. The Computer Journal, 28(3):298–308, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/298.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/299.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/300.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/301.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/302.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/303.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/304.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/305.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/306.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/307.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/308.tif.

Kollias:1981:CAS

[KH81] J. G. Kollias and M. Hatzopoulos. Criteria to aid in solving the problem of allocating copies of a file in a computer network. The Computer Journal, 24(1):29–30, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
01/tiff/29.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_01/tiff/
30.tif.

Klein:1987:CFD

[KH87] H. K. Klein and R. A.

Hirschheim. A comparative framework of data modelling paradigms and approaches. The Computer Journal, 30(1):8–15, February 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
01/tiff/10.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_01/tiff/
11.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_30/Issue_01/tiff/12.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_

                                        30/Issue_01/tiff/13.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        01/tiff/14.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_30/Issue_01/tiff/
                                        15.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_30/Issue_01/tiff/8.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        30/Issue_01/tiff/9.tif.

Khabaza:1963:ILS

[Kha63] I. M. Khabaza. An iterative least-square method suitable for solving large sparse matrices. The Computer Journal, 6(2):202–206, July 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/060202.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/tiff/202.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/tiff/203.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/tiff/204.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/tiff/205.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/tiff/206.tif.

Khanna:1991:LPS

[Kha91] S. Khanna. Logic programming for software verification and testing. The Computer Journal, 34(4):350–357, August 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/350.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/351.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/352.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/353.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/354.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/355.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/356.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/357.tif.

Khoong:1993:SPR

[Kho93]  C. M. Khoong.  Shortest-

path reconstruction algorithms. The Computer Journal, 36 (6):588–592, December 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_36/Issue_06/
Vol36_06.body.html#AbstractKhoong.

Kilburn:1961:MUA

[KHPS61] T. Kilburn, D. J. Howarth, R. B. Payne, and F. H. Sumner. The Manchester University Atlas operating system part I: internal organization. The Computer Journal, 4(3):222–225, October 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/040222.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/222.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/223.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/224.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/225.tif.

Kiessling:1988:APS

[Kie88] W. Kiessling. Access path selection in databases with intelligent disc subsystems. The Computer Journal, 31 (1):41–50, February 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
01/tiff/41.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_01/tiff/
42.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_31/Issue_01/tiff/43.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_

31/Issue_01/tiff/44.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
01/tiff/45.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_01/tiff/
46.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_31/Issue_01/tiff/47.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
31/Issue_01/tiff/48.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
01/tiff/49.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_01/tiff/
50.tif.

Kilgour:1981:BRI

[Kil81]  A. C. Kilgour.  Book re-

view: ILP: Intermediate Language for Pictures, by P. J. W. ten Hagen, T. Hagen, P. Klint, H. Noot, N. J. Sint, and A. H. Veen, 1981. The Computer Journal, 24(3):257, August 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/257.tif.

Kingsbury:1962:OEC

[Kin62] M. A. Kingsbury. Operating experience with COBOL in a service bureau. The Computer Journal, 5(3):157–158, October 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_

03/050157.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/157.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/158.tif.

King:1967:DT

[Kin67a] P. J. H. King. Decision tables. The Computer Journal, 10(2):135–142, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/100135.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/135.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/136.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/137.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/138.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/139.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/140.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/141.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/142.tif.

King:1967:DCS

[Kin67b] P. J. H. King. Discussion and correspondence: some comments on systematics. The Computer Journal, 10(1):116– 119, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
10/Issue_01/100116.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/116.tif;
 http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/117.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/118.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/119.tif.

King:1969:ILE

[Kin69a] P. J. H. King. The interpretation of limited entry decision table format and relationships among conditions. The Computer Journal, 12 (4):320–326, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/120320.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/320.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/321.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/322.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/323.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/324.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/325.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/326.tif. See corre-

spondence [Car70, Kin71].

King:1969:SAD

[Kin69b] P. J. H. King. Systems analysis documentation: Computeraided data dictionary definition. The Computer Journal, 12(1):6–9, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        01/120006.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        01/tiff/6.tif;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_12/Issue_01/tiff/
                                        7.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_12/Issue_01/tiff/8.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        12/Issue_01/tiff/9.tif.

King:1971:CIL

[Kin71] P. J. H. King. Correspondence: Interpretation of limited entry decision table format. The Computer Journal, 14(1):54, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/140054.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/tiff/54.tif. See [Car70,

Kin69a].

Kiossev:1981:CPP

[Kio81]  Nickolia Kiossev.  Corre-

spondence: Points and polygons. The Computer Journal, 24(1):94–95, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
01/tiff/94.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_01/tiff/

95.tif. See [dS80].

Kitney:1974:BRC

[Kit74]  R. I. Kitney.  Book re-

view: Computing Principles and Techniques, by B. L. Vickery, 1979. The Computer Journal, 17(4):383, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/383.tif.

Kizner:1966:ECL

[Kiz66]  W. Kizner.  Error curves

for Lanczos’ ‘selected points’ method. The Computer Journal, 8(4):372–382, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/080372.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/372.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/373.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/374.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/375.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/376.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/377.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/378.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/379.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/380.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/381.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_

04/tiff/382.tif. See com-

ment [Wri66a].

King:1975:CDT

[KJ75] P. J. H. King and R. G. Johnson. The conversion of decision tables to sequential testing procedures. The Computer Journal, 18(4):298–306, November 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/298.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/299.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/300.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/301.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/302.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/303.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/304.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/305.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/306.tif.

Kobayashi:1979:SMC

[KK79] Y. Kobayashi and T. L. Kunii. A simulation of a minicomputer controlled system and its use as a debugging tool. The Computer Journal, 22(1):41–49, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
01/tiff/41.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_01/tiff/
42.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_22/Issue_01/tiff/43.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
22/Issue_01/tiff/44.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
01/tiff/45.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_01/tiff/
46.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_22/Issue_01/tiff/47.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
22/Issue_01/tiff/48.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
01/tiff/49.tif.

Kinnersley:1994:EPT

[KK94] Nancy G. Kinnersley and William M. Kinnersley. An efficient polynomial-time algorithm for three-track gate matrix layout. The Computer Journal, 37(5):449–462, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

journal/Volume_37/Issue_05/
Vol37_05.body.html#AbstractKinnersley.

Kranakis:1996:MLL

[KKR96] Evangelos Kranakis, Danny Krizanc, and S. S. Ravi. On multi-label linear interval routing schemes. The Computer Journal, 39(2):133–139, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www.oup.co.uk/jnls/
list/comjnl/hdb/Volume_39/
Issue_02/390133.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_02/
Vol39_02.body.html#AbstractKranakis.

Kovacs:1971:NAM

[KL71] Zs. Kovács and Shirley A. Lill.

Note on Algorithm 46: A modified Davidon method for finding the minimum of a function using difference approximation for derivatives. The Computer Journal, 14(2):214–215, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/214.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/215.tif. See [Lil70,

Lil71].

Kaletzky:1977:POL

[KL77] A. Kaletzky and D. W. Lewin.

Problem orientated language for logic design. The Computer Journal, 20(2):163–169, May 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/163.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/164.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/165.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/166.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/167.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/168.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/169.tif.

Kafura:1989:IAB

[KL89] D. G. Kafura and K. H. Lee. Inheritance in actor based concurrent object-oriented languages. The Computer Journal, 32(4):297–304, August 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/297.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/298.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/299.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_32/Issue_
04/tiff/300.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/301.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/302.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/303.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/304.tif.

Kingdon:1991:HHD

[KLB91] H. Kingdon, D. R. Lester, and G. L. Burn. The HDGmachine: a highly distributed graph-reducer for a Transputer network. The Computer Journal, 34(4):290–301, August 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/290.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/291.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/292.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/293.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/294.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/295.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_34/Issue_
                                        04/tiff/296.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        04/tiff/297.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        04/tiff/298.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        04/tiff/299.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        04/tiff/300.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        04/tiff/301.tif.

Klein:1997:EOR

[Kle97] S. T. Klein. Efficient optimal recompression. The Computer Journal, 40(2/3):117–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www.oup.co.uk/computer_
                                        journal/Volume_40/Issue_02/
                                        Vol40_02.body.html#AbstractKlein;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_40/Issue_02/
                                        Vol40_03.body.html#AbstractKlein.

Kline:1975:CES

[Kli75]  John V. Kline.  Corre-

spondence: On ETAOIN SHRDLU. The Computer Journal, 18(3):286, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/286.tif.

Klerer:1965:UOP

[KM65] M. Klerer and J. May. A user oriented programming language. The Computer Journal, 8(2):103–109, July 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/080103.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/103.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/104.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/105.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/106.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/107.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/108.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/109.tif.

Keast:1966:ICN

[KM66] P. Keast and A. R. Mitchell.

On the instability of the Crank Nicholson formula under derivative boundary conditions. The Computer Journal, 9(1):110–114, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/090110.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/110.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/111.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/112.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/113.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/114.tif. See [PC64]

and correspondence [dD66].

Kardasz:1974:SSO

[KM74] J. H. Kardasz and G. Molnar. A SIMULA-based structure oriented language for the dynamic simulation of chemical plants. The Computer Journal, 17(1):28–37, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
01/tiff/28.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/
29.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_17/Issue_01/tiff/30.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
17/Issue_01/tiff/31.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_

01/tiff/32.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/
33.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_17/Issue_01/tiff/34.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
17/Issue_01/tiff/35.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
01/tiff/36.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/
37.tif.

Knight:1994:TDM

[KM94] Brian Knight and Jixin Ma.

A temporal database model supporting relative and absolute time. The Computer Journal, 37(7):588–597, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_37/Issue_07/
Vol37_07.body.html#AbstractKnight.

Kirstein:1995:COP

[KMK95] Peter Kirstein and Goli Montasser-Kohsari. The C-ODA project: Experiences and tools. The Computer Journal, 38(8):670– 680, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/Volume_
38/Issue_08/Vol38_08.body.
html#AbstractKirstein.

Kerber:1994:CTD

[KMR94] Gerrit Kerber, Helmut Meitner, and Friedemann Reim. The Comandos toolset for distributed systems management. The Computer Journal, 37(6):509–518, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_37/Issue_06/
Vol37_06.body.html#AbstractKerber.

Knight:1968:ASS

[Kni68] D. C. Knight. An algorithm for scheduling storage on a non-paged computer. The Computer Journal, 11(1):17–21, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/110017.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/17.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_01/tiff/
18.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_11/Issue_01/tiff/19.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/20.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/21.tif.

Knight:1993:ILT

[Kni93] B. Knight. Information loss in temporal knowledge representations. The Computer Journal, 36(2):127–136, April 1993. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_02/
Vol36_02.body.html#AbstractKnight.

Knott:1975:HF

[Kno75]  Gary D. Knott.  Hashing

functions. The Computer Journal, 18(3):265–278, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/265.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/266.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/267.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/268.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/269.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/270.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/271.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/272.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/273.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/274.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/275.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/276.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/277.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/278.tif. Also pub-

lished in/as: Stanford University Report, 1975.

Knuth:1984:LP

[Knu84]  Donald E. Knuth.  Liter-

ate programming. The Computer Journal, 27(2):97–111, May 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        02/tiff/100.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        02/tiff/101.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        02/tiff/102.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        02/tiff/103.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        02/tiff/104.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        02/tiff/105.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        02/tiff/106.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/107.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/108.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/109.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/110.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/111.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/97.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_02/tiff/
98.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_27/Issue_02/tiff/99.
tif.

Kollias:1978:EST

[Kol78] J. G. Kollias. An estimate of seek time for batched searching of random or index sequential structured files. The Computer Journal, 21(2):132–133, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/132.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/133.tif.

Kollias:1983:EMR

[Kol83]  J. G. Kollias.  The effect

of missing, retroactive, erroneous and/or late data in designing accounting applications. The Computer Journal, 26(2):117–123, May 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/117.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/118.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/119.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/120.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/121.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/122.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/123.tif.

Komor:1974:NLF

[Kom74] T. Komor. A note on left factored languages. The Computer Journal, 17(3):242–244, August 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/242.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/243.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_17/Issue_
03/tiff/244.tif.

Korver:1994:TS

[Kor94] Henri P. Korver. A theory for simulators. The Computer Journal, 37(4):279–287, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_37/Issue_04/
Vol37_04.body.html#AbstractKorver.

King:1960:AHE

[KP60] P. F. King and D. A. Peel. An analysis of a hydro-electric system. The Computer Journal, 3(3):161–163, October 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/030161.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/161.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/162.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/163.tif.

Korn:1979:ADC

[KP79] G. A. Korn and O. A. Palusinski. Application of a digital computer simulation language to control hybrid computer Monte Carlo studies. The Computer Journal, 22(2):136–141, May 1979. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/136.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/137.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/138.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/139.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/140.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/141.tif.

Kim:1992:ODA

[KP92] M. Ho Kim and S. Pramanik. Optimizing database accesses for parallel processing of multikey range searches. The Computer Journal, 35 (1):45–51, February 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_35/Issue_01/
Vol35_01.body.html#AbstractKim.

Kumar:1997:SIE

[KP97] Mohan Kumar and L. M. Patnaik. Special issue editorial: Automatic loop parallelization. The Computer Journal, 40(6):301, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www.oup.co.uk/computer_
journal/Volume_40/Issue_06/
AbstractEditorial;   http:
//www3.oup.co.uk/computer_
journal/Volume_40/Issue_06/
Vol40_06.body.html#AbstractEditorial.

Kwai:1998:TBD

[KP98] Ding-Ming Kwai and Behrooz Parhami. Tight bounds on the diameter of Gaussian cubes. The Computer Journal, 41(1):52–56, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_41/Issue_01/
Vol41_01.body.html#AbstractKwai.

Kapelios:1995:CRI

[KPP^{+}95] V. Kapelios, G. Panagopoulou, G. Papamichail, S. Sirmakessis, and A. Tsakalidis. The ‘cross’ rectangle intersection problem. The Computer Journal, 38(3):227–235, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_38/Issue_03/
Vol38_03.body.html#AbstractKapelios.

Koziris:1996:OTE

[KPT96] Nectarios Koziris, George Papakonstantinou, and Panayotis Tsanakas. Optimal time and efficient space free scheduling for nested loops. The Computer Journal, 39(5):439–448, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www.oup.co.uk/jnls/
list/comjnl/hdb/Volume_39/
Issue_05/390439.sgm.abs.html;
 http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_05/
Vol39_05.body.html#AbstractKoziris.

Katajainen:1989:AAS

[KR89] J. Katajainen and T. Raita. An approximation algorithm for space-optimal encoding of a text. The Computer Journal, 32(3):228–237, June 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/228.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/229.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/230.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/231.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/232.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/233.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/234.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/235.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_32/Issue_
03/tiff/236.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/237.tif.

Kral:1971:SPS

[Kra71] J. Kral. Some properties of the scatter storage technique with linear probing. The Computer Journal, 14(2):145–149, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/140145.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/145.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/146.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/147.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/148.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/149.tif.

Kohavi:1973:MDE

[KRK73] Z. Kohavi, J. A. Rivierre, and I. Kohavi. Machine distinguishing experiments. The Computer Journal, 16(2):141–147, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/160141.sgm.abs.html;

                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/141.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/142.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/143.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/144.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/145.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/146.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/147.tif.

Kronsjo:1970:AF

[Kro70a] T. O. M. Kronsjö. Algorithm 57: Find. The Computer Journal, 13(4):425–426, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/425.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/426.tif.

Kronsjo:1970:AIPa

[Kro70b] T. O. M. Kronsjö. Algorithm 58: An illustrative primal simplex linear program. The Computer Journal, 13 (4):426–428, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/426.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/427.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/428.tif. See note

[Kro71].

Kronsjo:1970:AIPb

[Kro70c] T. O. M. Kronsjö. Algorithm 59: An illustrative primal simplex linear program using multipliers. The Computer Journal, 13(4):428–429, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/428.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/429.tif. See note

[Kro71].

Kronsjo:1970:AIPc

[Kro70d] T. O. M. Kronsjö. Algorithm 60: An illustrative primal simplex linear program using multipliers and single dimensional arrays. The Computer Journal, 13(4):429–430, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/429.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/430.tif. See note

[Kro71].

Kronsjo:1970:AIS

[Kro70e] T. O. M. Kronsjö. Algorithm 61: An illustrative selfdual parametric simplex linear program using multipliers. The Computer Journal, 13(4):430–437, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/430.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/431.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/432.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/433.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/434.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/435.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/436.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/437.tif. See [Kro71].

Kronsjo:1971:NA

[Kro71] T. O. M. Kronsjö. Note on Algorithms 58 to 61. The Computer Journal, 14(2):215, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/215.tif. See [Kro70b,

Kro70c, Kro70d, Kro70e].

Keedy:1979:ISS

[KRR79] J. L. Keedy, K. Ramamohanarao, and J. Rosenberg. On implementing semaphores with sets. The Computer Journal, 22(2):146–150, May 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/146.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/147.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/148.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/149.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/150.tif.

Keedy:1982:SRW

[KRR82] J. L. Keedy, J. Rosenberg, and K. Ramamohanarao. On synchronizing readers and writers with semaphores. The Computer Journal, 25(1):121– 125, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_

                                        25/Issue_01/tiff/121.tif;
                                         http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/122.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/123.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/124.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/125.tif.

Knibb:1971:SPP

[KS71] D. Knibb and R. E. Scraton. On the solution of parabolic partial differential equations in Chebyshev series. The Computer Journal, 14(4):428–432, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/140428.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/428.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/429.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/430.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/431.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/432.tif.

Kieburtz:1982:SAM

[KS82] Richard Kieburtz and Abraham Silberschatz. On the static access-control mechanism in Concurrent Pascal. The Computer Journal, 25(1):17–21, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/17.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/
18.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_25/Issue_01/tiff/19.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/20.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/21.tif.

Kant:1985:EPR

[KS85] K. Kant and A. Silberschatz.

Error propagation and recovery in concurrent environments. The Computer Journal, 28(5):466–473, November 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/466.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/467.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/468.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/469.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        05/tiff/470.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        05/tiff/471.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        05/tiff/472.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        05/tiff/473.tif.

King:1991:DDI

[KS91] P. King and C. Small. Default databases and incomplete information. The Computer Journal, 34(3):239–244, June 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        03/tiff/239.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        03/tiff/240.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        03/tiff/241.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        03/tiff/242.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        03/tiff/243.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        03/tiff/244.tif.

Kulkarni:1997:LEL

[KS97] Dattatraya Kulkarni and Micheal Stumm. Linear and extended linear transformations for shared-memory multiprocessors. The Computer Journal, 40(6):373–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www.oup.co.uk/computer_
journal/Volume_40/Issue_06/
Vol40_06.body.html#AbstractKulkarni;
http://www3.oup.co.uk/computer_
journal/Volume_40/Issue_06/
Vol40_06.body.html#AbstractKulkarni.

Kollias:1977:IPI

[KSD77] J. G. Kollias, P. M. Stocker, and P. A. Dearnley. Improving the performance of an intelligent data management system. The Computer Journal, 20(4):302–307, November 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/302.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/303.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/304.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/305.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/306.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/307.tif.

Kirby:1973:CLC

[KTW^{+}73] G. H. Kirby, The Editor, R. Wane, I. D. Hill, and Brian Meek. Correspondence: Letters to The Computer Journal. The Computer Journal, 16(2):189–190, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/189.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/190.tif. See [Wal71,

GT72, MT72].

Khalil:1974:LSN

[KU74] H. M. Khalil and D. L. Ulery. LINEAL: A system for numerical linear algebra. The Computer Journal, 17(3):267–274, August 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/267.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/268.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/269.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/270.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/271.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_17/Issue_
03/tiff/272.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/273.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/274.tif.

Kayan:1999:ERT

[KU99] Ersan Kayan and Özgür Ulusoy. An evaluation of real-time transaction management issues in mobile database systems. The Computer Journal, 42(6):501– 510, 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
06/420501.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
06/pdf/420501.pdf.

Kuhn:1992:TAE

[Kuh92] D. R. Kuhn. A technique for analyzing the effects of changes in formal specifications. The Computer Journal, 35(6):574–578, December 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_06/
Vol35_06.body.html#AbstractKuhn.

Kumar:1999:SIE

[Kum99] Mohan Kumar. Special issue editorial: Mobile computing. The Computer Journal, 42(6): 441, 1999. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
06/420441.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
06/pdf/420441.pdf.

Kershaw:1962:EDE

[KV62] D. Kershaw and S. Vajda. The economics of dumping from electronic computers. The Computer Journal, 4(4):346–348, January 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/040346.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/346.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/347.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/348.tif.

Kitz:1963:PSE

[KV63] B. Kitz and S. Vajda. A program to study the effect of random delays on the ability of trains to runs to a schedule. The Computer Journal, 6(2):121–128, July 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/060121.sgm.abs.html;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/121.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/122.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/123.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/124.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/125.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/126.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/127.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/128.tif.

Kalvin:1984:TPR

[KV84] Alan D. Kalvin and Yaakov L.

Varol. On two problems reducible to topological sorting. The Computer Journal, 27(2):176–177, May 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/176.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/177.tif.

Koster:1995:HST

[KV95] C. H. A. Koster and Th. P.

Van Der Weide. Hairy search trees. The Computer Journal, 38(8):691–694, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_38/Issue_08/
                                        Vol38_08.body.html#AbstractKoster.

Kumar:1995:HDB

[KVKP95] M. J. Kumar, S. Venkatesh, D. Kieronska, and L. M. Patnaik. Hierarchical directorybased shared memory architecture. The Computer Journal, 38(3):207–216, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_38/Issue_03/
                                        Vol38_03.body.html#AbstractKumar.

Kirby:1989:RRH

[KVW89] G. H. Kirby, M. Visvalingam, and P. Wade. Recognition and representation of a hierarchy of polygons with holes. The Computer Journal, 32 (6):554–562, December 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        06/tiff/554.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        06/tiff/555.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        06/tiff/556.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_32/Issue_
06/tiff/557.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/558.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/559.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/560.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/561.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/562.tif.

Kouvatsos:1984:OBS

[KW84] D. D. Kouvatsos and S. K.

Wong. On optimal blocking of sequential files. The Computer Journal, 27(4):321–327, November 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/321.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/322.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/323.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/324.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/325.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/326.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        04/tiff/327.tif.

Lahodynskyj:1991:SC

[Lah91] Lu Lahodynskyj. Structure clashes. The Computer Journal, 34(6):566–568, December 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        06/tiff/566.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        06/tiff/567.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        06/tiff/568.tif.

Lai:1994:DDD

[Lai94] C.-H. Lai. Diakoptics, domain decomposition and parallel computing. The Computer Journal, 37(10):840–846, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_37/Issue_10/
                                        Vol37_10.body.html#AbstractLai.

Lalmas:1999:MRR

[Lal99]  Mounia Lalmas.  A model

for representing and retrieving heterogeneous structured documents based on evidential reasoning. The Computer Journal, 42(7):547–568, 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
07/420547.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
07/pdf/420547.pdf.

Lambourn:1963:RAM

[Lam63] S. Lambourn. Resource allocation and multi-project scheduling (RAMPS) — a new tool in planning and control. The Computer Journal, 5(4):300–304, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/050300.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/300.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/301.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/302.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/303.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/304.tif.

Landin:1964:MEE

[Lan64] Peter J. Landin. The mechanical evaluation of expressions. The Computer Journal, 6(4):308–320, January 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/060308.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/308.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/309.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/310.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/311.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/312.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/313.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/314.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/315.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/316.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/317.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/318.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/319.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/320.tif.

Langham:1965:USC

[Lan65] E. J. Langham. The underground storage of compressed air for gas turbines: a dynamic study on an analogue computer. The Computer Journal, 8(3):216–224, October 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/080216.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/216.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/217.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/218.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/219.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/220.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/221.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/222.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/223.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/224.tif.

Lansdowne:1971:CUS

[Lan71] R. John Lansdowne. Correspondence: On “The use of symbol-state tables”. The Computer Journal, 14(2):155, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/155.tif. See [Day70].

Land:1976:ESG

[Lan76] F. F. Land. Evaluation of systems goals in determining a design strategy for a computer based information system. The Computer Journal, 19(4):290–294, November 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        04/tiff/290.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        04/tiff/291.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        04/tiff/292.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        04/tiff/293.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        04/tiff/294.tif.

Land:1982:T

[Lan82] F. F. Land. Tutorial. The Computer Journal, 25(2):283–285, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_

02/tiff/283.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/284.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/285.tif.

Land:1985:ITE

[Lan85] Frank Land. Is an information theory enough? The Computer Journal, 28(3):211–215, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/211.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/212.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/213.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/214.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/215.tif.

Larkin:1964:CGI

[Lar64] F. M. Larkin. A combined graphical and iterative approach to the problem of finding zeros of functions in the complex plane. The Computer Journal, 7(3):212–219, October 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/070212.sgm.abs.html;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/212.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/213.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/214.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/215.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/216.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/217.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/218.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/219.tif.

Larkin:1967:STR

[Lar67] F. M. Larkin. Some techniques for rational interpolation. The Computer Journal, 10(2):178–187, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/100178.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/178.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/179.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_

02/tiff/180.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/181.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/182.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/183.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/184.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/185.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/186.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/187.tif.

Larson:1982:EWC

[Lar82]  Per-Åke Larson.  Expected

worst-case performance of hash files. The Computer Journal, 25(3):347–352, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/347.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/348.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/349.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/350.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        03/tiff/351.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        03/tiff/352.tif.

Larson:1985:PAS

[Lar85] Per-Åke Larson. Performance analysis of a single-file version of linear hashing. The Computer Journal, 28(3):319–329, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        03/tiff/319.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        03/tiff/320.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        03/tiff/321.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        03/tiff/322.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        03/tiff/323.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        03/tiff/324.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        03/tiff/325.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        03/tiff/326.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        03/tiff/327.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_28/Issue_
03/tiff/328.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/329.tif.

Larsson:1995:EAF

[Lar95] Mats Larsson. An engineering approach to formal digital system design. The Computer Journal, 38(2):101–110, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_38/Issue_02/
Vol38_02.body.html#AbstractLarsson.

Laski:1968:SVA

[Las68] J. G. Laski. Segmentation and virtual address topology — an essay in virtual research. The Computer Journal, 11(1):35–40, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/110035.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/35.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_01/tiff/
36.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_11/Issue_01/tiff/37.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/38.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/39.tif;   http://
www3.oup.co.uk/computer_journal/

                                        hdb/Volume_11/Issue_01/tiff/
                                        40.tif.

Lauer:1975:DPD

[Lau75] H. C. Lauer. Discussion on Ph.D. thesis proposals in computing science. The Computer Journal, 18(3):279–281, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/279.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/280.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/281.tif. See note

[HL76b].

Lau:1991:TSS

[Lau91] K. K. Lau. Top-down synthesis of sorting algorithm. The Computer Journal, 34(6):A001– A007, December 1991. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Laver:1964:UVD

[Lav64] F. J. M. Laver. A user’s view of A.D.P. systems design. The Computer Journal, 7(2):93–94, July 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        02/070093.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        02/tiff/93.tif;   http://

www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_02/tiff/
94.tif.

Laver:1970:JRD

[Lav70]  F. J. M. Laver.  JANUS

— reflections on DATAFAIR

69. The Computer Journal, 13(1):2–6, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/130002.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/tiff/2.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_13/Issue_01/tiff/
3.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/4.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
13/Issue_01/tiff/5.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/tiff/6.tif.

Lawrie:1969:ILP

[Law69] N. L. Lawrie. An integer linear programming model of a school timetabling problem. The Computer Journal, 12(4):307–316, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/120307.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_

04/tiff/307.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/308.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/309.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/310.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/311.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/312.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/313.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/314.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/315.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/316.tif.

Layzell:1985:HMP

[Lay85] P. J. Layzell. The history of macro processors in programming language extensibility. The Computer Journal, 28(1):29–33, February 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
01/tiff/29.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_01/tiff/
30.tif;   http://www3.oup.

co.uk/computer_journal/hdb/
Volume_28/Issue_01/tiff/31.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
28/Issue_01/tiff/32.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
01/tiff/33.tif.

Lazos:1981:ICS

[Laz81] C. Lazos. Increasing computer system productivity software and hardware methods: a comparative study. The Computer Journal, 24(3):230–234, August 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/230.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/231.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/232.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/233.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/234.tif.

Lim:1976:PVN

[LB76] R. T. C. Lim and J. D.

Blackburn. On program verification in numerical methods. The Computer Journal, 19(3):272–273, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/272.tif. See [Smi72].

Luker:1986:PGG

[LB86] P. A. Luker and A. Burns. Program generators and generation software. The Computer Journal, 29(4):315–321, August 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        04/tiff/315.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        04/tiff/316.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        04/tiff/317.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        04/tiff/318.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        04/tiff/319.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        04/tiff/320.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        04/tiff/321.tif.

Lee:1991:RDS

[LB91] W. W. L. Lee and D. S. Bowers. A relational database simulator. The Computer Journal, 34(6):A061–A072, December 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Ling:1992:MMT

[LB92] D. H. O. Ling and D. A.

Bell. Modelling and managing time in database systems. The Computer Journal, 35(4):332–341, August 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_35/Issue_04/
Vol35_04.body.html#AbstractLing.

Lamming:1994:DHM

[LBC^{+}94] M. Lamming, P. Brown, K. Carter, M. Eldridge, M. Flynn, G. Louie, P. Robinson, and A. Sellen. The design of a human memory prosthesis. The Computer Journal, 37(3):153–163, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_
03/tiff/153.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_
03/tiff/154.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_
03/tiff/155.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_
03/tiff/156.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_
03/tiff/157.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_
03/tiff/158.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_37/Issue_
03/tiff/159.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_
03/tiff/160.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_
03/tiff/161.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_
03/tiff/162.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_
03/tiff/163.tif.

Lin:1990:DIS

[LC90] F.-. C. Lin and K. Chen. On the design of an integrated systolic array for solving simultaneous linear equations. The Computer Journal, 33(3):252–260, June 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/252.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/253.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/254.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/255.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/256.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/257.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_

03/tiff/258.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/259.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/260.tif.

Lin:1993:GRM

[LC93] Min Sheng Lin and Deng Jyi Chen. General reduction methods for the reliability analysis of distributed computing systems. The Computer Journal, 36(7):631–644, ???? 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_07/
Vol36_07.body.html#AbstractLin.

Low:1994:SAP

[LC94] Marie Rose Low and Bruce Christianson. Self authenticating proxies. The Computer Journal, 37(5):422–428, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_37/Issue_05/
Vol37_05.body.html#AbstractLow.

Leung:1995:EFP

[LC95] C. H. C. Leung and E. Currie. The effect of failures on the performance of longduration database transactions. The Computer Journal, 38(6):471–477, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_38/Issue_06/
                                        Vol38_06.body.html#AbstractLeung.

Lin:1998:DMF

[LC98] Shun-Shii Lin and Shan-Tai Chen. On detecting multiple faults in baseline interconnection networks. The Computer Journal, 41(4):254–269, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_41/Issue_04/
                                        Vol41_04.body.html#AbstractLin.

Leung:1999:NPP

[LC99] H. Leung and T. Y. Chen. A new perspective of the proportional sampling strategy. The Computer Journal, 42(8):693–698, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        08/420693.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        08/pdf/420693.pdf.

Lin:1999:RAD

[LCH99] Min-Sheng Lin, Deng-Jyi Chen, and Maw-Sheng Horng. The reliability analysis of distributed computing systems with imperfect nodes. The Computer Journal, 42(2):129–141, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
02/420129.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
02/pdf/420129.pdf.

Lin:1992:RCD

[LCL92] C. H. Lin, C. C. Chang, and R. C. T. Lee. A record-oriented cryptosystem for database sharing. The Computer Journal, 35(6):658–660, December 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Lucido:1981:SMV

[LCP81] Tony P. Lucido, Rahul Chattergy, and Udo W. Pooch. Survey of microprogram verification and validation methods. The Computer Journal, 24(2):139–142, May 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/139.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/140.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/141.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/142.tif.

LeRiche:1969:ACP

[Le 69] P. J. Le Riche. Algorithm 41: A curve plotting procedure. The Computer Journal, 12(3):291–292, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/291.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/292.tif.

leQuesne:1988:IOF

[Le 88] P. N. Le Quesne. Individual and organisational factors and the design of IPSEs. The Computer Journal, 31(5):391–397, October 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/391.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/392.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/393.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/394.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/395.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/396.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/397.tif.

Levary:1986:AIA

[LE86] R. R. Levary and W. D. Edwards. Analysing the impact of adding a new software system on main memory usage. The Computer Journal, 29(6):522–526, December 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/522.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/523.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/524.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/525.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/526.tif.

Leavenworth:1977:NDP

[Lea77] B. M. Leavenworth. Nonprocedural data processing. The Computer Journal, 20(1):6– 9, February 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
20/Issue_01/tiff/6.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
01/tiff/7.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_20/Issue_01/tiff/
8.tif;   http://www3.oup.
co.uk/computer_journal/hdb/

Volume_20/Issue_01/tiff/9.
tif.

Lea:1978:TCA

[Lea78] R. M. Lea. Text compression with an associative parallel processor. The Computer Journal, 21(1):45–56, February 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
01/tiff/45.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_01/tiff/
46.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_21/Issue_01/tiff/47.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
21/Issue_01/tiff/48.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
01/tiff/49.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_01/tiff/
50.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_21/Issue_01/tiff/51.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
21/Issue_01/tiff/52.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
01/tiff/53.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_01/tiff/
54.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_21/Issue_01/tiff/55.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
21/Issue_01/tiff/56.tif.

Lea:1986:VWA

[Lea86] R. M. Lea. VLSI and WSI associative string processors for cost-effective parallel processing. The Computer Journal, 29(6):486–494, December 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/486.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/487.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/488.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/489.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/490.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/491.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/492.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/493.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/494.tif.

Lee:1972:FDB

[Lee72]  J. A. N. Lee.  The for-

mal definition of the BA-SIC language. The Computer Journal, 15(1):37–41, February 1972. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        01/150037.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        01/tiff/37.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_15/Issue_01/tiff/
                                        38.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_15/Issue_01/tiff/39.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        15/Issue_01/tiff/40.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        01/tiff/41.tif. See corre-

spondence [Bar72a].

Leech:1974:CSA

[Lee74]  J. Leech.  Correspondence:

On the spelling of ALGO-RITHM. The Computer Journal, 17(4):383, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        04/tiff/383.tif. See [Mar74b,

Mul75].

Lee:1978:UCS

[Lee78a] B. S. Lee. Use of case studies in teaching data processing. The Computer Journal, 21(4):366–372, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_

journal/hdb/Volume_21/Issue_
04/tiff/366.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/367.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/368.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/369.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/370.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/371.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/372.tif.

Lee:1978:RRB

[Lee78b]  P. A. Lee.  A reconsider-

ation of the recovery block scheme. The Computer Journal, 21(4):306–310, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/306.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/307.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/308.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/309.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/310.tif.

Lee:1982:DNS

[Lee82] Graham Lee. DOPLs: A new style of programming. The Computer Journal, 25(2):176– 182, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        25/Issue_02/tiff/176.tif;
                                         http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/177.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/178.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/179.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/180.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/181.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/182.tif.

Lehman:1973:CUC

[Leh73] M. M. Lehman. Computer usage control. The Computer Journal, 16(2):106–110, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/160106.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/106.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_16/Issue_
02/tiff/107.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/108.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/109.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/110.tif.

Leighton:1962:CGS

[Lei62] C. C. Leighton. Control gear simulation for an automatic car park. The Computer Journal, 4(4):313–316, January 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/040313.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/313.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/314.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/315.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/316.tif.

Leith:1983:HSP

[Lei83]  Philip Leith.  Hierarchi-

cally structured production rules. The Computer Journal, 26(1):1–5, February 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
01/tiff/1.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_26/Issue_01/tiff/
2.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_26/Issue_01/tiff/3.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
26/Issue_01/tiff/4.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
01/tiff/5.tif.

Leith:1986:FEL

[Lei86] P. Leith. Fundamental errors in legal logic programming. The Computer Journal, 29(6):545–552, December 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/545.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/546.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/547.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/548.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/549.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/550.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/551.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/552.tif.

Leith:1987:PSF

[Lei87]  P. Leith.  A programmed,

skeleton formal specification method: the OUFDM. The Computer Journal, 30(4):337– 342, August 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
30/Issue_04/tiff/337.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/338.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/339.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/340.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/341.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/342.tif.

Lemos:1980:DCM

[Lem80] R. S. Lemos. Discussion and correspondence: Measuring students’ attitudes towards structured walk-throughs. The Computer Journal, 23(4):377– 380, November 1980. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
23/Issue_04/tiff/377.tif;
http://www3.oup.co.uk/computer_

journal/hdb/Volume_23/Issue_
04/tiff/378.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/379.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/380.tif.

Lehman:1963:CCL

[LEN63] M. Lehman, R. Eshed, and Z. Netter. The checking of computer logic by simulation on a computer. The Computer Journal, 6(2):154–162, July 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/060154.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/154.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/155.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/156.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/157.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/158.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/159.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/160.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_

02/tiff/161.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/162.tif.

Leppanen:1995:PWO

[Lep95] Ville Leppänen. Performance of work-optimal PRAM simulation algorithms on coated meshes. The Computer Journal, 38(10):801–810, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_38/Issue_10/
Vol38_10.body.html#AbstractLeppanen.

Lerner:1969:FMD

[Ler69] A. J. Lerner. Formal methods of diagnosis in engineering and medicine. The Computer Journal, 12(1):29–32, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
01/120029.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
01/tiff/29.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_01/tiff/
30.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_12/Issue_01/tiff/31.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/32.tif.

Lescanne:1983:BCA

[Les83a] Pierre Lescanne. Behavioural categoricity of abstract data type specifications. The Computer Journal, 26(4):289–292, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
26/Issue_04/tiff/289.tif;
 http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/290.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/291.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/292.tif.

Lesuisse:1983:SLD

[Les83b] R. Lesuisse. Some lessons drawn from the history of the binary search algorithm. The Computer Journal, 26(2):154–163, May 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/154.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/155.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/156.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/157.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/158.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/159.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/160.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/161.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/162.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/163.tif.

Leung:1982:IOP

[Leu82a] C. H. C. Leung. An improved optimal-fit procedure for dynamic storage allocation. The Computer Journal, 25(2):199–206, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/199.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/200.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/201.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/202.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/203.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/204.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/205.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/206.tif.

Leung:1982:SMP

[Leu82b] C. H. C. Leung. A simple model for the performance analysis of disc storage fragmentation. The Computer Journal, 25(2):193–198, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/193.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/194.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/195.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/196.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/197.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/198.tif.

Leung:1983:ADF

[Leu83a]  C. H. C. Leung.  Analy-

sis of disc fragmentation using Markov chains. The Computer Journal, 26(2):113–116, May 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        02/tiff/113.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_

02/tiff/114.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/115.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/116.tif.

Leung:1983:MDL

[Leu83b] C. H. C. Leung. A model for disc locality referencing. The Computer Journal, 26 (3):196–198, August 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/196.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/197.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/198.tif.

Levison:1965:SF

[Lev65] M. Levison. The siting of fragments. The Computer Journal, 7(4):275–277, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/070275.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/275.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/276.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_07/Issue_
                                        04/tiff/277.tif.

Levene:1996:MCI

[Lev96]  Mark Levene.  Maintain-

ing consistency of imprecise relations. The Computer Journal, 39(2):114–123, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://www.oup.co.uk/jnls/
                                        list/comjnl/hdb/Volume_39/
                                        Issue_02/390114.sgm.abs.html;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_39/Issue_02/
                                        Vol39_02.body.html#AbstractLevene.

Levin:1999:RMI

[Lev99] Leonid. A. Levin. Robust measures of information. The Computer Journal, 42(4):284–286, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        04/420284.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        04/pdf/420284.pdf.

Lewis:1963:TSL

[Lew63] J. W. Lewis. Time sharing on LEO III. The Computer Journal, 6(1):24–28, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        01/060024.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_

journal/hdb/Volume_06/Issue_
01/tiff/24.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/
25.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_06/Issue_01/tiff/26.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
06/Issue_01/tiff/27.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/tiff/28.tif.

Lewis:1965:IAC

[Lew65a] C. D. Lewis. Iterative analogue computation applied to inventory policy simulation. The Computer Journal, 8(2):130–135, July 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/080130.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/130.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/131.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/132.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/133.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/134.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/135.tif.

Lewis:1965:MLC

[Lew65b] J. W. Lewis. The management of a large commercial computer bureau. The Computer Journal, 7(4):255–261, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        04/070255.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        04/tiff/255.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        04/tiff/256.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        04/tiff/257.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        04/tiff/258.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        04/tiff/259.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        04/tiff/260.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        04/tiff/261.tif.

Lewin:1971:AAA

[Lew71a] D. W. Lewin. Advanced aspects of asynchronous logic design. The Computer Journal, 14(3):254–259, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_

03/140254.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/254.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/255.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/256.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/257.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/258.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/259.tif.

Lewis:1971:SAI

[Lew71b] R. P. W. Lewis. The selection of ADI iteration parameters by numerical experiment for the solution of Poisson’s equation over a circular area. The Computer Journal, 14(1):73–74, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/140073.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/tiff/73.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_01/tiff/
74.tif.

Lewis:1972:CIN

[Lew72] B. A. Lewis. Correspondence:

On “Improved numerical methods for Volterra integral equations of the first kind”. The Computer Journal, 15(2):137, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        02/150137.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        02/tiff/137.tif. See [AW71].

Lew:1984:PCD

[Lew84] Art Lew. Proof of correctness of decision table programs. The Computer Journal, 27(3):230–232, August 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        03/tiff/230.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        03/tiff/231.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        03/tiff/232.tif.

Lee:1997:CMP

[LF97] Hyuk-Jae Lee and José A. B. Fortes. Communication-minimal partitioning and data alignment for affine nested loops. The Computer Journal, 40 (6):302–310, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www.oup.co.uk/computer_
                                        journal/Volume_40/Issue_06/
                                        Vol40_06.body.html#AbstractLee;
                                         http://www3.oup.co.uk/computer_

journal/Volume_40/Issue_06/
Vol40_06.body.html#AbstractLee.

Libera:1986:UBS

[LG86] F. D. Libera and F. Gosen. Using B-trees to solve geographic range queries. The Computer Journal, 29(2):176–181, April 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/176.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/177.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/178.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/179.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/180.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/181.tif.

Lam:1995:CCT

[LH95] Kam-Yiu Lam and Sheung-Lun Hung. Concurrency control for time-constrained transactions in distributed databases systems. The Computer Journal, 38(9):704–716, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_38/Issue_09/
Vol38_09.body.html#AbstractLam.

Li:1998:OAJ

[Li98] Keqin Li. Optimizing average job response time via decentralized probabilistic job dispatching in heterogeneous multiple computer systems. The Computer Journal, 41(4):223–??, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_41/Issue_04/
Vol41_04.body.html#AbstractLi.

Licker:1989:HRM

[Lic89]  P. S. Licker.  Human re-

source management and business success in small, hi-tech firms: some lessons for data processing. The Computer Journal, 32(6):533–540, December 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/533.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/534.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/535.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/536.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/537.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/538.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_32/Issue_
06/tiff/539.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/540.tif.

Liggins:1973:AAC

[Lig73] D. Liggins. Algorithm 79:

An algorithm for the capacitated transportation problem. The Computer Journal, 16(3):276–278, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/276.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/277.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/278.tif.

Lill:1970:AMD

[Lil70] Shirley A. Lill. Algorithm 46: A modified Davidon method for finding the minimum of a function using difference approximation for derivatives. The Computer Journal, 13(1):111–113, 106, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/tiff/111.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/tiff/112.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_

                                        01/tiff/113.tif. See note

[Lil71, KL71].

Lill:1971:NAM

[Lil71] Shirley A. Lill. Note on Algorithm 46: A modified Davidon method for finding the minimum of a function using difference approximation for derivatives. The Computer Journal, 14(1):106, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        01/tiff/106.tif. See [Lil70,

KL71].

Linz:1969:NMV

[Lin69] P. Linz. Numerical methods for Volterra integral equations of the first kind. The Computer Journal, 12(4):393–397, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/120393.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/393.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/394.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/395.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/396.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_12/Issue_
04/tiff/397.tif.

Lindsey:1972:AFT

[Lin72a] C. H. Lindsey. ALGOL 68 with fewer tears. The Computer Journal, 15(2):176–188, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/150176.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/176.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/177.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/178.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/179.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/180.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/181.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/182.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/183.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/184.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/185.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_15/Issue_
                                        02/tiff/186.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        02/tiff/187.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        02/tiff/188.tif.

Ling:1972:TCK

[Lin72b] R. F. Ling. On the theory and construction of kclusters. The Computer Journal, 15(4):326–332, November 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/150326.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/tiff/326.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/tiff/327.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/tiff/328.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/tiff/329.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/tiff/330.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/tiff/331.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/tiff/332.tif.

Lions:1967:OSS

[Lio67] J. Lions. The Ontario school scheduling program. The Computer Journal, 10(1):14–21, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/100014.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/14.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_01/tiff/
15.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_10/Issue_01/tiff/16.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/17.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/18.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_01/tiff/
19.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_10/Issue_01/tiff/20.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/21.tif.

Lipsky:1980:STS

[Lip80] L. Lipsky. A study of time sharing systems considered as queueing networks of exponential servers. The Computer Journal, 23(4):290–297, November 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/290.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_23/Issue_
                                        04/tiff/291.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        04/tiff/292.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        04/tiff/293.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        04/tiff/294.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        04/tiff/295.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        04/tiff/296.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        04/tiff/297.tif.

Lister:1974:VSP

[Lis74]  A. M. Lister.  Validation

of systems of parallel processes. The Computer Journal, 17(2):148–151, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        02/tiff/148.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        02/tiff/149.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        02/tiff/150.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        02/tiff/151.tif.

Littlewood:1970:APE

[Lit70] P. G. Littlewood. Algorithm 48: Procedure for the evaluation of an integral occurring in the mean square response analysis of linear systems. The Computer Journal, 13(2):207–208, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/206.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/207.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/208.tif. See note

[GE71].

Livesey:1959:SEU

[Liv59] P. B. Livesey. Symposium on experiences with the use of magnetic tape — 2: magnetic films on a National-Elliott 405. The Computer Journal, 2(3):120–121, October 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/020120.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/120.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/121.tif.

Livesley:1960:ALS

[Liv60] R. K. Livesley. The analysis of large structural systems. The Computer Journal, 3(1):34–39, April 1960. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        01/030034.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        01/tiff/34.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_03/Issue_01/tiff/
                                        35.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_03/Issue_01/tiff/36.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        03/Issue_01/tiff/37.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        01/tiff/38.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_03/Issue_01/tiff/
                                        39.tif.

Li:1998:CTD

[LJ98] Y. H. Li and A. K. Jain. Classification of text documents. The Computer Journal, 41(8):537–546, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_41/Issue_08/
                                        Vol41_08.body.html#AbstractLi.

Lucas:1976:SPE

[LK76] H. C. Lucas, Jr. and R. B. Kaplan. A structured programming experiment. The Computer Journal, 19(2):136–138, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_

journal/hdb/Volume_19/Issue_
02/tiff/136.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/137.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/138.tif.

Lyness:1977:TCA

[LK77] J. N. Lyness and J. J.

Kaganove. A technique for comparing automatic quadrature routines. The Computer Journal, 20(2):170–177, May 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/170.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/171.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/172.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/173.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/174.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/175.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/176.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/177.tif.

Lavington:1971:EPU

[LKK71] S. H. Lavington, D. J. Kinniment, and A. E. Knowles. An experimental paging unit. The Computer Journal, 14 (1):55–60, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        01/140055.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        01/tiff/55.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_14/Issue_01/tiff/
                                        56.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_14/Issue_01/tiff/57.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        14/Issue_01/tiff/58.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        01/tiff/59.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_14/Issue_01/tiff/
                                        60.tif.

Learner:1970:NTC

[LL70] A. Learner and A. L. Lim. A note on transforming contextfree grammars to Wirth-Weber precedence form. The Computer Journal, 13(2):142–144, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/130142.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_

journal/hdb/Volume_13/Issue_
02/tiff/142.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/143.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/144.tif.

Lim:1975:TSP

[LL75] A. L. Lim and G. R. Lewis.

Towards structured programs in APL. The Computer Journal, 18(2):140–143, May 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/140.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/141.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/142.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/143.tif.

Levene:1990:NRT

[LL90] M. Levene and G. Loizou. The nested relation type model: an application of domain theory to databases. The Computer Journal, 33(1):19–30, February 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
01/tiff/19.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_01/tiff/

20.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_33/Issue_01/tiff/21.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/22.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
01/tiff/23.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_01/tiff/
24.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_33/Issue_01/tiff/25.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/26.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
01/tiff/27.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_01/tiff/
28.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_33/Issue_01/tiff/29.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/30.tif.

Llewellyn:1973:DEF

[Lle73] J. A. Llewellyn. The deadly embrace — a finite state model approach. The Computer Journal, 16(3):223–225, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/160223.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/223.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_16/Issue_
03/tiff/224.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/225.tif.

Llewellyn:1983:FSM

[Lle83] John A. Llewellyn. Finite state models in the study of commafree codes. The Computer Journal, 26(4):303–311, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/303.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/304.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/305.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/306.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/307.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/308.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/309.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/310.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/311.tif.

Llewellyn:1985:CVC

[Lle85]  John A. Llewellyn.  The

construction of variable-length comma-free codes. The Computer Journal, 28(4):379– 390, August 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
28/Issue_04/tiff/379.tif;
 http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/380.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/381.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/382.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/383.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/384.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/385.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/386.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/387.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/388.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/389.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/390.tif.

Llewellyn:1987:DCS

[Lle87] J. A. Llewellyn. Data compression for a source with Markov characteristics. The Computer Journal, 30(2):149–156, April 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/149.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/150.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/151.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/152.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/153.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/154.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/155.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/156.tif.

Lloyd:1964:DR

[Llo64]  D. B. Lloyd.  Data re-

trieval. The Computer Journal, 7(2):110–113, July 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/070110.sgm.abs.html;

                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        02/tiff/110.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        02/tiff/111.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        02/tiff/112.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        02/tiff/113.tif.

Lloyd:1974:BRS

[Llo74]  Cliff Lloyd.  Book review:

Structured programming, by O.-J. Dahl, E. W. Dijkstra and C. A. R. Hoare. The Computer Journal, 17(2):186, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        02/tiff/186.tif.

Leaman:1973:DTP

[LLR73] R. J. Leaman, M. H. Lloyd, and C. S. Repton. The development and testing of a processor selftest program. The Computer Journal, 16(4):308–314, November 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        04/160308.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        04/tiff/308.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        04/tiff/309.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/310.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/311.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/312.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/313.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/314.tif.

Lambert:1963:UHD

[LM63a] J. D. Lambert and A. R.

Mitchell. The use of higher derivatives in quadrature formulae. The Computer Journal, 5(4):322–327, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/050322.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/322.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/323.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/324.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/325.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/326.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_05/Issue_
                                        04/tiff/327.tif.

Lyness:1963:IMH

[LM63b] J. N. Lyness and B. J. J. McHugh. Integration over multidimensional hypercubes I. A progressive procedure. The Computer Journal, 6 (3):264–270, October 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        03/060264.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        03/tiff/264.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        03/tiff/265.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        03/tiff/266.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        03/tiff/267.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        03/tiff/268.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        03/tiff/269.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        03/tiff/270.tif.

Lloyd:1968:BOR

[LM68] T. Lloyd and H. McCallion. Bounds for the optimum over-relaxation factor for the S.O.R. solution of Laplace type equations over irregular regions. The Computer Journal, 11(3):329–331, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/110329.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/329.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/330.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/331.tif.

Levy:1977:AAP

[LM77a] L. S. Levy and R. Melville.

The algebraic anatomy of programs. The Computer Journal, 20(4):340–345, November 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/340.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/341.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/342.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/343.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/344.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/345.tif.

Lowden:1977:SCD

[LM77b] B. G. T. Lowden and I. R. MacCallum. Some consequences of deferred binding in COBOL. The Computer Journal, 20(1):51–55, February 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        01/tiff/51.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_20/Issue_01/tiff/
                                        52.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_20/Issue_01/tiff/53.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        20/Issue_01/tiff/54.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        01/tiff/55.tif.

Lucas:1988:CMP

[LM88] C. Lucas and G. Mitra. Computer-Assisted Mathematical Programming (Modelling) Systems: CAMPS. The Computer Journal, 31(4):364– 375, August 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        31/Issue_04/tiff/364.tif;
                                         http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        04/tiff/365.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        04/tiff/366.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_31/Issue_
04/tiff/367.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/368.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/369.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/370.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/371.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/372.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/373.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/374.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/375.tif.

Loetzsch:1979:SIS

[LMP79] J. Loetzsch, M. Machura, and W. Paetzold. Sample implementation of a special purpose language by the DE-POT system. The Computer Journal, 22(4):357–362, November 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/357.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/358.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_22/Issue_
                                        04/tiff/359.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        04/tiff/360.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        04/tiff/361.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        04/tiff/362.tif.

Lozier:1973:BCP

[LMS73] D. W. Lozier, L. C. Maximon, and W. L. Sadowski. A bit comparison program for algorithm testing. The Computer Journal, 16(2):111–117, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/160111.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/111.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/112.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/113.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/114.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/115.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/116.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_16/Issue_
02/tiff/117.tif.

Leipala:1979:PSS

[LN79] T. Leipala and O. Nevalainen.

A plotter sequencing system. The Computer Journal, 22 (4):313–316, November 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/313.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/314.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/315.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/316.tif.

Lanzarone:1975:PCR

[LO75] G. A. Lanzarone and M. Ornaghi. Program construction by refinements preserving correctness. The Computer Journal, 18(1):55–62, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
01/tiff/55.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_01/tiff/
56.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_18/Issue_01/tiff/57.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
18/Issue_01/tiff/58.tif;

                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        01/tiff/59.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_18/Issue_01/tiff/
                                        60.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_18/Issue_01/tiff/61.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        18/Issue_01/tiff/62.tif.

Lo:1984:UMS

[Lo84] C. J. Lo. Unifying multiple semantic intentions for a syntactic construct. The Computer Journal, 27(1):83–84, February 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        01/tiff/83.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_27/Issue_01/tiff/
                                        84.tif.

Liu:1999:MAH

[LO99] Kecheng Liu and Tina Ong. A modelling approach for handling business rules and exceptions. The Computer Journal, 42(3):221–231, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        03/420221.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        03/pdf/420221.pdf.

LoDato:1970:PCC

[LoD70] V. A. LoDato. The permutation of a certain class of matrices. The Computer Journal, 13(4):405–410, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/130405.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/405.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/406.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/407.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/408.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/409.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/410.tif.

LoDato:1971:MDP

[LoD71] V. A. LoDato. The minimisation of distance in placement algorithms. The Computer Journal, 14(3):260–262, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/140260.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_

                                        03/tiff/260.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/tiff/261.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/tiff/262.tif.

Loizou:1968:EER

[Loi68] G. Loizou. An empirical estimate of the relative error of the computed solution x of Ax=b. The Computer Journal, 11(1):91–94, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        01/110091.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        01/tiff/91.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_11/Issue_01/tiff/
                                        92.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_11/Issue_01/tiff/93.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        11/Issue_01/tiff/94.tif.

Loizou:1972:QCJ

[Loi72] G. Loizou. On the quadratic convergence of the Jacobi method for normal matrices. The Computer Journal, 15(3):274–276, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/150274.sgm.abs.html;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/274.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/275.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/276.tif.

Loucif:1999:PMB

[LOKM99] S. Loucif, M. Ould-Khaoua, and L. M. Mackenzie. On the performance merits of bypass channels in hypermeshes and k-ary n-cubes. The Computer Journal, 42(1):62–72, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
01/420067.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
01/pdf/420062.pdf.

Lombardi:1988:ACD

[Lom88]  F. Lombardi.  Analysis of

comparison-based diagnosable systems using temporal criteria. The Computer Journal, 31(3):201–208, June 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/201.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/202.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_

03/tiff/203.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/204.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/205.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/206.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/207.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/208.tif.

Long:1961:DTS

[Lon61] P. A. Long. A data transmission survey. The Computer Journal, 4(2):95–102, July 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/040095.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/100.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/101.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/102.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/95.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_02/tiff/
96.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_04/Issue_02/tiff/97.

tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
04/Issue_02/tiff/98.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/99.tif.

Loopstra:1959:XC

[Loo59] B. J. Loopstra. The X-1 computer. The Computer Journal, 2(1):39–43, April 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
01/020039.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
01/tiff/39.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_01/tiff/
40.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_02/Issue_01/tiff/41.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
02/Issue_01/tiff/42.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
01/tiff/43.tif.

Lopriore:1994:SCM

[Lop94] Lanfranco Lopriore. Stack cache memory for block-structured programs. The Computer Journal, 37(7):610–620, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_37/Issue_07/
Vol37_07.body.html#AbstractLopriore.

Loretan:1976:TPE

[Lor76]  R. P. Loretan.  Timing

problems in electronic lockout chains. The Computer Journal, 19(3):250–253, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/250.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/251.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/252.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/253.tif.

Lowden:1977:NOS

[Low77] B. G. T. Lowden. A note on the oscillating sort. The Computer Journal, 20(1):92, February 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        01/tiff/92.tif.

Lowe:1995:SOM

[Low95]  Gavin Lowe.  Scheduling-

oriented models for real-time systems. The Computer Journal, 38(6):443–456, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_38/Issue_06/
                                        Vol38_06.body.html#AbstractLowe.

Lucas:1973:MSE

[LP73a] H. C. Lucas, Jr. and L. Presser.

A method of software evaluation: the case of programming language translators. The Computer Journal, 16(3):226–231, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/160226.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/226.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/227.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/228.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/229.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/230.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/231.tif.

Lynch:1973:PSP

[LP73b] M. F. Lynch and J. H. Petrie.

A program suite for the production of articulated subject indexes. The Computer Journal, 16(1):46–51, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_

                                        01/160046.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        01/tiff/46.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_16/Issue_01/tiff/
                                        47.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_16/Issue_01/tiff/48.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        16/Issue_01/tiff/49.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        01/tiff/50.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_16/Issue_01/tiff/
                                        51.tif.

Larsen:1992:SNE

[LP92] P. G. Larsen and N. Plat. Standards for non-executable specification languages. The Computer Journal, 35(6):567– 573, December 1992. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/Volume_
                                        36/Issue_06/Vol35_06.body.
                                        html#AbstractLarsen.

Lorentzos:1994:IUO

[LPS94] Nikos A. Lorentzos, Alexandra Poulovassilis, and Carol Small. Implementation of update operations for interval relations. The Computer Journal, 37(3):164–176, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_

journal/Volume_37/Issue_03/
Vol37_03.body.html#AbstractLorentzos.

Lam:1999:REC

[LPSC99] Kam-Yiu Lam, Chung-Leung Pang, Sang H. Son, and Jiannong Cao. Resolving executingcommitting conflicts in distributed real-time database systems. The Computer Journal, 42(8):674–692, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
08/420674.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
08/pdf/420674.pdf.

Loia:1996:OSD

[LQ96] Vincenzo Loia and Michel Quaggetto. The OPLA system: designing complex systems in an object-oriented logic programming framework. The Computer Journal, 39(1):20– 35, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www.oup.co.uk/jnls/
list/comjnl/hdb/Volume_39/
Issue_01/390020.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_01/
Vol39_01.body.html#AbstractLoia.

Lavington:1967:SFS

[LR67] S. H. Lavington and L. E.

Rosenthal. Some facilities for speech processing by computer. The Computer Journal, 9(4):330–339, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/090330.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/330.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/331.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/332.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/333.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/334.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/335.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/336.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/337.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/338.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/339.tif.

Lubran:1972:SOL

[LR72] J. F. Lubran and J. D. Roberts. Some observations on ‘least time to go’ scheduling. The Computer Journal, 15(1):32–36, February 1972.

CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/150032.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/tiff/32.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_01/tiff/
33.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_15/Issue_01/tiff/34.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
15/Issue_01/tiff/35.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/tiff/36.tif.

Lewis:1978:TPR

[LR78] B. A. Lewis and J. S. Robinson. Triangulation of planar regions with applications. The Computer Journal, 21 (4):324–332, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/324.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/325.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/326.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/327.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_

                                        04/tiff/328.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        04/tiff/329.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        04/tiff/330.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        04/tiff/331.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        04/tiff/332.tif.

Little:1978:CMF

[LS78] W. Little and D. A. Smith. Correspondence: On “Modifications to the Forsythe-Moler algorithm for solving linear algebraic systems”. The Computer Journal, 21(4):383, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        04/tiff/383.tif. See [SG78,

Hit78].

Liska:1993:AQE

[LS93a] Richard Liska and Stanly Steinberg. Applying quantifier elimination to stability analysis of difference schemes. The Computer Journal, 36 (5):497–503, October 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_36/Issue_05/
                                        Vol36_05.body.html#AbstractLiska.

Lougher:1993:DSS

[LS93b] Phillip Lougher and Doug Shepherd. The design of a storage server for continuous media. The Computer Journal, 36(1):32–42, February 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_36/Issue_01/
Vol36_01.body.html#AbstractLougher.

Leong:1997:DCA

[LS97] Hong Va Leong and Antonio Si. Database caching over the air-storage. The Computer Journal, 40(7):401–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_40/Issue_07/
Vol40_07.body.html#AbstractLeong.

Laenens:1989:BCO

[LSV89] E. Laenens, F. Staes, and D. Vermeir. Browsing a la carte in object-oriented databases. The Computer Journal, 32(4):333–340, August 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/333.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/334.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/335.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/336.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/337.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/338.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/339.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/340.tif.

Long:1961:DTP

[LT61] P. A. Long and E. H. Truslove. Data transmission — problems and prospects. The Computer Journal, 4(1):34–37, April 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/040034.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/tiff/34.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_01/tiff/
35.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_04/Issue_01/tiff/36.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
04/Issue_01/tiff/37.tif.

Leonard:1971:NFO

[LT71] A. J. Leonard and M. E. Tribe. Notable features of Orion. The Computer Journal, 14(4):344–347, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/140344.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/344.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/345.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/346.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/347.tif.

Lotfi:1979:MNN

[LT79] Z. M. Lotfi and A. J. Tosser.

Minimising the NAND-NOR-XOR network of modulo 2 sum of Boolean products. The Computer Journal, 22 (3):251–255, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/251.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/252.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/253.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/254.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/255.tif.

Lu:1989:IST

[Lu89] Xue-Miao M. Lu. An iterative solution for the 4-peg Towers of Hanoi. The Computer Journal, 32(2):187–189, April 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        02/tiff/187.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        02/tiff/188.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        02/tiff/189.tif.

Lukaszewicz:1967:ESM

[Luk67] L. Lukaszewicz. EOL — a symbol manipulation language. The Computer Journal, 10(1):53–59, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        01/100053.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        01/tiff/53.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_10/Issue_01/tiff/
                                        54.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_10/Issue_01/tiff/55.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        10/Issue_01/tiff/56.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        01/tiff/57.tif;   http://
                                        www3.oup.co.uk/computer_journal/

hdb/Volume_10/Issue_01/tiff/
58.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_10/Issue_01/tiff/59.
tif.

Luk:1999:NOB

[Luk99] R. W. P. Luk. Near optimal [beta] heap. The Computer Journal, 42(5):391–399, 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
05/420391.sgm.abs.html.

Lumb:1964:EPD

[Lum64]  O. S. Lumb.  Experience

of program development with FACT. The Computer Journal, 7(2):104–107, July 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/070104.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/104.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/105.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/106.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/107.tif.

Lunnon:1969:PSP

[Lun69] W. F. Lunnon. A postage stamp problem. The Computer Journal, 12(4):377–380, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/120377.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/377.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/378.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/379.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/380.tif. See corre-

spondence [Sel72].

Lunnon:1971:MDM

[Lun71]  W. F. Lunnon.  Multi-

dimensional map-folding. The Computer Journal, 14(1):75– 80, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
14/Issue_01/140075.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
14/Issue_01/tiff/75.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/tiff/76.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_01/tiff/
77.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_14/Issue_01/tiff/78.
tif;   http://www3.oup.co.

uk/computer_journal/hdb/Volume_
14/Issue_01/tiff/79.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/tiff/80.tif.

Lunnon:1975:CMP

[Lun75]  W. F. Lunnon.  Counting

multidimensional polyominoes. The Computer Journal, 18 (4):366–367, November 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/366.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/367.tif.

Lutz:1980:ART

[Lut80] R. K. Lutz. An algorithm for the real time analysis of digitised images. The Computer Journal, 23(3):262–269, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/262.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/263.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/264.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/265.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/266.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        03/tiff/267.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        03/tiff/268.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        03/tiff/269.tif.

Lazos:1979:MEI

[LV79] C. Lazos and A. Vafiadis. A method to estimate the I/O buffer size in a computer system. The Computer Journal, 22(4):323–327, November 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        04/tiff/323.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        04/tiff/324.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        04/tiff/325.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        04/tiff/326.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        04/tiff/327.tif.

Lance:1965:CPM

[LW65] G. N. Lance and W. T. Williams. Computer programs for monothetic classification (‘association analysis’). The Computer Journal, 8(3):246–249, October 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/080246.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/246.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/247.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/248.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/249.tif.

Lance:1966:CPH

[LW66a] G. N. Lance and W. T.

Williams. Computer programs for hierarchical polythetic classification (‘similarity analyses’). The Computer Journal, 9(1):60–64, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/090060.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/60.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_01/tiff/
61.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_09/Issue_01/tiff/62.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
09/Issue_01/tiff/63.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/64.tif.

Lindley:1966:NSS

[LW66b] J. Lindley and A. Wragg. The numerical solution of sequential decision problems involving parabolic equations with moving boundaries. The Computer Journal, 9(3):294–297, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        03/090294.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        03/tiff/294.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        03/tiff/295.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        03/tiff/296.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        03/tiff/297.tif.

Lance:1967:GTCa

[LW67a] G. N. Lance and W. T. Williams. A general theory of classificatory sorting strategies 1. Hierarchical systems. The Computer Journal, 9(4):373–380, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        04/090373.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        04/tiff/373.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_09/Issue_
04/tiff/374.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/375.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/376.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/377.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/378.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/379.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/380.tif.

Lance:1967:GTCb

[LW67b] G. N. Lance and W. T.

Williams. A general theory of classificatory sorting strategies II. Clustering systems. The Computer Journal, 10(3):271–277, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/100271.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/271.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/272.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/273.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_10/Issue_
                                        03/tiff/274.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        03/tiff/275.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        03/tiff/276.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        03/tiff/277.tif.

Lance:1967:NCM

[LW67c] G. N. Lance and W. T. Williams. Note on the classification of multi-level data. The Computer Journal, 9 (4):381–382, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        04/090381.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        04/tiff/381.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        04/tiff/382.tif.

Lance:1968:NNI

[LW68] G. N. Lance and W. T. Williams. Note on a new information-statistic classificatory program. The Computer Journal, 11(2):195, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        02/110195.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_

journal/hdb/Volume_11/Issue_
02/tiff/195.tif.

Lance:1971:NND

[LW71] G. N. Lance and W. T.

Williams. A note on a new divisive classificatory program for mixed data. The Computer Journal, 14(2):154–155, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/140154.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/154.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/155.tif.

Leung:1983:DDE

[LW83] C. H. C. Leung and K. Wolfenden.

Disc database efficiency: a scheme for detailed assessment based on semi-Markov models. The Computer Journal, 26(1):10–14, February 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
01/tiff/10.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_01/tiff/
11.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_26/Issue_01/tiff/12.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
26/Issue_01/tiff/13.tif;
http://www3.oup.co.uk/computer_

                                        journal/hdb/Volume_26/Issue_
                                        01/tiff/14.tif.

Leung:1985:AOD

[LW85a] C. H. C. Leung and K. Wolfenden. Analysis and optimisation of data currency and consistency in replicated distributed databases. The Computer Journal, 28(5):518–523, November 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        05/tiff/518.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        05/tiff/519.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        05/tiff/520.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        05/tiff/521.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        05/tiff/522.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        05/tiff/523.tif.

Leung:1985:EDA

[LW85b] C. H. C. Leung and K. Wolfenden. Estimating disc access patterns using diffusion models. The Computer Journal, 28(4):409–411, August 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        04/tiff/409.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/410.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/411.tif.

Leung:1985:MMF

[LW85c] C. H. C. Leung and K. Wolfenden.

Mathematical models of file growth. The Computer Journal, 28(2):179–183, May 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/179.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/180.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/181.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/182.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/183.tif.

Loos:1993:ALQ

[LW93] Rüdiger Loos and Volker Weispfenning. Applying linear quantifier elimination. The Computer Journal, 36(5):450– 462, October 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/Volume_
36/Issue_05/Vol36_05.body.
html#AbstractLoos.

Larmouth:1968:MTM

[LWS68] J. Larmouth and C. Whitby-Strevens. MLS — the Titan mixed language system. The Computer Journal, 11 (3):256–259, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/110256.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/256.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/257.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/258.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/259.tif. See cor-

respondence [PB69, LWS69b, GB69].

Landy:1969:TTS

[LWS69a] B. Landy and C. Whitby-Strevens. TSAS — the timeshared supervisor assembly system. The Computer Journal, 12(2):128–131, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/120128.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/128.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_12/Issue_
02/tiff/129.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/130.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/131.tif.

Larmouth:1969:CP

[LWS69b] J. Larmouth and C. Whitby-Strevens. Correspondence: Paging. The Computer Journal, 12(2):200, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/200.tif. See [LWS68,

PB69, GB69].

Lazos:1979:ICU

[LY79] C. Lazos and J. R. Yandle. Improving CPU utilisation in a multiprogramming system. The Computer Journal, 22(3):203–205, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/203.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/204.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/205.tif.

Lynn:1964:ESS

[Lyn64] M. S. Lynn. On the equivalence of SOR, SSOR and US-SOR as applied to σ_{1}-ordered systems of linear equations. The Computer Journal, 7(1):72–75, April 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
01/070072.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
01/tiff/72.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_01/tiff/
73.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_07/Issue_01/tiff/74.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
07/Issue_01/tiff/75.tif.

Lynch:1965:MCP

[Lyn65] W. C. Lynch. More combinatorial properties of certain trees. The Computer Journal, 7(4):299–302, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/070299.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/299.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/300.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/301.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/302.tif.

Lyness:1969:EIC

[Lyn69]  J. N. Lyness.  The ef-

fect of inadequate convergence criteria in automatic routines. The Computer Journal, 12(3):279–281, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/120279.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/279.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/280.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/281.tif. See corre-

spondence [GL70].

Lynn:1972:RDF

[Lyn72] P. A. Lynn. Recursive digital filters with linear-phase characteristics. The Computer Journal, 15(4):337–342, November 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/150337.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/337.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/338.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/339.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/tiff/340.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/tiff/341.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/tiff/342.tif.

Lyon:1985:AHT

[Lyo85] Gordon Lyon. Achieving hash table searches in one or two bucket probes. The Computer Journal, 28(3):313–318, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        03/tiff/313.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        03/tiff/314.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        03/tiff/315.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        03/tiff/316.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        03/tiff/317.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        03/tiff/318.tif.

Lowe:1995:RCS

[LZ95] Gavin Lowe and Hussein Zedan. Refinement of complex systems: A case study. The Computer Journal, 38(10):785–800, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_38/Issue_10/
Vol38_10.body.html#AbstractLowe.

McDonald:1989:DSC

[MA89] C. S. McDonald and L. Allison. Denotational semantics of a command interpreter and their implementation in Standard ML. The Computer Journal, 32(5):422–431, October 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/422.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/423.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/424.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/425.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/426.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/427.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/428.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/429.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/430.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_32/Issue_
                                        05/tiff/431.tif.

MacCallum:1970:SIG

[Mac70a] K. J. MacCallum. Surfaces for interactive graphical design. The Computer Journal, 13(4):352–358, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/130352.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/352.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/353.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/354.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/355.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/356.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/357.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/358.tif.

Macleod:1970:SFI

[Mac70b] I. A. Macleod. SP/1 — a FORTRAN integrated string processor. The Computer Journal, 13(3):255–260, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/130255.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/255.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/256.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/257.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/258.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/259.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/260.tif.

Macleod:1971:MFM

[Mac71]  I. A. Macleod.  MP/1 —

a FORTRAN macroprocessor. The Computer Journal, 14(3):229–231, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/140229.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/229.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/230.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/231.tif.

MacCallum:1973:SAO

[Mac73] I. R. MacCallum. A simple analysis of the n^{th} order polyphase sort. The Computer Journal, 16(1):16–18, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        01/160016.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        01/tiff/16.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_16/Issue_01/tiff/
                                        17.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_16/Issue_01/tiff/18.

tif. See errata [Cam73].

MacCallum:1974:CMI

[Mac74] K. J. MacCallum. Correspondence: On mixing interpretive and machine code. The Computer Journal, 17(2):188, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        02/tiff/188.tif. See [Daw73,

DP73, Bem74].

Macleod:1981:RMB

[Mac81] Ian A. Macleod. Relational model as a basis for document retrieval system design. The Computer Journal, 24(4):312–315, November 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/312.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/313.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/314.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/315.tif.

Macleod:1991:QLR

[Mac91] I. A. Macleod. A query language for retrieving information from hierarchic text structures. The Computer Journal, 34(3):254–264, June 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/254.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/255.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/256.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/257.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/258.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/259.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/260.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_34/Issue_
                                        03/tiff/261.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        03/tiff/262.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        03/tiff/263.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        03/tiff/264.tif.

Maddison:1980:FLH

[Mad80]  J. A. T. Maddison.  Fast

lookup in hash tables with direct rehashing. The Computer Journal, 23(2):188–189, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        02/tiff/188.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        02/tiff/189.tif. See corre-

spondence [Bre81, CF81].

Madria:1997:SCC

[Mad97] Sanjay Kumar Madria. A study of the concurrency control and recovery algorithms in nested transaction environment. The Computer Journal, 40(10):630–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_40/Issue_10/
                                        Vol40_10.body.html#AbstractMadria.

Madria:1998:HMC

[Mad98] Sanjay Kumar Madria. Handling of mutual conflicts in distributed databases using timestamps. The Computer Journal, 41(6):376–385, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_41/Issue_06/
Vol41_06.body.html#AbstractMadria.

Maes:1978:RPS

[Mae78] R. Maes. On the representation of program structures by decision tables: a critical assessment. The Computer Journal, 21(4):290–295, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/290.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/291.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/292.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/293.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/294.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/295.tif.

Mahjoub:1984:SED

[Mah84] A. Mahjoub. On the static evaluation of distributed systems performance. The Computer Journal, 27(3):201–208, August 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/201.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/202.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/203.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/204.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/205.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/206.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/207.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/208.tif.

Mahmoud:1986:EDD

[Mah86] H. M. Mahmoud. The expected distribution of degrees in random binary search trees. The Computer Journal, 29(1):36–37, February 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
01/tiff/36.tif;   http://
www3.oup.co.uk/computer_journal/

hdb/Volume_29/Issue_01/tiff/
37.tif.

Maibaum:1985:DIA

[Mai85] T. S. E. Maibaum. Database instances, abstract data types and database specification. The Computer Journal, 28(2):154– 161, May 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
28/Issue_02/tiff/154.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/155.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/156.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/157.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/158.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/159.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/160.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/161.tif.

Majumdar:1994:NIA

[Maj94] A. A. K. Majumdar. A note on the iterative algorithm for the Reve’s puzzle. The Computer Journal, 37(5):463–466, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_37/Issue_05/
Vol37_05.body.html#AbstractMajumdar.

Makinson:1968:AEC

[Mak68a] G. J. Makinson. Algorithm 36: Evaluation of the complex arcsin. The Computer Journal, 11(3):348–350, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/348.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/349.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/350.tif.

Makinson:1968:SHO

[Mak68b]  G. J. Makinson.  Stable

high order implicit methods for the numerical solution of systems of differential equations. The Computer Journal, 11(3):305–310, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/110305.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/305.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/306.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/307.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/308.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/309.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/310.tif.

Makinen:1989:ITA

[Mak89] E. Makinen. On implementing two adaptive data-compression schemes. The Computer Journal, 32(3):238–240, June 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/238.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/239.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/240.tif.

Makinen:1991:EGR

[Mak91a] E. Makinen. Efficient generation of rotational-admissible codewords for binary trees. The Computer Journal, 34(4): 379, August 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
34/Issue_04/tiff/379.tif.

Makinen:1991:SBT

[Mäk91b] E. Mäkinen. A survey on binary tree codings. The Computer Journal, 34(5):438–443, October 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/438.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/439.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/440.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/441.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/442.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/443.tif.

Makinen:2001:CFM

[Mäk01]  Erkki Mäkinen.  Comment

on ‘A Framework for Modelling Trojans and Computer Virus Infection’. The Computer Journal, 44(4):321–323, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_44/Issue_
04/440321.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_44/Issue_
04/pdf/440321.pdf.  See

[TAC98, TAC01, Coh01].

Mallinson:1959:PLA

[Mal59] C. W. Mallinson. Problems of local authorities in data processing. The Computer Journal, 2(3):105–107, October 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/020105.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/105.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/106.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/107.tif.

Maley:1961:EPR

[Mal61] C. E. Maley. The effect of parameters on the roots of an equation system. The Computer Journal, 4(1):62–63, April 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/040062.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/tiff/62.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_01/tiff/
63.tif.

Mallach:1977:SST

[Mal77] E. G. Mallach. Scatter storage techniques: A uniform viewpoint and a method for reducing retrieval times. The Computer Journal, 20(2):137–140, May 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        02/tiff/137.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        02/tiff/138.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        02/tiff/139.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        02/tiff/140.tif.

Man:1971:HOM

[Man71a] F. T. Man. A high-order method of solution for the Lyapunov matrix equation. The Computer Journal, 14 (3):291–292, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/140291.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/tiff/291.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/tiff/292.tif.

Manning:1971:AWM

[Man71b] J. R. Manning. Algorithm 68: White to move and mate in n moves. The Computer Journal, 14(2):209–213, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/209.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_14/Issue_
02/tiff/210.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/211.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/212.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/213.tif.

Manning:1974:CCS

[Man74]  J. R. Manning.  Conti-

nuity conditions for spline curves. The Computer Journal, 17(2):181–186, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/181.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/182.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/183.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/184.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/185.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/186.tif. See corre-

spondence [San75].

Manocha:1993:EAM

[Man93]  Dinesh Manocha.  Efficient

algorithms for multipolynomial resultant. The Computer Journal, 36(5):485–496, October 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_36/Issue_05/
                                        Vol36_05.body.html#AbstractManocha.

Martin:1958:RKM

[Mar58]  D. W. Martin.  Runge–

Kutta methods for integrating differential equations on high speed digital computers. The Computer Journal, 1(3):118–123, October 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        03/010118.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        03/tiff/118.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        03/tiff/119.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        03/tiff/120.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        03/tiff/121.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        03/tiff/122.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        03/tiff/123.tif.

Marin:1969:STN

[Mar69]  M. A. Marin.  Synthe-

sis of TANT networks using a Boolean analyser. The Computer Journal, 12(3):259– 267, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
12/Issue_03/120259.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
12/Issue_03/tiff/259.tif;
 http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/260.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/261.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/262.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/263.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/264.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/265.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/266.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/267.tif.

Marovac:1973:SDD

[Mar73] N. Marovac. A single datadisplay structure: a new view on interactive computer graphics in CAD. The Computer Journal, 16(2):152–156, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/160152.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/152.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/153.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/154.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/155.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/156.tif.

Marovac:1974:MDG

[Mar74a] N. Marovac. A method for defining general networks for CAD, using interactive computer graphics. The Computer Journal, 17(4):332–336, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        04/tiff/332.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        04/tiff/333.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        04/tiff/334.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        04/tiff/335.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        04/tiff/336.tif.

Marriott:1974:CSA

[Mar74b] H. A. Marriott. Correspondence: On the spelling of ALGORITHM. The Computer Journal, 17(2):187, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/187.tif. See [Lee74,

Mul75].

Marjasoo:1976:MHS

[Mar76] O. Marjasoo. A methodology for hardware and software availability analysis. The Computer Journal, 19(2):132–135, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/132.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/133.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/134.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/135.tif.

Marovac:1979:MCS

[Mar79] N. Marovac. Microcomputers in computer science programmes. The Computer Journal, 22(1):91–94, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

journal/hdb/Volume_22/Issue_
01/tiff/91.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_01/tiff/
92.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_22/Issue_01/tiff/93.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
22/Issue_01/tiff/94.tif.

Marovac:1982:RBB

[Mar82] Nenad Marovac. The rotating bus as a basis for interprocess communication in distributed systems. The Computer Journal, 25(1):22–31, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/22.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/
23.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_25/Issue_01/tiff/24.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/25.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/26.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/
27.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_25/Issue_01/tiff/28.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/29.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_

01/tiff/30.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/
31.tif.

Marshall:1991:DNT

[Mar91a] R. Marshall. Discourse and natural-language text analysis: functional grammar approach. The Computer Journal, 34(6): A035–A060, December 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Marshall:1991:MFS

[Mar91b]  R. Marshall.  Manipulating

full-text scientific databases: A logic-based semantico-prag-matic approach. The Computer Journal, 34(3):245–253, June 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/245.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/246.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/247.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/248.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/249.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/250.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_

                                        03/tiff/251.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        03/tiff/252.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        03/tiff/253.tif.

Mariani:1992:OOO

[Mar92] J. A. Mariani. Oggetto: An object oriented database layered on a triple store. The Computer Journal, 35(2):108–118, April 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_35/Issue_02/
                                        Vol35_02.body.html#AbstractMariani.

Marovac:1994:LAC

[Mar94] Nenad Marovac. Link associated computation in Hyper-Net. The Computer Journal, 37(2):88–95, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_37/Issue_02/
                                        Vol37_02.body.html#AbstractMarovac.

Massey:1963:CNS

[Mas63] R. G. Massey. Computers in a new steelworks. The Computer Journal, 5(4):271–275, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        04/050271.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_

journal/hdb/Volume_05/Issue_
04/tiff/271.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/272.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/273.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/274.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/275.tif.

Matthewman:1963:NSS

[Mat63]  J. H. Matthewman.  Note

on the selective summation of Fourier series. The Computer Journal, 6(3):248–249, October 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/060248.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/248.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/249.tif.

Matthewman:1967:CF

[Mat67] J. C. Matthewman. The crystallographers’ friend. The Computer Journal, 10(1):38–45, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/100038.sgm.abs.html;
http://www3.oup.co.uk/computer_

                                        journal/hdb/Volume_10/Issue_
                                        01/tiff/38.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_10/Issue_01/tiff/
                                        39.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_10/Issue_01/tiff/40.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        10/Issue_01/tiff/41.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        01/tiff/42.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_10/Issue_01/tiff/
                                        43.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_10/Issue_01/tiff/44.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        10/Issue_01/tiff/45.tif.

Maude:1973:IMG

[Mau73]  A. D. Maude.  Interpola-

tion — mainly for graph plotters. The Computer Journal, 16(1):64–65, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        01/160064.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        01/tiff/64.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_16/Issue_01/tiff/
                                        65.tif.

Mayers:1964:DAL

[May64] D. F. Mayers. The deferred approach to the limit in ordinary differential equations. The Computer Journal, 7(1):54–57, April 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
01/070054.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
01/tiff/54.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_01/tiff/
55.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_07/Issue_01/tiff/56.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
07/Issue_01/tiff/57.tif.

Mayne:1966:ACP

[May66] D. Mayne. An algorithm for the calculation of the pseudoinverse of a singular matrix. The Computer Journal, 9(3):312–317, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/090312.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/312.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/313.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/314.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/315.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_09/Issue_
                                        03/tiff/316.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        03/tiff/317.tif.

Mazlack:1978:PSS

[Maz78]  L. J. Mazlack.  Predicting

student success in an introductory programming course. The Computer Journal, 21 (4):380–381, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        04/tiff/380.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        04/tiff/381.tif.

Miller:1966:AER

[MB66] J. C. P. Miller and D. J. Spencer Brown. An algorithm for evaluation of remote terms in a linear recurrence sequence. The Computer Journal, 9(2):188–190, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        02/090188.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        02/tiff/188.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        02/tiff/189.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        02/tiff/190.tif.

Mukkamala:1990:ESE

[MB90] R. Mukkamala and S. C. Bruell. Efficient schemes to evaluate transaction performance in distributed database systems. The Computer Journal, 33(1):79–89, February 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
01/tiff/79.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_01/tiff/
80.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_33/Issue_01/tiff/81.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/82.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
01/tiff/83.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_01/tiff/
84.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_33/Issue_01/tiff/85.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/86.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
01/tiff/87.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_01/tiff/
88.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_33/Issue_01/tiff/89.
tif.

Morrissey:1996:MDT

[MB96] Joan M. Morrissey and W. T. Bealor. Minimizing data transfers in distributed query processing: A comparative study and evaluation. The Computer Journal, 39(8):675–??, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://www.oup.co.uk/jnls/
                                        list/comjnl/hdb/Volume_39/
                                        Issue_08/390675.sgm.abs.html;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_39/Issue_08/
                                        Vol39_08.body.html#AbstractMorrissey.

Mackay:1986:UBS

[MBC^{+}86] D. Mackay, G. Ball, M. Crowe, M. Hughes, D. Jenkins, and C. Nicol. A UNIX-based system for software configuration management. The Computer Journal, 29(6):527–530, December 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        06/tiff/527.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        06/tiff/528.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        06/tiff/529.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        06/tiff/530.tif.

Ma:1995:PAB

[MC95] Ruey-Liang Ma and Chung-Ping Chung. Periodic adaptive branch prediction and its application in superscalar processing in Prolog. The Computer Journal, 38(6):457–470, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_38/Issue_06/
Vol38_06.body.html#AbstractMa.

McCallien:1970:SRD

[McC70a] C. W. J. McCallien. The solution of reactor diffusion problems. The Computer Journal, 13(4):369–377, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/130369.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/369.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/370.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/371.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/372.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/373.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/374.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/375.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/376.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/377.tif.

McConalogue:1970:QIS

[McC70b] D. J. McConalogue. A quasiintrinsic scheme for passing a smooth curve through a discrete set of points. The Computer Journal, 13(4):392–396, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/130392.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/392.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/393.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/394.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/395.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/396.tif. See cor-

respondence [Ecc71, Smi71, Sab72, Coo72b].

McConalogue:1971:AAF

[McC71a] D. J. McConalogue. Algorithm 66: An automatic French-curve procedure for use with an incremental plotter. The Computer Journal, 14(2):207–209, May 1971. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/207.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/208.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/209.tif. See note

[McC72].

McConalogue:1971:AAI

[McC71b] D. J. McConalogue. Algorithm 67: An axis invariant procedure to provide slopes at points by local three-point parametric curve fitting. The Computer Journal, 14(2):209, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/209.tif. See note

[SM71].

McConalogue:1972:NAA

[McC72]  D. J. McConalogue.  Note

on Algorithm 66: An automatic French-curve procedure for use with an incremental plotter. The Computer Journal, 15(3):285, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/285.tif. See [McC71a].

McConalogue:1978:ACI

[McC78] D. J. McConalogue. Algorithm 102: Convolution integrals involving probability distribution functions. The Computer Journal, 21(3):270–272, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        03/tiff/270.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        03/tiff/271.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        03/tiff/272.tif.

McCulloch:1982:QDS

[McC82]  C. M. McCulloch.  Quick-

shunt — a distributive sorting algorithm. The Computer Journal, 25(1):102–104, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/102.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/103.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/104.tif.

McCallum:1993:SPS

[McC93] Scott McCallum. Solving polynomial strict inequalities using cylindrical algebraic decomposition. The Computer Journal, 36(5):432–438, October 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_05/
Vol36_05.body.html#AbstractMcCallum.

Morrison:1995:EPL

[MCC^{+}95] Ron Morrison, Richard C. H.

Connor, Quintin I. Cutts, Vivienne S. Dunstan, and Graham N. C. Kirby. Exploiting persistent linkage in software engineering environments. The Computer Journal, 38(1):1–16, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_38/Issue_01/
Vol38_01.body.html#AbstractMorrison.

McDonell:1977:III

[McD77] K. J. McDonell. An inverted index implementation. The Computer Journal, 20(2):116–123, May 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/116.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/117.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/118.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/119.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_

02/tiff/120.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/121.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/122.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/123.tif.

McEwan:1967:AAA

[McE67]  A. T. McEwan.  An At-

las Autocode to ALGOL 60 translator. The Computer Journal, 9(4):353–359, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/090353.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/353.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/354.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/355.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/356.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/357.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/358.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/359.tif.

McGettrick:1977:TMC

[McG77] A. D. McGettrick. Teaching mathematics by computer. The Computer Journal, 20(3):263–268, August 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/263.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/264.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/265.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/266.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/267.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/268.tif.

McIlroy:1982:NSD

[McI82] M. D. McIlroy. The number of states of a dynamic storage allocation system. The Computer Journal, 25(3):388–392, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/388.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/389.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_

                                        03/tiff/390.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        03/tiff/391.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        03/tiff/392.tif.

McLain:1974:DCA

[McL74]  D. H. McLain.  Drawing

contours from arbitrary data points. The Computer Journal, 17(4):318–324, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        04/tiff/318.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        04/tiff/319.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        04/tiff/320.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        04/tiff/321.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        04/tiff/322.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        04/tiff/323.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        04/tiff/324.tif.

McLain:1976:TDI

[McL76] D. H. McLain. Two-dimensional interpolation from random data. The Computer Journal, 19(2):178–181, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/178.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/179.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/180.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/181.tif.

McLain:1978:AVA

[McL78] D. H. McLain. Algorithm 100: Vector approximation to curves. The Computer Journal, 21(2):178–180, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/178.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/179.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/180.tif.

McLean:1985:MGS

[McL85] David McLean. A method of generating surfaces as a composite of cyclide patches. The Computer Journal, 28 (4):433–438, August 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_

04/tiff/433.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/434.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/435.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/436.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/437.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/438.tif.

Melhem:1989:SMW

[MCL89] R. G. Melhem, D. M. Chiarulli, and S. P. Levitan. Space multiplexing of waveguides in optically interconnected multiprocessor systems. The Computer Journal, 32(4):362–369, August 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/362.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/363.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/364.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/365.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/366.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_

04/tiff/367.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/368.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/369.tif.

McOustra:1966:LPC

[McO66]  C. J. C. McOustra.  Le-

gal protection for computer programs. The Computer Journal, 8(4):289–296, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/080289.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/289.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/290.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/291.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/292.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/293.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/294.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/295.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/296.tif.

Matthews:1971:CMN

[MD71] A. Matthews and D. Davies. A comparison of modified Newton methods for unconstrained optimisation. The Computer Journal, 14(3):293–294, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/140293.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/tiff/293.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/tiff/294.tif.

Macmillan:1984:FMI

[MD84] R. J. Macmillan and J. H. Davenport. Factoring mediumsized integers. The Computer Journal, 27(1):84–85, February 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        01/tiff/84.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_27/Issue_01/tiff/
                                        85.tif.

Mayhew:1987:APC

[MD87] P. J. Mayhew and P. A. Dearnley. An alternative prototyping classification. The Computer Journal, 30(6):481–484, December 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_

journal/hdb/Volume_30/Issue_
06/tiff/481.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/482.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/483.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/484.tif.

Morris:1972:SMO

[MDFS72] D. Morris, G. D. Detlefsen, G. R. Frank, and T. J. Sweeney. The structure of the MU5 operating system. The Computer Journal, 15(2):113–116, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/150113.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/113.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/114.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/115.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/116.tif.

Mitchell:1999:BR

[MdlIDS99] Richard Mitchell, B. de la Iglesia, J. C. W. Debuse, and Mark Staples. Book reviews. The Computer Journal, 42(7):623– 625, 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        07/420623.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        07/pdf/420623.pdf.

Messham:1980:CPC

[ME80] D. K. Messham and A. Elliott. Correspondence: Procedure calling and structured architecture. The Computer Journal, 23(4):382–383, November 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        04/tiff/382.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        04/tiff/383.tif. See [BB80b].

Megson:1988:IMP

[ME88a] G. M. Megson and D. J. Evans. Improved matrix product computation using double-pipeline systolic arrays. The Computer Journal, 31(6):567–569, December 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        06/tiff/567.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        06/tiff/568.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        06/tiff/569.tif.

Megson:1988:USD

[ME88b] G. M. Megson and D. J.

Evans. The unification of systolic differencing algorithms. The Computer Journal, 31 (1):83–91, February 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
01/tiff/83.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_01/tiff/
84.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_31/Issue_01/tiff/85.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
31/Issue_01/tiff/86.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
01/tiff/87.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_01/tiff/
88.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_31/Issue_01/tiff/89.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
31/Issue_01/tiff/90.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
01/tiff/91.tif.

Megson:1990:SAS

[ME90] G. M. Megson and D. J.

Evans. A systolic array solution for the assignment problem. The Computer Journal, 33(6):562–569, December 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/562.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/563.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/564.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/565.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/566.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/567.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/568.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/569.tif.

Moon:1998:PIE

[ME98] Soo-Mook Moon and Kemal Ebcioğlu. The performance impact of exploiting branch ILP with tree representation of ILP code. The Computer Journal, 41(1):26–44, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_41/Issue_01/
                                        Vol41_01.body.html#AbstractMoon.

Meadows:1973:AFM

[Mea73]  N. G. Meadows.  Auto-

matic frequency and Mikhailov locus plotting for systems with distributed lags: A new analogue computer technique. The Computer Journal, 16(3):266–268, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/160266.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/266.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/267.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/268.tif.

Meditch:1970:NMD

[Med70] J. S. Meditch. Newton’s method in discrete-time nonlinear data smoothing. The Computer Journal, 13(4):387–391, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/130387.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/387.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/388.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/389.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/390.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/391.tif.

Meek:1972:CSE

[Mee72] B. L. Meek. Correspondence: On “Suggested Extensions to FORTRAN IV”. The Computer Journal, 15(3):267, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/150267.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/267.tif. See [Fin72].

Megson:1990:EDP

[Meg90] G. M. Megson. An efficient double pipe systolic array for matrix product. The Computer Journal, 33(3):261–265, June 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        03/tiff/261.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        03/tiff/262.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        03/tiff/263.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        03/tiff/264.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        03/tiff/265.tif.

Meghini:1996:LIM

[Meg96] Carlo Meghini. Logical image modelling and retrieval. The Computer Journal, 39(3):173– 183, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www.oup.co.uk/jnls/
list/comjnl/hdb/Volume_39/
Issue_03/390173.sgm.abs.html;
 http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_03/
Vol39_03.body.html#AbstractMeghini.

Meister:1966:FCF

[Mei66] B. Meister. On a family of cubature formulae. The Computer Journal, 8(4):368–371, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/080368.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/368.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/369.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/370.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/371.tif.

Meltzer:1966:TPC

[Mel66] B. Meltzer. Theorem-proving for computers: some results on resolution and renaming. The Computer Journal, 8(4):341–343, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        04/080341.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        04/tiff/341.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        04/tiff/342.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        04/tiff/343.tif.

Melham:1995:SIE

[Mel95] T. F. Melham. Special issue editorial: Higher order logic theorem proving and its applications. The Computer Journal, 38(2):89–??, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_38/Issue_02/
                                        Vol38_02.body.html#AbstractMelham.

Merret:1970:CCP

[Mer70] T. H. Merret. Correspondence: Chain processing in PL/1. The Computer Journal, 13(2):221, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/tiff/221.tif.

Merrett:1990:SOC

[Mer90] T. H. Merrett. Scoping is not orthogonal to completeness. The Computer Journal, 33(4):381, August 1990.

CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/381.tif.

Morris:1997:SBC

[MES97] Derrick Morris, D. Gareth Evans, and Simon Schofield. Simulating the behaviour of computer systems: co-simulation of hardware/software. The Computer Journal, 40(10):617– ??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/Volume_
40/Issue_10/Vol40_10.body.
html#AbstractMorris.

Mor:1982:PGV

[MF82] M. Mor and A. S. Fraenkel.

Permutation generation on vector processors. The Computer Journal, 25(4):423–428, November 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/423.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/424.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/425.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/426.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_

                                        04/tiff/427.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        04/tiff/428.tif.

Miller:1988:MDD

[MF88] Mirka Miller and I. Fris. Minimum diameter of diregular digraphs of degree

2. The Computer Journal, 31(1):71–75, February 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        01/tiff/71.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_31/Issue_01/tiff/
                                        72.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_31/Issue_01/tiff/73.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        31/Issue_01/tiff/74.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        01/tiff/75.tif.

Milburn:1964:CI

[MFM64] J. F. Milburn, D. J. Fearnley, and C. G. Myers. A computer in insurance. The Computer Journal, 7(3):173–176, October 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        03/070173.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        03/tiff/173.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_07/Issue_
03/tiff/174.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/175.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/176.tif.

Maddison:1985:DDL

[MG85] R. N. Maddison and A. J.

Gawronski. A data dictionary for learning data analysis. The Computer Journal, 28(3):270–281, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/270.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/271.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/272.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/273.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/274.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/275.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/276.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/277.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_

                                        03/tiff/278.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        03/tiff/279.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        03/tiff/280.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        03/tiff/281.tif.

Moffat:1988:PPP

[MG88] D. S. Moffat and P. M. D. Gray. Perlog: a Prolog with persistence and modules. The Computer Journal, 31(2):110–115, April 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        02/tiff/110.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        02/tiff/111.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        02/tiff/112.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        02/tiff/113.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        02/tiff/114.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        02/tiff/115.tif.

Moore:1995:TUS

[MG95] Simon W. Moore and Brian T. Graham. Tagged up/down sorter — A hardware priority queue. The Computer Journal, 38(9):695–703, ????

1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_38/Issue_09/
Vol38_09.body.html#AbstractMoore.

Macnaghten:1977:FFT

[MH77] A. M. Macnaghten and C. A. R.

Hoare. Fast Fourier transform free from tears. The Computer Journal, 20(1):78–83, February 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
01/tiff/78.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_01/tiff/
79.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_20/Issue_01/tiff/80.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
20/Issue_01/tiff/81.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
01/tiff/82.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_01/tiff/
83.tif.

McFarlane:1994:ODR

[MH94] I. McFarlane and S. G. Hoggar. Optimal drivers for the ‘random’ iteration algorithm. The Computer Journal, 37(7):629–640, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_

                                        journal/Volume_37/Issue_07/
                                        Vol37_07.body.html#AbstractMcFarlane.

Moreau:1998:ELH

[MH98] Luc Moreau and Wendy Hall. On the expressiveness of links in hypertext systems. The Computer Journal, 41(7):459–473, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_41/Issue_07/
                                        Vol41_07.body.html#AbstractMoreau.

Millar:1993:MPU

[MHF93] R. J. Millar, M. Elizabeth C. Hull, and J. H. Frazer. The Millar polyhedron and its use in the construction of octrees. The Computer Journal, 36(2):186–194, April 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_36/Issue_02/
                                        Vol36_02.body.html#AbstractMillar.

Mehrez:1993:ISP

[MHLS93] Abraham Mehrez, Geoffry S. Howard, Yaakov Lugassi, and Peretz Shoval. Information system planning and selection: a multiattribute theoretic approach. The Computer Journal, 36(6):525–541, December 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_36/Issue_06/
                                        Vol36_06.body.html#AbstractMehrez.

Muir:1985:ALN

[MHS85] S. Muir, D. Hutchison, and D. Shepherd. ARCA: a local network file server. The Computer Journal, 28(3):243–249, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/243.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/244.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/245.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/246.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/247.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/248.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/249.tif.

Michie:1963:EMG

[Mic63] D. Michie. Experiments on the mechanization of gamelearning Part I. Characterization of the model and its parameters. The Computer Journal, 6(3):232–236, October 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/060232.sgm.abs.html;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/232.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/233.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/234.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/235.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/236.tif.

Michaelson:1964:PAS

[Mic64] R. L. Michaelson. Presidential address. Science and the non-scientist. The Computer Journal, 6(4):299–303, January 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/060299.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/299.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/300.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/301.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/302.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/303.tif.

Michie:1971:HS

[Mic71] D. Michie. Heuristic search.

The Computer Journal, 14 (1):96–102, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/140096.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/tiff/100.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/tiff/101.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/tiff/102.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/tiff/96.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_01/tiff/
97.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_14/Issue_01/tiff/98.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
14/Issue_01/tiff/99.tif.

Michie:1980:ES

[Mic80] D. Michie. Expert systems.

The Computer Journal, 23 (4):369–376, November 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/369.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_

04/tiff/370.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/371.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/372.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/373.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/374.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/375.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/376.tif. See corre-

spondence [BM81].

Michie:1981:TEC

[Mic81] Donald Michie. A theory of evaluative comments in chess with a note on minimaxing. The Computer Journal, 24(3):278–286, August 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/278.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/279.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/280.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/281.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_24/Issue_
03/tiff/282.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/283.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/284.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/285.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/286.tif.

Michie:1982:EMG

[Mic82] D. Michie. Experiments on the mechanization of gamelearning. II. rule-based learning and the human window. The Computer Journal, 25 (1):105–113, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/105.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/106.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/107.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/108.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/109.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/110.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/111.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/112.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/113.tif.

Michie:1991:MML

[Mic91] D. Michie. Methodologies from machine learning in data analysis and software. The Computer Journal, 34(6):559–565, December 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        06/tiff/559.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        06/tiff/560.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        06/tiff/561.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        06/tiff/562.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        06/tiff/563.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        06/tiff/564.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        06/tiff/565.tif.

Mickle:1997:ECC

[Mic97] Marlin H. Mickle. An electrical circuit and computer system analogy. The Computer Journal, 40(7):457–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_40/Issue_07/
Vol40_07.body.html#AbstractMickle.

Middleton:1976:UMI

[Mid76] A. G. Middleton. On the use of macros for iteration. The Computer Journal, 19(2):170–172, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/170.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/171.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/172.tif.

Middleton:1977:CTF

[Mid77] A. G. Middleton. A case for type and form flow analysis. The Computer Journal, 20(3):238–241, August 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/238.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/239.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/240.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/241.tif.

Middleton:1987:STR

[Mid87] A. G. Middleton. Some transformations of relevance to report generation. The Computer Journal, 30(1):37–42, February 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        01/tiff/37.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_30/Issue_01/tiff/
                                        38.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_30/Issue_01/tiff/39.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        30/Issue_01/tiff/40.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        01/tiff/41.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_30/Issue_01/tiff/
                                        42.tif.

Miller:1960:NNE

[Mil60] J. C. P. Miller. Note on the numerical evaluation of a first derivative from a table of a function satisfying a second order differential equation. The Computer Journal, 3(2):112–113, July 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        02/030112.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        02/tiff/112.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_03/Issue_
02/tiff/113.tif.

Mills:1963:OET

[Mil63] M. R. Mills. Operational experience of time sharing and parallel processing. The Computer Journal, 6(1):28–36, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/060028.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/tiff/28.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/
29.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_06/Issue_01/tiff/30.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
06/Issue_01/tiff/31.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/tiff/32.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/
33.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_06/Issue_01/tiff/34.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
06/Issue_01/tiff/35.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/tiff/36.tif.

Milner:1968:SHA

[Mil68] R. Milner. String handling in ALGOL. The Computer Journal, 10(4):321–324, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/100321.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/321.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/322.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/323.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/324.tif.

Millenson:1970:LLS

[Mil70] J. R. Millenson. Language and list structure of a compiler for experimental control. The Computer Journal, 13(4):340–343, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/130340.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/340.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/341.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/342.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/343.tif.

Millington:1981:SSA

[Mil81a]  Donald Millington.  Struc-

tured systems analysis and design using standard flowcharting symbols. The Computer Journal, 24(4):295–300, November 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/295.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/296.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/297.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/298.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/299.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/300.tif.

Milne:1981:FIS

[Mil81b] Wendy J. Milne. A framework for the investigation of a spatial data base. The Computer Journal, 24(1):52–55, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
01/tiff/52.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_01/tiff/
53.tif;   http://www3.oup.
co.uk/computer_journal/hdb/

                                        Volume_24/Issue_01/tiff/54.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        24/Issue_01/tiff/55.tif.

Mills:1995:LDW

[Mil95] Stella Mills. To live or drown: when information systems become critical. The Computer Journal, 38(6):413–417, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_38/Issue_06/
                                        Vol38_06.body.html#AbstractMills.

Mingers:1988:CCM

[Min88] J. Mingers. Comparing conceptual models and data flow diagrams. The Computer Journal, 31(4):376–379, August 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        04/tiff/376.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        04/tiff/377.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        04/tiff/378.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        04/tiff/379.tif.

Mirkin:1998:LSS

[Mir98] Boris Mirkin. Least-squares structuring, clustering and data processing issues. The Computer Journal, 41(8):518–536, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_41/Issue_08/
Vol41_08.body.html#AbstractMirkin.

Mitchem:1976:VAE

[Mit76]  John Mitchem.  On vari-

ous algorithms for estimating the chromatic number of a graph. The Computer Journal, 19(2):182–183, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/182.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/183.tif.

Mitchell:1989:APS

[Mit89a] C. Mitchell. Another postage stamp problem. The Computer Journal, 32(4):374–376, August 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/374.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/375.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/376.tif.

Mitchell:1989:MSE

[Mit89b] C. Mitchell. Multi-destination secure electronic mail. The Computer Journal, 32(1):13– 15, February 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
32/Issue_01/tiff/13.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
01/tiff/14.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_01/tiff/
15.tif.

Miyakoshi:1990:LCM

[Miy90] H. Miyakoshi. A literal code method for minimizing Boolean functions. The Computer Journal, 33(4):375–380, August 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/375.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/376.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/377.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/378.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/379.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/380.tif.

Mayne:1975:ICF

[MJ75] A. Mayne and E. B. James. Information compression by factorising common strings. The Computer Journal, 18(2):157– 160, May 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
18/Issue_02/tiff/157.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/158.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/159.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/160.tif.

Morris:1982:PCI

[MK82] R. J. T. Morris and J. S. Kaufman. Performance comparison of I/O access disciplines for transaction-processing systems. The Computer Journal, 25(1):74–83, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/74.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/
75.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_25/Issue_01/tiff/76.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/77.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/78.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/

79.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_25/Issue_01/tiff/80.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/81.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/82.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/
83.tif.

Markusz:1985:CCL

[MK85] Z. Markusz and A. A. Kaposi. Complexity control in logic-based programming. The Computer Journal, 28 (5):487–495, November 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/487.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/488.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/489.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/490.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/491.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/492.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/493.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_28/Issue_
05/tiff/494.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/495.tif.

Ma:1994:GTT

[MK94] Jixin Ma and Brian Knight.

A general temporal theory. The Computer Journal, 37 (2):114–123, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_37/Issue_02/
Vol37_02.body.html#AbstractMa.

Ma:1996:RTL

[MK96] Jixin Ma and Brian Knight.

A reified temporal logic. The Computer Journal, 39(9):800– ??, ???? 1996. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www.oup.co.uk/
computer_journal/Volume_39/
Issue_09/Vol39_09.body.html#
AbstractMa;   http://www3.
oup.co.uk/computer_journal/
Volume_39/Issue_09/Vol39_
09.body.html#AbstractMa.

Manolopoulos:1987:BIS

[MKB87] Y. P. Manolopoulos, J.-Y. G.

Kollias, and F. W. Burton. Batched interpolation search. The Computer Journal, 30 (6):565–568, December 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_

journal/hdb/Volume_30/Issue_
06/tiff/565.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/566.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/567.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/568.tif.

Manolopoulos:1986:SVB

[MKH86] Y. Manolopoulos, J. G. Kollias, and M. Hatzopoulos. Sequential vs. binary batched searching. The Computer Journal, 29(4):368–372, August 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/368.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/369.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/370.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/371.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/372.tif.

Morris:1971:F

[MKL71] D. Morris, T. G. Kennedy, and L. Last. Flocoder. The Computer Journal, 14 (3):221–223, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/140221.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/221.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/222.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/223.tif.

Manolopoulos:1987:MIF

[MKP87] Y. Manolopoulos, D. Kleftouris, and L. Petrou. A model for an ISAM file with multiple overflow chains. The Computer Journal, 30(6):529–534, December 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/529.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/530.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/531.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/532.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/533.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/534.tif.

Ma:1994:RTA

[MKP94] J. Ma, B. Knight, and M. Petridis.

A revised theory of action and time based on intervals and points. The Computer Journal, 37(10):847–857, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_
10/370858.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_
10/tiff/847.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_
10/tiff/848.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_
10/tiff/849.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_
10/tiff/850.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_
10/tiff/851.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_
10/tiff/852.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_
10/tiff/853.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_
10/tiff/854.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_
10/tiff/855.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_
10/tiff/856.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_
10/tiff/857.tif.

Murray:1965:COD

[ML65] W. A. Murray and M. S. Lynn.

A computer-oriented description of the Peaceman–Rachford ADI method. The Computer Journal, 8(2):166–175, July 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/080166.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/166.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/167.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/168.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/169.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/170.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/171.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/172.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/173.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/174.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/175.tif.

Mustard:1963:NQD

[MLB63] D. Mustard, J. N. Lyness, and J. M. Blatt. Numerical quadrature in n dimensions. The Computer Journal, 6(1):75–87, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/060075.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/tiff/75.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/
76.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_06/Issue_01/tiff/77.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
06/Issue_01/tiff/78.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/tiff/79.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/
80.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_06/Issue_01/tiff/81.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
06/Issue_01/tiff/82.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/tiff/83.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/
84.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_06/Issue_01/tiff/85.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_

06/Issue_01/tiff/86.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/tiff/87.tif.

Martin:1990:ESS

[MLR90] T. P. Martin, K. H. Lam, and Judy I. Russell. An evaluation of site selection algorithms for distributed query processing. The Computer Journal, 33(1):61–70, February 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
01/tiff/61.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_01/tiff/
62.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_33/Issue_01/tiff/63.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/64.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
01/tiff/65.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_01/tiff/
66.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_33/Issue_01/tiff/67.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/68.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
01/tiff/69.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_01/tiff/
70.tif.

Milledge:1960:FER

[MM60] D. Milledge and M. J. Mills. Forecasting election results. The Computer Journal, 2 (4):195–198, January 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_02/Issue_
                                        04/020195.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_02/Issue_
                                        04/tiff/195.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_02/Issue_
                                        04/tiff/196.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_02/Issue_
                                        04/tiff/197.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_02/Issue_
                                        04/tiff/198.tif.

Marliss:1969:WGB

[MM69] G. S. Marliss and W. A. Murray. William G. Bickley — an appreciation. The Computer Journal, 12(4):301–302, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/120301.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/301.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/302.tif.

McKee:1970:ADM

[MM70] S. McKee and A. R. Mitchell.

Alternating direction methods for parabolic equations in two space dimensions with a mixed derivative. The Computer Journal, 13(1):81–86, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/130081.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/tiff/81.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_01/tiff/
82.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/83.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
13/Issue_01/tiff/84.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/tiff/85.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_01/tiff/
86.tif.

McKee:1971:ADM

[MM71] S. McKee and A. R. Mitchell.

Alternating direction methods for parabolic equations in three space dimensions with mixed derivatives. The Computer Journal, 14(3):295–300, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_

03/140295.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/295.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/296.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/297.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/298.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/299.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/300.tif.

Mottl:1978:DPS

[MM78] J. Mottl and L. Mottlova. Description of a program for solving problems of nonlinear programming. The Computer Journal, 21(3):263–269, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/263.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/264.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/265.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/266.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/267.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/268.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/269.tif.

Mariani:1984:VMV

[MM84] John A. Mariani and Douglas R. McGregor. VUL-CAN: a modified viewdata system with improved indexing and formal hierarchy definition language. The Computer Journal, 27(2):121–126, May 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/121.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/122.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/123.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/124.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/125.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/126.tif.

Mackenzie:1987:MMS

[MMB87] L. M. Mackenzie, A. M.

Macleod, and D. J. Berry. A multiple microprocessor system for CPU-bound calculations. The Computer Journal, 30(2):110–118, April 1987.

CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/110.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/111.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/112.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/113.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/114.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/115.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/116.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/117.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/118.tif.

Morgan:1991:WDM

[MMG91] K. Morgan, R. L. Morris, and S. Gibbs. When does a mouse become a rat? or … comparing performance and preferences in direct manipulation and command line environment. The Computer Journal, 34(3):265–271, June 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

journal/hdb/Volume_34/Issue_
03/tiff/265.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/266.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/267.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/268.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/269.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/270.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/271.tif.

McGregor:1987:AAP

[MMH87] D. McGregor, S. McInnes, and M. Henning. An architecture for associative processing of large knowledge bases (LKBs). The Computer Journal, 30(5):404–412, October 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/404.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/405.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/406.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/407.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_30/Issue_
                                        05/tiff/408.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        05/tiff/409.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        05/tiff/410.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        05/tiff/411.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        05/tiff/412.tif.

Mazzeo:1998:FSC

[MMR^{+}98] Antonino Mazzeo, Nicola Mazzocca, Stefano Russo, Carlo Savy, and Valeria Vittorini. Formal specification of concurrent systems: A structured approach. The Computer Journal, 41(3):145–162, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_41/Issue_03/
                                        Vol41_03.body.html#AbstractMazzeo.

Merry:1961:CQS

[MN61] I. W. Merry and G. O. Norrie. Character quality and scanner organization. The Computer Journal, 4(2):137–144, July 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        02/040137.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_

02/tiff/137.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/138.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/139.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/140.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/141.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/142.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/143.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/144.tif.

Moltedo:1990:ICE

[MN90] L. Moltedo and A. Noferini.

An implementation of 7-bit character encoding for standard computer graphics environments. The Computer Journal, 33(5):429–436, October 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/429.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/430.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/431.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_

                                        05/tiff/432.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        05/tiff/433.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        05/tiff/434.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        05/tiff/435.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        05/tiff/436.tif.

Moffat:1997:ESI

[Mof97] A. Moffat. Editorial for special issue on lossless compression. The Computer Journal, 40(2/3):65–66, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www.oup.co.uk/computer_
                                        journal/Volume_40/Issue_02/
                                        Vol40_02.body.html#AbstractEditorial;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_40/Issue_02/
                                        Vol40_02.body.html#AbstractEditorial;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_40/Issue_02/
                                        Vol40_03.body.html#AbstractEditorial.

Mojena:1977:HGM

[Moj77] R. Mojena. Hierarchical grouping methods and stopping rules: An evaluation. The Computer Journal, 20(4):359–363, November 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        04/tiff/359.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_20/Issue_
04/tiff/360.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/361.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/362.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/363.tif.

Molnar:1971:SSE

[Mol71] G. Molnar. SEL — a selfextensible programming language. The Computer Journal, 14(3):238–242, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/140238.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/238.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/239.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/240.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/241.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/242.tif.

Moorhead:1959:ARA

[Moo59] W. G. Moorhead. Aircraft route analysis on a digital computer. The Computer Journal, 1(4):160–162, January 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/010160.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/tiff/160.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/tiff/161.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/tiff/162.tif.

Moore:1963:ASP

[Moo63] D. W. Moore. Accounting for the soldier’s pay. The Computer Journal, 5(4):249–257, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/050249.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/249.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/250.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/251.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/252.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/253.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/254.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/255.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/256.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/257.tif.

Moore:1979:DTJ

[Moo79] L. Moore. Design for a transportable job organisation language. The Computer Journal, 22(4):296–302, November 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/296.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/297.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/298.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/299.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/300.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/301.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/302.tif.

Morris:1989:RDT

[Mor89] R. A. Morris. Rendering digital type: a historical and economic view of technology.

The Computer Journal, 32 (6):524–532, December 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/524.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/525.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/526.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/527.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/528.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/529.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/530.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/531.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/532.tif.

Mottl:1979:DPN

[Mot79] J. Mottl. Description of a program for nonlinear programming. The Computer Journal, 22(3):256–261, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/256.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/257.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/258.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/259.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/260.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/261.tif.

Mottl:1980:DPN

[Mot80] J. Mottl. Description of a program for nonlinear programming: special problem of bivalent nonlinear programming. The Computer Journal, 23(4):360–368, November 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/360.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/361.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/362.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/363.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/364.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/365.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        04/tiff/366.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        04/tiff/367.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        04/tiff/368.tif.

Mottl:1985:DPN

[Mot85] J. Mottl. Description of a program for nonlinear programming: the centroid program. The Computer Journal, 28(1):89–94, February 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        01/tiff/89.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_28/Issue_01/tiff/
                                        90.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_28/Issue_01/tiff/91.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        28/Issue_01/tiff/92.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        01/tiff/93.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_28/Issue_01/tiff/
                                        94.tif.

Mitchell:1962:HAD

[MP62] A. R. Mitchell and R. P. Pearce. High accuracy difference formulae for the numerical solution of the heat conduction equation. The Computer Journal, 5(2):142–146, July 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/050142.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/142.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/143.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/144.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/145.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/146.tif.

Martin:1965:ECD

[MP65a] E. J. Martin, Jr. and P. C. Patton. Evaluation of certain definite integrals frequently encountered in radiational and diffraction problems involving circular geometry. The Computer Journal, 8(3):256–263, October 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/080256.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/256.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/257.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_

                                        03/tiff/258.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        03/tiff/259.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        03/tiff/260.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        03/tiff/261.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        03/tiff/262.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        03/tiff/263.tif. See erratum

[MP66].

Melbourne:1965:SCD

[MP65b] A. J. Melbourne and J. M. Pugmire. A small computer for the direction processing of FORTRAN statements. The Computer Journal, 8(1):24–27, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        01/080024.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        01/tiff/24.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_08/Issue_01/tiff/
                                        25.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_08/Issue_01/tiff/26.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        08/Issue_01/tiff/27.tif.

Martin:1966:EEC

[MP66] E. J. Martin, Jr. and P. C.

Patton. Erratum: “Evaluation of certain definite integrals frequently encountered in radiational and diffraction problems involving circular geometry”. The Computer Journal, 9(1):116, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/116.tif. See [MP65a].

Morgenstern:1967:NMS

[MP67] N. R. Morgenstern and V. E.

Price. A numerical method for solving the equations of stability of general slip surfaces. The Computer Journal, 9(4):388–393, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/090388.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/388.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/389.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/390.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/391.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_

                                        04/tiff/392.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        04/tiff/393.tif.

Miller:1968:ACP

[MP68] J. C. P. Miller and M. J. Prentice. Additive congruential pseudo-random number generators. The Computer Journal, 11(3):341–346, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        03/110341.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        03/tiff/341.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        03/tiff/342.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        03/tiff/343.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        03/tiff/344.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        03/tiff/345.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        03/tiff/346.tif.

Mavaddat:1979:DSF

[MP79] F. Mavaddat and B. Parhami. A data structure for family relations. The Computer Journal, 22(2):110–113, May 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/110.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/111.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/112.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/113.tif.

Maude:1985:SRR

[MP85] T. I. Maude and D. J. Pullinger.

Software for reading, refereeing and browsing in the BLEND system. The Computer Journal, 28(1):1–4, February 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
01/tiff/1.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_28/Issue_01/tiff/
2.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_28/Issue_01/tiff/3.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
28/Issue_01/tiff/4.tif.

Mancini:1989:PCP

[MP89] L. V. Mancini and G. P.

Pappalardo. Proving correctness properties of a replicated synchronous program. The Computer Journal, 32 (5):437–444, October 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        05/tiff/437.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        05/tiff/438.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        05/tiff/439.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        05/tiff/440.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        05/tiff/441.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        05/tiff/442.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        05/tiff/443.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        05/tiff/444.tif.

Macias:1995:MCP

[MP95] Benjamin Macias and Stephen G. Pulman. A method for controlling the production of specifications in natural language. The Computer Journal, 38(4):310–318, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_38/Issue_04/
                                        Vol38_04.body.html#AbstractMacias.

Moore:1965:CRF

[MR65] Roger H. Moore and R. H. Rosenbrock. Correspondence: On the Rosenbrock function. The Computer Journal, 8(3):224, October 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/224.tif.

Morris:1967:ACS

[MR67] D. Morris and J. S. Rohl.

The Atlas compiler system. The Computer Journal, 10 (3):227–230, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/100227.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/227.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/228.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/229.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/230.tif.

McBride:1976:MOW

[MR76a] W. E. McBride and A. K.

Rigler. A modification of the Osborne and Watson algorithm for nonlinear minimax approximation. The Computer Journal, 19(1):79–81, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_

                                        01/tiff/79.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_19/Issue_01/tiff/
                                        80.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_19/Issue_01/tiff/81.
                                        tif.

Mohn:1976:IPA

[MR76b] Karl-Heinz Mohn and Roger V. Roman. An interactive polynomial approximation algorithm. The Computer Journal, 19(1):74–78, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        01/tiff/74.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_19/Issue_01/tiff/
                                        75.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_19/Issue_01/tiff/76.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        19/Issue_01/tiff/77.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        01/tiff/78.tif.

Mamrak:1977:ASE

[MR77] S. Mamrak and J. M. Randal. An analysis of a software engineering failure. The Computer Journal, 20(4):316–320, November 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        04/tiff/316.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_20/Issue_
04/tiff/317.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/318.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/319.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/320.tif.

Missala:1982:CJS

[MR82] Marek Missala and Piotr Rudnicki. Correspondence: Jumping to some purpose. The Computer Journal, 25(2):286, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/286.tif. See [ASG79,

Hil80, Rob80, Tay80].

Marwick:1985:TM

[MR85] D. H. Marwick and C. R. G. B.

Rinaldi. The Tattoo on a micro-computer. The Computer Journal, 28(3):216–222, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/216.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/217.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/218.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_

                                        03/tiff/219.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        03/tiff/220.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        03/tiff/221.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        03/tiff/222.tif.

Mauw:1994:ASB

[MR94] S. Mauw and M. A. Reniers. An algebraic semantics of basic message sequence charts. The Computer Journal, 37(4):269–277, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_37/Issue_04/
                                        Vol37_04.body.html#AbstractMauw.

Mariani:1996:CIS

[MR96] John A. Mariani and Tom Rodden. Cooperative information sharing: developing a shared object service. The Computer Journal, 39(6):455–470, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://www.oup.co.uk/jnls/
                                        list/comjnl/hdb/Volume_39/
                                        Issue_06/390455.sgm.abs.html;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_39/Issue_06/
                                        Vol39_06.body.html#AbstractMariani.

Millard:1975:SES

[MRW75] G. E. Millard, D. J. Rees, and H. Whitfield. The standard EMAS subsystem. The Computer Journal, 18 (3):213–219, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/213.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/214.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/215.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/216.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/217.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/218.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/219.tif.

Mitchell:1990:SMA

[MRW90] C. Mitchell, D. Rush, and M. Walker. A secure messaging architecture implementing the X.400-1988 security features. The Computer Journal, 33(4):290–295, August 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/290.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/291.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        04/tiff/292.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        04/tiff/293.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        04/tiff/294.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        04/tiff/295.tif.

Murtagh:1970:CEQ

[MS70] B. A. Murtagh and R. W. H. Sargent. Computational experience with quadratically convergent minimisation methods. The Computer Journal, 13(2):185–194, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/130185.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/tiff/185.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/tiff/186.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/tiff/187.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/tiff/188.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/tiff/189.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/tiff/190.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/191.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/192.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/193.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/194.tif.

Mullin:1974:OBF

[MS74] J. K. Mullin and G. F. Stuart. Optimum blocking factors for a mix of sequentially processed jobs. The Computer Journal, 17(3):224–228, August 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/224.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/225.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/226.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/227.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/228.tif.

Mancini:1991:FRC

[MS91] L. Mancini and S. K. Shrivastava. Fault-tolerant reference counting for garbage collection in distributed systems. The Computer Journal, 34(6):503–513, December 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/503.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/504.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/505.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/506.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/507.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/508.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/509.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/510.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/511.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/512.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/513.tif.

Manner:1992:FTD

[MS92] R. Manner and O. Stucky. Fault-tolerant data transfer in a multiprocessor system by forward and backward hardware error recovery. The Computer Journal, 35(4):361–368, August 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_35/Issue_04/
Vol35_04.body.html#AbstractManner.

Muller:1997:TPL

[MS97] Olaf Müller and Konrad Slind.

Treating partiality in a logic of total functions. The Computer Journal, 40(10):640–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_40/Issue_10/
Vol40_10.body.html#AbstractMueller.

Martin:1999:BTE

[MSNW99] Keith M. Martin, Rei Safavi-Naini, and Huaxiong Wang. Bounds and techniques for efficient redistribution of secret shares to new access structures. The Computer Journal, 42(8):638–649, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
08/420638.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
08/pdf/420638.pdf.

Muchnick:1993:CIP

[MSS93] V. B. Muchnick, A. V. Shafarenko, and C. D. Sutton. F - code and its implementation: a portable software platform for data parallelism. The Computer Journal, 36(8):712–722, ???? 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_08/
Vol36_08.body.html#AbstractMuchnick.

Miles:1964:EDL

[MST64] G. A. Miles, K. L. Stewart, and G. J. Tee. Elementary divisors of the Liebmann process. The Computer Journal, 6(4):352–355, January 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/060352.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/352.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/353.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/354.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/355.tif.

Martin:1961:IML

[MT61] D. W. Martin and G. J. Tee. Iterative methods for linear equations with symmetric positive definite matrix. The Computer Journal, 4(3):242–254, October 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

journal/hdb/Volume_04/Issue_
03/040242.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/242.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/243.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/244.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/245.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/246.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/247.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/248.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/249.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/250.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/251.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/252.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/253.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/254.tif.

McLain:1970:MMA

[MT70] T. G. McLain and A. R. Trice.

The MINIMOP multi-access operating system. The Computer Journal, 13(3):237–242, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/130237.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/237.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/238.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/239.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/240.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/241.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/242.tif.

Moon:1972:CEP

[MT72] G. Moon and The Editor. Correspondence: On editorial policy. The Computer Journal, 15(2):124, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        02/150124.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        02/tiff/124.tif. See [Wal71,

GT72, KTW^{+}73].

McKenzie:1983:CSV

[MT83] B. J. McKenzie and T. Takaoko.

A control structure for a variable number of nested loops. The Computer Journal, 26(3):282–283, August 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/282.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/283.tif.

Mullender:1986:DCDb

[MT86] S. J. Mullender and A. S.

Tanenbaum. The design of a capability-based distributed operating system. The Computer Journal, 29(4):289–299, August 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/289.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/290.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/291.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/292.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/293.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/294.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/295.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/296.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/297.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/298.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/299.tif.

Morris:1988:HIR

[MTPL88] D. Morris, C. J. Theaker, R. Phillips, and W. Love. Human-computer interface recording. The Computer Journal, 31(5):437–444, October 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/437.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/438.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/439.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/440.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/441.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/442.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_31/Issue_
05/tiff/443.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/444.tif.

Morris:1988:SAS

[MTW88] D. Morris, C. J. Theaker, and P. B. Whitehead. Structured abstract schematics. The Computer Journal, 31(3):193–200, June 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/193.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/194.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/195.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/196.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/197.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/198.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/199.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/200.tif.

Muchsel:1982:PMP

[Muc82]  Reginald Muchsel.  Perfor-

mance measurements — a practical example from a German University Computing Centre. The Computer Journal, 25(2):188–192, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/188.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/189.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/190.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/191.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/192.tif.

Muir:1958:ASF

[Mui58]  A. Muir.  Automatic sales

forecasting. The Computer Journal, 1(3):113–116, October 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/010113.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/113.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/114.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/115.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/116.tif.

Mulholland:1969:STT

[Mul69]  K. A. Mulholland.  Soft-

ware to translate TELCOMP programs into KDF9ALGOL. The Computer Journal, 12 (3):221–224, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/120221.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/221.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/222.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/223.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/224.tif.

Mullins:1975:COW

[Mul75] Roy S. Mullins. Correspondence: On the origin of the word ‘algorithm’. The Computer Journal, 18(3):286, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/286.tif. See [Mar74b,

Lee74].

Mullin:1981:CAB

[Mul81] James K. Mullin. Change area B-trees: a technique to aid error recovery. The Computer Journal, 24(4):367–373, November 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/367.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/368.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/369.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/370.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/371.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/372.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/373.tif.

Mullin:1985:SSE

[Mul85]  James K. Mullin.  Spiral

storage: Efficient dynamic hashing with constant performance. The Computer Journal, 28(3):330–334, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/330.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/331.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/332.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_28/Issue_
03/tiff/333.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/334.tif.

Mullin:1987:ATD

[Mul87] J. K. Mullin. Accessing textual documents using compressed indexes of arrays of small Bloom filters. The Computer Journal, 30(4):343–348, August 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/343.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/344.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/345.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/346.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/347.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/348.tif.

Mullin:1992:HFH

[Mul92] J. K. Mullin. Hash functions for hash-based join methods. The Computer Journal, 35(6):A499– A503, December 1992. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Mumford:1980:SAS

[Mum80] Enid Mumford. Social aspects of systems analysis. The Computer Journal, 23(1):5–7, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
01/tiff/5.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/
6.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_23/Issue_01/tiff/7.
tif.

Mumford:1984:PSD

[Mum84] Enid Mumford. Participative systems design. The Computer Journal, 27(3):283, August 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/283.tif.

Mumford:1985:DSR

[Mum85] Enid Mumford. Defining system requirements to meet business needs: a case study example. The Computer Journal, 28(2):97–104, May 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/100.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/101.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/102.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/103.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/104.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/97.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_02/tiff/
98.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_28/Issue_02/tiff/99.
tif.

Murtagh:1983:SRA

[Mur83] F. Murtagh. A survey of recent advances in hierarchical clustering algorithms. The Computer Journal, 26(4):354–359, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/354.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/355.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/356.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/357.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/358.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/359.tif.

Murtagh:1985:SAC

[Mur85]  F. Murtagh.  A survey

of algorithms for contiguityconstrained clustering and related problems. The Computer Journal, 28(1):82–88, February 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        01/tiff/82.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_28/Issue_01/tiff/
                                        83.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_28/Issue_01/tiff/84.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        28/Issue_01/tiff/85.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        01/tiff/86.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_28/Issue_01/tiff/
                                        87.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_28/Issue_01/tiff/88.
                                        tif.

Murtagh:1998:FSI

[Mur98] F. Murtagh. Foreword to special issue on clustering and classification. The Computer Journal, 41(8):517, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_41/Issue_08/
                                        Vol41_08.body.html#AbstractEditorial.

Musk:1959:MCS

[Mus59] F. I. Musk. A Monte Carlo simulation of a production planning problem. The Computer Journal, 2(2):90–94, July 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
02/020090.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
02/tiff/90.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_02/tiff/
91.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_02/Issue_02/tiff/92.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
02/Issue_02/tiff/93.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
02/tiff/94.tif.

Musk:1966:OMMb

[Mus66a]  F. I. Musk.  One man’s

meat: part 2 — Now let’s pretend. The Computer Journal, 9(3):213–214, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/090213.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/213.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/214.tif.

Musk:1966:OMMa

[Mus66b]  F. I. Musk.  One man’s

meat: part I — The uses of adversity. The Computer Journal, 9(1):1–2, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        01/090001.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        01/tiff/1.tif;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_09/Issue_01/tiff/
                                        2.tif.

Musk:1967:OMM

[Mus67]  F. I. Musk.  One man’s

meat: part 3 — It’s all good grist that comes to our mill. The Computer Journal, 10(2):126–127, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        02/100126.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        02/tiff/126.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        02/tiff/127.tif.

Muxworthy:1970:NA

[Mux70]  D. T. Muxworthy.  Note

on Algorithm 47: A clustering algorithm. The Computer Journal, 13(2):219, May 1970. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/219.tif. See [Sal70c,

van70, vR70b].

Muxworthy:1973:CCT

[Mux73]  D. T. Muxworthy.  Corre-

spondence: On “Compile-Time Error Diagnostics in Syntax-Directed Compilers”. The Computer Journal, 16(3):285, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/285.tif. See [Bur72].

Maes:1988:RAI

[MV88] R. Maes and J. E. M. Van Dijk. On the role of ambiguity and incompleteness in the design of decision tables and rulebased systems. The Computer Journal, 31(6):481–489, December 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/481.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/482.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/483.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/484.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_31/Issue_
06/tiff/485.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/486.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/487.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/488.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/489.tif.

Misic:1992:FSD

[MVL92] V. Mišić, D. Velašević, and B. Lazarević. Formal specification of a data dictionary for an extended ER data model. The Computer Journal, 35(6):611–622, December 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_06/
Vol35_06.body.html#AbstractMisic.

See [Coo92a].

Mitra:1968:CTO

[MW68] G. Mitra and K. Wolfenden. A computer technique for optimizing the sites and heights of transmission line towers — a dynamic programming approach. The Computer Journal, 10(4):347–351, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/100347.sgm.abs.html;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/347.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/348.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/349.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/350.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/351.tif.

Memon:1997:RDC

[MW97] N. Memon and X. Wu. Recent developments in context-based predictive techniques for lossless image compression. The Computer Journal, 40(2/3):127–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www.oup.co.uk/computer_
journal/Volume_40/Issue_02/
Vol40_02.body.html#AbstractMemon;
http://www3.oup.co.uk/computer_
journal/Volume_40/Issue_02/
Vol40_03.body.html#AbstractMemon.

Morris:1970:SPG

[MWC70] D. Morris, I. R. Wilson, and P. C. Capon. A system program generator. The Computer Journal, 13(3):248–254, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/130248.sgm.abs.html;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/248.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/249.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/250.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/251.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/252.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/253.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/254.tif.

Majewski:1996:FPH

[MWHC96] Bohdan S. Majewski, Nicholas C. Wormald, George Havas, and Zbigniew J. Czech. A family of perfect hashing methods. The Computer Journal, 39(6):547– 554, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www.oup.co.uk/jnls/
list/comjnl/hdb/Volume_39/
Issue_06/390547.sgm.abs.html;
 http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_06/
Vol39_06.body.html#AbstractMajewski.

This paper claims the discovery of order-preserving perfect hashing methods that run in linear time.

Mitchell:1998:DAA

[MWR98] S. E. Mitchell, A. J. Wellings, and A. Romanovsky. Distributed atomic actions in Ada

95. The Computer Journal, 41(7):486–502, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_41/Issue_07/
Vol41_07.body.html#AbstractMitchell.

Mylander:1966:CNP

[Myl66] W. Charles Mylander. Correspondence: Nonlinear programming test problems. The Computer Journal, 8(4):391, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/391.tif. See com-

ments [Box66b].

MacCallum:1986:CST

[MZ86] K. J. MacCallum and J.- M. Zhang. Curve-smoothing techniques using B-splines. The Computer Journal, 29 (6):564–571, December 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/564.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/565.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/566.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_29/Issue_
                                        06/tiff/567.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        06/tiff/568.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        06/tiff/569.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        06/tiff/570.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        06/tiff/571.tif.

McDonald:1996:AVP

[MZ96] A. B. McDonald and T. Znati. ARP versus ES-IS: Performance evaluation of neighbourgreeting protocols. The Computer Journal, 39(10):854–??, ???? 1996. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www.oup.co.uk/computer_
                                        journal/Volume_39/Issue_10/
                                        Vol39_10.body.html#AbstractMcDonald;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_39/Issue_10/
                                        Vol39_10.body.html#AbstractMcDonald.

Naik:1997:DRT

[Nai97]  Kshirasagar Naik.  Design-

ing reliable test architectures for communication protocols. The Computer Journal, 40(7):441–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_40/Issue_07/
                                        Vol40_07.body.html#AbstractNaik.

Napper:1967:SPS

[Nap67]  R. B. E. Napper.  Some

proposals for SNAP, a language with formal macro facilities. The Computer Journal, 10(3):231–243, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/100231.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/231.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/232.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/233.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/234.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/235.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/236.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/237.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/238.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/239.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/240.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_10/Issue_
                                        03/tiff/241.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        03/tiff/242.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        03/tiff/243.tif.

Narayanan:1996:CST

[Nar96] T. S. Narayanan. A class of semi-X tree-based dictionary machines. The Computer Journal, 39(1):45–51, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://www.oup.co.uk/jnls/
                                        list/comjnl/hdb/Volume_39/
                                        Issue_01/390045.sgm.abs.html;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_39/Issue_01/
                                        Vol39_01.body.html#AbstractNaryanan.

Nash:1975:OTM

[Nas75]  J. C. Nash.  A one-sided

transformation method for the singular value decomposition and algebraic eigenproblem. The Computer Journal, 18(1):74–76, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        01/tiff/74.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_18/Issue_01/tiff/
                                        75.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_18/Issue_01/tiff/76.

tif. Cited in Åke Björck’s bibliography on least squares, which is available by anonymous ftp from math.liu.se in

pub/references.

Natarajan:1986:DSS

[Nat86] N. Natarajan. A distributed synchronisation scheme for communicating processes. The Computer Journal, 29(2):109– 117, April 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
29/Issue_02/tiff/109.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/110.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/111.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/112.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/113.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/114.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/115.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/116.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/117.tif.

Naur:1979:CID

[Nau79]  P. Naur.  Critique of the

ISO data processing vocabulary. The Computer Journal, 22(3):286, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/286.tif.

Naur:1993:UTU

[Nau93] P. Naur. Understanding Turing’s universal machine — personal style in program description. The Computer Journal, 36(4):351–372, August 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_04/
Vol36_04.body.html#AbstractNaur.

Navlakha:1987:SSC

[Nav87]  J. K. Navlakha.  A sur-

vey of system complexity metrics. The Computer Journal, 30(3):233–238, June 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/233.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/234.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/235.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/236.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/237.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/238.tif.

Nie:1995:IRC

[NBL95] Jian-Yun Nie, Martin Brisebois, and Françoise Lepage. Information retrieval as counterfactual. The Computer Journal, 38(8):643–657, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_38/Issue_08/
Vol38_08.body.html#AbstractNie.

Ng:1976:RTI

[NC76] T. S. Ng and A. Cantoni. Run time interaction with Fortran using mixed code. The Computer Journal, 19(1):91–92, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
01/tiff/91.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_01/tiff/
92.tif.

Nixon:1997:BPD

[NCR97] Paddy Nixon, Vinny Cahill, and Fethi Rabhi. Building parallel and distributed systems. The Computer Journal, 40(8):463–464, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_

journal/Volume_40/Issue_08/
Vol40_08.body.html#AbstractEditorial.

Neate:1959:SMS

[ND59] R. Neate and W. J. Dacey. A simulation of melting shop operations. The Computer Journal, 2(2):59–67, July 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
02/020059.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
02/tiff/59.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_02/tiff/
60.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_02/Issue_02/tiff/61.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
02/Issue_02/tiff/62.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
02/tiff/63.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_02/tiff/
64.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_02/Issue_02/tiff/65.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
02/Issue_02/tiff/66.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
02/tiff/67.tif.

Nash:1996:SSQ

[NDD96] Jonathan M. Nash, Peter M. Dew, and Martin E. Dyer. A scalable shared queue on a distributed memory machine. The Computer Journal, 39(6):483– 495, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www.oup.co.uk/jnls/
list/comjnl/hdb/Volume_39/
Issue_06/390483.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_06/
Vol39_06.body.html#AbstractNash.

Nathan:1967:ISM

[NE67] A. Nathan and R. K. Even.

The inversion of sparse matrices by a strategy derived from their graphs. The Computer Journal, 10(2):190–194, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/100190.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/190.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/191.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/192.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/193.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/194.tif.

Needham:1963:BRT

[Nee63] R. M. Needham. Book review: Textbook on Mechanized Information Retrieval, by Allen Kent, 1962. The Computer Journal, 6(3):249, October 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/249.tif.

Neff:1965:DIY

[Nef65] B. L. Neff. Do-it-yourself software experience. The Computer Journal, 8(3):201–207, October 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/080201.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/201.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/202.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/203.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/204.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/205.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/206.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/207.tif.

Nelson:1979:CSC

[Nel79] Sallie S. Nelson. CONSIM: a study of control issues in conversational simulation. The Computer Journal, 22(2):119–126, May 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/119.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/120.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/121.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/122.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/123.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/124.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/125.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/126.tif.

Newman:1961:SCC

[New61] E. A. Newman. Some comments on character recognition. The Computer Journal, 4(2):114–120, July 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/040114.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_

                                        02/tiff/114.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        02/tiff/115.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        02/tiff/116.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        02/tiff/117.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        02/tiff/118.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        02/tiff/119.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        02/tiff/120.tif.

Newman:1966:EPA

[New66]  W. M. Newman.  An ex-

perimental program for architectural design. The Computer Journal, 9(1):21–26, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        01/090021.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        01/tiff/21.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_09/Issue_01/tiff/
                                        22.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_09/Issue_01/tiff/23.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        09/Issue_01/tiff/24.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_

01/tiff/25.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_01/tiff/
26.tif.

Newman:1968:GTN

[New68] W. M. Newman. A graphical technique for numerical input. The Computer Journal, 11(1):63–64, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/110063.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/63.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_01/tiff/
64.tif.

Napper:1976:AUE

[NF76] R. B. E. Napper and R. N.

Fisher. ALEC — a user extensible scientific programming language. The Computer Journal, 19(1):25–31, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
01/tiff/25.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_01/tiff/
26.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_19/Issue_01/tiff/27.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
19/Issue_01/tiff/28.tif;
http://www3.oup.co.uk/computer_

journal/hdb/Volume_19/Issue_
01/tiff/29.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_01/tiff/
30.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_19/Issue_01/tiff/31.
tif.

Napper:1980:RUE

[NF80] R. B. E. Napper and R. N. Fisher. RCC — a userextensible systems implementation language. The Computer Journal, 23(3):212–222, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/212.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/213.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/214.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/215.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/216.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/217.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/218.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/219.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_

03/tiff/220.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/221.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/222.tif.

Nievergelt:1994:SIE

[NF94] Jurg Nievergelt and Mike Freeston. Special issue editorial: Other objects, or: What is unique about spatial data? The Computer Journal, 37(1):1–2, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_37/Issue_01/
Vol37_01.body.html#AbstractNievergelt.

Ngoh:1989:TPR

[NH89] L. H. Ngoh and T. P. Hopkins. Transport protocol requirements for distributed multimedia information systems. The Computer Journal, 32(3):252–261, June 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/252.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/253.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/254.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/255.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_32/Issue_
                                        03/tiff/256.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        03/tiff/257.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        03/tiff/258.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        03/tiff/259.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        03/tiff/260.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        03/tiff/261.tif.

Nievergelt:1970:BST

[NI70] J. Nievergelt and M. I. Irland. Bounce-and-skip. A technique for directing the flow of control in programs. The Computer Journal, 13(3):261–262, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/130261.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/261.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/262.tif. See note

[Osb71].

Nicholson:1966:FSR

[Nic66] T. A. J. Nicholson. Finding the shortest route between two points in a network. The Computer Journal, 9(3):275–280, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/090275.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/275.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/276.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/277.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/278.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/279.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/280.tif.

Nievergelt:1965:FVS

[Nie65]  J. Nievergelt.  Fixed ver-

sus selfmodifying programs. The Computer Journal, 8 (3):244–245, October 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/080244.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/244.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/245.tif.

Nipkow:1991:CR

[Nip91] T. Nipkow. Constructive rewriting. The Computer Journal, 34(1):34–41, February 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        01/tiff/34.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_34/Issue_01/tiff/
                                        35.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_34/Issue_01/tiff/36.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        34/Issue_01/tiff/37.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        01/tiff/38.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_34/Issue_01/tiff/
                                        39.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_34/Issue_01/tiff/40.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        34/Issue_01/tiff/41.tif.

Ng:1990:CMA

[NL90] K. W. Ng and H. F. Leung. Competition: A model of AND-Parallel execution of logic programs. The Computer Journal, 33(3):215–218, June 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        03/tiff/215.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_33/Issue_
03/tiff/216.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/217.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/218.tif.

Needham:1974:BRC

[NLW74] R. Needham, F. F. Land, and M. Wells. Book reviews: Computer Data Security, by H. Katzan, Jr., 1974; Information Systems in Management, by R. J. Radford, 1973; Data Transmission, by M. D. Bacon and G. M. Bull, 1973. The Computer Journal, 17(2):191, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/191.tif.

Nelder:1965:SMF

[NM65] J. A. Nelder and R. Mead. A simplex method for function minimization. The Computer Journal, 7(4):308–313, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/
content/7/4/308.full.pdf+
html.

Neelamkavil:1990:SGA

[NM90] F. Neelamkavil and O. Mullarney. Separating graphics from application in the design of user interfaces. The Computer Journal, 33(5):437–443, October 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/437.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/438.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/439.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/440.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/441.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/442.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/443.tif.

Naor:1991:PPQ

[NMM91] D. Naor, C. U. Martel, and N. S. Matloff. Performance of priority queue structures in a virtual memory environment. The Computer Journal, 34(5):428–437, October 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/428.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/429.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_

05/tiff/430.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/431.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/432.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/433.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/434.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/435.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/436.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/437.tif.

Nevill-Manning:1997:CEU

[NMW97] C. G. Nevill-Manning and I. H. Witten. Compression and explanation using hierarchical grammars. The Computer Journal, 40(2/3):103–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www.oup.co.uk/computer_
journal/Volume_40/Issue_02/
Vol40_02.body.html#AbstractNevill;
http://www3.oup.co.uk/computer_
journal/Volume_40/Issue_02/
Vol40_02.body.html#AbstractNevill;
http://www3.oup.co.uk/computer_
journal/Volume_40/Issue_02/
Vol40_03.body.html#AbstractNevill.

Novack:1997:RDL

[NN97] Steven Novack and Alexandru Nicolau. Resource-directed loop pipelining: Exposing just enough parallelism. The Computer Journal, 40(6):311–321, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_40/Issue_06/
Vol40_06.body.html#AbstractNovack.

Neelamkavil:1992:IEP

[NO92a] F. Neelamkavil and S. O’Shea. Interfacing Eiffel and PCTE. The Computer Journal, 35 (5):A439–??, October 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_35/Issue_05/
35000A439.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/Volume_35/Issue_05/
tiff/A439.tif.

Ng:1992:SNS

[NO92b] D. T. H. Ng and B. J. Oommen. Short notes: A short note on doubly-linked list re-organizing heuristics. The Computer Journal, 35(5):533–535, October 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_35/Issue_05/
350533.sgm.abs.html;   http:
//www3.oup.co.uk/computer_
journal/Volume_35/Issue_05/
tiff/533.tif;   http://www3.
oup.co.uk/computer_journal/
Volume_35/Issue_05/tiff/534.
tif;   http://www3.oup.co.

uk/computer_journal/Volume_
35/Issue_05/tiff/535.tif.

Noble:1991:SWD

[Nob91] F. Noble. Seven ways to develop office systems: A managerial comparison of office system development methodologies. The Computer Journal, 34(2):113–121, April 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/113.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/114.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/115.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/116.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/117.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/118.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/119.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/120.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/121.tif.

Noor:1984:FCM

[Noo84] A. El Sayed Noor. A framework for the creation and management of national computing strategies in developing countries. The Computer Journal, 27(3):193–200, August 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/193.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/194.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/195.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/196.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/197.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/198.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/199.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/200.tif.

Norton:1964:ISN

[Nor64]  H. J. Norton.  The itera-

tive solution of non-linear ordinary differential equations in Chebyshev series. The Computer Journal, 7(1):76–85, April 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
01/070076.sgm.abs.html;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
01/tiff/76.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_01/tiff/
77.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_07/Issue_01/tiff/78.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
07/Issue_01/tiff/79.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
01/tiff/80.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_01/tiff/
81.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_07/Issue_01/tiff/82.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
07/Issue_01/tiff/83.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
01/tiff/84.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_01/tiff/
85.tif.

Norman:1976:ESI

[Nor76] A. C. Norman. Expanding the solutions of implicit sets of ordinary differential equations in power series. The Computer Journal, 19(1):63–68, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
01/tiff/63.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_01/tiff/
64.tif;   http://www3.oup.

                                        co.uk/computer_journal/hdb/
                                        Volume_19/Issue_01/tiff/65.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        19/Issue_01/tiff/66.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        01/tiff/67.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_19/Issue_01/tiff/
                                        68.tif.

Notley:1970:CRL

[Not70]  M. G. Notley.  The cu-

mulative recurrence library. The Computer Journal, 13 (1):14–19, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        01/130014.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        01/tiff/14.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_13/Issue_01/tiff/
                                        15.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_13/Issue_01/tiff/16.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        13/Issue_01/tiff/17.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        01/tiff/18.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_13/Issue_01/tiff/
                                        19.tif.

Nicholson:1968:PPJ

[NP68] T. A. J. Nicholson and R. D. Pullen. A permutation procedure for job-shop scheduling. The Computer Journal, 11(1):48–56, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/110048.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/48.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_01/tiff/
49.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_11/Issue_01/tiff/50.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/51.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/52.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_01/tiff/
53.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_11/Issue_01/tiff/54.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/55.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/56.tif.

Neal:1990:YMC

[NP90] L. R. Neal and M. L. V.

Pitteway. Yet more circle generators. The Computer Journal, 33(5):408–411, October 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

                                        journal/hdb/Volume_33/Issue_
                                        05/tiff/408.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        05/tiff/409.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        05/tiff/410.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        05/tiff/411.tif.

Newton:1981:BRP

[NPCA81] R. W. Newton, F. Panzieri, C. C. Chang, and A. M. Addyman. Book reviews: Programming Language Standardisation, edited by I. D. Hill and B. L. Meek, 1980; Problem Solving and Structured Programming, by E. B. Koffmann, 1981; Data Base: Structured Techniques for Design, Performance and Management, by S. Atre, 1980; Pascal — The Language and its Implementation, edited by D. W. Barron, 1981. The Computer Journal, 24(4):384, November 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        04/tiff/384.tif.

Negri:1989:SPU

[NPS89] M. Negri, G. Pelagatti, and L. Sbattella. Semantics and problems of universal quantification in SQL. The Computer Journal, 32(1):90–91, February 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
01/tiff/90.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_01/tiff/
91.tif.

Noronha:1968:HCD

[NPW68] L. G. Noronha, C. Y. Po, and J. W. Womack. Hybrid computation of the dynamics of a distributed system. The Computer Journal, 11(2):196–205, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/110196.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/196.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/197.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/198.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/199.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/200.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/201.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/202.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/203.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_11/Issue_
                                        02/tiff/204.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        02/tiff/205.tif.

Newton:1980:BRS

[NQ80] Roy Newton and Harley Quilliam. Book reviews: Software Portability: an Advanced Course, edited by P. J. Brown, 1977; The Challenge of Microprocessors, by M. G. Hartley and Anne Buckley, 1979. The Computer Journal, 23(2):160, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        02/tiff/160.tif.

Nessett:1979:AAD

[NR79] D. M. Nessett and O. W. Rechard. Allocation algorithms for dynamically microprogrammable multiprocessor systems. The Computer Journal, 22(2):155–163, May 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/155.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/156.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/157.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/158.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/159.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/160.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/161.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/162.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/163.tif.

Nevalainen:1992:IHS

[NR92] O. Nevalainen and T. Raita.

An internal hybrid sort algorithm revisited. The Computer Journal, 35(2):177–183, April 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_35/Issue_02/
Vol35_02.body.html#AbstractNevalainen.

Nicklas:1977:ISI

[NS77] B. M. Nicklas and G. Schlageter.

Index structuring in inverted data bases by TRIES. The Computer Journal, 20(4):321– 324, November 1977. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
20/Issue_04/tiff/321.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/322.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_

04/tiff/323.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/324.tif.

Nash:1987:SAP

[NS87] J. C. Nash and S. Shlien. Simple algorithms for the partial singular value decomposition. The Computer Journal, 30(3):268–275, June 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/268.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/269.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/270.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/271.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/272.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/273.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/274.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/275.tif.

Newman:1987:HMP

[NSW87] I. A. Newman, R. P. Stallard, and M. C. Woodward. A hybrid multiple processor garbage collection algorithm. The Computer Journal, 30(2):119–127, April 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/119.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/120.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/121.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/122.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/123.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/124.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/125.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/126.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/127.tif.

Nevalainen:1979:PQA

[NT79] O. Nevalainen and J. Teuhola.

Priority queue administration by sublist index. The Computer Journal, 22(3):220–225, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/220.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_22/Issue_
                                        03/tiff/221.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        03/tiff/222.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        03/tiff/223.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        03/tiff/224.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        03/tiff/225.tif.

Nudds:1969:MCE

[Nud69] D. Nudds. Methods of computing event times in project networks. The Computer Journal, 12(1):38–40, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        01/120038.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        01/tiff/38.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_12/Issue_01/tiff/
                                        39.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_12/Issue_01/tiff/40.
                                        tif.

Nudds:1977:DMM

[Nud77]  D. Nudds.  The design

of the MAX macroprocessor. The Computer Journal, 20(1):30–36, February 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
01/tiff/30.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_01/tiff/
31.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_20/Issue_01/tiff/32.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
20/Issue_01/tiff/33.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
01/tiff/34.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_01/tiff/
35.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_20/Issue_01/tiff/36.
tif.

Nevalainen:1977:DBF

[NV77] O. Nevalainen and M. Vesterinen. Determining blocking factors for sequential files by heuristic methods. The Computer Journal, 20 (3):245–247, August 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/245.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/246.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/247.tif.

Needham:1974:DPM

[NW74] R. M. Needham and M. V.

Wilkes. Domains of protection and the management of processes. The Computer Journal, 17(2):117–120, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/117.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/118.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/119.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/120.tif.

Nedunuri:1987:DRS

[NW87] S. Nedunuri and N. E. Wiseman. Displaying random surfaces. The Computer Journal, 30(2):163–167, April 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/163.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/164.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/165.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/166.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/167.tif.

Nievergelt:1993:GFS

[NW93] J. Nievergelt and P. Widmayer. Guard files: Stabbing and intersection queries on fat spatial objects. The Computer Journal, 36(2):107–116, April 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_02/
Vol36_02.body.html#AbstractNievergelt.

OBeirne:1959:CGG

[O’B59a] Thos. H. O’Beirne. Correspondence: Gray or gros? The Computer Journal, 2(2):96, July 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
02/tiff/96.tif.

OBrien:1959:NDF

[O’B59b]  J. P. O’Brien.  Note on

decomposition into first order of multi-order linear differential equations with constant coefficients. The Computer Journal, 2(3):144, October 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/020144.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/144.tif.

OBrien:1967:CCO

[O’B67] W. O’Brien. COSMOS — Courtauld’s own system for matrix operations and statistics. The Computer Journal, 10(2):128–134, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/100128.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/128.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/129.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/130.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/131.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/132.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/133.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/134.tif.

OBrien:1968:TCF

[OB68a] F. O’Brien and R. C. Beckwith. A technique for computer flow chart generation. The Computer Journal, 11(2):138–140, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/110138.sgm.abs.html;
http://www3.oup.co.uk/computer_

journal/hdb/Volume_11/Issue_
02/tiff/138.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/139.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/140.tif.

OBrien:1968:CNM

[O’B68b] J. P. O’Brien. Correspondence: On “A new method for solving polynomial equations”. The Computer Journal, 11(2):194, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/194.tif. See [GJM68].

OBeirne:1970:CPW

[O’B70] T. H. O’Beirne. Computer programs which play music with microtones. The Computer Journal, 13(4):344–349, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/130344.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/344.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/345.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/346.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_

04/tiff/347.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/348.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/349.tif.

Osborn:1975:SPS

[OBLP75] R. A. Osborn, W. P. Bain, T. Lloyd, and P. H. Perring. SPG — a programming system for commercial transaction processing. The Computer Journal, 18(4):290–297, November 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/290.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/291.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/292.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/293.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/294.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/295.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/296.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/297.tif.

Obruca:1967:CBT

[Obr67] A. K. Obruca. Correspondence between two types of trees. The Computer Journal, 10(3):290–292, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/100290.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/290.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/291.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/292.tif.

Obruca:1968:STM

[Obr68]  A. K. Obruca.  Spanning

tree manipulation and the travelling salesman problem. The Computer Journal, 10 (4):374–377, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/100374.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/374.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/375.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/376.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_10/Issue_
                                        04/tiff/377.tif.

Obruca:1978:ENR

[Obr78a]  A. Obruča.  Algorithm 1:

Mintree. The Computer Journal, 21(3):276, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        03/tiff/276.tif. Reprinted

from The Computer Bulletin, September 1964, p. 67. See notes [Boo78, HO78, Obr78b, Pit78, RT78, WT79].

Obruca:1978:RAM

[Obr78b] Alex K. Obruča. Remark on Algorithm 1: Mintree. The Computer Journal, 21(3):277, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        03/tiff/277.tif. See [Obr78a].

OCarroll:1969:DCR

[O’C69] M. J. O’Carroll. Discussion and correspondence: S.O.R. and membranes. The Computer Journal, 12(3):296–297, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        03/120296.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        03/tiff/296.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/297.tif.

ODocherty:1991:MIS

[OD91] M. H. O’Docherty and C. N.

Daskalakis. Multimedia information systems — the management and semantic retrieval of all electronic data types. The Computer Journal, 34(3):225–238, June 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/225.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/226.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/227.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/228.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/229.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/230.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/231.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/232.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/233.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_

                                        03/tiff/234.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        03/tiff/235.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        03/tiff/236.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        03/tiff/237.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        03/tiff/238.tif.

Oommen:1993:FLA

[OF93] B. John Oommen and Chris Fothergill. Fast learning automaton-based image examination and retrieval. The Computer Journal, 36(6):542– 553, December 1993. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/Volume_
                                        36/Issue_06/Vol36_06.body.
                                        html#AbstractOommen.

Halsbury:1959:TYC

[oH59] Rt. Hon. The Earl of Halsbury. Ten years of computer development. The Computer Journal, 1(4):153–159, January 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        04/010153.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        04/tiff/153.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_

04/tiff/154.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/tiff/155.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/tiff/156.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/tiff/157.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/tiff/158.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/tiff/159.tif.

Ortega:1963:MST

[OK63] J. M. Ortega and H. F. Kaiser.

The LL^{T} and QR methods for symmetric tridiagonal matrices. The Computer Journal, 6(1):99–101, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/060099.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/tiff/100.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/tiff/101.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/tiff/99.tif.

OKeefe:1985:CAE

[O’K85] R. M. O’Keefe. Comment on ‘Complexity analyses of event set algorithms’. The Computer Journal, 28(5):496–497, November 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        05/tiff/496.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        05/tiff/497.tif.

Oka:1998:SMC

[Oka98] Ryuichi Oka. Spotting method for classification of real world data. The Computer Journal, 41(8):559–565, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_41/Issue_08/
                                        Vol41_08.body.html#AbstractOka.

Ould-Khaoua:1996:CEH

[OKMS96] M. Ould-Khaoua, L. M. Mackenzie, and R. Sotudeh. Comparative evaluation of hypermesh and multi-stage interconnection networks. The Computer Journal, 39(3):232–240, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://www.oup.co.uk/jnls/
                                        list/comjnl/hdb/Volume_39/
                                        Issue_03/390232.sgm.abs.html;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_39/Issue_03/
                                        Vol39_03.body.html#AbstractOuld.

Owen:1968:CAA

[OL68] David P. Owen and John Laski. Correspondence: An argument against paging hardware. The Computer Journal, 11(3):351, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/351.tif.

Osborne:1978:ISP

[OL78] M. Osborne and F. Little.

An implementation of structured programming in APL. The Computer Journal, 21 (4):311–315, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/311.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/312.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/313.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/314.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/315.tif.

Olerup:1991:DAC

[Ole91] A. Olerup. Design approaches:

A comparative study of information system design and architectural design. The Computer Journal, 34(3):215–224, June 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

journal/hdb/Volume_34/Issue_
03/tiff/215.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/216.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/217.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/218.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/219.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/220.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/221.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/222.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/223.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/224.tif.

Oliver:1969:EET

[Oli69] J. Oliver. An error estimation technique for the solution of ordinary differential equations in Chebyshev series. The Computer Journal, 12(1):57–62, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
01/120057.sgm.abs.html;
http://www3.oup.co.uk/computer_

journal/hdb/Volume_12/Issue_
01/tiff/57.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_01/tiff/
58.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_12/Issue_01/tiff/59.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/60.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
01/tiff/61.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_01/tiff/
62.tif.

Oliver:1971:EDI

[Oli71] J. Oliver. The evaluation of definite integrals using highorder formulae. The Computer Journal, 14(3):301–306, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/140301.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/301.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/302.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/303.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/304.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/305.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_14/Issue_
                                        03/tiff/306.tif.

Oliver:1972:DAC

[Oli72] J. Oliver. A doubly-adaptive Clenshaw–Curtis quadrature method. The Computer Journal, 15(2):141–147, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        02/150141.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        02/tiff/141.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        02/tiff/142.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        02/tiff/143.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        02/tiff/144.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        02/tiff/145.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        02/tiff/146.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        02/tiff/147.tif.

Osborne:1964:NSE

[OM64] M. R. Osborne and S. Michaelson. The numerical solution of eigenvalue problems in which the eigenvalue problems in which the eigenvalue parameter appears nonlinearly, with an application to differential equations. The Computer Journal, 7(1):66–71, April 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
01/070066.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
01/tiff/66.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_01/tiff/
67.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_07/Issue_01/tiff/68.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
07/Issue_01/tiff/69.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
01/tiff/70.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_01/tiff/
71.tif.

OwrangO:1988:QTB

[OM88] M. M. Owrang O. and L. L.

Miller. Query translation based on hypergraph models. The Computer Journal, 31(2):155–164, April 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/155.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/156.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/157.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        02/tiff/158.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        02/tiff/159.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        02/tiff/160.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        02/tiff/161.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        02/tiff/162.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        02/tiff/163.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        02/tiff/164.tif.

Oommen:1991:SAS

[OM91] B. J. Oommen and D. C. Y. Ma. Stochastic automata solutions to the object partitioning problem. The Computer Journal, 34(6):A105–A120, December 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Oliver:1979:PEA

[ON79] J. R. Oliver and R. S. Newton. Practical experience with Algol 68-Rt. The Computer Journal, 22(2):114–118, May 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/114.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_22/Issue_
02/tiff/115.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/116.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/117.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/118.tif.

Oommen:1990:GRP

[ON90] B. J. Oommen and D. T. H.

Ng. On generating random permutations with arbitrary distributions. The Computer Journal, 33(4):368–374, August 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/368.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/369.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/370.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/371.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/372.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/373.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/374.tif.

ONeill:1992:SNA

[O’N92] G. O’Neill. Short note: Automatic translation of VDM specifications into Standard ML programs. The Computer Journal, 35(6):623–624, December 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_35/Issue_
                                        06/tiff/623.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_35/Issue_
                                        06/tiff/624.tif.

Oommen:1994:NTE

[ON94] B. J. Oommen and D. T. H. Ng. A new technique for enhancing linked-list data retrieval: reorganize data using artificially synthesized queries. The Computer Journal, 37(7):598–609, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_37/Issue_07/
                                        Vol37_07.body.html#AbstractOommen.

Ongley:1969:EC

[Ong69] P. A. Ongley. Examinations by computer. The Computer Journal, 12(3):297–300, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        03/120297.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        03/tiff/297.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/298.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/299.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/300.tif.

Oommen:1993:TIC

[Oom93]  B. J. Oommen.  Trans-

forming ill-conditioned constrained problems using projections. The Computer Journal, 36(3):282–285, June 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_36/Issue_03/
Vol36_03.body.html#AbstractOommen.

Olariu:1991:MDP

[OOW91] S. Olariu, C. M. Overstreet, and Z. Wenn. A mergeable double-ended priority queue. The Computer Journal, 34 (5):423–427, October 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/423.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/424.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/425.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/426.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        05/tiff/427.tif.

Oliveira-Pinto:1971:ATC

[OP71] F. Oliveira-Pinto. Algorithm 69: Trigonometric curve fitting to equally or unequally spaced data. The Computer Journal, 14(2):213–214, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/213.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/214.tif.

Oliveira-Pinto:1973:GCP

[OP73a]  F. Oliveira-Pinto.  Gener-

alised Chebyshev polynomials and their use in numerical approximation. The Computer Journal, 16(4):375–379, November 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        04/160375.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        04/tiff/375.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        04/tiff/376.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        04/tiff/377.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        04/tiff/378.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/379.tif.

Oliveira-Pinto:1973:STA

[OP73b] F. Oliveira-Pinto. Simultaneous trigonometric approximation of the function and its first derivative. The Computer Journal, 16(1):73–76, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
01/160073.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
01/tiff/73.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_01/tiff/
74.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_16/Issue_01/tiff/75.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
16/Issue_01/tiff/76.tif.

Oliveira-Pinto:1975:UDO

[OP75] F. Oliveira-Pinto. The use of derivatives in the optimisation of the truncation error of polynomial approximations. The Computer Journal, 18(1):77–80, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
01/tiff/77.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_01/tiff/
78.tif;   http://www3.oup.

                                        co.uk/computer_journal/hdb/
                                        Volume_18/Issue_01/tiff/79.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        18/Issue_01/tiff/80.tif.

ORegan:1970:SSA

[O’R70] P. G. O’Regan. Step size adjustment at discontinuities for fourth order Runge–Kutta methods. The Computer Journal, 13(4):401–404, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/130401.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/401.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/402.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/403.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/404.tif. See correc-

tion [Whi72].

Orden:1964:SPQ

[Ord64] A. Orden. Stationary points of quadratic functions under linear constraints. The Computer Journal, 7(3):238–241, October 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        03/070238.sgm.abs.html;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/238.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/239.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/240.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/241.tif.

Orman:1988:FDC

[Orm88] L. Orman. Functional database constraints. The Computer Journal, 31(4):336–343, August 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/336.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/337.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/338.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/339.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/340.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/341.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/342.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/343.tif.

Orman:1991:CMD

[Orm91] L. Orman. Constraint maintenance as a data model design criterion. The Computer Journal, 34(1):73–79, February 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        01/tiff/73.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_34/Issue_01/tiff/
                                        74.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_34/Issue_01/tiff/75.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        34/Issue_01/tiff/76.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        01/tiff/77.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_34/Issue_01/tiff/
                                        78.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_34/Issue_01/tiff/79.
                                        tif.

Ortony:1971:SSV

[Ort71]  A. Ortony.  A system for

stereo viewing. The Computer Journal, 14(2):140–144, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/140140.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/140.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_14/Issue_
02/tiff/141.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/142.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/143.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/144.tif.

Ord-Smith:1965:EBD

[OS65] R. J. Ord-Smith. An extension of block design methods and an application in the construction of redundant fault reducing circuits for computers. The Computer Journal, 8(1):28–32, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
01/tiff/28.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_01/tiff/
29.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_08/Issue_01/tiff/30.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
08/Issue_01/tiff/31.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
01/tiff/32.tif.

OHara:1968:EEC

[OS68] H. O’Hara and F. J. Smith. Error estimation in the Clenshaw– Curtis quadrature formula. The Computer Journal, 11 (2):213–219, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/110213.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/213.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/214.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/215.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/216.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/217.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/218.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/219.tif.

OHara:1969:EDI

[OS69] H. O’Hara and Francis J. Smith. The evaluation of definite integrals by interval subdivision. The Computer Journal, 12(2):179–182, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/120179.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/179.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_

02/tiff/180.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/181.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/182.tif.

Ord-Smith:1970:GPS

[OS70] R. J. Ord-Smith. Generation of permutation sequences: part 1. The Computer Journal, 13(2):152–155, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/130152.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/152.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/153.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/154.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/155.tif.

Ord-Smith:1971:GPS

[OS71] R. J. Ord-Smith. Generation of permutation sequences: Part 2. The Computer Journal, 14(2):136–139, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/140136.sgm.abs.html;
http://www3.oup.co.uk/computer_

journal/hdb/Volume_14/Issue_
02/tiff/136.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/137.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/138.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/139.tif.

OShea:1988:CDU

[O’S88] G. O’Shea. Controlling the dependency of user access control mechanisms on correctness of user identification. The Computer Journal, 31(6):503–509, December 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/503.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/504.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/505.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/506.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/507.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/508.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/509.tif.

OShea:1994:SVV

[O’S94] Greg O’Shea. On the specification, validation and verification of security in access control systems. The Computer Journal, 37(5):437–448, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_37/Issue_05/
Vol37_05.body.html#AbstractO'
Shea.

Osborne:1963:IPS

[Osb63]  M. R. Osborne.  Iterative

procedures for solving finitedifference approximations to separable partial differential equations. The Computer Journal, 6(1):93–99, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/060093.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/tiff/93.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/
94.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_06/Issue_01/tiff/95.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
06/Issue_01/tiff/96.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/tiff/97.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/

                                        98.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_06/Issue_01/tiff/99.
                                        tif.

Osborne:1964:EAF

[Osb64a] M. R. Osborne. An error analysis of finite-difference methods for the numerical solution of ordinary differential equations. The Computer Journal, 7(3):232–237, October 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        03/070232.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        03/tiff/232.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        03/tiff/233.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        03/tiff/234.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        03/tiff/235.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        03/tiff/236.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        03/tiff/237.tif.

Osborne:1964:MFD

[Osb64b] M. R. Osborne. A method for finite-difference approximation to ordinary differential equations. The Computer Journal, 7(1):58–65, April 1964. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
01/070058.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
01/tiff/58.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_01/tiff/
59.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_07/Issue_01/tiff/60.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
07/Issue_01/tiff/61.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
01/tiff/62.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_01/tiff/
63.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_07/Issue_01/tiff/64.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
07/Issue_01/tiff/65.tif.

Osborne:1964:NMS

[Osb64c]  M. R. Osborne.  A new

method for the solution of eigenvalue problems. The Computer Journal, 7(3):228–232, October 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/070228.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/228.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_

                                        03/tiff/229.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        03/tiff/230.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        03/tiff/231.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        03/tiff/232.tif.

Osborne:1965:DMS

[Osb65] M. R. Osborne. Direct methods for the solution of finitedifference approximations to separable partial differential equations. The Computer Journal, 8(2):150–156, July 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        02/080150.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        02/tiff/150.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        02/tiff/151.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        02/tiff/152.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        02/tiff/153.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        02/tiff/154.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        02/tiff/155.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_08/Issue_
02/tiff/156.tif.

Osborne:1969:NSH

[Osb69]  M. R. Osborne.  The nu-

merical solution of the heat conduction equation subject to separated boundary conditions. The Computer Journal, 12(1):82–87, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
01/120082.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
01/tiff/82.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_01/tiff/
83.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_12/Issue_01/tiff/84.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/85.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
01/tiff/86.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_01/tiff/
87.tif.

Osborne:1971:CBS

[Osb71] G. W. Osborne. Correspondence: On “Bounce-and-skip. A technique for directing the flow of control in programs”. The Computer Journal, 14(2):155, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/155.tif. See [NI70].

Osborne:1993:PIL

[Osb93]  Chris Osborne.  Planning

for IT literacy in an institution of higher education: a case study. The Computer Journal, 36(4):343–350, August 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_36/Issue_04/
                                        Vol36_04.body.html#AbstractOsborne.

ONeill:1969:PCF

[OSS69] M. O’Neill, I. G. Sinclair, and F. J. Smith. Polynomial curve fitting when abscissas and ordinates are both subject to error. The Computer Journal, 12(1):52–56, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        01/120052.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        01/tiff/52.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_12/Issue_01/tiff/
                                        53.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_12/Issue_01/tiff/54.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        12/Issue_01/tiff/55.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        01/tiff/56.tif.

Osterby:1979:AES

[Øst79] Ole Østerby. Algorithm 108:

Efficient solution of tridiagonal linear systems. The Computer Journal, 22(3):283–285, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/283.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/284.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/285.tif.

Ou:1991:HSU

[OT91] Seng Fat Ou and Alan L.

Tharp. High storage utilisation for single-probe retrieval linear hashing. The Computer Journal, 34(5):455–468, October 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/455.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/456.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/457.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/458.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/459.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_34/Issue_
05/tiff/460.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/461.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/462.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/463.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/464.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/465.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/466.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/467.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/468.tif.

Oulsnam:1982:UUP

[Oul82] G. Oulsnam. Unravelling unstructured programs. The Computer Journal, 25(3):379– 387, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
25/Issue_03/tiff/379.tif;
 http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/380.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/381.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_25/Issue_
03/tiff/382.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/383.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/384.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/385.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/386.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/387.tif.

Oulsnam:1987:ATS

[Oul87] G. Oulsnam. The algorithmic transformation of schemas to structured form. The Computer Journal, 30(1):43–51, February 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
01/tiff/43.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_01/tiff/
44.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_30/Issue_01/tiff/45.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
30/Issue_01/tiff/46.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
01/tiff/47.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_01/tiff/
48.tif;   http://www3.oup.
co.uk/computer_journal/hdb/

                                        Volume_30/Issue_01/tiff/49.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        30/Issue_01/tiff/50.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        01/tiff/51.tif.

Oman:1992:UBS

[OV92] P. W. Oman and K. Van Houten. On the use of binary strings to represent tree structures. The Computer Journal, 35(5):A463–??, October 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_35/Issue_05/
                                        35000A463.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/Volume_35/Issue_05/
                                        tiff/A463.tif.

Overholt:1967:AGa

[Ove67a] K. J. Overholt. Algorithm 16: GOLD. The Computer Journal, 9(4):415, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        04/tiff/415.tif.

Overholt:1967:AGb

[Ove67b] K. J. Overholt. Algorithm 17: GOLDSEC. The Computer Journal, 9(4):415, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_

journal/hdb/Volume_09/Issue_
04/tiff/415.tif.

Overholt:1967:NAF

[Ove67c] K. J. Overholt. Note on Algorithm 2: Fibonacci search. The Computer Journal, 9(4): 414, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
09/Issue_04/tiff/414.tif.

See [PP79].

OHalloran:1966:NRI

[OW66] M. O’Halloran and W. M.

Waite. Note on rapid instruction analysis by table lookup. The Computer Journal, 9(3):248, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/090248.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/248.tif.

Osborne:1967:BLC

[OW67] M. R. Osborne and G. A.

Watson. On the best linear Chebyshev approximation. The Computer Journal, 10(2):172–177, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/100172.sgm.abs.html;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/172.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/173.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/174.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/175.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/176.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/177.tif.

Osborne:1968:NTM

[OW68] M. R. Osborne and G. A. Watson. Note on two methods of solving ordinary linear differential equations. The Computer Journal, 10(4):383–384, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/100383.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/383.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/384.tif.

Osborne:1969:AMA

[OW69] M. R. Osborne and G. A. Watson. An algorithm for minimax approximation in the nonlinear case. The Computer Journal, 12(1):63–68, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
01/120063.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
01/tiff/63.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_01/tiff/
64.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_12/Issue_01/tiff/65.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/66.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
01/tiff/67.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_01/tiff/

68.tif. See correspondence [Tem69].

Osborne:1971:ADN

[OW71] M. R. Osborne and G. A.

Watson. On an algorithm for discrete nonlinear L_{1} approximation. The Computer Journal, 14(2):184–188, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/140184.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/184.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/185.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/186.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/187.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/188.tif. See corre-

spondence [DW73].

Ottmann:1980:BTA

[OW80] T. Ottmann and D. Wood. 1 − 2 brother trees or AVL trees revisited. The Computer Journal, 23(3):248–255, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        03/tiff/248.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        03/tiff/249.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        03/tiff/250.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        03/tiff/251.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        03/tiff/252.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        03/tiff/253.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        03/tiff/254.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        03/tiff/255.tif.

Oliver:1983:OQE

[OW83a] M. A. Oliver and N. E. Wiseman. Operations on quadtree encoded images. The Computer Journal, 26(1):83–91, February 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
01/tiff/83.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_01/tiff/
84.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_26/Issue_01/tiff/85.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
26/Issue_01/tiff/86.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
01/tiff/87.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_01/tiff/
88.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_26/Issue_01/tiff/89.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
26/Issue_01/tiff/90.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
01/tiff/91.tif.

Oliver:1983:OQL

[OW83b] M. A. Oliver and N. E. Wiseman. Operations on quadtree leaves and related image areas. The Computer Journal, 26(4):375–380, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        04/tiff/375.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        04/tiff/376.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        04/tiff/377.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        04/tiff/378.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        04/tiff/379.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        04/tiff/380.tif.

Oxley:1985:SFT

[Oxl85]  A. Oxley.  Surface fitting

by triangulation. The Computer Journal, 28(3):335–339, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        03/tiff/335.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        03/tiff/336.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        03/tiff/337.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        03/tiff/338.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        03/tiff/339.tif.

Oxley:1989:CEV

[Oxl89] A. Oxley. Calculating earthworks volumes — an exercise in geometric algorithms. The Computer Journal, 32 (4):377–379, August 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/377.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/378.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/379.tif.

Philbrow:1990:EEH

[PA90] P. C. Philbrow and M. P. Atkinson. Events and exception handling in PS-algol. The Computer Journal, 33(2):108–125, April 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/108.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/109.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/110.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/111.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/112.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/113.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_33/Issue_
                                        02/tiff/114.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        02/tiff/115.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        02/tiff/116.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        02/tiff/117.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        02/tiff/118.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        02/tiff/119.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        02/tiff/120.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        02/tiff/121.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        02/tiff/122.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        02/tiff/123.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        02/tiff/124.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        02/tiff/125.tif.

Paakki:1991:PPC

[Paa91] J. Paakki. Prolog in practical compiler writing. The Computer Journal, 34(1):64– 72, February 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_
34/Issue_01/tiff/64.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
01/tiff/65.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_01/tiff/
66.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_34/Issue_01/tiff/67.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
34/Issue_01/tiff/68.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
01/tiff/69.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_01/tiff/
70.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_34/Issue_01/tiff/71.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
34/Issue_01/tiff/72.tif.

Pachl:1988:TLP

[Pac88] J. Pachl. Two livelock-free protocols for slotted rings. The Computer Journal, 31 (5):474–476, October 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/474.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/475.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/476.tif.

Padgett:1973:TDD

[Pad73] M. I. Padgett. Tree driven data input and its validation. The Computer Journal, 16(4):315–321, November 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        04/160315.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        04/tiff/315.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        04/tiff/316.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        04/tiff/317.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        04/tiff/318.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        04/tiff/319.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        04/tiff/320.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        04/tiff/321.tif.

Padawitz:1991:RNH

[Pad91] P. Padawitz. Reduction and narrowing for Horn clause theories. The Computer Journal, 34(1):42–51, February 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_

01/tiff/42.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_01/tiff/
43.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_34/Issue_01/tiff/44.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
34/Issue_01/tiff/45.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
01/tiff/46.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_01/tiff/
47.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_34/Issue_01/tiff/48.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
34/Issue_01/tiff/49.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
01/tiff/50.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_01/tiff/
51.tif.

Page:1958:NR

[Pag58] E. S. Page. A note on round-off.

The Computer Journal, 1(1):10, April 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
01/010011.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
01/tiff/10.tif.

Page:1960:TCR

[Pag60] E. S. Page. Theoretical considerations of routine maintenance. The Computer Journal, 2(4):199–204, January 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/020199.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/199.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/200.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/201.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/202.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/203.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/204.tif.

Page:1962:SJC

[Pag62] E. S. Page. On the scheduling of jobs by computer. The Computer Journal, 5 (3):214–220, October 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/050214.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/214.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/215.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_05/Issue_
03/tiff/216.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/217.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/218.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/219.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/220.tif.

Page:1963:NAP

[Pag63] E. S. Page. A note on assignment problems. The Computer Journal, 6(3):241–243, October 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/060241.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/241.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/242.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/243.tif.

Page:1971:SGO

[Pag71] E. S. Page. Systematic generation of ordered sequences using recurrence relations. The Computer Journal, 14(2):150–153, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

                                        journal/hdb/Volume_14/Issue_
                                        02/140150.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/150.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/151.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/152.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/153.tif.

Page:1973:EJW

[Pag73] E. S. Page. Editorial: the journal — why and what for. The Computer Journal, 16(4):290, November 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        04/160290.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        04/tiff/290.tif.

Pagan:1976:IDP

[Pag76] Frank G. Pagan. On interpreteroriented definitions of programming languages. The Computer Journal, 19(2):151–155, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/151.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/152.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/153.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/154.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/155.tif.

Pagan:1979:AMD

[Pag79] F. G. Pagan. Algol 68 as a metalanguage for denotational semantics. The Computer Journal, 22(1):63–66, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
01/tiff/63.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_01/tiff/
64.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_22/Issue_01/tiff/65.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
22/Issue_01/tiff/66.tif.

Pagan:1981:SWS

[Pag81] F. G. Pagan. Style for writing the syntactic portions of complete definitions of programming languages. The Computer Journal, 24(2):143–147, May 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/143.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_

                                        02/tiff/144.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        02/tiff/145.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        02/tiff/146.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        02/tiff/147.tif.

Pagan:1982:SSU

[Pag82a]  Frank G. Pagan.  Seman-

tic specification using two-level grammars: labels and GOTO statements. The Computer Journal, 25(1):63–67, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/63.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_25/Issue_01/tiff/
                                        64.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_25/Issue_01/tiff/65.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        25/Issue_01/tiff/66.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/67.tif.

Page:1982:OFA

[Pag82b]  I. P. Page.  Optimal fit

of arbitrary sized segments. The Computer Journal, 25 (1):32–33, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_25/Issue_
01/tiff/32.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/
33.tif.

Page:1984:ACP

[Pag84]  I. P. Page.  Analysis of

a cyclic placement scheme. The Computer Journal, 27 (1):18–26, February 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
01/tiff/18.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_01/tiff/
19.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_27/Issue_01/tiff/20.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
27/Issue_01/tiff/21.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
01/tiff/22.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_01/tiff/
23.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_27/Issue_01/tiff/24.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
27/Issue_01/tiff/25.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
01/tiff/26.tif.

Paine:1960:ACB

[Pai60]  R. M. Paine.  Automatic

coding for business applications. The Computer Journal, 3(3):144–149, October 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/030144.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/144.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/145.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/146.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/147.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/148.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/149.tif.

Paine:1965:GAV

[Pai65]  R. M. Paine.  The grad-

ual acceptance of a variety of commercial English languages. The Computer Journal, 8(3):208–215, October 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/080208.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/208.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/209.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/210.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/211.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/212.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/213.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/214.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/215.tif.

Paine:1966:POC

[Pai66]  R. M. Paine.  Preparation

for optical character recognition. The Computer Journal, 9(3):221–229, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/090221.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/221.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/222.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/223.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/224.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_

                                        03/tiff/225.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        03/tiff/226.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        03/tiff/227.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        03/tiff/228.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        03/tiff/229.tif.

Paine:1969:CRB

[Pai69] R. M. Paine. O.C.R. — benefits and pitfalls. The Computer Journal, 12(2):108–113, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        02/120108.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        02/tiff/108.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        02/tiff/109.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        02/tiff/110.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        02/tiff/111.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        02/tiff/112.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        02/tiff/113.tif. See corre-

spondence [WP69].

Paige:1999:IPD

[Pai99] Richard F. Paige. Integrating a program design calculus and a subset of UML. The Computer Journal, 42(2):82–99, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
02/420082.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
02/pdf/420082.pdf.

Pakes:1967:CID

[Pak67] H. W. Pakes. Correspondence:

On “An investigation into direct numerical methods for solving some calculus of variations problems”. The Computer Journal, 9(4):403, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/090403.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/403.tif. See [All66a].

Palmer:1962:SCI

[Pal62] D. R. Palmer. A stock-control and invoicing system using a Gamma 3 computer. The Computer Journal, 5(1):7–9, April 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
01/050007.sgm.abs.html;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
01/tiff/7.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_05/Issue_01/tiff/
8.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_05/Issue_01/tiff/9.
tif.

Palmer:1963:CCI

[Pal63] R. Palmer. Computer calculations on the initiation of high-explosive detonation. The Computer Journal, 6(1):39–43, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/060039.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/tiff/39.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/
40.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_06/Issue_01/tiff/41.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
06/Issue_01/tiff/42.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/tiff/43.tif.

Palmer:1969:IPO

[Pal69]  J. R. Palmer.  An im-

proved procedure for orthogonalising the search vectors in Rosenbrock’s and Swann’s direct search optimisation methods. The Computer Journal, 12(1):69–71, February 1969.

CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
01/120069.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
01/tiff/69.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_01/tiff/
70.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_12/Issue_01/tiff/71.
tif.

Palme:1972:CHL

[Pal72]  J. Palme.  Correspondence:

On high level languages. The Computer Journal, 15(1):4– 4, 36–36, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/tiff/36.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_01/tiff/

4.tif. See correspondence [Bra72, Hal72].

Palme:1973:CSN

[Pal73]  J. Palme.  Correspondence:

SIMPL/I — a new simulation language from IBM. The Computer Journal, 16(2):189, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/189.tif.

Pallo:1986:ERU

[Pal86] J. M. Pallo. Enumerating, ranking and unranking binary trees. The Computer Journal, 29(2):171–175, April 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        02/tiff/171.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        02/tiff/172.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        02/tiff/173.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        02/tiff/174.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        02/tiff/175.tif.

Pallo:1988:SPR

[Pal88]  J. Pallo.  Some properties

of the rotation lattice of binary trees. The Computer Journal, 31(6):564–565, December 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        06/tiff/564.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        06/tiff/565.tif.

Pankhurst:1970:CPG

[Pan70] R. J. Pankhurst. A computer program for generating diagnostic keys. The Computer Journal, 13(2):145–151, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/130145.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/145.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/146.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/147.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/148.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/149.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/150.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/151.tif.

Papworth:1960:CCR

[Pap60] D. G. Papworth. Computers and change-ringing. The Computer Journal, 3(1):47–50, April 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
01/030047.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
01/tiff/47.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_01/tiff/

                                        48.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_03/Issue_01/tiff/49.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        03/Issue_01/tiff/50.tif.

Pape:1968:CSC

[Pap68] Uwe Pape. Correspondence: On “Some computational notes on the shortest route problems”. The Computer Journal, 11(2):240, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        02/tiff/240.tif. See [Per65].

Papakonstantinou:1982:OEQ

[Pap82] G. Papakonstantinou. Optimal evaluation of queries. The Computer Journal, 25(2):239–241, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/239.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/240.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/241.tif.

Parlett:1967:LEF

[Par67] Beresford Parlett. Letter to the editor: On finding the eigenvalues of real symmetric tridiagonal matrices. The Computer Journal, 9(4):344–345, February 1967. URL http://

comjnl.oxfordjournals.org/
content/9/4/344.full.pdf.

Parker:1974:NAC

[Par74] Frank Parker. Note on Algorithm 78: Counting preferential votes in multi-member constituencies using absolute majority criteria. The Computer Journal, 17(4):380, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/380.tif. See [Van73,

HW74].

Parish:1975:AQT

[Par75a] C. C. M. Parish. On the application of queueing theory to analysing on-line computing systems. The Computer Journal, 18(2):117–121, May 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/117.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/118.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/119.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/120.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/121.tif.

Parkin:1975:BCF

[Par75b] A. Parkin. Bringing cost into file design decisions. The Computer Journal, 18(3):198–199, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/198.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/199.tif.

Parkin:1975:POA

[Par75c] A. Parkin. Programming objectives: an analysis. The Computer Journal, 18(4):307– 311, November 1975. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        18/Issue_04/tiff/307.tif;
                                         http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        04/tiff/308.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        04/tiff/309.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        04/tiff/310.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        04/tiff/311.tif.

Partridge:1975:DDW

[Par75d]  D. Partridge.  A dynamic

database which automatically removes unwanted generalisation for the efficient analysis of language features that exhibit a disparate frequency distribution. The Computer Journal, 18(1):43–48, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
01/tiff/43.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_01/tiff/
44.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_18/Issue_01/tiff/45.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
18/Issue_01/tiff/46.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
01/tiff/47.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_01/tiff/
48.tif.

Parkin:1977:POD

[Par77]  A. Parkin.  The probable

outcomes of a data processing project. The Computer Journal, 20(2):98–101, May 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/100.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/101.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/98.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_02/tiff/
99.tif.

Parkin:1978:CMA

[Par78] A. Parkin. Critique on multicriteria assessment applied to alternative data processing systems. The Computer Journal, 21(2):188–190, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        02/tiff/188.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        02/tiff/189.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        02/tiff/190.tif.

Parkin:1982:DAS

[Par82]  A. Parkin.  Data analysis

and system design by entityrelationship modelling — a practical example. The Computer Journal, 25(4):401–409, November 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        25/Issue_04/tiff/401.tif;
                                         http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        04/tiff/402.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        04/tiff/403.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        04/tiff/404.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        04/tiff/405.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/406.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/407.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/408.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/409.tif.

Parr:1988:EAR

[Par88] G. P. Parr. Enhanced address resolution in a multi-LAN Ethernet communications system. The Computer Journal, 31(4):296–303, August 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/296.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/297.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/298.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/299.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/300.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/301.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/302.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_31/Issue_
                                        04/tiff/303.tif.

Parker:1989:CFP

[Par89] J. R. Parker. On the conversion of filled polygonal regions from raster to vector representation. The Computer Journal, 32(6):549–553, December 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        06/tiff/549.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        06/tiff/550.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        06/tiff/551.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        06/tiff/552.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        06/tiff/553.tif.

Parhami:1992:SNR

[Par92] B. Parhami. Systolic number radix converters. The Computer Journal, 35(4):405–409, August 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_35/Issue_04/
                                        Vol35_04.body.html#AbstractParhami.

Parhami:1996:EVS

[Par96a] Behrooz Parhami. Extremevalue search and general selection algorithms for fully parallel associative memories. The Computer Journal, 39(3):241– 250, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www.oup.co.uk/jnls/
list/comjnl/hdb/Volume_39/
Issue_03/390241.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_03/
Vol39_03.body.html#AbstractParhami.

Parhami:1996:PTV

[Par96b]  Behrooz Parhami.  Parallel

threshold voting. The Computer Journal, 39(8):692–??, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www.oup.co.uk/jnls/
list/comjnl/hdb/Volume_39/
Issue_08/390692.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_08/
Vol39_08.body.html#AbstractParhami.

Paulson:1980:BRI

[Pau80] R. W. Paulson. Book review: An Introduction to Computational Fluid Mechanics, by Chuen-Yen Chow, 1979. The Computer Journal, 23(2):123, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/123.tif.

Paulson:1985:LLL

[Pau85] Lawrence C. Paulson. Lessons learned from LCF: a survey of natural deduction proofs. The Computer Journal, 28 (5):474–479, November 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/474.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/475.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/476.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/477.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/478.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/479.tif.

Penny:1986:DRM

[PAW86] J. P. Penny, P. J. Ashton, and A. L. Wilkinson. Data recording and monitoring for analysis of system response times. The Computer Journal, 29(5):396–403, October 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/396.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/397.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/398.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_

05/tiff/399.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/400.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/401.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/402.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/403.tif.

Pinkerton:1969:CP

[PB69] Tad Pinkerton and D. W. Barron. Correspondence: Paging. The Computer Journal, 12(1):105, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
01/tiff/105.tif. See [LWS68,

LWS69b, GB69].

Presser:1974:ECD

[PB74] L. Presser and J. Benson. Evaluation of compiler diagnostics. The Computer Journal, 17(2):121–123, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/121.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/122.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/123.tif.

Patel:1976:ASB

[PB76] M. R. Patel and K. H. Bennett. Analysis of speed of a binary multiplier using a variable number of shifts per cycle. The Computer Journal, 19(3):254–257, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/254.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/255.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/256.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/257.tif.

Patel:1978:ASB

[PB78] M. R. Patel and K. H. Bennett. Analysis of speed of a binary divider using a variable number of shifts per cycle. The Computer Journal, 21(3):246–252, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        03/tiff/246.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        03/tiff/247.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        03/tiff/248.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_

03/tiff/249.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/250.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/251.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/252.tif.

Patnaik:1986:TTI

[PB86] L. M. Patnaik and Julie Basu.

Two tools for interprocess communication in distributed dataflow systems. The Computer Journal, 29(6):506–521, December 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/506.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/507.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/508.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/509.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/510.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/511.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/512.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/513.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        06/tiff/514.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        06/tiff/515.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        06/tiff/516.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        06/tiff/517.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        06/tiff/518.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        06/tiff/519.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        06/tiff/520.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        06/tiff/521.tif.

Padget:1991:COP

[PBF91] J. Padget, R. Bradford, and J. Fitch. Concurrent objectoriented programming in Lisp. The Computer Journal, 34 (4):311–319, August 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        04/tiff/311.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        04/tiff/312.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        04/tiff/313.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_34/Issue_
04/tiff/314.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/315.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/316.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/317.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/318.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/319.tif.

Pitman:1986:BSS

[PBH86] N. M. Pitman, F. W. Burton, and E. W. Haddon. Buddy systems with selective splitting. The Computer Journal, 29(2):127–134, April 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/127.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/128.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/129.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/130.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/131.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_

                                        02/tiff/132.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        02/tiff/133.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        02/tiff/134.tif.

Plata:1990:ASP

[PBR^{+}90] O. G. Plata, J. D. Bruguera, F. F. Rivera, R. Doallo, and E. L. Zapata. ACLE: a software package for SIMD computer simulation. The Computer Journal, 33(3):194–203, June 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        03/tiff/194.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        03/tiff/195.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        03/tiff/196.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        03/tiff/197.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        03/tiff/198.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        03/tiff/199.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        03/tiff/200.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        03/tiff/201.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_

03/tiff/202.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/203.tif.

Parker:1964:PDE

[PC64] I. B. Parker and J. Crank.

Persistent discretization errors in partial differential equations of parabolic type. The Computer Journal, 7(2):163–167, July 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/070163.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/163.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/164.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/165.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/166.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/167.tif. See com-

ment [KM66] and correspondence [dD66].

Parkin:1978:EBD

[PC78] A. Parkin and R. B. Coats.

EDSIM-event based discrete simulation using general purpose languages such as FOR-TRAN. The Computer Journal, 21(2):122–127, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/122.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/123.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/124.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/125.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/126.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/127.tif.

Perry:1963:BRL

[PCSG63] G. H. Perry, J. C. Cluley, K. L. Smith, and R. L. Goodstein. Book reviews: Large-Capacity Memory Techniques for Computing Systems, edited by Marshall C. Yovits, 1962; Digital Techniques, by D. W. Davies, 1963; Theory and Design of Digital Machines, by Thomas C. Bartee, Irwin Lebow, [and] Irving S. Reed, 1962; A Survey of Mathematical Logic, by Hao Wang, 1963. The Computer Journal, 6(3):295–298, October 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/295.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/296.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/297.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/298.tif.

Price:1979:AWS

[PD79] W. L. Price and M. Dowson.

Algorithm 107: A weighted simplex procedure for the solution of simultaneous nonlinear equations. The Computer Journal, 22(3):282, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/282.tif. This algo-

rithm is published under the Journal’s agreement with the Institute of Mathematics and its Applications, and is associated with the authors’ article in the JIMA, Volume 24, Number 1, August 1979.

Pan:1996:FBR

[PD96a] Si Pan and R. Geoff Dromey.

A formal basis for removing goto statements. The Computer Journal, 39(3):203–214, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www.oup.co.uk/jnls/
list/comjnl/hdb/Volume_39/
Issue_03/390203.sgm.abs.html;
 http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_03/
Vol39_03.body.html#AbstractPan2.

Pan:1996:REL

[PD96b] Si Pan and R. Geoff Dromey. Re-engineering loops. The Computer Journal, 39(3):184–202, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://www.oup.co.uk/jnls/
                                        list/comjnl/hdb/Volume_39/
                                        Issue_03/390184.sgm.abs.html;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_39/Issue_03/
                                        Vol39_03.body.html#AbstractPan1.

Parsons:1974:DML

[PDY74] R. G. Parsons, A. G. Dale, and C. V. Yurkanan. Data manipulation language requirements for data base management systems. The Computer Journal, 17(2):99–103, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        02/tiff/100.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        02/tiff/101.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        02/tiff/102.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        02/tiff/103.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        02/tiff/99.tif.

Pearson:1963:SC

[Pea63] K. W. Pearson. Satellite communications. The Computer Journal, 5(4):308–312, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/050308.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/308.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/309.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/310.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/311.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/312.tif.

Pearson:1969:VMM

[Pea69]  J. D. Pearson.  Variable

metric methods of minimisation. The Computer Journal, 12(2):171–178, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/120171.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/171.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/172.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/173.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_12/Issue_
                                        02/tiff/174.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        02/tiff/175.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        02/tiff/176.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        02/tiff/177.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        02/tiff/178.tif.

Peckham:1970:NMM

[Pec70] G. Peckham. A new method for minimising a sum of squares without calculating gradients. The Computer Journal, 13 (4):418–420, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/130418.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/418.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/419.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/420.tif.

Penny:1966:ABT

[Pen66] J. P. Penny. An analysis, both theoretical and by simulation, of a time-shared computer system. The Computer Journal, 9(1):53–59, May 1966. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/090053.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/53.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_01/tiff/
54.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_09/Issue_01/tiff/55.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
09/Issue_01/tiff/56.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/57.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_01/tiff/
58.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_09/Issue_01/tiff/59.

tif. See correspondence [Bae66].

Pengilly:1976:ASD

[Pen76] P. J. Pengilly. An approach to systems design. The Computer Journal, 19(1):8–12, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
01/tiff/10.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_01/tiff/
11.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_19/Issue_01/tiff/12.
tif;   http://www3.oup.co.

                                        uk/computer_journal/hdb/Volume_
                                        19/Issue_01/tiff/8.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        01/tiff/9.tif.

Pengilly:1977:ASD

[Pen77]  P. J. Pengilly.  An ap-

proach to systems design. II. the popularisation of PRO-PLAN. The Computer Journal, 20(4):296–297, November 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        04/tiff/296.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        04/tiff/297.tif.

Perko:1965:SCN

[Per65] Aarni Perko. Some computational notes on the shortest route problems. The Computer Journal, 8(1):19–20, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        01/080019.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        01/tiff/19.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_08/Issue_01/tiff/

20.tif. See correction [Pap68].

Perlis:1976:ACS

[Per76] Donald Perlis. An application of compiler simulation at the source language level. The Computer Journal, 19(1):90, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
01/tiff/90.tif.

Perros:1980:RMP

[Per80] H. G. Perros. A regression model for predicting the response time of a disc I/O system. The Computer Journal, 23(1):34–36, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
01/tiff/34.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/
35.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_23/Issue_01/tiff/36.
tif.

PeytonJones:1989:PIF

[Pey89] Simon L. Peyton Jones. Parallel implementations of functional programming languages. The Computer Journal, 32(2): 175–186, April 1989. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
32/Issue_02/tiff/175.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/176.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_

02/tiff/177.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/178.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/179.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/180.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/181.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/182.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/183.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/184.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/185.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/186.tif.

Pezarro:1976:NEH

[Pez76]  M. T. Pezarro.  A note

on estimating hit ratios for direct-access storage devices. The Computer Journal, 19 (3):271–272, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/271.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/272.tif.

Pramanik:1986:IDM

[PF86] S. Pramanik and F. Fotouhi. Index database machine. The Computer Journal, 29(5):451–457, October 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/451.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/452.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/453.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/454.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/455.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/456.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/457.tif.

Pflug:1984:DMA

[Pfl84] G. Ch. Pflug. Dynamic memory allocation — a Markovian analysis. The Computer Journal, 27(4):328–333, November 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/328.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_

                                        04/tiff/329.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        04/tiff/330.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        04/tiff/331.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        04/tiff/332.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        04/tiff/333.tif.

Piper:1974:ARP

[PFW74] L. R. Piper, I. R. Franklin, and N. H. Westwood. Algorithm 83: A rejection procedure for sampling from the distribution of chi-squared (χ^{2}) or the sampling variance (s^{2}). The Computer Journal, 17(4):378, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        04/tiff/378.tif.

Prather:1981:DFC

[PG81] Ronald E. Prather and Shirla G. Giulieri. Decomposition of flow chart schemata. The Computer Journal, 24(3):258–262, August 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        03/tiff/258.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        03/tiff/259.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/260.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/261.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/262.tif.

Pitteway:1982:BAR

[PG82] M. L. V. Pitteway and A. J. R.

Green. Bresenham’s algorithm with run line coding shortcut. The Computer Journal, 25(1):114–115, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/114.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/115.tif.

Paton:1990:OED

[PG90] N. W. Paton and P. M. D.

Gray. Optimising and executing DAPLEX queries using Prolog. The Computer Journal, 33(6):547–555, December 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/547.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/548.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/549.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/550.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/551.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/552.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/553.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/554.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/555.tif.

Papazoglou:1984:ACP

[PGM84] M. P. Papazoglou, P. I. Georgiadis, and D. G. Maritsas. Architectural considerations of the parallel SIM-ULA machine. The Computer Journal, 27(3):254–259, August 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/254.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/255.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/256.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/257.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/258.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/259.tif.

Pawson:1973:CHT

[PH73] A. J. D. Pawson and F. R. A.

Hopgood. Correspondence: Hashing techniques for table searching. The Computer Journal, 16(3):285, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/285.tif. See [HD72].

Pham:1990:ACC

[Pha90]  B. Pham.  Algorithms for

calculating cubic Beta-splines. The Computer Journal, 33 (5):412–416, October 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/412.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/413.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/414.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/415.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/416.tif.

Phillips:1967:NRT

[Phi67a] C. S. E. Phillips. Networks for real-time programming. The Computer Journal, 10(1):46–52, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/100046.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/46.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_01/tiff/
47.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_10/Issue_01/tiff/48.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/49.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/50.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_01/tiff/
51.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_10/Issue_01/tiff/52.
tif.

Phillips:1967:NIT

[Phi67b] G. M. Phillips. Numerical integration in two and three dimensions. The Computer Journal, 10(2):202–204, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/100202.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/202.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_

02/tiff/203.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/204.tif.

Phillips:1967:NID

[Phi67c]  G. M. Phillips.  Numeri-

cal integration over an Ndimensional rectangular region. The Computer Journal, 10(3):297–299, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/100297.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/297.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/298.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/299.tif.

Phillips:1968:APS

[Phi68a] G. M. Phillips. Algorithms for piecewise straight line approximations. The Computer Journal, 11(2):211–212, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/110211.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/211.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_

02/tiff/212.tif. See corre-

spondence [BH69].

Phillips:1968:EME

[Phi68b] G. M. Phillips. Estimate of the maximum error in best polynomial approximations. The Computer Journal, 11(1):110–111, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/110110.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/110.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/111.tif.

Phillips:1970:ANI

[Phi70]  G. M. Phillips.  Analysis

of numerical iterative methods for solving integral and integrodifferential equations. The Computer Journal, 13 (3):297–300, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/130297.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/297.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/298.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/299.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_13/Issue_
03/tiff/300.tif.

Phillips:1974:PIF

[Phi74] D. A. Phillips. A preliminary investigation of function optimisation by a combination of methods. The Computer Journal, 17(1):75–79, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
01/tiff/75.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/
76.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_17/Issue_01/tiff/77.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
17/Issue_01/tiff/78.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
01/tiff/79.tif.

Pike:1967:NAF

[PHJ67] M. C. Pike, I. D. Hill, and F. D.

James. Note on Algorithm 2: Fibonacci search, and on Algorithm 7: MINX. The Computer Journal, 9(4):416–417, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/416.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/417.tif. See [PP79].

Pan:1996:SDM

[PHS96] Yi Pan, Mounir Hamdi, and Gurdip Singh. Selection on kdimensional meshes with multiple broadcasting. The Computer Journal, 39(2):140–144, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://www.oup.co.uk/jnls/
                                        list/comjnl/hdb/Volume_39/
                                        Issue_02/390140.sgm.abs.html;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_39/Issue_02/
                                        Vol39_02.body.html#AbstractHamdi.

Piessens:1971:SAG

[Pie71] R. Piessens. Some aspects of Gaussian quadrature formulae for the numerical inversion of the Laplace transform. The Computer Journal, 14(4):433–436, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/140433.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/433.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/434.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/435.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/436.tif.

Pierce:1974:SLD

[Pie74] R. H. Pierce. Source language debugging on a small computer. The Computer Journal, 17(4):313–317, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/313.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/314.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/315.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/316.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/317.tif.

Piessens:1982:AIL

[Pie82] Robert Piessens. Algorithm 113: Inversion of the Laplace transform. The Computer Journal, 25(2):278–282, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/278.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/279.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/280.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_

                                        02/tiff/281.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/282.tif.

Pike:1979:AR

[Pik79] M. C. Pike. Algorithm 5: RANCOMB. The Computer Journal, 22(1):88, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        01/tiff/88.tif. Reprinted

from The Computer Bulletin, September 1964, p. 67.

Pike:1981:TCU

[Pik81] J. Pike. Text compression using a 4-bit coding scheme. The Computer Journal, 24 (4):324–330, November 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        04/tiff/324.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        04/tiff/325.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        04/tiff/326.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        04/tiff/327.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        04/tiff/328.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        04/tiff/329.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/330.tif.

Pinkerton:1961:EDS

[Pin61] J. M. M. Pinkerton. The evolution of design in a series of computers, LEO I–III. The Computer Journal, 4(1):42–46, April 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/040042.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/tiff/42.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_01/tiff/
43.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_04/Issue_01/tiff/44.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
04/Issue_01/tiff/45.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/tiff/46.tif.

Piney:1977:GIS

[Pin77] C. Piney. A generalised interactive system for the acquisition of structured data. The Computer Journal, 20(1):37–44, February 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
01/tiff/37.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_01/tiff/
38.tif;   http://www3.oup.

co.uk/computer_journal/hdb/
Volume_20/Issue_01/tiff/39.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
20/Issue_01/tiff/40.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
01/tiff/41.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_01/tiff/
42.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_20/Issue_01/tiff/43.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
20/Issue_01/tiff/44.tif.

Pitteway:1965:ODL

[Pit65] M. L. V. Pitteway. The organization of a data link computing service. The Computer Journal, 8(3):190–200, October 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/080190.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/190.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/191.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/192.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/193.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/194.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_08/Issue_
03/tiff/195.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/196.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/197.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/198.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/199.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/200.tif.

Pitteway:1967:ADE

[Pit67] M. L. V. Pitteway. Algorithm for drawing ellipses or hyperbolae with a digital plotter. The Computer Journal, 10(3):282–289, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/100282.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/282.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/283.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/284.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/285.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_

                                        03/tiff/286.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        03/tiff/287.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        03/tiff/288.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        03/tiff/289.tif.

Pitcher:1974:CDR

[Pit74] H. H. W. Pitcher. Correspondence: Definition of random. The Computer Journal, 17(4):381, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        04/tiff/381.tif. See [SL73].

Pitman:1978:NAM

[Pit78] Anthony E. N. T. Pitman. Note on Algorithm 1: Mintree. The Computer Journal, 21(3): 276, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        21/Issue_03/tiff/276.tif.

See [Obr78a].

Prowse:1980:NLD

[PJ80] P. H. Prowse and R. G. Johnson. A natural language data base interface to the user. The Computer Journal, 23(1):22–25, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
01/tiff/22.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/
23.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_23/Issue_01/tiff/24.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
23/Issue_01/tiff/25.tif.

Pandya:1986:SDT

[PJ86] P. Pandya and M. Joseph. A structure-directed total correctness proof rule for recursive procedure calls. The Computer Journal, 29(6):531–537, December 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/531.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/532.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/533.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/534.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/535.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/536.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/537.tif.

Page:1989:SME

[PJ89] I. P. Page and R. T. Jacob. The solution of mutual exclusion problems which can be described graphically. The Computer Journal, 32(1):45–54, February 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        01/tiff/45.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_32/Issue_01/tiff/
                                        46.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_32/Issue_01/tiff/47.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        32/Issue_01/tiff/48.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        01/tiff/49.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_32/Issue_01/tiff/
                                        50.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_32/Issue_01/tiff/51.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        32/Issue_01/tiff/52.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        01/tiff/53.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_32/Issue_01/tiff/
                                        54.tif.

Pramanik:1985:HPM

[PK85] Sakti Pramanik and Chung-Ta King. A hardware pattern matching algorithm on a dataflow. The Computer Journal, 28(3):264–269, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/264.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/265.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/266.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/267.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/268.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/269.tif.

Papakonstantinou:1986:KRA

[PK86] George K. Papakonstantinou and John Kontos. Knowledge representation with attribute grammars. The Computer Journal, 29(3):241–245, June 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/241.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/242.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/243.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_

                                        03/tiff/244.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        03/tiff/245.tif.

Parker:1987:T

[PKK87] J. Parker, A. Kennard, and D. King. The ‘window’ terminal. The Computer Journal, 30(6):558–564, December 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        06/tiff/558.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        06/tiff/559.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        06/tiff/560.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        06/tiff/561.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        06/tiff/562.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        06/tiff/563.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        06/tiff/564.tif.

Poole:1968:DLC

[PL68] P. C. Poole and T. Lang. The development of on-line computing facilities for the KDF9 part 1; COSEC — a single on-line console. The Computer Journal, 11(1):5–11, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/110005.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/10.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_01/tiff/
11.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_11/Issue_01/tiff/5.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/6.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/7.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_11/Issue_01/tiff/
8.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_11/Issue_01/tiff/9.
tif.

Pai:1974:ASL

[PL74] D. Pai and D. Lewin. Analysis of sequential logic circuits. The Computer Journal, 17(1):64–68, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
01/tiff/64.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/
65.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_17/Issue_01/tiff/66.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
17/Issue_01/tiff/67.tif;

                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        01/tiff/68.tif.

Poix:1994:GRT

[PL94a] Christophe Poix and Robert Laurini. A geometric reasoning tool based on a successive improvement approach. The Computer Journal, 37(5):377–384, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_37/Issue_05/
                                        Vol37_05.body.html#AbstractPoix.

Poo:1994:OOS

[PL94b] Danny C. C. Poo and Shwu-Yi Lee. An object-oriented systems modelling method based on the Jackson approach. The Computer Journal, 37(8):669–682, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_37/Issue_08/
                                        Vol37_08.body.html#AbstractPoo.

Platt:1961:EAC

[Pla61]  A. J. Platt.  The experi-

ence of applying a commercial computer in a British organization. The Computer Journal, 3(4):185–197, January 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        04/030185.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_

journal/hdb/Volume_03/Issue_
04/tiff/185.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/186.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/187.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/188.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/189.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/190.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/191.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/192.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/193.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/194.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/195.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/196.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/197.tif.

Playfair:1964:CP

[Pla64] E. Playfair. Computers and psychology. The Computer Journal, 7(1):1–3, April 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        01/070001.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        01/tiff/1.tif;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_07/Issue_01/tiff/
                                        2.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_07/Issue_01/tiff/3.
                                        tif.

Proctor:1962:PSR

[PM62] W. G. Proctor and M. F. Mitchell. The PACE scaling routine for Mercury. The Computer Journal, 5(1):24–27, April 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        01/050024.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        01/tiff/24.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_05/Issue_01/tiff/
                                        25.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_05/Issue_01/tiff/26.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        05/Issue_01/tiff/27.tif.

Powell:1972:RCI

[PM72] D. R. Powell and J. R. Macdonald. A rapidly convergent iterative method for the solution of the generalised nonlinear least squares problem. The Computer Journal, 15(2):148–155, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/150148.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/148.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/149.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/150.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/151.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/152.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/153.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/154.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/155.tif.

Pan:1990:SCB

[PM90] Y. Pan and R. Melhem. Short circuits in buffered multistage interconnection networks. The Computer Journal, 33 (4):323–329, August 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
04/tiff/323.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        04/tiff/324.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        04/tiff/325.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        04/tiff/326.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        04/tiff/327.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        04/tiff/328.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        04/tiff/329.tif.

Polley:1960:PRI

[Pol60] D. W. Polley. A progress report on the introduction of A.D.P. for recording contributions paid under the new graduated pensions scheme. The Computer Journal, 3(3):117–120, October 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        03/030117.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        03/tiff/117.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        03/tiff/118.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        03/tiff/119.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        03/tiff/120.tif.

Pong:1991:IIG

[Pon91] M.-C. Pong. I-pigs: an interactive graphical environment for concurrent programming. The Computer Journal, 34(4):320–330, August 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/320.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/321.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/322.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/323.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/324.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/325.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/326.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/327.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/328.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/329.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/330.tif.

Poo:1991:RBP

[Poo91]  C.-C. D. Poo.  Represent-

ing business policies in the Jackson system development method. The Computer Journal, 34(2):122–131, April 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        02/tiff/122.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        02/tiff/123.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        02/tiff/124.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        02/tiff/125.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        02/tiff/126.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        02/tiff/127.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        02/tiff/128.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        02/tiff/129.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        02/tiff/130.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        02/tiff/131.tif.

Poulovassilis:1992:IFF

[Pou92] A. Poulovassilis. The implementation of FDL, a functional database language. The Computer Journal, 35(2):119–128, April 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_35/Issue_02/
Vol35_02.body.html#AbstractPoulovassilis.

Poulovassilis:1993:PMA

[Pou93] A. Poulovassilis. A patternmatching algorithm for functional databases. The Computer Journal, 36(2):195–199, April 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_02/
Vol36_02.body.html#AbstractPoulovassilis.

Powell:1962:IMF

[Pow62] M. J. D. Powell. An iterative method for finding stationary values of a function of several variables. The Computer Journal, 5(2):147–151, July 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/050147.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/147.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/148.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/149.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_05/Issue_
02/tiff/150.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/151.tif.

Powell:1964:EMF

[Pow64] M. J. D. Powell. An efficient method for finding the minimum of a function of several variables without calculating derivatives. The Computer Journal, 7(2):155–162, July 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/070155.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/155.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/156.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/157.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/158.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/159.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/160.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/161.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/162.tif.

Powell:1965:MMS

[Pow65] M. J. D. Powell. A method for minimizing a sum of squares of non-linear functions without calculating derivatives. The Computer Journal, 7(4):303–307, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/070303.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/303.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/304.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/305.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/306.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/307.tif.

Powell:1967:MEP

[Pow67] M. J. D. Powell. On the maximum errors of polynomial approximations defined by interpolation and by least squares criteria. The Computer Journal, 9(4):404–407, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/090404.sgm.abs.html;
http://www3.oup.co.uk/computer_

                                        journal/hdb/Volume_09/Issue_
                                        04/tiff/404.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        04/tiff/405.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        04/tiff/406.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        04/tiff/407.tif.

Powell:1968:COV

[Pow68] M. J. D. Powell. On the calculation of orthogonal vectors. The Computer Journal, 11(3):302–304, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        03/110302.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        03/tiff/302.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        03/tiff/303.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        03/tiff/304.tif. See corre-

spondence [And69].

Powell:1969:TRO

[Pow69] M. J. D. Powell. A theorem on rank one modifications to a matrix and its inverse. The Computer Journal, 12(3):288–290, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_

journal/hdb/Volume_12/Issue_
03/120288.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/288.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/289.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/290.tif.

Pitteway:1969:NAB

[PP69] M. L. V. Pitteway and R. W.

Parry. Note on Algorithm 35: Best rational approximation to a real number. The Computer Journal, 12(3):293, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/293.tif. See [Pre68,

Cob69].

Pike:1979:AFS

[PP79] M. C. Pike and J. Pixner. Algorithm 2: Fibonacci search. The Computer Journal, 22(1): 86, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
22/Issue_01/tiff/86.tif. Reprinted

from The Computer Bulletin, vol. 8, p. 147, 1965. See [Boo79c, Ove67c, PHJ67].

Panholzer:1998:TMP

[PP98] Alois Panholzer and Helmut Prodinger. Towards a more precise analysis of an algorithm to generate binary trees: A tutorial. The Computer Journal, 41(3):201–204, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_41/Issue_03/
Vol41_03.body.html#AbstractPanholzer.

Prakash:1983:DDF

[PPB83] Naveen Prakash, Parimala N., and N. Bolloju. Data definition facilities in Admin. The Computer Journal, 26 (4):329–335, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/329.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/330.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/331.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/332.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/333.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/334.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/335.tif.

Parimala:1984:NCN

[PPB84a] N. Parimala, Naveen Prakash, and N. Bolloju. New corecs and new cosets in Admin. The Computer Journal, 27 (4):310–314, November 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/310.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/311.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/312.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/313.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/314.tif.

ParimalaN:1984:CCA

[PPB84b] Parimala N., Naveen Prakash, and N. Bolloju. Concurrency control in Admin. The Computer Journal, 27(1):62– 66, February 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
27/Issue_01/tiff/62.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
01/tiff/63.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_01/tiff/
64.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_27/Issue_01/tiff/65.
tif;   http://www3.oup.co.

                                        uk/computer_journal/hdb/Volume_
                                        27/Issue_01/tiff/66.tif.

Prakash:1984:SIC

[PPB84c] Naveen Prakash, N. Parimala, and N. Bolloju. Specifying integrity constraints in a network DBMS. The Computer Journal, 27(3):209–217, August 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        03/tiff/209.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        03/tiff/210.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        03/tiff/211.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        03/tiff/212.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        03/tiff/213.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        03/tiff/214.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        03/tiff/215.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        03/tiff/216.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        03/tiff/217.tif.

Parimala:1989:QFN

[PPRB89] N. Parimala, N. Prakash, B. L. N. Rao, and N. Bolloju. A query facility to a network DBMS. The Computer Journal, 32(1):55–62, February 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
01/tiff/55.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_01/tiff/
56.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_32/Issue_01/tiff/57.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
32/Issue_01/tiff/58.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
01/tiff/59.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_01/tiff/
60.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_32/Issue_01/tiff/61.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
32/Issue_01/tiff/62.tif.

Parker-Rhodes:1964:CA

[PR64] A. F. Parker-Rhodes. The communication of algorithms. The Computer Journal, 7(1):28–35, April 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
01/070028.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
01/tiff/28.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_01/tiff/
29.tif;   http://www3.oup.

                                        co.uk/computer_journal/hdb/
                                        Volume_07/Issue_01/tiff/30.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        07/Issue_01/tiff/31.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        01/tiff/32.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_07/Issue_01/tiff/
                                        33.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_07/Issue_01/tiff/34.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        07/Issue_01/tiff/35.tif.

Pierce:1977:TCS

[PR77] R. H. Pierce and J. Rowell. A transformation-directed compiling system. The Computer Journal, 20(2):109–115, May 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        02/tiff/109.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        02/tiff/110.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        02/tiff/111.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        02/tiff/112.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        02/tiff/113.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        02/tiff/114.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_20/Issue_
02/tiff/115.tif.

Prasse:1998:WCT

[PR98] Michael Prasse and Peter Rittgen. Why Church’s Thesis still holds. some notes on Peter Wegner’s tracts on interaction and computability. The Computer Journal, 41(6):357–362, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_41/Issue_06/
Vol41_06.body.html#AbstractPrasse.

Prather:1984:ATS

[Pra84] Ronald E. Prather. An axiomatic theory of software complexity measure. The Computer Journal, 27(4):340–347, November 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/340.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/341.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/342.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/343.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/344.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/345.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_27/Issue_
                                        04/tiff/346.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        04/tiff/347.tif.

Prather:1988:CET

[Pra88] R. E. Prather. Comparison and extension of theories of Zipf and Halstead. The Computer Journal, 31(3):248–252, June 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        03/tiff/248.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        03/tiff/249.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        03/tiff/250.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        03/tiff/251.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        03/tiff/252.tif.

Prescott:1968:ABR

[Pre68] R. J. Prescott. Algorithm 35: Best rational approximation to a real number. The Computer Journal, 11(3):347–348, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        03/tiff/347.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_

03/tiff/348.tif. See note

[PP69, Cob69].

Price:1964:DCF

[Pri64] V. E. Price. Desirable computing facilities for a College of Advanced Technology. The Computer Journal, 7(3):177–179, October 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/070177.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/177.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/178.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/179.tif.

Price:1968:OVI

[Pri68] C. F. Price. An offset vector iteration method for solving two-point boundary-value problems. The Computer Journal, 11(2):220–228, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/110220.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/220.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/221.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_11/Issue_
                                        02/tiff/222.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        02/tiff/223.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        02/tiff/224.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        02/tiff/225.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        02/tiff/226.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        02/tiff/227.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        02/tiff/228.tif.

Price:1977:CRS

[Pri77] W. L. Price. A controlled random search procedure for global optimisation. The Computer Journal, 20(4):367–370, November 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        04/tiff/367.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        04/tiff/368.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        04/tiff/369.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        04/tiff/370.tif.

Price:1978:SSD

[Pri78] W. L. Price. Simulation studies of data communication networks operating in datagram mode. The Computer Journal, 21(3):219–223, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/219.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/220.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/221.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/222.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/223.tif.

Pringle:1981:JFH

[Pri81]  Alison M. Pringle.  Justi-

fication with fewer hyphens. The Computer Journal, 24 (4):320–323, November 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/320.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/321.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/322.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/323.tif.

Priami:1995:SC

[Pri95]  Corrado Priami.  Stochas-

tic π-calculus. The Computer Journal, 38(7):578–589, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_38/Issue_07/
                                        Vol38_07.body.html#AbstractPriami.

Pliskin:1993:PVA

[PRLW93] N. Pliskin, T. Romm, A. S. Lee, and Y. Weber. Presumed versus actual organizational culture: managerial implications for implementation of information systems. The Computer Journal, 36(2):143–152, April 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_36/Issue_02/
                                        Vol36_02.body.html#AbstractPliskin.

Proskurowski:1974:OMA

[Pro74]  W. Proskurowski.  Order-

ing method to accelerate the solution of mate-in-two chess problems by computer. The Computer Journal, 17 (1):80–81, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        01/tiff/80.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_17/Issue_01/tiff/
                                        81.tif.

Perrott:1980:SEU

[PRO80a] R. H. Perrott, A. K. Raja, and P. C. O’Kane. A simulation experiment using two languages. The Computer Journal, 23(2):142–146, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/142.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/143.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/144.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/145.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/146.tif.

Prowse:1980:DBA

[Pro80b] P. Prowse. The data base approach. The Computer Journal, 23(1):9–12, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
01/tiff/10.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/
11.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_23/Issue_01/tiff/12.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
23/Issue_01/tiff/9.tif.

Prosser:1983:NCS

[Pro83] C. J. Prosser. A note on computing the square root of an integer. The Computer Journal, 26(2):187–188, May 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        02/tiff/187.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        02/tiff/188.tif.

Puliafito:1996:EPP

[PRS96] A. Puliafito, S. Riccobene, and M. Scarpa. Evaluation of performability parameters in client-server environments. The Computer Journal, 39(8):647– ??, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://www.oup.co.uk/jnls/
                                        list/comjnl/hdb/Volume_39/
                                        Issue_08/390647.sgm.abs.html;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_39/Issue_08/
                                        Vol39_08.body.html#AbstractPuliafito.

Poole:1974:CSO

[PS74] F. Poole and P. M. Stocker. Correspondence: On “Selforganising data management systems”. The Computer Journal, 17(1):95, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        01/tiff/95.tif. See [SD73].

Price:1990:SPT

[PS90] Camille C. Price and M. A.

Salama. Scheduling of precedenceconstrained tasks on multiprocessors. The Computer Journal, 33(3):219–229, June 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/219.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/220.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/221.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/222.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/223.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/224.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/225.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/226.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/227.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/228.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/229.tif.

Paschos:1992:EEC

[PS92] V. T. H. Paschos and A. N. Stafylopatis. Evaluation of the execution cost of recursive definitions. The Computer Journal, 35(5):A429–??, October 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_35/Issue_05/
                                        35000A429.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/Volume_35/Issue_05/
                                        tiff/A429.tif.

Pemberton:1993:EMI

[PS93] Jonathan D. Pemberton and Joanne M. Smailes. Educating the masses — is IT a tricky business? The Computer Journal, 36(4):335–342, August 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_36/Issue_04/
                                        Vol36_04.body.html#AbstractPemberton.

Peng:1996:APD

[PS96] Shietung Peng and Stanislav Sedukhin. Array processors design for division-free linear system solving. The Computer Journal, 39(8):713–??, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://www.oup.co.uk/jnls/
                                        list/comjnl/hdb/Volume_39/
                                        Issue_08/390713.sgm.abs.html;
                                         http://www3.oup.co.uk/computer_

journal/Volume_39/Issue_08/
Vol39_08.body.html#AbstractPeng.

Prakash:1981:SDB

[PSD^{+}81] Naveen Prakash, P. Sadanandan, S. M. Desai, K. Rama Rao, and P. N. Gejji. Samhita — a data base management system. The Computer Journal, 24(1):48–51, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
01/tiff/48.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_01/tiff/
49.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_24/Issue_01/tiff/50.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
24/Issue_01/tiff/51.tif.

Paakki:1996:AHS

[PSK96] Jukka Paakki, Airi Salminen, and Jussi Koskinen. Automated hypertext support for software maintenance. The Computer Journal, 39(7):577– 597, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www.oup.co.uk/jnls/
list/comjnl/hdb/Volume_39/
Issue_07/390577.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_07/
Vol39_07.body.html#AbstractPaakki.

Puliti:1990:ITD

[PT90] P. Puliti and G. Tascini.

Interpreting technical drawings. The Computer Journal, 33(5):417–423, October 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/417.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/418.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/419.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/420.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/421.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/422.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/423.tif.

Pan:1997:SEA

[PTV97] Yi Pan, Jerry L. Trathan, and Ramachandran Vaidyanathan. A scalable and efficient algorithm for computing the city block distance transform on reconfigurable meshes. The Computer Journal, 40(7):435–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_40/Issue_07/
Vol40_07.body.html#AbstractPan.

Pullin:1964:FAT

[Pul64]  D. Pullin.  A FORTRAN

to ALGOL translator. The Computer Journal, 7(1):24–27, April 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
01/070024.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
01/tiff/24.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_01/tiff/
25.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_07/Issue_01/tiff/26.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
07/Issue_01/tiff/27.tif.

Pullen:1967:CAT

[PW67] H. G. M. Pullen and M. H. J.

Webb. A computer application to a transport scheduling problem. The Computer Journal, 10(1):10–13, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/100010.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/10.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_01/tiff/
11.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_10/Issue_01/tiff/12.
tif;   http://www3.oup.co.

                                        uk/computer_journal/hdb/Volume_
                                        10/Issue_01/tiff/13.tif.

Peters:1969:EBS

[PW69] G. Peters and J. H. Wilkinson. Eigenvalues of Ax = λBx with band symmetric A and B. The Computer Journal, 12(4):398–404, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/120397.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/398.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/399.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/400.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/401.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/402.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/403.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/404.tif.

Peters:1970:LSP

[PW70] G. Peters and J. H. Wilkinson. The least squares problem and pseudo-inverses. The Computer Journal, 13(3):309–316, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/130309.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/309.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/310.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/311.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/312.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/313.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/314.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/315.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/316.tif.

Page:1974:IAS

[PW74a] E. S. Page and L. Waller. An interactive approach for the solution of a class of discrete optimisation problems. The Computer Journal, 17(1):69–74, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
01/tiff/69.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/
70.tif;   http://www3.oup.

                                        co.uk/computer_journal/hdb/
                                        Volume_17/Issue_01/tiff/71.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        17/Issue_01/tiff/72.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        01/tiff/73.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_17/Issue_01/tiff/
                                        74.tif.

Proll:1974:COD

[PW74b] L. G. Proll and F. P. Wyman. Correspondence: Observations on a decision rule for binary programming. The Computer Journal, 17(2):187–188, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        02/tiff/187.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        02/tiff/188.tif. See [Wym73].

Page:1981:EAM

[PW81a] I. P. Page and R. T. Wood. Empirical analysis of a moving head disc model with two heads separated by a fixed number of tracks. The Computer Journal, 24(4):339–342, November 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        04/tiff/339.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        04/tiff/340.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/341.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/342.tif.

Page:1981:SSA

[PW81b] I. P. Page and R. T. Wood.

Shaded-picture scanning attachment for incremental plotters. The Computer Journal, 24(3):287, August 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/287.tif.

Petre:1990:LMP

[PW90] Marian Petre and R. Winder.

On languages, models and programming styles. The Computer Journal, 33(2):173–180, April 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/173.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/174.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/175.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/176.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/177.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_33/Issue_
                                        02/tiff/178.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        02/tiff/179.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        02/tiff/180.tif.

Pavey:1993:DES

[PW93] D. J. Pavey and L. A. Winsborrow. Demonstrating equivalence of source code and PROM contents. The Computer Journal, 36(7):654–667, ???? 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_36/Issue_07/
                                        Vol36_07.body.html#AbstractPavey.

Patterson:1994:CAA

[PW94] J. W. Patterson and P. J. Willis. Computer assisted animation: 2D or not 2D? The Computer Journal, 37(10):829–839, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_37/Issue_10/
                                        Vol37_10.body.html#AbstractPatterson.

Quinlan:1983:ICA

[Qui83] J. R. Quinlan. INFERNO: a cautious approach to uncertain interference. The Computer Journal, 26(3):255–269, August 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_

journal/hdb/Volume_26/Issue_
03/tiff/255.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/256.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/257.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/258.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/259.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/260.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/261.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/262.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/263.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/264.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/265.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/266.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/267.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/268.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/269.tif.

Qian:1996:EFC

[QW96] Haifeng Qian and Jie Wu. Enhanced Fibonacci cubes. The Computer Journal, 39(4):331– 345, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://www.oup.co.uk/jnls/
                                        list/comjnl/hdb/Volume_39/
                                        Issue_04/390331.sgm.abs.html;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_39/Issue_04/
                                        Vol39_04.body.html#AbstractQian.

Quin:1989:DCS

[QY89] A. Quin and Y. Yanagisawa. On data compaction of scanning curves. The Computer Journal, 32(6):563–566, December 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        06/tiff/563.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        06/tiff/564.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        06/tiff/565.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        06/tiff/566.tif.

Rachman:1965:CRE

[Rac65]  A. Rachman.  Computa-

tion of rotational energy levels of rigid asymmetric top molecules. The Computer Journal, 8(2):147–149, July 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/080147.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/147.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/148.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/149.tif. See [Duk66,

Jon66].

Raita:1987:ASF

[Rai87]  T. Raita.  An automatic

system for file compression. The Computer Journal, 30 (1):80–86, February 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
01/tiff/80.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_01/tiff/
81.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_30/Issue_01/tiff/82.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
30/Issue_01/tiff/83.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
01/tiff/84.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_01/tiff/
85.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_30/Issue_01/tiff/86.
tif.

Ralston:1961:STC

[Ral61] A. Ralston. Some theoretical and computational matters relating to predictor-corrector methods of numerical integration. The Computer Journal, 4(1):64–67, April 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        01/040064.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        01/tiff/64.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_04/Issue_01/tiff/
                                        65.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_04/Issue_01/tiff/66.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        04/Issue_01/tiff/67.tif.

Ramsay:1970:FGM

[Ram70] J. O. Ramsay. A family of gradient methods for optimization. The Computer Journal, 13(4):413–417, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/130413.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/413.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/414.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_13/Issue_
04/tiff/415.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/416.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/417.tif.

Ramsay:1985:ECS

[Ram85] Allan Ramsay. On efficient context switching. The Computer Journal, 28(4):375–378, August 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/375.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/376.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/377.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/378.tif.

Ramsey:1999:ESE

[Ram99] Norman Ramsey. Eliminating spurious error messages using exceptions, polymorphism, and higher-order functions. The Computer Journal, 42(5):360– 372, 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
05/420360.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
05/Appendix.html.

Randall:1968:NGF

[Ran68a] T. J. Randall. Note on a general finite-difference formula for the solution of axially symmetric fields. The Computer Journal, 11(3):332–333, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        03/110332.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        03/tiff/332.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        03/tiff/333.tif.

Randall:1968:NEO

[Ran68b] T. J. Randall. A note on the estimation of the optimum successive overrelaxation parameter for Laplace’s equation. The Computer Journal, 10(4):400–401, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/100400.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/tiff/400.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/tiff/401.tif.

Randell:1971:LAM

[Ran71] B. Randell. Ludgate’s analytical machine of 1909. The Computer Journal, 14(3):317–326, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/140317.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/317.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/318.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/319.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/320.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/321.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/322.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/323.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/324.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/325.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/326.tif. See corre-

spondence [WR72].

Randall:1974:IDF

[Ran74] F. E. Randall. Interrogating date-sensitive files. The Computer Journal, 17(4):302– 305, November 1974. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        17/Issue_04/tiff/302.tif;
                                         http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        04/tiff/303.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        04/tiff/304.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        04/tiff/305.tif.

Randle:1979:CAE

[Ran79] J. Randle. Certification of Algorithm 3: Easter. The Computer Journal, 22(1):87, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        01/tiff/87.tif. See [Hil79c].

Randell:1986:SDS

[Ran86] B. Randell. System design and structuring. The Computer Journal, 29(4):300–306, August 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        04/tiff/300.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        04/tiff/301.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        04/tiff/302.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_

04/tiff/303.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/304.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/305.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/306.tif.

Robinson:1985:DDS

[RB85] J. Robinson and A. Burns.

A dialogue development system for the design and implementation of user interfaces in Ada. The Computer Journal, 28(1):22–28, February 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
01/tiff/22.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_01/tiff/
23.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_28/Issue_01/tiff/24.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
28/Issue_01/tiff/25.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
01/tiff/26.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_01/tiff/
27.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_28/Issue_01/tiff/28.
tif.

Rom:1986:NTA

[RB86] M. Rom and S. Bergman. A new technique for automatic contouring and contour representation from machine-readable spatial data. The Computer Journal, 29(5):467–471, October 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        05/tiff/467.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        05/tiff/468.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        05/tiff/469.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        05/tiff/470.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        05/tiff/471.tif.

Rakotonirainy:1997:DOD

[RBCM97] Andry Rakotonirainy, Andrew Berry, Stephen Crawley, and Zoran Milosevic. Describing open distributed systems:A foundation. The Computer Journal, 40(8):479–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_40/Issue_08/
                                        Vol40_08.body.html#AbstractRakotonirainy.

Rohl:1970:LPF

[RC70] J. S. Rohl and G. Cordingley. List processing facilities in Atlas Autocode. The Computer Journal, 13(1):20–24, February 1970. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/130020.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/tiff/20.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_01/tiff/
21.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/22.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
13/Issue_01/tiff/23.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/tiff/24.tif.

Riddell:1980:CRS

[RD80] I. J. Riddell and L. M.

Delves. The comparison of routines for solving Fredholm integral equations of the second kind. The Computer Journal, 23(3):274–285, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/274.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/275.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/276.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/277.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_

                                        03/tiff/278.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        03/tiff/279.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        03/tiff/280.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        03/tiff/281.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        03/tiff/282.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        03/tiff/283.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        03/tiff/284.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        03/tiff/285.tif.

Ruiz-Delgado:1995:ROO

[RDPS95] Antonio Ruiz-Delgado, David Pitt, and Colin Smythe. A review of object-oriented approaches in formal methods. The Computer Journal, 38 (10):777–784, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_38/Issue_10/
                                        Vol38_10.body.html#AbstractDelgado.

Rabel:1980:SNN

[RDT80] M. Rabel, D. Dubus, and A. Tosser. Saving of NAND/ NOR gates by inhibition or d-inhibition. The Computer Journal, 23(2):165–185, May 1980. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/165.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/166.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/167.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/168.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/169.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/170.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/171.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/172.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/173.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/174.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/175.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/176.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/177.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/178.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        02/tiff/179.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        02/tiff/180.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        02/tiff/181.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        02/tiff/182.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        02/tiff/183.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        02/tiff/184.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        02/tiff/185.tif.

Reader:1970:RDM

[Rea70] A. V. Reader. Random digits by a mincing process. The Computer Journal, 13(1):118, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        01/130118.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        01/tiff/118.tif.

Redfern:1960:EUD

[Red60] P. Redfern. Experience in using a Deuce computer for the family expenditure survey. The Computer Journal, 2(4):164–169, January 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/020164.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/164.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/165.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/166.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/167.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/168.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/169.tif.

Redish:1971:CNA

[Red71] K. A. Redish. Correspondence: (note on) Algorithm

44. The Computer Journal, 14(1):64, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/140064.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/tiff/64.tif. See [Bro69c,

Fie70, HL70].

Reeves:1967:DSD

[Ree67] C. M. Reeves. Description of a syntax-directed translator. The Computer Journal, 10(3):244–255, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/100244.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/244.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/245.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/246.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/247.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/248.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/249.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/250.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/251.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/252.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/253.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/254.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/255.tif.

Rees:1975:ED

[Ree75] D. J. Rees. The EMAS director. The Computer Journal, 18(2):122–130, May 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/122.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/123.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/124.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/125.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/126.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/127.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/128.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/129.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/130.tif.

Reeves:1979:AS

[Ree79a] C. M. Reeves. Algorithm 8:

SUD. The Computer Journal, 22(1):89, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_22/Issue_
                                        01/tiff/89.tif. Reprinted

from The Computer Bulletin,

September 1964, p. 67.

Reeves:1979:AB

[Ree79b]  C. M. Reeves.  Algorithm

9: Boyserf. The Computer Journal, 22(1):89–90, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        01/tiff/89.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_22/Issue_01/tiff/

90.tif. Reprinted from The Computer Bulletin, September 1964, p. 67.

Reeves:1979:FSD

[Ree79c] C. M. Reeves. Free store distribution under random fit allocation. I. The Computer Journal, 22(4):346–351, November 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        04/tiff/346.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        04/tiff/347.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        04/tiff/348.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        04/tiff/349.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_

04/tiff/350.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/351.tif.

Reeves:1980:FSD

[Ree80a] C. M. Reeves. Free store distribution under random fit allocation: Part 2. The Computer Journal, 23(4):298–306, November 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/298.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/299.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/300.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/301.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/302.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/303.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/304.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/305.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/306.tif.

Reeves:1980:CUD

[Ree80b] R. Reeves. Correspondence: Using data abstraction for logical design. The Computer Journal, 23(4):381–382, November 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/381.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/382.tif. See [Wel80,

JL80].

Reeves:1983:FSD

[Ree83] C. M. Reeves. Free store distribution under random fit allocation. III. The Computer Journal, 26(1):25–35, February 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
01/tiff/25.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_01/tiff/
26.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_26/Issue_01/tiff/27.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
26/Issue_01/tiff/28.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
01/tiff/29.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_01/tiff/
30.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_26/Issue_01/tiff/31.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
26/Issue_01/tiff/32.tif;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
01/tiff/33.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_01/tiff/
34.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_26/Issue_01/tiff/35.
tif.

Reeves:1984:CAE

[Ree84a]  C. M. Reeves.  Complex-

ity analyses of event set algorithms. The Computer Journal, 27(1):72–79, February 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
01/tiff/72.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_01/tiff/
73.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_27/Issue_01/tiff/74.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
27/Issue_01/tiff/75.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
01/tiff/76.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_01/tiff/
77.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_27/Issue_01/tiff/78.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
27/Issue_01/tiff/79.tif.

Reeves:1984:LMC

[Ree84b] C. M. Reeves. A lumped-state model of clustering in dynamic storage allocation. The Computer Journal, 27(2):135–142, May 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/135.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/136.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/137.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/138.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/139.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/140.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/141.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/142.tif.

Reeves:1987:CMF

[Ree87] C. M. Reeves. Chainmail: a model of first-fit memory allocation. The Computer Journal, 30(3):201–213, June 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/201.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/202.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/203.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/204.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/205.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/206.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/207.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/208.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/209.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/210.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/211.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/212.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/213.tif.

Reid:1966:MFO

[Rei66] J. K. Reid. A method for finding the optimum successive over-relaxation parameter. The Computer Journal, 9(2):200–204, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_09/Issue_
                                        02/090200.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        02/tiff/200.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        02/tiff/201.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        02/tiff/202.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        02/tiff/203.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        02/tiff/204.tif.

Reid:1967:NLS

[Rei67] J. K. Reid. A note on the least squares solutions of a band system of linear equations by Householder reductions. The Computer Journal, 10(2):188–189, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        02/100188.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        02/tiff/188.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        02/tiff/189.tif.

Reid:1971:NAP

[Rei71] J. K. Reid. A note on the approximation of plane regions. The Computer Journal, 14(3):307–308, August 1971. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/140307.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/307.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/308.tif.

Reingold:1981:CEP

[Rei81] Edward M. Reingold. A comment on the evaluation of Polish postfix expressions. The Computer Journal, 24(3):288, August 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/288.tif.

Repko:1972:ACD

[Rep72] Marya Goldman Repko. Animated cartoon design with a CRT. The Computer Journal, 15(4):316–320, November 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/150316.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/316.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/317.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/318.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/tiff/319.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/tiff/320.tif.

Rettelbach:1995:PBM

[Ret95] Michael Rettelbach. Probabilistic branching in Markovian process algebras. The Computer Journal, 38(7):590–599, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_38/Issue_07/
                                        Vol38_07.body.html#AbstractRettelbach.

Reynolds:1971:CPI

[Rey71a] C. F. Reynolds. CODIL, part 1. The importance of flexibility. The Computer Journal, 14(3):217–220, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/140217.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/tiff/217.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/tiff/218.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/tiff/219.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/tiff/220.tif.

Reynolds:1971:CPC

[Rey71b] C. F. Reynolds. CODIL, part 2. The CODIL language and its interpreter. The Computer Journal, 14(4):327–332, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/140327.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/327.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/328.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/329.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/330.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/331.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/332.tif.

Reynolds:1977:CSH

[Rey77]  R. A. Reynolds.  Charac-

ter string handling in Fortran. The Computer Journal, 20(4):325–329, November 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/325.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_

                                        04/tiff/326.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        04/tiff/327.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        04/tiff/328.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        04/tiff/329.tif.

Reynolds:1983:HPI

[Rey83] R. A. Reynolds. Human performance in interactive graphics operations. The Computer Journal, 26(1):93–94, February 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        01/tiff/93.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_26/Issue_01/tiff/
                                        94.tif.

Reynolds:1990:CAI

[Rey90]  C. F. Reynolds.  CODIL:

The architecture of an information language. The Computer Journal, 33(2):155–163, April 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        02/tiff/155.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        02/tiff/156.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        02/tiff/157.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_33/Issue_
02/tiff/158.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/159.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/160.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/161.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/162.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/163.tif.

Rohl:1986:RP

[RG86] J. S. Rohl and T. D. Gedeon.

The Reve’s puzzle. The Computer Journal, 29(2):187–188, April 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/187.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
02/tiff/188.tif.

Reiter:1995:SCR

[RG95] Michael Reiter and Li Gong.

Securing causal relationships in distributed systems. The Computer Journal, 38(8):633–642, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_38/Issue_08/
Vol38_08.body.html#AbstractReiter.

Robinson:1998:PSD

[RHHR98] Hugh Robinson, Pat Hall, Fiona Hovenden, and Janet Rachel. Postmodern software development. The Computer Journal, 41(6):363–375, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_41/Issue_06/
                                        Vol41_06.body.html#AbstractRobinson.

Robison:1989:EPC

[RHS89] A. D. Robison, B. J. Hafner, and S. S. Skiena. Eight pieces cannot cover a chess board (search technique). The Computer Journal, 32(6):567– 570, December 1989. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        32/Issue_06/tiff/567.tif;
                                         http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        06/tiff/568.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        06/tiff/569.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        06/tiff/570.tif.

Ribaudo:1995:ATS

[Rib95] Marina Ribaudo. On the aggregation techniques in stochastic Petri nets and stochastic process algebras. The Computer Journal, 38(7):600–611, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_38/Issue_07/
Vol38_07.body.html#AbstractRibaudo.

Richens:1958:IMT

[Ric58] R. H. Richens. Interlingual machine translation. The Computer Journal, 1(3):144– 147, October 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
01/Issue_03/010144.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
01/Issue_03/tiff/144.tif;
 http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/145.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/146.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/147.tif.

Richardson:1964:UEC

[Ric64] M. Richardson. User’s experience of CLEO. The Computer Journal, 7(2):102–103, July 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/070102.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/102.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/103.tif.

Ricci:1971:ARH

[Ric71] A. Ricci. An algorithm for the removal of hidden lines in 3D scenes. The Computer Journal, 14(4):375–377, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/140375.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/375.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/376.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/377.tif.

Ricci:1973:CGC

[Ric73] A. Ricci. A constructive geometry for computer graphics. The Computer Journal, 16(2):157–160, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/160157.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/157.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/158.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/159.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_16/Issue_
02/tiff/160.tif.

Richards:1988:WAW

[Ric88] D. Richards. On the worstpossible analysis of weighted comparison-based algorithms. The Computer Journal, 31(3): 276–278, June 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
31/Issue_03/tiff/276.tif;
 http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/277.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/278.tif.

Rigg:1964:OLC

[Rig64]  F. A. Rigg.  The opera-

tion of large computer systems. The Computer Journal, 7(3):169–172, October 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/070169.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/169.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/170.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/171.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/172.tif.

Ringland:1975:SFV

[Rin75] Gill Ringland. Software functional variability. The Computer Journal, 18(4):312–317, November 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        04/tiff/312.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        04/tiff/313.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        04/tiff/314.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        04/tiff/315.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        04/tiff/316.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        04/tiff/317.tif.

Rissanen:1999:HST

[Ris99a] J. Rissanen. Hypothesis selection and testing by the MDL principle. The Computer Journal, 42(4):260–269, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        04/420260.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        04/pdf/420260.pdf.

Rissanen:1999:DPM

[Ris99b] Jorma Rissanen. Discussion of paper ‘Minimum Message Length and Kolmogorov Complexity’ by C. S. Wallace and D. L. Dowe. The Computer Journal, 42(4):327–329, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
04/420327.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
04/pdf/420327.pdf.

Rissanen:1999:R

[Ris99c]  Jorma Rissanen.  Rejoin-

der. The Computer Journal, 42(4):343–344, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
04/420343.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
04/pdf/420343.pdf.

Ritchie:1989:HTL

[Rit89] I. Ritchie. Hypertext-moving towards large volumes. The Computer Journal, 32(6):516– 523, December 1989. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
32/Issue_06/tiff/516.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/517.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_32/Issue_
06/tiff/518.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/519.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/520.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/521.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/522.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/523.tif.

Rivlin:1968:PAT

[Riv68] T. J. Rivlin. Polynomial approximation and the τ-method. The Computer Journal, 11 (3):337–338, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/110337.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/337.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/338.tif.

Rosenberg:1992:AML

[RKA92] J. Rosenberg, J. L. Keedy, and D. Abramson. Addressing mechanisms for large virtual memories. The Computer Journal, 35(4):369–375, August 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_35/Issue_04/
Vol35_04.body.html#AbstractRosenberg.

Australia.

Rohl:1972:NCA

[RL72] J. S. Rohl and J. A. Linn. A note on compiling arithmetic expressions. The Computer Journal, 15(1):13–14, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/150013.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/tiff/13.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_01/tiff/
14.tif.

Reid:1973:BRS

[RL73] J. K. Reid and D. W. Lewin.

Book reviews: Sparse Matrices and their Applications, edited by D. L. Rose and R. A. Willoughby, 1972; Introduction to Digital Computer Technology, by Louis Nashelsky, 1973. The Computer Journal, 16(3):222, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/222.tif.

Rokne:1975:ACI

[RL75] J. Rokne and P. Lancaster. Algorithm 86: Complex interval arithmetic. The Computer Journal, 18(1):83–85, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
01/tiff/83.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_01/tiff/
84.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_18/Issue_01/tiff/85.
tif.

Ramamohanarao:1982:DHS

[RL82] K. Ramamohanarao and John W. Lloyd. Dynamic hashing schemes. The Computer Journal, 25(4):478–485, November 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/478.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/479.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/480.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/481.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/482.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/483.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/484.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/485.tif.

Raj:1989:CMS

[RL89] R. K. Raj and H. M. Levy. A compositional model for software reuse. The Computer Journal, 32(4):312–322, August 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/312.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/313.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/314.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/315.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/316.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/317.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/318.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/319.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/320.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/321.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_32/Issue_
                                        04/tiff/322.tif.

Ranai:1994:SMF

[RLD94] Kisenchand Ranai, Koon Seng Lim, and R. H. Deng. Simulating a multimedia FDDI backbone network. The Computer Journal, 37(3):188–198, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_37/Issue_03/
                                        Vol37_03.body.html#AbstractRanai.

Rook:1999:BR

[RLT99] Jim Rook, Ian J. Lloyd, and Kenneth J. Turner. Book reviews. The Computer Journal, 42(2):150–152, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        02/420150.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        02/pdf/420150.pdf.

Robertson:1969:NNI

[RM69] H. H. Robertson and M. J. Mc-Cann. A note on the numerical integration of conservative systems of first-order ordinary differential equations. The Computer Journal, 12(1):81, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_

01/120081.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
01/tiff/81.tif.

Rahman:1989:EBS

[RM89] M. Rahman and S. Misbahuddin. Effect of a binary symmetric channel on the synchronisation recovery of variable length codes. The Computer Journal, 32(3):246–251, June 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/246.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/247.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/248.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/249.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/250.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/251.tif.

Robinson:1959:DIP

[Rob59] C. Robinson. DEUCE interpretive programs. The Computer Journal, 1(4):172–176, January 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/010172.sgm.abs.html;

                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        04/tiff/172.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        04/tiff/173.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        04/tiff/174.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        04/tiff/175.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        04/tiff/176.tif.

Robinson:1963:BPA

[Rob63] F. D. Robinson. The background of the PERT algorithm. The Computer Journal, 5(4):297–300, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        04/050297.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        04/tiff/297.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        04/tiff/298.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        04/tiff/299.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        04/tiff/300.tif.

Robinson:1975:DIB

[Rob75] K. A. Robinson. Database — the ideas behind the ideas. The Computer Journal, 18(1):7–11, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
01/tiff/10.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_01/tiff/
11.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_18/Issue_01/tiff/7.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
18/Issue_01/tiff/8.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
01/tiff/9.tif.

Robson:1977:WCF

[Rob77] J. M. Robson. Worst case fragmentation of first fit and best fit storage allocation strategies. The Computer Journal, 20(3):242–244, August 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/242.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/243.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/244.tif.

Robinson:1979:EED

[Rob79] K. A. Robinson. An entity/ event data modelling method. The Computer Journal, 22 (3):270–279, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/270.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/271.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/272.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/273.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/274.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/275.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/276.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/277.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/278.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/279.tif. See discus-

sion [Var79].

Robinson:1980:CJS

[Rob80] G. L. Robinson. Correspondence: Jumping to some purpose. The Computer Journal, 23(3):288, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_

journal/hdb/Volume_23/Issue_
03/tiff/288.tif. See [ASG79,

Hil80, Tay80, MR82].

Roberts:1988:FSQ

[Rob88] W. T. Roberts. A formal specification of the QMC message system: the underlying abstract model. The Computer Journal, 31(4):313–324, August 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/313.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/314.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/315.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/316.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/317.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/318.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/319.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/320.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/321.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/322.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        04/tiff/323.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        04/tiff/324.tif.

Roberts:1990:ESC

[Rob90] H. J. Roberts. Expert systems clubs: Design methods. The Computer Journal, 33(6):556–561, December 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/556.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/557.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/558.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/559.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/560.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/561.tif.

Rodabaugh:1970:SC

[Rod70] D. J. Rodabaugh. On stable correctors. The Computer Journal, 13(1):98–100, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        01/130098.sgm.abs.html;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/tiff/100.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/tiff/98.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_01/tiff/
99.tif.

Rohl:1968:NBN

[Roh68] J. S. Rohl. A note on Backus Naur Form. The Computer Journal, 10(4):336–337, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/100336.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/336.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/337.tif.

Rohlf:1973:AHC

[Roh73] F. James Rohlf. Algorithm 76: Hierarchical clustering using the minimum spanning tree. The Computer Journal, 16(1):93–95, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
01/tiff/93.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_01/tiff/
94.tif;   http://www3.oup.
co.uk/computer_journal/hdb/

Volume_16/Issue_01/tiff/95.
tif.

Rohlf:1974:ADP

[Roh74]  R. James Rohlf.  Algo-

rithm 81: Dendrogram plot. The Computer Journal, 17 (1):89–91, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
01/tiff/89.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/
90.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_17/Issue_01/tiff/91.
tif.

Rohlf:1975:NAC

[Roh75] F. J. Rohlf. A new approach to the computation of the Jardine-Sibson B_{k} clusters. The Computer Journal, 18(2):164–168, May 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/164.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/165.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/166.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/167.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/168.tif.

Rohl:1976:PIF

[Roh76]  J. S. Rohl.  Programming

improvements to Fike’s algorithm for generating permutations. The Computer Journal, 19(2):156–159, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/156.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/157.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/158.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/159.tif.

Rohl:1978:GPC

[Roh78]  J. S. Rohl.  Generating

permutations by choosing. The Computer Journal, 21 (4):302–305, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/302.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/303.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/304.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/305.tif.

Rohl:1987:THD

[Roh87] J. S. Rohl. Towers of Hanoi: the derivation of some iterative versions. The Computer Journal, 30(1):70–76, February 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
01/tiff/70.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_01/tiff/
71.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_30/Issue_01/tiff/72.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
30/Issue_01/tiff/73.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
01/tiff/74.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_01/tiff/
75.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_30/Issue_01/tiff/76.
tif.

Rohl:1990:RIL

[Roh90] J. S. Rohl. Reversing is not inherent in lexicographical permutation generation. The Computer Journal, 33(3):274–278, June 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/274.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/275.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/276.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/277.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/278.tif.

Rohl:1991:OPP

[Roh91] J. S. Rohl. Ord-Smith’s pseudolexicographical permutation procedure is the Tompkins-Paige algorithm. The Computer Journal, 34(6):569–570, December 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/569.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/570.tif.

Rojas:1995:CCS

[Roj95] Isabel Rojas. Compositional construction of SWN models. The Computer Journal, 38(7):612–621, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_38/Issue_07/
Vol38_07.body.html#AbstractRojas.

Romanovsky:1999:SCC

[Rom99] Alexander Romanovsky. On structuring cooperative and competitive concurrent systems. The Computer Journal, 42 (8):627–637, 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
42/Issue_08/420627.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
42/Issue_08/pdf/420627.pdf.

Rosenbrock:1960:AMF

[Ros60] H. H. Rosenbrock. An automatic method for finding the greatest or least value of a function. The Computer Journal, 3(3):175–184, October 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/030175.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/175.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/176.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/177.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/178.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/179.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/180.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/181.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/182.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/183.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/184.tif.

Rosenbrock:1963:SGI

[Ros63a] H. H. Rosenbrock. Some general implicit processes for the numerical solution of differential equations. The Computer Journal, 5(4):329–331, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/050329.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/329.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/330.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/331.tif.

Ross:1963:CDL

[Ros63b] H. McG Ross. A convention to distinguish letter O from numeral zero. The Computer Journal, 6(1):49, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/060049.sgm.abs.html;
http://www3.oup.co.uk/computer_

                                        journal/hdb/Volume_06/Issue_
                                        01/tiff/49.tif.

Rose:1964:CSE

[Ros64a] M. J. Rose. Classification of a set of elements. The Computer Journal, 7(3):208–211, October 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        03/070208.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        03/tiff/208.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        03/tiff/209.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        03/tiff/210.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        03/tiff/211.tif.

Ross:1964:CC

[Ros64b] H. McG Ross. The I.S.O. character code. The Computer Journal, 7(3):197–202, October 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        03/070197.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        03/tiff/197.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        03/tiff/198.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_

03/tiff/199.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/200.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/201.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/202.tif.

Rosin:1966:AES

[Ros66] R. F. Rosin. An approach to executive system maintenance in disk-based systems. The Computer Journal, 9(3):242–247, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/090242.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/242.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/243.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/244.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/245.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/246.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/247.tif.

Rosenthal:1967:AAM

[Ros67]  L. E. Rosenthal.  Ana-

lytic algebraic manipulation. The Computer Journal, 10 (3):265–270, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/100265.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/265.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/266.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/267.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/268.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/269.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/270.tif.

Ross:1970:BSD

[Ros70] H. McGregor Ross. The British Standard Data Code and how to exploit it. The Computer Journal, 13(3):223–229, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/130223.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/223.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_

03/tiff/224.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/225.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/226.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/227.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/228.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/229.tif.

Rosenquist:1982:ELC

[Ros82] C. J. Rosenquist. Entity life cycle models and their applicability to information systems development life cycles, a framework for information systems design and implementation. The Computer Journal, 25(3):307–315, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/307.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/308.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/309.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/310.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/311.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        03/tiff/312.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        03/tiff/313.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        03/tiff/314.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        03/tiff/315.tif.

Rossi:1993:PDE

[Ros93] G. Rossi. Programs as data in an extended Prolog. The Computer Journal, 36(3):217–226, June 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_36/Issue_03/
                                        Vol36_03.body.html#AbstractRossi.

Rousell:1962:PRN

[Rou62] A. R. Rousell. A progress report on NEBULA. The Computer Journal, 5(3):162–163, October 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        03/050162.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        03/tiff/162.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        03/tiff/163.tif.

Roy:1978:EPA

[Roy78]  M. K. Roy.  Evaluation

of permutation algorithms.

The Computer Journal, 21 (4):296–301, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/296.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/297.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/298.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/299.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/300.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/301.tif.

Richards:1963:BRA

[RP63] B. Richards and E. S. Page.

Book reviews: Annual Review in Automatic Programming, Volume 3, edited by R. Goodman; Proceedings of a Harvard Symposium on Digital Computers and Their Applications, edited by A. G. Oettinger, 1962. The Computer Journal, 6(3):263–263, 270–270, October 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/263.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/270.tif.

Rechard:1969:ETA

[RS69] O. W. Rechard and R. H. Stark. Equivalence of two algorithms for Cooper’s generalised factorial function. The Computer Journal, 12(1):33–37, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        01/120033.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        01/tiff/33.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_12/Issue_01/tiff/
                                        34.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_12/Issue_01/tiff/35.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        12/Issue_01/tiff/36.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        01/tiff/37.tif.

Redish:1987:EMP

[RS87] K. A. Redish and W. F. Smyth. Evaluating measures of program quality. The Computer Journal, 30(3):228–232, June 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        03/tiff/228.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        03/tiff/229.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_

03/tiff/230.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/231.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/232.tif.

Razaz:1988:TPM

[RS88] M. Razaz and J. L. Schonfelder. Test procedures for measurement of floating-point characteristics of computing environments. The Computer Journal, 31(1):12–16, February 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
01/tiff/12.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_01/tiff/
13.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_31/Issue_01/tiff/14.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
31/Issue_01/tiff/15.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
01/tiff/16.tif.

Ramshaw:1987:IRP

[RSS87] M. Ramshaw, A. P. Steward, and D. Stoves. An intelligent retrieval and presentation system for management information. The Computer Journal, 30(6):485–486, December 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

                                        journal/hdb/Volume_30/Issue_
                                        06/tiff/485.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        06/tiff/486.tif.

Robinson:1976:EAF

[RT76] S. K. Robinson and I. S. Torsun. An empirical analysis of FORTRAN programs. The Computer Journal, 19 (1):56–62, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        01/tiff/56.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_19/Issue_01/tiff/
                                        57.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_19/Issue_01/tiff/58.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        19/Issue_01/tiff/59.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        01/tiff/60.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_19/Issue_01/tiff/
                                        61.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_19/Issue_01/tiff/62.
                                        tif.

Rodgers:1978:CAM

[RT78] M. H. Rodgers and R. H. Thomason. Certification of Algorithm 1: Mintree. The Computer Journal, 21 (3):276–277, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/276.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/277.tif. See [Obr78a].

Rudall:1967:AC

[Rud67a]  B. H. Rudall.  Algorithm

19: Complex. The Computer Journal, 10(1):112–113, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/112.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/113.tif.

Rudall:1967:NAC

[Rud67b] B. H. Rudall. Note on Algorithm 19: Complex. The Computer Journal, 10(2):208– 209, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
10/Issue_02/tiff/208.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/209.tif.

Russell:1963:PDT

[Rus63] D. G. Russell. The problems of data transmission systems in a general manufacturing dat processing installation. The Computer Journal, 6(3):210–213, October 1963.

CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/060210.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/210.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/211.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/212.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/213.tif.

Rushby:1986:KEA

[Rus86] N. Rushby. A knowledge engineering approach to instructional design. The Computer Journal, 29(5):385–389, October 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/385.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/386.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/387.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/388.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/389.tif.

Ruzicka:1991:NEI

[Ruž91] P. Ružička. A note on the efficiency of an interval routing algorithm. The Computer Journal, 34(5):475–476, October 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/475.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/476.tif.

Romao:1978:AMN

[RV78] J. C. Romao and R. Vilela Mendes. Algorithms for multidimensional numerical integration with singularities. The Computer Journal, 21 (4):355–358, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/355.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/356.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/357.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/358.tif.

Rennhackkamp:1988:UAM

[Rv88] M. H. Rennhackkamp and S. H.

von Solms. On using an abstract model of distributed database concurrency control methods. The Computer Journal, 31(6):570–572, December 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/570.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/571.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/572.tif.

Ramanathan:1993:FTI

[RV93] S. Ramanathan and P. Venkat Rangan. Feedback techniques for intra-media continuity and inter-media synchronization in distributed multimedia systems. The Computer Journal, 36(1):19–31, February 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_01/
Vol36_01.body.html#AbstractRamanathan.

Rollett:1961:ESC

[RW61] J. S. Rollett and J. H. Wilkinson. An efficient scheme for the co-diagonalization of a symmetric matrix by Givens’ method in a computer with a two-level store. The Computer Journal, 4(2):177–180, July 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/040177.sgm.abs.html;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/177.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/178.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/179.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/180.tif.

Rollett:1962:LEG

[RW62] J. S. Rollett and J. H. Wilkinson. Letter to the editor: [givens and householder processes]. The Computer Journal, 4(4):279, January 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/
cgi/reprint/4/4/279.

Ranyard:1967:OAT

[RW67] J. C. Ranyard and A. Wren.

The optimum arrangement of towers in an electric power transmission line. The Computer Journal, 10(2):157–161, August 1967. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/100157.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/157.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/158.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_10/Issue_
                                        02/tiff/159.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        02/tiff/160.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        02/tiff/161.tif.

Riley:1968:GTD

[RW68] J. L. Riley and P. H. Walker. Generation of time delays on analogue computers. The Computer Journal, 11(1):72–76, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        01/110072.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        01/tiff/72.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_11/Issue_01/tiff/
                                        73.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_11/Issue_01/tiff/74.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        11/Issue_01/tiff/75.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        01/tiff/76.tif.

Rozenberg:1974:GMP

[RW74] G. Rozenberg and D. Wood. Generative models for parallel processes. The Computer Journal, 17(4):344–348, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_

journal/hdb/Volume_17/Issue_
04/tiff/344.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/345.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/346.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/347.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/348.tif.

Ridgion:1964:ACS

[RWT64] J. M. Ridgion, A. J. Willmott, and J. H. Thewlis. An analogue computer simulation of a Cowper Stove. The Computer Journal, 7(3):188–196, October 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/070188.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/188.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/189.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/190.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/191.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/192.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_

                                        03/tiff/193.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        03/tiff/194.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        03/tiff/195.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        03/tiff/196.tif.

Ryan:1963:LFT

[Rya63]  W. S. Ryan.  LEAPS —

the first three years. The Computer Journal, 6(1):6–13, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        01/060006.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        01/tiff/10.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_06/Issue_01/tiff/
                                        11.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_06/Issue_01/tiff/12.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        06/Issue_01/tiff/13.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        01/tiff/6.tif;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_06/Issue_01/tiff/
                                        7.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_06/Issue_01/tiff/8.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        06/Issue_01/tiff/9.tif.

Ryan:1966:DIA

[Rya66]  J. T. Ryan.  A direction-

independent algorithm for determining the forward and backward compute point for a term or subscript during compilation. The Computer Journal, 9(2):157–160, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/090157.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/157.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/158.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/159.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/160.tif.

Ryder:1964:NAC

[Ryd64] K. L. Ryder. Note on an ALGOL 60 compiler for Pegasus I. The Computer Journal, 6(4):336–338, January 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/060336.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/336.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_

                                        04/tiff/337.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        04/tiff/338.tif.

Rzevski:1985:CAI

[Rze85] George Rzevski. On criteria for assessing an information theory. The Computer Journal, 28(3):200–202, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        03/tiff/200.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        03/tiff/201.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        03/tiff/202.tif.

Schneck:1973:FFO

[SA73] P. B. Schneck and Ellinor Angel. A FORTRAN to FORTRAN optimising compiler. The Computer Journal, 16(4):322–330, November 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        04/160322.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        04/tiff/322.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        04/tiff/323.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        04/tiff/324.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/325.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/326.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/327.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/328.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/329.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/330.tif.

Sandys:1979:CPP

[SA79] M. A. Sandys and Aleph Null. Correspondence: Point in polygon. The Computer Journal, 22(3):287–288, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/287.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/288.tif.

Stecher:1981:ACC

[SA81] P. Stecher and Volker Allenstein. Application controller concept: a first experience. The Computer Journal, 24(2):97–106, May 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_

journal/hdb/Volume_24/Issue_
02/tiff/100.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/101.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/102.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/103.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/104.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/105.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/106.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/97.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_02/tiff/
98.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_24/Issue_02/tiff/99.
tif.

Suraweera:1988:AMA

[SAA88] F. Suraweera and J. M. Al-Anzy. Analysis of a modified address calculation sorting algorithm. The Computer Journal, 31(6):561–563, December 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/561.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_

06/tiff/562.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/563.tif.

Sabin:1972:CCS

[Sab72] M. A. Sabin. Correspondence:

Comments on some algorithms for the representation of curves by straight line segments. The Computer Journal, 15(2):104, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/150104.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/104.tif. See [Ecc71,

McC70b, Smi71, Coo72b].

Sabin:1976:MDI

[Sab76]  M. A. Sabin.  A method

for displaying the intersection curve of two quadric surfaces. The Computer Journal, 19(4):336–338, November 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/336.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/337.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/338.tif.

Sack:1962:NCT

[Sac62] R. A. Sack. Newton–Cotes type quadrature formulas with terminal corrections. The Computer Journal, 5(3):230–237, October 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/050230.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/230.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/231.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/232.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/233.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/234.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/235.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/236.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/237.tif.

Sadri:1992:DDD

[Sad92]  F. Sadri.  Data dependen-

cies in deductive databases. The Computer Journal, 35 (5):A415–??, October 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_35/Issue_
05/tiff/A415.tif;   http:

//www3.oup.co.uk/computer_
journal/Volume_35/Issue_05/
35000A415.sgm.abs.html.

Scowen:1971:SPW

[SAHS71] R. S. Scowen, D. Allin, A. L.

Hillman, and M. Shimell. SOAP

— a program which documents and edits ALGOL 60 programs. The Computer Journal, 14(2):133–135, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/140133.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/133.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/134.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/135.tif.

Sale:1968:CMS

[Sal68] A. H. J. Sale. The calculation of e to many significant digits. The Computer Journal, 11(2):229–230, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/110229.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/229.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/230.tif.

Sale:1970:AAS

[Sal70a] A. H. J. Sale. Algorithm 54: Approximation of straight lines. The Computer Journal, 13(4):422–424, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/422.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/423.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/424.tif.

Sale:1970:NAI

[Sal70b] A. H. J. Sale. Note on Algorithm 42: Interpolation by certain quintic splines. The Computer Journal, 13(1):115, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        01/tiff/115.tif. See [Spä69b].

Sale:1970:NAC

[Sal70c] A. H. J. Sale. Note on Algorithm 47: A clustering algorithm. The Computer Journal, 13(3):326, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/326.tif. See [van70,

vR70b, Mux70].

Sale:1971:AIC

[Sal71a] A. H. J. Sale. Algorithm 65:

An improved clustering algorithm. The Computer Journal, 14(1):104–106, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/tiff/104.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/tiff/105.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/tiff/106.tif.

Sale:1971:CFS

[Sal71b] A. H. J. Sale. The classification of FORTRAN statements. The Computer Journal, 14(1):10–12, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/140010.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/tiff/10.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_01/tiff/
11.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_14/Issue_01/tiff/12.
tif.

Salzer:1972:LIC

[Sal72] H. E. Salzer. Lagrangian interpolation at the Chebyshev points x_{n,ν} ≡ cos(νπ/n), ν = 0(1)n; some unnoted advantages. The Computer Journal, 15(2):156–159, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/150156.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/156.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/157.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/158.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/159.tif. See corre-

spondence [Sal73].

Salzer:1973:CLI

[Sal73] H. E. Salzer. Correspondence: On “Lagrangian interpolation at the Chebyshev points x_{n,ν} ≡ cos(νπ/n), ν = 0(1)n”. The Computer Journal, 16(2):190, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/190.tif. See [Sal72].

Salomon:1974:EAW

[Sal74] Kenneth B. Salomon. An efficient algorithm which determines the output from a sequential machine for regular inputs. The Computer Journal, 17(4):349–354, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/349.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/350.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/351.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/352.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/353.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/354.tif.

Salomon:1987:DEN

[Sal87] D. Salomon. A design for an efficient NOR-gate only, binary-ripple adder with carrycompletion-detection logic. The Computer Journal, 30(3):283– 285, June 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
30/Issue_03/tiff/283.tif;
 http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/284.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/285.tif.

Samet:1963:CHR

[Sam63] Paul A. Samet. Correspondence: A hardware representation for ALGOL 60 using Creed Teleprinter equipment. The Computer Journal, 6(3):286, October 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        03/tiff/286.tif. See corre-

spondence [Bar64, Tay64].

Samet:1965:EAB

[Sam65] P. A. Samet. The efficient administration of blocks in AL-GOL. The Computer Journal, 8(1):21–23, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        01/080021.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        01/tiff/21.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_08/Issue_01/tiff/
                                        22.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_08/Issue_01/tiff/23.
                                        tif.

Samet:1966:CES

[Sam66] Paul A. Samet. Correspondence: On “The Egdon system for the KDF9”. The Computer Journal, 9(2):174, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_

02/tiff/174.tif. See [BHJV66]

and correspondence [Bar66].

Samet:1968:DCS

[Sam68]  P. A. Samet.  Discussion

and correspondence: software requirements of universities. The Computer Journal, 11(2):236–240, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/110236.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/236.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/237.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/238.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/239.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/240.tif.

Samet:1972:CDS

[Sam72a] P. A. Samet. Calculation of a double-length square root from double-length number using single precision techniques. The Computer Journal, 15(2):116, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/116.tif.

Samet:1972:ING

[Sam72b] P. A. Samet. Insight, not numbers. The Computer Journal, 15(1):88–92, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        01/150088.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        01/tiff/88.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_15/Issue_01/tiff/
                                        89.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_15/Issue_01/tiff/90.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        15/Issue_01/tiff/91.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        01/tiff/92.tif.

Samson:1976:TOA

[Sam76] W. B. Samson. Testing overflow algorithms for a table of variable size. The Computer Journal, 19(1):92, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        01/tiff/92.tif.

Samson:1981:HTC

[Sam81] W. B. Samson. Hash table collision handling on storage devices with latency. The Computer Journal, 24(2):130–131, May 1981. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/130.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/131.tif.

Subieta:1988:OCC

[SAM88] K. Subieta, K. Anacki, and M. Missala. Overlay clusters for C programs. The Computer Journal, 31(3):285–286, June 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/285.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
03/tiff/286.tif.

Sandys:1975:CCC

[San75] M. A. Sandys. Correspondence:

On “Continuity Conditions for Spline Curves”. The Computer Journal, 18(3):282–283, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/282.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/283.tif. See [Man74].

Sanderson:1999:BR

[San99] Mark Sanderson. Book reviews. The Computer Journal, 42(8):713–715, ???? 1999.

CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
08/420713.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
08/pdf/420713.pdf.

Sarda:1990:AQL

[Sar90] N. L. Sarda. Algebra and query language for A historical data model. The Computer Journal, 33(1):11–18, February 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
01/tiff/11.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_01/tiff/
12.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_33/Issue_01/tiff/13.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/14.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
01/tiff/15.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_01/tiff/
16.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_33/Issue_01/tiff/17.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/18.tif.

Sarbo:1996:BSK

[Sar96] J. J. Sarbo. Building subknowledge bases using concept lattices. The Computer Journal, 39(10):868–??, ???? 1996. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www.oup.co.uk/computer_
journal/Volume_39/Issue_10/
Vol39_10.body.html#AbstractSarbo;
 http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_10/
Vol39_10.body.html#AbstractSarbo.

Saunders:1972:DAC

[Sau72] R. P. Saunders. A description and analysis of character maps. The Computer Journal, 15(2):160–169, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/150160.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/160.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/161.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/162.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/163.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/164.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/165.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_

                                        02/tiff/166.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        02/tiff/167.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        02/tiff/168.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        02/tiff/169.tif.

Savage:1969:JOT

[Sav69] M. J. Savage. Junction optimisation technique. The Computer Journal, 12(3):268–272, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        03/120268.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        03/tiff/268.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        03/tiff/269.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        03/tiff/270.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        03/tiff/271.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        03/tiff/272.tif. See erratum

[Sav70].

Savage:1970:EJO

[Sav70]  M. J. Savage.  Erratum:

“Junction optimisation technique”. The Computer Journal, 13(1):120, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/tiff/120.tif. See [Sav69].

Silvey:1970:HCT

[SB70] T. I. Silvey and J. R. Barker.

Hybrid computing techniques for solving parabolic and hyperbolic partial differential equations. The Computer Journal, 13(2):164–170, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/130164.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/164.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/165.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/166.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/167.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/168.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/169.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/170.tif.

Stephenson:1981:CIB

[SB81] J. Stephenson and G. Brown.

Correspondence: Implementation of BEDSOCS: An interactive simulation language. The Computer Journal, 24(2):191–192, May 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        02/tiff/191.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        02/tiff/192.tif. See [ER80].

Samson:1985:ROD

[SB85] W. B. Samson and A. Bendell. Rank order distributions and secondary key indexing. The Computer Journal, 28(3):309–312, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        03/tiff/309.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        03/tiff/310.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        03/tiff/311.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        03/tiff/312.tif.

Schieder:1999:ACL

[SB99] Birgit Schieder and Manfred Broy. Adapting calculational logic to the undefined. The Computer Journal, 42(2):73–81, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
02/420073.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
02/pdf/420073.pdf.

Sommerville:1994:CSD

[SBRS94] Ian Sommerville, Richard Bentley, Tom Rodden, and Peter Sawyer. Cooperative systems design. The Computer Journal, 37(5):357–366, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_37/Issue_05/
Vol37_05.body.html#AbstractSommerville.

Squire:1971:CBT

[SC71] William Squire and M. G. Cox.

Correspondence: On “A bracketing technique for computing a zero of a function”. The Computer Journal, 14(1):72, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/tiff/72.tif. See [Cox70].

Sentance:1981:DES

[SC81] W. A. Sentance and I. P. Cliff.

The determination of eigenvalues of symmetric quindiagonal matrices. The Computer Journal, 24(2):177–179, May 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

journal/hdb/Volume_24/Issue_
02/tiff/177.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/178.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/179.tif.

Sernadas:1987:RCM

[SC87] Cristina Sernadas and R. Carapuca. The role of conceptual modelling abstractions in compiler development. The Computer Journal, 30(4):328–336, August 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/328.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/329.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/330.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/331.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/332.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/333.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/334.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/335.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_30/Issue_
04/tiff/336.tif.

Sherrell:1995:FIS

[SC95] Linda B. Sherrell and Doris L.

Carver. FunZ: An intermediate specification language. The Computer Journal, 38 (3):193–206, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_38/Issue_03/
Vol38_03.body.html#AbstractSherrell.

Scarrott:1985:ILB

[Sca85]  G. G. Scarrott.  Informa-

tion, the life blood of organisation. The Computer Journal, 28(3):203–205, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/203.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/204.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/205.tif.

Scarrott:1989:NI

[Sca89] G. G. Scarrott. The nature of information. The Computer Journal, 32(3):262–266, June 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_

03/tiff/262.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/263.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/264.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/265.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/266.tif.

Seed:1996:OOG

[SCC^{+}96] G. M. Seed, D. E. R. Clark, J. R. Corney, R. Tuttle, and G. Little. Object-oriented graph-based geometric feature recognition. The Computer Journal, 39(9):808–811, ???? 1996. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www.oup.co.uk/computer_
journal/Volume_39/Issue_09/
Vol39_09.body.html#AbstractSeed;
 http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_09/
Vol39_09.body.html#AbstractSeed.

Schorr:1965:ADU

[Sch65] H. Schorr. Analytic differentiation using a syntax-directed compiler. The Computer Journal, 7(4):290–298, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/070290.sgm.abs.html;
http://www3.oup.co.uk/computer_

journal/hdb/Volume_07/Issue_
04/tiff/290.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/291.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/292.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/293.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/294.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/295.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/296.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/297.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/298.tif.

Schlageter:1976:PLV

[Sch76]  G. Schlageter.  The prob-

lem of lock by value in large data bases. The Computer Journal, 19(1):17–20, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
01/tiff/17.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_01/tiff/
18.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_19/Issue_01/tiff/19.
tif;   http://www3.oup.co.

                                        uk/computer_journal/hdb/Volume_
                                        19/Issue_01/tiff/20.tif.

Scheuermann:1977:MIS

[Sch77a] P. Scheuermann. Modelling the information space in physical storage at different levels of detail. The Computer Journal, 20(3):248–256, August 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        03/tiff/248.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        03/tiff/249.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        03/tiff/250.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        03/tiff/251.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        03/tiff/252.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        03/tiff/253.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        03/tiff/254.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        03/tiff/255.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        03/tiff/256.tif.

Schwarz:1977:GCT

[Sch77b] Jerald Sydney Schwarz. Generic commands — A tool for partial correctness formalisms. The Computer Journal, 20(2):151–155, May 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/151.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/152.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/153.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/154.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/155.tif.

Schagen:1982:ACS

[Sch82a] I. P. Schagen. Automatic contouring from scattered data points. The Computer Journal, 25(1):7–11, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/10.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/
11.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_25/Issue_01/tiff/7.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/8.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/9.tif.

Schmidt:1982:SDS

[Sch82b] B. Schmidt. Simulation of discrete systems using GPSS-Fortran. The Computer Journal, 25(1):84–86, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/84.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_25/Issue_01/tiff/
                                        85.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_25/Issue_01/tiff/86.
                                        tif.

Schlorer:1983:ILP

[Sch83]  Jan Schlörer.  Information

loss in partitioned statistical databases. The Computer Journal, 26(3):218–223, August 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        03/tiff/218.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        03/tiff/219.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        03/tiff/220.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        03/tiff/221.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        03/tiff/222.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_26/Issue_
03/tiff/223.tif.

Schneider:1985:WDT

[Sch85] M. L. Schneider. Weighted decision tables — an alternative solution for ambiguity. The Computer Journal, 28(4):366–371, August 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/366.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/367.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/368.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/369.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/370.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/371.tif.

Schorn:1994:DGC

[Sch94] Peter Schorn. Degeneracy in geometric computation and the perturbation approach. The Computer Journal, 37(1):35–42, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_37/Issue_01/
Vol37_01.body.html#AbstractSchorn.

Scowen:1969:NAS

[Sco69] R. S. Scowen. Note on Algorithm 40. spline interpolation of degree three. The Computer Journal, 12(4):409, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/408.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/409.tif. See [Spä69a].

Scott:1970:NSU

[Sco70] M. R. Scott. Numerical solution of unstable initial value problems by invariant imbedding. The Computer Journal, 13(4):397–400, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/130397.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/397.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/398.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/399.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/400.tif.

Scott:1974:CDC

[Sco74] Jean V. Scott. Computer design of cascade logic. The Computer Journal, 17(3):253–260, August 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/253.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/254.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/255.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/256.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/257.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/258.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/259.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/260.tif.

Scoltock:1982:SLC

[Sco82] Julie Scoltock. A survey of the literature of cluster analysis. The Computer Journal, 25(1):130–134, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/130.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/131.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/132.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/133.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/134.tif.

Scraton:1964:ETE

[Scr64a] R. E. Scraton. Estimation of the truncation error in Runge–Kutta and allied processes. The Computer Journal, 7(3):246–248, October 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        03/070246.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        03/tiff/246.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        03/tiff/247.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        03/tiff/248.tif.

Scraton:1964:NSS

[Scr64b] R. E. Scraton. The numerical solution of second-order differential equations not containing the first derivative explicitly. The Computer Journal, 6(4):368–370, January 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        04/060368.sgm.abs.html;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/368.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/369.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/370.tif.

Scraton:1965:SLD

[Scr65] R. E. Scraton. The solution of linear differential equations in Chebyshev series. The Computer Journal, 8(1):57–61, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
01/080057.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
01/tiff/57.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_01/tiff/
58.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_08/Issue_01/tiff/59.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
08/Issue_01/tiff/60.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
01/tiff/61.tif.

Scraton:1970:MIC

[Scr70] R. E. Scraton. A method for improving the convergence of Chebyshev series. The Computer Journal, 13(2):202–203, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/130202.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/tiff/202.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/tiff/203.tif.

Scraton:1972:CGE

[Scr72] R. E. Scraton. Correspondence: The generalised Euler transformation. The Computer Journal, 15(3):258, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/258.tif. See [Wyn71].

Swinnerton-Dyer:1962:CPS

[SD62] H. P. F. Swinnerton-Dyer. The calculation of power spectra. The Computer Journal, 5(1):16–23, April 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        01/050016.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        01/tiff/16.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_05/Issue_01/tiff/
                                        17.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_05/Issue_01/tiff/18.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        05/Issue_01/tiff/19.tif;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
01/tiff/20.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_01/tiff/
21.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_05/Issue_01/tiff/22.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
05/Issue_01/tiff/23.tif.

Stocker:1973:SOD

[SD73] P. M. Stocker and P. A. Dearnley. Self-organising data management systems. The Computer Journal, 16(2):100–105, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/160100.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/100.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/101.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/102.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/103.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/104.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/105.tif. See corre-

spondence [PS74].

Samson:1978:STU

[SD78] W. B. Samson and R. H. Davis. Search times using hash tables for records with nonunique keys. The Computer Journal, 21(3):210–214, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        03/tiff/210.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        03/tiff/211.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        03/tiff/212.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        03/tiff/213.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        03/tiff/214.tif.

Sacks-Davis:1982:ARN

[SD82] R. Sacks-Davis. Applications of redundant number representations to decimal arithmetic. The Computer Journal, 25(4):471–477, November 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        04/tiff/471.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        04/tiff/472.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        04/tiff/473.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/474.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/475.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/476.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/477.tif.

Smith:1986:GCC

[SD86] G. W. Smith and J. B. H. Du Boulay. The generation of cryptic crossword clues. The Computer Journal, 29(3):282–284, June 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/282.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/283.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/284.tif.

Shanehchi:1982:MBA

[SE82] J. Shanehchi and D. J. Evans. A modified bisection algorithm for the determination of the eigenvalues of a symmetric quindiagonal matrix. The Computer Journal, 25(4):493–494, November 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/493.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        04/tiff/494.tif.

Seaton:1969:DCS

[Sea69]  M. J. Seaton.  Diagonali-

sation of complex symmetric matrices using a modified Jacobi method. The Computer Journal, 12(2):156–157, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        02/120156.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        02/tiff/156.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        02/tiff/157.tif.

Seaks:1971:EEE

[Sea71] Terry G. Seaks. Erratum: The evaluation of eigenvalues and eigenvectors of real symmetric matrices by simultaneous iteration. The Computer Journal, 14(4):351, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/140351.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/351.tif. See [CJ70].

Segev:1987:OTJ

[Seg87] A. Segev. Optimising twoway joins in horizontally partitioned database systems.

The Computer Journal, 30 (5):458–468, October 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/458.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/459.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/460.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/461.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/462.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/463.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/464.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/465.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/466.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/467.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/468.tif.

Seldon:1972:CPS

[Sel72]  J. L. Seldon.  Correspon-

dence: The postage stamp problem. The Computer Journal, 15(4):361, November 1972.

CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/tiff/361.tif. See [Lun69].

Selim:1988:SVV

[Sel88] S. M. Selim. Split vertices in vertex colouring and their application in developing a solution to the faculty timetable problem. The Computer Journal, 31(1):76–82, February 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        01/tiff/76.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_31/Issue_01/tiff/
                                        77.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_31/Issue_01/tiff/78.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        31/Issue_01/tiff/79.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        01/tiff/80.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_31/Issue_01/tiff/
                                        81.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_31/Issue_01/tiff/82.
                                        tif.

Sideri:1989:SDP

[SEP89] M. Sideri, S. Efremidis, and G. Papakonstantinou. Semantically driven parsing of contextfree languages. The Computer Journal, 32(1):91–93, February 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
01/tiff/91.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_01/tiff/
92.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_32/Issue_01/tiff/93.
tif.

Sernadas:1981:SSSa

[Ser81a]  Amilcar Sernadas.  SYS-

TEMATICS: its syntax and semantics as a query language. The Computer Journal, 24(1):56–61, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
01/tiff/56.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_01/tiff/
57.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_24/Issue_01/tiff/58.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
24/Issue_01/tiff/59.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
01/tiff/60.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_01/tiff/
61.tif.

Sernadas:1981:SSSb

[Ser81b] Amilcar Sernadas. SYSTEM-ATICS: Its syntax and semantics as a query language (2). The Computer Journal, 24(2):125–129, May 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        02/tiff/125.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        02/tiff/126.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        02/tiff/127.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        02/tiff/128.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        02/tiff/129.tif.

Sereno:1995:TPF

[Ser95] Matteo Sereno. Towards a product form solution for stochastic process algebras. The Computer Journal, 38(7):622–632, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_38/Issue_07/
                                        Vol38_07.body.html#AbstractSereno.

Strachey:1961:RMC

[SF61] C. Strachey and J. G. F. Francis. The reduction of a matrix to codiagonal form by eliminations. The Computer Journal, 4(2):168–176, July 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_

02/040168.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/168.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/169.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/170.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/171.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/172.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/173.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/174.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/175.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/176.tif.

Semmens:1992:ISA

[SFD92] L. T. Semmens, R. B. France, and T. W. G. Docker. Integrated structured analysis and formal specification techniques. The Computer Journal, 35(6):600–610, December 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_06/
Vol35_06.body.html#AbstractSemmens.

See [Coo92a].

Smith:1978:MFA

[SG78] D. A. Smith and P. S. Guipre. Modifications to the Forsythe-Moler algorithm for solving linear algebraic systems. The Computer Journal, 21(2):174–177, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        02/tiff/174.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        02/tiff/175.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        02/tiff/176.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        02/tiff/177.tif. See corre-

spondence [LS78, Hit78].

Sekharan:1996:STT

[SGR96] Chandran Sekharan, John G. Del Greco, and Sailesh Rathi. Space-time tradeoffs in the relative unranking of binary trees. The Computer Journal, 39(1):36– 44, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://www.oup.co.uk/jnls/
                                        list/comjnl/hdb/Volume_39/
                                        Issue_01/390036.sgm.abs.html;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_39/Issue_01/
                                        Vol39_01.body.html#AbstractSekharan.

Smith:1972:ROS

[SH72] J. L. M. Smith and T. S.

Holden. Restart of an operating system having a permanent file structure. The Computer Journal, 15(1):25–31, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/150025.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/tiff/25.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_01/tiff/
26.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_15/Issue_01/tiff/27.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
15/Issue_01/tiff/28.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/tiff/29.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_01/tiff/
30.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_15/Issue_01/tiff/31.
tif.

Storer:1997:LIC

[SH97] J. A. Storer and H. Helfgott.

Lossless image compression by block matching. The Computer Journal, 40(2/3):137–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www.oup.co.uk/computer_
journal/Volume_40/Issue_02/
Vol40_02.body.html#AbstractStorer;
 http://www3.oup.co.uk/computer_

                                        journal/Volume_40/Issue_02/
                                        Vol40_03.body.html#AbstractStorer.

Shackleton:1958:MWS

[Sha58] P. Shackleton. A model for weekly shop loading. The Computer Journal, 1(2):87–89, July 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        02/010087.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        02/tiff/87.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_01/Issue_02/tiff/
                                        88.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_01/Issue_02/tiff/89.
                                        tif.

Sharman:1965:CR

[Sha65]  F. H. Sharman.  Charac-

ter recognition. The Computer Journal, 8(2):89–94, July 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        02/080089.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        02/tiff/89.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_08/Issue_02/tiff/
                                        90.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_08/Issue_02/tiff/91.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        08/Issue_02/tiff/92.tif;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
02/tiff/93.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_02/tiff/
94.tif.

Shain:1972:ATR

[Sha72a] M. Shain. Acceptance trials for a real time system. The Computer Journal, 15 (4):298–301, November 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/150298.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/298.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/299.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/300.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/301.tif.

Shanks:1972:RTS

[Sha72b] J. A. Shanks. Romberg tables for singular integrands. The Computer Journal, 15 (4):360–361, November 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/150360.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_

                                        04/tiff/360.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/tiff/361.tif.

Shave:1981:EFB

[Sha81] M. J. R. Shave. Entities, functions, and binary relations: Steps to a conceptual schema. The Computer Journal, 24(1):42–47, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        01/tiff/42.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_24/Issue_01/tiff/
                                        43.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_24/Issue_01/tiff/44.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        24/Issue_01/tiff/45.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        01/tiff/46.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_24/Issue_01/tiff/
                                        47.tif.

Shackel:1982:BSP

[Sha82]  B. Shackel.  The BLEND

system: programme for the study of some ‘electronic journals’. The Computer Journal, 25(2):161–168, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/161.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/162.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/163.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/164.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/165.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/166.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/167.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/168.tif.

Shackel:1986:IMU

[Sha86] B. Shackel. IBM makes usability as important as functionality. The Computer Journal, 29(5):475–476, October 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/475.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/476.tif.

Sharman:1994:SRO

[Sha94] R. A. Sharman. Speech recognition in the office: how the technology supports dictation. The Computer Journal, 37(9):735–744, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_37/Issue_09/
                                        Vol37_09.body.html#AbstractSharman.

Shafarenko:1995:SDP

[Sha95] A. Shafarenko. Symmetries in data parallelism. The Computer Journal, 38(5):365–380, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_38/Issue_05/
                                        Vol38_05.body.html#AbstractShafarenko.

Stewart:1991:ADF

[SHCP91] A. Stewart, C. Holt, M. Clint, and R. Perrott. Algorithm 128: A DAP FORTRAN subroutine for thinning digitised pictures. The Computer Journal, 34(4):377–379, August 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        04/tiff/377.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        04/tiff/378.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        04/tiff/379.tif.

Shepherd:1971:ASP

[She71]  R. F. Shepherd.  Algo-

rithms supplement: Previously published algorithms. The Computer Journal, 14(4):444, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
14/Issue_04/tiff/444.tif.

Shepherd:1972:ASP

[She72a]  R. F. Shepherd.  Algo-

rithms supplement: Previously published algorithms. The Computer Journal, 15 (3):280–281, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/280.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/281.tif.

Shepherd:1972:ASS

[She72b] R. F. Shepherd. Algorithms supplement: Statement of policy. The Computer Journal, 15(4):384, November 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/384.tif.

Shepherd:1973:ASSa

[She73a] R. F. Shepherd. Algorithms supplement — statement of policy. The Computer Journal, 16(1):95, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_

journal/hdb/Volume_16/Issue_
01/tiff/95.tif.

Shepherd:1973:ASSb

[She73b] R. F. Shepherd. Algorithms supplement — statement of policy. The Computer Journal, 16(3):279–280, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/279.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/280.tif.

Shepherd:1974:ASS

[She74] R. F. Shepherd. Algorithms supplement: Statement of policy. The Computer Journal, 17(1):91–92, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
01/tiff/91.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/
92.tif.

Shepherd:1979:ASS

[She79] R. F. Shepherd. Algorithms supplement: Statement of policy. The Computer Journal, 22(1):90, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
01/tiff/90.tif.

Shebini:1985:FFE

[She85] Nader Shebini. FEDOS: a Finite Element Database of Offshore Structures. The Computer Journal, 28(5):498–500, November 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/498.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/499.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/500.tif.

Shepherd:1989:FBF

[She89] M. A. Shepherd. A fixed-size Bloom filter for searching textural documents. The Computer Journal, 32(3):212–219, June 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/212.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/213.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/214.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/215.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/216.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_

                                        03/tiff/217.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        03/tiff/218.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        03/tiff/219.tif.

Shen:1992:SAM

[She92] H. Shen. Self-adjusting mapping: A heuristic mapping algorithm for mapping parallel programs on to Transputer networks. The Computer Journal, 35(1):71–80, February 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_35/Issue_01/
                                        Vol35_01.body.html#AbstractShen.

Shen:1995:EPS

[She95] Hong Shen. Efficient parallel k-set chain range-join in hypercubes. The Computer Journal, 38(3):217–225, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_38/Issue_03/
                                        Vol38_03.body.html#AbstractShen.

Shen:1999:DKC

[She99] Alexander Shen. Discussion on Kolmogorov complexity and statistical analysis. The Computer Journal, 42(4):340–342, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_

journal/hdb/Volume_42/Issue_
04/420340.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
04/pdf/420340.pdf.

Strachey:1981:AIP

[SHH81] C. Strachey, Velma R. Huskey, and Harry D. Huskey. Anecdotes: An impossible program; Charles Babbage and Lady Lovelace. Annals of the History of Computing, 3(4):414– 415, October/December 1981. CODEN AHCOE5. ISSN 0164-

1239.  URL http://dlib.
computer.org/an/books/an1981/
pdf/a4414.pdf.

Shields:1985:CM

[Shi85] M. W. Shields. Concurrent machines. The Computer Journal, 28(5):449–465, November 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/449.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/450.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/451.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/452.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/453.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_

05/tiff/454.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/455.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/456.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/457.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/458.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/459.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/460.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/461.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/462.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/463.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/464.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
05/tiff/465.tif.

Shields:1989:ISS

[Shi89] M. W. Shields. Implicit system specification and the interface equation. The Computer Journal, 32(5):399–412, October 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

journal/hdb/Volume_32/Issue_
05/tiff/399.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/400.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/401.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/402.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/403.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/404.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/405.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/406.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/407.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/408.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/409.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/410.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/411.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/412.tif.

Shizgal:1990:AS

[Shi90]  I. Shizgal.  The Amoeba-

Prolog system. The Computer Journal, 33(6):508–517, December 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/508.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/509.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/510.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/511.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/512.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/513.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/514.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/515.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/516.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/517.tif.

Short:1981:PDT

[Sho81]  K. W. Short.  Protection

in data type abstractions using constraints on data values. The Computer Journal, 24(2):118–124, May 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/118.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/119.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/120.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/121.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/122.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/123.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/124.tif.

Shrivastava:1975:VCP

[Shr75] S. K. Shrivastava. A view of concurrent process synchronisation. The Computer Journal, 18(4):375–379, November 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/375.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/376.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/377.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/378.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_18/Issue_
                                        04/tiff/379.tif.

Sarda:1980:AAD

[SI80] N. L. Sarda and J. R. Isaac. Automatic analysis and design of information processing systems. The Computer Journal, 23(1):90–94, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        01/tiff/90.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_23/Issue_01/tiff/
                                        91.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_23/Issue_01/tiff/92.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        23/Issue_01/tiff/93.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        01/tiff/94.tif.

Sibson:1971:SOP

[Sib71]  R. Sibson.  Some observa-

tions on a paper by Lance and Williams. The Computer Journal, 14(2):156–157, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/140156.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/156.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/157.tif.

Sibson:1973:SOE

[Sib73] R. Sibson. SLINK: an optimally efficient algorithm for the single-link cluster method. The Computer Journal, 16 (1):30–34, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
01/160030.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
01/tiff/30.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_01/tiff/
31.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_16/Issue_01/tiff/32.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
16/Issue_01/tiff/33.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
01/tiff/34.tif.

Sibson:1978:LET

[Sib78] R. Sibson. Locally equiangular triangulations. The Computer Journal, 21(3):243–245, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/243.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/244.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/245.tif.

Silber:1972:CDD

[Sil72]  David Silber.  Correspon-

dence: On “Dynamic debugging in BASIC”. The Computer Journal, 15(3):258, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/258.tif. See [Bul72].

Silberschatz:1979:SIP

[Sil79] A. Silberschatz. On the safety of the I/O primitive in Concurrent PASCAL. The Computer Journal, 22(2):142–145, May 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/142.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/143.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/144.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/145.tif.

Silberschatz:1981:PDC

[Sil81] A. Silberschatz. Port directed communication. The Computer Journal, 24(1):78–82, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        01/tiff/78.tif;   http://

www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_01/tiff/
79.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_24/Issue_01/tiff/80.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
24/Issue_01/tiff/81.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
01/tiff/82.tif.

Silberschatz:1982:PQS

[Sil82] Abraham Silberschatz. Priority and queuing specification in ‘distributed processes’. The Computer Journal, 25(1):34–36, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/34.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/
35.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_25/Issue_01/tiff/36.
tif.

Simpson:1969:CFO

[Sim69]  H. R. Simpson.  A com-

pact form of one-track syntax analyser. The Computer Journal, 12(3):233–243, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/120233.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_

03/tiff/233.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/234.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/235.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/236.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/237.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/238.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/239.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/240.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/241.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/242.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/243.tif.

Sinharoy:1997:OTC

[Sin97] Balaram Sinharoy. Optimized thread creation for processor multithreading. The Computer Journal, 40(6):388–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www.oup.co.uk/computer_
journal/Volume_40/Issue_06/
Vol40_06.body.html#AbstractSinharoy;
 http://www3.oup.co.uk/computer_

journal/Volume_40/Issue_06/
Vol40_06.body.html#AbstractSinharoy.

Sipala:1969:FDA

[Sip69] P. Sipala. Formatting and display of ALGOL expressions. The Computer Journal, 12(4):365–367, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/120365.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/365.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/366.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/367.tif.

Sipala:1988:EIB

[Sip88] P. Sipala. Efficient implementation of binary trees in LISP systems. The Computer Journal, 31(4):330–335, August 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/330.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/331.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/332.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/333.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/334.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/335.tif.

Simpson:1979:PSM

[SJ79] H. R. Simpson and K. Jackson. Process synchronisation in MASCOT. The Computer Journal, 22(4):332–345, November 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/332.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/333.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/334.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/335.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/336.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/337.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/338.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/339.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/340.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_

04/tiff/341.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/342.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/343.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/344.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/345.tif.

Shacham:1976:DMS

[SK76] M. Shacham and E. Kehat. A direct method for the solution of large sparse systems of linear equations. The Computer Journal, 19(4):353–359, November 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/353.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/354.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/355.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/356.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/357.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/358.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/359.tif.

Santoro:1985:LIR

[SK85] Nicola Santoro and Ramez Khatib. Labelling and implicit routing in networks. The Computer Journal, 28(1):5–8, February 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        01/tiff/5.tif;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_28/Issue_01/tiff/
                                        6.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_28/Issue_01/tiff/7.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        28/Issue_01/tiff/8.tif.

Schwederski:1993:SIE

[SK93] Thomas Schwederski and Rudolph Kober. Special issue editorial: Parallel processing. The Computer Journal, 36(8):683, ???? 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_36/Issue_08/
                                        Vol36_08.body.html#AbstractSchwederski.

Sutton:1995:IIF

[SK95] D. R. Sutton and P. J. H. King. Incomplete information and the functional data model. The Computer Journal, 38(1):31–42, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_38/Issue_01/
                                        Vol38_01.body.html#AbstractSutton.

Skillicorn:1991:PCP

[Ski91] D. B. Skillicorn. Practical concurrent programming for parallel machines. The Computer Journal, 34(4):302–310, August 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/302.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/303.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/304.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/305.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/306.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/307.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/308.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/309.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/310.tif.

Skillicorn:1996:UDH

[Ski96] D. B. Skillicorn. Using distributed hypermedia for collaborative learning in universities. The Computer Journal, 39(6):471–482, ???? 1996. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://www.oup.co.uk/jnls/
                                        list/comjnl/hdb/Volume_39/
                                        Issue_06/390471.sgm.abs.html;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_39/Issue_06/
                                        Vol39_06.body.html#AbstractSkillicorn.

Sklenar:1981:AS

[Skl81]  I. Sklenář.  Affine slices.

The Computer Journal, 24 (4):374–377, November 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        04/tiff/374.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        04/tiff/375.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        04/tiff/376.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        04/tiff/377.tif.

Styne:1985:PSR

[SKW85] B. A. Styne, T. R. King, and N. E. Wiseman. Pad structures for the Rainbow Workstation. The Computer Journal, 28(1):68–72, February 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        01/tiff/68.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_28/Issue_01/tiff/
                                        69.tif;   http://www3.oup.

co.uk/computer_journal/hdb/
Volume_28/Issue_01/tiff/70.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
28/Issue_01/tiff/71.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
01/tiff/72.tif.

Spyropoulos:1973:DAN

[SKY73] K. Spyropoulos, E. Kiountouzis, and A. Young. Discrete approximation in the L_{1} norm. The Computer Journal, 16(2):180–186, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/160180.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/180.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/181.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/182.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/183.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/184.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/185.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/186.tif.

Stentiford:1973:EAC

[SL73] F. W. M. Stentiford and D. W. Lewin. An evolutionary approach to the concept of randomness. The Computer Journal, 16(2):148–151, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/160148.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/148.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/149.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/150.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/151.tif. See corre-

spondence [Pit74].

Swift:1978:CCM

[SL78] A. Swift and G. R. Lindfield. Comparison of a continuation method with Brent’s method for the numerical solution of a single nonlinear equation. The Computer Journal, 21(4):359–362, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        04/tiff/359.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        04/tiff/360.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/361.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/362.tif.

Slater:1960:PIL

[Sla60] L. D. Slater. Problems of the introduction of large scale data processing into the Royal Army Pay Corps. The Computer Journal, 3(3):120–123, October 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/030120.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/120.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/121.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/122.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/123.tif.

Slater:1962:RA

[Sla62] L. J. Slater. Regression analysis. The Computer Journal, 4(4):287–291, January 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/040287.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_

                                        04/tiff/287.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        04/tiff/288.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        04/tiff/289.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        04/tiff/290.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        04/tiff/291.tif.

Slater:1964:DPP

[Sla64a]  L. J. Slater.  A dynamic

programming process. The Computer Journal, 7(1):36–39, April 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        01/070036.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        01/tiff/36.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_07/Issue_01/tiff/
                                        37.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_07/Issue_01/tiff/38.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        07/Issue_01/tiff/39.tif.

Slater:1964:SEP

[Sla64b] L. J. Slater. Some experiences in price mapping. The Computer Journal, 6(4):348–351, January 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_

journal/hdb/Volume_06/Issue_
04/060348.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/348.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/349.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/350.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/351.tif.

Slater:1966:CSE

[Sla66] L. J. Slater. Computing the state of the economy. The Computer Journal, 9(1):11–15, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/090011.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/11.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_01/tiff/
12.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_09/Issue_01/tiff/13.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
09/Issue_01/tiff/14.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/15.tif.

Slater:1971:BRM

[Sla71] L. J. Slater. Book review:

Mathematical Model Building in Economics and Industry (Second Series), by M. G. Kendall (editor), 1970. The Computer Journal, 14(4):436, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/436.tif.

Smith:1983:IUI

[SLH83] A. M. R. Smith, L. S. Lee, and D. J. Hand. Interactive userfriendly interfaces to statistical packages. The Computer Journal, 26(3):199–204, August 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        03/tiff/199.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        03/tiff/200.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        03/tiff/201.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        03/tiff/202.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        03/tiff/203.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        03/tiff/204.tif.

Sabin:1971:NAA

[SM71] M. A. Sabin and D. J. Mc-Conalogue. Note on Algorithm 67: An axis invariant procedure to provide slopes at points by local three-point parametric curve fitting. The Computer Journal, 14(4):451, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/451.tif. See [McC71b].

Stojmenovic:1988:ASA

[SM88] I. Stojmenovic and M. Miyakawa.

Applications of a subsetgenerating algorithm to base enumeration, knapsack and minimal covering problems. The Computer Journal, 31 (1):65–70, February 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
01/tiff/65.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_01/tiff/
66.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_31/Issue_01/tiff/67.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
31/Issue_01/tiff/68.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
01/tiff/69.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_01/tiff/
70.tif.

Stephen:1994:WST

[SM94] Graham A. Stephen and Paul Mather. What is SP? (theory of computing). The Computer Journal, 37(9):745–752, ???? 1994. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_37/Issue_09/
                                        Vol37_09.body.html#AbstractStephen.

Singhai:1997:PLF

[SM97] Sharad K. Singhai and Kathryn S. McKinley. A parametrized loop fusion algorithm for improving parallelism and cache locality. The Computer Journal, 40(6):340–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www.oup.co.uk/computer_
                                        journal/Volume_40/Issue_06/
                                        Vol40_06.body.html#AbstractSinghai;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_40/Issue_06/
                                        Vol40_06.body.html#AbstractSinghai.

Small:1988:IES

[Sma88] Carol Small. The implementation of the Exegesis system. The Computer Journal, 31(2):125–132, April 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        02/tiff/125.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        02/tiff/126.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        02/tiff/127.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        02/tiff/128.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/129.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/130.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/131.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/132.tif.

Smeaton:1992:PAN

[Sme92] Alan F. Smeaton. Progress in the application of natural language processing to information retrieval tasks. The Computer Journal, 35(3):268–278, June 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_35/Issue_03/
Vol35_03.body.html#AbstractSmeaton.

Smith:1963:VDT

[Smi63] K. L. Smith. The views of the data transmission committee. The Computer Journal, 6(3):222–227, October 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/060222.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/222.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/223.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_06/Issue_
                                        03/tiff/224.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        03/tiff/225.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        03/tiff/226.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        03/tiff/227.tif.

Smith:1964:CDU

[Smi64] C. R. Smith. Computer development from a user’s point of view. The Computer Journal, 7(2):89–92, July 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        02/070089.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        02/tiff/89.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_07/Issue_02/tiff/
                                        90.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_07/Issue_02/tiff/91.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        07/Issue_02/tiff/92.tif.

Smith:1966:LEC

[Smi66a] A. C. Smith. A lower estimate of the cumulative truncation error in Milne’s method. The Computer Journal, 8 (4):395–397, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/080395.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/395.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/396.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/397.tif.

Smith:1966:SAP

[Smi66b] N. Smith. Systems analysis problems encountered during a large computer application. The Computer Journal, 9(3):230–237, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/090230.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/230.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/231.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/232.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/233.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/234.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/235.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_09/Issue_
                                        03/tiff/236.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        03/tiff/237.tif.

Smith:1968:SCE

[Smi68] B. Smith. SPECOL — a computer enquiry language for the non-programmer. The Computer Journal, 11(2):121– 127, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        11/Issue_02/110121.sgm.abs.
                                        html;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        11/Issue_02/tiff/121.tif;
                                         http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        02/tiff/122.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        02/tiff/123.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        02/tiff/124.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        02/tiff/125.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        02/tiff/126.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        02/tiff/127.tif.

Smillie:1970:ISM

[Smi70a] K. W. Smillie. Algorithm 49: Indexing subarrays in multidimensional arrays. The Computer Journal, 13(2):208, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/208.tif.

Smith:1970:DSR

[Smi70b] B. Smith. Developments in SPECOL — a retrieval language for the non-programmer. The Computer Journal, 13 (1):10–13, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/130010.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/tiff/10.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_01/tiff/
11.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/12.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
13/Issue_01/tiff/13.tif.

Smith:1971:DEH

[Smi71] L. B. Smith. Drawing ellipses, hyperbolas or parabolas with a fixed number of points and maximum inscribed area. The Computer Journal, 14(1):81–86, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/140081.sgm.abs.html;

                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        01/tiff/81.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_14/Issue_01/tiff/
                                        82.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_14/Issue_01/tiff/83.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        14/Issue_01/tiff/84.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        01/tiff/85.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_14/Issue_01/tiff/

86.tif. See [McC70b, Ecc71,

Sab72, Coo72b].

Smith:1972:PVP

[Smi72] J. Meredith Smith. Proof and validation of program correctness. The Computer Journal, 15(2):130–131, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        02/150130.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        02/tiff/130.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        02/tiff/131.tif. See note

[LB76].

Smith:1983:XCC

[Smi83] P. D. Smith. XENO: Computerassisted compilation of crossword puzzles. The Computer Journal, 26(4):296–302, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/296.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/297.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/298.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/299.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/300.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/301.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/302.tif.

Smith:1986:ISP

[Smi86] Joan M. Smith. The implications of SGML for the preparation of scientific publications. The Computer Journal, 29(3):193–200, June 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/193.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/194.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/195.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_29/Issue_
                                        03/tiff/196.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        03/tiff/197.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        03/tiff/198.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        03/tiff/199.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        03/tiff/200.tif.

Sanna:1998:FAM

[SMR98] Andrea Sanna, Paolo Montuschi, and Massimo Rossi. A flexible algorithm for multiprocessor ray tracing. The Computer Journal, 41(7):503–516, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_41/Issue_07/
                                        Vol41_07.body.html#AbstractSanna.

Smyth:1991:MMS

[Smy91] W. F. Smyth. Mu-balancing Mway search trees. The Computer Journal, 34(5):406–414, October 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        05/tiff/406.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        05/tiff/407.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        05/tiff/408.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/409.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/410.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/411.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/412.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/413.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/414.tif.

Sun:1992:QOD

[SMY92] W. Sun, W. Meng, and C. Yu. Query optimisation in distributed object-oriented database systems. The Computer Journal, 35(2):98–107, April 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_35/Issue_02/
Vol35_02.body.html#AbstractSun.

Stoyenko:1995:LCR

[SMY95] Alexander D. Stoyenko, Thomas J.

Marlowe, and Mohamed F. Younis. A language for complex real-time systems. The Computer Journal, 38(4):319–338, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_38/Issue_04/
Vol38_04.body.html#AbstractStoyenko.

Slagle:1975:ATT

[SN75] J. R. Slagle and L. M. Norton. Automated theorem-proving for the theories of partial and total ordering. The Computer Journal, 18(1):49–54, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        01/tiff/49.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_18/Issue_01/tiff/
                                        50.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_18/Issue_01/tiff/51.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        18/Issue_01/tiff/52.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        01/tiff/53.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_18/Issue_01/tiff/
                                        54.tif.

Sneath:1966:MCS

[Sne66] P. H. A. Sneath. A method for curve seeking from scattered points. The Computer Journal, 8(4):383–391, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        04/080383.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        04/tiff/383.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_

04/tiff/384.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/385.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/386.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/387.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/388.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/389.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/390.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/391.tif.

Snell:1970:SCS

[Sne70]  J. Snell.  The solution in

Chebyshev series of systems of linear differential equations with general boundary conditions. The Computer Journal, 13(1):103–106, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/130103.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/tiff/103.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/tiff/104.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_

                                        01/tiff/105.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        01/tiff/106.tif.

Snell:1980:CEP

[Sne80] J. Snell. Correspondence: On “An efficient predictor-corrector algorithm”. The Computer Journal, 23(4):382, November 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        04/tiff/382.tif. See [Wes80a,

Ano80b].

Sakaki:1976:AFL

[SNH76] T. Sakaki, K. Nakashima, and Y. Hattori. Algorithms for finding in the lump both bounds of the chromatic number of a graph. The Computer Journal, 19(4):329–332, November 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        04/tiff/329.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        04/tiff/330.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        04/tiff/331.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        04/tiff/332.tif.

Solomon:1960:CPN

[Sol60]  E. W. Solomon.  A com-

prehensive program for network problems. The Computer Journal, 3(2):89–97, July 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
02/030089.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
02/tiff/89.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_02/tiff/
90.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_03/Issue_02/tiff/91.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
03/Issue_02/tiff/92.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
02/tiff/93.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_02/tiff/
94.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_03/Issue_02/tiff/95.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
03/Issue_02/tiff/96.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
02/tiff/97.tif.

Solomon:1968:CBD

[Sol68]  Eric W. Solomon.  Corre-

spondence: Basic data input routines. The Computer Journal, 11(3):352, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

                                        journal/hdb/Volume_11/Issue_
                                        03/tiff/352.tif.

Solomonoff:1999:TKP

[Sol99] Ray Solomonoff. Two kinds of probabilistic induction. The Computer Journal, 42(4):256– 259, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        42/Issue_04/420256.sgm.abs.
                                        html;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        42/Issue_04/pdf/420256.pdf.

Southwell:1976:FDN

[Sou76] W. H. Southwell. Fitting data to nonlinear functions with uncertainties in all measurement variables. The Computer Journal, 19(1):69–73, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        01/tiff/69.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_19/Issue_01/tiff/
                                        70.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_19/Issue_01/tiff/71.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        19/Issue_01/tiff/72.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        01/tiff/73.tif.

Soule:1977:NNT

[Sou77] S. Soule. A note on the nonrecursive traversal of binary trees. The Computer Journal, 20(4):350–352, November 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/350.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/351.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/352.tif.

Sowa:1987:MSS

[Sow87]  M. Sowa.  A method for

speeding up serial processing in dataflow computers by means of a program counter. The Computer Journal, 30 (4):289–294, August 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/289.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/290.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/291.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/292.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/293.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/294.tif.

Skordalakis:1982:CSV

[SP82] E. Skordalakis and G. Papakonstantinou. A control structure for a variable number of nested loops. The Computer Journal, 25(1):48–51, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/48.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_25/Issue_01/tiff/
                                        49.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_25/Issue_01/tiff/50.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        25/Issue_01/tiff/51.tif.

Son:1991:ADP

[SP91] N. T. Son and Y. Paker. Adaptive deadlock-free packet routing in transputer-based multiprocessor interconnection networks. The Computer Journal, 34(6):493–502, December 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        06/tiff/493.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        06/tiff/494.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        06/tiff/495.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        06/tiff/496.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/497.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/498.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/499.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/500.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/501.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/502.tif.

Spath:1968:AFD

[Spä68]  H. Späth.  Algorithm 33:

Fitting data to an exponential with a straight line as background. The Computer Journal, 11(1):114–115, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/114.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/115.tif.

Spath:1969:ASI

[Spä69a]  H. Späth.  Algorithm 40:

Spline interpolation of degree three. The Computer Journal, 12(2):198–199, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_12/Issue_
                                        02/tiff/198.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        02/tiff/199.tif. See note

[Sco69].

Spath:1969:AIC

[Spä69b] H. Späth. Algorithm 42: Interpolation by certain quintic splines. The Computer Journal, 12(3):292–293, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        03/tiff/292.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        03/tiff/293.tif. See note

[Sal70b].

Shackel:1983:BLP

[SPMD83] B. Shackel, D. J. Pullinger, T. I. Maude, and W. P. Dodd. The BLEND-LINC project on ‘electronic journals’ after two years. The Computer Journal, 26(3):247–254, August 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        03/tiff/247.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        03/tiff/248.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        03/tiff/249.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_26/Issue_
03/tiff/250.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/251.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/252.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/253.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/254.tif.

Spoletini:1985:AGP

[Spo85] Enrico Spoletini. Algorithm 122: Generation of permutations in direct lexicographic ordering by arithmetic methods. The Computer Journal, 28(4):444–445, November 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/444.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/445.tif.

Sprugnoli:1992:PBT

[Spr92] R. Sprugnoli. Properties of binary trees related to position. The Computer Journal, 35(4):395–404, August 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_35/Issue_04/
Vol35_04.body.html#AbstractSprugnoli.

Spurr:1988:DDA

[Spu88] Kathleen Spurr. A data dictionary approach to teaching information systems analysis. The Computer Journal, 31(2):182–189, April 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/182.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/183.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/184.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/185.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/186.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/187.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/188.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/189.tif.

Smyth:1974:SSH

[SR74] W. F. Smyth and E. Radaceanu. A storage scheme for hierarchic structures. The Computer Journal, 17(2):152–156, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

journal/hdb/Volume_17/Issue_
02/tiff/152.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/153.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/154.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/155.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/156.tif.

Siskind:1976:STD

[SR76] V. Siskind and J. Rosenhead.

Seek times for disc file processing: Some results from probability theory. The Computer Journal, 19(4):301–305, November 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/301.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/302.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/303.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/304.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/305.tif.

Srinivasan:1991:AOT

[Sri91] B. Srinivasan. An adaptive overflow technique to defer splitting in B-trees. The Computer Journal, 34(5):397–405, October 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/397.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/398.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/399.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/400.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/401.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/402.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/403.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/404.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/405.tif.

Seshadri:1994:OAC

[SRS94] Sridhar Seshadri, Doron Rotem, and Arie Segev. Optimal arrangements of cartridges in carousel type mass storage systems. The Computer Journal, 37(10):873–887, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_

journal/Volume_37/Issue_10/
Vol37_10.body.html#AbstractSeshadri.

Scraton:1963:NNS

[SS63a] R. E. Scraton and J. W.

Searle. Note on the numerical solution of linear differential equations with constant coefficients. The Computer Journal, 6(2):206–207, July 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/060206.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/206.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/207.tif.

Stroud:1963:MPF

[SS63b] A. H. Stroud and D. Secrest.

A multiple-precision floatingpoint interpretive program for the Control Data 1604. The Computer Journal, 6(1):62–66, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/060062.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/tiff/62.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/
63.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_06/Issue_01/tiff/64.
tif;   http://www3.oup.co.

                                        uk/computer_journal/hdb/Volume_
                                        06/Issue_01/tiff/65.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        01/tiff/66.tif.

Schlechtendahl:1972:CRF

[SS72a] E. G. Schlechtendahl and U. Schumann. Correspondence: Remarks on “Faults in functions, in ALGOL and FORTRAN”. The Computer Journal, 15(2):147, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        02/150147.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        02/tiff/147.tif. See [Hil71].

Smith:1972:NSD

[SS72b] I. R. Smith and L. A. Snider. Numerical solution of the differential equations of a synchronous generator. The Computer Journal, 15 (4):350–355, November 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/150350.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/tiff/350.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/tiff/351.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_

04/tiff/352.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/353.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/354.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/355.tif.

Stoy:1972:OEOa

[SS72c] J. E. Stoy and C. Strachey.

OS6 — an experimental operating system for a small computer. Part 1: general principles and structure. The Computer Journal, 15(2):117–124, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/150117.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/117.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/118.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/119.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/120.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/121.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/122.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_

                                        02/tiff/123.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        02/tiff/124.tif.

Stoy:1972:OEOb

[SS72d] J. E. Stoy and C. Strachey. OS6 — an experimental operating system for a small computer. Part 2: input/output and filing system. The Computer Journal, 15(3):195–203, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/150195.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/195.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/196.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/197.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/198.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/199.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/200.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/201.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/202.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_15/Issue_
03/tiff/203.tif.

Smillie:1975:CDT

[SS75] K. W. Smillie and M. J. R.

Shave. Converting decision tables to computer programs. The Computer Journal, 18(2):108–111, May 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/108.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/109.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/110.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/111.tif.

Schmidt:1980:TCE

[SS80] G. Schmidt and T. Ströhlein.

Timetable construction — an annotated bibliography. The Computer Journal, 23 (4):307–316, November 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/307.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/308.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/309.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_23/Issue_
04/tiff/310.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/311.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/312.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/313.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/314.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/315.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/316.tif.

Smith:1981:PCC

[SS81] P. D. Smith and S. Y. Steen. Prototype crossword compiler. The Computer Journal, 24(2):107–111, May 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/107.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/108.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/109.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/110.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/111.tif.

Sondergaard:1992:NDF

[SS92] Harald Søndergaard and Peter Sestoft. Non-determinism in functional languages. The Computer Journal, 35(5):514– 523, October 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/Volume_
35/Issue_05/Vol35_05.body.
html#AbstractSondergaard.

Accepted — Melbourne Technical Report 88/18.

Saiedian:1996:EAC

[SS96a] Hossein Saiedian and Thomas Spencer. An efficient algorithm to compute the candidate keys of a relational database schema. The Computer Journal, 39(2):124–132, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www.oup.co.uk/jnls/
list/comjnl/hdb/Volume_39/
Issue_02/390124.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_02/
Vol39_02.body.html#AbstractSaiedian.

Sekerinski:1996:TPC

[SS96b] Emil Sekerinski and Kaisa Sere. A theory of prioritizing composition. The Computer Journal, 39(8):701–??, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www.oup.co.uk/jnls/
list/comjnl/hdb/Volume_39/
Issue_08/390701.sgm.abs.html;

 http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_08/
Vol39_08.body.html#AbstractSekerinski.

Schipler:1989:EIR

[SSDS89] A. Schipler, R. Simon, P. Desarzens, and J.-A. Sengstag. Efficient implementation of rendezvous. The Computer Journal, 32(3):267–272, June 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/267.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/268.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/269.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/270.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/271.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/272.tif.

Sibson:1981:SQE

[ST81] Robin Sibson and Graeme D. Thomson. Seamed quadratic element for contouring. The Computer Journal, 24(4):378– 382, November 1981. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
24/Issue_04/tiff/378.tif;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/379.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/380.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/381.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/382.tif.

Symons:1982:SPA

[ST82] C. R. Symons and P. Tijsma. A systematic and practical approach to the definition of data. The Computer Journal, 25(4):410–422, November 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/410.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/411.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/412.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/413.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/414.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/415.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/416.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_25/Issue_
04/tiff/417.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/418.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/419.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/420.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/421.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/422.tif.

Sommerville:1989:ASS

[ST89] I. Sommerville and R. Thomson. An approach to the support of software evolution. The Computer Journal, 32(5):386–398, October 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/386.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/387.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/388.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/389.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/390.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_

05/tiff/391.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/392.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/393.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/394.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/395.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/396.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/397.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/398.tif.

Stark:1971:LA

[Sta71] R. H. Stark. A language for algorithms. The Computer Journal, 14(1):40–44, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/140040.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/tiff/40.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_01/tiff/
41.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_14/Issue_01/tiff/42.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
14/Issue_01/tiff/43.tif;

                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        01/tiff/44.tif.

Stacey:1974:FIC

[Sta74] G. M. Stacey. A Fortran interface to the CODASYL Database Task Group specifications. The Computer Journal, 17(2):124–129, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        02/tiff/124.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        02/tiff/125.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        02/tiff/126.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        02/tiff/127.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        02/tiff/128.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        02/tiff/129.tif.

Stamper:1977:LPS

[Sta77] R. K. Stamper. The LEGOL 1 prototype system and language. The Computer Journal, 20(2):102–108, May 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        02/tiff/102.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_20/Issue_
02/tiff/103.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/104.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/105.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/106.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/107.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/108.tif.

Stamper:1985:TTI

[Sta85] R. K. Stamper. Towards a theory of information. information: Mystical fluid or a subject for scientific enquiry? The Computer Journal, 28(3):195–199, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/195.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/196.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/197.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/198.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/199.tif.

Stavridou:1994:FMV

[Sta94] Victoria Stavridou. Formal methods and VLSI engineering practice. The Computer Journal, 37(2):96–113, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_37/Issue_02/
                                        Vol37_02.body.html#AbstractStavridou.

Stern:1967:OQF

[Ste67] M. D. Stern. Optimal quadrature formulae. The Computer Journal, 9(4):396–403, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        04/090396.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        04/tiff/396.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        04/tiff/397.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        04/tiff/398.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        04/tiff/399.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        04/tiff/400.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        04/tiff/401.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        04/tiff/402.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
04/tiff/403.tif.

Stephens:1974:ILC

[Ste74] P. D. Stephens. The IMP language and compiler. The Computer Journal, 17(3):216– 223, August 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
17/Issue_03/tiff/216.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/217.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/218.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/219.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/220.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/221.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/222.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/223.tif.

Stecher:1977:PIS

[Ste77a] P. Stecher. Proposal for an interface system between the business and data processing systems. The Computer Journal, 20(3):194–201, August 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/194.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/195.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/196.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/197.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/198.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/199.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/200.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/201.tif.

Steele:1977:CNA

[Ste77b]  K. A. Steele.  CPM net-

work analysis with a storage tube terminal. The Computer Journal, 20(4):330–334, November 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/330.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/331.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/332.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_20/Issue_
04/tiff/333.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/334.tif.

Stewart:1986:QSS

[Ste86]  I. P. Stewart.  Quadtrees:

Storage and scan conversion. The Computer Journal, 29(1):60–75, February 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
01/tiff/60.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_01/tiff/
61.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_29/Issue_01/tiff/62.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
29/Issue_01/tiff/63.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
01/tiff/64.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_01/tiff/
65.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_29/Issue_01/tiff/66.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
29/Issue_01/tiff/67.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
01/tiff/68.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_01/tiff/
69.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_29/Issue_01/tiff/70.

                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        29/Issue_01/tiff/71.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        01/tiff/72.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_29/Issue_01/tiff/
                                        73.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_29/Issue_01/tiff/74.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        29/Issue_01/tiff/75.tif.

Storey:1968:CFB

[Sto68] S. H. Storey. The convergence of Fourier–Bessel expansions. The Computer Journal, 10(4):402–405, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/100402.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/tiff/402.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/tiff/403.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/tiff/404.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/tiff/405.tif.

Storey:1972:OFB

[Sto72] S. H. Storey. Optimal fixed batch sizes for real-time inquiry systems. The Computer Journal, 15(2):132–133, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/150132.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/132.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/133.tif.

Stojak:1973:HDC

[Sto73] P. F. Stojak. Hybrid and digital computation results in multicomponent distillation simulation. The Computer Journal, 16(4):368–374, November 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/160368.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/368.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/369.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/370.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/371.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/372.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/373.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_16/Issue_
                                        04/tiff/374.tif.

Stone:1986:CGP

[Sto86] R. G. Stone. On the choice of grammar and parser for the compact analytical encoding of programs. The Computer Journal, 29(4):307–314, August 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        04/tiff/307.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        04/tiff/308.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        04/tiff/309.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        04/tiff/310.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        04/tiff/311.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        04/tiff/312.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        04/tiff/313.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        04/tiff/314.tif.

Strachey:1959:TSR

[Str59]  C. Strachey.  On taking

the square root of a complex number. The Computer Journal, 2(2):89, July 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
02/020089.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
02/tiff/89.tif.

Strachey:1960:TCT

[Str60]  C. Strachey.  Two contri-

butions to the techniques of queuing problems. The Computer Journal, 3(2):114–116, July 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
02/030114.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
02/tiff/114.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
02/tiff/115.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
02/tiff/116.tif.

Stringer:1961:PCR

[Str61] J. B. Stringer. The place of character recognition, data transmission and document handling in A.D.P. systems. The Computer Journal, 4(2): 161–167, July 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
04/Issue_02/040161.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
04/Issue_02/tiff/161.tif;

 http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/162.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/163.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/164.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/165.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/166.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/167.tif.

Strachey:1965:GPM

[Str65a] C. Strachey. A general purpose macrogenerator. The Computer Journal, 8(3):225–241, October 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/080225.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/225.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/226.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/227.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/228.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/229.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/230.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/231.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/232.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/233.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/234.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/235.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/236.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/237.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/238.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/239.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/240.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/241.tif.

Strachey:1965:LEI

[Str65b] C. Strachey. Letter to the Editor: An impossible program. The Computer Journal, 7(4):313, January 1965. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Stross:1972:ODD

[Str72] C. O. M. Stross. Operation of a disc data base. The Computer Journal, 15(4):290–297, November 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/150290.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/tiff/290.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/tiff/291.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/tiff/292.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/tiff/293.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/tiff/294.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/tiff/295.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/tiff/296.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        04/tiff/297.tif.

Strawson:1974:CH

[Str74]  Joy Strawson.  Correspon-

dence: On “Himmelbett”. The Computer Journal, 17(2):190, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_

journal/hdb/Volume_17/Issue_
02/tiff/190.tif. See [Bel73,

Fre74].

Sturman:1971:FPM

[Stu71] J. N. Sturman. The functional partitioning of memoryits application to interactive computing. The Computer Journal, 14(4):352–360, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/140352.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/352.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/353.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/354.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/355.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/356.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/357.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/358.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/359.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/360.tif.

Stubbs:1984:ERS

[Stu84] Michael Stubbs. An examination of the resolution of structure clashes by structure inversion. The Computer Journal, 27(4):354–361, November 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        04/tiff/354.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        04/tiff/355.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        04/tiff/356.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        04/tiff/357.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        04/tiff/358.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        04/tiff/359.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        04/tiff/360.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        04/tiff/361.tif.

Subieta:1985:SMD

[Sub85] Kazimierz Subieta. A simple method of data correction. The Computer Journal, 28(4):372–374, August 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_

journal/hdb/Volume_28/Issue_
04/tiff/372.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/373.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/374.tif.

Suffern:1990:QAC

[Suf90] K. G. Suffern. Quadtree algorithms for contouring functions of two variables. The Computer Journal, 33(5):402–407, October 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/402.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/403.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/404.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/405.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/406.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
05/tiff/407.tif.

Suraweera:1986:UDC

[Sur86] F. Suraweera. Use of doubly chained tree structures in file organisation for optimal searching. The Computer Journal, 29(1):52–59, February 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
01/tiff/52.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_01/tiff/
53.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_29/Issue_01/tiff/54.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
29/Issue_01/tiff/55.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
01/tiff/56.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_01/tiff/
57.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_29/Issue_01/tiff/58.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
29/Issue_01/tiff/59.tif.

Sutton:1960:FYE

[Sut60] R. L. Sutton. The first year’s experience with a large computer in a life assurance office. The Computer Journal, 3(1):2–9, April 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
01/030002.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
01/tiff/2.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_03/Issue_01/tiff/
3.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_03/Issue_01/tiff/4.
tif;   http://www3.oup.co.

uk/computer_journal/hdb/Volume_
03/Issue_01/tiff/5.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
01/tiff/6.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_03/Issue_01/tiff/
7.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_03/Issue_01/tiff/8.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
03/Issue_01/tiff/9.tif.

Sutcliffe:1976:ADC

[Sut76a]  D. C. Sutcliffe.  An algo-

rithm for drawing the curve F (X, Y) = 0. The Computer Journal, 19(3):246–249, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/246.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/247.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/248.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/249.tif.

Sutcliffe:1976:RCA

[Sut76b] D. C. Sutcliffe. A remark on a contouring algorithm. The Computer Journal, 19 (4):333–335, November 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_

journal/hdb/Volume_19/Issue_
04/tiff/333.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/334.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/335.tif.

Smeaton:1983:REQ

[Sv83] A. F. Smeaton and C. J. van Rijsbergen. The retrieval effects of query expansion on a feedback document retrieval system. The Computer Journal, 26(3):239–246, August 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/239.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/240.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/241.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/242.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/243.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/244.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/245.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/246.tif.

Shepherd:1968:CAA

[SW68] M. J. Shepherd and A. J.

Willmott. Cluster analysis on the Atlas computer. The Computer Journal, 11(1):57–62, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/110057.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/57.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_01/tiff/
58.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_11/Issue_01/tiff/59.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/60.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/61.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_01/tiff/
62.tif.

Szwarcfiter:1978:SPT

[SW78] J. L. Szwarcfiter and L. B.

Wilson. Some properties of ternary trees. The Computer Journal, 21(1):66–72, February 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
01/tiff/66.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_01/tiff/
67.tif;   http://www3.oup.

                                        co.uk/computer_journal/hdb/
                                        Volume_21/Issue_01/tiff/68.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        21/Issue_01/tiff/69.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        01/tiff/70.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_21/Issue_01/tiff/
                                        71.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_21/Issue_01/tiff/72.
                                        tif.

Spengler:1979:ADR

[SW79] C. B. Spengler and J. B. Westwood. Assessing data retrieval capabilities: a case study. The Computer Journal, 22(1):17–21, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        01/tiff/17.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_22/Issue_01/tiff/
                                        18.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_22/Issue_01/tiff/19.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        22/Issue_01/tiff/20.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        01/tiff/21.tif.

Slape:1983:CFA

[SW83] John K. Slape and Peter J. L. Wallis. Conversion of Fortran to Ada using an intermediate tree representation. The Computer Journal, 26(4):344–353, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/344.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/345.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/346.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/347.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/348.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/349.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/350.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/351.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/352.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
04/tiff/353.tif.

Six:1984:SRA

[SW84] H. W. Six and L. Wegner.

Sorting a random access file in situ. The Computer Journal, 27(3):270–275, August 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

                                        journal/hdb/Volume_27/Issue_
                                        03/tiff/270.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        03/tiff/271.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        03/tiff/272.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        03/tiff/273.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        03/tiff/274.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        03/tiff/275.tif.

Slape:1991:AMS

[SW91a] J. K. Slape and P. J. L. Wallis. Algorithm 127: A modification of Sale’s algorithm to accommodate FOR-TRAN 77. The Computer Journal, 34(4):373–376, August 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        04/tiff/373.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        04/tiff/374.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        04/tiff/375.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        04/tiff/376.tif.

Sutcliffe:1991:IHC

[SW91b] A. G. Sutcliffe and I. Wang. Integrating human computer interaction with Jackson system development. The Computer Journal, 34(2):132–142, April 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/132.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/133.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/134.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/135.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/136.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/137.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/138.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/139.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/140.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/141.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/142.tif.

Sjoberg:1997:SCL

[SWA97] Dag I. K. Sjöberg, Ray Welland, and Malcolm P. Atkinson. Software constraints for large application systems. The Computer Journal, 40(10):598–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_40/Issue_10/
                                        Vol40_10.body.html#AbstractSjoberg.

Sommerville:1987:DSD

[SWB87] I. Sommerville, R. Welland, and S. Beer. Describing software design methodologies. The Computer Journal, 30(2):128–133, April 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        02/tiff/128.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        02/tiff/129.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        02/tiff/130.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        02/tiff/131.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        02/tiff/132.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        02/tiff/133.tif.

Schneider:1993:SLS

[SWZ93] M. Schneider, U. Wever, and Q. Zheng. Solving large and sparse linear equations in analog circuit simulation on a cluster of workstations. The Computer Journal, 36(8):685–689, ???? 1993. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_08/
Vol36_08.body.html#AbstractSchneider.

Smith:1972:GRE

[SY72] L. W. Smith and S. S. Yau.

Generation of regular expressions for automata by the integral of regular expressions. The Computer Journal, 15(3):222–228, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/150222.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/222.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/223.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/224.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/225.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/226.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/227.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/228.tif.

Symons:1971:CAF

[Sym71] C. R. Symons. A cost accounting formula for multiprogramming computers. The Computer Journal, 14(1):13–20, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        01/140013.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        01/tiff/13.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_14/Issue_01/tiff/
                                        14.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_14/Issue_01/tiff/15.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        14/Issue_01/tiff/16.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        01/tiff/17.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_14/Issue_01/tiff/
                                        18.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_14/Issue_01/tiff/19.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        14/Issue_01/tiff/20.tif.

Symvonis:1995:OSM

[Sym95] Antonios Symvonis. Optimal stable merging. The Computer Journal, 38(8):681–690, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_38/Issue_08/
                                        Vol38_08.body.html#AbstractSymvonis.

Shaftel:1974:ACR

[SZ74] T. L. Shaftel and R. W.

Zmud. Allocation of computer resources through flexible pricing. The Computer Journal, 17(4):306–312, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/306.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/307.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/308.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/309.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/310.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/311.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/312.tif.

Szanser:1973:BTE

[Sza73]  A. J. Szanser.  Bracketing

technique in elastic matching. The Computer Journal, 16(2):132–134, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/160132.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/132.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/133.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/134.tif.

Sousa:1994:OPH

[SZNM94] Pedro Sousa, Andre Zuquete, Nuno Neves, and Jose Alves Marques. Orthogonal persistence in a heterogeneous distributed object-oriented environment. The Computer Journal, 37(6):531–541, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_37/Issue_06/
                                        Vol37_06.body.html#AbstractSousa.

Thimbleby:1998:FMT

[TAC98] Harold Thimbleby, Stuart Anderson, and Paul Cairns. A framework for modelling Trojans and computer virus infection. The Computer Journal, 41(7):445–458, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_41/Issue_07/
                                        Vol41_07.body.html#AbstractThimbleby.

See comments [Mäk01, TAC01,

Coh01].

Thimbleby:2001:RCS

[TAC01] H. Thimbleby, S. O. Anderson, and P. A. Cairns. Reply to “Comment on ‘A Framework for Modelling Trojans and Computer Virus Infection’ by E. Mäkinen”. The Computer Journal, 44(4):324–325, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_44/Issue_
04/440324.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_44/Issue_
04/pdf/440324.pdf.  See

[TAC98, Mäk01, Coh01].

Takahashi:1959:DJD

[Tak59] S. Takahashi. Development of Japanese digital computers. The Computer Journal, 2(3):122–129, October 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/020122.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/122.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/123.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/124.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/125.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/126.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/127.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_02/Issue_
                                        03/tiff/128.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_02/Issue_
                                        03/tiff/129.tif.

Takaoka:1986:SNW

[Tak86]  Tadao Takaoka.  The se-

mantics of new while loop. The Computer Journal, 29 (1):33–35, February 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        01/tiff/33.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_29/Issue_01/tiff/
                                        34.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_29/Issue_01/tiff/35.
                                        tif.

Takaoka:1996:LRP

[Tak96] Tadao Takaoka. A left-to-right preprocessing computation for the Boyer–Moore string matching algorithm. The Computer Journal, 39(5):413–416, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://www.oup.co.uk/jnls/
                                        list/comjnl/hdb/Volume_39/
                                        Issue_05/390413.sgm.abs.html;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_39/Issue_05/
                                        Vol39_05.body.html#AbstractTakaoka.

Takaoka:1999:TAC

[Tak99] Tadao Takaoka. O(1) time algorithms for combinatorial generation by tree traversal. The Computer Journal, 42(5):400– 408, 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
05/420400.sgm.abs.html.

Talbot:1984:ILO

[Tal84]  Steve Talbot.  An investi-

gation into logical optimization of relational query languages. The Computer Journal, 27(4):301–309, November 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/301.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/302.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/303.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/304.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/305.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/306.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/307.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/308.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_

                                        04/tiff/309.tif.

Tanenbaum:1978:CPA

[Tan78] A. S. Tanenbaum. A comparison of PASCAL and AL-GOL 68. The Computer Journal, 21(4):316–323, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        04/tiff/316.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        04/tiff/317.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        04/tiff/318.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        04/tiff/319.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        04/tiff/320.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        04/tiff/321.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        04/tiff/322.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        04/tiff/323.tif.

Tsanakas:1989:FDM

[TAP89] P. Tsanakas, N. Alexandridis, and G. Papakonstantinou. An FP-based design methodology for problem-oriented architectures. The Computer Journal, 32(5):453–460, October 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/453.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/454.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/455.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/456.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/457.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/458.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/459.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
05/tiff/460.tif.

Taylor:1961:EDC

[Tay61] B. R. Taylor. Experience with a digital computer in an aeroplane testing establishment. The Computer Journal, 4(1):25–29, April 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/040025.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/tiff/25.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_01/tiff/
26.tif;   http://www3.oup.
co.uk/computer_journal/hdb/

                                        Volume_04/Issue_01/tiff/27.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        04/Issue_01/tiff/28.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        01/tiff/29.tif.

Taylor:1964:CRA

[Tay64] P. Taylor. Correspondence: On the representation of AL-GOL. The Computer Journal, 6(4):335–335, 367–367, January 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        04/tiff/335.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        04/tiff/367.tif. See [Sam63,

Bar64].

Taylor:1968:COF

[Tay68] R. Taylor. Correspondence: Orion FORTRAN compiler. The Computer Journal, 10(4): 388, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        10/Issue_04/100388.sgm.abs.
                                        html;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        10/Issue_04/tiff/388.tif.

See [TH64, Wil67c].

Taylor:1970:BVP

[Tay70a] J. C. Taylor. Boundary value problems for singular ordinary differential equations. A widely applicable method for calculating eigenvalues. The Computer Journal, 13(3):291–296, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/130291.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/291.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/292.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/293.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/294.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/295.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/296.tif.

Taylor:1970:SDF

[Tay70b] P. J. Taylor. The stability of the Du Fort-Frankel method for the diffusion equation with boundary conditions involving space derivatives. The Computer Journal, 13(1):92–97, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/130092.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/tiff/92.tif;   http://

                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_13/Issue_01/tiff/
                                        93.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_13/Issue_01/tiff/94.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        13/Issue_01/tiff/95.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        01/tiff/96.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_13/Issue_01/tiff/
                                        97.tif.

Taylor:1976:FIO

[Tay76] R. Taylor. Fast input/output of variable-length arrays in FORTRAN IV. The Computer Journal, 19(3):273, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/273.tif.

Taylor:1980:CJS

[Tay80]  N. B. Taylor.  Correspon-

dence: Jumping to some purpose. The Computer Journal, 23(2):189–190, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        02/tiff/189.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        02/tiff/190.tif. See [ASG79,

Hil80, Rob80, MR82].

Taylor:1981:CGP

[Tay81] D. J. Taylor. Correspondence:

GOTOs in practice. The Computer Journal, 24(1):95, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
01/tiff/95.tif.

Taylor:1990:LOD

[Tay90] M. C. Taylor. Logical optimisation of distributed knowledge base queries. The Computer Journal, 33(1):49–60, February 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
01/tiff/49.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_01/tiff/
50.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_33/Issue_01/tiff/51.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/52.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
01/tiff/53.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_01/tiff/
54.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_33/Issue_01/tiff/55.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/56.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_

                                        01/tiff/57.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_33/Issue_01/tiff/
                                        58.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_33/Issue_01/tiff/59.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        33/Issue_01/tiff/60.tif.

Taylor:1996:SFF

[Tay96] Dr. John M. Taylor, OBE, F. Eng. Special feature: Forward with foresight. The Computer Journal, 39(7):561–565, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://www.oup.co.uk/jnls/
                                        list/comjnl/hdb/Volume_39/
                                        Issue_07/390561.sgm.abs.html;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_39/Issue_07/
                                        Vol39_07.body.html#AbstractTaylor.

Thorby:1963:AEC

[TB63] R. P. Thorby and B. Benjamin. The application of the electronic computer to the 1961 population census of Great Britain. The Computer Journal, 5(4):264–270, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        04/050264.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        04/tiff/264.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        04/tiff/265.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/266.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/267.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/268.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/269.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/270.tif.

Thomas:1982:CDD

[TB82] R. C. Thomas and A. Burns.

The case for distributed decision making systems. The Computer Journal, 25(1):148–152, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/148.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/149.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/150.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/151.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/152.tif.

Taylor:1986:LCB

[TB86] D. J. Taylor and J. P. Black.

A locally correctable B-tree implementation. The Computer Journal, 29(3):269–276, June 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        03/tiff/269.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        03/tiff/270.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        03/tiff/271.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        03/tiff/272.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        03/tiff/273.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        03/tiff/274.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        03/tiff/275.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        03/tiff/276.tif.

Tindell:1994:FPS

[TB94] K. Tindell and A. Burns. Fixed priority scheduling of hard realtime multi-media disk traffic. The Computer Journal, 37(8):691–697, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_37/Issue_08/
                                        Vol37_08.body.html#AbstractTindell.

Turtle:1992:CTR

[TC92] Howard R. Turtle and W. Bruce Croft. A comparison of text retrieval models. The Computer Journal, 35(3):279–290, June 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_35/Issue_03/
Vol35_03.body.html#AbstractTurtle.

Tomlin:1971:BRI

[TCB71] J. A. Tomlin, D. C. Cooper, and D. A. Bell. Book reviews: Integer and Dynamic Programming, by J. Abadie (editor), 1970; Algorithms and Recursive Functions, by A. L. Mal’cev, 1970; Information Theory for Systems Engineers (Econometrics and Operations Research XVII), by L. P. Hyvärinen, 1970. The Computer Journal, 14(4):343, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/343.tif.

Tangney:1994:RPP

[TCCH94] Brendan Tangney, Andrew Condon, Vinny Cahill, and Neville Harris. Requirements for parallel programming in object-oriented distributed systems. The Computer Journal, 37(6):499–508, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_37/Issue_06/
Vol37_06.body.html#AbstractTangney.

Tan:1989:AID

[TCL89] K. P. Tan, T. S. Chua, and P. T. Lee. AUTO-DFD: an intelligent data flow processor. The Computer Journal, 32(3):194–201, June 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        03/tiff/194.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        03/tiff/195.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        03/tiff/196.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        03/tiff/197.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        03/tiff/198.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        03/tiff/199.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        03/tiff/200.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        03/tiff/201.tif.

Taylor:1994:COO

[TCM94] Paul Taylor, Vinny Cahill, and Michael Mock. Combining object-oriented systems and open transaction processing. The Computer Journal, 37(6):487–498, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_37/Issue_06/
Vol37_06.body.html#AbstractTaylor.

Tosser:1977:AON

[TD77] A. Tosser and D. Dubus. Algebraic optimisation of NAND-NOR switching circuits. The Computer Journal, 20(1):73– 77, February 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
20/Issue_01/tiff/73.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
01/tiff/74.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_01/tiff/
75.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_20/Issue_01/tiff/76.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
20/Issue_01/tiff/77.tif.

Tomov:1999:AET

[TDW^{+}99] Neven Tomov, Euan Dempster, M. Howard Williams, Peter J. B. King, and Albert Burger. Approximate estimation of transaction response time. The Computer Journal, 42(3):241–250, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
03/420241.sgm.abs.html;
http://www3.oup.co.uk/computer_

journal/hdb/Volume_42/Issue_
03/pdf/420241.pdf.

Tee:1963:ESR

[Tee63a] G. J. Tee. Eigenvectors of the successive over-relaxation process, and its combination with Chebyshev semiiteration. The Computer Journal, 6(3):250–263, October 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/060250.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/250.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/251.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/252.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/253.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/254.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/255.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/256.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/257.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/258.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/259.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/260.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/261.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/262.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
03/tiff/263.tif.

Tee:1963:NFD

[Tee63b] G. J. Tee. A novel finitedifference approximation to the biharmonic operator. The Computer Journal, 6(2):177–192, July 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/060177.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/177.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/178.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/179.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/180.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/181.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_

                                        02/tiff/182.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/tiff/183.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/tiff/184.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/tiff/185.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/tiff/186.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/tiff/187.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/tiff/188.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/tiff/189.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/tiff/190.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/tiff/191.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        02/tiff/192.tif.

Tee:1967:CCV

[Tee67] G. H. Tee. Correspondence: On coding values of mathematical constants. The Computer Journal, 9(4):393, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        04/090393.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_

journal/hdb/Volume_09/Issue_
04/tiff/393.tif.

Temes:1969:CNA

[Tem69] G. C. Temes. Correspondence:

Note on “An algorithm for minimax approximation in the nonlinear case”. The Computer Journal, 12(4):410, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/410.tif. See [OW69].

Templeman:1972:CDC

[Tem72]  J. A. Templeman.  Cor-

respondence: On “A data collection and display system for a large-scale simulation”. The Computer Journal, 15(4):359, November 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/150359.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/359.tif. See [BPSS72].

Terrine:1973:CGP

[Ter73] G. Terrine. Coordinate grammars and parsers. The Computer Journal, 16(3):232– 244, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
16/Issue_03/160232.sgm.abs.

html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
16/Issue_03/tiff/232.tif;
 http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/233.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/234.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/235.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/236.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/237.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/238.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/239.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/240.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/241.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/242.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/243.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/244.tif.

Teskey:1977:KME

[Tes77] N. Teskey. KATE: a macroprocessor for extending command languages. The Computer Journal, 20(2):187–189, May 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/187.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/188.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/189.tif.

Teufel:1987:SDR

[Teu87] B. Teufel. System design for the remote execution of library routines. The Computer Journal, 30(3):254–257, June 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/254.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/255.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/256.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/257.tif.

Tewarson:1967:RCP

[Tew67] R. P. Tewarson. Row-column permutation of sparse matrices. The Computer Journal, 10(3):300–305, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

journal/hdb/Volume_10/Issue_
03/100300.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/300.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/301.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/302.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/303.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/304.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
03/tiff/305.tif.

Tewarson:1968:CME

[Tew68]  R. P. Tewarson.  A com-

putational method for evaluating generalized inverses. The Computer Journal, 10 (4):411–413, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/100411.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/411.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/412.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/413.tif.

Tewarson:1969:CRS

[Tew69a] R. P. Tewarson. The Crout reduction for sparse matrices. The Computer Journal, 12(2):158–159, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/120158.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/158.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/159.tif.

Tewarson:1969:LSI

[Tew69b]  R. P. Tewarson.  A least

squares iterative method for singular equations. The Computer Journal, 12(4):388–392, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/120388.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/388.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/389.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/390.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/391.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_12/Issue_
04/tiff/392.tif.

Tewarson:1969:PMS

[Tew69c] R. P. Tewarson. Projection methods for solving sparse linear systems. The Computer Journal, 12(1):77–80, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
01/120077.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
01/tiff/77.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_01/tiff/
78.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_12/Issue_01/tiff/79.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/80.tif.

Tewarson:1972:MSL

[Tew72] R. P. Tewarson. On minimax solutions of linear equations. The Computer Journal, 15(3):277–279, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/150277.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/277.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/278.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_15/Issue_
03/tiff/279.tif.

Tsutsui:1987:DPP

[TF87] S. Tsutsui and Y. Fujimoto.

Deadlock prevention in process control computer systems. The Computer Journal, 30(1):20–26, February 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
01/tiff/20.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_01/tiff/
21.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_30/Issue_01/tiff/22.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
30/Issue_01/tiff/23.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
01/tiff/24.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_01/tiff/
25.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_30/Issue_01/tiff/26.
tif.

Theel:1996:BRC

[TF96] Oliver E. Theel and Brett D.

Fleisch. The boundaryrestricted coherence protocol for scalable and highly available distributed shared memory systems. The Computer Journal, 39(6):496–510, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://www.oup.co.uk/jnls/
                                        list/comjnl/hdb/Volume_39/
                                        Issue_06/390496.sgm.abs.html;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_39/Issue_06/
                                        Vol39_06.body.html#AbstractTheel.

Tubella:1996:MAP

[TGE96] J. Tubella, A. Gonzalez, and Eduard Elias. The multipath architecture for Prolog programs. The Computer Journal, 39(9):780–??, ???? 1996. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www.oup.co.uk/computer_
                                        journal/Volume_39/Issue_09/
                                        Vol39_09.body.html#AbstractTubella;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_39/Issue_09/
                                        Vol39_09.body.html#AbstractTubella.

Taylor:1964:FSO

[TH64] R. Taylor and D. A. Harragan. The FORTRAN system for Orion. The Computer Journal, 7(2):114–116, July 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        02/070114.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        02/tiff/114.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        02/tiff/115.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_

02/tiff/116.tif. See corre-

spondence [Wil67c, Tay68].

Tan:1982:BSL

[TH82] K. P. Tan and L. S. Hsu.

Block sorting of a large file in external storage by a 2component key. The Computer Journal, 25(3):327–330, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/327.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/328.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/329.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/330.tif.

Tsaur:1998:ESE

[TH98] Woei-Jiunn Tsaur and Shi-Jinn Horng. Establishing secure Ethernet LANs using intelligent switching hubs in Internet environments. The Computer Journal, 41(2):125–133, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_41/Issue_02/
Vol41_02.body.html#AbstractTsaur.

Thacher:1962:IMQ

[Tha62] H. C. Thacher, Jr. An iterative method for quadratures. The Computer Journal, 5(3):228–229, October 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/050228.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/228.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/229.tif.

Thacher:1966:CRI

[Tha66] H. C. Thacher, Jr. Closed rational integration formulas. The Computer Journal, 8(4):362–367, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/080362.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/362.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/363.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/364.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/365.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/366.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
04/tiff/367.tif.

Halsbury:1991:IFS

[The91a] The Rt. Hon. The Earl Of Halsbury. Innovation for failure — some reflections on the work of the NRDC relevant to the early history of the computer industry in the UK. The Computer Journal, 34(3):272–274, June 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/272.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/273.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/274.tif.

Halsbury:1991:TYC

[The91b] The Rt. Hon. The Earl Of Halsbury. Ten years of computer development. The Computer Journal, 34(3):275–279, June 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/275.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/276.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/277.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/278.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_34/Issue_
                                        03/tiff/279.tif.

Tawbi:1993:VCS

[THHS93] W. Tawbi, F. Horn, E. Horlait, and J.-B. Stefani. Video compression standards and quality of service. The Computer Journal, 36(1):43–54, February 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_36/Issue_01/
                                        Vol36_01.body.html#AbstractTawbi.

Thimbleby:1981:WBA

[Thi81] H. W. Thimbleby. A word boundary algorithm for text processing. The Computer Journal, 24(3):249–255, August 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        03/tiff/249.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        03/tiff/250.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        03/tiff/251.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        03/tiff/252.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        03/tiff/253.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        03/tiff/254.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_24/Issue_
03/tiff/255.tif.

Thimbleby:1986:ELP

[Thi86] H. Thimbleby. Experiences of ‘literate programming’ using Cweb (A variant of Knuth’s Web). The Computer Journal, 29(3):201–211, June 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/201.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/202.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/203.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/204.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/205.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/206.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/207.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/208.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/209.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/210.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_29/Issue_
                                        03/tiff/211.tif.

Thimbleby:1995:NCW

[Thi95] Harold Thimbleby. A new calculator and why it is necessary. The Computer Journal, 38(6):418–433, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_38/Issue_06/
                                        Vol38_06.body.html#AbstractThimbleby.

Thompson:1958:FYA

[Tho58a]  T. R. Thompson.  Four

years of automatic office work. The Computer Journal, 1 (3):106–112, October 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        03/010106.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        03/tiff/106.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        03/tiff/107.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        03/tiff/108.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        03/tiff/109.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        03/tiff/110.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        03/tiff/111.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/112.tif.

Thomson:1958:MCM

[Tho58b] W. E. Thomson. A modified congruence method of generating pseudo-random numbers. The Computer Journal, 1(2):83, July 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
02/010083.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
02/tiff/83.tif.

Thompson:1960:PAC

[Tho60] T. R. Thompson. Problems of auditing computing data: internal audit practice and external audit theory. The Computer Journal, 3(1):10–11, April 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
01/030010.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
01/tiff/10.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_01/tiff/
11.tif.

Thompson:1961:IAC

[Tho61]  J. G. Thompson.  An in-

troduction to analogue computer methods. The Computer Journal, 3(4):211–219, January 1961. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/030211.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/211.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/212.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/213.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/214.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/215.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/216.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/217.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/218.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/219.tif.

Thompson:1962:FPE

[Tho62]  T. R. Thompson.  Funda-

mental principles of expressing a procedure for a computer application. The Computer Journal, 5(3):164–169, October 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_

03/050164.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/164.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/165.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/166.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/167.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/168.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/169.tif.

Thomas:1964:RIU

[Tho64] D. G. W. Thomas. Record identification using variable alphanumeric names. The Computer Journal, 7(1):19–23, April 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
01/070019.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
01/tiff/19.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_01/tiff/
20.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_07/Issue_01/tiff/21.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
07/Issue_01/tiff/22.tif;
http://www3.oup.co.uk/computer_

                                        journal/hdb/Volume_07/Issue_
                                        01/tiff/23.tif.

Thomson:1968:CGT

[Tho68a] W. E. Thomson. Correspondence: Generation of time delays on analogue computers. The Computer Journal, 11(2):172, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        02/tiff/172.tif.

Thomson:1968:CSL

[Tho68b] W. E. Thomson. Correspondence: Solution of linear differential equation. The Computer Journal, 10(4):417–418, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/tiff/417.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/tiff/418.tif.

Thornton:1972:RLQ

[Tho72] B. S. Thornton. Reduction and linearisation of quadratic constrained optimisation in linked systems. The Computer Journal, 15(1):84–87, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        01/150084.sgm.abs.html;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/tiff/84.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_01/tiff/
85.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_15/Issue_01/tiff/86.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
15/Issue_01/tiff/87.tif.

Thoma:1980:CUB

[Tho80] Martyn Thoma. Correspondence: The unconditional branch — a modest proposal. The Computer Journal, 23(4):380–381, November 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/380.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/381.tif. See corre-

spondence [Wig81].

Thompson:1993:CIP

[Tho93]  Peter Thompson.  Concur-

rent interconnect for parallel systems. The Computer Journal, 36(8):778–784, ???? 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_08/
Vol36_08.body.html#AbstractThompson.

Thomas:1998:MY

[Tho98] Martyn Thomas. Myths of the Year 2000. The Computer Journal, 41(2):67–70, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_41/Issue_02/
Vol41_02.body.html#AbstractThomas.

terHofstede:1999:CSV

[tHO99] A. H. M. ter Hofstede and M. E. Orlowska. On the complexity of some verification problems in process control specifications. The Computer Journal, 42(5):349–359, 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
05/420349.sgm.abs.html.

Throsby:1969:FEA

[Thr69] P. W. Throsby. A finite element approach to surface definition. The Computer Journal, 12(4):385–387, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/120385.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/385.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/386.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/387.tif.

Treleaven:1982:CDF

[THR82] Philip C. Treleaven, Richard P.

Hopkins, and Paul W. Rautenbach. Combining data flow and control flow computing. The Computer Journal, 25(2):207–217, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/207.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/208.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/209.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/210.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/211.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/212.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/213.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/214.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/215.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/216.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/217.tif.

Tabata:1975:ETI

[TI75] T. Tabata and R. Ito. Effective treatment of the interpolation factor in Marquardt’s nonlinear least-squares fit algorithm. The Computer Journal, 18(3):250–251, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/250.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/251.tif.

Tily:1969:AFD

[Til69] C. N. J. Tily. Algorithm 37: Fitting data to an exponentially damped linear function. The Computer Journal, 12(1):100–101, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        01/tiff/100.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        01/tiff/101.tif.

Timbleby:1987:OSP

[Tim87] H. Timbleby. Optimising selfreplicating programs. The Computer Journal, 30(5):475– 476, October 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        30/Issue_05/tiff/475.tif;

 http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
05/tiff/476.tif.

Tizard:1960:CBA

[Tiz60] R. H. Tizard. Conversion between analogue and digital measures. The Computer Journal, 3(1):51–59, April 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
01/030051.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
01/tiff/51.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_01/tiff/
52.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_03/Issue_01/tiff/53.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
03/Issue_01/tiff/54.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
01/tiff/55.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_01/tiff/
56.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_03/Issue_01/tiff/57.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
03/Issue_01/tiff/58.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
01/tiff/59.tif.

Tahar:1995:FSV

[TK95] Sofiène Tahar and Ramayya Kumar. Formal specification and verification techniques for RISC pipeline conflicts. The Computer Journal, 38(2):111–120, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_38/Issue_02/
                                        Vol38_02.body.html#AbstractTahar.

Tsuchiya:1999:AEQ

[TK99] Tatsuhiro Tsuchiya and Tohru Kikuno. Availability evaluation of quorum-based mutual exclusion schemes in general topology networks. The Computer Journal, 42(7):613– 622, 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        07/420613.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        07/pdf/420613.pdf.

Triance:1985:LEF

[TL85a] J. M. Triance and P. J. Layzell. Language enhancement facility for COBOL — its design and implementation. The Computer Journal, 28(2):128–133, May 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        02/tiff/128.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        02/tiff/129.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/130.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/131.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/132.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/133.tif.

Triance:1985:MPE

[TL85b] J. M. Triance and P. J. Layzell.

Macro processors for enhancing high-level languages — some design principles. The Computer Journal, 28(1):34–43, February 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
01/tiff/34.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_01/tiff/
35.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_28/Issue_01/tiff/36.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
28/Issue_01/tiff/37.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
01/tiff/38.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_01/tiff/
39.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_28/Issue_01/tiff/40.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
28/Issue_01/tiff/41.tif;

                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        01/tiff/42.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_28/Issue_01/tiff/
                                        43.tif.

Turner:1985:IFS

[TL85c] R. Turner and B. G. T. Lowden. An introduction to the formal specification of relational query languages. The Computer Journal, 28(2):162–169, May 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        02/tiff/162.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        02/tiff/163.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        02/tiff/164.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        02/tiff/165.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        02/tiff/166.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        02/tiff/167.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        02/tiff/168.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_28/Issue_
                                        02/tiff/169.tif.

Thompson:1992:CMC

[TL92] S. J. Thompson and R. D. Lins. The Categorical Multi-Combinator Machine: CMCM. The Computer Journal, 35(2): 170–176, April 1992. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/Volume_
35/Issue_02/Vol35_02.body.
html#AbstractThompson.

Tsai:1997:MPR

[TL97] W.-J. Tsai and S.-Y. Lee. Multipartition RAID: A new method for improving performance of disk arrays under failure. The Computer Journal, 40(1):30–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www.oup.co.uk/computer_
journal/Volume_40/Issue_01/
Vol40_01.body.html#AbstractTsai;
 http://www3.oup.co.uk/computer_
journal/Volume_40/Issue_01/
Vol40_01.body.html#AbstractTsai.

Tse:1998:MEI

[TL98] Savio S. H. Tse and Francis C. M. Lau. More on the efficiency of interval routing. The Computer Journal, 41(4):238–??, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_41/Issue_04/
Vol41_04.body.html#AbstractTse.

TerHofstede:1996:CDM

[TLF96] A. H. M. Ter Hofstede, E. Lippe, and P. J. M. Frederiks. Conceptual data modelling from a categorical perspective. The Computer Journal, 39(3):215– 231, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://www.oup.co.uk/jnls/
                                        list/comjnl/hdb/Volume_39/
                                        Issue_03/390215.sgm.abs.html;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_39/Issue_03/
                                        Vol39_03.body.html#AbstractHofstede.

Thomson:1994:SPP

[TM94] Helen E. Thomson and Pam Mayhew. The software process: A perspective on improvement. The Computer Journal, 37(8):683–690, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_37/Issue_08/
                                        Vol37_08.body.html#AbstractThomson.

Turpin:1995:PLL

[TM95] Andrew Turpin and Alistair Moffat. Practical lengthlimited coding for large alphabets. The Computer Journal, 38(5):339–347, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_38/Issue_05/
                                        Vol38_05.body.html#AbstractTurpin.

Tompa:1967:GNI

[Tom67]  H. Tompa.  Gaussian nu-

merical integration of a function depending on a parameter. The Computer Journal, 10(2):204–205, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/100204.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/204.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/205.tif.

Tomlinson:1982:AHS

[Tom82] D. J. Tomlinson. An aid to hidden surface removal in real time CGI systems. The Computer Journal, 25(4):429–441, November 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/429.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/430.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/431.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/432.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/433.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/434.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/435.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_25/Issue_
                                        04/tiff/436.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        04/tiff/437.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        04/tiff/438.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        04/tiff/439.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        04/tiff/440.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        04/tiff/441.tif.

Topor:1982:FPG

[Top82]  R. W. Topor.  Functional

programs for generating permutations. The Computer Journal, 25(2):257–263, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/257.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/258.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/259.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/260.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/261.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/262.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/263.tif.

Townsend:1977:SPA

[Tow77] H. R. A. Townsend. SST — a program for the automatic generation of symbol state tables. The Computer Journal, 20(2):124–127, May 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/124.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/125.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/126.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/127.tif.

Townsend:1978:AE

[Tow78] H. R. A. Townsend. Analysing electroencephalograms. The Computer Journal, 21(3):259– 262, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
21/Issue_03/tiff/259.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/260.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/261.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_21/Issue_
03/tiff/262.tif.

Tse:1989:TFF

[TP89] T. H. Tse and L. Pong. Towards a formal foundation for DeMarco data flow diagrams. The Computer Journal, 32(1):1–12, February 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
01/tiff/1.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_32/Issue_01/tiff/
10.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_32/Issue_01/tiff/11.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
32/Issue_01/tiff/12.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
01/tiff/2.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_32/Issue_01/tiff/
3.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_32/Issue_01/tiff/4.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
32/Issue_01/tiff/5.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
01/tiff/6.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_32/Issue_01/tiff/
7.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_32/Issue_01/tiff/8.
tif;   http://www3.oup.co.

uk/computer_journal/hdb/Volume_
32/Issue_01/tiff/9.tif.

Tse:1991:ERS

[TP91] T. H. Tse and L. Pong. An examination of requirements specification languages. The Computer Journal, 34(2):143–152, April 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/143.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/144.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/145.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/146.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/147.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/148.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/149.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/150.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/151.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/152.tif.

TerHofstede:1996:QFI

[TPvdW96] A. H. M. Ter Hofstede, H. A. Proper, and Th. P. van der Weide. Query formulation as an information retrieval problem. The Computer Journal, 39(4):255–274, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://www.oup.co.uk/jnls/
                                        list/comjnl/hdb/Volume_39/
                                        Issue_04/390255.sgm.abs.html;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_39/Issue_04/
                                        Vol39_04.body.html#AbstractHofstede.

Teuhola:1993:AFS

[TR93] Jukka Teuhola and Timo Raita. Application of a finitestate model to text compression. The Computer Journal, 36(7):607–614, ???? 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_36/Issue_07/
                                        Vol36_07.body.html#AbstractTeuhola.

Travis:1982:APR

[Tra82] A. J. B. Travis. An aid to pattern recognition. The Computer Journal, 25(1):37–44, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/37.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_25/Issue_01/tiff/
                                        38.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_25/Issue_01/tiff/39.

tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/40.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/41.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/
42.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_25/Issue_01/tiff/43.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/44.tif.

Travis:1984:APR

[Tra84] A. J. B. Travis. An aid to pattern recognition. II. The Computer Journal, 27 (4):362–367, November 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/362.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/363.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/364.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/365.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/366.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/367.tif.

Treble:1970:CCC

[Tre70] D. P. Treble. Computers and composition in change ringing. The Computer Journal, 13(4):350–351, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/130350.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/350.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        04/tiff/351.tif.

Treleaven:1982:WRV

[Tre82] Philip C. Treleaven. Workshop report: VLSI Machine Architectures and Very High Level Languages. The Computer Journal, 25(1):153–157, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/153.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/154.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/155.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/156.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/157.tif.

Trenouth:1991:SES

[Tre91] J. Trenouth. A survey of exploratory software development. The Computer Journal, 34(2):153–163, April 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/153.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/154.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/155.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/156.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/157.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/158.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/159.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/160.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/161.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/162.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
02/tiff/163.tif.

Tredoux:1992:MES

[Tre92] G. Tredoux. Mechanizing execution sequence semantics in HOL. The Computer Journal, 7(??):81–86, ???? 1992. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). Proceedings of the 7th Southern African Computer Research Symposium, Johannesburg, South Africa.

Triance:1974:HRV

[Tri74]  J. M. Triance.  Handling

records with a variable structure in COBOL. The Computer Journal, 17(1):93–94, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        01/tiff/93.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_17/Issue_01/tiff/
                                        94.tif.

Triance:1976:SCS

[Tri76] J. M. Triance. The significance of the 1974 COBOL standard. The Computer Journal, 19(4):295–300, November 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        04/tiff/295.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        04/tiff/296.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        04/tiff/297.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        04/tiff/298.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_

04/tiff/299.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/300.tif.

Triance:1978:SCP

[Tri78]  J. M. Triance.  A study

of COBOL portability. The Computer Journal, 21(3):278– 281, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
21/Issue_03/tiff/278.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/279.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/280.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/281.tif.

Triance:1980:SPC

[Tri80] J. M. Triance. Structured programming in COBOL — the current options. The Computer Journal, 23(3):194–200, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/194.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/195.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/196.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_

03/tiff/197.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/198.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/199.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/200.tif.

Triance:1983:PCS

[Tri83]  J. M. Triance.  The pro-

posed COBOL Standard — its significance. The Computer Journal, 26(1):60–67, February 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
01/tiff/60.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_01/tiff/
61.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_26/Issue_01/tiff/62.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
26/Issue_01/tiff/63.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
01/tiff/64.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_01/tiff/
65.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_26/Issue_01/tiff/66.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
26/Issue_01/tiff/67.tif.

Trundle:1966:LAL

[Tru66] R. W. L. Trundle. LITHP — an ALGOL list processor. The Computer Journal, 9(2):167–172, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/090167.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/167.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/168.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/169.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/170.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/171.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/172.tif.

Taylor-Smith:1964:ACI

[TS64] A. Taylor-Smith. Authorization and control of input in the Royal Army Pay Corps Computer Centre application. The Computer Journal, 7(1):14–18, April 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
01/070014.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
01/tiff/14.tif;   http://
www3.oup.co.uk/computer_journal/

                                        hdb/Volume_07/Issue_01/tiff/
                                        15.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_07/Issue_01/tiff/16.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        07/Issue_01/tiff/17.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        01/tiff/18.tif.

Temperly:1968:GPP

[TS68] J. F. Temperly and B. W. Smith. A grading procedure for PL/1 student exercises. The Computer Journal, 10(4):368–373, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/100368.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/tiff/368.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/tiff/369.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/tiff/370.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/tiff/371.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/tiff/372.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/tiff/373.tif.

Trapp:1975:SNV

[TS75] G. E. Trapp, Jr. and W. Squire.

Solving nonlinear Vandermonde systems. The Computer Journal, 18(4):373–374, November 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/373.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/374.tif.

Tse:1987:IPU

[Tse87a] T. H. Tse. The identification of program unstructuredness: a formal approach. The Computer Journal, 30(6):507–511, December 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/507.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/508.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/509.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/510.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
06/tiff/511.tif.

Tse:1987:TSC

[Tse87b]  T. H. Tse.  Towards a

single criterion for identifying program unstructuredness. The Computer Journal, 30(4):378–380, August 1987. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        04/tiff/378.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        04/tiff/379.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        04/tiff/380.tif.

Tyler:1973:CE

[TT73] C. R. Tyler and The Editor. Correspondence: Editorial. The Computer Journal, 16(3):286–287, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        03/tiff/286.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        03/tiff/287.tif. See [Ham73,

CT74, Fis74b].

Turkan:1996:EMM

[TU96] Yasemin Turkan and Özgür Ulusoy. An efficient memory management algorithm for client-server database management systems. The Computer Journal, 39(5):374–385, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://www.oup.co.uk/jnls/
                                        list/comjnl/hdb/Volume_39/
                                        Issue_05/390374.sgm.abs.html;
                                         http://www3.oup.co.uk/computer_

journal/Volume_39/Issue_05/
Vol39_05.body.html#AbstractTurkan.

Tully:1985:IHA

[Tul85] C. J. Tully. Information, human activity and the nature of relevant theories. The Computer Journal, 28(3):206–210, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/206.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/207.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/208.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/209.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/210.tif.

Turski:1968:SDA

[Tur68] W. M. Turski. SODA — a dual activity operating system. The Computer Journal, 11(2):148–156, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/110148.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/148.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/149.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/150.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/151.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/152.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/153.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/154.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/155.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/156.tif.

Turner:1975:GCS

[Tur75]  L. F. Turner.  The on-

ground compression of satellite data. The Computer Journal, 18(3):243–247, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/243.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/244.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/245.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/246.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_18/Issue_
03/tiff/247.tif.

Turski:1978:MND

[Tur78] W. M. Turski. A model of nonmorphological data collection based on sentential calculus. The Computer Journal, 21(4):337–340, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/337.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/338.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/339.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/340.tif. See note

[Hay79].

Turner:1979:CDW

[Tur79] J. C. Turner. The conditional distribution of waiting time given queue length in a computer system. The Computer Journal, 22(1):57–62, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
01/tiff/57.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_01/tiff/
58.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_22/Issue_01/tiff/59.
tif;   http://www3.oup.co.

                                        uk/computer_journal/hdb/Volume_
                                        22/Issue_01/tiff/60.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        01/tiff/61.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_22/Issue_01/tiff/
                                        62.tif.

Turega:1992:CAS

[Tur92] M. Turega. A computer architecture to support neural net simulation. The Computer Journal, 35(4):353–360, August 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_35/Issue_04/
                                        Vol35_04.body.html#AbstractTurega.

Tayeb:1998:QBD

[TUW98] Jamel Tayeb, Özgür Ulusoy, and Ouri Wolfson. A quadtreebased dynamic attribute indexing method. The Computer Journal, 41(3):185–200, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_41/Issue_03/
                                        Vol41_03.body.html#AbstractTayeb.

Thorp:1964:PAG

[TW64] E. O. Thorp and W. E. Walden. A partial analysis of Go. The Computer Journal, 7(3):203–207, October 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_

journal/hdb/Volume_07/Issue_
03/070203.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/203.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/204.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/205.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/206.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/207.tif.

Ting:1977:MRS

[TW77] T. C. Ting and Y. W. Wang.

Multiway replacement selection sort with dynamic reservoir. The Computer Journal, 20(4):298–301, November 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/298.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/299.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/300.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/301.tif.

Twizell:1975:NSS

[Twi75] E. H. Twizell. The numerical solution of second order hyperbolic partial differential equations with unequally spaced initial conditions. The Computer Journal, 18(3):252–257, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/252.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/253.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/254.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/255.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/256.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        03/tiff/257.tif.

Tischer:1993:CBL

[TWMG93] P. E. Tischer, R. T. Worley, A. J. Maeder, and M. Goodwin. Context-based lossless image compression. The Computer Journal, 36(1):68–77, February 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_36/Issue_01/
                                        Vol36_01.body.html#AbstractTischer.

Tyldesley:1988:EUE

[Tyl88]  D. A. Tyldesley.  Employ-

ing usability engineering in the development of office products. The Computer Journal, 31(5):431–436, October 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/431.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/432.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/433.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/434.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/435.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
05/tiff/436.tif.

Tzafestas:1973:SEA

[Tza73] S. G. Tzafestas. State estimation algorithms for nonlinear stochastic sequential machines. The Computer Journal, 16(3):245–253, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/160245.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/245.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/246.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/247.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_16/Issue_
                                        03/tiff/248.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        03/tiff/249.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        03/tiff/250.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        03/tiff/251.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        03/tiff/252.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        03/tiff/253.tif.

Ueda:1990:DKL

[UC90] Kazunori Ueda and Takashi C. Chikayama. Design of the kernel language for the parallel inference machine. The Computer Journal, 33(6):494–500, December 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/494.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/495.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/496.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/497.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/498.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_

06/tiff/499.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/500.tif.

Uckan:1985:RSD

[Uck85] Yuksel Uckan. A relational schema description and manipulation facility. The Computer Journal, 28(2):170–178, May 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/170.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/171.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/172.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/173.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/174.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/175.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/176.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/177.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/178.tif.

Ullmann:1985:CAS

[UHS85] Julian R. Ullmann, Robert M.

Haralick, and Linda G. Shapiro.

Computer architecture for solving consistent labelling problems. The Computer Journal, 28(2):105–111, May 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/105.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/106.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/107.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/108.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/109.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/110.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/111.tif.

Ulery:1974:SLS

[UK74] Dana L. Ulery and H. M. Khalil. A survey of languageoriented systems for numerical linear algebra. The Computer Journal, 17(1):82–88, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
01/tiff/82.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/
83.tif;   http://www3.oup.

co.uk/computer_journal/hdb/
Volume_17/Issue_01/tiff/84.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
17/Issue_01/tiff/85.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
01/tiff/86.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/
87.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_17/Issue_01/tiff/88.
tif.

Ullmann:1965:BAP

[Ull65]  J. R. Ullmann.  A basic

approach to pattern recognition. The Computer Journal, 7(4):282–289, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/070282.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/282.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/283.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/284.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/285.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/286.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/287.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        04/tiff/288.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        04/tiff/289.tif.

Ullmann:1967:AIP

[Ull67]  J. R. Ullmann.  Algebraic

inference of pattern similarity. The Computer Journal, 10(3):256–264, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        03/100256.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        03/tiff/256.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        03/tiff/257.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        03/tiff/258.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        03/tiff/259.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        03/tiff/260.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        03/tiff/261.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        03/tiff/262.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        03/tiff/263.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_10/Issue_
03/tiff/264.tif.

Ullmann:1973:BRI

[Ull73]  J. R. Ullmann.  Book re-

view: Introduction to Mathematical Techniques in Pattern Recognition, by Harry C. Andrews, 1972. The Computer Journal, 16(3):225, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/225.tif.

Ullmann:1977:BGT

[Ull77] J. R. Ullmann. A binary ngram technique for automatic correction of substitution, deletion, insertion and reversal errors in words. The Computer Journal, 20(2):141–147, May 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/141.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/142.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/143.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/144.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/145.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_

                                        02/tiff/146.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        02/tiff/147.tif.

Ullmann:1988:FIR

[Ull88]  J. R. Ullmann.  Fast im-

plementation of relational operations via inverse projections. The Computer Journal, 31(2):147–154, April 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        02/tiff/147.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        02/tiff/148.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        02/tiff/149.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        02/tiff/150.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        02/tiff/151.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        02/tiff/152.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        02/tiff/153.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        02/tiff/154.tif.

Underhill:1963:GCG

[Und63] L. H. Underhill. The growth of complexity of a generalpurpose program. The Computer Journal, 6(1):37–38, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/060037.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/tiff/37.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/
38.tif.

Umphress:1989:FPG

[UPT89] D. A. Umphress, U. W. Pooch, and M. Tanik. Fast prototyping of a goal-oriented simulation environment system. The Computer Journal, 32(6):541–548, December 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/541.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/542.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/543.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/544.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/545.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/546.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
06/tiff/547.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        06/tiff/548.tif.

Unnikrishnan:1987:CCL

[UVS87] A. Unnikrishnan, Y. V. Venkatesh, and Priti Shankar. Connected component labelling using quadtrees — A bottomup approach. The Computer Journal, 30(2):176–182, April 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        02/tiff/176.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        02/tiff/177.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        02/tiff/178.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        02/tiff/179.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        02/tiff/180.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        02/tiff/181.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        02/tiff/182.tif.

Valentine:1971:CAN

[Val71] S. H. Valentine. Correspondence: “On the automatic numerical evaluation of definite integrals”. The Computer Journal, 14(4):365, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/365.tif. See [CP71].

Valentine:1974:CNA

[Val74] S. H. Valentine. Comparative notes on ALGOL 68 and PL/I. The Computer Journal, 17(4):325–331, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/325.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/326.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/327.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/328.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/329.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/330.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/331.tif.

vanRijsbergen:1970:CA

[van70] C. J. van Rijsbergen. Algorithm 47: A clustering algorithm. The Computer Journal, 13(1):113–115, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

                                        journal/hdb/Volume_13/Issue_
                                        01/tiff/113.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        01/tiff/114.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        01/tiff/115.tif. See notes

[vR70b, Mux70, Sal70c].

vanRijsbergen:1971:AIS

[van71] C. J. van Rijsbergen. An algorithm for information structuring and retrieval. The Computer Journal, 14(4):407–412, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/140407.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/407.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/408.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/409.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/410.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/411.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/412.tif.

VanHoa:1973:ACP

[Van73] Tran Van Hoa. Algorithm 78: Counting preferential votes in multi-member constituencies using absolute majority criteria. The Computer Journal, 16(3):273–276, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/273.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/274.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/275.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/276.tif. See note

[HW74, Par74].

vanRijsbergen:1986:NLI

[van86]  C. J. van Rijsbergen.  A

non-classical logic for information retrieval. The Computer Journal, 29(6):481–485, December 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/481.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/482.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/483.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
06/tiff/484.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_

                                        06/tiff/485.tif.

vanderWeide:1988:ITT

[van88] Th. P. van der Weide. Iterative tree traversal using refinements and pointer rotation. The Computer Journal, 31(3):269–275, June 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        03/tiff/269.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        03/tiff/270.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        03/tiff/271.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        03/tiff/272.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        03/tiff/273.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        03/tiff/274.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        03/tiff/275.tif.

Vanthienen:1991:ALP

[Van91]  J. Vanthienen.  Algorithm

124: A longest path algorithm to display decision tables. The Computer Journal, 34(4):358–362, August 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        04/tiff/358.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/359.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/360.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/361.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/362.tif.

vanRijsbergen:1992:PRR

[van92] C. J. van Rijsbergen. Probabilistic retrieval revisited. The Computer Journal, 35(3):291– 298, June 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/Volume_
35/Issue_03/Vol35_03.body.
html#AbstractRijsbergen.

vanRijsbergen:1993:W

[van93] Keith van Rijsbergen, Editorin-Chief. Welcome. The Computer Journal, 36(1):1, ???? 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_01/
Vol36_01.body.html#AbstractRijsbergen.

vanderGaag:1996:BBN

[van96] Linda C. van der Gaag.

Bayesian belief networks: Odds and ends. The Computer Journal, 39(2):97–113, ???? 1996. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www.oup.co.uk/jnls/
list/comjnl/hdb/Volume_39/
Issue_02/390097.sgm.abs.html;
 http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_02/
Vol39_02.body.html#AbstractGaag.

Various:1962:DSA

[Var62] Various. Discussion: Session 2: 17 April 1962. The Computer Journal, 5(3):170–176, October 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/170.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/171.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/172.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/173.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/174.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/175.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/176.tif.

Various:1967:DCO

[Var67] Various. Discussion and correspondence: One-day symposium on PL/1 held at National Physical Laboratory, 18 May 1967. The Computer Journal, 10(2):211–215, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/211.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/212.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/213.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/214.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/215.tif.

Various:1968:AS

[Var68] Various. Algorithms supplement. The Computer Journal, 11(2):231–235, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/231.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/232.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/233.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/234.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/235.tif.

Varney:1976:TUO

[Var76a] R. C. Varney. Toward the understandability of an operating system. The Computer Journal, 19(3):213–215, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/213.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/214.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        03/tiff/215.tif.

Varol:1976:SRC

[Var76b]  Y. L. Varol.  Some re-

marks on computer acquisition. The Computer Journal, 19(2):127–131, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/127.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/128.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/129.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/130.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/131.tif.

Varol:1977:GPP

[Var77] Y. L. Varol. On the generation of the pseudo-remainder in polynomial division. The Computer Journal, 20(2):178–180, May 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/178.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/179.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/180.tif.

Various:1979:DEE

[Var79] Various. Discussion on “An entity/event data modelling method. The Computer Journal, 22(3):279–281, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/279.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/280.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/281.tif. See [Rob79].

Various:1980:DNA

[Var80a] Various. Discussion on ‘New approaches to systems analysis and design’. The Computer Journal, 23(1):28–33, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
01/tiff/28.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/
29.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_23/Issue_01/tiff/30.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
23/Issue_01/tiff/31.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
01/tiff/32.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/
33.tif.

Varol:1980:SAP

[Var80b] Y. L. Varol. On a subset of all the permutations of n marks. The Computer Journal, 23(4):344–346, November 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/344.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/345.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
04/tiff/346.tif.

Vaughan:1995:HPD

[Vau95] John G. Vaughan. A hierarchical protocol for decentralising information dissemination in distributed systems. The Computer Journal, 38(1):57–70, ???? 1995. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_38/Issue_01/
Vol38_01.body.html#AbstractVaughan.

vanBommel:1996:OOM

[vBFvdW96] P. van Bommel, P. J. M.

Frederiks, and P. van der Weide. Object-oriented modelling based on logbooks. The Computer Journal, 39 (9):793–799, ???? 1996. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www.oup.co.uk/computer_
journal/Volume_39/Issue_09/
Vol39_09.body.html#AbstractBommel.

Vin:1993:CMD

[VCB93] H. M. Vin, M.-S. Chen, and T. Barzilai. Collaboration management in DiCE. The Computer Journal, 36 (1):87–96, February 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_36/Issue_01/
Vol36_01.body.html#AbstractVin.

vandeVel:1970:NAP

[vdV70] H. van de Vel. A note on the automatic pretreatment of polynomials. The Computer Journal, 13(3):289–290, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_

03/130289.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/289.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/290.tif.

vanderWeide:1998:ISS

[vdWHvB98] Th. P. van der Weide, T. W. C. Huibers, and P. van Bommel. The incremental searcher satisfaction model for information retrieval. The Computer Journal, 41(5):311–318, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_41/Issue_05/
Vol41_05.body.html#AbstractVanDerWeide.

Velasevic:1982:RCG

[Vel82] Dusan M. Velasevic. Rightto-left code generation for arithmetic expressions. The Computer Journal, 25(3):316– 326, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
25/Issue_03/tiff/316.tif;
 http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/317.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/318.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/319.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_

03/tiff/320.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/321.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/322.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/323.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/324.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/325.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/326.tif.

Verkamo:1987:PQA

[Ver87] A. I. Verkamo. Performance of Quicksort adapted for virtual memory use. The Computer Journal, 30(4):362–371, August 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/362.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/363.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/364.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/365.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/366.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_30/Issue_
                                        04/tiff/367.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        04/tiff/368.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        04/tiff/369.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        04/tiff/370.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        04/tiff/371.tif.

Verrall:1991:UDI

[Ver91] M. S. Verrall. Unity doesn’t imply unification or overcoming heterogeneity problems in distributed software engineering environments. The Computer Journal, 34(6):522–533, December 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        06/tiff/522.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        06/tiff/523.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        06/tiff/524.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        06/tiff/525.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        06/tiff/526.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        06/tiff/527.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_34/Issue_
06/tiff/528.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/529.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/530.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/531.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/532.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
06/tiff/533.tif.

vanSinderen:1992:PDI

[vFV92] M. van Sinderen, L. Ferreira Pires, and C. A. Vissers. Protocol design and implementation using formal methods. The Computer Journal, 35(5):478–491, October 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_35/Issue_05/
Vol35_05.body.html#AbstractSinderen.

Varney:1973:SFO

[VG73] R. C. Varney and M. H. Gotterer. The structural foundation for an operating system. The Computer Journal, 16(4):357–359, November 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/160357.sgm.abs.html;

                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        04/tiff/357.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        04/tiff/358.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        04/tiff/359.tif.

Vovk:1999:CAP

[VG99] Vladimir Vovk and Alexander Gammerman. Complexity approximation principle. The Computer Journal, 42(4):318–322, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        04/420318.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        04/pdf/420318.pdf.

Vranesic:1972:TLP

[VH72] Z. G. Vranesic and V. C. Hamacher. Ternary logic in parallel multipliers. The Computer Journal, 15(3):254–258, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/150254.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/254.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/255.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_15/Issue_
03/tiff/256.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/257.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/258.tif. See corre-

spondence [BVH74].

Vee:1999:ACP

[VH99] Voon-Yee Vee and Wen-Jing Hsu. Applying Cilk in provably efficient task scheduling. The Computer Journal, 42(8):699–712, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
08/420699.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
08/pdf/420699.pdf.

Vine:1959:ACA

[Vin59] J. Vine. Application of a combination of analogue and digital computers to electron trajectory tracing. The Computer Journal, 2(3):134–144, October 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/020134.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/134.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_

03/tiff/135.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/136.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/137.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/138.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/139.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/140.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/141.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/142.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/143.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/144.tif.

Vince:1964:DSD

[Vin64]  P. H. Vince.  Documenta-

tion for systems design. The Computer Journal, 7(1):8–14, April 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
01/070008.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
01/tiff/10.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_01/tiff/

11.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_07/Issue_01/tiff/12.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
07/Issue_01/tiff/13.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
01/tiff/14.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_01/tiff/
8.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_07/Issue_01/tiff/9.
tif.

Vit:1975:RIE

[Vit75]  K. Vit.  Rational interpo-

lation and extrapolation for SUMT. The Computer Journal, 18(4):368–370, November 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/368.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/369.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/370.tif.

Vazirgiannis:1993:OOM

[VM93] M. Vazirgiannis and C. Mourlas. An object-oriented model for interactive multimedia presentations. The Computer Journal, 36(1):78–86, February 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

                                        journal/Volume_36/Issue_01/
                                        Vol36_01.body.html#AbstractVazirgiannis.

Vadera:1994:EFS

[VM94] Sunil Vadera and Farid Meziane. From English to formal specifications. The Computer Journal, 37(9):753–763, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_37/Issue_09/
                                        Vol37_09.body.html#AbstractVadera.

vanOosterom:1978:THT

[vO78] A. van Oosterom. Triangulating the human torso. The Computer Journal, 21(3):253–258, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        03/tiff/253.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        03/tiff/254.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        03/tiff/255.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        03/tiff/256.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        03/tiff/257.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        03/tiff/258.tif.

vonWright:1995:RHO

[von95] J. von Wright. Representing higher-order logic proofs in HOL. The Computer Journal, 38(2):171–179, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_38/Issue_02/
Vol38_02.body.html#AbstractWright.

Vallance:1980:BRP

[VP80] D. M. Vallance and Peter H.

Prowse. Book reviews: Programming in Standard Fortran 77, by A. Balfour and D. Marwick, 1979; Issues in Data Base Management, by Herbert Weber and Anthony I. Wasserman, 1979. The Computer Journal, 23(2):152, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/152.tif.

vanRijsbergen:1970:AFH

[vR70a] C. J. van Rijsbergen. Algorithm 52: A fast hierarchic clustering algorithm. The Computer Journal, 13(3):324–326, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/324.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/325.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/326.tif.

vanRijsbergen:1970:NAC

[vR70b] C. J. van Rijsbergen. Note on Algorithm 47: A clustering algorithm. The Computer Journal, 13(2):219, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        02/tiff/219.tif. See [van70,

Mux70, Sal70c].

Varol:1981:AGA

[VR81a] Y. L. Varol and D. Rotem. An algorithm to generate all topological sorting arrangements. The Computer Journal, 24(1):83–84, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        01/tiff/83.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_24/Issue_01/tiff/
                                        84.tif.

Varol:1981:SEJ

[VR81b] Yaakov L. Varol and Doron Rotem. Some experiences with the Johnson-Trotter permutation generation algorithm. The Computer Journal, 24(2):173– 176, May 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        24/Issue_02/tiff/173.tif;
                                         http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_

02/tiff/174.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/175.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
02/tiff/176.tif.

Vassallo:1992:ABS

[VR92] M. Vassallo and A. Ralston.

Algorithms for De Bruijn sequences — A case study in the empirical analysis of algorithms. The Computer Journal, 35(1):88–90, February 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_35/Issue_01/
Vol35_01.body.html#AbstractVassallo.

Verrijn-Stuart:1987:TTI

[VS87] A. Verrijn-Stuart. Themes and trends in information systems: TC8, 1975–1985. The Computer Journal, 30(2):97–109, April 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/100.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/101.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/102.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
02/tiff/103.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_

                                        02/tiff/104.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        02/tiff/105.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        02/tiff/106.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        02/tiff/107.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        02/tiff/108.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        02/tiff/109.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        02/tiff/97.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_30/Issue_02/tiff/
                                        98.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_30/Issue_02/tiff/99.
                                        tif.

vanSteen:1998:ADG

[vSHBT98] Maarten van Steen, Franz J. Hauck, Gerco Ballintijn, and Andrew S. Tanenbaum. Algorithmic design of the globe widearea location service. The Computer Journal, 41(5):297–310, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_41/Issue_05/
                                        Vol41_05.body.html#AbstractVanSteen.

vanLeeuwen:1987:IR

[VT87] J. Van Leeuwen and R. B. Tan. Interval routing. The Computer Journal, 30(4):298–307, August 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/298.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/299.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/300.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/301.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/302.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/303.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/304.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/305.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/306.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
04/tiff/307.tif.

vanderWeide:1992:UDS

[vtv92] Th. P. van der Weide, A. H. M.

ter Hofstede, and P. van Bommel. Uniquest: Determining the semantics of complex uniqueness constraints. The Computer Journal, 35(2):148–156, April 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_35/Issue_02/
                                        Vol35_02.body.html#AbstractWeide.

Varley:1994:RET

[VV94] D. A. Varley and M. Visvalingam. Road extraction and topographic data validation using area topology. The Computer Journal, 37(1):3–15, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_37/Issue_01/
                                        Vol37_01.body.html#AbstractVarley.

Visvalingam:1996:CAR

[VVW96] M. Visvalingam, D. Varley, and C. Wright. A cognitive approach to road recognition with novel feature indicators. The Computer Journal, 39(9):751–??, ???? 1996. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www.oup.co.uk/computer_
                                        journal/Volume_39/Issue_09/
                                        Vol39_09.body.html#AbstractVisvalingham;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_39/Issue_09/
                                        Vol39_09.body.html#AbstractVisvalingham.

vanLeeuwen:1993:IH

[vW93] J. van Leeuwen and D. Wood. Interval heaps. The Computer Journal, 36(3):209–216, June 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_

journal/Volume_36/Issue_03/
Vol36_03.body.html#AbstractLeeuwen.

Verdon:1995:CBU

[VW95] Frank P. Verdon and Mike Wells. Computing in British universities: the Computer Board 1966–1991. The Computer Journal, 38(10):822–830, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_38/Issue_10/
Vol38_10.body.html#AbstractVerdon.

Verco:1996:PMC

[VW96] K. L. Verco and M. J. Wise.

Plagiarism à la mode: a comparison of automated systems for detecting suspected plagiarism. The Computer Journal, 39(9):741–??, ???? 1996. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www.oup.co.uk/computer_
journal/Volume_39/Issue_09/
Vol39_09.body.html#AbstractVerco;
http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_09/
Vol39_09.body.html#AbstractVerco.

Vyugin:1999:ACS

[V’y99]  V. V. V’yugin.  Algorith-

mic complexity and stochastic properties of finite binary sequences. The Computer Journal, 42(4):294–317, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

journal/hdb/Volume_42/Issue_
04/420294.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
04/pdf/420294.pdf.

Wah:1991:SMM

[WA91] W. Wah and S. G. Akl. Simulating multiple memory accesses in logarithmic time and linear space. The Computer Journal, 34(6):A085–A088, December 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Waite:1970:BMP

[Wai70]  W. M. Waite.  Building

a mobile programming system. The Computer Journal, 13(1):28–31, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/130028.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/tiff/28.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_01/tiff/
29.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/30.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
13/Issue_01/tiff/31.tif.

Wakeman:1993:PVO

[Wak93]  Ian Wakeman.  Packetized

video: options for interaction between the user, the network and the codec. The Computer Journal, 36(1):55–67, February 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_01/
Vol36_01.body.html#AbstractWakeman.

Wallace:1964:CRE

[Wal64]  C. S. Wallace.  Correlated

round-off errors in digital integrating differential analyzers. The Computer Journal, 7(2):131–134, July 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/070131.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/131.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/132.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/133.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/134.tif.

Walwyn:1971:CHL

[Wal71] P. R. Walwyn. Correspondence:

On “High level languages for low level users”. The Computer Journal, 14(4):425, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_

                                        04/140425.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/425.tif. See [ER71]

and correspondence [Bac71, Fel72, MT72, GT72, Hil72, KTW^{+}73].

Walters:1976:BTC

[Wal76] R. E. Walters. Benchmark techniques: a constructive approach. The Computer Journal, 19(1):50–55, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        01/tiff/50.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_19/Issue_01/tiff/
                                        51.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_19/Issue_01/tiff/52.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        19/Issue_01/tiff/53.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        01/tiff/54.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_19/Issue_01/tiff/
                                        55.tif.

Wallach:1977:SAC

[Wal77] Y. Wallach. Scheduling algorithms for concurrent execution. The Computer Journal, 20(2):132–136, May 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_

02/tiff/132.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/133.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/134.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/135.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/136.tif.

Wallis:1982:PGP

[Wal82]  Peter J. L. Wallis.  The

preparation of guidelines for portable programming in highlevel languages. The Computer Journal, 25(3):375–378, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/375.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/376.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/377.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/378.tif.

Wallace:1998:ICS

[Wal98] C. S. Wallace. Intrinsic classification of spatially correlated data. The Computer Journal, 41(8):602–611, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_41/Issue_08/
                                        Vol41_08.body.html#AbstractWallace.

Wand:1980:DRA

[Wan80] I. C. Wand. Dynamic resource allocation and supervision with the programming language MODULA. The Computer Journal, 23(2):147–152, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        02/tiff/147.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        02/tiff/148.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        02/tiff/149.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        02/tiff/150.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        02/tiff/151.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        02/tiff/152.tif.

Warmington:1960:FYP

[War60]  C. B. Warmington.  The

first year’s production on a computer, and future plans. The Computer Journal, 3 (3):124–127, October 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        03/030124.sgm.abs.html;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/124.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/125.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/126.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/127.tif.

Ward:1974:VDM

[War74]  R. G. Ward.  A variable

delay method for improving recognition of parallel processable code in computer programs. The Computer Journal, 17(2):157–164, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/157.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/158.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/159.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/160.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/161.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/162.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/163.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        02/tiff/164.tif.

Ward:1986:PCC

[War86] R. K. Ward. Parity check codes for logic processors. The Computer Journal, 29 (1):12–16, February 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        01/tiff/12.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_29/Issue_01/tiff/
                                        13.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_29/Issue_01/tiff/14.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        29/Issue_01/tiff/15.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        01/tiff/16.tif.

Ward:1994:RET

[War94] M. P. Ward. Reverse engineering through formal transformation: Knuth’s “Polynomial Addition” algorithm. The Computer Journal, 37(9):795–813, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_37/Issue_09/
                                        Vol37_09.body.html#AbstractWard.

Ward:1996:PAF

[War96] M. P. Ward. Program analysis by formal transformation. The Computer Journal, 39(7):598– 618, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www.oup.co.uk/jnls/
list/comjnl/hdb/Volume_39/
Issue_07/390598.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_07/
Vol39_07.body.html#AbstractWard.

Watt:1959:NET

[Wat59] J. M. Watt. A note on the evaluation of trigonometric series. The Computer Journal, 1(4):162, January 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/010162.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
04/tiff/162.tif.

Watt:1963:RAP

[Wat63] J. M. Watt. The realization of ALGOL procedures and designational expressions. The Computer Journal, 5 (4):332–337, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/050332.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/332.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/333.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/334.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/335.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/336.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/337.tif.

Watt:1967:ASP

[Wat67] J. M. Watt. Algorithm 21: Solution of the polynomial equation using the method of Bairstow. The Computer Journal, 10(2):207–208, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/207.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
02/tiff/208.tif.

Watt:1968:CSD

[Wat68]  J. M. Watt.  Convergence

and stability of discretization methods for functional equations. The Computer Journal, 11(1):77–82, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/110077.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/77.tif;   http://

www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_01/tiff/
78.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_11/Issue_01/tiff/79.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/80.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
01/tiff/81.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_01/tiff/
82.tif.

Waters:1971:BSP

[Wat71] S. J. Waters. Blocking sequentially processed magnetic files. The Computer Journal, 14(2):109–112, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/140109.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/109.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/110.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/111.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/112.tif.

Waters:1972:EFD

[Wat72a] S. J. Waters. Erratum: “File design fallacies”. The Computer Journal, 15(3):258, August 1972. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/258.tif. See [Wat72b].

Waters:1972:FDF

[Wat72b] S. J. Waters. File design fallacies. The Computer Journal, 15(1):1–4, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        01/150001.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        01/tiff/1.tif;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_15/Issue_01/tiff/
                                        2.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_15/Issue_01/tiff/3.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        15/Issue_01/tiff/4.tif. See

erratum [Wat72a].

Watson:1973:EBA

[Wat73]  G. A. Watson.  On esti-

mating best approximations of functions defined by integral equations. The Computer Journal, 16(1):77–80, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        01/160077.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_

01/tiff/77.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_01/tiff/
78.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_16/Issue_01/tiff/79.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
16/Issue_01/tiff/80.tif.

Waters:1974:CAM

[Wat74a] S. J. Waters. Computer-aided methodology of computer systems design. The Computer Journal, 17(3):211–215, August 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/211.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/212.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/213.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/214.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
03/tiff/215.tif.

Waters:1974:MCS

[Wat74b]  S. J. Waters.  Methodol-

ogy of computer systems design. The Computer Journal, 17(1):17–24, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_

01/tiff/17.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/
18.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_17/Issue_01/tiff/19.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
17/Issue_01/tiff/20.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
01/tiff/21.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/
22.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_17/Issue_01/tiff/23.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
17/Issue_01/tiff/24.tif.

Waters:1975:ASD

[Wat75a]  S. J. Waters.  Analysis

of self-indexing, disc files. The Computer Journal, 18 (3):200–205, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/200.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/201.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/202.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/203.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/204.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/205.tif.

Waters:1975:EMD

[Wat75b] S. J. Waters. Estimating magnetic disc seeks. The Computer Journal, 18(1):12–17, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
01/tiff/12.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_01/tiff/
13.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_18/Issue_01/tiff/14.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
18/Issue_01/tiff/15.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
01/tiff/16.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_01/tiff/
17.tif.

Waters:1976:CPA

[Wat76a] S. J. Waters. CAM 01: a precedence analyser. The Computer Journal, 19(2):122–126, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/122.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
02/tiff/123.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_

                                        02/tiff/124.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/125.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        02/tiff/126.tif.

Waters:1976:HR

[Wat76b]  S. J. Waters.  Hit ra-

tios. The Computer Journal, 19(1):21–24, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_19/Issue_
                                        01/tiff/21.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_19/Issue_01/tiff/
                                        22.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_19/Issue_01/tiff/23.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        19/Issue_01/tiff/24.tif.

Waters:1977:CSP

[Wat77a] S. J. Waters. CAM02: a structured precedence analyser. The Computer Journal, 20(1):2–5, February 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        01/tiff/2.tif;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_20/Issue_01/tiff/
                                        3.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_20/Issue_01/tiff/4.
                                        tif;   http://www3.oup.co.

uk/computer_journal/hdb/Volume_
20/Issue_01/tiff/5.tif.

Waters:1977:MVM

[Wat77b] S. J. Waters. Majority verdicts in multi-processing — any two from three. The Computer Journal, 20(3):207–212, August 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/207.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/208.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/209.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/210.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/211.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/212.tif.

Waters:1979:TCS

[Wat79]  S. J. Waters.  Towards

comprehensive specifications. The Computer Journal, 22 (3):195–199, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/195.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/196.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        03/tiff/197.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        03/tiff/198.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        03/tiff/199.tif.

Watson:1981:CDT

[Wat81] D. F. Watson. Computing the n-dimensional Delaunay tessellation with application to Voronoĭ polytopes. The Computer Journal, 24(2):167–172, May 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        02/tiff/167.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        02/tiff/168.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        02/tiff/169.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        02/tiff/170.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        02/tiff/171.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        02/tiff/172.tif.

Watt:1982:CPI

[Wat82]  J. Watt.  Correspondence:

The possible impact of minicomputers and microprocessors [on] mainframe computer manufacturers. The Computer Journal, 25(2):286, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/286.tif. See [Hea81].

Watson:1991:EDS

[Wat91a]  David Watson.  Editorial:

Data structures. The Computer Journal, 34(5):385, October 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
05/tiff/385.tif.

Watson:1991:ECP

[Wat91b] Ian Watson. Editorial: Concurrent programming. The Computer Journal, 34(4):289, August 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/289.tif.

Watt:1991:MDP

[Wat91c] D. A. Watt. Modular description of programming languages. The Computer Journal, 34(6):A009–A028, December 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Watkins:2008:QAR

[Wat08] David S. Watkins. The QR algorithm revisited. SIAM Review, 50(1):133–145, ???? 2008. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See [Fra61, Fra62, Wat11].

Watkins:2011:FA

[Wat11] David S. Watkins. Francis’s algorithm. American Mathematical Monthly, 118(5):387–403, May 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/info/
10.4169/amer.math.monthly.

118.05.387. See [Fra61, Fra62, Wat08].

Wallace:1968:IMC

[WB68] C. S. Wallace and D. M. Boulton. An information measure for classification. The Computer Journal, 11(2):185–194, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/110185.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/185.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/186.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/187.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/188.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/189.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_

02/tiff/190.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/191.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/192.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/193.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/194.tif.

Wellings:1996:PRS

[WB96] A. J. Wellings and A. Burns.

Programming replicated systems in Ada 95. The Computer Journal, 39(5):361–373, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www.oup.co.uk/jnls/
list/comjnl/hdb/Volume_39/
Issue_05/390361.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_05/
Vol39_05.body.html#AbstractWellings.

Witter:1998:DLS

[WB98] Dian I. Witter and Michael W.

Berry. Downdating the latent semantic indexing model for conceptual information retrieval. The Computer Journal, 41(8):589–601, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_41/Issue_08/
Vol41_08.body.html#AbstractWitter.

Ward:1999:RRI

[WB99] Martin P. Ward and Keith H. Bennett. Recursion removal/ introduction by formal transformation: An aid to program development and program comprehension. The Computer Journal, 42(8):650–673, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
08/420650.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
08/pdf/420650.pdf.

Wetherell:1972:DKV

[WBB72] C. S. Wetherell, T. J. Buckholtz, and K. S. Booth. A director for Kriegspiel, a variant of chess. The Computer Journal, 15(1):66–70, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/150066.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/tiff/66.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_01/tiff/
67.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_15/Issue_01/tiff/68.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
15/Issue_01/tiff/69.tif;
http://www3.oup.co.uk/computer_

journal/hdb/Volume_15/Issue_
01/tiff/70.tif.

Witten:1994:SGM

[WBM^{+}94] Ian H. Witten, Timothy C.

Bell, Alistair Moffat, Craig G. Nevill-Manning, Tony C. Smith, and Harold Thimbleby. Semantic and generative models for lossy text compression. The Computer Journal, 37(2):83–87, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_37/Issue_02/
Vol37_02.body.html#AbstractWitten.

Williams:1985:RPP

[WC85] M. H. Williams and G. Chen.

Restructuring Pascal programs containing GOTO statements. The Computer Journal, 28(2): 134–137, May 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
28/Issue_02/tiff/134.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/135.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/136.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
02/tiff/137.tif.

Williams:1986:TPE

[WC86] M. H. Williams and G. Chen.

Translating Pascal for execution on a Prolog-based system. The Computer Journal, 29(3):246–252, June 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/246.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/247.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/248.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/249.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/250.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/251.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/252.tif.

Wu:1991:NNS

[WC91] J.-. L. Wu and S.-T. Cheng. A neural network simulating system. The Computer Journal, 34(1):92–93, February 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
01/tiff/92.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_01/tiff/
93.tif.

Wade:1994:DCG

[WC94] Greg A. Wade and Jiang-Hsing Chu. Drawability of complete graphs using a minimal slope set. The Computer Journal, 37(2):139–142, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_37/Issue_02/
Vol37_02.body.html#AbstractWade.

Wu:1998:DBG

[WC98] I-Chen Wu and Loon-Been Chen. On detection of bounded global predicates. The Computer Journal, 41(4):231–??, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_41/Issue_04/
Vol41_04.body.html#AbstractWu2.

Wiseman:1978:MGA

[WCH78] N. E. Wiseman, C. I. C. Campbell, and J. Harradine. On making graphic arts quality output by computer. The Computer Journal, 21(1):2–6, February 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
01/tiff/2.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_21/Issue_01/tiff/
3.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_21/Issue_01/tiff/4.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
21/Issue_01/tiff/5.tif;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
01/tiff/6.tif.

Williams:1971:GSD

[WCL71] W. T. Williams, H. T. Clifford, and G. N. Lance. Groupsize dependence: a rationale for choice between numerical classifications. The Computer Journal, 14(2):157–162, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/140157.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/157.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/158.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/159.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/160.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/161.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/162.tif.

Wells:1974:DEP

[WD74] M. Wells and A. Denson. Direct execution of programming languages. The Computer Journal, 17(2):130–134, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/130.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/131.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/132.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/133.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/134.tif.

Wichmann:1979:PCG

[WD79] B. A. Wichmann and J. Du Croz. A program to calculate the GAMM measure. The Computer Journal, 22 (4):317–322, November 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/317.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/318.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/319.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/320.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/321.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
04/tiff/322.tif.

Wise:1980:BRM

[WD80] A. H. Wise and A. Colin Day. Book reviews: Machine and Assembly Language Programming of the PDP-11, by Arthur Gill, 1979; Assembly Language Fundamentals, by Rina Yarmish and J. Yarmish, 1978. The Computer Journal, 23(2):164, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        02/tiff/164.tif.

Wallace:1999:MML

[WD99a] C. S. Wallace and D. L. Dowe. Minimum message length and Kolmogorov complexity. The Computer Journal, 42(4):270–283, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        04/420270.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        04/pdf/420270.pdf.

Wallace:1999:RMM

[WD99b] C. S. Wallace and D. L. Dowe. Refinements of MDL and MML coding. The Computer Journal, 42(4):330–337, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        04/420330.sgm.abs.html;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
04/pdf/420330.pdf.

Wallace:1999:R

[WD99c] C. S. Wallace and D. L.

Dowe. Rejoinder. The Computer Journal, 42(4):345–347, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
04/420345.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
04/pdf/420345.pdf.

Webb:1970:BRE

[Web70] M. J. H. Webb. Book review: The Economics of Computers, by William F. Sharpe, 1969. The Computer Journal, 13(2):222, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/222.tif.

Webb:1972:CPP

[Web72] M. H. J. Webb. Computer procurement policies for universities and similar users. The Computer Journal, 15(2):170–175, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/150170.sgm.abs.html;
http://www3.oup.co.uk/computer_

journal/hdb/Volume_15/Issue_
02/tiff/170.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/171.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/172.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/173.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/174.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/175.tif.

Wegbreit:1972:GCG

[Weg72] B. Wegbreit. A generalised compactifying garbage collector. The Computer Journal, 15(3):204–208, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/150204.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/204.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/205.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/206.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/207.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/208.tif.

Wegbreit:1974:PCE

[Weg74] Ben Wegbreit. Procedure closure in EL1. The Computer Journal, 17(1):38–43, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
01/tiff/38.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/
39.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_17/Issue_01/tiff/40.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
17/Issue_01/tiff/41.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
01/tiff/42.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/
43.tif.

Wegener:1993:SMX

[Weg93] I. Wegener. A simple modification of Xunrang and Yuzhang’s HEAPSORT variant improving its complexity significantly. The Computer Journal, 36(3):286–288, June 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_36/Issue_03/
Vol36_03.body.html#AbstractWegener.

Weiss:1991:ESE

[Wei91] M. A. Weiss. Empirical study of the expected running time of Shellsort. The Computer Journal, 34(1):88–91, February 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        01/tiff/88.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_34/Issue_01/tiff/
                                        89.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_34/Issue_01/tiff/90.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        34/Issue_01/tiff/91.tif.

Weiss:1993:FHB

[Wei93] M. A. Weiss. On finding the height of a binary search tree. The Computer Journal, 36(3):280–281, June 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_36/Issue_03/
                                        Vol36_03.body.html#AbstractWeiss.

Wells:1970:CTF

[Wel70] M. Wells. Correspondence: Towards Fortran VI? The Computer Journal, 13(1):120, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        01/tiff/120.tif. See [HL70].

Wells:1972:FCU

[Wel72] M. Wells. File compression using variable length encodings. The Computer Journal, 15(4):308–313, November 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/150308.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/308.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/309.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/310.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/311.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/312.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
04/tiff/313.tif. See corre-

spondence [GW73].

Weldon:1980:UDB

[Wel80] Jay Louise Weldon. Using data base abstractions for logical design. The Computer Journal, 23(1):41–45, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
01/tiff/41.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/
42.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_23/Issue_01/tiff/43.

                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        23/Issue_01/tiff/44.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_23/Issue_
                                        01/tiff/45.tif. See corre-

spondence [JL80, Ree80b].

Wensley:1959:CNA

[Wen59] J. H. Wensley. A class of non-analytical iterative processes. The Computer Journal, 1(4):163–167, January 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        04/010163.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        04/tiff/163.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        04/tiff/164.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        04/tiff/165.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        04/tiff/166.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        04/tiff/167.tif.

Westreich:1980:EPA

[Wes80a]  D. Westreich.  An effi-

cient predictor-corrector algorithm. The Computer Journal, 23(2):186, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_23/Issue_
02/tiff/186.tif. See erratum

[Ano80b] and correspondence [Sne80].

Westreich:1980:IPE

[Wes80b] D. Westreich. Improving polynomial evaluation at an approximate root. The Computer Journal, 23(2):187, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
02/tiff/187.tif. See corre-

spondence [Bre81, CF81].

Wetherfield:1966:TPM

[Wet66] M. R. Wetherfield. A technique for program monitoring by interruption. The Computer Journal, 9(2):161–166, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/090161.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/161.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/162.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/163.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/164.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_

02/tiff/165.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/166.tif.

Wettstein:1981:LOM

[Wet81] H. Wettstein. Locking operations for maximum concurrency. The Computer Journal, 24(3):243–248, August 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/243.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/244.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/245.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/246.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/247.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/248.tif.

Weyuker:1982:TNP

[Wey82]  Elaine J. Weyuker.  On

testing non-testable programs. The Computer Journal, 25 (4):465–470, November 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/465.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_25/Issue_
04/tiff/466.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/467.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/468.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/469.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
04/tiff/470.tif.

Williams:1978:DCA

[WF78] M. H. Williams and E. V. C.

Fielding. Discussion and correspondence: automatic generation of payroll programs. The Computer Journal, 21 (4):378–380, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/378.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/379.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/380.tif.

Wray:1989:NSL

[WF89] S. C. Wray and J. Fairbairn.

Non-strict languages — programming and implementation. The Computer Journal, 32(2):142–151, April 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/142.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/143.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/144.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/145.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/146.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/147.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/148.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/149.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/150.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
02/tiff/151.tif.

Wu:1998:OFS

[WFD98] Jie Wu, Eduardo B. Fernandex, and Donglai Dai. Optimal faultsecure scheduling. The Computer Journal, 41(4):207–??, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_41/Issue_04/
Vol41_04.body.html#AbstractWu1.

Wampler:1983:CI

[WG83] S. B. Wampler and R. E.

Griswold. Co-expressions in Icon. The Computer Journal, 26(1):72–78, February 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
01/tiff/72.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_01/tiff/
73.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_26/Issue_01/tiff/74.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
26/Issue_01/tiff/75.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
01/tiff/76.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_01/tiff/
77.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_26/Issue_01/tiff/78.
tif.

Wiseman:1968:RSP

[WH68] N. E. Wiseman and J. O. Hiles.

A ring structure processor for a small computer. The Computer Journal, 10(4):338–346, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/100338.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/338.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/tiff/339.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/tiff/340.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/tiff/341.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/tiff/342.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/tiff/343.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/tiff/344.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/tiff/345.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        04/tiff/346.tif.

Wade:1987:NAB

[WH87] S. Wade and J. Henstridge. Note on Algorithm 117: Buffered output in FORTRAN. The Computer Journal, 30(3):276, June 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        03/tiff/276.tif. See [HPB84].

Whale:1990:IPS

[Wha90] G. Whale. Identification of program similarity in large populations. The Computer Journal, 33(2):140–146, April 1990. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/140.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/141.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/142.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/143.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/144.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/145.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
02/tiff/146.tif.

Wheeler:1959:NRK

[Whe59]  D. J. Wheeler.  Note on

the Runge–Kutta method of integrating ordinary differential equations. The Computer Journal, 2(1):23, April 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
01/020023.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
01/tiff/23.tif.

Wood-Harper:1982:TCA

[WHF82] A. T. Wood-Harper and G. Fitzgerald. A taxonomy of current approaches to systems analysis. The Computer Journal, 25(1):12–16, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/12.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_25/Issue_01/tiff/
                                        13.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_25/Issue_01/tiff/14.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        25/Issue_01/tiff/15.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        01/tiff/16.tif.

Wood-Harper:1983:ALT

[WHF83] A. T. Wood-Harper and D. J. Flynn. Action learning for teaching information systems. The Computer Journal, 26 (1):79–82, February 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        01/tiff/79.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_26/Issue_01/tiff/
                                        80.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_26/Issue_01/tiff/81.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        26/Issue_01/tiff/82.tif.

Whiteley:1972:CSS

[Whi72]  E. Whiteley.  Correspon-

dence: On “Step size adjustment at discontinuities for fourth order Runge–Kutta methods”. The Computer Journal, 15(1):24, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/150024.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/tiff/24.tif. See [O’R70].

Wells:1971:EOS

[WHM71] M. Wells, D. Holdsworth, and A. P. McCann. The Eldon 2 operating system for KDF9. The Computer Journal, 14(1):21–24, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/140021.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/tiff/21.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_01/tiff/
22.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_14/Issue_01/tiff/23.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
14/Issue_01/tiff/24.tif.

Wong:1984:PSR

[WI84] F. S. Wong and M. R. Ito. Parallel sorting on a re-circulating systolic sorter. The Computer Journal, 27(3):260–269, August 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/260.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/261.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/262.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/263.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/264.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/265.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/266.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/267.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/268.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/269.tif.

Wichmann:1972:FAC

[Wic72] B. A. Wichmann. Five AL-GOL compilers. The Computer Journal, 15(1):8–12, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/150008.sgm.abs.html;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/tiff/10.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_01/tiff/
11.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_15/Issue_01/tiff/12.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
15/Issue_01/tiff/8.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
01/tiff/9.tif.

Wichmann:1982:CPA

[Wic82] B. A. Wichmann. A comparison of PASCAL and Ada. The Computer Journal, 25(2):248–252, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/248.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/249.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/250.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/251.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/252.tif.

Wichmann:1987:NAR

[Wic87] B. A. Wichmann. Note on Algorithm 121: RSA key calculation in Ada. The Computer Journal, 30(3):276, June 1987. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        03/tiff/276.tif. See [Hun85].

Wichmann:1989:TFS

[Wic89] B. A. Wichmann. Towards a formal specification of floating point. The Computer Journal, 32(5):432–436, October 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        05/tiff/432.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        05/tiff/433.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        05/tiff/434.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        05/tiff/435.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        05/tiff/436.tif.

Wichmann:1992:NUF

[Wic92] B. A. Wichmann. A note on the use of floating point in critical systems. The Computer Journal, 35(1):41–44, February 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_35/Issue_01/
                                        Vol35_01.body.html#AbstractWichmann.

Wiederhold:1979:TSS

[Wie79] J. F. A. Wiederhold. Tuning to service standards without monitors. The Computer Journal, 22(1):28–36, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
01/tiff/28.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_01/tiff/
29.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_22/Issue_01/tiff/30.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
22/Issue_01/tiff/31.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
01/tiff/32.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_01/tiff/
33.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_22/Issue_01/tiff/34.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
22/Issue_01/tiff/35.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
01/tiff/36.tif.

Wieringa:1995:CSD

[Wie95]  R. J. Wieringa.  Combin-

ing static and dynamic modelling methods: a comparison of four methods. The Computer Journal, 38(1):17–30, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_38/Issue_01/
                                        Vol38_01.body.html#AbstractWieringa.

Wight:1975:EAP

[Wig75] A. S. Wight. The EMAS archiving program. The Computer Journal, 18(2):131–134, May 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        02/tiff/131.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        02/tiff/132.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        02/tiff/133.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        02/tiff/134.tif.

Wigg:1981:CUB

[Wig81] J. D. Wigg. Correspondence: Unconditional branches. The Computer Journal, 24(2):192, May 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        02/tiff/192.tif. See [Tho80].

Wilkes:1958:SDC

[Wil58a]  M. V. Wilkes.  The sec-

ond decade of computer development. The Computer Journal, 1(3):98–105, October 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/010098.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/100.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/101.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/102.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/103.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/104.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/105.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/98.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_03/tiff/
99.tif.

Wilkinson:1958:CEM

[Wil58b] J. H. Wilkinson. The calculation of eigenvectors by the method of Lanczos. The Computer Journal, 1(3):148– 152, October 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
01/Issue_03/010148.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
01/Issue_03/tiff/148.tif;
 http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_

                                        03/tiff/149.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        03/tiff/150.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        03/tiff/151.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        03/tiff/152.tif.

Wilkinson:1958:CEC

[Wil58c] J. H. Wilkinson. The calculation of the eigenvectors of codiagonal matrices. The Computer Journal, 1(2):90–96, July 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        02/010090.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        02/tiff/90.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_01/Issue_02/tiff/
                                        91.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_01/Issue_02/tiff/92.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        01/Issue_02/tiff/93.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        02/tiff/94.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_01/Issue_02/tiff/
                                        95.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_01/Issue_02/tiff/96.
                                        tif.

Wilson:1959:SCL

[Wil59] L. B. Wilson. Solution of certain large sets of equations on Pegasus using matrix methods. The Computer Journal, 2(3):130–133, October 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/020130.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/130.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/131.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/132.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/133.tif.

Wilkinson:1960:HMS

[Wil60] J. H. Wilkinson. Householder’s method for the solution of the algebraic eigenproblem. The Computer Journal, 3(1):23–27, April 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
01/030023.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
01/tiff/23.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_01/tiff/
24.tif;   http://www3.oup.
co.uk/computer_journal/hdb/

Volume_03/Issue_01/tiff/25.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
03/Issue_01/tiff/26.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
01/tiff/27.tif.

Wilkes:1961:DTN

[Wil61a] M. V. Wilkes. Data transmission and the new outlook for the computer field. The Computer Journal, 4(1):1–9, April 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/040001.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/tiff/1.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_04/Issue_01/tiff/
2.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_04/Issue_01/tiff/3.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
04/Issue_01/tiff/4.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/tiff/5.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_04/Issue_01/tiff/
6.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_04/Issue_01/tiff/7.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
04/Issue_01/tiff/8.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/tiff/9.tif.

Wilkinson:1961:REB

[Wil61b] J. H. Wilkinson. Rigorous error bounds for computed eigensystems. The Computer Journal, 4(3):230–241, October 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/040230.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/230.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/231.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/232.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/233.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/234.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/235.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/236.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/237.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/238.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/239.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_

                                        03/tiff/240.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        03/tiff/241.tif.

Williams:1961:PFF

[Wil61c] M. B. Williams. Present and future facilities for data transmission. The Computer Journal, 4(2):88–95, July 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        02/040088.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        02/tiff/88.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_04/Issue_02/tiff/
                                        89.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_04/Issue_02/tiff/90.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        04/Issue_02/tiff/91.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        02/tiff/92.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_04/Issue_02/tiff/
                                        93.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_04/Issue_02/tiff/94.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        04/Issue_02/tiff/95.tif.

Wilkinson:1962:BRS

[Wil62a] J. H. Wilkinson. Book review: Solutions Numériques des Équations Algébraiques, by E. Durand (Paris: Masson et Cie., 1960, pp. 328, 65 NF). The Computer Journal, 5(1):32, April 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/
content/5/1/28.full.pdf.

Wilkinson:1962:IEM

[Wil62b] J. H. Wilkinson. Instability of the elimination method of reducing a matrix to tridiagonal form. The Computer Journal, 5(1):61–70, April 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
01/050061.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
01/tiff/61.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_01/tiff/
62.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_05/Issue_01/tiff/63.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
05/Issue_01/tiff/64.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
01/tiff/65.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_01/tiff/
66.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_05/Issue_01/tiff/67.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
05/Issue_01/tiff/68.tif;
http://www3.oup.co.uk/computer_

                                        journal/hdb/Volume_05/Issue_
                                        01/tiff/69.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_05/Issue_01/tiff/
                                        70.tif.

Wilkinson:1963:BRS

[Wil63a] J. H. Wilkinson. Book review: Solutions Numériques des Équations Algébraiques, Volume II, by E. Durand (Paris: Masson et Cie., 1961, pp. 445, 90 NF). The Computer Journal, 5(4):263, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.org/
                                        content/5/4/258.full.pdf.

Wilson:1963:ETA

[Wil63b] J. F. Wilson. Experience in transmitting accounting data. The Computer Journal, 5 (4):305–307, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        04/050305.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        04/tiff/305.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        04/tiff/306.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        04/tiff/307.tif.

Wilkes:1964:PUF

[Wil64a] Maurice V. Wilkes. A programmer’s utility filing system. The Computer Journal, 7(3):180–184, October 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/070180.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/180.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/181.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/182.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/183.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/184.tif.

Williams:1964:PES

[Wil64b] J. W. J. Williams. E.S.P.

the Elliott simulator package. The Computer Journal, 6(4):328–331, January 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/060328.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/328.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/329.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/330.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        04/tiff/331.tif.

Wilkes:1965:LWT

[Wil65a] M. V. Wilkes. Lists and why they are useful. The Computer Journal, 7(4):278–281, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        04/070278.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        04/tiff/278.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        04/tiff/279.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        04/tiff/280.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        04/tiff/281.tif.

Wilkinson:1965:CLP

[Wil65b] J. H. Wilkinson. The calculation of Lamé polynomials. The Computer Journal, 8(3):273–277, October 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        03/080273.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        03/tiff/273.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        03/tiff/274.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/275.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/276.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/277.tif.

Wilkinson:1965:CRA

[Wil65c]  J. H. Wilkinson.  Conver-

gence of the LR, QR and related algorithms. The Computer Journal, 8(1):77–84, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
01/080077.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
01/tiff/77.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_01/tiff/
78.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_08/Issue_01/tiff/79.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
08/Issue_01/tiff/80.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
01/tiff/81.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_01/tiff/
82.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_08/Issue_01/tiff/83.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
08/Issue_01/tiff/84.tif.

Wilkinson:1965:ARS

[Wil65d] J. H. Wilkinson. The QR algorithm for real symmetric matrices with multiple eigenvalues. The Computer Journal, 8(1):85–87, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        01/080085.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_08/Issue_
                                        01/tiff/85.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_08/Issue_01/tiff/
                                        86.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_08/Issue_01/tiff/87.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        08/Issue_01/tiff/88.tif.

Wilkes:1967:DMA

[Wil67a]  M. V. Wilkes.  The de-

sign of multiple-access computer systems. The Computer Journal, 10(1):1–9, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        01/100001.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        01/tiff/1.tif;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_10/Issue_01/tiff/
                                        2.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_10/Issue_01/tiff/3.

tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/4.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/5.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_10/Issue_01/tiff/
6.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_10/Issue_01/tiff/7.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/8.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/9.tif.

Wilkinson:1967:BRM

[Wil67b] J. H. Wilkinson. Book review: The Matrix Analysis of Vibration, by R. E. D. Bishop, G. M. L. Gladwell, and S. Michaelson, 1965; 404 pages (London: Cambridge University Press, 100s). The Computer Journal, 10(1):77, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/
content/10/1/74.full.pdf.

Willey:1967:DCE

[Wil67c]  E. L. Willey.  Discussion

and correspondence: an established U.K. software development. The Computer Journal, 10(3):313–314, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_

                                        03/100313.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        03/tiff/313.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        03/tiff/314.tif. See corre-

spondence [Tay68] and [TH64].

Wilkes:1968:OIS

[Wil68a] M. V. Wilkes. The outer and inner syntax of a programming language. The Computer Journal, 11(3):260–263, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        03/110260.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        03/tiff/260.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        03/tiff/261.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        03/tiff/262.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        03/tiff/263.tif.

Williams:1968:SLN

[Wil68b] H. P. Williams. The synthesis of logical nets consisting of NOR units. The Computer Journal, 11(2):173–176, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        02/110173.sgm.abs.html;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/173.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/174.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/175.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/176.tif.

Wilkinson:1970:BRC

[Wil70] J. H. Wilkinson. Book review: A Collection of Matrices for Testing Computational Algorithms, by Robert T. Gregory, D. L. Karney, 1969; 154 pages. (John Wiley and Sons Ltd., £4.75). The Computer Journal, 13(4):391, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/
content/13/4/387.full.pdf.

Wilkes:1972:PID

[Wil72]  M. V. Wilkes.  On pre-

serving the integrity of data bases. The Computer Journal, 15(3):191–194, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/150191.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/191.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_

                                        03/tiff/192.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/193.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        03/tiff/194.tif.

Wilkes:1973:DP

[Wil73] M. V. Wilkes. The dynamics of paging. The Computer Journal, 16(1):4–9, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        01/160004.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        01/tiff/4.tif;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_16/Issue_01/tiff/
                                        5.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_16/Issue_01/tiff/6.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        16/Issue_01/tiff/7.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        01/tiff/8.tif;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_16/Issue_01/tiff/
                                        9.tif.

Willers:1976:SDG

[Wil76a] I. M. Willers. The syntax directed graph algorithm for the input of equations to the Taylor series system for solving ordinary differential equations. The Computer Journal, 19(4):344–347, November 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/344.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/345.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/346.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/347.tif.

Williams:1976:DE

[Wil76b] M. R. Williams. The difference engines. The Computer Journal, 19(1):82–89, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
01/tiff/82.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_01/tiff/
83.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_19/Issue_01/tiff/84.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
19/Issue_01/tiff/85.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
01/tiff/86.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_01/tiff/
87.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_19/Issue_01/tiff/88.
tif;   http://www3.oup.co.

                                        uk/computer_journal/hdb/Volume_
                                        19/Issue_01/tiff/89.tif.

Wilkie:1977:TLE

[Wil77a] A. D. Wilkie. Table look-up (with examples in COBOL). The Computer Journal, 20 (3):202–206, August 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        03/tiff/202.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        03/tiff/203.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        03/tiff/204.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        03/tiff/205.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        03/tiff/206.tif.

Williams:1977:GSF

[Wil77b] M. H. Williams. Generating structured flow diagrams: the nature of unstructuredness. The Computer Journal, 20(1):45–50, February 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        01/tiff/45.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_20/Issue_01/tiff/
                                        46.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_20/Issue_01/tiff/47.

tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
20/Issue_01/tiff/48.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
01/tiff/49.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_01/tiff/
50.tif.

Willis:1977:RTH

[Wil77c] P. J. Willis. A real time hidden surface technique. The Computer Journal, 20(4):335– 339, November 1977. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
20/Issue_04/tiff/335.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/336.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/337.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/338.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/339.tif.

Wilson:1977:BSI

[Wil77d] J. M. Wilson. Boolean simplification and integer inequalities. The Computer Journal, 20(4):356–358, November 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/356.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/357.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/358.tif.

Williams:1978:SSF

[Wil78a] M. H. Williams. Static semantic features of Algol60 and Basic. The Computer Journal, 21(3):234–242, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/234.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/235.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/236.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/237.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/238.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/239.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/240.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/241.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/242.tif.

Williams:1978:BDH

[Wil78b] P. W. Williams. Balancing demand in a hierarchical structure. The Computer Journal, 21(4):341–346, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/341.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/342.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/343.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/344.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/345.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
04/tiff/346.tif.

Williams:1978:CCS

[Wil78c]  P. W. Williams.  Criteria

for choosing subsets to obtain maximum relative entropy. The Computer Journal, 21(1):57–62, February 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
01/tiff/57.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_01/tiff/
58.tif;   http://www3.oup.
co.uk/computer_journal/hdb/

                                        Volume_21/Issue_01/tiff/59.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        21/Issue_01/tiff/60.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        01/tiff/61.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_21/Issue_01/tiff/
                                        62.tif.

Williams:1978:RBE

[Wil78d] P. W. Williams. The relationship between entropy and file search times. The Computer Journal, 21(1):63–65, February 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        01/tiff/63.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_21/Issue_01/tiff/
                                        64.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_21/Issue_01/tiff/65.
                                        tif.

Williams:1979:CEO

[Wil79] M. H. Williams. Conditions for extended operator precedence parsing. The Computer Journal, 22(2):164–168, May 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/164.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/165.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_22/Issue_
02/tiff/166.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/167.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/168.tif.

Williams:1982:CDF

[Wil82a] M. H. Williams. A comment on the decomposition of flowchart schemata. The Computer Journal, 25(3):393–396, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/393.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/394.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/395.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/396.tif.

Williams:1982:PLB

[Wil82b] M. H. Williams. The programming language BPL. The Computer Journal, 25(3):289– 306, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
25/Issue_03/tiff/289.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/290.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_25/Issue_
03/tiff/291.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/292.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/293.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/294.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/295.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/296.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/297.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/298.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/299.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/300.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/301.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/302.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/303.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/304.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/305.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/306.tif.

Williams:1983:FSP

[Wil83a] M. H. Williams. Flowchart schemata and the problem of nomenclature. The Computer Journal, 26(3):270–276, August 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/270.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/271.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/272.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/273.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/274.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/275.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/276.tif.

Wilson:1983:AIP

[Wil83b]  J. M. Wilson.  Algorithm

114: Interrupted permutations in lexicographic order. The Computer Journal, 26(1):92, February 1983. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Wilkes:1984:SMP

[Wil84] Maurice V. Wilkes. Security management and protection — a personal approach. The Computer Journal, 27(1):3–7, February 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
01/tiff/3.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_27/Issue_01/tiff/
4.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_27/Issue_01/tiff/5.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
27/Issue_01/tiff/6.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
01/tiff/7.tif.

Williams:1985:LAC

[Wil85a] M. H. Williams. A linear algorithm for colouring planar graphs with five colours. The Computer Journal, 28 (1):78–81, February 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
01/tiff/78.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_01/tiff/
79.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_28/Issue_01/tiff/80.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
28/Issue_01/tiff/81.tif.

Williams:1985:STG

[Wil85b] M. H. Williams. Structuring two-level grammar specifications. The Computer Journal, 28(3):250–256, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/250.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/251.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/252.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/253.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/254.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/255.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
03/tiff/256.tif.

Wilson:1986:WMA

[Wil86] J. M. Wilson. A weighted markscaling algorithm. The Computer Journal, 29(5):477, October 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
05/tiff/477.tif.

Williams:1988:GQ

[Wil88]  R. Williams.  The Gob-

lin quadtree. The Computer Journal, 31(4):358–363, August 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/358.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/359.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/360.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/361.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/362.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
04/tiff/363.tif.

Wilson:1989:CCU

[Wil89]  J. M. Wilson.  Crossword

compilation using integer programming. The Computer Journal, 32(3):273–275, June 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/273.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/274.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/275.tif.

Wills:1990:FPS

[Wil90] H. Wills. Fundamentals of pricing and scheduling computer services and investment in computer equipment. The Computer Journal, 33(3):266–273, June 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/266.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/267.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/268.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/269.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/270.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/271.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/272.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
03/tiff/273.tif.

Wilkes:1991:MVC

[Wil91a] M. V. Wilkes. The maturing of VLSI — and its consequences for the computer industry. The Computer Journal, 34(3):193–194, June 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/193.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        03/tiff/194.tif.

Wilkinson:1991:CPO

[Wil91b] B. Wilkinson. Comparative performance of overlapping connectivity multiprocessor interconnection networks. The Computer Journal, 34(3):207–214, June 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        03/tiff/207.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        03/tiff/208.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        03/tiff/209.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        03/tiff/210.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        03/tiff/211.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        03/tiff/212.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        03/tiff/213.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        03/tiff/214.tif.

Windley:1959:ISA

[Win59a] P. F. Windley. The influence of storage access time on merging processes in a computer. The Computer Journal, 2(2):49–53, July 1959. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
02/020049.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
02/tiff/49.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_02/tiff/
50.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_02/Issue_02/tiff/51.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
02/Issue_02/tiff/52.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
02/tiff/53.tif.

Windley:1959:TMD

[Win59b] P. F. Windley. Transposing matrices in a digital computer. The Computer Journal, 2(1):47–48, April 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
01/020047.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
01/tiff/47.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_01/tiff/
48.tif.

Windley:1960:TFR

[Win60] P. F. Windley. Trees, forests and rearranging. The Computer Journal, 3(2):84–88, July 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        02/030084.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        02/tiff/84.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_03/Issue_02/tiff/
                                        85.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_03/Issue_02/tiff/86.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        03/Issue_02/tiff/87.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        02/tiff/88.tif.

Winrich:1969:NCE

[Win69]  L. B. Winrich.  Note on

a comparison of evaluation schemes for the interpolating polynomial. The Computer Journal, 12(2):154–155, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        02/120154.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        02/tiff/154.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        02/tiff/155.tif.

Winn:1975:DAE

[Win75]  G. C. E. Winn.  A dig-

ital approach to the efficient synthesis of threshold gates. The Computer Journal, 18(3):239–242, August 1975. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/239.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/240.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/241.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/242.tif.

Wiseman:1964:ALP

[Wis64] N. E. Wiseman. Application of list-processing methods to the design of interconnections for a fast logic system. The Computer Journal, 6(4):321–327, January 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/060321.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/321.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/322.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/323.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/324.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/325.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_

                                        04/tiff/326.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_06/Issue_
                                        04/tiff/327.tif.

Wiseman:1968:NCD

[Wis68] N. E. Wiseman. A note on compiling display file from a data structure. The Computer Journal, 11(2):141–147, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        02/110141.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        02/tiff/141.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        02/tiff/142.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        02/tiff/143.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        02/tiff/144.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        02/tiff/145.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        02/tiff/146.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        02/tiff/147.tif.

Witten:1980:AAL

[Wit80]  I. H. Witten.  Algorithms

for adaptive linear prediction. The Computer Journal, 23(1):78–84, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
01/tiff/78.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/
79.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_23/Issue_01/tiff/80.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
23/Issue_01/tiff/81.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
01/tiff/82.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/
83.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_23/Issue_01/tiff/84.
tif.

Woodward:1961:AL

[WJ61] P. M. Woodward and D. P.

Jenkins. Atoms and lists. The Computer Journal, 4(1):47–53, April 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/040047.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/tiff/47.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_01/tiff/
48.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_04/Issue_01/tiff/49.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
04/Issue_01/tiff/50.tif;

                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_04/Issue_
                                        01/tiff/51.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_04/Issue_01/tiff/
                                        52.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_04/Issue_01/tiff/53.
                                        tif.

Wagle:1974:CPM

[WJ74] B. Wagle and P. M. Jenkins. Corporate planning, models and computer systems. The Computer Journal, 17(3):194–200, August 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        03/tiff/194.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        03/tiff/195.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        03/tiff/196.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        03/tiff/197.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        03/tiff/198.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        03/tiff/199.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        03/tiff/200.tif.

Wang:1995:PAC

[WJW95] Hu Jun Wang, Yuan-Da Jiang, and Hui-Zhong Wu. Performance analysis of clusterbased PPMB multiprocessor systems. The Computer Journal, 38(5):355–364, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_38/Issue_05/
Vol38_05.body.html#AbstractWang.

Williams:1977:DRU

[WK77] P. W. Williams and M. T.

Khallaghi. Document retrieval using a substring index. The Computer Journal, 20(3):257–262, August 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/257.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/258.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/259.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/260.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/261.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
03/tiff/262.tif.

Windley:1960:DPU

[WKRJ60] P. F. Windley, L. R. Kay, and A. Rowland-Jones. Data processing in university administration. The Computer Journal, 3(1):15–20, April 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
01/030015.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
01/tiff/15.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_01/tiff/
16.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_03/Issue_01/tiff/17.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
03/Issue_01/tiff/18.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
01/tiff/19.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_01/tiff/
20.tif.

Willcox:1972:ACD

[WL72] W. R. Willcox and S. P. Lapage. Automatic construction of diagnostic tables. The Computer Journal, 15(3):263–267, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/150263.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/263.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/264.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_

03/tiff/265.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/266.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/267.tif.

Wey:1987:NSA

[WL87] C.-L. Wey and F. Lombardi.

On a novel self-test approach to digital testing. The Computer Journal, 30(3):258–267, June 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/258.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/259.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/260.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/261.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/262.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/263.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/264.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/265.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_
03/tiff/266.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_30/Issue_
                                        03/tiff/267.tif.

Wang:1996:BAC

[WL96] Hu-Jun Wang and Jian Li. Bandwidth analysis for A class of bus-based systems. The Computer Journal, 39(4):346–352, ???? 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://www.oup.co.uk/jnls/
                                        list/comjnl/hdb/Volume_39/
                                        Issue_04/390346.sgm.abs.html;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_39/Issue_04/
                                        Vol39_04.body.html#AbstractWang.

Williams:1971:CCC

[WLDC71] W. T. Williams, G. N. Lance, M. B. Dale, and H. T. Clifford. Controversy concerning the criteria for taxonometric strategies. The Computer Journal, 14(2):162–165, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/140162.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/162.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/163.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/164.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/165.tif.

Williams:1969:WLC

[WM69] P. E. Williams and D. C.

Menzies. The Windscale on-line computing system. The Computer Journal, 12 (3):244–250, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/120244.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/244.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/245.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/246.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/247.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/248.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/249.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
03/tiff/250.tif.

Watt:1983:EAG

[WM83] David A. Watt and Ole Lehrmann Madsen. Extended attribute grammars. The Computer Journal, 26(2):142–153, May 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

journal/hdb/Volume_26/Issue_
02/tiff/142.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/143.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/144.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/145.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/146.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/147.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/148.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/149.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/150.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/151.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/152.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
02/tiff/153.tif. See also: re-

port 10, Computer Science Department, University of Glasgow (July 1977) and report DAIMI PB-105, Computer Science Department, Aarhus University (November 1979).

Williams:1984:PAC

[WM84] M. H. Williams and K. T.

Milne. The performance of algorithms for colouring planar graphs. The Computer Journal, 27(2):165–170, May 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/165.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/166.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/167.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/168.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/169.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/170.tif.

Wilson:1995:IAC

[WM95] Stephen P. Wilson and John A.

McDermid. Integrated analysis of complex safety critical systems. The Computer Journal, 38(10):765–776, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_38/Issue_10/
Vol38_10.body.html#AbstractWilson.

Wilkes:1968:DMA

[WN68] M. V. Wilkes and R. M.

Needham. The design of multiple-access computer systems: part 2. The Computer Journal, 10(4):315–320, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/100315.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/315.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/316.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/317.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/318.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/319.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/320.tif.

Witten:1981:ILF

[WN81] I. H. Witten and Y. H. Ng. An ideographic language front end processor for accessing English language computer systems. The Computer Journal, 24(1):62–70, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
01/tiff/62.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_01/tiff/

63.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_24/Issue_01/tiff/64.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
24/Issue_01/tiff/65.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
01/tiff/66.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_01/tiff/
67.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_24/Issue_01/tiff/68.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
24/Issue_01/tiff/69.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
01/tiff/70.tif.

Wiseman:1986:CRF

[WN86] N. E. Wiseman and S. Nedunuri. Computing random fields. The Computer Journal, 29(4):373–377, August 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/373.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/374.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/375.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/376.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/377.tif.

Williams:1988:AHI

[WN88] M. H. Williams and K. A. Nicholson. An approach to handling incomplete information in databases. The Computer Journal, 31(2):133–140, April 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        02/tiff/133.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        02/tiff/134.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        02/tiff/135.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        02/tiff/136.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        02/tiff/137.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        02/tiff/138.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        02/tiff/139.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        02/tiff/140.tif.

Wheeler:1973:CWU

[WNT73] D. Wheeler, R. Needham, and The Editor. Correspondence: On word use. The Computer Journal, 16(1):18, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_

journal/hdb/Volume_16/Issue_
01/tiff/18.tif. See corre-

spondence [FT73].

Williams:1975:CMI

[WO75] M. H. Williams and H. L.

Ossher. A computer model for instructional purposes. The Computer Journal, 18 (4):333–341, November 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/333.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/334.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/335.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/336.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/337.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/338.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/339.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/340.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
04/tiff/341.tif.

Williams:1978:CUF

[WO78] M. H. Williams and H. L.

Ossher. Conversion of unstructured flow diagrams to structured form. The Computer Journal, 21(2):161–167, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        02/tiff/161.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        02/tiff/162.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        02/tiff/163.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        02/tiff/164.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        02/tiff/165.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        02/tiff/166.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        02/tiff/167.tif.

Wodon:1978:CNT

[Wod78] P. Wodon. Correspondence: On “A note on the Towers of Hanoi problem”. The Computer Journal, 21(2):191, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        02/tiff/191.tif. See [Hay77].

Wolfe:1969:NSN

[Wol69] M. A. Wolfe. The numerical solution of non-singular integral and integrodifferential equations by iteration with Chebyshev series. The Computer Journal, 12(2):193–196, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/120193.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/193.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/194.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/195.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/196.tif.

Wolfe:1971:NSI

[Wol71]  M. A. Wolfe.  The nu-

merical solution of implicit first order ordinary differential equations with initial conditions. The Computer Journal, 14(2):173–178, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/140173.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/173.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/174.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/175.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/176.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/177.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        02/tiff/178.tif.

Wolff:1978:RNL

[Wol78] J. G. Wolff. Recoding of natural language for economy of transmission or storage. The Computer Journal, 21(1):42–44, February 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_21/Issue_
                                        01/tiff/42.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_21/Issue_01/tiff/
                                        43.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_21/Issue_01/tiff/44.
                                        tif.

Wolff:1990:SPU

[Wol90] J. G. Wolff. Simplicity and power — some unifying ideas in computing. The Computer Journal, 33(6):518–534, December 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/518.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/519.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_33/Issue_
06/tiff/520.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/521.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/522.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/523.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/524.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/525.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/526.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/527.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/528.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/529.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/530.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/531.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/532.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/533.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/534.tif.

Wong:1999:SLS

[Won99] W. F. Wong. Source level static branch prediction. The Computer Journal, 42(2):142–149, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        02/420142.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        02/pdf/420142.pdf.

Woolner:1959:TPH

[Woo59] A. D. Woolner. Test programs for HEC. The Computer Journal, 2(1):44–47, April 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_02/Issue_
                                        01/020044.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_02/Issue_
                                        01/tiff/44.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_02/Issue_01/tiff/
                                        45.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_02/Issue_01/tiff/46.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        02/Issue_01/tiff/47.tif.

Woodger:1960:IA

[Woo60] M. Woodger. An introduction to ALGOL 60. The Computer Journal, 3(2):67–75, July 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
02/030067.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
02/tiff/67.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_02/tiff/
68.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_03/Issue_02/tiff/69.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
03/Issue_02/tiff/70.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
02/tiff/71.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_02/tiff/
72.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_03/Issue_02/tiff/73.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
03/Issue_02/tiff/74.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
02/tiff/75.tif.

Woodward:1966:SMS

[Woo66] P. M. Woodward. The synthesis of music and speech. The Computer Journal, 9 (3):257–262, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/090257.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
03/tiff/257.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_09/Issue_
                                        03/tiff/258.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        03/tiff/259.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        03/tiff/260.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        03/tiff/261.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        03/tiff/262.tif.

Wood:1968:SCU

[Woo68] D. C. Wood. A system for computing university examination timetables. The Computer Journal, 11(1):41–47, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        01/110041.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        01/tiff/41.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_11/Issue_01/tiff/
                                        42.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_11/Issue_01/tiff/43.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        11/Issue_01/tiff/44.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        01/tiff/45.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_11/Issue_01/tiff/
                                        46.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/

Volume_11/Issue_01/tiff/47.
tif.

Wood:1969:NFT

[Woo69a]  D. Wood.  The normal

form theorem — another proof. The Computer Journal, 12(2):139–147, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/120139.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/139.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/140.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/141.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/142.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/143.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/144.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/145.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/146.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
02/tiff/147.tif.

Wood:1969:TLF

[Woo69b]  D. Wood.  The theory of

left factored languages: Part

1. The Computer Journal, 12 (4):349–356, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/120349.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/349.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/350.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/351.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/352.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/353.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/354.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/355.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/356.tif.

Wood:1969:TCG

[Woo69c] D. C. Wood. A technique for colouring a graph applicable to large scale timetabling problems. The Computer Journal, 12(4):317–319, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

journal/hdb/Volume_12/Issue_
04/120317.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/317.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/318.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/319.tif.

Woodall:1969:ALR

[Woo69d] A. D. Woodall. Algorithm 43: A listed radix sort. The Computer Journal, 12(4):406, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/406.tif.

Wood:1970:GNF

[Woo70a] D. Wood. A generalised normal form theorem for contextfree grammars. The Computer Journal, 13(3):272–277, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/130272.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/272.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/273.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/274.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_13/Issue_
03/tiff/275.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/276.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/277.tif.

Wood:1970:TLF

[Woo70b]  D. Wood.  The theory of

left factored languages: Part

2. The Computer Journal, 13(1):55–62, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/130055.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/tiff/55.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_01/tiff/
56.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/57.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
13/Issue_01/tiff/58.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/tiff/59.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_01/tiff/
60.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/61.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
13/Issue_01/tiff/62.tif.

Woodall:1970:AIS

[Woo70c] A. D. Woodall. Algorithm 45: An internal sorting procedure using a two-way merge. The Computer Journal, 13 (1):110–111, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/tiff/110.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
01/tiff/111.tif.

Woodall:1970:AIM

[Woo70d] A. D. Woodall. Algorithm 55:

An internal merge sort giving ranks of items. The Computer Journal, 13(4):424–425, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/424.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/425.tif.

Woodall:1970:ADC

[Woo70e]  A. D. Woodall.  Algo-

rithm 56: To disentangle a chain. The Computer Journal, 13(4):425, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/425.tif.

Woodall:1970:NA

[Woo70f] A. D. Woodall. Note on Algorithm 43: A listed radix sort. The Computer Journal, 13(3):326, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/326.tif.

Woodall:1970:NAS

[Woo70g] A. D. Woodall. Note on Algorithms 25, 26. 25: Sort a section of the elements of an array by determining the rank of each element. 26: Order the subscripts of an array section according to the magnitudes of the elements. The Computer Journal, 13(3):326, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/326.tif. See [Boo67f,

Boo67g, Hil69].

Wood:1971:FNT

[Woo71a] D. Wood. A further note on top-down deterministic languages. The Computer Journal, 14(4):396–403, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/140396.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_

04/tiff/396.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/397.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/398.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/399.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/400.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/401.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/402.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/403.tif.

Woodall:1971:ART

[Woo71b] A. D. Woodall. Algorithm 63:

A recursive tree sort. The Computer Journal, 14(1):103, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
01/tiff/103.tif.

Woodall:1971:ANR

[Woo71c] A. D. Woodall. Algorithm 64: A non-recursive tree sort. The Computer Journal, 14(1):104, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_

journal/hdb/Volume_14/Issue_
01/tiff/104.tif.

Wood:1977:TVN

[Woo77a] D. Wood. Two variations on N-parallel finite state generators. The Computer Journal, 20(1):63–67, February 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
01/tiff/63.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_01/tiff/
64.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_20/Issue_01/tiff/65.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
20/Issue_01/tiff/66.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
01/tiff/67.tif.

Woodall:1977:GPS

[Woo77b]  A. D. Woodall.  Genera-

tion of permutation sequences. The Computer Journal, 20 (4):346–349, November 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/346.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/347.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
04/tiff/348.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_20/Issue_
04/tiff/349.tif.

Woodwark:1982:EQT

[Woo82] J. R. Woodwark. The explicit quad tree as a structure for computer graphics. The Computer Journal, 25(2):235–238, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/235.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/236.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/237.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/238.tif.

Woodwark:1984:CQT

[Woo84] J. R. Woodwark. Compressed quad trees. The Computer Journal, 27(3):225–229, August 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/225.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/226.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/227.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/228.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_27/Issue_
                                        03/tiff/229.tif.

Woodall:1986:TPD

[Woo86] A. D. Woodall. Techniques of program design. The Computer Journal, 29(5):434–435, October 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        05/tiff/434.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        05/tiff/435.tif.

Woodcock:1992:RAR

[Woo92]  J. C. P. Woodcock.  The

rudiments of algorithm refinement. The Computer Journal, 35(5):441–450, October 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_35/Issue_05/
                                        Vol35_05.body.html#AbstractWoodcock.

See [Coo92a].

Worboys:1994:UMS

[Wor94] Michael F. Worboys. A unified model for spatial and temporal information. The Computer Journal, 37(1):26–34, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/Volume_37/Issue_01/
                                        Vol37_01.body.html#AbstractWorboys.

Welsh:1967:UBC

[WP67] D. J. A. Welsh and M. B.

Powell. An upper bound for the chromatic number of a graph and its application to timetabling problems. The Computer Journal, 10(1):85–86, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/100085.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/85.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_01/tiff/

86.tif. See comment [Col67].

Wolfe:1968:SMS

[WP68] M. A. Wolfe and G. M. Phillips.

Some methods for the solution of non-singular Volterra integro-differential equations. The Computer Journal, 11 (3):334–336, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/110334.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/334.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/335.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/336.tif.

Whitby:1969:COB

[WP69] W. P. L. Whitby and R. M. Paine. Correspondence: OCR

— benefits and pitfalls. The Computer Journal, 12(4):410– 411, November 1969. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        12/Issue_04/tiff/410.tif;
                                         http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/411.tif. See [Pai69].

White:1975:SLT

[WP75] J. R. White and L. Presser. A structured language for translator construction. The Computer Journal, 18(1):34–42, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        01/tiff/34.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_18/Issue_01/tiff/
                                        35.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_18/Issue_01/tiff/36.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        18/Issue_01/tiff/37.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        01/tiff/38.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_18/Issue_01/tiff/
                                        39.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_18/Issue_01/tiff/40.
                                        tif;   http://www3.oup.co.

uk/computer_journal/hdb/Volume_
18/Issue_01/tiff/41.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
01/tiff/42.tif.

Wallace:1967:CMA

[WR67] C. S. Wallace and B. G.

Rowswell. Competition for memory access in the KDF9. The Computer Journal, 10(1): 64–68, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
10/Issue_01/100064.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/64.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/65.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_01/tiff/
66.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_10/Issue_01/tiff/67.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/68.tif.

Williams:1972:CLA

[WR72] M. R. Williams and B. Randell. Correspondence: On “Ludgate’s analytical machine of 1909’. The Computer Journal, 15(2):131, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/131.tif. See [Ran71].

From The Times (London, Saturday, 21 October, 1871), under the heading ‘Deaths’: ‘On the 18th Oct, at 1, Dorset-street Manchester Square, Charles Babbage, Esq. in the 80th year of his age.’.

Williams:1986:SID

[WR86] M. H. Williams and J. Riddall. The segregation of incompatible duties. The Computer Journal, 29(1):9–11, February 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_29/Issue_
                                        01/tiff/10.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_29/Issue_01/tiff/
                                        11.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_29/Issue_01/tiff/9.
                                        tif.

Wragg:1966:ULM

[Wra66] A. Wragg. The use of Lanczos τ-methods in the numerical solution of a Stefan problem. The Computer Journal, 9(1):106–109, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        01/090106.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        01/tiff/106.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_09/Issue_
                                        01/tiff/107.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/108.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/109.tif.

Wren:1971:BRC

[Wre71]  Anthony Wren.  Book re-

view: Case Exercises in Operations Research, by J. J. C. Martin and R. A. Denison, 1971. The Computer Journal, 14(4):377, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/377.tif.

Wright:1959:BAD

[Wri59] A. G. Wright. A business application of a digital computer. The Computer Journal, 2(3):103–104, October 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/020103.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/103.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
03/tiff/104.tif.

Wright:1960:MLI

[Wri60] M. A. Wright. Mechanizing a large index. The Computer Journal, 3(2):76–83, July 1960. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
02/030076.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
02/tiff/76.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_02/tiff/
77.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_03/Issue_02/tiff/78.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
03/Issue_02/tiff/79.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
02/tiff/80.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_02/tiff/
81.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_03/Issue_02/tiff/82.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
03/Issue_02/tiff/83.tif.

Wright:1961:DCT

[Wri61a] E. P. G. Wright. Data collection and transmission. The Computer Journal, 4(2):103–109, July 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/040103.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/103.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/104.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/105.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/106.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/107.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/108.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/109.tif.

Wright:1961:MII

[Wri61b] M. A. Wright. Matching inquiries to an index. The Computer Journal, 4(1):38–41, April 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/040036.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/tiff/38.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_01/tiff/
39.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_04/Issue_01/tiff/40.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
04/Issue_01/tiff/41.tif.

Wrigley:1963:AJM

[Wri63] H. E. Wrigley. Accelerating the Jacobi method for solving simultaneous equations by Chebyshev extrapolation when the eigenvalues of the iteration matrix are complex. The Computer Journal, 6(2):169–176, July 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/060169.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/169.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/170.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/171.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/172.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/173.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/174.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/175.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
02/tiff/176.tif.

Wright:1964:CCM

[Wri64] K. Wright. Chebyshev collocation methods for ordinary differential equations. The Computer Journal, 6(4):358–365, January 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/060358.sgm.abs.html;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/358.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/359.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/360.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/361.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/362.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/363.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/364.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
04/tiff/365.tif.

Wright:1966:CEC

[Wri66a] K. Wright. Correspondence: On “Error curves for Lanczos ‘selected points’ method”. The Computer Journal, 9(1):115, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
01/tiff/115.tif. See [Kiz66].

Wright:1966:SMI

[Wri66b] K. Wright. Series methods for integration. The Computer Journal, 9(2):191–199, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/090191.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/191.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/192.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/193.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/194.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/195.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/196.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/197.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/198.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_
02/tiff/199.tif.

Wright:1974:ASE

[Wri74] W. E. Wright. An axiomatic specification of Euclidean analysis. The Computer Journal, 17(4):355–364, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/355.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_

04/tiff/356.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/357.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/358.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/359.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/360.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/361.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/362.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/363.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/364.tif.

Wright:1982:SVD

[Wri82] William E. Wright. Single versus double buffering in constrained merging. The Computer Journal, 25(2):227–230, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/227.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/228.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/229.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/230.tif.

Wright:1983:SFS

[Wri83]  William E. Wright.  Some

file structure considerations pertaining to magnetic bubble memory. The Computer Journal, 26(1):43–51, February 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        01/tiff/43.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_26/Issue_01/tiff/
                                        44.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_26/Issue_01/tiff/45.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        26/Issue_01/tiff/46.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        01/tiff/47.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_26/Issue_01/tiff/
                                        48.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_26/Issue_01/tiff/49.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        26/Issue_01/tiff/50.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_26/Issue_
                                        01/tiff/51.tif.

Walsh:1973:BRC

[WS73] Joan Walsh and P. A. Samet. Book reviews: Computer Applications of Numerical Methods, by Shan S. Kuo, 1972; Introduction to Computational Methods for Students of Calculus, by S. S. McNeary, 1972. The Computer Journal, 16(3):261, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/261.tif.

Whitby-Strevens:1974:CMS

[WS74]  C. Whitby-Strevens.  Cor-

respondence: Microprogramming and systems architecture. The Computer Journal, 17(4):381, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
04/tiff/381.tif. See [Bro74a].

Wong:1986:SRD

[WS86] E. Wong and W. B. Samson.

The specification of a relational database (PRECI) as an abstract data type and its realisation in HOPE. The Computer Journal, 29(3):261–268, June 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/261.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/262.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/263.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_

03/tiff/264.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/265.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/266.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/267.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
03/tiff/268.tif.

Wilkes:1984:RW

[WSG^{+}84] A. J. Wilkes, D. W. Singer, J. J. Gibbons, T. R. King, P. Robinson, and N. E. Wiseman. The Rainbow Workstation. The Computer Journal, 27(2):112–120, May 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/112.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/113.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/114.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/115.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/116.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/117.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_

02/tiff/118.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/119.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/120.tif.

Walker:1979:AAM

[WT79] A. N. Walker and The Editor. Algorithms: On “Algorithm 1: Mintree”. The Computer Journal, 22(3):288, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/288.tif. See [Obr78a].

Ward:1984:ECD

[WT84] R. K. Ward and M. Tabandeh. Error correction and detection, a geometric approach. The Computer Journal, 27(3):246–253, August 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/246.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/247.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/248.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/249.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
03/tiff/250.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        03/tiff/251.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        03/tiff/252.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_27/Issue_
                                        03/tiff/253.tif.

Wicovsky:1991:IRH

[WTW91] K. Wicovsky, P. F. Thanisch, and M. H. Williams. Improved recursion handling through integrity constraints. The Computer Journal, 34(3):282–285, June 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        03/tiff/282.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        03/tiff/283.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        03/tiff/284.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        03/tiff/285.tif.

Wu:1996:BDR

[Wu96]  Xindong Wu.  A Bayesian

discretizer for real-valued attributes. The Computer Journal, 39(8):688–??, ???? 1996. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://www.oup.co.uk/jnls/
                                        list/comjnl/hdb/Volume_39/
                                        Issue_08/390688.sgm.abs.html;
                                         http://www3.oup.co.uk/computer_

journal/Volume_39/Issue_08/
Vol39_08.body.html#AbstractWu.

Wu:1997:HIN

[Wu97] X. Wu. Hybrid interpretation of ‘no match’ and ‘multiple match’ in induction. The Computer Journal, 40(1):50–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www.oup.co.uk/computer_
journal/Volume_40/Issue_01/
Vol40_01.body.html#AbstractWu;
http://www3.oup.co.uk/computer_
journal/Volume_40/Issue_01/
Vol40_01.body.html#AbstractWu.

Waliuzzaman:1970:DMV

[WV70] K. M. Waliuzzaman and Z. G.

Vranesic. On decomposition of multi-valued switching functions. The Computer Journal, 13(4):359–362, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/130359.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/359.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/360.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/361.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/362.tif.

Whitfield:1973:EEM

[WW73] H. Whitfield and A. S. Wight. EMAS — the Edinburgh Multi-Access System. The Computer Journal, 16(4):331–346, November 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/160331.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/331.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/332.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/333.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/334.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/335.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/336.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/337.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/338.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/339.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/340.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_

04/tiff/341.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/342.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/343.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/344.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/345.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/346.tif.

Walker:1976:LBB

[WW76] A. Walker and D. Wood. Locally balanced binary trees. The Computer Journal, 19 (4):322–325, November 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/322.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/323.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/324.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
04/tiff/325.tif.

Williams:1979:CAA

[WW79] P. W. Williams and D. Woodhead. Computer assisted analysis of cryptic crosswords. The Computer Journal, 22(1):67–70, February 1979.

CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
01/tiff/67.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_01/tiff/
68.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_22/Issue_01/tiff/69.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
22/Issue_01/tiff/70.tif.

Wilkes:1982:SCS

[WW82] A. J. Wilkes and N. E. Wiseman. A soft-edged character set and its derivation. The Computer Journal, 25 (1):140–147, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/140.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/141.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/142.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/143.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/144.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/145.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_

01/tiff/146.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
01/tiff/147.tif.

Woods:1983:MAM

[WW83] J. V. Woods and A. J. T.

Wheen. MU6P: an advanced microprocessor architecture. The Computer Journal, 26(3):208–217, August 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/208.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/209.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/210.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/211.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/212.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/213.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/214.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/215.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/216.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_
03/tiff/217.tif.

Waugh:1990:DSE

[WWK^{+}90] K. G. Waugh, M. H. Williams, Q. Kong, S. Salvini, and G. Chen. Designing SQUIR-REL: an extended SQL for a deductive database system. The Computer Journal, 33 (6):535–546, December 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/535.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/536.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/537.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/538.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/539.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/540.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/541.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/542.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/543.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        06/tiff/544.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_

06/tiff/545.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_
06/tiff/546.tif.

Wang:1991:SGM

[WWN91] C. J. Wang, C. H. Wu, and V. P. Nelson. A study of the generalized multiple busconnected parallel computer. The Computer Journal, 34(6): A089–A094, December 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Watt:1962:ASL

[WY62] J. M. Watt and A. Young. An attempt to simulate the liver on a computer. The Computer Journal, 5(3):221–227, October 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/050221.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/221.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/222.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/223.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/224.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
03/tiff/225.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_

                                        03/tiff/226.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        03/tiff/227.tif.

Watt:1971:BRC

[WY71] J. M. Watt and J. Z. Young. Book reviews: Computation and Theory in Ordinary Differential Equations, by James W. Daniel and Ramon E. Moore, 1970; Progress in Brain Research, Vol. 33. Computers and Brains, by J. P. Schade and J. Smith, 1970. The Computer Journal, 14(4):412, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        04/tiff/412.tif.

Wyman:1973:BPD

[Wym73] F. P. Wyman. Binary programming: a decision rule for selecting optimal vs heuristic techniques. The Computer Journal, 16(2):135–140, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/160135.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/135.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/136.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/137.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/138.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/139.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/140.tif. See corre-

spondence [PW74b].

Wynn:1971:NGE

[Wyn71] P. Wynn. A note on the generalised Euler transformation. The Computer Journal, 14(4):437–441, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/140437.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/437.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/438.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/439.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/440.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/441.tif. See corre-

spondence [Scr72].

Williams:1999:CIF

[WZ99] Hugh E. Williams and Justin Zobel. Compressing integers for fast file access. The Computer Journal, 42(3):193–201, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        03/420193.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        03/pdf/420193.pdf.

Xiang:1997:GOE

[XTU97] L. Xiang, C. Tang, and K. Ushijima. Grammaroriented enumeration of binary trees. The Computer Journal, 40(5):278–??, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www.oup.co.uk/computer_
                                        journal/Volume_40/Issue_05/
                                        Vol40_05.body.html#AbstractXiang;
                                         http://www3.oup.co.uk/computer_
                                        journal/Volume_40/Issue_05/
                                        Vol40_05.body.html#AbstractXiang.

Xunrang:1990:NHA

[XY90] G. Xunrang and Z. Yuzhang. A new HEAPSORT algorithm and the analysis of its complexity. The Computer Journal, 33(3):281–282, June 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        03/tiff/281.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_33/Issue_
                                        03/tiff/282.tif.

Young:1961:PDS

[YA61] A. Young and G. Almond. Predicting distributions of staff. The Computer Journal, 3 (4):246–250, January 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/030246.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/246.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/247.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/248.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/249.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/250.tif.

Yannakoudakis:1992:CNL

[YAB92] E. J. Yannakoudakis and H. A.

Attar-Bashi. Constructing natural language sentences from database tuples. The Computer Journal, 35(5):A445–??, October 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_35/Issue_05/
35000A445.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/Volume_35/Issue_05/
tiff/A445.tif.

Yannokoudakis:1980:CCB

[YAH80] E. J. Yannokoudakis, F. H. Ayres, and J. A. W. Huggill. Character coding for bibliographical record control. The Computer Journal, 23(1):53–60, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
01/tiff/53.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/
54.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_23/Issue_01/tiff/55.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
23/Issue_01/tiff/56.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
01/tiff/57.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/
58.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_23/Issue_01/tiff/59.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
23/Issue_01/tiff/60.tif.

Yates:1962:CRP

[Yat62] F. Yates. Computers in research — promise and performance. The Computer Journal, 4(4):273–279, January 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_

04/040273.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/273.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/274.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/275.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/276.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/277.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/278.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
04/tiff/279.tif.

Yang:1992:MSF

[YCC92] J. S. H. Yang, Y. H. Chin, and C. G. Chung. Many-sorted first-order logic database language. The Computer Journal, 35(2):129–137, April 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_35/Issue_02/
Vol35_02.body.html#AbstractYang.

Yelland:1989:FST

[Yel89] P. M. Yelland. First steps towards fully abstract semantics for object-oriented languages. The Computer Journal, 32(4):290–296, August 1989. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/290.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/291.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/292.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/293.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/294.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/295.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
04/tiff/296.tif.

Young:1971:BRC

[YFM71] A. Young, L. Fox, and D. F. Mayers. Book reviews: Convexity and Optimization in Finite Dimensions I, by Josef Stoer and Christoph Witzgall, 1970; On Round-off Errors in Linear Programming, by H. Müller-Merbach, 1970; Approximation Theory, edited by A. Talbot, 1970. The Computer Journal, 14(4):390, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/390.tif.

Yau:1993:PLD

[YG93] S. S. Yau and J. V. Gore.

Paradigm lost: discovering the intersection of assertionguided program construction and constraint-driven re-evaluation. The Computer Journal, 36(2): 137–142, April 1993. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/Volume_
36/Issue_02/Vol36_02.body.
html#AbstractYau.

Yates:1963:SAA

[YGS63] F. Yates, J. C. Gower, and H. R. Simpson. A specialized autocode for the analysis of replicated experiments. The Computer Journal, 5 (4):313–319, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/050313.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/313.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/314.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/315.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/316.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/317.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        04/tiff/318.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_05/Issue_
                                        04/tiff/319.tif.

Yue:1991:ESS

[YJ91] K-B. Yue and R. T. Jacob. An efficient starvationfree semaphore solution for the graphical mutual exclusion problem. The Computer Journal, 34(4):345–349, August 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        04/tiff/345.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        04/tiff/346.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        04/tiff/347.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        04/tiff/348.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_34/Issue_
                                        04/tiff/349.tif.

Yue:1994:OAR

[YJ94] Kwok-Bun Yue and R. T. Jacob. An optimal algorithm for reducing edge-solvable mutual exclusion graphs. The Computer Journal, 37(2):129–138, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_

journal/Volume_37/Issue_02/
Vol37_02.body.html#AbstractYue.

Yokoo:1997:DCU

[Yok97]  H. Yokoo.  Data compres-

sion using a sort-based similarity measure. The Computer Journal, 40(2/3):94–102, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_40/Issue_02/
Vol40_02.body.html#AbstractYokoo;
http://www3.oup.co.uk/computer_
journal/Volume_40/Issue_02/
Vol40_03.body.html#AbstractYokoo.

Youle:1960:SFS

[You60] P. V. Youle. Simulation of full-scale multi-stage batchwise chemical plant. The Computer Journal, 3(3):150–157, October 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/030150.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/150.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/151.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/152.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/153.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_

03/tiff/154.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/155.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/156.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
03/tiff/157.tif.

Young:1963:IUE

[You63]  A. Young.  The impact

on universities of the expansion in their computer facilities. The Computer Journal, 5(4):294–296, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/050294.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/294.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/295.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
04/tiff/296.tif.

Young:1986:UEF

[YP86] S. J. Young and C. Proctor. UFL: an experimental frame language based on abstract data types. The Computer Journal, 29(4):340–347, August 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_

04/tiff/340.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/341.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/342.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/343.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/344.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/345.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/346.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_
04/tiff/347.tif.

Yates:1958:UEC

[YR58] F. Yates and D. H. Rees.

The use of an electronic computer in research statistics: four years’ experience. The Computer Journal, 1(2):49–58, July 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
02/010049.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
02/tiff/49.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_02/tiff/
50.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_01/Issue_02/tiff/51.
tif;   http://www3.oup.co.

                                        uk/computer_journal/hdb/Volume_
                                        01/Issue_02/tiff/52.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        02/tiff/53.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_01/Issue_02/tiff/
                                        54.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_01/Issue_02/tiff/55.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        01/Issue_02/tiff/56.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_01/Issue_
                                        02/tiff/57.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_01/Issue_02/tiff/
                                        58.tif.

Yates:1960:GPA

[YS60] F. Yates and H. R. Simpson. A general program for the analysis of surveys. The Computer Journal, 3(3):136–140, October 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        03/030136.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        03/tiff/136.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        03/tiff/137.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        03/tiff/138.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_03/Issue_
                                        03/tiff/139.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_03/Issue_
03/tiff/140.tif.

Yates:1961:ASP

[YS61] F. Yates and H. R. Simpson. The analysis of surveys: processing and printed the basic tables. The Computer Journal, 4(1):20–24, April 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/040020.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/tiff/20.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_01/tiff/
21.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_04/Issue_01/tiff/22.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
04/Issue_01/tiff/23.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/tiff/24.tif.

Yannakoudakis:1991:CEM

[YS91] E. J. Yannakoudakis and A. A.

Shahid. A controlled environment for meta-message processing. The Computer Journal, 34(4):331–344, August 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/331.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_

04/tiff/332.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/333.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/334.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/335.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/336.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/337.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/338.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/339.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/340.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/341.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/342.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/343.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
04/tiff/344.tif.

Yau:1979:GST

[YT79] S. S. Yau and Y. S. Tang. Generation of shortest test sequences for detecting individual faults of sequential circuits. The Computer Journal, 22(2):169–172, May 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/169.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/170.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/171.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/172.tif.

Yule:1968:EHA

[Yul68] A. P. Yule. Extensions to the heuristic algorithm for University timetables. The Computer Journal, 10(4):360–364, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/100360.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/360.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/361.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/362.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/363.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/364.tif.

Yannakoudakis:1982:QGG

[YW82] E. J. Yannakoudakis and A. K. P. Wu. Quasiequifrequent group generation and evaluation. The Computer Journal, 25(2):183–187, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/183.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/184.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/185.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/186.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_25/Issue_
                                        02/tiff/187.tif.

Yu:1989:DEB

[YY89] K. W. Yu and T. L. Yu. Data encryption based upon time reversal transformations. The Computer Journal, 32(3):241– 245, June 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        32/Issue_03/tiff/241.tif;
                                         http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        03/tiff/242.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_32/Issue_
                                        03/tiff/243.tif;   http:/
                                        /www3.oup.co.uk/computer_

journal/hdb/Volume_32/Issue_
03/tiff/244.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
03/tiff/245.tif.

Zafarullah:1971:SSI

[Zaf71]  A. Zafarullah.  Some sta-

ble implicit difference methods for heat equation with derivative boundary condition. The Computer Journal, 14(3):309–311, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/140309.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/309.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/310.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/311.tif.

Zahran:1981:UDM

[Zah81] F. S. Zahran. Universal data model holder for data dictionary systems. The Computer Journal, 24(3):193–199, August 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/193.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
03/tiff/194.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_24/Issue_
                                        03/tiff/195.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        03/tiff/196.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        03/tiff/197.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        03/tiff/198.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        03/tiff/199.tif.

Zakrajsek:1977:GEP

[Zak77] E. Zakrajsek. Algorithm 93: Generalised eigenvalue problem. The Computer Journal, 20(1):86–91, February 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        01/tiff/86.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_20/Issue_01/tiff/
                                        87.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_20/Issue_01/tiff/88.
                                        tif;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        20/Issue_01/tiff/89.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_20/Issue_
                                        01/tiff/90.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_20/Issue_01/tiff/
                                        91.tif.

Zaki:1984:ASS

[Zak84] Ahmed S. Zaki. Algorithm 119: A space saving insertion algorithm for 2-3 trees. The Computer Journal, 27 (4):368–372, November 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/368.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/369.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/370.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/371.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/372.tif.

Zambardino:1970:ADP

[Zam70] R. A. Zambardino. Algorithm 53: Decomposition of positive definite symmetric band matrices. The Computer Journal, 13(4):421–422, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/421.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/422.tif.

Zambardino:1974:SSL

[Zam74] R. A. Zambardino. Algorithm 82: Solutions of systems of linear equations with partial pivoting and reduced storage requirements. The Computer Journal, 17(4):377–378, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        04/tiff/377.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_17/Issue_
                                        04/tiff/378.tif. The algo-

rithm number was omitted in the original journal issue.

Zangwill:1967:MFC

[Zan67] W. I. Zangwill. Minimizing a function without calculating derivatives. The Computer Journal, 10(3):293–296, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        03/100293.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        03/tiff/293.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        03/tiff/294.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        03/tiff/295.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_10/Issue_
                                        03/tiff/296.tif.

Zaremba:1969:SAI

[Zar69] W. A. Zaremba. A syntax for ALGOL input/output formats. The Computer Journal, 12(4):342–348, November 1969. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/120342.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/342.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/343.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/344.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/345.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/346.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/347.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/348.tif.

Zaremba:1980:DCA

[Zar80] W. A. Zaremba. Discussion and correspondence: An alternative for the ‘completer’ function. The Computer Journal, 23(3):286–287, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/286.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
03/tiff/287.tif. See corre-

spondence [Hun81].

Zdrahal:1981:RCP

[ZBS81] Z. Zdrahal, I. Bratko, and A. Shapiro. Recognition of complex patterns using cellular arrays. The Computer Journal, 24(3):263–270, August 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        03/tiff/263.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        03/tiff/264.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        03/tiff/265.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        03/tiff/266.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        03/tiff/267.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        03/tiff/268.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        03/tiff/269.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        03/tiff/270.tif.

Zissos:1971:NNO

[ZD71] D. Zissos and F. G. Duncan. NOR and NAND operators in Boolean algebra applied to switching circuit design. The Computer Journal, 14(4):413–417, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/140413.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/413.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/414.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/415.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/416.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
04/tiff/417.tif.

Zissos:1973:BM

[ZD73] D. Zissos and F. G. Duncan. Boolean minimisation. The Computer Journal, 16(2):174–179, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/160174.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/174.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/175.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/176.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
02/tiff/177.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/178.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_16/Issue_
                                        02/tiff/179.tif.

Zaki:1988:ARM

[ZE88] M. Zaki and M. M. Elboraey. Analysis of reliability models for interconnecting MIMD systems. The Computer Journal, 31(4):304–312, August 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        04/tiff/304.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        04/tiff/305.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        04/tiff/306.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        04/tiff/307.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        04/tiff/308.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        04/tiff/309.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        04/tiff/310.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        04/tiff/311.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_31/Issue_
                                        04/tiff/312.tif.

Zhou:1999:MRR

[ZG99] Wanlei Zhou and Andrzej Goscinski. Managing replicated remote procedure call transactions. The Computer Journal, 42(7):592–608, 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
07/420592.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
07/pdf/420592.pdf.

Zviran:1993:CPT

[ZH93] M. Zviran and W. J. Haga. A comparison of password techniques for multilevel authentication mechanisms. The Computer Journal, 36(3):227–237, June 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_36/Issue_03/
Vol36_03.body.html#AbstractZviran.

Zheng:1994:RSS

[ZHS94] Yuliang Zheng, T. Hardjono, and J. Seberry. Reusing shares in secret sharing schemes. The Computer Journal, 37 (3):199–205, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/Volume_37/Issue_03/
Vol37_03.body.html#AbstractZheng.

Zakian:1973:NIL

[ZL73] V. Zakian and R. K. Littlewood. Numerical inversion of Laplace transforms by weighted least-squares approximation. The Computer Journal, 16(1):66–68, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
01/160066.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
01/tiff/66.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_01/tiff/
67.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_16/Issue_01/tiff/68.
tif.

Zhang:1995:VGV

[ZM95] Kang Zhang and Gaurav Marwaha. Visputer — a graphical visualization tool for parallel programming. The Computer Journal, 38(8):658–669, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_38/Issue_08/
Vol38_08.body.html#AbstractZhang.

Zmud:1975:ETE

[Zmu75a]  R. W. Zmud.  Erratum:

“Two Enhancements to a Flexible Pricing Control System for Allocating Computer Resources”. The Computer Journal, 18(3):287, August 1975.

CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
03/tiff/287.tif. See [Zmu75b].

Zmud:1975:TEF

[Zmu75b] R. W. Zmud. Two enhancements to a flexible pricing control system for allocating computer resources. The Computer Journal, 18(2):112–114, May 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/112.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/113.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/114.tif. See erratum

[Zmu75a].

Zivani:1985:AIS

[ZOG85] Nivio Zivani, Henk J. Olivié, and Gaston H. Gonnet. The analysis of an improved symmetric binary B-tree algorithm. The Computer Journal, 28(4):417–425, August 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/417.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/418.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/419.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/420.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/421.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/422.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/423.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/424.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_
04/tiff/425.tif.

Zhao:1988:OOD

[ZR88] Liping Zhao and S. A. Roberts. An object-oriented data model for database modelling, implementation and access. The Computer Journal, 31(2):116– 124, April 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/Volume_
31/Issue_02/tiff/116.tif;
 http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/117.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/118.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/119.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_31/Issue_
02/tiff/120.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/121.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/122.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/123.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
02/tiff/124.tif.

Zhang:1996:ANR

[ZS96] N. Zhang and Q. Shi. Achieving non-repudiation of receipt. The Computer Journal, 39(10):844–??, ???? 1996. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www.oup.co.uk/computer_
journal/Volume_39/Issue_10/
Vol39_10.body.html#AbstractZhang;
 http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_10/
Vol39_10.body.html#AbstractZhang.

Zhang:1999:FAS

[ZSM99] N. Zhang, Q. Shi, and M. Merabti. A flexible approach to secure and fair document exchange. The Computer Journal, 42(7):569–581, 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_
07/420569.sgm.abs.html;
http://www3.oup.co.uk/computer_

                                        journal/hdb/Volume_42/Issue_
                                        07/pdf/420569.pdf.

Zhou:1999:ELM

[ZSP99] Sihui Zhou, Aruna Seneviratne, and Terry Percival. An efficient location management scheme for hybrid wireless asynchronous transfer mode networks. The Computer Journal, 42(6):487–500, 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        06/420487.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_42/Issue_
                                        06/pdf/420487.pdf.

Zvegintzov:1975:PAD

[Zve75] N. Zvegintzov. Problems with analytic description of text formats. The Computer Journal, 18(2):148–149, May 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        02/tiff/148.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_18/Issue_
                                        02/tiff/149.tif.

Zvegintzov:1980:PRI

[Zve80] N. Zvegintzov. Partial-match retrieval in an index sequential directory. The Computer Journal, 23(1):37–40, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_
01/tiff/37.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/
38.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_23/Issue_01/tiff/39.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
23/Issue_01/tiff/40.tif.